A Complete Bibliography of Publications in SIAM Journal on Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

30 May 2024 Version 3.41

Title word cross-reference

#13843 [Coo81]. #2759 [HN79]. #BIS [GSVY16, JKP20]. #BIS-Hard [JKP20]. #BIS-hardness [GSVY16]. #CSP [CCL16, CLX17, CF22]. #P [LSS13]. #P-Complete [LSS13].

#BIS [GGJ16]. #P [Sto85]. (1 + ϵ, β) [EP04]. (1 + ε) [FFKP18]. (2 + ε) [AGH17]. (2, 1) [Sch16]. (2n − 1) [Çam03]. (8, 5) [Bro99]. (α, β) [Efr05]. (Δ + 1) [BEK14, CLP20]. (≤ k) [Cha00]. (log n)^{2−o(1)} [GH98]. (N, +) [RS07]. (MODp − MODm) [GT00]. 0 [CKR04, KKMR09, Lip78, MS04, SSSS11, Shi79b]. 1 [BG81a, BP02, CM13a, HHH05, Lip78, McI74, MS04, SSSS11, Shi79b]. 1.5 [BMKM07]. 1.6 [GL87]. 1.998 [Xia13]. 1/2 [BFG23, KMZ18, BFNS15, DFM23]. 1/2 + ϵ [BFG23]. 1/8 [Tho18]. 12 [KS17a]. 2 [BHIK97, BGA23, CHKZ03, DH13, Fre97, FG82, GR05, GI92, GLW82, Hås98, Hås14b, HHH05, HJW84, HJW85, IKW12, Kap00, KC96, KMW75, KN08, KS17a, KP98, LW98a, La 00, LW80, MPRS79, NZ01, STP94, Sit14, Ste97, YS93, ZH05]. 2, 3 [RS78]. 2 − 3 [BH81b]. 2 − o(1) [GMPT10]. 2.5n [Pau77]. 2^{(log n)^{Ω(1)}} [KS17a]. 2 × n [AT77]. 3 [AES99, ASSS16, ABS07, AK97a, AK11, AS97, BHI^{+}13, CF86, CPPW12, Efr12, ES22, GI93, GN16, GLW82, GKKS16, Her14, Hwa80, La 00, MPRS79, Pel94, SS12a, Sch13, SCY00]. 3 × 3 [JM86]. 4 [FLMS15, GB82, Jac90, KMSV13, KS17b]. 4n [Zwi91]. 5 [BDD^{+}16, HCD89]. 8 [ENZ00].

0 [BT20, RS10b]. _{1} [GN16]. _{w} [CG01]. A [BG81a]. A2\frac{1}{2} [Swe99]. ABC [yCLZ00]. AC [Wil90]. AC^{0} [LRR17, AHM^{+}08, She09, BS92]. α [CW83, Zha95]. b [GS21a, GHT76]. β [CW83, Zha95]. BP [TW89]. c [FJ90a]. c^{k}n [BDD^{+}16]. N P [Szk99]. N P_{C} [MM98]. D [ASSS16, CH06, DL90, DJZC00]. D0L [JS81]. D_{∞} [Wad76]. Δ [BEK14]. Δ + poly(log Δ) [HMR98]. ℓ_{p} [DDH^{+}09]. ϵ [AES10, KLX10, RS10a]. f [GS21a, GS21b]. \frac{3}{2} [MRT07]. G(n, d/n) [Eft16]. GF (2) [Lem75, RB91]. GF(22m) [Cus98]. GF (2^{n}) [Lem75]. GF(q) [MvV92]. GF(q)[x] [Nor89]. H [GKP04, CPPT24, DHV22, GGJ16]. H_{1}(Γ_{h}, Z) [HO02]. hol^{c} [Bac21]. j [ABK^{+}98]. K [JM78, KIM81, SW85, Sch04a, GS17a, SWCP96, AM06, ASSS16, ANFSW20, ABK^{+}98, AKPW95, AV09, AGK^{+}04, AABV98, BGNV18, BC22, BD97, BCC^{+}09, CKT99, CC07, Che09, CKT93, CT00, CVV03, CV14, CH06, CL91, CAKM19, CO10, CSY87, CS24, DJZC00, ERS20, Epp98, Eve75, FL12, FSS20, FW98, Fre97, Fri86a, FGK05, FRS19, Fus88, GG12, GLLON18, HTC13, KYN07, KRR94, Kho06, KS91, Ko89, KR07, KN05, LS16, Liu22, Mao93, MR96, MTZC81, MBCV98, Mit99, NN17, NS08, Pio04, Rod10, Ros14, SZ00, SV95, SS90, Sch94, SBI04, Tho05, Tro78, Vai88, WZ21, YY82, Yao82]. kd [AdBS10]. L [HOS92, FGLM98]. L^{1} [FKSV02]. L_{1} [KR09, LW80, NS07, KN08]. L_{∞} [LW80]. L_{p} [AKOT03, JW21]. λ [DCdP98, NR13]. log_{d}(N, m, p) [HHW05]. log log N [Wil85b]. log n [CMSS00]. m [HHH05, Nic88, Pfa83, RS85a, Sut91]. m + 1 [HHH05]. F_{p}^{n} [HL14]. R^{3} [AEZ20, KRS10a, The02]. VP ≠ VNP [BLMW11]. N C [DR95a, HM87]. N P [HN10, KM13]. O(log ⇕) [CMS18]. P =?N P [BGS75]. AC^{0}[⊕] [LSS^{+}21]. LWE [BV14]. NP [Hir23]. TFNP [GHJ^{+}24]. m × n × (mn − 2) [AL83]. m × n × (mn − k) [Bsh90]. N [Rob84, AM96, SS89a, ACR^{+}10, And99, FW76, Fri86a, Gal84a, Gal84b, Hwa80]. n > 3t [GM98a]. N^{1/2+o(1)} [ACR^{+}10, KSS20]. n^{2} + o(1) [Ull16]. n^{5/2} [HK73]. NC [Wil90, KM97a, VV89]. NC^{0} [AIK06]. NP [BH77, Pla78]. NP − P [Kur85]. N × N [CSS90]. O [Sto79]. O(1) [FN93, GHW10]. O(1/ϵ) [BI06]. O(Δm) [Sch98b]. O(Distance + 1/SessionRate) [AFHB^{+}00]. O(E log E + I) [Mye85]. O(EV log V) [GMG86]. O(log^{2} k) [FL12]. O(log k) [AR98b, Fil19]. O(log log n) [AB96, BG90, LPSPP05]. O(log n) [BGS15, BGS18, Hag90, JM96b, KR88, LW88b]. O(log n log log n) [Tri08]. O(m) [BCCK19]. O(M · N) [Hsu95]. O(m log n) [CHT93]. O(mn) [HKN16]. O(N) [CSS90, Kun93, CS16]. O(N + M) [RLDL96]. O(n^{2}) [DF18]. O(n^{2} log^{2} n) [Spi73]. O(n^{2} log n) [MT87, SCH92]. O(n^{2} log n log^{∗} n) [Blo83]. o(n^{3}) [CHM96]. O(n^{6}) [Raj96]. O(nα) [KS06]. O(n · log n) [MT78]. O(n log^{2}(n)) [Rei83]. O(n log^{2} h) [ES91]. O(n log^{2} n) [AW96, HJ85, MTZC81]. O(n log log n) [TV88a, TV88b]. O(n log n) [AR98a, BY98, CFCH^{+}00, Fri86a, GKK13, WZ21, ALS22]. O(nm) [KS06]. O(\sqrt{log n}) [AHK10b]. O^{∗}(n^{3}) [CV18]. oAC^{0} [BH12]. ω [RY87, Saf06]. Ω(D log(N/D)) [KM98]. Ω(log n) [GR05]. Ω(n log n) [FMP82]. Ω(\sqrt{log log n}) [GJM98]. GF(2) [AKW90]. GF(2^{n}) [Sti90]. Poly(log log N) [AM96]. P [Mer01, AR88, Fab77, Lau83, Loo83, MT83]. P^{∗} [BG93b]. P^{A} ≠ NP^{A} ≠ co−NP^{A} [BG81a]. P_{4} [JO92]. P_{6} [CSZ24b, CSZ24a]. ϕ [VCJS22]. p → q [BGG^{+}23]. R^{3} [Meg83a]. s [BBRS98, BKRU94, KS91, Rei83, TV20]. S = O(2^{n/4}) [SS81]. Sample(x) = (a ∗ x ≤ t) [Tho18]. \sqrt{n} [HSSS22]. st [BE98, Edm98, EPA99, Tri08]. t [BBRS98, BKRU94, CHJT04, CC07, Coh98, CKP86, KS91, Rei83, Rus78, TV20]. t + 1 [GM98a]. t/(t + 1) [DTA94]. T = O(2^{n/2}) [SS81]. ASPACE(o(log log n)) [Iwa93]. BPP [Lut93]. Dover [KS95]. D^{p} [CM87a]. LR(k) [AU73]. PV_{chunk} [HZ77]. Õ(M(|V |)) [Che97]. Õ(n^{2}) [AHK10b]. Õ(n^{5/3}) [GKM^{+}21]. V [BR79]. ε [Iba78]. W [P] [AR08]. x [ES88]. X + Y [JM78]. X_{1} + X_{2} + ⋯ + X_{m} [JM78]. Z [MW99b].

* [IQ19]. *-Algebras [IQ19].

-Adic [Lau83, Loo83]. -Approximation [BDD^{+}16, BFNS15, BFG23, BGA23, ENZ00, LR21, MRT07, NZ01, FL12, Swe99]. -Approximations [NR13]. -Ary [Rus78, Tro78]. -Atomicity [GLLON18]. -Automata [Saf06]. -Block [CMSS00]. -Body [And99]. -Calculus [DCdP98]. -Center [Tho05, WZ21]. -Centre [MT83]. -Chains [FW98]. -Clique [Ros14]. -Cliques [ERS20]. -CNFs [ABS07]. -Colorable [AK97a, KS17a]. -Coloring [GKP04, BEK14, CLP20]. -Colorings [BC22, GGJ16]. -Colouring [HCD89]. -Compatible [CPPW12]. -Competitive [VCJS22, CMS18]. -Complete [Szk99, GN16, CM87a]. -Complexes [STP94]. -complexity [TW89]. -Concave [YS93]. -Connected [CT00, ZH05, GB82]. -Connectivity [BE98, Edm98, EPA99, LW98a, Sch13, Tri08, KS91]. -connectors [Ngo05]. -Covered [Efr05]. -Cube [FW76]. -d [Kap00, NS08]. -decomposable [FJ90a]. -Diagnosability [CC07]. -Diagnosable [DTA94]. -Difference [FKSV02]. -Dimensional [AES99, HJW84, KMW75, Pfa83, Sch04a, Vai88, DL90, HJW85, LW80, Yao82]. -Discrepancy [AKOT03]. -distributed [Fus88]. -DNF [SBI04]. -Dominant [HTC13]. -Edge [GG12, GI92]. -Edge-Connected [GI93, La 00]. -Edge-Connectivity [Fre97]. -Embedding [FFKP18]. -Extension [CKR04, KKMR09]. -factors [GS21a, GS21b]. -Fit [Mao93]. -Flow [Jac90]. -Free [BDK00, Iba78, CSZ24b, CSZ24a, CPPT24]. -Hard [CG01, Pla78, GGJ16]. -Hardness [KM13]. -Hop [CHKZ03]. -hulls [CSY87]. -Levels [Cha00]. -Lower [Pau77]. -machines [RY87]. -matching [GS21a]. -Matchings [GHT76]. -Means [ANFSW20, AV09, Che09, CAKM19, FSS20, FRS19]. -Median [ANFSW20, AGK^{+}04, Che09, CAKM19, KR07, LS16, Tho05]. -Minor [DHV22]. -modal [MR96]. -Models [Wad76]. -Monotonic [BHIK97]. -MST [MBCV98, Mit99]. -Nearest [CS24]. -Net [RS10a]. -Nets [AES10]. -Node [CV14, KN05]. -Opt [BHZ23, CKT99]. -Optimal [BR79]. -Orders [Sch16]. -Oriented [MW99b]. -Partite [BCC^{+}09]. -Planarity [CM13a]. -Printable [FGLM98, AR88]. -Private [GR05]. -probe [SS90]. -Processor [AM96]. -Query [DH13, Efr12, IKW12, KSS20]. -Random [Mer01]. -rational [Gal84a, Gal84b]. -rays [ES88]. -Resilient [CHJT04]. -Sat [AGH17, AM06, CO10, FGK05, GS17a, GN16, Her14]. -satisfiability [CF86]. -Satisfy- [ABK^{+}98]. -sequences [Sut91]. -Server [Sit14, AKPW95, KRR94]. -Set [SZ00]. -Shaped [HOS92]. -Solution [Nic88]. -Sorters [SS89a]. -Soundness [KLX10]. -Space [AS97, ES22, Pel94, SCY00]. -Spanner [EP04]. -Spanners [KP98, Coh98]. -Sparse [JO92]. -Stage [Çam03]. -Steiner [BD97]. -Subgraph [Kho06]. -Sum [BGNV18]. -tape [FG82]. -terminal [SW85, SWCP96]. -Time [CHM96, TV88b, AM96, RLDL96, WZ21, AB96, CHT93, TV88a]. -Tours [TV20]. -transform [Sut91]. -Trees [AdBS10, AABV98, MPRS79, RS78]. -Uniform [KS17a]. -versus- [HHH05]. -Vertex [CVV03, FL12, CKT93]. -Way [BP02, Fab77, KYN07]. -Well-Supported [DFM23].

0.2ex [BV06].

1/6 [AM05]. 13 [FJ90b]. 14 [Lan91, Lon88, SWCP96]. 16 [Vai90]. 17 [Col93, Kad91].

2 [Aus10]. 2-CSP [Aus10]. 2-Variable [KKMO07]. 20 [BX92]. 21 [RS94a]. 22 [CCG^{+}97]. 24 [BG96]. 2EXPTIME [Koz09].

3-Colorability [KLR03]. 3-Dimensional [KS03a]. 3-Stage [CG07b]. 3CNF [BSHR05].

48th [ACHCP21].

5-Independent [TZ12]. 57 [HN79]. 58 [Coo81].

6 [HN79].

7 [Coo81].

8 [HH80]. 80h [HH80]. 85c [FJ90b]. 86d [Lan91]. 86m [SWCP96]. 87h [Lon88]. 88c [Vai90]. 89m [Col93].

90a [Kad91]. 92h [BX92]. 93f [RS94a]. 94b [CCG^{+}97]. 95j [BG96].

= [Mei13].

a.k.a [Sch16]. ABD [Bro99]. Abduction [CZ06]. Abelian [Ili89a, Ili89b, MC87]. Absolute [Duv91, Ste97]. Abstract [CS05, Gär95, Mer02, FL83b]. Abstraction [DGGW20, JM96a]. AC [BT20, RS10b]. Accelerated [MPR03]. Accelerating [BŠVV08]. Acceleration [WP03]. Acceptance [Saf06, KR88]. Accepted [Boo72a]. Accepting [Mor82]. Acceptors [BNP74, FG82]. Access [BAG01, BLR^{+}15, DK15, DKR96, DGLV10, HLR96, LL92, PY91, SZ76, CDR86, IP87, KvLP88, KMST87, Lon85, Lon88, SV84, Wil86].

Accesses [BKPY18]. accessing [Wil89]. Accommodating [BLN01]. Account [Hen76]. Accounting [GMR98]. Accurate [CK01, LMM99]. Achieve [DFL^{+}22]. Achieving [BADTS22, GM18, IKK^{+}11]. Achilles [Mar99]. Acknowledgment [GJKP04]. Acknowledgment-Based [GJKP04]. ACM [AGK^{+}09, ACHCP21, AIM16, Bab06, CCRZWN22, CDG11, CR17, DKI18, DKSS22, EvMP^{+}12, FGKO08, IKM^{+}12, MS16, PKS^{+}13, SK18, VNA20, AKS09, Vit85b]. Action [ABR90]. Acyclic [BKS92, Dur10, HPT99, HP99b, HL96, Pro86, TY84, TY85]. Acyclicity [BRV16, GP04, TY84, TY85]. Adams [AM13, TZ17]. Adaptive [AAT97, AF01, BNS^{+}21, BN89, CLP16, CL21, CS03b, CG94, DKKP99, FRV10, Gam03, LW98b, LZ00, Mes10, SRR95, GMR88]. Adaptively [CPV22]. Addendum [BX92, TY85]. Adding [CM13a]. Addition [DL80, DLS81, FR75, FK90, KK77, Thu99]. Additional [Mer01]. Additions [BC76, Kal95]. Additive [ABP18, ADL18, BKN10, GK98b, LS13, NPSW23, Pro76, Ris85]. Addressable [MO74]. Addressing [Hos75, Yao78]. Adiabatic [AvDK^{+}07, ATS07]. Adic [Lau83, Loo83]. adjacency [ACM84b]. Adjoining [PSW90, Raj96]. Adjunct [Lev76]. Adjunction [KOWZ23]. Adjusting [KS20, ST86]. admit [Tom82]. Advance [Fis05]. Advanced [FRS03]. Advantage [Amb16]. Adversarial [ÀBS04, AKU05, AKK08, BGL04, BE17, Gam03, LPSR04, MNS11]. Adversary [ACH10, Mac97]. Advice [AD14, CEEP16, HU22, HS03, KT94]. Affine [BSK12, BSRZ15, BGWZ20, HL14, MRRS07]. Affordable [FK00]. Again [KPS94]. Against [KQT99, AN93, AL18, CW04, CS03b, Fis05, GMR88, TS11, Wat09]. Agents [DDDR11]. Aggregate [CP12].

Aggregation [CTX10]. Agnostic [FGKP09, FGRW12]. Agnostically [KKMS08]. Agreement [AK97b, AF01, CFCH^{+}00, DS83, DGP07, FM97, FBGSV15, GGP11, GM98a, HNO^{+}20, HMW01, MRRT08, SZ00, TPS87, WBZ13]. Ahead [Smi76]. Ajtai [Mic04]. Algebra [BPSV04, BT83a, BKO88, Hal75, HRB87, HS87, Sto79]. Algebraic [AL18, ALRS98, AEZ20, BAG01, BOC92, BMM21, BG21, BCE95, CC86, CCC^{+}13, GRRV21, GSY95, GK98b, GKSS22, Har09, HHN04, JB90, Lan85, Pri86, PM03, Rei86, Ren92, Rie99, Rie02, Sem09, SH96, Yao91, Yun79, vzG84, ACM84a, ACM84b, GJ82, Lan91, Len87]. Algebras [BS00, Blä04, Bsh94b, BJK05, Che08, CS07, IMM^{+}10, IQ19, KT88, Pig91, Gal84a, Gal84b, Har85, Lic89, dG83]. Algorithm [AT77, AAEZ21, AFB94, AM96, Amb07, ALW08, AV09, AS10b, ABH98, AR98b, Bak86, BHI^{+}13, BDG19, BBRS98, BST23, BGK20, BMKM07, BFG03, Ber76, BKHL00, BC00, Bod96, BDD^{+}16, BSW14, Bra98, BT94, BFP89, BCM^{+}15, BD06, BS82, BCGR92, BA10, CF22, CFJ^{+}10, CL77, CBCG05, CCK^{+}22, CW75, CKT99, Cha92b, CMS18, CH95, CHM96, CVV03, CKS17, mCL04, CZ23, Cla88, CP05, CO10, CSSS89, Col94, CFCH^{+}00, CH02, Col74, CPS85, CDH13, CGH^{+}98, CPPW12, DKLR00, DDY16, DHR20, DK15, DR95a, DFM23, Dow90, DLR95, DCH21, ET93, Eft16, EW03, EK05, EiC17, ESY17, EL86, Eve75, ENZ00, FL12, FKSV02, FBL03, FW74, Fil19, Fin20, FV13, FR94, GP90, GM91a, GKP98, GMSV15, GW77, Gär95, Gaz91b, GL80, GM91b, GHW10, Gol08]. Algorithm [GJ13, GKPS08, GJ19, GR17, HHKL21, HL05a, HC99, HU75, HT98, HM95, HSS01, HKN21, HS99, HSY22, Hli05, HL72, Iwa03, IK22b, Jai07, JPZ08, Jus92, KMSZ05, KYN07, KMR87, KW97, KST99a, KM97a, Kar99b, KKL^{+}93, KM94a, KLP96, KK02, KM08, KS86, KMMT09, KRT03, KS95, KM77, Kup05, LP94, LS02, LS98, LY05, LPT98, LS93, LST99, Mal05, MM02, MPP22, MT78, MTZC81, MV13, ML00, MBCV98, Moi16, MRŚ10, MRT07, NGM97, NZ01, NSS00, NPSW23, NS09, Oht76, PS91, Par80, Par98, PR02, PR05, Pla90, PVG04, RY96, Ram96, Rey77, RLDL96, RZ16, RS99, Sak90, Sav74, SW99, STU97, Sit14, ST13, Spi73, Ste97, Swe99, Tan78, TV85, TV22, Tri08, Tsa92, TIAS77, Tze92, VP03, Vég16, VCJS22, WP03, WZ21, WZ06]. Algorithm [WZ15, dBBKBK23, AKR95, AB96, ACM84a, ACM84b, AS80, BéT89, BOFKT88, BN89, Bit82, Blo83, BG90, CHT93, Cha84, CKT93, CNS82, CL91, CL79, CG73, ES91, FG82, Fri84, FK88, Fri86b, GMG86, GGT89, GS89, GL87, GO80, HT82a, HJ85, HY88, HK73, Hor87, JS89a, KM96, Kar79, KIM81, Kin86, KM89, Kun93, Liu81, Lub86, MN88, MNS85, May84, MD73, Meg83b, Mil88, MT87, Mye85, PST88, Ryt80, SW85, SWCP96, Sch88, Sch78, SS81, Sha79, SCH92, Shi79a, Shi80, SW89, Sys81, TV88b, TV88a, VR89, dG83, BKPS06]. Algorithmic [ACOH^{+}10, AMT07, AHLM07, BNS^{+}21, BKL99, CR00, DD13, DR14, DX21, FGS85, GR11a, Hal08, HV20, HNR08, KV10, KV12, Kol18, LLRS01, RTL76, SMW07, SW10, She20, The02]. Algorithmically [Boo94, TT13]. Algorithms [Aar03, ABW18, AB24, AIK19, AG94a, AS90, AW01, AS05a, AP06, ANFSW20, AOST94, ACC^{+}11, AAELvZ12, Alb98, AS05b, AAT97, Amb16, AK97b, ASS17, AW97, AKK08, AALM90, ABHK11, ABD^{+}07, ABP19, AKSW22, AF01, Aur87, ABSD^{+}98, BRS08, Bak97, BCS08, BBN12, BGNV18, BNL04, BYGNR98, BBR04, BK10b, BHL18, BCN^{+}20, Bei90, BHNW23, BCK^{+}07, BH98, BS93, BGG22, BDK00, BMS13, BPP23, BGSW19, Bro78, BGKP99, BM02, BGK^{+}08, BDH^{+}05, BT05, BY98, CLX17, CKR04, Cha10, CHPJ22, CGH13, Cha98a, CK09, CKR01, CG05, CLM05, CHKS10, CS18, CKL^{+}09, Che97, CN85, CT90, CCH^{+}16, CC82, CS03a, CMS14, CL89, CM94, CK04b, COS99, CV18, Csa76, CPP16, CLM^{+}20, DGL23, DKK^{+}19a, DDH^{+}09, DGV05, DNS81, DHH96, DS94, DKR96, DS83, DH04, DKM06a, DKM06b]. Algorithms [DKM06c, DKS88, EGH07, EGS10, EV21, EvS05, ENRS99, FHL08, FW98, FBGSV15, FFK^{+}98, FM03, FW76, FKTV08, FLM^{+}22, Fra91, FHK78, FS89, FNR15, Gab77, GK82, GT89, GG12, GS21a, GS21b, GG10, GI92, GS17a, GPT00, GK07, Gat92, Gav72b, GK12a, GLM^{+}99, GMR98, Gie95, GHSZ08, GKKS93, GPST92, Gol95, GJLR97, Gop08, Gra94, GM00, GKS90, GQ17, GLN02, GPRS11, GL21, GMFB87, HJLS14, Hal02, HPQ17, HT84a, HKM08, Har20, Har09, Hav24, HKSS17, HL86, HN94, Hol73, Hol74, HU73, HM99, IL04, IKS10, IQ19, IL98, IWY16, JC91, JV08, JT10, JNS06, JK82, JKP20, JS93, JDU^{+}74, KKL09, KQT99, KW00, KR95, KNP99, Ken76, KMV92, KKW04, KS94, KPST94, Kle96, KY15, Kli79, KS02, KR02, KM99, KMMS06, KT77, LS77, LY13]. Algorithms [Lin92, Liu72, LRS01, LRWY99, MS10, Mac97, MS99, MSS07, MR99, MYZ06, MHSR98, MS91, MPS92, MD76, MP05, MvV92, MW99b, MRRS07, MMSA93, MR07b, MO74, NNS89, NSV94, PT87a, Pan80, PY90, PR17, PSWvZ17, Pul84, PM03, Rab80, RV98, Rei14, RS92, RS94b, RS91b, Riv76, Riv77, RZ08, SLC91, SSW94, Sho92, Sho97, Smi76, Sti90, SF11, SS12b, Tar72, Tet99, TN95, Tro78, UY91, Val08, VY23, Ver97, VC85, WLL08, Wes94, WBZ13, WW75, Wil76, XLXJ09, YYI12, ZS02, dBBKB^{+}20, vDHI06, vzG84, AA88, ALM96, ACG89, BS83, BC84, BW79, Bol81, CM89, CMST85, CGJT80, Coh98, CB81, Cun86, Dye84, Fed87, FL83a, FJ81, Gol81, Gus87b, HJLS89, Ham83, Hoc82, Hsu85, Hsu95, Ili89a, Ili89b, IA86, IKP96, JS82, JM86, KB79, Kao93]. algorithms [KS93, KRR94, KS91, LLK80, LP83b, Meg83a, Nor89, PA79, PA80, RR89, RS94a, RS91a, Rei85, RH78, Sag83, SJ81, Smi86, Sto85, TY84, TY85, TN82, Tom82, TC84, WT79, WT89, Yao85b, ZS89]. Aligned [Tót08]. Alignment [NSH^{+}08, WJG00, KM96]. All-Ones [CLWZ04]. All-or-Nothing [CKS13, iKK18]. All-Pairs [CL77, Cha10, DHZ00, HKN16, KKP93, LY13, RZ12, Wil18, BK10b]. All-Terminal [GJ19, Kar99a]. allocated [BC84]. Allocating [KRT00]. Allocation [ADK^{+}08, AS10b, AST99, AMRU22, BNCK^{+}99, BKK^{+}04, CW75, CFW93a, CKV10, FP86, Hol73, Hol74, KK06, LL95, LL00a, Set75, CL79, CKS85, Ked85, Yao81]. Allocations [ANSS21, ABKU99, BCSV06, CG10, HHR^{+}16, JV08]. Allowed [MV75]. Almost [ACER22, BATS11, BM99, BHV99, CKPR02, CLL08, CKMS21, CP12, DHIP07, DHZ00, FGLS14, GM91a, GP01, HT17, HKN21, HR82, KPV14, LS18, Mic04, Moi16, MR08, RT82, RZ16, SS11a, Sch87, FK89]. Almost-Linear [LS18, MR08, RT82]. Almost-Minimum [BM99]. Almost-Optimal [CP12]. Almost-Optimally [HT17]. Almost-Ramanujan [BATS11]. Almost-Tight [HKN21]. Along [JB90]. Alphabet [ABF94, GP96, Iba78]. Alphabet-Independent [GP96]. Alphabetic [CKP86, Ita76, KK85, LP94]. Alphabets [Jac72]. Alternating [BCG^{+}21, Gas87, IJ91, KLL90, LLS84, LR96]. alternative [Lic89]. Alternatives [CK14, FKKN11]. Always [Mak97]. AM [SU09]. ambiguity [RI89]. Ambiguous [Leu98]. Ambivalent [Fre97]. amid [CHW13]. Amidst [MS97, Sha87, AS96]. Among [CKT00, HJLS14, HSY22, JLTX20, ASU79, HJLS89, IST12, JJLZ22, KPR^{+}13, MS08b, Pel96, YLW95, adH86]. Amortization [AARV21]. Amortized [FKNN95, MT86, Saw01, WT89]. Amount [CC97, GR05]. Amplification [AARV21, BBR88, CRS14, DLWZ14, DZ97, GKS93, GM20, GM91c, IJK09, SV10, SB13]. Amplify [HVV06]. Analog [Maa97]. Analogues [Gel75a]. Analyses [GLM^{+}99, Nor89]. Analysing [MMR95]. Analysis [ABKPM09, AV09, BNS^{+}21, BCGL17, BW99, BS93, Bro78, Bro79, BP81, BCM^{+}15, BP93b, BT05, CW75, CV83, CM89, Che89, CB74, CCRM86, CKR72, CPW82, CS03b, DH20, DRT92, DHR97, Fat74, FK88, FMM11, GM01, Gus90, HLR96, HU75, HT98, HL86, IL04, JKLAR90, Ken76, Kin86, KP00a, KLL17, KLR21, LS77, Lav73, LSH05, LS10, Liu72, MT86, Mil76, MSV87, MR02, PY90, RT82, Rey77, Ros80, Ros82a, SS05a, Smi76, Tsa92, WC92, Wir79, AMS84, AL88, BT80c, CF86, CL79, CL89, CK80, DL90, HR87, Meh82, Pap81, PB83b, RSL77, VR89, WT89, Yao81]. Analysts [MO12]. Analytic [Kao97]. Analytic-Invariant [Kao97]. analyze [HR86]. Analyzing [COF14, Ver97]. Anarchy [ADG^{+}11, BBM13]. Ancestor [AAK^{+}06, Mai79, YR92]. Ancestor-Controlled [YR92]. Ancestors [AHU76, HT84a, ASSU81, SV88]. Anchored [BVY22]. AND-OR [ACR^{+}10]. AND/OR [GL95, MSS04]. Angular [ACK^{+}99]. Angular-Metric [ACK^{+}99]. Anisotropic [BWY15b, BRLW19, CNVW08, CNVW10]. Annealing [BŠVV08]. Annual [AGK^{+}09, AKS09, ACHCP21, AIM16, Bab06, BBS13, BSO14, BCAKT23, CCRZWN22, CDG11, CR17, DKI18, DMK19, DKSS22, EvMP^{+}12, FGKO08, HIMW16, IKM^{+}12, KRS20, LMOT14, MS16, MV18, PKS^{+}13, Rou15, SK18, VNA20]. Anomalies [Cha89]. Anonymous [SY99, SY06]. Answer [BKN09]. Answering [Cha86, Ull16]. Anti [ESV20]. Anti-Concentration [ESV20]. Any [ACR^{+}10, Aus10, HNO^{+}09, Hai13, LH00, Bie90, HILL99]. Apex [FLM^{+}22]. Apex-Minor-Free [FLM^{+}22]. Application [AFB96, ABD^{+}07, AV09, Asa87, AKOT03, Bak97, BX91, BFL06, BGV00, BGS07, BvMR^{+}00, CGJ78, CO82, Doe04, Dye86, Gav72a, GS00, GR17, HOS00, Maz76, NSH^{+}08, Ram94, Sch98a, ST13, SOAD09, ASSU81, AKPW95, BX92, PCHM85, Pla80]. Application-Controlled [BGV00]. Application-Specific [GR17]. Applications [AS90, Aga92, AAS97, AAOS97, AES99, AS05a, AAEZ21, AFG^{+}14, AEZ20, Aur87, BCS10, BYGNR98, BG09, BBR04, BSGH^{+}06, BG22, CGH^{+}89, Cha06, CHPJ20, CC07, CRS95, Cha98b, CC99, Che09, CLL05, CLS19, CPS16, CRS14, Coh16, CV88, DGL23, DHR97, DKSS13b, DN07, DS00, EN19, ELRS03, FSZ21, FM03, FMS10, FPS16, FW07, GR02, GZ05, GVY96, GKK^{+}09, Gel75b, GIP20, GL23, Gia95, GV05, Han96, HPM06, HN10, HC99, HRZW20, HS90, IIM85, IQ19, JR13, KLR19, KRS12, KM97b, KLP96, KLL^{+}15, KPST94, KR78, LV22a, LP77, LLRS01, LWW21, LT80, LATV17, MTZC81, Mic04, MR91, NN93, NSV94, NR13, PW11, RS94b, RW00, RS96, SS11b, SJ05, Sri06, TZ12, Ver97, AB79, AK88, BCD^{+}89a, BCD^{+}89b, CW83, EW86, Fed87, Fre85, GGT89, GI80, Iba78, IKM85]. applications [LW80, EOS86]. Applied [GR93, HM73]. Applying [DG93]. Approach [AT98, ABF94, BLMW11, CDW21, Cha86, Cha98b, CPT96, Duv91, EC95, GRRV21, GR17, GPV94, KS97, KK02, LZ22b, LV95, LNRW10, ŁN06, MS01, Pre81, PM03, Sub94, WY07, Cha88, Gus88a, HS88]. Approaches [AW21, EGOS07, FGS85]. Approaching [KMP14]. Approximability [ABF^{+}99, BK11, CG10, DFHS08, JKK06, KTW99, KMSV98, Tre00, Von13]. Approximate [AC09, AA89, AS10a, AF03, AdFM18, AR98b, Bar18, BK10b, BFKL21, BSRZ15, BLR80, BP15, BUV18, BT20, CR10, CNVW08, CNVW10, CJ14, Coh95, CV88, CH02, DY10, DMR09, DKM06c, EN19, EKR01, ENRS99, FFTC20, FKSV02, FMS10, FMdR10, Fri90b, GP90, GVY96, GS73, GLPS12, GR21, GHS02, HP99a, HPK13, HK11, HKN16, IJK09, JP05, KS96b, KKM^{+}11, KOR00, LPSR09, LP13, Mat91, Mit99, Nat91, Nat92, Nat95, NTZ16, OPS04, Rei99, RZ12, Sch98a, SY12, TU93, Vid16, Wad78, Vid20]. Approximately [BDGJ99, CDR10, DLM22, DFJ98, ELRS17, GGJ16, KM13]. Approximating [AN06, AO12, AHP08, BNGNS01, Bar18, BDKR05, BBC^{+}22, BI06, BT15, CM99, CG12, CRT05, CGYZ06, CT00, CV14, CG18, CFLS97, CEF^{+}05, DKM06a, ERS20, EP09, FMS02, FHKS02, GG10, HSSS22, HPK15, HKN21, JS89b, KRY95, KR05, KN05, KMR22, LS16, LNW19, Man08, MR12, NS00, NSH^{+}08, Nut10, RRSS09, Ren92, SV18, TY97, VA00, VVYZ07, Ren89]. Approximation [AB24, AS98, AW01, AAELvZ12, AM04, AdW22, ALW08, ABHK11, AK03, AHK10b, AS10b, AR98b, AABV98, BRS^{+}21, BRS08, BCS08, BCS10, BYGNR98, BBR04, BGK16, BHL18, BGK20, BCNW09, BPU92, BMKM07, BK15, BGG^{+}19, BHNW23, BCK^{+}07, BDD^{+}16, BSW14, Bra98, BKV11, BHZ23, BFNS15, BFG23, BA10, BGA23, CKR04, CHD21, CGL97, CMNS01, CK05, CHKS10, CS18, CM01, CVV03, CS03a, CPPT24, CAKM19, CAdVdM21, CFF^{+}12, CKST99, CS24, DDDR11, DKM06b, DLR95, EW03, EK05, Elk06, EJS05, ENZ00, FL12, FK02, FHL08, FV16, FKK^{+}08, FHK78, FJ81, FNR15, GK12a, GW95, GMY12, GMM09, GJ19, GPRS11, GHM^{+}11, Hal02, HPQ17, Har20, HL04a, HL05a, Hoc82, JKLS05, JT10, Jan10, JR20, JS93, JL95a, JS20, KKL09, KW00, Kar99a, KM08, KPST94, Kle08]. Approximation [KY15, KM81, KMMT09, KS02, KR07, KR02, KKL04, LS13, LR21, LS16, LNRW10, MR99, MYZ06, Man95, MHSR98, MBCV98, Mit99, MW99b, MR07b, MRT07, MRR15, NZ01, NSS00, PW04, PR17, QS06, RV98, RR04, RB91, SHK^{+}12, SSW94, Sit21, Sri99, SVV12, SF11, SS12b, Swe99, TV22, TSSW00, Vel85, WJG00, WLL08, WLB^{+}99, YYI12, Zim98, AKR95, CNS82, GL87, Hal75, HS88, Lag85, Mur87, Pro88, Sto85]. Approximations [BLR14, CLL19, EGS03, HY17, NR13, Pri86, Rot16, Wan92, Wan98, ORS86]. Apriori [PVG04]. APTAS [EL08]. Arbitrarily [ACG^{+}16, BGL04]. Arbitrary [AKU05, BY86, BX91, BL21, BDD^{+}07, CCWZ18, Cha92a, CKP14, FR94, Jus92, Pla90, RS03a, Sav73, Vio07, BX92, GJST81, Sim83]. Arborescence [SS05b]. Arborescences [NSV94]. Arboricity [CN85]. Arc [APS93, BF03a, Cha98a, DHH96, LSW90, Tuc80, VP03, Hsu85, Hsu95, MN88, SCH92]. Arc-Disjoint [BF03a]. Architecture [PY90, PST88]. Architecture-Independent [PY90]. Architectures [ABR90]. Arcs [Spi73]. Area [BG82, EMSV12, Ked85]. Area-Time [BG82]. Area-Universal [EMSV12]. Areas [EGS03]. Argument [Hai13]. Arguments [Aar06, BG09, HNO^{+}09, LM08a, SU77, YH90]. Arising [BLMW11, IMM^{+}10]. Arithmetic [AIK14, BHH95, BC98, DL91, DKSS13a, DSY10, Ebe89, GKKS16, Hoo90, KLSS17, KM75, KS15, KS17b, KS19, MVW04, NP22, RSY08, Shp09, AB79, BJR89, GJ82, HH79, HH80, MRK88, Ren89]. arm [AK88]. Armed [ACBG^{+}17, BSS14]. arms [HJW85]. Arrangement [AMS11, CEG^{+}93, EiC17, HP00b, SSSV01, Shi79a]. Arrangements [AdBMS98, AMS98, AES99, AS05a, EOS86, ESS93, Mat91, PS91, Rie99, Rie02, EW86]. Array [AHHP00, Kos75, Kni88, RL88]. Arrays [AB98, Bol76, CMS97, FH11, GV05, IJ87, Lou83, MM93, Ros75a, Ros75b, Sei77, Bak78b, IP87]. Arrow [Jai07]. Arthur [CCM^{+}19, KLN19, San09]. Ary [Rus78, Tro78]. Ascending [NP97]. Aspects [BKL99, Cyp93b, GR11a, Hal08, LLRS01, MS76a, RTL76]. Assembled [SW07]. Assemblies [CGR12]. Assembly [AKK^{+}09, ACG^{+}05, CD17, Dot10, SW10]. Assertion [Mis77]. assertions [AM80]. Assignment [DR06, ELRS03, FKR95, FS07b, JR20, Sot13, Wal79, AL88]. Assignments [CVY19, KB76, Gus88b]. Associative [BP90, Bro99, Bsh94b, FSW87, Har85]. Associativity [LL00b]. Assumptions [AHR96, CLP16, FLS99]. Asteroidal [COS99]. Asymmetric [FKR95, FS07b, FSS13, Sch04a, TV22]. Asymmetry [She18c]. Asymptotic [Chi76, CKL98, CW82, Gav72a, GHK96, KMST87, KS00, MM83, Szp93, Tsa92, Fat74, Mit85, Sto79]. Asymptotically [CSW98, HM91, HP98, Bit82]. Asynchronous [AG91, CIL94, CN99, CP05, FYO^{+}15, HS06, JT73, Mal05, MPS92, MMSA93, MRRT08, Nis94]. Asynchronously [DPV15]. ATM [SV00b]. Atomic [AR98a, DGLV10, GHKT12, Hav04, LAB01, WS91]. Atomicity [GLLON18]. Attack [CS03b]. Attacks [AL18, Wat09, GMR88]. Attribute [May81, RU81, Alb85, Eng86]. Auctions [Ala14, AKSW22, DFL^{+}22, Dob21, FRGH^{+}23, GK18, KQT99, LB17]. Augmentation [BGA23, ET76, FGG08, Gus87a, HR93c, SV18, FJ81, Vel85]. Augmentations [RG77]. Augmented [BHMO22, CCKM13, GM82]. Augmented-Set [GM82]. Augmenting [KL15]. Authenticated [DS83]. Automata [BE98, BBR^{+}99, BV96, BJKP05, Boo78, BP02, BGW00, CHPW98, CPY89, CK94, DS02, DS90, Edm98, EJ99, Epp90, EWS05, GL13, Gra90b, Kar92, Kos75, LLS84, Leu98, LMPP15, Mac98, NP97, Saf06, Sei90, Tze92, aBC08, HM81, IKM85, RI89, RH78, SH85, YI83]. Automata-Theoretic [LMPP15]. Automatic [ALMZ99, BW99, Mis77, PS81, Sto79, VV17]. Automatizable [AR08]. Automatization [BPR00]. Automaton [KM94b, WS78]. Automorphism [ABL00, CB81]. Autonomous [AP06, CP08]. Autoreducibility [BFvMT00]. Autoreducible [BFS06]. Auxiliary [MV75, Rev75]. Availability [MRW00, NW98]. Average [ABN15, ANIM96, ARS97, AH08, BYCDM92, BP10, BG95b, BT06, BP81, BPP89, BP93b, CS99, CMNS01, CR22, DRST14, EP98, Fei06, Fra91, GGK92, Hir23, HL86, KRT17, Lev86, LST99, McD88, MR07a, MRSG05, PB85, PVG04, Spi73, VY23, Vio20, AA88, BS86, Iwa89, YY82]. Average-Case [ARS97, BT06, CR22, EP98, Hir23, KRT17, McD88, MR07a, PVG04, Vio20, YY82]. Average-Time [CS99]. Averaging [Dun73]. AVL [Vai90, Li86, MT86, Vai87]. AVL-trees [Vai90, Li86, MT86, Vai87]. Avoiding [DTCR08]. Aware [CKS11]. Awesome [Knu19]. Axiom [Coo78, Coo81]. Axiomatising [Ace94, Hen88]. Axiomatization [GM87, Pig91]. Axiomatizations [HvdMV04]. Axioms [Rév́85]. Axis [AES10, Tót08]. Axis-Aligned [Tót08]. Axis-Parallel [AES10].

B [BDFC05]. B-Trees [BDFC05]. Backoff [GJKP04, HLR96]. Backtracking [AIGT94, BP81, BP93b, FW74, LW98a, Nic88, Pur78, PH97, PB83b, WT89]. Backup [AK11]. Backward [FRS03, WZ15]. Bad [GHK91]. Balance [BP10, NR73]. Balanced [ANSS21, AMRU22, ABKU99, BCSV06, Bro79, CL94, CK04b, MR01b, Tsa92, Meh82]. Balanced-Replication [CK04b]. Balancing [AKU05, AKK08, AHKL07, BNFN01, BFG^{+}07, FGS12, GLM^{+}99, RS94b]. Balls [CRSW13]. Banaszczyk [BDG19]. band [BC87]. Bandit [ACBFS02, BSS14]. Bandits [ACBG^{+}17]. Bandwidth [ABK00, BNCK^{+}99, KRT00, KS96a, LVW84, Mil88, Tur86]. Bang [BKN09]. Bang-Jensen [BKN09]. Barrier [BCCK19, BGA23, Dob21, DGGW20, GHRU13, HSSS22, HKN16, HSS09, LWY20, She18a]. Base [BFP89]. Based [AAE^{+}23, AM01, ASS17, BF03b, BCHK07, BDMI14, BCF^{+}14, CDW21, CGMO14, CP12, CDH13, CadHS00, DSSS08, DCH21, EGIS98, GM01, GJKP04, GK10, GPV94, IQ19, KS97, LSH05, MR99, MR07a, MV13, MPR22, OL74, RS08, SSSS11, SS12b, TZ12, XLXJ09, FMS05, FN80, SF11]. Bases [GM82, LS95, Lem75, MN87, Dub90, SL80]. Basic [BGWZ20, CV88, DF95, Gal84a, ACM84a]. Basing [AWY18]. Basis [HKM08, HT98, Hor87, Zwi91]. Batch [Che93]. batched [BW79]. Bath [SMW07]. Bayesian [Ala14, BYGNR98, BCGL17, CDW21, DX21, FSZ21]. Be [ACR^{+}10, BI18, DFHS08, NS95, Nor09, Val02, CGG01]. Beating [GHM^{+}11, HSSS22, TV20]. Beats [KRTV18, KMZ18]. Before [KNPS24]. Behaved [ADG08]. Behavior [BBM13, KS00, OC75, Riv77, HT82b]. Behaviors [BV96, Szp93]. Being [IL98]. belts [EW86]. Bendix [KN85]. bends [Tam87]. Bennett [LS90]. Bentley [PS91]. Best [AH89, HK11, IKK^{+}11, Meh77, RT89a, RT96, YY82]. Best-Fit [RT96]. Better [ANFSW20, Alb99, ALW08, CLNS15, Che93, CO10, DH04, FM03, GW93a, GMWW17, HL05a, Rot16, CCG^{+}97, ĎG84]. Better-Than-Greedy [HL05a]. Between [CG07a, CR05, FS07b, HSV05, IY23, KRS10b, KS06, MW99a, MRS18, SG04, Wad76, AAvK^{+}13, BCG^{+}21, CKP19, FRW88, HR78, PCHM85, VW85b, ZS89].

Beyond [BPS07, KP00a, ST23, Ver09, MRSV21, PR14]. Bézout [BG96, BG95a]. bi [Kal85b]. bi- [Kal85b]. Bias [ACG^{+}16, HO14, NN93]. Biased [BST85]. Biclique [AMS11]. Biclustering [WLL08]. Biconnect [HR93c, RG77]. Biconnectivity [Hen00, TV85]. Bicriteria [FFFV19, GKPS08]. Bidders [Dob21]. Bidding [KQT99]. Bidimensionality [FLST20]. Bifactor [BA10]. Big [FSS20, Sto79]. big- [Sto79]. Bilinear [Blä04, Bsh94b, GS12, HM73, JáJ79, MV75, BLR80, JáJ80, JM86, Pro78]. Bimatrix [DFM23]. Bin [BBGR08, BCS10, CGJ78, CGJ83, EL08, FL86, GPT00, IL98, NPSW23, RT89c, Rhe90, RT93, Rot16, SvSE03, Mao93, Mur87, RT89d, RT89e]. Bin-Packing [CGJ78, GPT00]. Binary [AGMV00, AH89, Ara86, BYCDM92, CBGWW13, CFCH^{+}00, CG97, DP94a, DR95c, DR95b, Dev02, DN04, FP86, Gar74, GW77, GRS93, HF98, HST05, HR82, HW87, LMP02, LW98b, Lar87, LP94, McI74, Meh77, Meh79, NR73, PR99, RS94b, RH77, ST94, Tót03, Tót08, LW86, Sut91, Wil89]. Binary-Search-Tree [ST94]. Binomial [Bro78, Sch90]. Bins [CRSW13, NPSW23, FK88]. Biobjective [DY10]. Biometrics [DORS08]. Bipartite [AOST94, AAELvZ12, AFN07, CH82b, FGT21, GS21a, GKK13, GPST92, GMFB87, HC99, KLST01, Kho06, MSS18, NV03, Sch98b, SSSV01, Yan81a, Yao88, GK82, HK73]. Bipartiteness [KKR04]. Birkhoff [KLR19]. Bisection [CLP^{+}19, FK02, Kho06, JKLS05]. Bisimulation [FPP11, Sén05]. Bisimulations [Dob05]. Bisubmodularity [HKP14]. Bit [CK17, FNV13, GRS06, KZ07, Mul99, Vio12a, AAG^{+}89]. Bit-Complexity [FNV13]. Bit-Fixing [GRS06, KZ07]. Bit-Probe [Vio12a]. bitonic [BN89]. Bits [Ara86, BGS98, BV10, HS03, Lu14, NRS99, Vio07, BM84, CG88, LW88b]. Bitvectors [BMRV02]. Biweighted [HF98]. Black [BP15, CKPR02, CPS16, DR14, HIK^{+}11, HP10, Hir23, Vin04]. Black-Box [CKPR02, DR14, HIK^{+}11, Vin04]. Black-White [HP10]. Blackbox [SS12a]. Block [Bro99, CGS22, CMSS00, GP18, LCC90]. Blocking [BD93]. Blow [RRW01]. Blow-Up [RRW01]. Blue [AS90]. BMMC [CSW98]. Boas [Wil00]. Bodies [KN08]. Body [And99]. Boltzmann [BFKV11]. Boolean [Kad91, Aar03, BS76, BNT98, BT15, Blu94, BHIK97, BEM10, BG21, CGH^{+}88, CGH^{+}89, CF22, CS16, CK96a, CK91, DKR96, DGJ09, EP98, FMP82, Fri86a, GJW24, GKMP09, GM91c, GM92, GLV13, HR97, HW02, HS90, Kad88, KMS18, Kut91, LWY20, LW18, MI97, MiHI99, MP77, MTV10, Pau77, Pra75b, PRS97, Val86, Vio20, Zwi91]. Bootstrapping [BST95]. border [BC87]. Both [IKK^{+}11, CHPW98]. Bottleneck [EKS00, WL95]. Bottlenecks [Sch79]. Bottom [BJK^{+}94, CCH98, Jon87, SW76]. Bottom-Up [Jon87, SW76]. Bound [AG94b, AM05, BBGR08, BDG19, BHK^{+}19, BB21, Blu94, BG92b, Bsh89, BT20, CR10, CR12, Chi76, CKL98, Cyp93a, Elk06, Ezr16, FKR95, Gil98, GJM98, GJKP04, Gra90b, HR03, Joh88, Kam05, KLX10, KLSS17, Kla07b, KRT17, KM98, Lin84, LP13, Mac97, Meh77, NS22, PR14, Pau77, PR00, Pon98, Ram94, Rao11, RSY08, RV01, SS11a, Sch81a, ST23, Ste97, VZ13, Wal98, Zhu21, Cha84, Gra88, Kal85a, KK86, Lic89, MNA88, Zwi91]. Boundary [PT02, BH87]. Bounded [ABP19, BKN10, BK10a, BST23, BFG^{+}19, BH98, BNP74, BL93, BEG^{+}02, BGK^{+}18, BGSW19, BvMR^{+}00, BL02, BFL02, BOBP^{+}04, CFLY11, CG12, Cha92a, CGL97, CKLW01, CJL03, CXY04, CGYZ06, Che05, CW07, CDR10, CSS09, DFSS08, DFK^{+}19, DGJ^{+}10, DS97, DHV22, DMW05, DSY10, DPPU88, Ezr16, FRR^{+}14, FFKP18, FNR15, Gal77, GKRdW09, GKL13, Gra74, GM91c, HLV18, HB94, Hit07, KS11, KMSV13, iKK18, Kel06, Ko91, KR02, KSS20, LS13, LMS98b, LL90, LPPS17, Lu14, MS99, Mer02, MSZ11, NT05, NR73, OW91, OR21, Pat13, RS03a, Sag88, SS12a, SZ76, Wag90, vMP19, HJW85, KF80]. Bounded-Curvature [GKL13]. Bounded-Degree [CSS09, LMS98b]. Bounded-Depth [BOBP^{+}04]. Bounded-Difference [BGSW19]. Bounded-Error [GKRdW09]. Bounded-Genus [Lu14]. Bounded-Quantum-Storage [DFSS08]. Bounded-Round [BGK^{+}18]. Bounded-Treewidth [OR21]. Boundedness [HW93, Mar99]. Bounding [CI98]. Bounds [Aar06, ABP18, AB24, ACER22, AG91, ASSS16, AOT89, Alb99, ABS07, ABC^{+}13, AvMSS12, AO95, AHHP00, AdBES14, ACHAH16, AKL21, AGM13, ACH10, ACHP19, Aza92, BRS^{+}21, BBO22, BNBK^{+}89, BE98, BFL06, BC91, BPS07, BBI16, BHMO22, BAG01, BC22, BGLL99, BDL14, BKRS00, BDFP86, BGK^{+}18, BKS15, BOBP^{+}04, CW04, CK17, CG10, CEEP16, CW94, Cha94, Cha98b, CFKX07, CR22, CLS19, CFHM20, CS80, Col94, CHPZ95, CH97, CSW98, CG07b, CK96b, DP94a, DvM06, DKM^{+}94, DIM97, EPA99, Eri99, EP98, FX15, FV16, FGLS14, FLMS15, Fre76, FL83c, GR05, GG10, GG91, GG92, GG75, GMWW17, Gaz91a, GMP98, GGKT05, GadHW96, GIS77, GP18, GR21, GM18, GKS08, GT00, HHRS15, HS79, Her14, Hit07, HY20, Hya76, IKKP19, Jaf80, JTT00, JMP11, JDU^{+}74, KKP93, KKR04]. Bounds [Kel06, KLL^{+}15, Kla07a, KRT17, KR05, KMR22, KR09, KM75, KMRZ98, Kut02, Lan14, LM08a, LWY20, LM99, LWW21, Maa97, MST91, McD88, McI74, Mul99, MW20, NH90, PS97, PY91, Par77, PD06, PT09, Păt11, PVZ16, PS12, PSWvZ17, QS06, RRSS09, RS03a, RKS12, RC03, RSZ02, San09, SWK09, SLW^{+}98, SBI04, SvSE03, Sel96, Shp03, Sim97b, SS95, SF11, Tom92, Vio12a, Vio20, Wil13, Yao80, Yao91, Yun79, Zha09, dBBKB^{+}20, AARV21, AV87, BD86, CGJT80, CR80, CDR86, Cun86, FMP82, Fre81, Fri84, Fri87a, Gol81, Ili89a, Ili89b, KRR94, Mao93, Sta80, Wil89, Zel88]. Box [BP15, CKPR02, CPS16, DR14, HIK^{+}11, Hir23, Vin04]. Boxes [AES10]. Boyer [AG86, Col94, GO80, Ryt80, Sch88, TU93]. BP [EHS^{+}19]. BPP [ACRT99, BvMR^{+}00, GPW20]. Bracelets [Saw01]. Branch [FT06, Hli05, HiO08]. Branch-Decompositions [HiO08]. Branch-Width [FT06, Hli05]. Branched [KS94]. Branching [BDFP86, BRRY14, BCG20, CL94, ESY17, GL13, LSV07, New02, Pon98, IWY16]. Branching-Time [GL13]. Breaching [BGA23]. Breakers [Coh16]. Breaking [BCCK19, CIRR16, Dob21, HKN16, HSS09, KLX10, MW99a, MOOY02, She18a]. Breakpoint [AP07]. Breaks [Jay98]. Bridge [BCG^{+}21]. Broadcast [ACK^{+}98, BCS08, BNL03, De 10, EK05, GKS08, KL00, KM98]. Broadcasting [BF94, KP04]. Brother [OPR^{+}84]. Bruijn [BF94, Wil73]. Bubbles [DKKP99]. Büchi [EWS05]. Buckets [CGS99, LM94a]. Buddy [Rus77b]. Budget [DSSS08]. Budgeted [BHL18, CG10]. Buffer [ACER22, EW12, KLM^{+}04]. Buffering [AS05b, VZ13]. Bufferless [BMIM07]. Build [AN19]. Bulk [CEKP18, CHKS10]. Bundle [MSV06]. burst [JS89a]. Bursty [KRT00]. Bus [CLLS96]. Butterflies [LNRS92]. Butterfly [MS99, Tam98, WG05]. Buy [CEKP18, CHKS10, CKLW01]. Buy-and-Hold [CKLW01]. Buy-at-Bulk [CEKP18, CHKS10]. Buyer [Ala14, CDO^{+}22]. Buyers [Ala14]. Buying [BK11]. Byzantine [DS83, FM97, GM98a, HMW01, MRW00].

Cables [GJ00]. Cache

[ABD^{+}07, BGV00, BDFC05, LL00b]. Cache-Oblivious [ABD^{+}07, BDFC05]. Caching [BBN12, GKP22, KK00]. Calculation [FP78]. calculations [Sto79]. Calculus [ABSRW02, DCdP98, FLN^{+}15, Wad76, Wad78, vT04]. Call [LMSPR01, Hen80, Oya93]. call-by-name [Hen80]. call-by-need [Oya93]. call-by-value [Hen80]. Can [ACR^{+}10, Baz09, CGG01, DFHS08, KLN^{+}11, Koc92, NS95, Val02]. Candidate [GGH^{+}16]. Cannot [BI18]. Canonical [Gie95, KKLV11, Bab80, Ili89a, Ili89b]. Canonization [LPPS17]. Capabilities [CHN^{+}18]. Capacitated [ASS17, Ben95, BK15, CM99, CKR01, GKPS08, KNPS24]. Capacities [CN06]. Capacity [GH17, GJKP04, KPST94, KR78, NW98, Sch90, Hul86]. Capture [CGW16, CLX17, Imm87]. Carathéodory [Bar18]. Carlo [DKLR00, DKM06a, DKM06b, DKM06c, KKL^{+}93, SS77b, SS78]. Cartier [BGS07]. Cascade [CKP14]. Cascading [ACG89]. Case [ARS97, AV09, AF03, AMMW07, BCGL17, BCSV06, BG95b, BT06, CW75, Che08, CR22, CMS97, DMV04, DV00, EP98, Fra91, GS21b, GW93a, GHK96, Hir23, HL86, KRT17, KV98, LS77, LL00b, Lev86, McD88, MR07a, PVG04, RS91b, Riv77, Vio20, Blu86, CCG^{+}97, Cha84, Coh83, Ili89a, Ili89b, JDU^{+}74, KK86, Mao93, Pap81, RS91a, Ren89, Sta80, SS89b, Vit85b, YY82]. Cases [Gol78, NTZ16]. Catastrophically [BG98]. Categorical [YH90]. Category [Lut90, SP82]. Category-theoretic [SP82]. Catenable [BST95, KOT00]. Cauchy [CDMI^{+}16]. Caused [LS10]. Cavity [KLST00]. CDCL [MPR22]. CDCL-Based [MPR22]. Cell [CR10, CDR10, MV13, PD06, Păt11, PT92].

Cell-Bounded [CDR10]. Cell-Complex [PT92]. Cell-Probe [PD06, Păt11]. Cellular [CPY89, DS02, ELRS03, IJ87, Kar92, IKM85, Ron82, YI83]. Center [Tho05, WZ21]. Centers [BGKP99, CHPJ22]. Centrality [BPP23]. Centre [Dye86, MT83]. Certain [Sho92, ACGS88, Bin84, SS81, Yes83, Zwi91]. Certifiably [Zuc19]. Certificate [Pra75a]. certificates [CKT93, FK89]. Certified [Mal05, AW97]. Certifying [CDH13, DCH21, KMMS06, NW06, Sch13]. CFLs [KR88]. Chain [BRS98, HS82, HS84, LS93, LRS01, MNS10, Ram96, SW19, CBH84, DL80, HT82b]. Chaining [Cha94, DDMM05, VC85]. Chains [BH75, CDG^{+}06, DLS81, FW98, Kut02, Thu99]. Challenge [AMT07]. Change [BFP89, HN02]. changeover [BD78b]. Channel [BKPY18, DK15, JC91, PRS94, Wil86]. Channels [GH17, HLR96]. Character [AFB94, DLWZ14]. Characterization [AD14, APS74, AS08a, AFNS09, ÀBS04, BP90, Blä04, HMW01, HR73, HZ77, HS06, KC96, LMPP15, Lut93, Ram84, TT01, dG83, BT83a, Car79, Pri79]. Characterizations [BP02, GS02, IL81, RS96, IKM85, JS89a, KP82]. characterizes [VT89]. Characterizing [BSS14, LTW18]. Characters [BPWY99]. Charity [CKMS21]. Chase [GMP23]. Chasing [FKT17]. Chasm [GKKS16]. Chatter [BDP23]. Cheap [BK11]. Cheaply [KKN15]. Cheating [KQT99]. Checked [HS95]. Checkers [Rob84]. Checking [EKR01, FG01, GL13, GP04, KRT03, MM18, MS17]. Checkpoint [CFW93b, TB84]. Cheeger [KLL17]. Chemistry [GL23]. Chernoff [Gil98, PS97]. Chervonenkis [DSS18, DHV22]. chip [Ked85]. choice [Rei84]. Choosing [GKP04]. Chord [DDY16]. Chordal [BJ82, Gav72b, HM99, KST99b, Kle96, NNS89]. chordality [TY84, TY85]. Chosen [BCHK07, CS03b, RS10c, GMR88]. Chosen-Ciphertext [BCHK07, RS10c]. chosen-message [GMR88]. Chow [BGP^{+}23, OS11]. Chromatic [FLS04, HSV07, McD79, CG73]. Chromosomal [WW06]. Church [Ott86]. Ciphertext [BCHK07, CS03b, RS10c]. Circle [Kel06, Hsu85]. Circuit [AG94b, AGvM^{+}18, AM05, AvMSS12, AIK^{+}20, AIKW15, BFL06, BH12, BMPT97, CCH98, CW04, CR22, CH85, FY96, GJT76, IR78, Jac90, LR86, MW20, RY96, San09, Ven92, Vio14, Wil18, GB82, KS85]. Circuit- [FY96]. circuit-size [KS85]. Circuits [AGKS15, ASSS16, AHM^{+}08, AKW90, AIK14, AGS21, ALS22, BCH86, BP90, BT15, DR95a, DS07, DSY10, FRR^{+}14, GJW24, GGH^{+}16, GH98, GK98a, GK98b, GT00, GKKS16, Hås14a, Hoo90, IK22a, IPS97, ILPR81, Joh75, KMSV13, KW98, KS17b, KS19, Mah20, MD76, MVW04, MTV10, MT99, PRS97, RS03a, RSY08, Rei86, RT90, RT92, Ros18, SW93, SS12a, She09, She18a, She18c, Shp09, Tar73, Val02, Vio07, MRK88, Rap89, SV84]. Circular [APS93, Cha98a, DHH96, LSW90, Tuc80, VP03, Hsu85, Hsu95, MN88, SCH92]. Circular-Arc [Cha98a, DHH96, LSW90, Tuc80, Hsu85, Hsu95, MN88, SCH92]. circulation [BéT89]. Circumference [CXY04]. Circumventing [BBM13]. cities [CP84]. CLAP [CZ23]. Class [HZ77, KHNS05, MT78, MS08b, QS06, Rus77b, Vég16, Ye90, Yun79, KMS83, LP83b, TC84]. Classes [Bas72, BG98, BN99a, BB80, BLS84, Boo91, BL99, BFvMT00, BC03, CS99, CD72, CKST99, Dur10, FL94, Gas87, Iga77, Imm87, JS08, KAI79, KV10, KV12, KLL90, Lut90, Lut03, Mac99, Mol90, Nat92, San09, SXB83, Sie04, TO92, Ven92, Wag90, BN78, GG86, RS81, TW89, VT89, Wey79]. Classical [AA18, AGS21, BYJK08, Dru15, Gel75a, KŠdW07, Mah20, Mah22, SG04, SZ08, SS89b]. Classically [Val02]. Classification [CZ06, CM16, Dun73, HO02]. Classifiers [GMP00]. Classifying [BJK05]. Claus [Sve12]. Clause [BP93b, MMR95, Sag88]. Clauses [Gal77]. Claw [BDK00]. Client [AHKL07]. Client-Server [AHKL07]. Clique [ABW18, AM05, BY86, BX91, BHK^{+}19, CCK^{+}20, CR05, CPP16, CDP21, FGLS10, FGLS14, Gav72b, GP04, Kho06, Pul84, Ros14, BX92, Hsu85]. Clique-Width [CR05, FGLS10, FGLS14, GP04]. Cliques [Akk73, ERS20, FRS^{+}20, Gav72b, KS96a]. Clock [AHR96]. Clocks [BNS18]. Cloning [KKW04]. Clos [CG07b, DGHK98, LDHX99, NV03]. Closed [CK12, GGL^{+}13, KM08, KSS20, LS75, Imm88, KMST87, MY91]. Closeness [FL94]. Closer [CK96a]. Closest [Cla88, GRSS98, KS96b, KPV14, MS10, Mar08, Pel94, Sha85]. Closest-Pair [GRSS98, Sha85]. Closest-Point [Cla88, KS96b]. Closure [AARV21, BN99b, BGJ^{+}12, BB95, Boo79, GKV06, KMPHT14, TV91, UY91, Sch78, Tom82]. closures [BN78]. Cloud [BLNZ15]. Cluster [KNPS24]. Clustered [PSZ17]. Clustering [AC11, AAELvZ12, BL16, CCFM04, Che09, CHLX17, FSS20, GKK^{+}12, LNRW10, PSZ17, ST13]. Clusters [CJM^{+}06, Pfa83]. CM [LS19]. CM-Orders [LS19]. CMSO [FTV15]. CNF [GGGY21, Iwa89]. CNF-satisfiability [Iwa89]. CNFs [ABS07]. Co [EKLS16]. Co-location [EKLS16]. coalesced [CV83, PY88]. Coalitions [AN93]. Cocomparability [CDH13, DS94, DCH21]. Code [BGH^{+}15, FR80, Rei80, CW83, MRK88]. Codes [ADL18, AE15, AL07, BSGK^{+}10, BUV18, CGL20, CW07, CGV13, CGS22, DS07, DGY11, Efr12, Fre78, GY96, GGR11, GLS^{+}24, GHH^{+}17, HRZW20, HL04b, IJK09, JWW80, KLR19, KY09, KS13, KRZSW23, Kou82, Lar89, Mei09, Mei13, ML00, MN01, PR99, ST23]. Coding [BSGH^{+}06, BE17, CLL13, DGL23, DFVW99, DG93, GMY12]. Coefficients [BGS07, Gan95, Lip78, Str74, Fri87b]. Coffman [BT94, ST00]. Cographs [CPS85]. Cohomology [CEN12, GQ17]. Coin [ACG^{+}16, AN93, BHMO22, CK17, HO14, HT17, LSS^{+}21, LT82, PTW11, RSZ02]. Coin-Flipping [AN93, HT17, RSZ02]. Collapse [HHH98, HHH05, KW98]. Collapses [GHJ^{+}24, HNOS96, KvM02, Ogi98, Kad88, Kad91]. Collapsibility [Che08]. Collecting [ABHK11, AABV98]. Collection [Lar77]. Collections [BBD^{+}86, LSH05]. Collective [RSZ02]. collide [AKR95]. Collision [AS90, Pel94]. Collision-Free [Pel94]. Collisions [HHRS15]. Collusion [GKW20]. Color [VP03]. Colorability [KLR03]. Colorable [AK97a, KS17a]. Colorations [FW76]. Colored [KRSV08]. Colorful [DLM22]. Coloring [AK97a, BX91, BEK14, CLP20, CFK^{+}07, CSZ24b, CSZ24a, CH82b, CPPW12, CDP21, DMR09, Gav72b, GKP04, GHS02, GHH^{+}17, HMR98, Hol81a, KS17a, KLS^{+}24, NV03, PS97, Sch98b, BX92, GK82, MT84]. Colorings [BCG^{+}21, BC22, BDGJ99, CV10, Eft16, ELRS03, GGJ16, GLLZ19, Mol04]. Colors [GMP05, HMR98, KS17a]. Colouring [HCD89]. Column [BDMI14]. Column-Based [BDMI14]. Combination [DFHS08]. Combinational [HS79, Pau77, Zwi91]. Combinatorial [Ala14, AFNS09, AKSW22, AR02, BATS11, BŠVV08, BK07, BK11, CG05, CL79, CS05, DHR20, DR06, Dob21, EK05, EGH07, FGL16, FW98, FW74, GS00, GMRZ13, JWW80, Lar14, LL00a, LB17, Mei09, MMM22, Mur90, Mur95, OPS04, SH96, SC79b, KP82, SS89b, ZR79]. Combinatorics [ADL18, NPSW23]. Combinators [YH90]. Combined [BGWZ20, MRRT08]. Combining [JNS06].

Comilexity [Iga77]. Comments [AI77, Fat74]. Commit [GHKT12]. Commitment [CK17, WS91]. Commitments [HNO^{+}09, HHRS15, LPS20, PR08a]. Common [AHU76, CHK03, HSSS22, HT84a, JL95a, KMR87, ASSU81, MS84, Nor89, SV88]. Common-Face [CHK03]. Communicating [BOHP14, BKO88, GR85]. Communication [ACT00, AKSW22, BGKL03, BYII93, BYJK08, BBCR13, BTY94, BPS07, BH12, BNT^{+}19, BE17, BGK^{+}18, CR12, CCM^{+}19, CFK^{+}21, CKL98, CD18, DF92, FKNN95, FX15, GKR21, GKRdW09, GKK^{+}09, GH17, GJLR97, Goo99, GP18, GPW18, GPW20, GR21, HHRS15, HR93a, HS03, Hu74, HY17, HPZZ21, KLL^{+}15, Kla07a, Kla07b, KOR98, LPSPP05, LT91, NW93, PE86, PU87, PVZ16, PR20, PRS97, RT10, She12, She16, She18b, SZ08, Val82, WS91, dW03, AMS84, CG88, EMC85, HCAL89, Vit88]. Communication-Efficient [Goo99]. Communication-Space [BTY94]. Communication-Time [PU87]. communications [RS85b]. Community [BKM19, BCN^{+}20]. Commutativity [Kol18, JáJ80]. Compact [AAK^{+}06]. Compaction [Vik03, Vik04, BC84]. Compactness [GP01]. Comparability [EHSS99, HM99, Spi85]. Comparable [Ogi95]. compared [Lon85, Lon88]. Comparing [LL00a, LL00b]. Comparison [AFG^{+}14, FT97, GGK92, HS02, Ken76, LMS98a, Val75, AA88, AV87, BH81b, RS78]. Comparison-Driven [AFG^{+}14]. comparison-sorting [AA88]. Comparisons [ES81, HH81, SW79]. Compasses [ISK^{+}12]. Compatibility [KL72]. Compatible [BT80a, CPPW12]. Competing [FKT17]. Competitive [BBN12, BLN01, CMS18, DP94b, DXSC94, FFK^{+}98, FKK^{+}08, GRS05, GGK16, IKL^{+}04, KP00a, LRWY99, Sit14, VCJS22, IKP96]. Competitiveness [DM97]. Complementary [GMSV15]. complementation [BCD^{+}89a, BCD^{+}89b, Imm88]. Complete [AG91, BH77, Blä04, BG95b, BCC^{+}09, BHT98, CM87a, CGW16, CZ06, CGGR16, DT97, FF93, For79, GH92, GJT76, GK12a, GLT80, GGM11, GPSS06, GPSZ08, GS02, HvdMV04, HR97, KAI79, KLN15, Kou77, Koz09, KD79, Lev86, LSS13, LZ00, RT81, RT89a, RT89b, Rob84, Set75, STP94, SS05b, Sin97, SOAD09, Szk99, WC90, Zuc96, Gra88, HM81, KMZ87, Lub81, Rap89, SS81, Yes83, DT97, GN16, Gra94, KST93]. Completely [ABP19]. Completeness [AM80, Coo78, GKV06, HP10, HN77, Hol81a, KKMO00, KM81, MP87, PS02, Coo81, Dob89, DF95, HN79, Hol81b, Pet83]. Completion [BSS21, CMNS01, IKS10, JK86, KST99b, LP03, Li20, Sch04b, VHPT09, KS96a, KN85]. Completions [BN82]. Complex [BCGW93, BGGS20, CCL13, Eps79, Pan94, PT92, Rei99]. Complexes [Hav04, STP94]. Complexities [ACG^{+}05, HHRS15, dW03]. Complexity [ABSRW02, ABSRW04, ANIM96, ABKPM09, ARS97, AFL08, ABL00, AF03, AKL21, AKSW22, AHLM07, ACHP19, ABSD^{+}98, BGKL03, BYJK08, Bas72, BDKR05, BC91, BNT98, BPS07, BH12, BG94, BSS08, BS00, BV97, BG92a, Blä04, BV99, BEGJ00, BB80, BLS84, Boo91, BL93, BCHS94, Bra15, BGK^{+}18, BT10, BT80b, BL99, BD78b, Bsh94b, BFvMT00, BFL02, BJK05, Bür00b, Bür00a, BC03, BLMW11, CS99, yCLZ00, CLX11, CCL16, CRZ22, CR12, CM13b, CGR12, CKP19, CBGWW13, eCY91, CG07a, Che08, CDO^{+}22, CGS22, CK12, CT90, CK95, CCC^{+}13, Col94, CHPZ95, CH97, CK91, CLLS96, CW82, CZ06, CM16, DJP^{+}94, DGP09, DO04, DGR99, DFG13, DF92, DFVW99, DHHM06, DGP07, DRTC21, DS90, DF88, DGH98, DGJ09, Efr05, EJ99, EY10, EIS76, FKNN95, FX15, FPV18, FNV13, FLN^{+}15, FRGH^{+}23]. Complexity [FG04, Fre93, Fre76, Fre78, FL94, GGP11, GG91, GG92, GJ75, GT01, GW93b, GKRdW09, GKK^{+}09, GGK92, Gil77, GL77, Gli03, GKP04, GGJT10, GJ14, GOP98, GS06, GK05, GJW18, Grä90a, GKS94, GK98b, GS88, GLV13, HF98, HSV07, HS79, HT98, HHKW07, HJ90, Hir80, HS06, Hon82, HM73, HY17, HPZZ21, HY20, HS90, HMRS98, Imm87, Imm89, IL86, JLL12, JK82, JR94, JS81, JL95b, Kam05, KP85, KY10, KKR06, KLL^{+}15, KL72, KH83, KSS93, Kla07a, Kla07b, Ko91, KT94, KKR17, Kum96, KR23, KLL90, Lad77, LM08a, LDRT24, LMM99, LSSV08, LV95, LRR17, LW18, LT91, Lut90, Lut03, Maa86, Maa97, Mac99, MM18, MV75, MNS10, MT83, MNR90, Mil75, MT99, Mil93, MS07b, Mol90, Muc76, MRS18, MS01, MS08b, NHL82]. Complexity [NS09, NT05, NW93, OR21, PT02, PST00, Pan00, PS77, PY87, Pap92, Pau77, PR00, PVZ16, PR20, Pro76, PB83a, PRS97, Raz03, Ren92, Ris85, RH93, Ros14, SV08, SXB83, She12, She16, SZ08, SI87, SW07, Str76, Str83, SSS91, Tho97, TS81, Tra72, Tun91, Vad01, Val79, Ven92, Ver09, Ver05, Vik03, Vin04, Vio12b, Wal98, Whi90, Wil18, WS91, Yao85a, Yao89, adHW87, vzGS00, AAG^{+}89, AA88, AV87, Bab80, BM88, Blu86, BCN81, CI83, CG88, CSS90, Eng86, Fre81, GMR89, Gra84, Har85, HCS80, HRB87, Huỳ86a, IL82a, Ili89a, Ili89b, JáJ80, KS85, Kut91, Lag85, Lic87, Lic89, MT84, MC87, MS84, MNA88, PE86, Pri79, Pro86, RS81, RL88, Ren89, RY87, SS90, SSY82, SS86b, TW89, VT89, Yao79]. complexity [YY82, Zwi91]. Complexity-Restricted [KT94]. Complexity-Theoretical [KY10]. Compliant [FHS96]. component [Ham83]. Components [FRT93, GI93, Gaz91b, HT73, KNP99, Kei85, La 00, Pac74, SW99, Kao93, NS80, vzGS98]. Composable [CLP16]. Composing [GJ85]. composite [FS86a]. Composites [CP19b, Gop08]. Composition [BT80b, DDPY08, DH13, GMWW17, GK96, Jay98, KU11, KLR10, PTW11]. Compress [AIKW15, BBCR13]. Compressed [BLR^{+}15, DKM06c, GV05, Loh06]. Compressibility [BL02, HN10]. Compressible [CKL^{+}13]. Compressing [She18b]. Compression [BST95, CLM^{+}20, Dru15, FFFV19, FNV13, GS91, IW22, KM99, Lu14, RS08, SS05a, SR97, Yac98, Yao85b]. Compressions [KLST00]. Comput [BX92, BG96, CCG^{+}97, Col93, Coo81, FJ90b, HH80, HN79, Kad91, Lan91, Lon88, RS94a, SWCP96, Vai90]. Computability [ADH97, DM97, DHN02, LV22a]. Computable [Bra08, CI83]. Computation [AvDK^{+}07, ABO08, ALMZ99, ABT21, AL10, Aza92, Bac21, Ben89, BG95a, Bis78, BCL75, CKM^{+}14, CPV22, CCM^{+}19, CP12, CGP^{+}20, DS76, FRGH^{+}23, GR02, GR05, GP96, GGL98, HIK^{+}11, HRTS03, HHT97, HR97, HS75, HS82, HS84, IKOS09, IKK^{+}11, KW97, LC89, LS90, MS08a, MN01, MW93, Nis94, Reg96, Reg04, RT92, RRR21, Sav74, SW19, Sei77, Sim97a, Vaz97, Wra78, Yao91, vT04, BG96, BC82, FRW88, GS87, Sut91, VSBR83, VW85b]. Computational [AK10, AHLM07, BN10, Bür00b, Bür00a, CCH98, CLX11, CP09, CK95, DW82, DGR99, DKK^{+}19b, FV98, GT01, GW93b, Gil77, GOP98, GKMP09, GKS94, HNO^{+}20, HSV07, HR78, KMSV98, Kim97, KS19, Lad77, Maa97, MV75, Maz76, Mil75, MT99, Muc76, RS03b, Ren92, SJ05, SS86b, Str76, Str83, Tra72, Vad06, Ver05, Vik03, Wad76, WW75, Wil00, BLR80, Gra84, HCS80, HRB87, Lag85, Lic87, Pro78]. Computationally [FRR^{+}14, Mic00, Sah74]. Computations [AC78, BBD^{+}97, BL98, BO10, BC03, CGGK95, CLLS96, DHV22, EKR01, GJMM00, Gef91, KKMO00, Mah22, MST91, MV13, PP95, Pan00, SLC91, SL96, Wal98, CC86, Hel84, Ked85, KF80, Oya93, Pro78, Pro86, Rei84, vzG86]. Compute [BC76, FHOS97, GR15, Str74]. Computed [BI18, NS95]. Computer [AEM^{+}11, AK07, ACR^{+}10, ANR09, BBS13, BSO14, BCAKT23, CR17, DT08, DMK19, GS73, HIMW16, KRS20, KR78, LMOT14, Lav73, LU10, MS87b, MV18, Rou15, Sho97, NS80, Sto79]. Computers [CK96b, GJLR97, MPS92, AHMP87, CSS90, adH86]. Computing [AGK^{+}09, AA18, AKK^{+}09, AAS97, AMS98, AOS06, AAvK^{+}13, AAKS14, AKS09, AM05, ACHCP21, AGKM16, AIM16, Bab06, BOC92, BBBV97, BJK^{+}94, BP98, CL13, CCRZWN22, CDG11, CDL86, CEG^{+}93, CJL03, CHW13, CS07, CK93, CFPS12, Col74, CK94, DGP09, DKI18, DLS81, DHP13, DKM06b, DKM06c, DKSS22, DLR95, DF88, DGH98, EV21, EKL10, EvMP^{+}12, ESY17, FGKO08, FMP^{+}03, FRPU94, FLMS15, FIM^{+}14, GKM^{+}21, GRS05, GM91b, GLLON18, GJ14, GSY95, GK98b, Gro03, Gus91, HKRT95, Har99, HM95, HNOS96, HHKW07, IKM^{+}12, Jai07, JKLAR90, KMR87, KWY98, KW00, KM94b, LY05, Loo83, MW99a, MS16, McK79, MMS17, Moi16, Mur90, Mur95, Pac74, PKS^{+}13, PB83a, Rap89, SK18, SZ99, VNA20, WZ06, dOS94, dBG08, vzG87, vzGS98, Ili89a, Ili89b, KB79, PE86, SW85]. computing [SWCP96]. Concave [GMSV15, HC99, YS93]. Concentration [ESV20, Rao11]. Concentrators [Sch90]. Concepts [BGGM98, KMMS22, LV91]. Concerning [BS00, IJTW95, IM83, RY87]. Concurrency [KP85, Pap83, PY87]. Concurrent [Ace94, CKPR02, CIRR16, DCdP98, DS97, EHS12, HS85b, Hen88, IISI05, KPST94, LPS20, PR08a, CRS87, FRW88, Man86, Mit85, RSL84]. concurrent-write [FRW88]. Condensing [RSW06]. Condition [AF98, Saf06, Var97, Oya93]. Conditional [AARV21, CRS15, CFGM16, DMR09, GHK96]. Conditions [ACFW98, Hen92, MRRT08, Par80, Par98, SWPL11]. Conductance [CHS12, CHHKM17]. cone [MD73]. Confidence [LS75]. Conflict [CFK^{+}07, DK15, ELRS03]. Conflict-Free [CFK^{+}07, ELRS03]. Confluently [KOT00]. Congested [CDP21]. Congestion [ADG^{+}11, AZ07, BLNZ15, CKS09, CGGS19, Chu16]. Congruence [BK10a]. Congruences [Dob05, FHK^{+}88]. Conjecture [AWY18, BDG19, BKN09, BP20, BMM21, CS04, CMSS00, Col00, DT97, FFK96, GL23, JT86, Juk06, Lov21, ST00, Ver09]. Conjunctions [GHRU13]. Conjunctive [BLR14, CvdM00, CD16, GS17b, JK83]. Conjunctively [BHL95]. Connected [CT00, CVV03, CV14, CFHM20, FL12, GG12, GI93, Gaz91b, KNP99, KSS93, KN05, La 00, PB83a, RZ10, SW99, ZH05, CSS90, DM88, GB82, Ham83, Kao93, NS80]. Connectedness [SSS91]. Connecting [CPR11]. Connection [CK96a, Mic04, RS94b]. Connections [KRT00]. Connectivities [CLL13, LY13]. Connectivity [BBRS98, BE98, BGA23, Cha06, CPR11, CC23, DV00, DP20, Edm98, EPA99, EGIS98, Eve75, ET75, Fre97, FNR15, Gal80, GI92, GIP20, GH98, GKMP09, HRW20, Kao95, KKKP04, KSS08, KKL04, LW98a, MRSV21, NGM97, Pat13, Ros18, Sch13, SV18, Tri08, BKRU94, CKT93, KS91, LW88a, Sch79]. Connectors [Pip78, Ngo05]. connects [RW05]. coNP [BCMW17]. Conquer [CKL^{+}09, LR89, ACG89, SR83]. Consecutive [ABH98, Kou77, SW94, Vel85]. Consecutive-Retrieval [SW94]. Consensus [AT98, ACH10, CHJT04, CIL94, DDR07, GHKT12, HY87, Jay98, KWY98, MR02, Rup00, SWK09, Zhu21, AW96].

Consequences [KW98, NS22]. Considering [EW12]. Consistency [AFL08, ACFW98, BG22, FRS03, GS00, GS17b, Koz21, RSL84, SWPL11]. Constant [ABN15, ABO08, AIKW15, BGK20, BMKM07, BOC92, BKPY18, BH80, BP90, BSW14, BM99, CPV22, CSV84, CKS09, CL19, CVY19, CGS22, Chu16, CGG^{+}97, CGH^{+}98, CDG^{+}06, CDP21, DSS90, EN19, FR06, FSS20, GGK16, GMM09, HH81, HO14, HW93, Koc92, LW98b, MOOY02, Meh18, MBCV98, MR08, NPSW23, Pag01, PP08, PPR09, Rao09, Rei99, RRR21, SW93, Saw01, She18a, She18c, Sie04, Vio07, WROM86, YYI12]. Constant-Competitive [GGK16]. Constant-Depth [She18a, She18c, Vio07]. Constant-Factor [BMKM07, BSW14, MBCV98]. Constant-Query [CVY19]. Constant-Round [CPV22, CDP21, RRR21]. Constant-Size [FSS20]. Constant-Time [CGG^{+}97, Sie04, YYI12]. Constant-Time-Maintainable [HW93]. Constrained [AMS84, AW01, ABL^{+}02, BJM00, CW98, EMSV12, GW95, HL04a, HL05b, MR07b, QS06, Sve11, Dev12]. Constraint [BYGNR98, BK16, BP20, BL00, BMPP19, BGWZ20, BG21, BGG22, BM14, CCPV11, CRZ22, CM13b, Che08, CVY19, DR13, FV98, FH06, GJK^{+}22, GS17b, KOWZ23, MS07a, Mol03, MS07b, Vik04]. Constraints [Ado77, BC10, BGG22, BJK05, BD06, BH13, CFW93a, EHPR17, EHJM00, GG75, GJ75, GMN09, GL95, GGM11, HLL95, JL77, LNSS09, LP03, LR21, LS97, MSS04, PT87b, SLC91, SU05, WL95, Jaf85]. construct [LR88b]. Constructibility [Gef91]. Constructible [KKPV07]. Constructing [AdBMS98, BPWY99, EW86, Fri86a, GLN02, HRV13, HP99a, HP00a, HSS09, KM00, LP95, LS18, Coh98, HY88, JM96b, Yao82, EOS86]. Construction [ABCP98, BATS11, BFU94, BFSU98, Cha00, CW83, CH03b, ES05, HPM06, HO02, LPSPP05, Mei09, ML00, PR99, PR00, Plo76, RS10a, RS22, DMT88, MR80]. Constructions [BCF^{+}14, EP04, GGKT05, HIK^{+}11, Lau83, NN93, PR08b, GI80, Gus83, Kla84, KMZ87]. Constructive [AEZ20, LM15, MY91, PPS00, Rot17]. contain [JK83]. Containment [CKS02, CvdM00, SH85]. Contention [BKPY18, CVZ14, CadHS00, FSZ21, GMR98, RU98]. Context [AP72, Boo72b, GL77, Gli03, Gra74, Gre73, Gre74, HR78, JM96a, Kos75, Sav73, SS77a, Tai80, HC83, Hum89, ORW85]. Context-Free [AP72, GL77, Gli03, Gre73, Gre74, HR78, Kos75, Sav73, SS77a, Tai80, HC83, ORW85]. Context-Sensitive [Boo72b]. Contexts [LSV07]. Contextual [LS22]. Contingency [CDG^{+}06, CDR10]. Continued [Jus92, Str83]. Continuous [CCWZ18, DDR07, Dun73, FPP11, HY20, LV22a, LM82]. Contracted [AP07]. Contracting [MS08a]. Contraction [CLV23, MR91]. Contractions [CCH^{+}16, Kao98, Sch13]. Contracts [DRTC21]. Control [BBR04, FMS05, JT73, KP85, KCH82, LMSPR01, Pap83, PY87, SW10, HR87]. Controlled [BGV00, YR92]. Convergence [Ald75, CP05, CP08, EHS^{+}19, FRV10, GM00, RT89a, RT89b, RT96]. Conversion [Rév́85]. Convex [AS96, ABL^{+}02, AFB96, AS97, AST97, BD92, BDD^{+}07, BY98, Cha84, Cha92b, CHW13, CMS14, DHR20, EW03, Eri99, GKL13, GAOPL^{+}21, GHK91, IS90, KRS10a, KM97b, KN08, KS86, PR99, Pfa83, PT92, RS92, Rot17, Sha87, Vég16, ES88, ES91, MD73, RS94a]. Convexity [PRR03]. Cook [Sol76, dB84]. Cook-Reducible [Sol76]. Cooling [CV18, SMW07]. Cooperative [GRS05].

Coordinated [DFK^{+}19]. Coordinatewise [CMS14]. Coordination [Alb09, IR96]. Cops [AGG^{+}19]. Core [EHS^{+}19]. Cores [EIM01]. Coresets [Che09, DDH^{+}09, FSS20]. Corner [LV06]. Corners [KKN15]. Correct [DS76, Ryt80]. Correctable [CGS22]. Corrected [BGS18, VR89]. Correcting [AP72, CGdW13, Mei13, WS78]. Correction [Col93, EiC17, SS89a, MR80]. Correctness [AF98, FS81, Ger76, GK10, Rus77a, Rut90]. Correctors [CGR18]. Correlated [RS10c]. Correlation [AAvK^{+}13, AAELvZ12, BGG^{+}19, Coh16, Hås14a, HNO^{+}20, HL14, KLS^{+}24]. Correlations [RT10, Sut91]. Corrigenda [HN79]. Corrigendum [AFGN23, Coo81, DCH21, Fri02, HH80, KV12, PA80, RS22, Rie02]. Coset [FIM^{+}14]. Cosine [PM03]. Cost [AHK^{+}10a, AAZ16, ADK^{+}08, BY98, CCKM13, CVV03, CV14, CS24, DM97, DPV15, DH04, EL08, EHJM00, FKLS10, GW77, GPRS11, HW95, IMP14, JV08, KLSvZ08, LW98b, MMP08, OPR^{+}84, SV18, Vég16, WLB^{+}99, BéT89, BS86, Kni88, Ron82, SS86b]. Cost-Distance [MMP08]. Cost-Sharing [GPRS11]. Costs [ABD^{+}05, CFLY11, CW94, GY96, GMY12, LMP02, Sch04a, BD78b]. Counter [GL13, IJTW95, Var00, Vit85a, WS78, Vit85b]. Counter-Automaton-Recognizable [WS78]. Counterexample [Che12, Raz11, Car79]. Counterexamples [Ben95]. Counterfeit [LT82]. Countermeasures [AL18, BD81]. Counters [IJ91]. Counting [APS93, AS10a, AM13, BŠVV08, BKPS06, BDGJ99, BC03, CDR10, DLM22, Dye91, DFJ98, DFJ02, DR13, ELRS17, FG04, FR24, GGJ16, GGGY21, GGM12, GGM^{+}15, GLLZ19, HMRS98, IL86, KRSV08, KT15, Lad89, LRS17, LSS13, PST00, PT07, PB83a, Rus81, SVV12, TO92, Ull16, Vad01, Vin04, WW13, BCD^{+}89a, BCD^{+}89b, Iwa89]. Cover [AAA^{+}09, AKL21, BBFT23, BHI18, BHNW23, CDH13, CPP16, DF18, DGKR05, DCH21, EW03, GMPT10, GKPS08, Hal02, HL05a, KKMS11, MRR15, RV98, Tho97, Hoc82]. Coverage [DFHS08, GL17]. Covered [CEG^{+}94, Efr05]. Covering [AK11, ABD^{+}05, BCS10, CN06, FLM^{+}22, Gav72b, GGL^{+}13, HOS00, ILPR81, KY15, LC05, Maa86, Mic04, RV98, RT89c, Rhe90, Sri99, BM88, Hoc82, LP83b, Pul84]. Coverings [BI06]. Covers [ABCP98, Bie90, Jac90, KS94, Man08, TT13]. CREW [Kut91, Nis91, vzGS00]. Criteria [GM00]. Criterion [HN77, HN79]. Critical [GP18, KN05]. Cross [AM06, Kao97, RW05, Sut91]. Cross-connects [RW05]. cross-correlations [Sut91]. Cross-Tabulated [Kao97]. Crossing [CM13a, EV21, LWY20, SW06a]. Crossing-Free [SW06a]. Crossings [EGS10, EGS03]. Crossword [Maz76]. Cryptographic [BDP02, CK17, CHN^{+}18, GGKT05, HHT97, HN10, HY20, PR08b]. cryptographically [BM84]. Cryptography [AIK06, BOV07, BCF^{+}14, CGMO14, DFSS08, DDN00, GKK^{+}09, KZ07, RS08]. Cryptosystems [GS88, NS12, MRS88]. CSP [Aus10, BKN09, DGJ09, GHM^{+}11, JKK06]. CSPs [CZ23, DKK^{+}19a, HKP14, KKMO07, KTZ15, KKR17, KMR22, Koz21, TZ17]. Cube [DMS24, FW76]. Cubes [MS04]. Cubic [AK11]. Cuckoo [FPS13, FMM11, KMW10]. CUR [BW17]. Curing [KMP^{+}20]. Current [ABW18]. Curvature [AW01, ABL^{+}02, GKL13]. Curvature-Constrained [AW01, ABL^{+}02]. Curve [AM01, JB90, KS05, O’R86]. Curve-Sensitive [KS05]. Curved [HSY22]. Curves [AHPY13, AEZ20, Sho92, Wal98, vzGSS03, MY91, vzGS98]. Cut [AN06, AMS11, AI77, AdW22, Ben95, BNS18, CHM13, CCH^{+}16, GVY96, Had75, NZ01, Pol95, Tre12, Rei83, Shi80, AHK10b, CMST85, Kar99b, KKMO07, AR98b]. Cut-Norm [AN06]. Cut-Trees [Ben95]. Cuts [BK15, BD93, CEN12, CEFN23, CKRV15, CX18, DJP^{+}94, GSV99, KYN07, KM97a, KPST94, LSW90, PB83a, SV95]. Cutset [CHM13, MR14]. cutsets [Sha79]. Cutting [BEGJ00, KKN15]. Cuttings [HP00a, KS05]. Cycle [BFKV11, CGP^{+}20, DS94, FMS02, HKM08, IS90, Man08, Tho97, Hor87, SCH92, Sys81]. Cycles [CGYZ06, CH87, DVW21, DHK11, DFJ98, Gab07, SW06a, Tom92, Vik03, WG05, dVE10, MNS86, SSY82, Sys81]. Cyclic [SI87, Ata82]. Cylinders [ECM03]. Cylindrical [ACM84a, ACM84b, Pri86].

d [Kap00, NS08, AK11, BHI^{+}13, BMKM07]. Daily [CKLW01]. Dantzig [KY15]. Dart [CFSZ02]. Dart-Free [CFSZ02]. Data [AAK^{+}12, AHS92, ABL09, AFG^{+}14, And99, ALRS98, BRS08, BJS97, BNS^{+}21, BV93, BHI18, Bit79, BST95, CJM^{+}06, CGdW13, CTX10, DORS08, DHR97, FFFV19, FKSV02, FKM^{+}09, FSS20, Fre85, Fre97, GRSS98, GPV94, HU75, Hen00, IJ94, JW21, JR13, KMR88a, KV86, Ken76, KMV91, KKW04, KS20, LWY20, LM82, LNW19, LKS97, LP13, LRWY99, MS87b, NSW11, PT07, Ros80, Ros82a, Ros72, Sco76, Sed78, SR97, TV99, Vio12a, Wil85a, Yac98, BT83a, BT80c, Cha88, Don79, Fre84, FS89, Fre81, Gus88a, RD81, RS83]. Data-Stream [FKM^{+}09]. Data-Structural [BST95]. Database [BMSU81, CM87b, HW93, KKPV07, NTZ16, Sag88, TA99, CBH84, Hul86, RSL84]. Databases [BRV16, GKV06, LSH05, MW95, HITK79, IIK85]. Datalog [FV98, Mar99]. Davis [BKPS02]. Deadline [CLT05, LP03]. Deadlines [FR94, VCJS22, BD78b, GJST81, Sim83]. Deadlines* [GJ77]. Deadlock [CG94, Gol78, TU81, TS81, Yan82]. Deadlock-Free [CG94, TU81, TS81]. Debaters [CFLS97]. Debreu [Jai07]. deBruijn [GH97]. Decay [BGG^{+}19, KLS^{+}24]. decaying [CTX10]. Decentralized [BEGL24]. Decidability [Ges02, IJTW95, Wey79]. Decidable [BJKP05, QY21, CK87, Gur82, Oya93]. Deciding [BV99, CJK^{+}22, GGP11, GGM11, HR93b, Sei90, Web93, FG82]. Decision [BG94, BM02, DLM22, DS02, DGP07, EP98, FPP11, FR75, HITK79, JS05, KM93, MNS10, MPR22, Nis91, NRS99, OS07, YR80, GHY89, IR96]. Decodability [CW07, CGV13, GLS^{+}24]. Decodable [DH13, DS07, Efr12, KY09]. Decoding [DHK^{+}21, GGG18, GGR11, GL21, IJK09, KY10, KS13, KRZSW23, ST23]. Decomposability [Wil90]. Decomposable [BRV18, DK14, FJ90a]. Decomposing [CD16, Kei85, Web93]. Decomposition [AGG^{+}19, AES99, BD92, BKRS00, DT97, DKM06c, GZ05, KLST01, KT77, Sky76, ACM84a, ACM84b, KMR88b, WT79]. Decompositions [AN19, AFGN22, AFGN23, BEGL24, BCG^{+}21, Bod96, BW17, CGP^{+}20, GRS16, HiO08, Pri86]. Decreasing [RT89a, RT89b, RT96]. Dedicated [Bab06, FGKO08]. Deferral [GGK16]. Deferred [KMR88a]. Defies [MRS08]. Definability [FG01, RS07, Zim98]. Definable [AF23, EM03]. Defined [BFvRV15, MiHI99, SZ76, LW88a]. Definitions [ACFW98, Ven92, Car84, Coh83]. Degeneracies [EC95]. Degenerate [BGP^{+}22]. Degree [ASSS16, ACOH^{+}10, BKN10, BFG^{+}19, BCH97, BDGJ99, BT20, CFLY11, CXY04, CKP86, CSS09, DF18, DPPU88, Fei06, FNR15, GJW24, GNS20, GHH^{+}17, Hås88, HB94, KS11, KRY96, KB76, KLR03, KR02, KR05, KP98, KSS20, Kun74, LNSS09, LS13, LMS98b, LPT98, MOP22, Mol04, MR08, Mur95, Sak90, Sch81a, Sén05, She13, Tom92, CMST85, KMZ87, Mil88]. Degree- [ASSS16]. Degree-Bounded [BKN10, FNR15, KR02]. Degree-Driven [LPT98]. degree-three [Mil88]. Degrees [AS86, CJL03, CGYZ06]. Delaunay [BWY15b, BFvRV15, CD03, CW98, LM12, Xia13]. Delay [AFHB^{+}00, AHR96, AH08, Lar89, Ron82]. Delays [GKP22, LS10, RU98]. Delegating [RRR21]. Deletion [AHS92, Asa87, KD79, MSV87, Yan81b, Yan81a]. Deletions [Ber16, Meh82]. Delivery [CM99]. Demand [BNL04, BN10, CDO^{+}22, Smi76]. Deniable [SW21]. Denotational [Wad76]. Dense [Bar18, DFJ98, GM91a, GR11a, GRS16, Kho06, SU77]. Density [ABS07, BH77, BG22, BH13, mCL04, DHN02, Fri90a, HPK14, HPQ17, KM81, LM94b, LZ00, Mor82, RS99]. Dependence [CHHKM17, CKS17]. Dependencies [BH81a, FMUY83, BV84, CV85, FT83, HITK79]. dependency [FT83]. Dependent [BNMSS98]. Depth [AAK90, ASSS16, AHP08, BCH86, BPU92, BP90, BT15, Bor77, BCE95, BOBP^{+}04, CCK^{+}22, CHPJ22, CSV84, CP19b, CLS19, Coh95, CH85, DS07, DSY10, FLMS15, Gar74, GH98, GM91c, GKKS16, Hås14a, JS05, Kao95, KMSV13, KLSS17, KS15, KS17b, KS19, KR23, LSS^{+}21, MP77, Pio04, RS03a, Rei86, SW93, SS11a, SS12a, She09, She18a, She18c, Shp09, Tar72, Vio07, dOS94, dBG08, Hag90, HY88, JM96b, Smi86, VW85b, IPS97]. Depth- [ASSS16, FLMS15, KMSV13, SS12a, ASSS16]. Depth-2 [BT15, She09]. Depth-3 [SS11a, Shp09]. Depth-First [AAK90, Kao95, Hag90, Smi86]. Depth-Universal [CH85]. Depths [Dev98, GH97]. Dequantizing [GL23]. Derandomization [BOV07, BCM99, CR22, DM97, GM20, GKSS22, Han96, HKN16, MRSV21, Wil16, vMS05]. Derandomized [IJKW10]. Derandomizing [HU22, MR99, SW06b, vMP19]. Derivation [GL77, LS75]. Derivatives [Zuc19]. Descending [CJ14, NP97]. Descent [AK09, BCLL21]. Description [FW74, GHSZ08, FL83b]. Descriptions [GW93b, JL77, SS77a]. Design [Alb09, AAZ16, ADK^{+}08, BKN10, BCGL17, BKV11, Bro99, BT80c, CDW21, CEKP18, Cha89, CFLY11, CHKS10, CS03b, DLL^{+}03, DD13, DR14, EGOS07, FNR15, GKLV14, GKR12, KKL04, KNPS24, LNSS09, LS13, LZ22a, LZ22b, LPT98, MMP08, Tet99, TS81, YR92]. Designing [CRV10, CCH^{+}16, CS19, Fus88, GMP00, ACG89]. Destructive [LM99]. detecting [AB96]. Detection [AS90, AT98, BKM19, BCN^{+}20, BCR16, DVW21, HK00, KSS08]. Detectors [GHKT12]. Determinant [Bjö14, CS07, DPPR20, HT15]. Determinant-Preserving [DPPR20]. Determinantal [Mur90]. Determinants [MR12, Mur95]. Determine [MPS^{+}94]. Determining [ARS97, Chv77, Eve75, McD79, Pat13, Rup00, BOFKT88, CG73, Kun93]. Determinism [DF92, GGM12]. Deterministic [AM96, AHMP87, BEG17, BN99b, BHI18, BHNW23, BFG23, CGH13, CKP19, CK13, CCRM86, CC23, CD18, CDP21, DK15, GRS06, Gas87, GKM^{+}21, GLPW16, Gli03, GPW18, GS02, Gre74, HH72, HKN21, HB94, Hon82, IKS10, JPZ08, KYN07, KZ07, KMSV13, KP04, LS02, Mac98, Mal05, MV13, MRS18, PR17, PPS00, SVV12, Vis91, AA88, FG82, Gur82, KR88, RH78]. Determinization [BGW00, Saf06]. Device [AFRV19]. Device-Independent [AFRV19]. DFS [BCCK19]. Diagnosability [CC07]. Diagnosable [DTA94]. Diagnosis [DTA94, SW99]. Diagram [CW98]. Diagrams [AdBMS98, AMT07, Aur87, BRLW19, CHLX17, GKM^{+}21, HPK15, IIM85, KS03a, LW80, LD81]. Diameter [ACIM99, BRS^{+}21, BEGL24, DHV22, GKM^{+}21, HLCW91, KN08]. Dichotomies [GKMP09]. Dichotomy [Bac21, BKN09, BP20, BG21, CCL13, CGW16, CCL16, CM13b, DR13, GGJT10, HNO^{+}20, ŁN06]. Dictionaries [AO95, Pag01]. Dictionary [MNR90, Yu22]. Difference [BGSW19, FKSV02]. Different [AHR96, BW05, HR93a, Har80, JKS84]. Differential [BUV18, Dav86, NTZ16, Ull16]. Differentially [BHMO22, BN14, FX15, KMMS22]. Difficult [Gol78, SC79b]. Difficulty [Fis05, GMP00]. Diffie [MW99a]. Diffusive [Cha15]. Digital [AKOT03, Doe04, FS86b, KPS94, KS00, LST99, Pro92, GMR88]. Digits [Che05]. Digraph [Fin20, KRY95]. Digraphs [AST76, BKN09, BDMT98, BF03a, CDTT95, HR73, HC99, Kle74, VTL82, HL96, CDTT95]. Dihedral [Kup05]. Dimension [AM96, AHLM07, CKM^{+}14, DSS18, DHV22, Ezr16, FFKP18, Hit04, Hit07, KMSZ05, KRS12, KW93, Lut03, Sny92, Bak78b]. Dimensional [AES99, AKS14, ABF94, AB98, ANSS21, BDD^{+}07, Cha05, Cha92b, CK95, CGH^{+}98, GP96, HPM06, HPK13, Har99, HJW84, Kar92, KU99, KMW75, KS03a, KOR00, LR86, LC90, MW99b, Pfa83, RS92, Sch04a, Vai88, BS83, CGJT80, DL90, EW86, ES91, Gol81, HJW85, IJ94, JDU^{+}74, LW80, RS94a, Yao82]. Dimensionality [BGK16, GOS^{+}11]. Dimensions [AS96, AAS97, AST97, BCR80, Cha00, DKK^{+}19b, Eri99, HP99a, Jus92, LS02, LM08b, MV00, MS97, ZS02]. Diophantine [CC82, Ili89b, Lag85]. Direct [DDG^{+}17, IJK09, IJKW10, IKW12, JT86, KŠdW07, She12, VC85]. Direct-Product [IKW12]. Directed [AAK90, AGU72, ADS86, Bak78a, BKN10, BBRS98, BE98, BKS92, Ben95, Ber16, BKM^{+}17, CHM13, EK05, FR06, FSS13, GM78, GIP20, HKK^{+}07, HKM08, HPT99, HP99b, Joh75, KMRS18, MPP22, NZ01, Ned99, Pac74, PPSS04, RZ08, Rod10, RZ16, Tar73, Tar74, Kao93, KS93, Sch94]. Direction [BV99, FRS03]. Directions [GLPW16, Tót03]. Directly [MO74]. Dis [AAZ16]. Disclosure [AARV21]. Discounted [BCK^{+}07]. Discounted-Reward [BCK^{+}07]. Discrepancy [AKOT03, CFK^{+}21, Ezr16, Lar14, LM15, Rot16, Rot17]. Discrete [AAKS14, BV06, BRLW19, Bra98, Che05, CCC^{+}13, DHR97, FOS08, FA77, GR93, GKM18, Hal08, Iba73, LY05, MW99a, MMP87, Pfa83, PR20, PM03, Sho97, LW88b, MR96, Mor80]. Discretization [RW00]. Discretized [BGGM98]. discs [Sha85]. Discussion [BBR88]. Disjoint [AKW00, BH19, BF03a, BS93, BFU94, BFSU98, CKS09, CKN22, Fri02, GH97, GSSZ04, HL72, IST12, IY23, KMV92, Luc90, NSV94, RZ10, Blu86, KS91, RLWW97, Rom82, Sch94]. Disjointness [BGK^{+}18, She16]. Disjunctions [EIM01]. Disjunctive [AHM^{+}08, CD16]. Disk [CKR72, CSW98, GZ05, GLP75, HLP17, HT84b, KL00]. Disks [CH82a, HSV05, KKMS11]. Dispenser [GKR05]. Dispersers [BSK12, Coh21]. Dispersion [CLL19]. Disproving [Juk06]. dissection [Kun93]. Dissemination [BSH94a, SCH81c]. Distance [AAKS14, AP72, AK10, AO12, AWZ00, BI18, Bod21, BGSW19, CR12, CW07, CHKZ03, DM88, DHP13, DHR97, DHV22, HLL95, KKMR09, Liu22, MMP08, PR14, Ros18, WWW87, PE86, ZS89]. Distance-hereditary [DM88]. Distances [ABF^{+}99, DTCR08, DN04, FKM^{+}09, Mar08, LW86]. Distinct [PT07, RRSS09, Nos82]. Distinctness [Amb07, BFadH^{+}87, BDH^{+}05, Yao94]. Distinguisher [Tho18]. Distinguishing [Gel75b, Hie10]. Distortion [ABN15, ACKS15, CKPT19, Fil19, KRS10b]. Distributed [AG94a, ANIM96, ALMZ99, AK09, BBO22, BD81, BEK14, BBR04, BCN^{+}20, BFKL21, BEG17, BFG^{+}07, CHD21, CLP20, CT90, CFPS12, CDDS89, De 10, DGP07, DGLV10, Elk06, FRS03, GKP98, GKK^{+}18, GKPS08, GM92, Har20, HKN21, Hie10, HY17, Jaf85, KP85, Lin92, LPSR09, LRWY99, MW93, NR95, PS97, PR00, SHK^{+}12, SY99, SY06, TPS87, CBH84, Fus88, IR96, RS85b]. Distribution [BS15, Blu94, BJ99, CFGM16, CK04b, FRS^{+}20, HK07, JS05, JL95b, LM94a, PU89b, RRSS09, RT96, HT82b]. Distribution-Free [Blu94, FRS^{+}20, HK07]. Distributional [Sch04a, Wan99]. Distributions [ACOH^{+}10, BGP^{+}23, CRS15, CK09, FOS08, GK18, LV22a, LV91, Nat92, She18b, Val11, Vio12b, Wat14, BCG20]. distributive [KMZ87]. Distributivity [BK10a]. Divide [CKL^{+}09, DY75, LR89, SR83, ACG89]. Divide-and-Conquer [CKL^{+}09, LR89, SR83, ACG89]. Divide/Sort/Merge [DY75]. Dividing [HT73]. divisibility [Pla78]. Division [BCH86, BP93a, MP05, PR20, RT90, Lic87, Lic89, dG83]. Divisions [Rev75]. Divisor [KMR87, Nor89]. Divisors [BMS86]. DNA [KW94]. DNF [ABK^{+}98, Baz09, BJ99, HY87, SBI04]. Do [aBC08]. Document [NN17]. Does [AH08, Bis78]. Doesn’t [SS12a]. Dog [DHP13]. Dogleg [Szk99]. Dogleg-Free [Szk99]. Domain [BP20, Blu94, CF22, SP82]. Domains [FOS08, GHSZ08, JKN08, Yao91]. Domatic [FHKS02]. Dominant [HTC13]. Dominating [BJ82, CCK^{+}20, CL19, COS99, FT06, RLDL96]. Domination [Cha98a, DM88]. Dominators [AHLT99, BGK^{+}08, Tar74]. Domino [Grä90a]. Double [BST95, CHH02, DHK^{+}21, JV75].

Double-Ended [BST95]. Doubled

[EMS03]. Doubling

[CLNS15, CHJ18, FRS19]. Doubly

[GJLR97, Lub87]. Down

[AIS08, HPK14, Jay98, MGHK09, SS05a]. Downward [BG98, HHH98, HHH05]. Drawing [FHH^{+}93, HL96]. Drawings [AABC^{+}17, CDTT95, DLV98, EGS03, SSSV01]. Driven [AFG^{+}14, DIM97, LPT98]. Drugs [DLL^{+}03]. Dual [ANFSW20, BEG^{+}02, GKPS08, JV08, KRTV18, KR05, KS20, RV98, HS88]. Dual-Bounded [BEG^{+}02]. Dual-Type [JV08]. Duality [AS05a, Asa87, BSRZ15, CDW21, CFKX07, GK18, GM18, HM73, HSV05, Pro78]. Duality-Based [CDW21]. Dualization [BEM10, EGM03, YS93]. Duals [He93]. Duplication [Bak97]. Duplications [AP07]. Dutch [Bit82]. Dyadic [Wir79, Zwi91]. Dyck [Lie03]. Dynamic [AAK^{+}12, AHS92, ABL09, AFHB^{+}00, AKK08, ABOK20, BGS15, BGS18, BCCK19, BHI18, BHNW23, BKK^{+}04, CHK13, Cha06, CPR11, CN18, CCFM04, CJ92, CPT96, CGJ83, CDTT95, CKM10, CMSS00, Col00, CH05, DK15, DHIP07, DKM^{+}94, DKKP99, FT97, FVY21, Fre97, GI92, GRS05, GRSS98, GT98, HSS01, Hen00, HK01, HKN16, HR03, IL98, JS20, KS96b, Kap00, KLM^{+}17, KMV91, LW98a, LWY20, LNRS92, LL95, Li89, Lou83, LP13, Meh79, MT99, Mor06, NN14, RZ08, RZ12, RZ16, Ros82b, SV00a, VZ13, CKS85, Man86, PT89]. Dynamical [MP05]. Dynamically [Bit79]. Dynamics [BBM13, CLV23, GJW24, Mol04].

Early [GGP11, HK00, CV83, PY88].

early-insertion [CV83, PY88].

Earthmover [KKMR09, NS07]. Ease [KMP^{+}20]. Easily [HS95]. Easy [BJK^{+}94, DFKL20, Gol78, KRT10, LL99, MW20, PVZ16, HH79, HH80]. Eccentricities [BRS^{+}21, DHV22, OL74]. economies [AAZ16]. Edge [AMS11, Asa87, BFU94, BFSU98, CM13a, CKS09, CH82b, CPP16, DV00, EGIS98, EGS10, FGG08, Fre97, Fri02, GG12, GI92, GI93, GMM09, HRW20, Hol81a, IY23, Kle08, KLS^{+}24, La 00, LY13, MM02, NGM97, NSV94, NV03, PS97, Pat13, RZ10, Sch98b, Yan81b, GK82, Hol81b, Lue81, Sch79]. Edge-Coloring [Hol81a, NV03]. Edge-Connectivities [LY13]. Edge-Connectivity [DV00]. Edge-Deletion [Asa87, Yan81b]. Edge-Disjoint [BFU94, BFSU98, CKS09, Fri02, IY23, NSV94]. edge-partition [Hol81b]. Edge-Weighted [MM02]. Edge-Weights [Kle08]. Edges [LM15, MM02]. Edit [AK10, AO12, BI18, BGSW19]. Editing [AALM90, ZS89]. effect [MT84]. Effective [AHLM07, BO10, CCL16, DR13, HT21, SS11c, TT01]. Effects [AZ04, DLL^{+}03]. Efficiency [BKPS02, CK96b, FMS05, GGKT05, HRV13]. Efficient [AAEZ21, AFN07, AK97b, AIK^{+}20, AALM90, AK88, AF01, BGNV18, BNL04, BLR14, BADTS22, BCLR92, BL98, BRS98, BV14, CFJ^{+}10, Cha98a, CK09, CLL19, CDK^{+}11, CS05, EL86, FP78, Fid85, Fin20, FH11, FJ89, GHSZ08, GM79, Goo99, Gop08, GS12, HOS92, HS02, HL04a, HH91, IA86, KRSV08, KS91, KMV92, Kle96, KNPS24, KOR00, LY13, LW88a, LRS17, LY05, MS10, Mac98, MZ08, Mai79, MR80, MSV06, MRK88, Mor80, NN93, PR93, PT07, PR06, PT92, RY96, Ram96, Reg96, RT88, RS99, SF90, SV00a, TV99, Tam98, TV85, Thu99, TN95, TA99, TC84, Tuc80, WJG00, XLXJ09, adH86, BHV99, CL79, FJ90a, Gus88b, IP87, MNS85, Mur87, SJ81, Sys81, WE80]. Efficiently [BWY15a, DHW98, FHS96, FGK05, KM00]. Eigenvalue [Kel06, Tre12, Hal75]. Eigenvalues [BP98]. Eigenvectors [BRV18]. Eighth [KRS20, AGK^{+}09]. Elastic [BGP^{+}22, LM94a]. Elastic-Degenerate [BGP^{+}22]. Election [AG91, BN00, GHW10, HP98, RSZ02, Sin97, AAG^{+}89]. Element [Amb07, BFadH^{+}87, BDH^{+}05, CRS95, Che08, JM78, Yao94]. Elementary [Gro03, Joh75, Kon75, Tar73]. Elements [Bea91, Hwa80, Maz76, PT07, RRSS09, Nos82]. Elimination [Fra91, HR73, Kle74, Oht76, RTL76, XLXJ09]. Ellipsoid [AF23]. Elliptic [Sho92]. Else [BT83b, GM87, MN87, Pig91]. Elusive [HW04, Tri94]. Embedded [CCE13, CAdVdM21, ZH05]. Embedding [ABN15, AFGN22, AFGN23, AKR18, Cha91, DMS24, EFN18, FFKP18, Lou83, Tam98, WC90, CS89, Tam87]. Embeddings [AHPY13, AL07, BCLR92, BGLL99, CDG^{+}09, CMM10, CT99, CHK03, ECM03, KR09, LNRS92, Mer02, MS08b, VW85a]. Emde [Wil00]. Emptiness [Eri00, SS11b]. Empty [BFvRV15, CDL86]. Enclosing [ZS02]. Encoding [AIKW15, Gel75a, Gel75b]. Encodings [BCG^{+}18, HKL00]. Encryption [AS16, ANSS21, BF03b, BCHK07, BV14, CS03b, GGH^{+}16, LATV17, Mah20, SW21]. End [KOR98, SU09]. End-to-End [KOR98]. Ended [BST95]. Energy [AIK^{+}20, KNPS24]. Enough [GJK^{+}22]. Enriched [PRS23]. Ensemble [FG92]. Entangled [KRT10, KKM^{+}11, QY21, Vid16, Vid20]. Entities [EKLS16]. Entropy [BDKR05, BADTS22, Coh21, KM99, LV22a, Rao09]. Enumerability [KS01]. Enumerable [Kum96]. Enumerating [BKPS06, KR95, MD76]. Enumeration [Akk73, CGM^{+}22, JV75, KW97, KV10, KV12, Sel78, Tar73, Val79, WW75, Gus88b]. Enumerative [The02]. Envelopes [AAS97, KS03b]. environment [Hen80, Man86]. Environments [AH00, CN99, MNS11]. Envy [CKMS21, CGV11, CFF^{+}12]. Envy-Free [CGV11, CFF^{+}12]. Envy-Freeness [CKMS21]. EOL [MSW78]. Epsilon [RS22, HY17, LR21]. epsilon-Approximations [HY17]. Epsilon-Net [RS22]. Equal [CJST07, Sed77]. Equal-Length [CJST07]. Equalities [CM22]. Equality [BC10, BB80, HPZZ21, ST94, BV84, Pet83, Pig91]. Equality-Test [Pig91]. Equation [Dav86, LT91, BEW80a, JS89a]. Equational [MN87, Sén05]. Equations [BPSV04, CKS17, EKR01, GJMM00, Hås88, KN08, KM13, MPR03, Mur90, NS22, PCHM85, Rub99, Sem09, SP82, XLXJ09, CC82, Ili89b, JK86]. Equilateral [BFvRV15]. Equilibria [Bar18, CRV10, CS19, DFM23, EN16, EY10, FRV10, GR21, LB17]. Equilibrium [DGP09, FGL16, FRGH^{+}23, GMSV15, HK11, Jai07, Rub18]. Equiprobable [GY96]. Equitable [JV08]. Equivalence [AHOW92, BMSU81, BV96, BG84, BL99, CHJT04, FMdR10, JT73, KMPHT14, Rub91, Sei90, Tze92, Web93, Coh83, CK87, FG82, Gur82, HCS80, Iba78, IL82a, SH85]. Equivalences [ASU79, SG04]. Equivalent [AvDK^{+}07, AKR18, Boo78, KRY95, LM12, Rei14, SPLK98, FS86a]. Equivocating [CPV22]. Erasure [DRTV18, GH17]. Erasure-Resilient [DRTV18]. EREW [KNP99]. Errata [Bas76, BD78a]. Erratum [BG96, BCD^{+}89a, FJ90b, HJLS14, Hol74, Kad91, Lan91, Lon88, Par98, RS94a, SS78, SY06, TV88b, Vai90, Vid20]. Error [ABO08, AP72, AS08a, AK23, BL21, BCG20, CGdW13, DH13, FK00, GKRdW09, HL14, KS13, Mei13, MR08, SL96, SR97, AB79]. Error-Correcting [AP72, CGdW13, Mei13]. Error-Resilient [SR97]. Errors [CJL03, GKW20, KW00, KL93]. Escott [GGG18]. establishing [Pet83]. Estimate [LL95]. Estimating [AK10, CS09, Fei06, GJ13, KKL^{+}93, SS17].

Estimation [ACIM99, BPP23, CDK^{+}11, DKLR00, FN07, GM09, HPK14, TZ12]. Estimators [DKK^{+}19b, LS75, Hum89]. ETH [dBBKBK23]. ETH-Tight [dBBKBK23]. ETOL [ERV81]. Euclid [Tre00]. Euclidean [AHPY13, ANFSW20, AKS14, BHZ23, CHP08, Che09, CAKM19, Col74, CEF^{+}05, DHR20, Dye86, HT82a, HS99, IW22, KR07, KS03a, NS00, Sak90, SCY00, SPLK98, SS95, SR83, WP03, dBBKBK23]. Euclidean-Norm [DHR20]. Eugene [Ano95]. Euler [CS18, CHH02]. Eulerian [GH73]. Evaluate [Bür00b, Lip78, PS73]. Evaluated [ACR^{+}10]. Evaluating [ASU75, BS86, DR95a, Rev75, SU77]. Evaluation [AL10, BS76, BMPT97, BGK^{+}08, BCGR92, CRSW13, Chi76, Gan95, GJ82, JáJ79, KM75, NP06, Pip80, RU81, Rei99, Yao76, DL80, Hya79, MRK88]. Evasive [Yao88]. Evasiveness [CKS02]. Even [FR80, Ned99, Sed78, ZH05]. Event [Che12, Kin86]. Events [IISI05]. Eventual [HMW01]. Every [AS08b, GHM^{+}11, MOP22, NS13, Pra75a]. Evolution [HHKL21, WW06]. Evolutionary [AK97b, BPWY99, CGG01, CK01, FT97, KW94, Kao98, KLST00]. Evolutionary-Tree [FT97]. Exact [ADG^{+}11, Amb16, BGGM98, CHPZ95, CH97, Dot10, FKTV08, GG91, GG92, GKS93, KS96b, PRS23, Sny92, dBBKBK23, Bak78b, Fri87b]. exact-match [Bak78b]. Exact-Size [PRS23]. Exactly [Che76, Ko89]. Examining [Wil00]. Example [BJ99, Epp03]. Examples [GM92, Sim97b]. Excellent [CSZ24b, CSZ24a]. Exchange [Çam03, EL86]. Exchangeability [Li89]. Excluded [GM15]. Excluding [GNW23]. Exclusion [BHK09, GM01, KMRZ98]. Exclusive [DKR96, FW90, HS02, KR88, Mit85]. Exclusive-Write [DKR96, KR88].

Execution [ACFW98, GM98b, LL95]. Exegesis [GMS81]. Exhaustive [Wil13]. Existence [ACG^{+}16, AMT07, BCPR16, BFU94, DMT88, GKL93, Kun74, dB84]. Existentially [FS20]. EXP [BvMR^{+}00, Fu95]. expanded [FS89]. Expander [BF03a, BFU94, CLL13, Fri02, Gil98, JKP20, KS97]. Expander-Based [KS97]. Expanders [AK23, BK07, FGRV14, KPS13]. Expanding [Ala14, Kla84]. Expansion [BFK^{+}14, Chi77, HPQ17, KS11, KPS13, KLL17, MS17, Pat13, Loo83]. Expansions [FS92, KMST87, MM83]. Expectations [Jac72]. Expected [AF03, AMMW07, CH82a, FP78, Gol08, GS87, KU99, Nic88, Sch90, Szp93, Wal79, AW96, Blo83, MT87, Mye85, Sch88, Sch78, Yao85b]. Expected-Case [AF03, AMMW07]. Expensive [BK11]. Experimental [LZ22a]. Explicit [KRS12, MOP22, MS08b, RS10a, RS22, Kla84]. Exploration [HIKK01]. Exploring [AH00]. Exponent [ABL09, Hås98]. Exponential [BYJK08, BNS18, BPP89, CCK^{+}20, CKP19, Che89, DHR20, FT06, GKR21, GKRdW09, GKK^{+}09, KLSS17, KMR22, MV13, NPSW23, PS12, Saf06, dBBKB^{+}20, BORW88, Hel84, RC79, Sta80]. exponential-time [BORW88]. Exponential-Time-Hypothesis [dBBKB^{+}20]. Exponentially [CKS17, ES15, Leu98, HR86]. Exponentials [GSY95]. Exponentiation [Sti90, Yac98]. Exposure [KZ07]. Exposure-Resilient [KZ07]. Expressibility [Imm89, Mol90]. Expressing [KKPV07]. Expressions [ASU79, BS76, GGM12, KM75, MMN14, MNS10, SU77, AB79, ASSU81, GJ82, Sag83, SH85]. EXPTIME [CGGR16, GGM11, Rob84]. EXPTIME-Complete [CGGR16, GGM11]. Extend [Che01]. Extended [MGHK09, SL94, FL83b]. Extendible [Bol76, Ros75a]. Extending [HHH05, Bak78b, CSZ24a]. Extension [AAE^{+}23, CKR04, Dur10, GJW18, KKMR09, PS97, Sch81a, Sri06, Lic87, Lic89, Pro78, SSC85]. Extension-Based [AAE^{+}23]. Extensional [MTV10]. Extensions [ADG08, CGL20, CFHM20, DKSS13b, GQ17, HLP17, MiHI99, Pan00, PR94, SW76, BEW80a]. External [AV03, GKR21, Pro92, VZ13]. Extinction [ESY17]. Extra [CLL08]. Extractable [BCPR16]. Extracting [BIW06, GK98b, RSW06]. Extraction [BL04, BEG17]. Extractor [DPVR12]. Extractors [BADTS22, BSRZ15, CGL20, CRS14, Coh16, DORS08, DLWZ14, DW11, GRS06, KZ07, Rao09, TS11, Vio14]. Extrapolation [LR88a]. Extremal [AB24, BDIG85, McI74]. extreme [MD73]. Extremely [AWY18, Sie04]. Eyes [GGL^{+}13].

f [Ngo05]. F. [AI77]. Face

[Bie90, CEG^{+}93, CHK03]. Faces [AMS98, GTV21, BM88]. Facility [ASS17, AGK^{+}04, BA10, CG05, CS03a, MYZ06, Tho05]. Factor [BGK20, BMKM07, BSW14, FGG08, GMM09, KPV14, KH83, Koc92, Mic04, MBCV98, NS00, Xia13, KMS83]. Factored [Bac88]. Factoring [BMS86, BCGW93, Lan85, Lan91, Len87, NRR02, PST88]. Factorization [AGKM16, BGS07, Duv91, GL80, KU11, Lem75, Pan94, SL80, Sho97, Kal85b]. Factors [KY09, GS21a, GS21b]. Fail [AAE^{+}23, CKS11, PP97]. Fail-Aware [CKS11]. Fail-Stop [PP97]. Failed [DTCR08]. Fails [BG98, BGG^{+}19]. Failure [AT98, GHKT12, KSS08]. Failures [DP20, GIP20, MRRT08, SWK09, BKO88]. Fair [ADK^{+}08, AS10b, BNMSS98, EWS05, HT17, IM20, PR20]. Fairness [BP10, KK06, Saf06, RY87]. Faithful [Esw75]. False [BI18]. familiar [Tom82]. Families [BS15, BDL14, CRSW13, MSS18, SZ76]. Family [Esw75, FS89, PA79, PA80]. Fan [CCH98, KMSV13, KS15]. Fan-In [KS15, CCH98, KMSV13]. Fanin [SS12a]. Farms [CKV10]. Fast [AC09, ACIM99, ADK04, BLS97, BBR04, BGRV98, BGP^{+}22, BG95a, BG96, Bra98, BGSW19, CHS12, Chi76, Cla89, Coh98, Csa76, CK01, DL91, DKSS13a, DK15, DMS24, Dow90, DKM06a, DKM06b, DKM06c, DHR97, DLR95, DGLV10, DS00, Ebe89, ES81, ENRS99, Fed87, FA77, FRV10, GS80, GM98b, GPST92, GKS90, GLN02, GMFB87, HM91, Han96, HPM06, HT84a, HKL00, HN94, HM99, KW97, KNP99, KL15, KvLP88, KU11, KMP77, Kru92, KT77, LP94, LL90, MS91, NNS89, NGM97, Oht76, Pan80, PR93, PR94, PM03, SJ81, SS77b, SS78, Sti90, TPS87, Val82, VSBR83, Vis91, XLXJ09, Yac98, BOFKT88, GGT89, Gus87b, HT82a, HH79, HH80, SW89, Sut91, ZS89, CDMI^{+}16]. Faster [ALW08, BGNV18, BK10b, CRSW13, CGP^{+}20, CDMI^{+}16, CH02, CH03b, DVW21, Für09, GT89, GK07, GNS20, HKM08, HRW20, Her14, Iwa03, KST99a, KPST94, NPSW23, Wil18]. Fat [AGMV00, AdBES14, MPS^{+}94, PT02, Tót08, dBG08]. father [LW86]. father-son [LW86]. Fault [AP06, ABO08, AL07, BCR18, BCH97, CLPR10, CK93, DPPU88, GGL98, KP05, KP00b, LMS98b, LM99, MMSA93, Sch87, YY85, vDMMS07]. Fault-Tolerant [AP06, ABO08, BCR18, BCH97, KP05, LM99, MMSA93, YY85, vDMMS07, GGL98]. Faults [CMS97, DHW98, HK00, LM99, Pio04, Tam98]. Feasibility [KC96]. Feasible [DKR96, Hoo90]. Feedback [ACBG^{+}17, BYGNR98, Ber76, CB74, ENZ00, Gel75a, Gel75b, GH17]. Ferromagnetic [GSVY16, GJ13]. Few [BIW06, IMM^{+}10, Vio07, CP84, VSBR83].

Fewer [GMP05]. Fibonacci [GR93]. Field [SS12a]. Fields [BW05, BCR16, BMS13, BK90, Che05, Cus98, Eps79, GKS10, GKS90, HSS13, Kam05, KR06, KY09, Lan85, Lov21, Rab80, RS08, Sem09, Sho92, Str76, vzGSS03, Lan91, Len87, Lic87, Lic89, SL80, vzGS98]. FIFO [BGL04]. Fifth [AIM16, CR17]. Fiftieth [BBS13, VNA20]. Fifty [KRS20, BSO14, BCAKT23, CCRZWN22, CR17, DMK19, HIMW16, LMOT14, MV18, Rou15]. Fifty-Eighth [KRS20]. Fifty-Fifth [CR17]. Fifty-First [LMOT14]. Fifty-Fourth [HIMW16]. Fifty-Ninth [BCAKT23]. Fifty-Second [BSO14, CCRZWN22]. Fifty-Seventh [DMK19]. Fifty-Sixth [MV18]. Fifty-Third [Rou15]. File [CH87, CGJL85, NR95, Whi90]. files [Wil85b]. Fill [BT01, FKTV08, FV13, NSS00]. Fill-In [FKTV08, FV13, NSS00, BT01]. Filled [Pan00]. Filter [DCdP98]. Filtering [Cha86]. final [BT83a]. Financial [CKLW01, Zuc19]. Find [BCN^{+}20, Hor87, MNA88, SW79]. Finding [AHU76, Alb85, AA89, AI77, Ata82, BY86, BX91, Bea91, BH03, Bod96, BDIG85, BGKP99, CHH02, CT76, CW98, CLY06, CL09, Epp98, FBGSV15, FRS^{+}20, Fri90a, FRT93, GM78, Gab07, Gal80, GSV99, Gaz91b, GM82, HJLS14, Had75, HHRS15, HOS00, HT84a, He93, HiO08, HR93c, IR78, IY23, Joh75, KYN07, KKP93, KMV92, KS94, KPST94, Kre90, MM02, MW95, MvV92, NS80, Ned99, NT05, Oht76, PR02, RLDL96, SV95, Sch98a, Sch94, SW99, Spi73, SP75, Tar74, TT77, Tho97, WW13, vzGSS03, BX92, GMG86, HJLS89, KIM81, KS91, MN88, MNS85, Pro88, SV88, SSY82, Sha79, CSZ24b]. Fine [CS99, OR21]. Fine-Grained [OR21]. Finger [CMSS00, Col00, DN04]. Fingerprinting [BUV18]. Finite [Ace94, ADG08, BBP00, BMPT97, BCR16, Boo78, BP02, BK90, BGW00, BJK05, BCL75, Che05, CC83, CN99, CHPW98, CK94, Cus98, Dur10, DS90, GHSZ08, GKS10, GKS90, Hen88, Iba73, Jac72, Kam05, KY09, KR78, Leu98, PVV98, Rab80, RT10, RS91b, RS08, Sei90, Sem09, Sén05, Sho92, Str76, Web93, Yao91, vzGSS03, AM80, BS86, Che86, ERV81, FG82, GR85, Ili89a, Lic89, PS81, RI89, SL80, SSC85, SH85, vzGS98]. Finite-State [DS90]. Finite-Valued [Web93]. Firing [CDDS89, DHM12, GK12b]. First [AAK90, ABI97, Coo74, FR75, FRGH^{+}23, Kao95, KMPHT14, KKPV07, PST00, PVV98, RT89b, RS07, Tar72, Car84, CKT93, Gra84, Hag90, HY88, Smi86, IKM^{+}12, LMOT14]. First-Order [ABI97, KMPHT14, KKPV07, PST00, PVV98, RS07, Car84, Gra84]. First-Price [FRGH^{+}23]. Fit [Mao93, RT89a, RT89b, RT96]. Fitting [ABF^{+}99, AC11, HPW04]. Five [KS19]. Fixed [AFB94, ASU75, AM96, Bie90, CKM^{+}14, CCK^{+}20, CHM13, CFHM20, CAdVdM21, CLP^{+}19, DF95, EKL10, EY10, FBGSV15, FG01, GLP75, GKS93, Hav24, KS03a, LSS^{+}21, LPPS17, MR14, RS94b, RS99, SOAD09, VHPT09, WBZ13, WWW87, BH87, NHL82]. Fixed-Depth [LSS^{+}21]. Fixed-Head [GLP75]. Fixed-Parameter [CHM13, CLP^{+}19, FBGSV15, FG01, Hav24, LPPS17, MR14, WBZ13, DF95]. Fixed-Points [SOAD09]. FixedPoint [MS76a]. Fixing [GRS06, KZ07]. Flag [Bit82]. Flat [DW85]. Flatto [ST00]. Flip [BKPS06]. Flipping [ACG^{+}16, AN93, BHMO22, CK17, DSS90, HO14, HT17, RSZ02]. Floor [He99, MST91, YS93]. Floor-Plan [He99]. Floor-Planning [YS93]. Flow [AC82, APS74, AOT89, AOST94, AALR02, AAE13, BPS10, BBR04, BGK20, BSW14, BKM^{+}17, CLL08, CKS13, CH95, CHM96, Coh95, DD81, EV21, ET75, EIS76, FMS05, GVY96, GK07, GKK^{+}18, GMFB87, Gus90, HU72, HU75, HRW20, IS79, Jac90, KKKP04, iKK18, KTW99, Ken76, KPST94, KS02, MN95, Ros80, Ros82a, Shi79b, Vég16, CM89, GGT89, Gus83, HJ85, KK86, LY89, AR98b]. Flow-Time [BGK20, KK86]. Flows [BK15, CEN12, CKRV15, CDR10, FS07a, MNS85, MNS86]. Flushing [Var00]. FOCS [BBS13, BSO14, CR17, DMK19, HIMW16, KRS20, LMOT14, MV18, Rou15]. Focused [AIK19]. Folded [KRZSW23]. Folding [BGSW19, FW07, RL88]. Follow [BPS07]. Fool [Baz09]. Fools [DGJ^{+}10, HLV18]. Forbidden [AFN07, AK23, KSS20, MS07a]. Forcing [HSV07]. Forest [CHJ18, FKLS10, GW95, HLP17, KLSvZ08, NSH^{+}08, PR02]. Forests [HK01, WBZ13]. Forgetron [DSSS08]. Form [AHM^{+}08, Ben77, VL00, BLR80]. Formal [AW76, BJM00, BV84, FL83b, Kal93, KK02, LV95]. Formal-Language-Constrained [BJM00]. Formation [FYO^{+}15, SY99, SY06]. Forms [ABP19, BG84, Bsh94b, Gie95, GL77, HM73, JáJ79, MV75, MSW78, MSW81, Sav74, Ili89a, JáJ80, KB79]. Formula [AT00, ACR^{+}10, BCGR92, CLS19, DDPY08, Fid85, GMWW17, Kla07b, KRT17, KR23, Kal85a]. Formulae [CL94, GM91c, Lic82, Ren92, Fri86a]. Formulas [ART09, ASSS16, AHM^{+}08, Baz09, BOC92, BHH95, BCE95, BC98, CD16, Che76, COF14, DZ97, FLMS15, GGGY21, GKS93, Hås98, HW02, HS90, KLSS17, KS15, LL00b, LTW18, Ros18, RS07, RV01, FMP82]. Forrelation [AA18]. Fortieth [CDG11]. Forty [AIM16, DKI18, DKSS22, EvMP^{+}12, IKM^{+}12, MS16, PKS^{+}13, SK18]. Forty-Fifth [AIM16]. Forty-First [IKM^{+}12]. Forty-Fourth [MS16]. Forty-Ninth [DKSS22]. Forty-Second [PKS^{+}13]. Forty-Seventh [DKI18]. Forty-Sixth [SK18]. Forty-Third [EvMP^{+}12]. Forward [SV00a, WZ15].

Forward-Backward [WZ15]. Forwarding [HK00, Whi90]. Foundations [AEM^{+}11, AK07, ANR09, BBS13, BSO14, BCAKT23, CR17, DT08, DMK19, HIMW16, KRS20, LMOT14, LU10, MV18, Rou15]. Four [AAS97, CSZ24b, CSZ24a, CLY06, KLSS17, KLR03, Gus87b]. Four-Coloring [CSZ24b, CSZ24a]. Fourier [BC91, CGV13, Cla89, GOS^{+}11, GKM18, KM93, MRRS07, MRS08, Mor80]. Fourth [Ber97, HIMW16, MS16]. FPT [BDT18, CCH^{+}16, IWY16]. Fractals [LM08b]. Fraction [Chi77, Jus92, KT77]. Fractional [BI06, GK07, KPR^{+}13]. Fractionality [iKK18]. Fractions [LC89, Str83, CC86]. Fragmentation [Rus77b, CKS85]. Frame [CS04]. Framework [FSZ22, FHHP19, GJ00, LZ22a, The02, dBBKB^{+}20]. Fréchet [AAKS14, DHP13]. Free [AGG^{+}19, AP72, AR02, BGS98, Blu94, BDK00, BT10, BDD^{+}07, CHJT04, CFK^{+}07, CGV11, CV14, CIL94, CSZ24b, CSZ24a, CPPT24, CFSZ02, CAKM19, CFF^{+}12, CG94, DHV22, EV21, ELRS03, FLM^{+}22, FRS^{+}20, GL77, Gli03, Gre73, Gre74, HK07, HOS00, Hav04, HR78, Kos75, LH00, Nat99, Pel94, SZ00, Sav73, SS77a, SW06a, Szk99, Tai80, TU81, TS81, WROM86, COS99, Gal84a, HC83, Hum89, Iba78, IL82a, ORW85, Ott86]. Freedom [Yan82]. Freeness [CKMS21]. Frege [BPU92, BPR00, BOBP^{+}04]. Frequencies [DLR95]. Frequency [ELRS03]. Frequent [GLPW16]. Frictional [RS03b]. Friendly [DH20]. front [OH87]. Frozen [ART09]. Frustration [BT10]. Frustration-Free [BT10]. Full [AD14, Bac21, FH06, GGL98, HSS09, JM96a]. Full-Text [HSS09]. Fully [BGS15, BGS18, BHI18, BV14, GI92, GM98a, HSS01, Hen00, IL98, Kar99a, LATV17, Mor06, PT89, RZ16]. Function [Aar03, ABC^{+}13, AM05, AALP06, BKM19, BGG18, Ber76, BLN01, CCPV11, CVZ14, GMWW17, GJ13, HILL99, HNO^{+}09, Lar77, MiHI99, NS93, Sit14, Fri86a]. Functional [AS16, BH81a, EKR01, GGH^{+}16, Rub99, CV85, Cha88, JS89a]. Functions [AAvK^{+}13, App13, AIKW15, AL18, ALRS98, Aza92, BH18, Bas72, BRV18, BBC^{+}22, BBD^{+}86, BCPR16, BT15, BWY15a, BHIK97, Bra08, CS16, CPS16, CK91, CFLS97, CK94, Dev02, DRST14, DKR96, Fei09, FMV11, FV16, FN99, GJ85, GGJT10, GKS93, GSY95, GM92, HRV13, HO14, HHL18, HHN04, Hoo90, HN02, IL81, IMP14, Iwa03, Jac72, JR13, KW93, KY10, KK77, KT94, KMN^{+}22, LL00a, Liu22, MI97, MV75, MP77, MZ13, NRR02, Nat91, Nat99, PR17, Pau77, PW11, RS10a, RS22, Rei86, RS91c, Rev75, Sie04, Sim97b, SOAD09, Val86, Yao91, ACGS88, BS86, BS92, CI83, Che86, Kal85a, Kut91, LR88b, RC79, SS90, SSY82, Zwi91, vzG86]. Fundamental [DFVW99, WZ06]. Further [GHJ^{+}24, MR91, PT09]. Fusion [Wil00]. Fuzzy [DORS08].

G [Kon75]. G/G/1 [Kon75]. Gadgets [CFK^{+}21, TSSW00]. Galerkin [HOS00]. Galil [AG86]. Galled [JNS06]. Galois [McK79]. Galton [Dev12]. Game [KLSvZ08, MNR90, AKPW95, VT89]. Games [ADG^{+}11, AN93, BVY22, CJK^{+}22, CGGS19, DFM23, EN16, EWS05, Grä90a, JPZ08, KKL09, KAI79, KRT10, KKM^{+}11, KLN19, LSZ11, Meh18, QY21, Rao11, SC79b, Vid16, Vid20]. Gap [CR12, CCK^{+}20, DS90, MW93, SB13, CG10]. Gap-Exponential [CCK^{+}20]. Gap-Hamming-Distance [CR12]. Gaps [GMPT10, JLTX20, JJLZ22, PS12]. Garbage [Lar77]. Garble [AIK14]. Garbled [AIKW15]. Gates [AM05, RS03a, Shp09, Vio07, vDMMS07]. Gather [BDP23]. Gathering [AP06, CFPS12, EHJM00, ISK^{+}12]. Gauss [Sak90]. Gaussian [CV18, MR07a]. GCD [SPLK98]. Gene [MLZ00]. General [AAK90, ACOH^{+}10, AFHB^{+}00, AST76, BCR18, Bea91, Bür00b, CRZ22, CLPR10, CM01, CPW82, CGG01, DV00, EP04, EC95, FLS99, FHHP19, GW95, HKL00, Her14, IMP14, JP05, JS81, KSV08, KKR04, KR06, KH83, KTZ15, KKR17, LNRW10, Liu22, Pan00, QS06, RY96, SW06b, SD88, Sny92, TZ17, Urq11, Ver97, BC82, GMG86, May84]. General-Valued [CRZ22, KTZ15, KKR17, TZ17]. Generalization [BLN01, BvMR^{+}00, Gia95, LD81]. Generalizations [DSS18, GS21a, GS21b]. Generalized [Abr87, AAEZ21, ACG^{+}05, Bak78a, BBN12, BBFT23, BC91, BEG17, BT80b, CIRR16, Cha76, CHJT04, Chi76, EL08, FJ84, FJ90b, HPK15, HS95, JV75, JKN08, KLN19, Lau83, Lie03, Lie76, LST99, MR96, NV03, Pip78, RS07, ST23, Sit14, SOAD09, Szp93, AKR95, KMS83]. Generalizing [Jus92]. Generate [Bac88, DORS08, BM84]. generated [Eng86, Wil85b]. Generating [BEG^{+}02, EGM03, EL86, Fre78, Gab77, KP98, LLK80, PR94, RH77, Rus78, RS99, Saw01, TIAS77, Wor87, ZR79, BV84, FL83a]. Generation [ATS07, BH80, BWY15b, DR95c, GW17, Lie03, MMM22, Mis77, MV00, RT88, SF90, Thu99, WROM86, HC83]. Generator [BBS86b, EKS00, HILL99, Fus88]. Generators [ABSRW04, App13, BRRY14, GKL93, GMRZ13, HRV13, HO02, HU22, MZ13, NS93]. Generic [GGKT05]. Genericity [Wan98]. Genetic [DLL^{+}03, GM00, Hol73, Hol74, Kim97]. Genome [AP07, BP96]. Genomes [EMS03]. Genomic [NSH^{+}08]. genuinely [Meg83b]. Genus [CS18, Kel06, Lu14, Pat13]. Geodesic [AdBS19, LV22b, MMP87]. Geometric [AS98, BRS97, BCC^{+}09, BDMS13, BGGM98, BLMW11, Cha03, Cha06, CHPJ22, Cha98b, CLM05, CO82, Duv91, EHPR17, EGS10, EJS05, ELRS03, FGG08, FW07, GLN02, JKLS05, KPX10, KMMS22, KS97, MS92, MBCV98, Mit99, MS01, MS08b, MRR15, NR13, RS94b, RKS12, SY99, SY06, Tre00, WY07, dBBKB^{+}20, IA86, MS84, Pap81]. Geometrical [Sel96]. Geometries [Fri93]. Geometry [BP14, CP09, CPR11, CDRR07, FPS16, IIM85, NTZ16, SJ05, SD88, The02, Vai89a, Wil00, SS86b]. GF [KLX10]. Ghosts [FKT17]. Gibbard [FKKN11]. Gibbs [Eft16]. Girth [CL13, CV10, Mol04]. Given [AHM^{+}08, BHIK97, DLR95, Fis05, Jac72, Kun74]. Glauber [CLV23, Mol04]. Global [ALN^{+}12, AHKL07, BGG22, CG12, CMM10, HU75, Ken76, Mer01, MR10]. Gluing [BCLL21]. GM [Lov21]. GM-MDS [Lov21]. Gödel [MWY78]. Goedel [HB75]. Goemans [Kar99b]. Goes [FZ18]. Goldreich [TW14]. Good [Bak86, BBM13, CRV10, CS19, DDY16, DFL^{+}22, GMP00, GHK91, Kar99b, Lin84, MSW78, Rey77, TT13, VR89]. Goods [AS10b]. Gossip [CP12, GK10]. Gossip-Based [CP12, GK10]. Gossiping [BF94, BGRV98, FY96, Kru92, KCV92, WS91]. goto [Rus77a]. Gowers [ST09a]. Gradient [AK09]. Graham [CCG^{+}97, BT94, GW93a]. Grained [OR21]. Grammar [BLR^{+}15, GL77, GPV94, MSW81]. Grammar-Based [GPV94]. Grammar-Compressed [BLR^{+}15]. Grammars [Ben77, Boo72b, Coo74, Gra74, May81, RU81, Sav73, Tai80, Alb85, CRGM81, Eng86, Hum89, SH85]. Graph [AGU72, AABC^{+}17, AGvM^{+}18, AS08a, AS08b, AFNS09, ACBG^{+}17, AI77, AdW22, ABD^{+}07, AC78, ABL00, AF23, BES80, BRS^{+}21, BY86, BFK^{+}14, BRV16, BE98, BFL06, BBR^{+}99, BEGL24, Bra08, BPP23, CM87a, CCL13, CL13, CLP20, CLL13, CKP14, CK12, CGP^{+}20, Chv77, CDTT95, Coo74, CS24, DVW21, DNS81, DV00, DT97, DLR95, DHHM06, Edm98, Eve75, ET75, ENRS99, EGS03, Fei06, FKM^{+}09, FFK^{+}98, Fis05, FN07, FM08, FHHP19, GI93, GZ05, Gav72b, Gaz91b, Gus87a, Gus91, Had75, HS85a, HU72, HZ77, Hen92, HT73, HW04, HR93c, IR78, Joh75, KLR19, KM00, Kho06, KH83, KT96, KSS08, KvM02, KK98, KLL17, Lin92, MM02, MD76, McD79, MW95, Mil76, Mol04, MRŚ10, Ned99, Oht76, OR21, Pac74, PR17, PB83a, RLDL96, RG77, STP94, SS11c, ST13, Spi73, Tar72]. Graph [Tar73, Tho97, Tor04, Tri94, VY23, Wei80, Yao88, YS93, AKPW95, BM88, CG73, CK80, FS86a, FJ81, Gus88a, HW86, Hor87, IA86, JS82, Lub81, MN88, SJ81, Sch94, Sys81, Tam87, TC84]. Graph-Centrality [BPP23]. Graph-Structured [ACBG^{+}17]. Graph-Theoretic [HZ77, AKPW95]. Graphical [FVY21]. Graphs [ABN15, AGG^{+}19, APS74, AAK90, Akk73, AP07, AK97a, AFN07, ACOH^{+}10, ABR90, AK11, AK23, BX91, BKM19, BGG18, BST23, BK10b, BCCK19, BCG^{+}21, BKS92, BFG^{+}19, BATS11, BK15, Ber16, BC22, BMPP19, BB88, Boo78, BJ82, BD97, BKM^{+}17, BCC^{+}09, BDK00, BFU94, BFSU98, BDGJ99, BLTV99, BP92, CM13a, CCE13, CRRZ24, CHS12, CEFN23, Cha10, CCWZ18, Cha98a, Cha73, CLPR10, CKS09, CHH02, CHK03, CXY04, CGYZ06, CV14, CLL13, CV10, CSZ24b, CSZ24a, CPPT24, Chu16, CFSZ02, CAdVdM21, Coh21, CH82b, COS99, CDH13, CSS09, DHH96, DS94, DFP13, DW85, DHR97, DP20, DHV22, DMW05, DCH21, DFJ98, DFJ02, EP04, EP09, EGS10, EV21, EJS05, EHSS99, FMS02, FLS04, FFKP18, FT06, FLM^{+}22, FHH^{+}93, Fri02, GM78, GS21a, Gal80, GW17, GSV99, GP01, GKM^{+}21]. Graphs [GH97, GIP20, GM91b, Gil98, GGM^{+}15, GKK13, GMP05, GR11a, GH73, GH74, GM15, GNS20, GNW23, GRS16, HCD89, Hal02, HPQ17, HKM08, Har20, He93, He99, HKL00, HR92, HPT99, HP99b, HSS01, HK01, HW95, HR82, HM99, ILPR81, IPS82, Jac90, JO92, JKLS05, JKP20, KS11, KPX10, KST99b, KMRS18, KR95, KL15, KKR04, KM94a, iKK18, Kel06, Kle96, Kle08, KKLV11, KLR03, KT15, KN05, KMMS06, KSS20, Kun74, La 00, LSW90, LPPS17, Lu14, MSS18, MPP22, MSZ11, MN95, MOP22, NNS89, NS00, NS13, NV03, NT05, OR21, Pat13, PSZ17, PR05, Pla90, RZ10, RZ08, Rod10, RZ12, RZ16, RTL76, Ros72, Ros14, Sch87, Sch98a, Seb97, Sén05, Set76, SW99, STU97, ST11, ST13, Spi85, Tar74, Tho05, Tom92, TV20, Tuc80, Vad01, VP03, Ver05]. Graphs [VW85a, WG05, Wei80, Wor87, Yan81a, ZH05, dBBKB^{+}20, Bab80, BX92, BH87, CS89, CNS82, CB81, DM88, Fed87, FS86a, GK82, GMG86, GB82, HY88, HK73, Hsu85, Hsu95, HCAL89, Kao93, KS93, KS96a, KS91, Kla84, KM89, LW88a, Lue81, Pro81, Pro86, Pul84, RLWW97, Sha79, SCH92, Shi80, Smi86, TY84, TY85]. Graustein [MY91]. Gravitational [CP05]. Gray [Fre78, FR80, JWW80]. Greatest [KMR87, Nor89]. Greedy [AS05b, CKMP95, CK01, FS20, FMR90, GLN02, HL05a, KMZ18, LB17, PSWvZ17, RT81, BKPS06]. Grid [GJ00, Ges02, IPS82, KM94a, Sch98a, Tam87]. Grids [Cha91, CC07, ECM03, HL04b, KV86]. Gröbner [Dub90]. Grothendieck [AN06]. Group [ABR90, DGV05, DP94b, DXSC94, Dun73, EGH07, FV98, FY96, GV19, GQ17, GS07, HRTS03, HKK^{+}07, IQ19, KK02, KLSvZ08, Lar14, MRS08, Vin04, WW75, MC87]. Group-Strategyproof [KLSvZ08]. Grouping [BT01]. Groups [BFP89, Bür00b, CKM^{+}14, Cla89, CK91, GRS05, GS12, McK79, MRRS07, SW06b, Wan99, BLS97, Ili89a, Ili89b]. Growth [CG12, SS89b]. Grzegorczyk [BN99a]. Guarantee [EW03]. Guaranteed [CDRR07]. Guaranteeing [BNMSS98]. Guarantees [ANFSW20, AFHB^{+}00, AABV98, BKN10, CDG^{+}09, CKMS21, Sri99]. Guarding [BMKM07, KK11]. Guillotine [Mit99].

Hadlock [AI77]. Half [IWY16]. Half-integrality [IWY16]. Halfspace [AS10a, Cha00]. Halfspaces [DGJ^{+}10, FGKP09, FGRW12, GR09, KKMS08, MORS10, MRR15, SSSS11, She13]. Halftoning [AKOT03, Doe04]. Hallucinate [KNPS24]. Hallucination [AIK^{+}20]. Hamilton [DFJ98, IPS82]. Hamiltonian [BDK00, CM16, DS94, GJT76, GH74, GB82, GS87, KKR06, MT84, SCH92]. Hamiltonians [BT10, HHKL21, KMP^{+}20]. Hamiltonicity [Bjö14]. Hamming [AL07, CR12, DMS24, PE86, Tre00]. Hand [PVZ16]. Handling [ER80]. Hankel [BG96, BG95a, LCC90]. Hanoi [CS04]. Haplotype [CCC^{+}09]. Haplotypes [XLXJ09]. Hard [BKM19, BSHR05, BT00, CM13a, CN06, CG01, DFHS08, EHS^{+}19, FGRW12, GGJ16, GJK^{+}22, GHM^{+}11, HK11, Hit07, JKP20, JR93, JL95b, KKM^{+}11, KK11, Lip78, LM94b, Lut95, MHSR98, NP22, SY91, Str74, Tod91, TO92, Ull16, Vid16, ACGS88, AGH17, FT83, HR86, HS87, Huỳ86b, KMS83, Meh18, Pla78, Vid20, FL94]. Hard-Core [EHS^{+}19]. Hardcore [KY10]. Harder [CL09]. Hardest [Gre73]. Hardness [AWY18, AK10, AZ07, AGS21, BCR16, BDV21, BGG22, BG22, CLP16, CN07, CKN22, CFLS97, DVW21, DGKR05, DMR09, DSY10, GJW24, GGG18, GL23, GKS10, GKSS22, GHS02, GR09, GL17, GHH^{+}17, Hås14b, HVV06, HY20, IJK09, KPV14, KS17a, KMP^{+}20, KKL04, KZ20, SHK^{+}12, SU09, SV10, Sve11, Tor04, GSVY16, LLK80, KM13]. Hardness-Randomness [DSY10]. Hardware [CIL94]. Harmonic [BHI^{+}13, CK01]. Hash [CRSW13, Dev99, GM79, Sie04, SS90]. Hashing [AHS92, BBD^{+}86, BK07, CKPT19, CO82, CadHS00, DMV04, DKM^{+}94, FPS13, FMM11, GadHW96, GM98b, KMW10, Knu19, MSV87, PP08, TZ12, Wil00, CV83, PY88]. Having [KW98, Lav73, Kni88, Ko89]. Head [GLP75]. Heap [BST95]. Heap-Ordered [BST95]. Heaps [CGS99, GM86, HST11, Jon87, KTZ13, KS20, SJ05, ST86]. Heat [SMW07]. Heat-Bath [SMW07]. Heaviest [CL09]. Heavily [BCSV06, CEG^{+}94]. heavy [GJ86]. Hecht [APS74]. Hecht-Ullman [APS74]. Height [Bro79, DSR92, DR95b, Dev99, KS00, Lar87, LP94, Meh82]. Height-Balanced [Bro79, Meh82]. Height-Limited [LP94]. Heilbronn [BKHL00, LS02]. Hell [BKN09]. Hellman [MW99a]. Helly [Hal08]. Helly-Type [Hal08]. Help [CJST07, MS92, NRS99]. Helps [AIK^{+}20, IL98, Kao95, Mak97, MMS17, Vai89a]. Hennessy [Dob05]. Henriksen [Kin86]. Hereditary [CSS09, FR24, DM88]. Hermite [Ili89a, KB79]. Heterogeneous [BNGNS00]. Heuristic [BHZ23, Che92, Che93, GW93a, Pap92, Pla90, RT81, WC92, WL95, AL88, CCG^{+}97]. Heuristics [AGK^{+}04, Bit79, LW98b, CF86, CK80, Fre84, RSL77, SR83, Tur86]. Hidden [ABK^{+}04, BGKS12, BS82, DISW13, FIM^{+}14, HRTS03, KKP93, Kup05, MRRS07, vDHI06]. Hidden-Line [BS82]. hides [LW88b]. Hiding [ALMZ99, HNO^{+}09, HHRS15]. Hierarchical [AC11, BNFN01, CD17, GPV94, LNRW10, RH93, ST09b, YR92]. Hierarchically [MHSR98, LW88a]. Hierarchies [Ald75, Gre74, HR97, LR21, LH00, Wat14, Ko89, RY87]. Hierarchy [ABP18, CW04, CK96a, CP19a, DvM06, GMPT10, GJK^{+}22, HNOS96, HHH98, Iba74, KvM02, Lav73, LSS^{+}21, Mor82, Ogi98, SL94, TT01, Tod91, TO92, BORW88, CGH^{+}88, CGH^{+}89, Kad88, Kad91, KS85]. High [AKS14, AF98, BL93, CV10, DKK^{+}19b, DKKP99, FHH^{+}93, HL14, HRZW20, KS13, KOR00, Mol04, She13, UY91]. High-Dimensional [AKS14]. High-Dimensions [DKK^{+}19b]. High-Performance [AF98]. High-Probability [UY91]. High-Rate [HRZW20]. High-Speed [DKKP99]. Higher [AdBMS98, KU99, MV00, PT09]. Higher-Dimensional [KU99]. Highest [Sch98a]. Highness [BBS86a]. Highway [FFKP18, GR16]. Hiring [BKK^{+}10]. Hirsch [Ver09]. History [KW94]. Hit [LV06]. Hit-and-Run [LV06]. Hitchcock [TN95]. Hits [ASSS16]. Hitting [AGKS15, ASSS16, ACRT99, BST23, HZ20, Tet99]. Hitting-Sets [AGKS15]. Hoeffding [PS97]. Holant [CLX11, GLV13, LW18]. Hold [CKLW01, Her14]. Holds [BKN09, FFK96]. Hole [HPK14]. Holes [MPS^{+}94]. Holographic [CLX17, CF22, Val08, vMS05]. Holyer [DT97]. HOM [CGGR16]. Homeomorph [KMV92]. Homeomorphism [STP94]. Homogeneous [BMPP19, KLSS17, KS17b, Kun81]. Homologous [DHK11]. Homology [BCL75, CEN12]. Homomorphic [BV14, LATV17, Mah20]. Homomorphism [OR21, SW06b]. Homomorphisms [CCL13]. Homotopic [KM94a]. Homotopy [CKM^{+}14, Hav04]. Hood [DMV04]. Hop [CHKZ03]. Hopbound [EN19]. Hopsets [EN19]. Horn [BBC^{+}22, EIM01, MiHI99, Sag88]. Horn-Clause [Sag88]. Horner [Dow90]. Hough [CSS90]. HP [FW07]. HP-Model [FW07]. Huffman [GMY12, Kou82, LP95, ML00, Par80, Par98, PR99]. Huge [FLS04, GGN10]. Hull [Eri99, KS86, ES91]. Hulls [CMS14, Har99, RS92, CSY87, RS94a]. Hybrid [AT98, FK88]. Hypercube [AL07, CS16, JMP11, Kru92, PU89a, WC90, CS89]. Hypercubes [BCLR92, Cha91, ECM03, FY96, LNRS92]. Hypercubic [LP98]. Hyperfinite [NS13].

Hypergraph

[DGKR05, EG95, EGM03, FMS10, GM18, GLLZ19, GHS02, GHH^{+}17, HT21, HNR08]. Hypergraphs [ADS86, BKL99, CX18, CR00, FPS16, GP04, Hal02, Har20, KS17a, TY84, TY85]. Hyperplane [DKM09, ESS93]. Hyperplanes [EOS86]. Hyperspace [BS82]. Hypertext [LL99]. Hypertree [GP04]. Hypertree-Width [GP04]. Hypertrees [Fri91]. Hypothesis [CCK^{+}20, dBBKB^{+}20].

I/O [MZ08]. I/O-Efficient [MZ08]. ICP [AV09]. Ideal [KV10, KV12, Dob89]. idealized [AHMP87]. ideals [Dub90, Huỳ86a]. Ideas [IL04]. Identical [Sve11]. Identification [GKRdW09, GKS93, MI97]. Identifying [CJM^{+}06, EG95, HL04b]. Identities [CS07, HT15, RS00, SS11a, GJ82]. Identity [BF03b, BCHK07, DS07, IQ19, KMSV13, SS12a, VV17]. Identity-Based [BF03b, BCHK07]. IEEE [BBS13, BSO14, BCAKT23, DMK19, HIMW16, LMOT14, MV18, Rou15, KRS20]. If [ABW18, BT83b, GM87, HS85b, Kad88, Kad91, MN87, Pig91]. II [AMS13, ACM84b, CGH^{+}89, CN99, CSZ24b, Col00, DKM06b, EGIS98, GS21b, Gal84b, Gel75b, HP99b, HS84, KS93, La 00, MS08b, RS91a, RT89d, RT93, SS86b]. III [DKM06c, RT89e]. Illustration [LPT98]. Im [KMN^{+}22]. Image [GJ85]. Immanants [Bür00a]. Immerman [Gef93]. Immune [AN93]. Immunity [SB84]. Impact [RW03, RRW01]. Impartial [FK15]. Imperative [NW06]. Implementation [Bro78, GGN10, PPS00, Rus77a, Mor80]. Implementations [CHJT04, CW94, JTT00, Tom82]. implementing [Yao81]. implication [CV85]. Implications [AKK^{+}09]. Implicit [FN93, HH91, Fre84]. Implies [BGK16, BK10a, HO14, Wil13, FT83]. Impossibility [BP15, BN14, BCF^{+}14, MRŚ10, SWK09]. Impossible [SZ00]. Imprecise [SLC91, SL96, vKLM10]. Improve [BSS21]. Improved [AOT89, AOST94, AAELvZ12, Alb98, ABHK11, AdBES14, BCS08, BHL18, BC22, BP93a, CG10, CG05, CKL^{+}09, CKS17, CS03a, CM94, Coh21, CG07b, Cun86, DY75, EGH07, EGS03, FHL08, GP90, GLS^{+}24, Hal02, Hen00, IL04, KRT17, KR02, KRZSW23, KR09, KLL17, PY91, PR08b, RZ08, RKS12, RC03, SSW94, Sri99, TV22, YYI12, CKT93]. Improvements [HRV13]. Improves [EW12, Fra91]. Improving [ACC^{+}11, BC00, CT90, CMS14, FGG08, Wil13]. In-Tree [PT87b]. Inaccurate [BJS97, CP08]. Inapproximability [AMS11, Aus10, BGGS20, BGG^{+}23, CCK^{+}20, CL19, DSS18, KKMO07, PSWvZ17, Rub18]. Incidence [Pri86]. Inclusion [BHK09, KM95, CV85]. Inclusion-Exclusion [BHK09]. Incomplete [PPSS04]. Incompressibility [BLTV99, BHV99]. Increase [EM03, Koc92]. Increasing [GG10, Koc92, NGM97, SS17]. Incremental [CCFM04, DV00, LMS98a, LNRW10, KM89]. Independence [AH11, Baz09, CV14, DGJ^{+}10, HLV18, KLR19, NT05, PPR09]. Independence-Free [CV14]. Independent [ABF94, AFRV19, BGG18, BIW06, BCG^{+}21, BGGS20, CM87b, CPPT24, CLY06, Dob84, DFJ02, FP78, FK03, Fei06, GRS06, GP96, Gav72b, GJW18, GS02, Jac72, KS77, MRT07, PY90, Rao09, TT77, TZ12, TA99, TIAS77, WL95, CNS82, GS89, LLK80, Lub86, Lue81, MN88, Rei84]. Index [BCHS94, CCKM13, Ata82, BCN81]. Indexed [Li20]. Indexing [AFG^{+}14, GV05]. Indices [HSS09]. Indistinguishability [AS16, AM13, BCG^{+}18, GGH^{+}16, SW21]. Indivisible [AS10b, CG18]. Induced [AFN07, CHLX17, DVW21, FR24, FTV15].

Induces [CM22]. Inductive

[JS08, BCD^{+}89a, BCD^{+}89b]. Inequalities [BEG^{+}02, CM94, DFKL20, Pan94, AS80, Gal84b]. Inequality [AN06, CM94, GR93, HN94, KLL17, LMR90, VV17, AS80, MR96]. Inference [ACKS15, BYGNR98, CCC^{+}09, JS08, Pla80]. Inferring [ASSU81, JLL12, KW94]. Infinite [AKK^{+}09, AJ22, BP20, Fri91, SL94, BS86, Ili89b]. Infinitely [BFS06]. Infinitely-Often [BFS06]. Infix [Rei72]. Influence [Cha15, MR10, ST09a]. Information [AHLM07, AHKL07, BSH94a, BN00, Bra15, CHS12, CCKM13, CCFM04, CK14, DPVR12, EHJM00, FRPU94, GKR21, GGL98, HKL00, JKS84, Kao97, KLL^{+}15, Kou77, KLR10, Lin84, Lu14, Mai79, MRRT08, RSZ02, SCH81c, VL00, WY07, CW83, Hul86, Jaf85, Man86]. Information-Theoretic [Lin84, WY07]. Information-Theoretically [HKL00, KLR10, Lu14]. Infrequent [Fra91]. Inherent [EHS12]. Inhomogeneity [AS86]. Initial [BT83a]. Inner [MV75]. Input [AKU05, IL82b, Iba78, MS87a, Sta80]. Inputs [CG97, DFKL20]. Insertion [BC84, FPS13, CV83, PY88]. Insertions [MT86, Meh82]. Insight [BT94]. Inspired [Bac21]. Instability [BGL04]. Installation [GMM09]. Instance [Dru15, Kum96, VV17, IIK85]. Instances [FGK05, GGM11, HN10, SF90]. Insufficiency [Rei72]. Integer [BEG^{+}02, BGS07, CCK^{+}22, DKSS13a, Für09, HJLS14, HS02, Har99, HSW92, HN94, JV75, Pol95, Pon98, RV98, RT90, SPLK98, Sri99, Sri06, FL83a, Fri87b, FHK^{+}88, HJLS89, Ili89a, KB79, Pla78]. Integers [CGGK95, KMR87, McI74, Pan00, PST88, WT79]. Integral [Che89, CDR10, Che86, Kal85b, Loo83]. Integrality [GMPT10, HKK^{+}07, PS12, TV20, IWY16]. Integration [Che86, SSC85]. Interactive [BBCR13, BOHP14, BNT^{+}19, Bra15, BE17, BCL75, CN99, GH17, Hai13, HHRS15, KLN15, LPS20, RRR21, She18b, GHY89, GMR89, MM83]. Interchange [AHK^{+}10a, ORW85]. interconnect [Vit88]. Interconnection [CR91]. Interdomain [LSZ11]. interesting [SPH84b]. Interior [GPST92]. Interior-Point [GPST92]. Interlacing [MSS18]. Interleaved [GGR11, GV19]. Interleaving [GMN09]. intermediate [AM80]. Internal [KPS94, Pro92, Rus77b]. Internet [EHJM00]. Interpolating [BC98]. Interpolation [BPR00, Chi76, DG93, Gop08, GKS90, GKS94, LS95, Man95, RB91, Shp09]. Interpretations [MSW81]. interprocessor [HCAL89]. Intersecting [BP00, Cha92b, Goo91]. Intersection [AS90, BYCDM92, Bas72, EJS05, FSZ22, HL04a, HPZZ21, NSV94, Sha85, She13, dBBKB^{+}20, BN78, Cun86, FS89, Mye85, SW79]. Intersections [APS93, BNP74]. Interval [AV03, Cha98a, DHH96, GP01, HSS01, HHKW07, HM99, KST99b, KKLV11, KMMS06, PY79, VHPT09, CB81, FLST86, JS89a, KS96a, KM89]. Interval-Ordered [PY79]. Intervals [BYHN^{+}06, CFK^{+}07, DCV90, Esw75, LS75, NHL82]. Intractability [DKK^{+}19b, FGLS10]. Intrinsic [LS22]. Introduction [Gol98, Vaz97]. Intruder [SY92]. Invariance [CK91]. Invariant [CCK^{+}22, HL14, Kao97, MM02, TZ22, WW75]. Invariants [BG84, Ver05]. Inverse [Chi77, KS98]. Inverses [BC91, LCC90]. Inversion [Csa76, vzGS00, Lic89]. Inverting [FN99, Kal93]. Investment [ADK04]. Involving [Rus77a]. IP [Mei13]. Irrational [CK00]. Irreducibility [CHJT04]. irreducible [Ata82]. Irrelevant [BP20]. Ising [GJ13, JS93]. Isolating [GTV21]. Isolation [CRS95, vMP19]. Isometry [CGV13]. Isomorph [Per74, Wil73]. Isomorphism [AT00, ABI97, AGvM^{+}18, ABC^{+}13, AF23, BES80, BWY15a, Boo78, CG07a, FFK96, Fis05, FM08, GQ17, GM15, GNS20, GNW23, IQ19, KT88, KR23, LS19, LRR17, LPPS17, MRŚ10, STP94, Tor04, CK80, FS86a, Hsu95, Lub81]. Isomorphisms [BH77, MWY78]. Isoperimetric [KMS18]. Isoperimetric-type [KMS18]. Isotropic [BCG^{+}21]. Issue [AGK^{+}09, Bab06, DT08, FGKO08, GS06]. Issues [BLMW11, GKR05]. Item [DFL^{+}22]. Items [CG18, RT89c, Rhe90, RT93, RT89d, RT89e]. Iterate [HHR11]. Iterated [CLS19, FLMS15, GG12]. Iteration [BEW80b, BÉ88, BEW80a]. Iterations [BI06, KY15, FSW87]. Iterative [BEW80b, Fri90a, FNR15, IM20, Sei77, Tra72]. IV [BST23, MSS18, Pul84].

J [BX92, BG96, CCG^{+}97, Col93, Coo81, FJ90b, HH80, HN79, Kad91, Lan91, Lon88, RS94a, SWCP96, Vai90]. Jackson [KM08]. Jacobian [ASSS16]. Jaywalking [DHP13]. Jensen [BKN09]. Job [Ado77, CM01]. Jobs [BH75, CJST07, DGBL00, DD81, FR94, HLL95, JT10, Sch04b, Sve12, Sim83, SW89]. Johnson [AC09, ST23]. join [FT83]. Joins [AGM13, Sag83]. Joint [LM94a]. Jump [Gre74, KST93]. Juntas [FV16, GLM^{+}16, KMR22].

Kakeya [DW11, DKSS13b]. Karp [AS86]. Katona [OW13]. Kautz [GH97]. Kernel [CFKX07, DSSS08, SSSS11]. Kernel-Based [DSSS08, SSSS11]. Kernelization [CFKX07]. Kernels [FLST20]. Kernighan [Pap92]. Key [BBC^{+}22, BH81b, CS03b, GS88, HNO^{+}20, NS12]. Key-Agreement [HNO^{+}20]. Keys [AFG^{+}14, AIKW15, ALS22, DORS08, Sed77]. Kinds [Sky76]. Kinetic [AdBS10]. Kings [SSW03]. Kinodynamic [RW00]. KKL [OW13]. Klee [Gaz91a]. Knapsack [CK05, Jan10, MS04, SVV12]. Knee [JC91]. Kneser [Hav24]. Knock [JC91]. Knock-Knee [JC91]. Knots [LDRT24]. Knowledge [ACK^{+}98, BDMP91, BCH^{+}20, BJSW20, BG22, CKPR02, DDPY08, FLS99, GK96, GOP98, HNO^{+}09, HvdMV04, IKOS09, KP04, PTW11, RW03, SZ00, Vad06, Wat09, GHY89, GMR89]. Known [AW21, CPP16]. Knuth [KN85, MMM22]. Kolmogorov [BL93, BLTV99, BFL02, Gli03, Kum96, LM08a, LV95]. Komlós [BDG19]. Kraft [BBP00, GR93, MR96]. Kruskal [OW13]. Kuratowski [KMV92].

L [Ano95]. Label [KLST00]. Labeled [DFK^{+}19]. Labeling [AAK^{+}06, AST99, BKS15, Che12, CN07, CKM10, KKMR09, KKKP04, Bab80]. labelings [Pro81]. Labels [CBGWW13, CHKZ03]. Lagarias [Fri86b]. Lagrangian [ML00]. Laguerre [IIM85]. Lake [BKK^{+}10]. Lambda [Rév́85, Wad76, Wad78, vT04]. Lambda-Calculus [Wad76]. Lambda-Conversion [Rév́85]. Landscape [Păt11]. Langevin [GJW24]. Language [BJM00, BGSW19, Cas99, CCKM13, Gre73, LV95, Lie03, Raj96, Rus77a, Rut90, SF90, FL83b, HC83]. Languages [AP72, BN99b, Boo72a, BNP74, BL93, Boo94, DW82, DDPY08, DS76, GMN09, Gli03, Gra74, Gre74, HH72, HM81, HR78, Iga77, Imm87, Kos75, KKPV07, Lev76, LMPP15, LM94b, New02, PSW90, RS07, SS77a, Sky76, SZ76, WS78, aBC08, BCN81, BN78, CK87, Eng86, Har80, ORW85]. Laplacian [AdW22, MRSV21]. Lapse [DHPW99]. Large [Akk73, CHS12, CM22, FTV15, LSH05, Mor80, PST88]. Larger [JKN08]. Largest [CDL86]. Latency [ALMZ99, ALW08, AK03, FSS13, MI97, Sit21]. lateness [LS84]. Latent [ACKS15]. Lattice [AF01, Dye91, GMP05, HHKL21, HT98, LRS01, Mer02, MV13, PR99, Reg04, TV99]. Lattices [GTV21, HSW92, LS19, Mic04, Sco76, SOAD09, HRB87]. Lawler [Ano95]. Laws [Dev98, Dev02, HN02]. Layered [FFK^{+}98, MR02]. Laying [HR92]. Layout [AKW00, DMW05, EGS03, BH87]. Layouts [EMSV12, HPT99, HP99b, LR86, SV00b]. Lazy [AHS92, Ann19, MSV87]. LCA [CH05]. LCCs [DGGW20]. LCF [Coh83]. LDFS [CDH13, DCH21]. LDFS-Based [CDH13, DCH21]. Leader [BN00, GHW10, RSZ02, Sin97, AAG^{+}89]. Leading [BBM13]. Leaf [Vai87, CW83, Vai90]. Leakage [FRR^{+}14, GR15, NS12]. Learn [KLN^{+}11]. Learnability [BH18, IMM^{+}10, KW93, SG04]. Learned [CGG01]. Learning [ABK^{+}98, ABK^{+}04, AGS21, BJS97, BS15, BV96, BGP^{+}23, Blu94, BHH95, BGGM98, BJ99, BM02, Cas99, CK09, FOS08, FGKP09, FGRW12, FX15, FS92, GRS93, GKW20, GM92, GR09, HSW92, Hit07, JS05, KKMS08, KMMS22, KL93, Ko91, KM93, LV91, Maa97, Nat91, Nat92, Nat99, NSW11, OS07, Ser02, SSSS11, Sim97b]. Least [EKL10, Eve75, HL87, TO92]. Left [BEM10, Nor89]. left-shift [Nor89]. Left-to-Right [BEM10]. Legged [BDL00]. Lemma [AIK19, CGH13, CR00, FPS16, GS00, GLLZ19, HV20, JM96a, Kol18, LLRS01, MW20, RRW01, SBI04, ORW85, Sri06]. Lemmas [EPR81]. Lempel [FNV13, KS00, KM99, LST99]. Length [AFG^{+}14, BH03, CGS22, CJST07, DP94a, ET93, Efr12, Gab07, GG10, GLP75, Jac72, KPS94, Meh77, ML00, Szk99, Thu99, FMP82, Gab88, LY89, SW89, Vit88]. Lengths [AST76]. Lens [BDV21]. Less [Xia13]. Letter [GY96, GMY12]. Level [Juk06, MO74, CGJT80]. level-oriented [CGJT80]. Levelability [Wan92]. levelable [ORS86]. Levels [AdBMS98, AES99, Cha00, Mat91, Ko89]. Levin [TW14]. Lexical [Lub87]. Lexicographic [Hal08]. Lexicographical [Lie03]. Lexicographically [RH77, Rus78, ZR79]. Lift [BSS21]. Lift-and-Round [BSS21]. Lifting [CFK^{+}21, GPW20]. Light [ES15]. Lighter [ES15]. Lightweight [KPX10]. Like [Pan00, PS12, Sav73]. likelihood [Hum89]. Limit [CPY89, Dev98, Dev02, HN02]. Limitations [CK17, Cyp93b, JS20, Kla84, Mis77, RS81]. Limited [ABK00, BNT98, BG98, BS03, BH75, CK14, HS03, LP94, Tót03]. Limiting [MP87, SW93]. Limits [AW21, AS16, Dru15, KS15, SMW07, VZ13]. Lin [Pap92]. Lindenstrauss [AC09]. Line [AS05a, AU72, Alb98, AGL99, BNFN01, BGLL99, BDD^{+}07, BS82, CEG^{+}93, DT96, FG98, FMR90, GW93a, GI93, Goo91, IL82b, KRS10a, KS95, LS98, LSH05, LL92, LRWY99, MM93, Mat91, PS91, RT93, SL96, SWW95, Tet99, The02, Tót03, ALM96, BFL06, CCG^{+}97, CL91, Fre85, LMSPR01, MR80, MRK88, Rap89, SC79a]. Linear [AT77, AAvK^{+}13, AM96, ABS07, AKW90, AN93, AHLT99, AMS11, AO12, ANSS21, ABCP98, AK09, Bak78a, BEK14, BBR04, Bie90, BP90, Bod96, Bod21, BNP74, BN78, BEG^{+}02, BKM^{+}17, BGS07, BGWZ20, BDK00, BFNS15, BST95, BGK^{+}08, Bür00b, BY98, CCK^{+}17, Cha05, CCC^{+}09, CCWZ18, CGR12, CL13, CHH02, CGV13, CGS22, CKS17, CW98, CDMI^{+}16, CAdVdM21, CM94, CB81, CH03c, CPS85, COS99, DHH96, DHK11, DRT92, DS02, Dye84, DS00, EM03, EGS10, EiC17, EL86, FZ18, Fid85, GM91a, GMSV15, Gol08, GMY12, GMS81, Gra94, GS02, HSSS22, HKRT95, HN94, IK22b, Jai07, JO92, JP05, Kao93, KS93, KW00, KM97b, KM94a, KN08, KM13, Kle08, Knu19, KR07, KTZ15, KS13, KMP14, KM77, KZ20, LS18, LMM99, Li20, Lu14, Luc90, LMR90, MM02, Meg83a, MR08]. Linear [Nat95, NW06, PPR09, PSW90, PR93, PRS23, Pol95, PR94, RS10a, RS22, RR04, Reg96, RT82, RZ16, Sav74, Sch13, Sei77, SSSV01, SPLK98, ST13, SY80, Tar72, TZ12, TZ22, TSSW00, XLXJ09, Ye90, YY14, AS80, Bin84, BW79, CC82, FHK^{+}88, Ili89b, KP82, KM89, Meg83b, Mil88, PS81, SW85, SWCP96, Sch78, Sha79, Shi79a, TY84, TY85]. Linear-Invariant [TZ22]. Linear-processor [Kao93, KS93]. Linear-Sized [AN93, LS18]. Linear-Time [Bie90, Bod96, BGK^{+}08, CCC^{+}09, DHH96, EGS10, GM91a, GMY12, KW00, KM94a, Kle08, KR07, Lu14, Sei77, Meg83a, KM89, SW85, SWCP96, TY84, TY85]. Linear/Polynomial [NW06]. Linearity [KLX10, GO80]. Linearization [GKV06]. Linearly [HL72, MPS^{+}94]. Lines [AMS98, BDD^{+}07, EOS86, KS03a, Pel94]. Linguistics [Maz76]. Link [BT05, CBGWW13, DTCR08, SWK09]. Linkage [KT96, LW98b]. Linkages [HJW84]. Links [CH03b]. Liouville [SSC85]. Lipschitz [JR13]. List [Alb98, BE17, CM22, CW07, CGV13, CV10, CS80, CV88, CPPW12, DHK^{+}21, GW93a, GGM^{+}15, GGR11, GLS^{+}24, HRZW20, IJK09, KY10, KS13, KRZSW23, OH87, ST23, BD86, CCG^{+}97]. List-Decodability [GLS^{+}24]. List-Decoding [DHK^{+}21, IJK09, KS13, ST23]. List-Recoverability [GLS^{+}24]. Listing [CN85, KK98, Rus81, Tro78, Sys81]. Lists [CGS99, HK87, KOT00, Wil76, BT80c]. Literal [BPP89, PB85]. Little [CKMS21]. Liu [BGP^{+}23]. LL [LC05]. LLL [NS09]. Load [AKU05, AKK08, AHKL07, BNFN01, BFG^{+}07, FGS12, GLM^{+}99, MRW00, NW98, RS94b, BJR89, BKK^{+}04]. Loaded [BCSV06]. Local [Aar06, AIK19, Ann19, App13, AL18, ALN^{+}12, AGK^{+}04, AST99, BFvRV15, BPP23, BG22, CGH13, CMM10, CV10, CAKM19, Coh16, CM16, DGL23, DF18, DGP07, FW14, FRS19, FRT93, GS17a, GLM^{+}99, GKKS93, GS17b, GLLZ19, GL21, Har20, HV20, HRZW20, HRW20, HY20, JL77, KWY98, KKR06, KBP03, KMP^{+}20, Kol18, KR05, KS13, KP04, Koz21, KLL17, LZ22a, LSV13, LLRS01, Lev76, LS10, Mes10, MR10, NS93, OPS04, PS77, Pol95, SS10, SY91, ST13, Sri06, Zha09, Jaf85, CKP19, CP19a]. Locality [AvMSS12, App13, ANSS21, CHPJ20, CKPT19, KP00b, Lin92, ST72, Vit88, IKP96]. Locality-Sensitive [CHPJ20]. Localization [GMR97, KMMT09]. Localized [AHKL07]. Localizing [DRW98]. Locally [AE15, AdBES14, BSGK^{+}10, CGS22, DS07, Efr12, KY09, KM94b, Kre90, Mei09, NS95]. Locating [BH13, KM96, Pre79]. Location [ASS17, AGK^{+}04, AMMW07, BA10, CP09, CN18, CG05, CJ92, CL21, CPT96, CS03a, EGS86, GT98, LP77, MYZ06, MTZC81, Pel96, Pre81, PT92, Sny92, TV91, Tho05, EKLS16, MS84, Pap81, PT89]. Lock [LPS20]. Locking [Pap83, Yan82, CRS87]. locks [Mit85]. Log [BCH86, BKPY18, JS05, KOR98, LMR90, Wil86, AM05, HM81]. Log-Depth [JS05]. Log-logarithmic [Wil86]. Log-Logstar [BKPY18]. Log-Space [KOR98]. log-tape [HM81]. Logarithm [Che05, LW88b]. Logarithmic [BADTS22, CV88, Dob21, EW03, EK05, GJLR97, LWY20, MRSV21, PD06, Rei86, Che86, Wil86]. Logarithmically [CKPR02]. Logarithms [GSY95, MW99a, Sho97]. Logic [GM18, KMPHT14, KT15, Lad77, LFS00, LMPP15, MS07a, PST00, SOAD09, TT01, Car84, Gal84b, PCHM85, RL88, SPH84b]. Logical [Zim98]. Logics [AM13, EJ99]. Logspace [HU22, KKLV11, Mac99, RY87]. Logstar [BKPY18]. Long [App13, BGH^{+}15, GHH^{+}17]. Longest [AdBS10, CGYZ06, FMS02, GG10, HSSS22, JL95a, KS00, MTZC81, SS17].

Longest-Side [AdBS10]. Look

[Sav73, Smi76]. Look-Ahead [Smi76]. Loop [Mis77, Yao78, IL82a]. loop-free [IL82a]. Loops [CCC^{+}09, MS90]. Loopy [EHS^{+}19]. Losers [BV06]. Loss [DKK^{+}19a, GK10, SSSS11, VL00]. Lossy [PW11]. Lottery [CDO^{+}22]. Lovász [BKM19, BGG18, BPS07, CGH13, FK03, GMPT10, HV20, Kol18, PS12, Sri06]. Low [ACKS15, AN19, Aga92, App13, BGK16, BEGL24, BN99a, BGL04, Cha05, CFK^{+}21, CDK^{+}20, DH13, DKM06b, FK00, FFKP18, FLM^{+}22, GJW24, GHH^{+}17, Hås88, HPM06, HPK13, HPQ17, KRS10b, KRY96, KP98, KM99, MR08, Pag01, SU09, SL94, Tom92, KS85]. Low-Degree [GJW24, GHH^{+}17, KRY96, KP98]. Low-Density [HPQ17]. Low-Dimensional [Cha05, HPM06, HPK13]. Low-Dimensionality [BGK16]. Low-Discrepancy [CFK^{+}21]. Low-Distortion [ACKS15]. Low-End [SU09]. Low-Error [DH13]. Low-Rank [DKM06b]. Low-Stretch [BEGL24]. Lower [Aar06, ABP18, ASSS16, AG94b, AM05, AvMSS12, AARV21, ACHAH16, ACH10, Aza92, BBO22, BBGR08, BHK^{+}19, BE98, BFL06, BC91, BPS07, BBI16, BB21, BAG01, BG92b, Bsh89, BKS15, BT20, BOBP^{+}04, CW04, CR10, CR12, CG10, CEEP16, Cha98b, CFKX07, CR22, CLS19, CHPZ95, Cyp93a, DP94a, DvM06, DKM^{+}94, EPA99, Elk06, EEST08, Eri99, EP98, FGLS14, FLMS15, Fre81, GR05, GG10, GG91, GMWW17, GadHW96, GJM98, GP18, GR21, GKS08, Gra90b, GT00, HHRS15, HS79, Hit07, HY20, HR03, JTT00, Joh88, Kam05, KRR94, KLSS17, KLL^{+}15, Kla07a, Kla07b, KRT17, KMR22, KR09, KM98, KMRZ98, Kut02, Lan14, LM08a, LWY20, LP13, Mac97, MST91, McD88, Mul99, MW20, NH90, PY91, PD06, PT09, Păt11, Pau77, PR00, PVZ16, PS12, Pon98]. Lower [Ram94, RRSS09, RS03a, RSY08, RV01, RSZ02, San09, SWK09, SBI04, Sel96, Shp03, SF11, Tom92, VZ13, Vio12a, Vio20, Wil89, Wil13, Yao91, dBBKB^{+}20, Cha84, CR80, CDR86, FMP82, Gra88, Kal85a, Lic89, MNA88, Sta80, Zwi91]. Lower-Bound [Mac97]. Lower-bounds [AARV21]. Lower-Stretch [EEST08]. Lowest [AHU76, ASSU81, SV88]. Lowness [BBS86a, BORW88, KHNS05]. LP [ASS17, Hal08, IWY16, KMR22, Kun81, LR21]. LP-Based [ASS17]. LP-branching [IWY16]. LP-scheduling [Kun81]. LP-Type [Hal08]. LPs [KRTV18]. LPT [Fri87a, GIS77]. LRF [KK86]. Lubricant [Ros82a]. Lucas [Kut02]. Lucid [AW76].

Machine [Ado77, Ann19, BGK20, Che93, CMS18, DH04, ELMS^{+}12, GS02, KTW99, Maz76, QS06, SU05, ST00, IKM85]. Machine-Independent [GS02]. Machines [AST76, BSS21, BNGNS01, BAG01, Bis78, CLL08, CK13, CFW93a, CG97, DvM06, DKR96, FK90, Gil77, IJTW95, Ko91, KLL90, LP03, LL90, LL92, PY91, PR06, San75, Sch80, SWW95, SZ76, Sve11, Sve12, Vit85a, BN89, CR80, CDR86, ĎG84, GR85, IP87, KvLP88, MS87a, NHL82, Pri79, RY87, SV84, Vit85b]. Macro [EM03]. Made [DFKL20, PVZ16]. Maintainable [HW93]. Maintaining [Ber16, DGIM02, GI93, GGK16, HF98, HK01, LY13, Yao85a, Man86]. Maintenance [Bak97, Cha03, DV00, Kap00, La 00]. Major [Sib98]. Majority [ARS97, GK98a, SV10, She09]. Make [GJ00, Sav73]. Makes [CM13a, KST93]. Makespan [CFF^{+}12, EÖW14, LR21]. Making [BGH^{+}15, HMR18, KLN15, RA00, IR96]. Malicious [KL93]. Malign [Mil93]. Malleable [ADL18, LPS20, MRT07]. Mal’tsev [BD06]. Management [ACER22, AV03, EW12, KLM^{+}04, LRWY99, BLS97].

Managing [AFG^{+}14, Ros75a]. Manhattan [Szk99]. Manin [BGS07]. Many [AMS98, Ala14, CKPR02, CGL20, CK14, ECM03, EGS10, MPS^{+}94, CP84]. Map [Sak90]. Mapping [BLNZ15, IL04, WG05]. Mappings [Yun79]. Maps [CPT96, HP99a, KRS10b]. Market [GMSV15, Jai07]. Markets [AGM^{+}11, CKLW01]. Marking [HL86]. Markov [BMS13, CGG01, CDG^{+}06, FPP11, HT82b, KPR00, LRS01]. Marriage [Kav14, NH90, Gus87b]. Marriages [IL86]. Martingales [BHMO22]. Mass [CPW82]. Massive [FKSV02, PT07, ST13]. Massively [CC23]. Match [CH03a, CH06, Riv76, Sch04a, Bak78b]. Matchgates [CLX17, CF22]. Matching [AWY18, Abr87, AGM^{+}11, ABK^{+}04, ABF94, AALP06, AK11, BDG19, BGS15, BGS18, BGP^{+}22, BHI18, BDL14, BG92b, BY98, BCC^{+}13, CRS95, CC99, CKL^{+}09, Che97, CT00, Col94, CHPZ95, CH97, CH02, CH03c, CGG^{+}97, CGH^{+}98, CLM^{+}20, DGY11, DKS88, EV21, FGT21, FG98, FMR90, GS21a, GS21b, GS80, GP90, GG91, GG92, GPST92, GV05, Har20, Har09, KU99, KLP96, KMP77, KRT17, LSV13, LPSR09, MS91, MT78, RT81, SRR95, Sub94, TU93, Vai89a, Vis91, Bak78b, BG90, GMG86, Liu81, SR83, SRP83, Yao79]. Matchings [AIKM07, GKK13, GHT76, HHR^{+}16, KLST00, KLST01, RRW01, SW06a, YYI12, HK73, MNS86]. Mathematical [BLMW11, May81, MR84, RD81]. Mathematics [ER80]. Mating [CCC^{+}09]. Matrices [BHPT24, BG95a, BG22, CCK^{+}22, CGV13, Cop82, DKM06a, DKM06b, DKM06c, DPPR20, FJ84, GL80, Gia95, HM91, HC99, JV75, LCC90, Lov21, Lub87, Pan00, Pra75b, BC87, BG96, FJ90b, SL80]. Matrix [ACIM99, AW21, AGKM16, AKOT03, BCG^{+}21, BGP^{+}22, BGG^{+}23, BP98, BG95b, BDMI14, BW17, BRS98, BD76, BD78a, Bsh89, CLS19, CW82, Csa76, CL09, DNS81, DKM06a, DKM06b, DKM06c, FLMS15, GLPW16, Gie95, GGM^{+}15, HKSS17, HS97, HM73, HS82, HS84, IKS10, Kal95, KLR21, LC89, Lan14, Lem75, LNW19, Pan80, PP95, Pro76, Ram96, Raz03, RS03a, Sch81b, SW19, She11, Shp03, Ata82, Ili89a, JM86, KB79, Rom82, Vel85]. Matrix-Chain-Product [Ram96]. Matroid [BFG23, CCPV11, DK14, FW14, GJ13, Har09, HL04a, Hli05, IK22b, JK82, LSV13, NSV94, Sot13, Cun86, FS89]. Matroids [DK14, FSZ22]. Matter [KR02, SS12a]. MAX [Kar99b, AS10b, BFK^{+}14, CHZ07, Coh95, EP09, GVY96, Hås14b, Pol95, Tre12, JKLS05, JKK06, JKN08, KKMO07, AR98b]. MAX-BISECTION [JKLS05]. Max-Cut [Pol95, KKMO07]. Max-Flow [Coh95, GVY96, AR98b]. Max-Min [AS10b]. Max-Not- [Hås14b]. Max-Stretch [EP09]. Maxima [ABSD^{+}98, CMS14, GKKS93, Kap00, LKS97]. Maximal [Akk73, BGS15, BGS18, Bsh90, CGM^{+}22, Gar74, TIAS77, CHT93, GMG86, GS89, LLK80, Lub86]. Maximally [AOS06, KLR19, QY21]. Maximin [LR89]. Maximization [BFNS15, BFG23, CG10, CHK13, CVZ14, ESV20, FW14, KMZ18, Von13]. Maximizing [CCPV11, Fei09, FMV11, GRS12]. Maximum [AOT89, AHS92, AA89, AMS11, AK97b, AI77, ACM^{+}99, BY86, BX91, BKM^{+}17, CC99, CH95, CHM96, CPPT24, mCL04, CFCH^{+}00, EV21, FK03, FW98, Gav72b, GKK^{+}18, GAOPL^{+}21, Had75, Har20, Hsu85, IS79, IY23, KLST01, KMV91, KLR03, LV22a, LS93, MI97, Mol04, PSWvZ17, TT77, WBZ13, YYI12, BX92, BH81b, CM89, CNS82, GGT89, HJ85, HK73, Hum89, KM96, LS84, MN88, MNS86]. Maximum-Density [mCL04]. Maximum-Flow [CH95, CHM96]. maximum-likelihood [Hum89]. May [Nor09]. maze [CR80]. MCMC [EHS^{+}19]. McMillan [BBP00]. MDS [Lov21]. Mean [AC82, DD81, KK86]. Means [ANFSW20, AV09, Che09, CAKM19, FSS20, FRS19]. Measure [AKOT03, Bis78, BvMR^{+}00, BL02, Gaz91a, Lut90, LM94b, BCN81]. Measurement [SZ08]. Measurements [GLPS12]. Measures [BL99, Cha03, GS88, KK06, MR07a, Mil93, Pat13]. Mechanical [RS03b]. Mechanism [BCGL17, BKV11, CDW21, DD13, DR14, GKLV14, KLSvZ08]. Mechanisms [Ala14, BSS14, Bar84, BN14, CDO^{+}22, DSS18, GRS12, JLTX20, JJLZ22, FL83b]. Media [BNL04]. Media-on-Demand [BNL04]. Median [ANFSW20, AGK^{+}04, Che09, CAKM19, DZ99, FMP^{+}03, KR07, LS16, MP03, Tho05]. Medians [BGKP99]. medium [Fri87b]. Meet [DPV15]. Meeting [RRW01]. Meets [Ann19, KR05, Tre00]. Mellin [Pro92]. Membership [AdFM18, BM99, BCMW17, GS07, GK10, HH91, JS81, Koz09, Loh06, LP13, New02, Ogi95, VZ13, Huỳ86a]. Memoriam [Ano95]. Memories [GK97, KS77, MO74, LS84]. Memory [AV03, ACFW98, CadHS00, DGLV10, GJMM00, GMR99, LL00b, MPS92, MW88, Nis94, PPS00, Reg96, SWPL11, VZ13, BN89, FSW87, Ord89, Yao81]. Memory-Efficient [Reg96]. Menger [RLWW97]. Merge [Col88, DY75, FR80, SY80, Col93]. Mergers [Coh16, DW11, DKSS13b]. Merging [BNL04, Fab77, FKS93, GGK92, HU73, HL72, Hwa80, Kon72, LM99, Lin84, Par77, Sed78, Tan78]. Merlin [CCM^{+}19, KLN19, San09]. Mersenne [KY09]. Mesh [AKW00, BWY15b, CLLS96, HOS00, MV00, CSS90, NS80]. mesh-connected [NS80]. Mesh-Free [HOS00]. Meshes [BCH97, PPS00, Sib98]. Meshing [CD03, CDRR07]. Message [AG91, BNGNS00, CT90, CK96b, DIM97, FJ89, GK10, HK00, LAB01, FN80, FJ90a, GMR88]. message-based [FN80]. Message-Driven [DIM97]. Message-Optimal [CT90]. Message-Passing [CK96b]. Messages [BGKL03, BGRV98]. Metabelian [Cla89]. Meters [KBP03]. Method [AM04, AV09, AF23, BCS10, Ber97, Bra75, BLTV99, BHV99, CN18, Che92, CK93, DH20, DKSS13b, Fri90b, HNR08, KSV08, Mai79, MM93, MMR95, She11, Ver09, JM96b, WE80]. Methodology [HKL00]. Methods [ALMZ99, FRV10, Fri90a, GPST92, Gus90, HOS00, Jac72, KM97b, KLN15, DL80]. Metric [AFGN22, AFGN23, ACK^{+}99, ALN^{+}12, BM04a, BT80a, BA10, CDG^{+}09, CFLY11, CMM10, Che09, CGV11, CN07, CS09, CS24, DMS24, EFN18, GZ05, HS97, IW22, KKMR09, MYZ06, MMP08]. Metrical [BCLL21, FM03]. Metrics [ABN15, ABF^{+}99, AC11, AK97b, BTW00, CG12, CHJ18, CAKM19, FPP11, FRS19, HPM06, LW80]. Migration [AALR02, KKW04, KDA00, Wes94]. Milner [Dob05]. Min [AS10b, BFK^{+}14, BBFT23, BADTS22, BGSW19, CHZ07, GVY96, Kho06, Rao09, CMST85, AR98b]. Min- [GVY96]. Min-Bisection [Kho06]. Min-Cut [AR98b]. Min-Entropy [BADTS22, Rao09]. Min-Max [BFK^{+}14, CHZ07]. Min-Plus [BGSW19]. Minima [BJK^{+}94]. Minimal [ADS86, Blä04, BEG^{+}02, BT01, CM87a, EG95, GJMM00, GK12b, HKRT95, JR93, KK98, KCV92, NP97, Ord89, OPR^{+}84, RS78, Thu99, Alb85, KMR88b, dG83]. Minimal-comparison [RS78]. Minimal-Cost [OPR^{+}84]. Minimal-Time [GK12b]. Minimally [JLL12]. Minimax [CKP86, KK85]. Minimean [Tan78]. Minimization [AZ07, CMS18, FGS85, HPK15, HW02, KC94, LM15, RS91c, Rot17, ST00, Spr02, LVW84, Tur86]. Minimize [AC82, DD81, GJ85, GKK^{+}12, Li20, MRSG05, SL96, LS84]. Minimizing [AHM^{+}08, AALR02, EKLS16, Iwa03, KTW99, Lar77, LY89, LP03, RU81, WW13, AK88, Sag83]. Minimum [AP72, AGvM^{+}18, AMS11, AAZ16, ALW08, AK03, AABV98, BX91, BX92, BLNZ15, BBC^{+}22, BT01, BTW00, BM99, BY98, CEFN23, CRT05, CX18, CLWZ04, CT76, CT00, CVV03, CV14, CT90, CDH13, CLP^{+}19, CEF^{+}05, CS09, DF18, DHR20, DRT92, DRW98, DCH21, EW03, Elk06, EP09, EÖW14, EiC17, FK02, FHL08, FH11, FKTV08, FV13, GHY89, GKP98, GW77, Gav72b, HKM08, HL04a, HL05a, HK01, HLCW91, HW95, IR78, KYN07, KM97a, KRY95, Ko91, KR02, Kou82, Lar89, LSW90, LPSPP05, MMS17, NSS00, OL74, PR00, PR02, Rei83, RLDL96, SV18, Sit21, Szk99, Tho97, Vai88, Vég16, WLB^{+}99, YY14, ZS02, BéT89, Fre85, KIM81, KMZ87, LY89, Sha79, Shi79a, Shi80, SR83, SRP83, Tam87, Yao82]. Minimum-Cost [AAZ16, BY98, CVV03, CV14, Vég16, BéT89]. Minimum-knowledge [GHY89]. Minimum-Size [CT00]. Minimum-Weight [AABV98, CT90, GKP98, LPSPP05, PR00, RLDL96, YY14]. Minmax [ET93]. Minor [AGG^{+}19, CK12, CAKM19, DHV22, EV21, FLM^{+}22, KSS20]. Minor-Closed [CK12]. Minor-Free [AGG^{+}19, CAKM19]. Minors [BST23, GNW23, KSS20]. Minsky [She18a]. Mirror [BCLL21]. Miss [FP78]. Missing [CH03b]. Mistake [Blu94]. Mistake-Bound [Blu94]. Mitotic [JS08]. Mixed [ALRS98, DGH98, GGJT10, Gus87a]. Mixing [AJ22, BGG^{+}19, CLV23, CDG^{+}06, GMP05, Mor08]. Mixture [KSV08]. Mixtures [CK09, FOS08]. Mobile [AP06, CFPS12, CP08, FYO^{+}15, SY92, SY99, SY06]. Modal [Lad77, MR96]. Model [Ald75, ÀBS04, BGL04, CLP16, CPV22, CKP19, CP19a, CGG01, DFSS08, DCdP98, EPA99, EHS^{+}19, FP78, FKM^{+}09, FG01, FRS^{+}20, FW07, FY96, GSVY16, GPT00, GMR98, GadHW96, GJ13, GGL98, GR11a, GL13, GP04, Hal08, JC91, JS93, KRTV18, Lar14, LAB01, MM18, MMN14, Mul99, Nis94, OC75, PD06, RSZ02, Sot13, SH96, Szk99, VZ13, VC85, AB79, BC82, CKS85, Dob89, FN80, HT82b, Liu81, MR84]. Model-Checking [FG01]. Modeling [Kim97, KR78]. Models [ACG^{+}05, BE98, BFKL21, Blu94, CKL^{+}13, FVY21, GS73, GPV94, KSV08, Mol03, Wad76, Wad78, FRW88, GJ86, Mit85, adH86]. Moderately [RZ10]. Modes [HR93a]. Modification [Bra75, Sch80]. Modifications [DXSC94]. Modular [ABL00, BGG22, DL91, DKSS13a, GS07, Hås88, KU11, vzGS00]. Modularity [FFTC20]. Module [Bak86]. Modules [YS93, Har85]. Modulo [KN08, RS85a, Win80, JK86]. Molecules [GJ99]. Moment [Ber97, BCMW17, TZ12]. Moments [AM06, MM83]. Monadic [BMM21, FV98, LSSV08]. Mondshein [Sch16]. Monoids [BMPT97, Ros80, Ott86]. Monomials [FGKP09, FGRW12, Pip80]. Monotone [AS08b, BW05, BMM21, BEG^{+}02, BEM10, CCPV11, CGS99, DDPY08, DR95a, EGS86, FV98, FW14, GH98, GM91c, GM92, OS07, OW13, RY96, Ros14, Yao88, FMV11, Fri86a, PT89, EGM03]. Monotonic [BHIK97, Epp90, MRT07]. Monotonicity [BB21, CS16, KMS18, SS10]. Monte [DKLR00, DKM06a, DKM06b, DKM06c, KKL^{+}93, SS77b, SS78]. Monte-Carlo [KKL^{+}93, SS77b, SS78]. Moore [AG86, Col94, GO80, Ryt80, Sch88, TU93]. Morgan [Hås98, KRT17]. Morph [AABC^{+}17]. Most [CKP86, HMR18, MV13, AM05, LP83b]. Mostly [IL98]. Motion [AS90, AS97, BDL00, DFK^{+}19, FG92, FHS96, HOS92, Pel96, Rei80, RW00, HW86, KRR94]. motion-planning [KRR94]. move [OH87]. move-to-front [OH87]. move-to-rear [OH87]. Movement [HJW84, MS87b, Sed78, HJW85]. Movements [CP08]. Moving [AHPY13, EKLS16]. MPC [CDP21]. MPLS [GKR05]. MR [BX92, BG96, CCG^{+}97, Col93, Coo81, FJ90b, HH80, HN79, Kad91, Lan91, Lon88, RS94a, SWCP96, Vai90]. MSO [EM03]. MST [MSVW16, MBCV98, Mit99]. MSTs [KR05]. Multi [ACBG^{+}17, BSS14, CCRM86, DFL^{+}22, GVY96, Got81, Shi79b, HITK79]. Multi-Armed [ACBG^{+}17, BSS14]. Multi-Item [DFL^{+}22]. Multi-Pass [CCRM86]. Multi-Terminal [Shi79b]. multi-valued [HITK79]. Multi-Way [Got81]. Multiarmed [ACBFS02]. Multicast [KK02]. Multicasting [BNGNS00]. Multicommodity [CKS13, CKRV15, EIS76, GK07, iKK18, MNS85, MNS86]. multicomputers [Vit88]. Multicut [BDT18, MR14]. Multicuts [CAdVdM21]. multidestination [Jaf85]. Multidimensional [AFB96, BCS10, CK04a, DL76, Dye86, DS00, EvS05, GP83, Sei77, Cha88]. Multiembedding [BM04a]. Multifit [FL83c, Fri84]. Multiflows [Seb97]. multigraphs [GK82]. Multihop [BYII93, CGK07]. Multikey [LATV17]. Multilayered [DGKR05]. Multilevel [Lav73]. Multilinear [AGKS15, CVZ14, CLS19, KMSV13, RSY08]. Multiobjective [GKLV14]. Multiparty [BPS07, BH12, CPV22, DF92, IKOS09, IKK^{+}11, PVZ16, She16, BHMO22]. Multiple [BJS97, BNGNS01, BYII93, BKM^{+}17, CCE13, CK05, DK15, FLS99, HLR96, HPZZ21, IY23, Jan10, KQT99, MN95, Muc76, YH90, HW86, KMST87, Wil86]. Multiple-Access [DK15]. Multiple-Object [KQT99]. Multiple-Sink [BKM^{+}17]. Multiple-Source [BKM^{+}17, CCE13]. Multiplex [ACKS15]. Multiplication [ACIM99, AW21, BGP^{+}22, BEM10, BD76, BD78a, Bsh89, BK90, CLS19, CW82, Cop82, CL09, DKSS13a, DKM06a, Fat74, FLMS15, Für09, HM73, Kal95, Kam05, Kli79, Lan14, Pon98, Pro76, Sch81b, Sho92, Shp09, Win80, JM86, Rom82]. Multiplications [BD76, BD78a, Gro03, PS73, Rev75]. Multiplications/Divisions [Rev75]. Multiplicative [NR13, Har85]. Multiplicities [DKSS13b]. Multiplicity [BV96, KRZSW23]. Multiplying [Pra75b]. Multiprocess [LL00a]. Multiprocessor [CC07, CM01, CGJ78, GG75, GJ75, JP05, KS77, KDA00, LL95, LL00a, Sim83, Wes94, Kun81, LS84, SW89]. Multiprocessors [AF98, CLT05, DGBL00]. Multiprogrammed [GS73, Lav73]. Multiprogramming [Bas73, Bas76, Gav72a]. Multiprover [BOHP14]. Multirate [CR91, DGHK98, HHW05, LDHX99, MT89, NV03]. multiset [Nos82]. Multisets [MS76b]. Multistage [SS12b]. Multiterminal [DJP^{+}94, GL87, Gus83, Shi80]. Multithreaded [BL98]. Multitype [ESY17]. Multiunit [JJLZ22]. Multivalued [CW94, DDR07, SOAD09, FT83, IK22a]. Multivariate [BM02, GKS10, GKS90, HSS13, MPR03, RB91, Hya79, Kal85b, Len87, Pla80]. Multiway [BNS18, CHM13, LM94a, NZ01]. Multiway-Cut [BNS18]. Multiwriter [SWPL11]. must [HS85b]. Mutual [GM01, KMRZ98]. Myopic [IL98].

n [AM05]. NAE [GS17a]. NAE- [GS17a]. name [Hen80]. Namespace [ACHP19]. Narrow [Nor09]. Nash [Bar18, CG18, DGP09, DFM23, EY10, GR21, HK11, Rub18]. National [Bit82]. Natural [AS08a, BFG03, BG81b, Eps79, Sel88, Wil16, Gra88]. Nature [CV10, Sky76]. Navigating [BRS97]. Navigation [AWZ00, BC00, LL99]. NC [EV21, FGT21, FR94, Kao93, KS93, SPLK98]. Near [ABL09, AO12, ABCP98, BFKL21, BADTS22, BHNW23, BGP^{+}23, BKM^{+}17, BDMI14, BGK^{+}18, BCG20, CCK^{+}17, CAdVdM21, DKK^{+}19a, FBL03, GKK^{+}18, GHJY91, GR21, HKSS17, HSY22, JMP11, KMW75, KK00, KMMT09, MOP22, PT09, PR00, Rei99, Urq11, Yao94, BW79, RL88]. Near-Linear [AO12, ABCP98, BKM^{+}17, CCK^{+}17, CAdVdM21]. Near-Logarithmic [BADTS22]. Near-Neighbor [PT09]. Near-Optimal [ABL09, BFKL21, BHNW23, BGP^{+}23, BDMI14, BGK^{+}18, BCG20, GKK^{+}18, GR21, HKSS17, HSY22, KMW75, KK00, Urq11, Yao94, BW79, RL88]. Near-Optimally [CCK^{+}17]. Near-Perfect [FBL03]. Near-Ramanujan [MOP22]. Near-Testable [GHJY91]. Near-Tight [KMMT09, PR00]. Near-Unanimity [DKK^{+}19a]. Nearest [AC09, AF03, CR10, CS24, HPK13, HT84a, KOR00, Liu22]. Nearly [ABC^{+}13, BHK^{+}19, BT20, CGS22, Fin20, Gie95, GKP04, GL17, KR07, Liu22, MRSV21, RS91b, SC79a, ST13, Yu22, HY88, RS91a]. Nearly-Linear [CGS22]. Necessary [PS73]. Neciporuk [Kla07b]. Necklaces [RS99]. Need [BDP23, Oya93]. Needed [Sim97b]. Negation [AM05, BNT98, Val86]. Negation-Limited [BNT98]. Negations [SW93]. Negative [Pra75b, MNS86]. Negligible [AK23]. Neighbor [AF03, CR10, CS24, HPK13, KOR00, PT09]. Neighborhood [ABCP98]. Neighbors [AKS14, AC09, Liu22]. Nerves [TT13]. Nested [Lan92]. Net [BDK00, RS10a, RS22, May84]. Nets [AES10, HPM06, Lie76, Maa97, OW13, GL87]. Network [ACKS15, AOST94, Alb09, AKU05, AZ04, AAZ16, ADK^{+}08, BKN10, BD93, CEKP18, CFLY11, CHKS10, CRV10, CLL13, EGOS07, EL86, ET75, FGG08, FR06, FNR15, GT89, GK05, GJ19, GKR12, GMFB87, Gus90, JNS06, JR94, Kar99a, KSS08, Koc92, KKL04, KNPS24, LNSS09, LS13, LZ22b, LS75, LSS13, MMP08, MW88, PRS94, SS89a, ALM96, CM89, CR86, Gus83, KMZ87, Rei83]. Networks [AG91, AG94a, AL10, BNGNS00, BYII93, BNT98, BF94, BTW00, BHV99, Çam03, CPR11, CKRV15, CS19, CGK07, CR91, Coh95, CG07b, Cyp93a, CD18, De 10, DKKP99, DY75, DGHK98, DPPU88, ELRS03, FRS^{+}20, FJ89, Gam03, GJM98, GK12b, HB94, HHW05, IS79, KBP03, KM08, KP04, KM98, LMS98b, LP98, LM99, LDHX99, MS99, MS08a, MT89, MR10, NR95, NV03, Ngo05, Nut10, Pio04, Pip91, RS94b, SW99, Sin97, SV00b, Tam98, TU81, TS81, Yao80, YY85, AKR95, EMC85, FSW87, FJ90a, GR85, HJ85, JM96b, MNS85, MR84, Ron82, SW85, SWCP96, Sch79]. Neural [Maa97, MW88]. Neurones [Ald75]. Newton [Fri90b]. Next [Mao93]. Next- [Mao93]. NFA [JR93]. NGSM [Iba78]. Nice [Che12, KY09]. Nilpotency [HJ90, Kar92]. Ninth [AKS09, BCAKT23, DKSS22]. NNJAG [EPA99]. No [BKN09, CHHKM17, CG01, CPPW12, BX92, BJR89, BG96, CCG^{+}97, Col93, Coo81, FJ90b, HH80, HN79, Kad91, Lan91, Lon88, RS94a, SWCP96, Tom82, Vai90, BDP23]. Node [AKW00, BHL18, Ben95, CV14, DTCR08, FNR15, HLP17, KD79, KNPS24, MR07b, Nut10, SSS91, Yan81a, KN05]. Node-Capacitated [KNPS24]. Node-Connectivity [FNR15]. Node-Deletion [KD79, Yan81a]. Node-Disjoint [AKW00]. Node-weighted [BHL18, HLP17]. Node-Weights [Nut10]. Nodes [Pro92]. Noise [BL21, BE17, DRST14, GR09, HLV18, KPS13, Nat99]. Noise-Free [Nat99]. Noisy [BNT^{+}19, DORS08, EP98, FRR^{+}14, FRPU94, GKS08, Nat99, NSW11, PRS94, QY21]. Noisy-Channel [PRS94]. Non [ADL18, BP15, CPS16, FMV11, FW14, Hir23, LPS20, Hum89]. Non-Black-Box [BP15, CPS16, Hir23]. non-context-free [Hum89]. Non-Interactive [LPS20]. Non-Malleable [ADL18, LPS20]. Non-monotone [FMV11]. Non-Oblivious [FW14]. Nonabelian [MR12]. Nonadaptive [DK15, Gam03]. Nonapproximability [BGS98, KTW99]. Nonblocking [CR91, CG07b, GHKT12, HHW05, JTT00, MT89, Ngo05, RW05, ALM96]. Noncanonical [SW76]. Noncommutative [JM86, LTW18, MV75]. Nonconvex [ESV20, Maa86, VA00]. noncounting [CRGM81]. Nondeterminism [Bal85, BG98, BG93b, CC97, CIRR16, DF92, HVV06, RA00, SB84, KF80]. Nondeterministic [CHPW98, Gef91, HS03, Imm88, Jay98, LH00, MW20, Sei77, Ven92, WM97, dW03, ĎG84, Hen80, RS81]. nonexclusive [Mit85]. Nonexpanding [CSS09]. Nonindependent [Doe04]. Noninduced [DVW21]. Noninteractive [BDMP91, FLS99, GS12]. Nonisomorphism [KvM02]. Nonlinear [HS87]. Nonlocal [BVY22, QY21, SZ08]. Nonmalleability [GRRV21]. Nonmalleable [BADTS22, CGL20, CRS14, DLWZ14, DDN00, PR08a, PR08b]. Nonmigratory [CLT05]. Nonnegative [AGKM16, CCL16, GLM^{+}16, JV75, LW18, Moi16, Ata82]. Nonoblivious [FMS05]. nonoverlapping [KM96]. Nonpreemptive [Kun81]. Nonreducible [APS74]. Nonregular [HR93b]. Nonscalar [PS73]. Nonsequential [ACFW98]. nonserial [Ros82b]. Nonsimple [dVE10]. Nonstandard [LS95]. Nonstochastic [ACBG^{+}17, ACBFS02, DFKL20]. nonsymmetric [Kar79]. Nontrivial [CR22, Gra90b, Mal05, Gra88]. Nontruthful [AKSW22, DFL^{+}22]. Nonuniform [CHK13, CHKS10, EHPR17, KR05, LMP02, Mol90, RW00]. nonuniformly [Wil85b]. nonuniversal [Pri79]. Nonzero [BW99]. Norm [AN06, DHR20]. Normal [AHM^{+}08, Ben77, BG84, VL00, Ili89a, KB79]. Normed [DH92]. Norms [AKR18, BGG^{+}23, LNW19, SZ08, BS92]. Note [APS74, BéT89, Bas72, CL77, DSR92, FR80, For79, Fre93, Fri93, Hav04, JS81, Jon87, Kle74, Kon72, Kur87, LS90, Rhe90, SWCP96, CCG^{+}97, Pre79]. Nothing [CKS13, iKK18]. notion [MRS88]. Notions [PS02]. Novel [SJ05]. Nowhere [Jac90]. NP [Vid20, MS01, AGH17, BT06, BCMW17, Cha92a, CG01, DT97, FT83, FL94, GGG18, GJT76, GSSZ04, GPSS06, GPSZ08, Gra88, Gra90b, Gra94, GL17, Hås14b, Hol81a, Hol81b, KK11, KS85, KW98, KST93, KD79, LLK80, Lub81, Mor82, OW91, PS02, Rap89, SS81, SS05b, Vid16, WC90, Zim98, Zuc96]. NP-Complete [DT97, Gra94, KST93, GJT76, GPSS06, GPSZ08, KD79, SS05b, WC90, Zuc96, Gra88, Lub81, Rap89, SS81]. NP-Completeness [Hol81a, PS02, Hol81b]. NP-hard [Vid20, FL94, KK11, Vid16, AGH17, FT83]. NP-Hardness [GGG18, GL17, Hås14b, LLK80]. NP-Pairs [GSSZ04]. NPDTOL [CK87]. Nucleation [SW10]. Number [Aga92, AFB94, BFG^{+}19, BOC92, BBS86b, BC76, Bra08, BS03, BD76, BD78a, CM13a, CFHM20, Chv77, CAdVdM21, CDG^{+}06, DFP13, ERS20, FHKS02, FR06, Goo91, GPW18, Gro03, HSV07, HS03, KLR19, KY15, KS03a, Lan85, Lu14, Mal05, McD79, McI74, Mou90, NPSW23, NT05, NS93, PW04, PS73, PVZ16, RU81, Rao09, RT88, Rev75, SW06a, Sim97b, Tót03, Tun91, WP03, CG73, Fus88, Lan91, Len87, MT84, Nos82, SW79, Tam87]. Number-in-Hand [PVZ16]. Numberings [HB75, MWY78]. Numbers [BMS86, Bac88, BCGW93, CK00, FLS04, KY09, Rup00, HJLS89, HH79, HH80, HJLS14]. Numeric [BC03]. Numerical [ABF^{+}99, ABKPM09, GMN09, KW00, Mil75]. Numerically [LMM99]. NV [Oya93]. NV-sequentiality [Oya93].

O-Efficient [MZ08]. Obfuscation

[AS16, BP15, BCG^{+}18, BDV21, GGH^{+}16, KMN^{+}22, SW21]. Object [CvdM00, HOS92, KQT99, Rut90, HW86]. Object-Oriented [Rut90]. Object-Preserving [CvdM00]. Objectives [Vég16]. Objects [AdBES14, ACHAH16, CIRR16, Efr05, EHS12, GGN10, YR92, dBG08, ZR79]. Oblivious [ABD^{+}07, BDFC05, Cha94, FMS05, FW14, FYO^{+}15, GM98b, GR11b, HHMW17, ISK^{+}12, Moi13, NP06, PU90, SS90]. Observations [Boo91, BMS13, Fre78, Sti90, Huỳ86b]. Observing [LSV07]. Obstacles [CKT00, HSY22, KM00, YLW95, Pro88]. Obstructions [MS08b]. Obtaining [GRS06]. Occam [Nat99]. Occur [ASSS16]. Occur- [ASSS16]. Occurrence [KST93]. Odd [CH87, DV00, Eri99, FR80, Sed78]. Odd-Even [FR80, Sed78]. Odlyzko [Fri86b]. Off [AKW00, LY13, PU90, AB79, Elk06]. offline [IM83]. Offs [Ben89, Vai89b, BBI16, BN20, FN99, VW85b]. Often [BFS06]. Old [CKT99, DW11, EGK^{+}14]. On-Line [Alb98, BNFN01, BGLL99, DT96, FG98, GW93a, GI93, KS95, LS98, LL92, LRWY99, MM93, RT93, SL96, SWW95, Tet99, ALM96, BFL06, LMSPR01, CCG^{+}97, CL91, Fre85, MR80]. Once [BCG20, BHH95, BC98, DZ97, GKS93, GM91c, Pon98]. One [AS08a, App13, BYJK08, Bis78, BCPR16, BH75, CM13a, CPS16, Dye86, ECM03, EV21, GKK^{+}09, GL13, HILL99, HNO^{+}09, HRV13, HO14, IL82b, IJ87, IJ91, IQ19, JDU^{+}74, Kar92, Kla07b, KMN^{+}22, KST93, KSS20, LL90, NS22, YI83, Bak78b, ĎG84, HT82a, HM81, MS87a]. One-Centre [Dye86]. One-Counter [GL13]. One-Crossing-Minor-Free [EV21]. One-Dimensional [Kar92, JDU^{+}74]. One-Reversal [IJ91]. One-Sided [AS08a, KSS20]. One-Tape [Bis78, LL90]. One-Variable [NS22]. One-Way [App13, BYJK08, BCPR16, CPS16, GKK^{+}09, HNO^{+}09, HRV13, HO14, IJ87, Kla07b, KMN^{+}22, HILL99, HM81]. Ones [ABH98, CLWZ04, ES15, LL90, AHMP87, NS80, Vel85, JKN08]. Online [Alb99, AAA^{+}09, AFG^{+}14, AIKW15, BBGR08, BC00, BM04b, BKS15, CEKP18, Cha03, CLT05, CFK^{+}07, CMS18, CLL19, CJST07, DH04, EHM^{+}12, EN16, EÖW14, EvS05, FSZ21, FKK^{+}08, GPT00, GGK16, GKR12, HLP17, HKSS17, Hit07, HL05b, IMP14, IKKP19, KRTV18, KMZ18, KLS^{+}24, MSVW16, MP03, MRSG05, RC03, SvSE03]. Only [LL99]. Open [AC82]. Operation [MST91, BEW80a, Blu86]. Operations [AR98a, BP90, KB76, Mul99, Pan80, TV99, AW96, BJR89, FL83b, OH87]. Operator [BGS07, CRGM81, KLR21, LMPP15]. Opt [BHZ23, CKT99, BKK^{+}04]. Optical [BFL06, GJLR97, GJM98, GMR99]. Optical-Communication [GJLR97]. Optima [Cha03]. Optimal [ABW18, AAK^{+}12, ABL09, AV03, ABD^{+}07, ABD^{+}05, AMMW07, ANSS21, AHR96, AIS08, ACK^{+}98, BR79, BNL04, Bas73, Bas76, BST23, BCR18, BFKL21, BKS92, BAFN99, BMKM07, BDMT98, BHNW23, BGP^{+}23, BW79, BH98, Bol81, BN00, BFvRV15, BDMI14, BW17, BGK^{+}18, BCG20, BN14, BFSU98, BMRV02, BT20, BCGR92, BA10, CK17, CR10, CR12, Cha91, CN18, CCK^{+}22, CHPJ22, CRS95, Cha92b, CKLW01, CGV11, CP12, CDO^{+}22, CR86, CT90, mCL04, CFW93a, CDK^{+}11, CV88, CSSS89, CGH^{+}98, CPP16, DKLR00, DGV05, DHK11, DLV98, DKR96, DHM12, DS00, EGS86, EvS05, EHJM00, FM97, FV16, FG98, FS20, FK15, FKK^{+}08, FGLS14, FY96, Gar74, GSV99, Gaz91b, GL80, GKK^{+}18, Gie95, GKKS93, GR21, Got81, GL17, Gus87a, HCD89, HF98, HSS13, HKSS17, HKL00, HM95, HS99, HSY22]. Optimal [HP98, Hol73, Hol74, HZ20, Hwa80, IKL^{+}04, IW22, Ita76, JáJ79, JMP11, JR94, KQT99, KKMS11, KU99, KMW75, Ked85, KKMO07, KK00, Kir83, KŠdW07, KRT03, KS95, Kre90, KV98, Lar87, LP83a, LL95, LL00a, LL00b, LRS17, LY05, Liu22, LS93, Lou83, LL92, Lu14, Mal05, MS76a, MPS92, MR01a, MPRS79, MO12, Moi16, NR95, NN14, NN17, NS22, Oht76, ORS86, PP08, PSW90, Par77, PU89a, PR02, Pio04, RR89, RT90, RS92, RS94a, RS91b, RT89c, RT89d, RT89e, Rhe90, SV95, SS11a, SW99, STU97, Sho92, Smi76, SR97, Urq11, VV17, Vit85a, YHC87, Yao94, Yu22, Zav77, Zel88, AB96, BH87, BN89, BH81b, Bit82, BG90, Cha84, CW83, CL91, Fre81, HT82a, HY88, HR87, MN88, RL88, RS91a, Ros82b, Vit85b]. Optimal-Time [CSSS89]. Optimality [DHIP07, GK18, KMP14, Par80, Par98, SY80, KMZ87, Zha95]. Optimally [AA18, CCK^{+}17, HT17, NGM97]. Optimistic [FMS05]. Optimization [AFB96, AU72, ABL09, BH18, BV06, CvdM00, CFW93b, CCC^{+}13, DFKL20, GJW24, Gär95, GKL13, GKLV14, GPRS11, Iba73, KM97b, KS97, KY15, Lue81, MS92, OPS04, PR99, Ron82, SS12b, Wei80, YYI12, Zim98, ASSU81, HS87, KP82, Ros82b, SS89b]. Optimized [IJKW10]. Optimizing [CBH84, GMN09, GLPS12, JK83]. Optimum [EMC85, Hos75, Hu74, LP94, MYSH77, TB84, VC85, AMS84]. Oracle [BG81a, BHIK97, BJ99, DLM22, DHV22, FFK96, Lut93]. Oracles [DTCR08, DP20, FHOS97, Gas87, HV20, Mer01, PR14, Lon85, Lon88]. Order [AdBMS98, ABI97, BP93b, CFJ^{+}10, CBCG05, Coo74, FR75, Gab77, GM09, HHW98, KRTV18, KMPHT14, KM94b, KMZ18, KKPV07, LNSS09, LSSV08, PST00, PVV98, PH97, RS07, YHC87, Yao89, Car84, Gra84]. Ordered [BN82, BST95, BCL75, FT88, GHSZ08, HL72, KM95, MPR22, PY79, Kni88]. Ordering [BRS98, GM79, GHM^{+}11, MMSA93, Oht76, Opa79, RR04]. Orderings [BT80a, CHPJ20, Lub87]. Orderly [CLL05]. Orders [GRS93, KW95, KV10, KV12, LS19, Sch16, dOS94, dBG08, FLST86]. Organ [KKM84]. Organize [Bit79]. Organizing [GMS81, Hen76, Fre84, OH87, TN82]. Orientation [BTW00, SV18]. Orientations [BB88, KS03a, WWW87]. Oriented [HT84b, MW99b, Rut90, CGJT80, Gol81]. Orienteering [BCK^{+}07, CHP08]. Orthogonal [BCR80, DLV98, HST05, KRSV08, Lem75, MW99b, Mor06, Vai89b, Wil85a, Gol81, SL80]. Orthogonally [BRV18]. Other [Aga92, BH77, BGK^{+}08, CRZ22, DY10, DORS08, EY10, FL94, GK07, HM73, KKMO07, MV75, Pat13, Sit21, Tom92, YYI12, ZR79]. Ottmann [PS91]. Our [GGL^{+}13]. Out-Degree [Sén05]. Outliers [HPW04]. Output [BP98, ES05, Gav72a, GM91b, Goo91, KB76, Iba78]. Output- [Goo91]. Output-Sensitive [ES05, GM91b]. Overflow [KLM^{+}04]. Overlays [KS03b]. Overloaded [KS95]. Overview [BLMW11].

P} [MS01, KR88, Rub91]. P-RAM [KR88]. PAC [GR17, Ser02]. Package [DFG13]. Packet [AS05b, AFHB^{+}00, AH08, BMIM07, CKMP95, CKL98, CG94, Gam03, PRS94, TU81, TS81]. Packets [BSH94a, EW12, VCJS22]. Packing [BX91, BBGR08, BCS10, BHI^{+}13, CK04a, CKL^{+}09, CGJ78, CGJ83, EHM^{+}12, EHPR17, EvS05, EL08, FL86, GPT00, GK07, IL98, JDU^{+}74, KRTV18, KY15, LC90, MW99b, NPSW23, Rhe90, RT93, Rot16, SvSE03, Sri99, Ste97, BS83, BX92, CGJT80, CL89, FK88, Gol81, Mao93, Mur87, RT89d, RT89e]. Packing-Covering [KY15]. Packings [BCR80, BI06, GLS^{+}24, Kel06]. Padded [AGG^{+}19, HK87]. Padé [CC86, LC89]. Page [CK93, ST72, Wes94]. Page-Fault [CK93]. Paged [MO74]. Paging [BGV00, KPR00, KL00, Smi76, IKP96]. Paintings [HLP17]. Pair [GZ05, GRSS98, Pel94, Sha85]. Pairing [BF03b, HST11]. Pairs [CL77, Cha10, COS99, DHZ00, GSSZ04, Gus90, HKN16, JáJ79, KKP93, LY13, RZ12, Wil18, BK10b, MT87]. Paper [AI77]. Papert [She18a]. Parallel [ABK00, AAK90, ALMZ99, ABR90, AALM90, Ara86, AC78, AES10, AH08, Aza92, BSH94a, BS76, BVY22, BV93, Bin84, BP93a, BG95a, Bis78, BH98, BS03, BG92b, BC98, BCGR92, Che93, CDTT95, CV88, Col88, Col93, CK91, CSW98, CC23, CGG^{+}97, CGH^{+}98, Csa76, CK96b, CLM^{+}20, DL91, DNS81, DGBL00, DKR96, Dow90, DKS88, DS00, Ebe89, Fin20, FS81, Gaz91b, GGK92, GMR98, GM98b, GPST92, GJLR97, Goo91, Goo99, GKS90, HCD89, HH81, Hai13, HS02, HT98, HSV05, Imm89, JS82, JC91, JT10, KMR87, Kao95, KLP96, KMV92, KK00, KSS93, KR88, Kle96, KM75, LP95, LMM99, LP03, Mac97, MPS92, Mil72, MR91, MT99, MN01, Mul99, MS90, NNS89, NSV94, Nis94, PSW90, PR93, PP95, Pan00, PY90, PY91, PR02, QS06, RY96, Ram96, Rao11]. Parallel [Raz98, Raz11, RT88, RS92, Rut90, Sch04b, SWW95, Sky76, Smi86, Sni85, Sti90, TV91, TV85, Tsa92, UY91, VTL82, Val82, WC92, YHC87, adHW87, vzG84, vzG87, AA88, AHMP87, AB96, ACG89, AV87, BOFKT88, BJR89, BN89, BG96, BG90, CKT93, CDR86, EMC85, FLST86, FRW88, GS89, Hag90, HY88, Hya79, KS91, Lub86, MC87, MRK88, NS80, RR89, Rei84, RS94a, Rei85, SW85, SWCP96, SJ81, SV84, TC84, VSBR83, VT89, VW85b, adH86, vzG86]. Parallelism [CD17, Mak97, Val75]. parallelization [SV88]. Parallels [HR78]. Parameter [CCK^{+}20, CHM13, CLP^{+}19, DDDR11, FBGSV15, FG01, Hav24, LPPS17, MR14, VHPT09, WBZ13, DF95]. Parameterizations [FGLS10]. Parameterized [Bak97, BM14, CCK^{+}22, CG07a, CL19, CHM13, FG04, FV13, FGLS14, FLM^{+}22, Jan10, KST99b, LDRT24, LMS18, MR14, MPP22]. Parameters [Epp03, OS11, WM97]. Parametric [AM93, CFKX07, GGT89]. Parametrized [DFVW99, Hli05]. paramodulation [Sta80]. Parentheses [MR01b]. Parenthesized [MMN14]. Pareto [DY10, MO12]. Parities [FGKP09]. Parity [CJK^{+}22, EWS05, GLV13, Hås14a, IK22b, JPZ08]. Parser [ABW18, AP72, PSW90]. Parsers [AU73]. Parsing [Raj96, SW76]. Part [MR91, SPH84b, CV88, CMS97, CMSS00, Col00, HPT99, HP99b, HS82, HS84, RS91b]. Partial [ACK^{+}98, Bro79, CFJ^{+}10, CH03a, CH06, Chi77, DCV90, Ger76, HF98, HR03, KW95, KT77, Pur78, Riv76, Sch04a, Sch81b, Yao85a, Yao89, Meh82, Pla80]. Partially [BN82, BWY15a, BCL75, MiHI99]. Partite [BCC^{+}09]. Partition [CC99, CPPW12, GGJT10, GJ13, GPW18, KS03b, PT87a, PR17, Cun86, Hol81b]. Partitioned [VL00]. Partitioning [AAEZ21, AEZ20, BFK^{+}14, BHK09, BNS18, BP92, ENRS99, FMS10, HRW20, Joh88, Kel06, KM77, KLL17, LLRS01, PSZ17, ST13, YDEP89, CR86]. Partitions [AS98, AGMV00, GGM^{+}15, HST05, Hos75, SW06a, Tót03, Tót08, WW75, Wil76, Cha84, FL83a]. parts [ACGS88]. Party [HT17, HNO^{+}20, SV08]. Pass [AKL21, CK09, CCRM86, KLM^{+}17, Alb85]. Pass-Efficient [CK09]. Passes [RU81]. Passing [CK96b, LAB01]. Passive [Pio04]. patching [Kar79]. Path [AFGN22, AFGN23, BJM00, BH03, BCM^{+}15, BST95, BGK^{+}08, CL77, CKT00, CKL^{+}09, CDH13, DP94a, DCH21, Fre76, FSS13, Gol08, GI80, HP99a, HS97, HY87, KKP93, KPS94, MTZC81, Meh77, Ned99, PR05, SS05a, SCY00, SV00b, TVZ22, YLW95, ALM96, Blo83, GS87, MT87, Yao85b]. Path-Evaluation [BGK^{+}08]. Paths [AW01, ABL^{+}02, AKW00, ACIM99, BK10b, BFKL21, Ber16, BH19, BF03a, BSW14, BFU94, BFSU98, CCE13, Cha10, CKS09, CHW13, CNVW08, CNVW10, CJ14, CPT96, CKN22, DY10, DHZ00, DFJ98, EN19, Epp98, Fri02, Gab07, GS21b, GKL13, Gol95, GR16, HC99, HKN16, HKN21, HS99, HSY22, HW95, IPS82, IY23, KKP93, KL15, KMV92, MS92, MW95, Mou90, NT05, RZ10, Rod10, RZ12, Sch98a, SS86a, Sha87, Spi73, SP75, VA00, Wil18, WZ15, dVE10, Coh98, Fed87, KS91, MT84, Pro81, Sch94]. Pathwidth [KS96a]. Patience [AMRU22]. Patricia [NS08]. Pattern [ABF94, BCC^{+}13, CH03c, CGH^{+}98, DVW21, Dun73, FLM^{+}22, FYO^{+}15, KU99, KMP77, SRR95, She11, Vis91, FL83b, Liu81, Yao79]. Patterns [CDK^{+}20, DVW21, MS07a, SY99, SY06]. Pauli [BL21]. PCA [FSS20]. PCP [DGKR05, DR06, GL23, GS00]. PCPs [BGS98, BSGH^{+}06, BSS08, BHPT24, DH13, IKW12, ST09a]. PDA [Gre74]. PDL [SPH84b]. Pebble [KAI79, VT89]. Pebbled [Dev99]. Pebbling [GLT80, HP10]. Pedigree [XLXJ09]. Pedigrees [CCC^{+}09]. Peeling [CCK^{+}17]. Per-Packet [AFHB^{+}00]. Perceptron [CBCG05, DSSS08, Ser02]. Perfect [AFB94, AJ22, ABT21, BMS86, BBD^{+}86, BC22, BN00, CRS95, CFSZ02, DDPY08, DKM^{+}94, EV21, FGT21, FBL03, GKK13, HR73, JW21, KW97, KM08, Kle74, Mic04, PPSS04, RSZ02, KMN^{+}22]. Perfect-Information [BN00]. Perfectly [KLN15]. Perfectness [HW04]. Performance [AS05b, AZ04, AF98, BC00, Che92, CGJT80, CH82a, EW03, Fra91, GS17a, Gol81, HT84b, JDU^{+}74, Koc92, LL00a, Nic88, PVG04, Ste97, BD86, LP83b, Mao93, Tur86, Yao85b]. Performing [CSW98, DHW98]. perimeter [KMR88b]. Periodic [HW95, SSY82, YI83]. Periodically [MHSR98]. Periodicity [AB98]. Permanent [AG94b, BŠVV08, IST12, JS89b, KKL^{+}93, MR12]. Permutation [BFP89, KKM84, KMMS06, RLDL96, Spi85, vzG91, BLS97, MC87, Ron82]. Permutations [AB94, CSW98, KST99a, RKS12, Vio20, LR88b]. Permuter [Ram84, Car79]. Permuting [FMP95]. Persistent [BNMSS98, KOT00]. Perspective [KPS94, Wil00]. Persuasion [DX21]. Persuasively [DSS90]. Perturbation [BL16]. Petal [AN19]. Petal-Decompositions [AN19]. Petri [Lie76, May84]. Pez [GKR05]. Phase [HN02]. Phenomena [HTC13]. Phrase [KS00]. Phylogenetic [CJL03, JNS06, JLL12, LSS13, SY12]. Phylogenies [KW97]. Phylogeny [AFB94, AC11, FBL03, PPSS04]. Physical [CK17, SMW07, Sel96]. Picture [OL74]. Piecewise [AAvK^{+}13, GMSV15]. Piecewise-Linear [AAvK^{+}13, GMSV15]. Pigeonhole [BOBP^{+}04]. Pin [Cyp93b]. Pipe [KKM84]. pipeline [PST88]. pipelines [Gab88]. Pivot [GMSV15]. PL [Ogi98, SPH84b]. Place [BCN^{+}20, FMP95]. Placement [BRS08, CW75, JR94, KMW75, KBP03, LP83a]. Placements [Pel94]. Plain [CLP16, CPV22]. Plan [He99]. Planar [AABC^{+}17, AI77, AMMW07, BFG^{+}19, BKM^{+}17, BP92, CM13a, CLX17, CF22, CN18, CL13, CKS09, CHH02, CHK03, CJ92, CL21, CPT96, CFHM20, CDTT95, DR95a, DFP13, FW98, FT06, FLM^{+}22, FJ89, GJT76, GSV99, GKM^{+}21, Had75, HCD89, Hag90, HP00a, HP00b, He93, HMRS98, IS79, JKLS05, Kao95, iKK18, Kir83, KS86, Kle08, LP77, Lic82, LT80, Liu22, LRS01, MZ08, MPP22, MNS86, MS92, MN95, NS07, Ned99, Pre81, RY96, Seb97, SPLK98, Smi86, TV91, Vad01, Ver05, BH87, BM88, CHT93, CNS82, CB81, Fed87, GB82, HJ85, HY88, HL96, Kao93, KS93, MNS85, MY91, Mye85, Pre79, Pro86, Rei83, RLWW97, Sch94, Sha85, Shi80, Sys81]. Planarity [BDMT98, CM13a, DT96, GT01, JS82]. Plane [AOS06, AdBES14, ABSD^{+}98, BGGS20, BP00, CKT00, FHH^{+}93, FHS96, He99, HM95, HS99, HSY22, IST12, MBCV98, O’R86, ACM84b, LD81]. Planes [BEGJ00]. Planning [AS90, AS97, BDL00, DFK^{+}19, FG92, FHS96, HOS92, Pel96, RW00, YS93, HW86, KRR94]. Plans [AGM13]. Planted [BHK^{+}19, FPV18]. Player [Meh18, Vid16, Vid20]. Playing [KKL09]. Plottable [Bra08]. Plural [WM97]. Plus [BGSW19, HLV18]. Point [AMMW07, CP09, CN18, CJ92, CL21, CPT96, Cla88, DHR20, EGS86, EKL10, FW98, Fil19, GPST92, GT98, HOS00, KS96b, Kap00, KRS10b, LP77, Pel96, Pre81, PT92, Rei99, TV91, VVYZ07, PT89]. Pointed [BEW80b, BKPS06, MD73]. Pointing [BFKV11]. Points [AdBS19, AHPY13, ASU75, CEG^{+}94, Dye91, EY10, GKL13, GKS93, JB90, KKN15, KKMS11, LM08b, MS91, OL74, Sny92, SOAD09, vKLM10, vzGSS03, Pre79]. Policies [CKR72, FKT17, Gam03, Yan82]. Poly [LMR90]. Poly-Log [LMR90]. Pólya [BFKV11]. Polygon [ABL^{+}02, CW98, HIKK01, KKMS11, KM00, Kei85, MS91, Ram94, Wil82, TV88b, TV88a]. Polygonal [Mit99, SY92]. Polygons [BDIG85, GHK91, IST12, IS90, ES88]. Polyhedra [AS96, AST97, BD92, BDD^{+}07, Cha92b, Dye91, KRS10a, Sha87, Cha84]. Polyhedral [AdBS19, AS96, SS86a, YR80, MD73]. Polyhedron [AS97, DF88, Mou90, VA00].

Polylog [BSS08, BRS98]. Polylogarithmic [Baz09, Coh21, FK02, GHH^{+}17, SS17]. Polymatroidal [CKRV15]. Polymorphic [BPWY99, Dob89]. Polynomial [AAEZ21, AFB94, ADG^{+}11, AM01, AI77, AEZ20, BL04, BBRS98, BGK16, BS15, BB21, BGGS20, BG92a, BP93a, BG95a, BP98, BH19, Boo72a, Boo79, BHIK97, BGS07, BGG22, BS92, BCC^{+}13, CKM^{+}14, CW04, CK96a, CK05, CM01, CS07, Chi76, CPPT24, CMST85, CGG01, CPPW12, DKK^{+}19a, DFM23, DS94, DRST14, DvM06, DPV15, DHR97, DS07, Ebe89, EJS05, EKL10, ESY17, Fat74, FBL03, FLN^{+}15, FPS16, Fu95, GH92, GM98a, GLT80, GLLON18, GJ13, GJ14, Gop08, GKS90, GJ19, HJLS89, Had75, HPQ17, HL04a, HNOS96, HHH98, Iba74, IKS10, IQ19, Jai07, JKLS05, JR20, JS93, Kal85b, Kam05, KB79, Kar99a, KM97b, KU11, KPV14, KS94, Kli79, KvM02, Kou77, Kur87, KOR98, Lad89, LR88a, LS95, LS02, LT91, MVW04, MZ13, Mit99, MPR03, MRR15, NP06, NSS00]. Polynomial [NS93, NW06, OS07, Ogi95, OW91, Pan94, PR17, PB85, RT92, Rei99, Rev75, RH78, Rub91, Sel78, Sho97, Sit21, SVV12, Tod91, TO92, Tze92, Val02, Vég16, Wal98, Wan92, Wan98, Win80, WLB^{+}99, AS80, BOFKT88, BG96, CJK^{+}22, Dub90, FG82, GG86, GB82, Hel84, HS88, Hor87, Huỳ86a, Iwa89, Kad88, Kad91, KF80, Ko89, LLK80, Liu81, Meg83b, Pla78, RY87, Tom82, Ukk83, WT79, Yes83, HJLS14]. Polynomial-Expansion [HPQ17]. Polynomial-Space [Iba74]. Polynomial-Time [AFB94, BL04, BHIK97, CKM^{+}14, CW04, DFM23, DvM06, EJS05, FBL03, GJ13, GJ19, JS93, KvM02, LS02, Mit99, Ogi95, OW91, PR17, Sho97, SVV12, Tod91, TO92, Tze92, Wan98, WLB^{+}99, Kal85b, Hor87, KF80, LLK80, Yes83]. Polynomially [Boo78, Leu98, Rao09, ORS86]. Polynomials [ASU75, BCGW93, BSK12, BCR16, BV10, BC76, BK90, BM02, Cus98, Duv91, EKR01, Fri90b, GJW24, Gan95, GKS10, Gro03, HSS13, HS75, IQ19, Kal93, KW93, KR06, Lan85, Lip78, Man95, PR17, PS73, Ris85, RB91, RS96, She20, Str74, Win80, vzG91, DL80, Hya79, KMS83, Lan91, Len87, Loo83, Pla80, PA79, PA80, Ren89, SS86b, VSBR83]. Polyphase [Zav77]. polytime [MW20]. Polytope [AAOS97, AdFM18, DHR20, KLR19]. Polytopes [BCMW17, GJW18, GTV21, LV22b, Ver09]. Popular [AWY18, AIKM07, LMS98b]. Popularity [Kav14]. Poset [Fri93]. Posets [DKM^{+}11]. Position [BCF^{+}14, CGMO14, MR80]. Position-Based [BCF^{+}14, CGMO14]. Positioning [CDK^{+}20]. Positive [AK09, BBR04, BKN09, BHIK97, CJM^{+}06, EKL10, MI97, SXB83, Spi73]. Possible [Meh77]. Postman [Jac90]. Postnikov [CKM^{+}14]. Postorder [Luc90]. Potatoes [CCK^{+}17]. Potential [AM04, EKLS16]. Potentials [Seb97]. Potts [GSVY16]. Power [ABK^{+}06, AS16, AIS08, Aur87, Bar84, BW05, BG81b, BN10, BCH^{+}20, BGKS12, BGWZ20, BT80b, CC97, CCH98, CI98, Cas99, CFGM16, CEEP16, CHPW98, CG97, DW82, DD13, EÖW14, GS17b, GGK16, GL21, HHR11, HHMW17, IJ91, Kal93, KTZ15, KDA00, KS17b, KS19, LSV13, LLV92, LFS00, Maa97, MGHK09, MSVW16, Mer01, MRRS07, MRRT08, Per74, Pra75b, RS03b, She18c, Sim97a, TZ17]. Power-Down [AIS08]. Powerdomain [Plo76]. Powerful [CLL08, LC05]. powering [PA79, PA80]. Powerless [Val86]. Powers [Fat74, HS75, Pip80, Yao76, vzG87]. PP [Tod91]. PPAD [Meh18]. PPAD-hard [Meh18]. PPSZ [Her14]. Practical [CS03b, Gol08]. Practice [HP00a]. PRAM [GMR98, KNP99, vzGS00]. PRAMs [DS00, HS02, HB94, Kut91, Nis91]. Precedence [Ado77, GL95, HL05b, LR21, LS97, LMPP15, MSS04, PT87b, QS06, SU05, Sve11, CRGM81, HCAL89]. Precise [Nor89, Pri79]. Precisely [CLX17]. Precision [CKS17, SCY00, BS86]. Precision-Sensitive [SCY00]. Precoloring [CSZ24b, CSZ24a]. preconditioning [Zel88]. Predicates [Wra78]. Predicting [Che92, Nos82]. Prediction [Gol78, HKSS17, KV98]. Predictors [Tai80]. Preemption [BNCK^{+}99, BH75, CI98]. Preemptions [BS03]. Preemptive [Che93, DGBL00, FG92, LS84, Sch04b]. Preface [CP76, Gal88, Lew72]. Prefetching [HSV05, KK00, KV98]. Prefix [BJK^{+}94, BP90, GY96, KLP96, Rei72, CW83]. preflow [CM89]. Preprocessing [AST99, FH11, KPV14, vKLM10, Ryt80]. Presburger [Che76, IL81, NP22]. Presence [BPWY99, BE17, DPVR12, DHW98, GR15, KL93]. present [Ott86]. Preservation [ADG08, Ben77, Dur10]. Preservers [AB24, Bod21]. Preserving [BH81a, Bol76, CvdM00, DPPR20, LL78, PP95, Ros75b]. Price [ADG^{+}11, ADK^{+}08, AAE13, BBM13, BUV18, CGGS19, FRGH^{+}23]. Pricing [BK11, CHK13, CGV11, CDO^{+}22, DFKL20, GR16]. Primal [ANFSW20, Ezr16, GKPS08, JV08, KRTV18, KR05, RV98]. Primal-Dual [ANFSW20, GKPS08, KR05, RV98]. Primality [Leh82, SS77b, SS78, AL82]. Prime [HSS13, KY09, Pra75a, Sho97]. Primitive [BN99a, RC79]. Primitives [CK17, HZ77]. Principle [KL72]. Principles [BOBP^{+}04]. Printability [Rub91]. Printable [FGLM98, AR88]. Priori [SS95]. Priorities [BP10]. Prioritized [EFN18]. Priority [ABD^{+}07, AB94, CGS99, GR85, McC85, Tho00]. Privacy [BBR88, BUV18, CRS14, DLWZ14, GRS12, JR13, NTZ16, Ull16]. Private [BCNW09, BHMO22, BN14, CGGK95, EGOS07, FX15, GR02, GR05, HHR^{+}16, KMMS22, KKMO00, WY07]. Privately [GHRU13, KLN^{+}11]. Prize [ABHK11, AABV98]. Prize-Collecting [ABHK11, AABV98]. Probabilistic [BS93, BK07, BD86, CF86, CHPW98, DS90, FM97, FKS93, FS07b, Gil77, Gol98, LSH05, ŁN06, LSV07, Mac98, Mac99, Mit85, Rab80, Rei85, SPH84a, Tze92, AL88, CG88, CL89, HS85b, MRS88, Pap81, Rei84]. Probabilities [ESY17, GHK96, Jac72, LW86]. Probability [BG81a, CRS15, NN93, PB83a, Tho18, UY91, HT82a]. Probable [FK03, LP83b, Tur86]. Probably [CPP16, Nat91, Nat92]. Probe [CR10, FN93, PD06, Păt11, PH97, Vio12a, Kni88, SS90]. Probing [ES88, PPR09, TZ12]. Problem [AA18, AKK^{+}09, AT77, AFB94, ACK^{+}99, AT00, AOT89, Alb98, AAA^{+}09, AZ07, AGL99, AIGT94, ALW08, ACM^{+}99, AKL21, ABH98, AF23, ACBFS02, Bak86, BBGR08, BHI^{+}13, BYGNR98, BHK^{+}19, BGK16, BKHL00, BB95, BGKS12, BKK^{+}10, BHZ23, BNS18, BvMR^{+}00, BKS15, BA10, yCLZ00, CKR04, CL77, Cha89, CHJ18, CKT99, CK05, CKS13, CC99, CLWZ04, CHP08, CL19, CLL19, Che05, Che12, Chi77, CGK07, CS03a, CPPT24, mCL04, CDDS89, CFW93b, CFCH^{+}00, CDH13, CGGR16, CPPW12, Dav86, DFHS08, DH92, DCH21, Dye86, DR13, EK05, Elk06, EGH07, ENZ00, FL12, FMS02, FR06, FZ18, FSZ22, FMRT96, Fre93, Fre76, FKR95, GM91a, GS17a, GGG18, GJT76, Gaz91a, GLT80, Gol95, GRSS98, GM18, GKS08, Gra90b, GR16, GMM09, GMR97, GS07, HRTS03, Hal02, HL04a, HL05a, HW02]. Problem [HL87, HIKK01, ISK^{+}12, Jan10, JR20, Joh88, Kar99a, Kar92, KMW75, KM94a, Kav14, KS98, iKK18, KKR06, KPV14, KMV92, KSS93, KMP^{+}20, KPST94, KLR03, KMMT09, KR07, Koz09, KC94, Kup05, LS02, LWW21, LSS^{+}21, LT82, LS93, ŁN06, MM18, MSS07, Mal05, MPP22, MNR90, MMM22, MP03, Mic04, MMP87, MBCV98, MW99b, NZ01, NSS00, NH90, NV03, NSH^{+}08, OS11, OR21, Opa79, PS77, Pap92, PR99, PU89b, PSWvZ17, RY96, Ram94, Ram96, RRSS09, Rey77, RW03, Rod10, RC03, Sén05, SSSV01, SS05b, Sit14, Sot13, SSY90, SSS91, SW94, TN95, TB84, TV22, Wal79, Wan99, WZ21, Yao78, YR80, AKR95, AKPW95, AL88, BLR80, Bit82, Blu86, CHT93, CV85, CF86, CNS82, CMST85, CK80, CP84, CK87, ES91, FG82, Fri86b, GS89, GB82]. problem [Gra88, Gur82, GS87, Gus88b, Hal75, HT82a, Ham83, Huỳ86a, Iba78, IL82a, Kar79, KK86, LVW84, Lub86, MT84, May84, MNA88, Mye85, Pap81, RLWW97, RSL77, SCH92, SRP83, VV89, Vel85, VR89, AW97]. Problems [ANIM96, AGvM^{+}18, AALM90, AHP08, AK03, AGK^{+}04, Asa87, BRS08, BCS10, BGNV18, BNL03, BJM00, BFL06, BK16, BP20, BHL18, BCH86, BCNW09, BS00, BŠVV08, BD93, BG93a, BJKP05, BKRS00, BMPP19, BT06, Boo78, BCHS94, BEG^{+}02, BGWZ20, BGG22, BDK00, BK11, BGK^{+}08, CLX11, CRZ22, CM99, CM13b, Cha98a, CRS95, CG05, CEG^{+}93, CK04a, CS18, CVY19, Che97, CCH^{+}16, CN06, CO82, CM16, DGV05, DISW13, DFG13, DY10, DL76, DFVW99, Dun73, DHHM06, DS00, EG95, EvS05, Eri99, ET76, EIS76, ES22, FT88, FH06, FPV18, FSZ21, FG04, FGLS14, FHK78, FS07b, FSS13, GT89, GJW24, GH92, GK07, Gär95, GK12a, GHSZ08, GJK^{+}22, GW95, GPST92, GK12b, Gra94, GS02, GS17b, GLV13, GPRS11, HS85a, Har09, Hav24, HU75, HLCW91, Hon82, HJW84, HMRS98]. Problems [HR03, Iba73, IJ94, IKS10, IS90, JKP20, JR93, JR94, JL95b, KST99b, KS96b, KAI79, KPR^{+}13, KH83, KKL04, Kou77, KD79, Lad89, Lev86, LW18, Loh06, LMS18, Lut95, LZ00, MS10, MYZ06, MHSR98, Mar99, MNS10, Mar08, MT78, MTZC81, MT83, MS92, MV13, Mit99, MS97, Mol03, MS07b, MR07b, MS01, PT09, QS06, RR04, Reg04, RS92, Sah74, SY91, Sel96, Set75, SSW94, Sit21, SH96, Sub94, The02, Val75, Val79, Vég16, Vin04, Von13, WLL08, Wei80, WWW87, YLW95, Yan81b, Yan81a, YYI12, Zim98, Zuc96, vDHI06, vzG84, AK88, BS83, BCD^{+}89a, BCD^{+}89b, CSY87, FJ81, FS89, Fri87b, FK89, GHY89, Gus87b, HITK79, Hoc82, Hol81b, Hsu95, Huỳ86b, IA86, KS96a, KP82, Lag85, LW88a, Lub81, Lue81, MC87, Meg83a, MS84]. problems [Pla78, RS94a, Ros82b, RY87, SJ81, SS81, Sha85, SH85, SS89b, TC84, Yao82, ZS89]. Procedure [DS02, FR75, KN85]. Procedures [BKPS02, PS12]. process [PCHM85, SPH84b]. Processes [Ace94, BDP02, ESY17, FPP11, GL13, Hen88, Tra72, BKO88]. Processing [RH93, BW79, Mit85]. Processor [AM96, Bas73, Bas76, BS03, BH75, CKL98, GK99, GJ77, GS73, GJ86, HM87, KMV92, LS98, Tsa92, AW96, BD81, CL79, Fri84, Kao93, KS93, MM83, SC79a, VV89]. Processor-Ring [CKL98]. Processor-shared [GJ86]. Processors [CS80, Dob84, FL83c, GM98a, GIS77, Goo91, Mal05, MS90, Set76, Fri87a, Gab88, HS88, KMZ87, VSBR83]. Product [BATS11, BGSW19, BCM99, FOS08, IJK09, IJKW10, IKW12, KŠdW07, Ram96, Raz03, RS03a, She12, She18b, Shp03]. Production [CFJ^{+}10]. Productions [Yao89]. Productivity [Wan92]. Products [GGR11, GV19, HLV18, HS82, HS84, MV75, NRS99, RS10c, SW19, TT01]. Profile [KC94, KLL17, LS97, LST99]. Profiles [PHNS09]. Program [AM75, BGS72, Cha76, CD72, Coo78, Mil72, OC75, RT82, RS96, YY14, AM80, Coo81, HT82b, HCS80, Wey79]. Programmable [SW10, RL88]. Programming [AM96, ABL09, ACFW98, BBR04, BGWZ20, Cha05, CCWZ18, CCK^{+}22, DHK11, DS00, FT97, JS20, KTZ15, Li89, Li20, LMR90, MWY78, MR99, SPLK98, Sri06, TSSW00, Ye90, Meg83a, Meg83b, PS81, Ros82b]. Programs [AU72, AW76, AK09, BW05, BCG^{+}18, BDFP86, BRRY14, BCG20, CL94, CLD73, Che92, Che76, CS05, EJ99, FS81, HR93b, HN94, IL82b, Mar99, Mis77, New02, NW06, Pol95, Pon98, RV98, Rei14, Sag88, SPH84a, Sri99, SOAD09, Car84, CI83, Dye84, HS85b, HR86, IL82a, PS81]. Projection [Rao11]. projections [vzGS98]. Projective [FSS20, Ye90]. Promise [BGWZ20, BG21, CZ23, KOWZ23]. Proof [ACG^{+}16, ABSRW04, BMM21, BEGJ00, BDP02, BJSW20, Col00, DR06, DT97, Ger76, GK96, Gol98, GS12, HV20, Lov21, MPR22, FN80, GMR89, GO80]. Proofs [AAE^{+}23, AFRV19, AF23, BPU92, BOHP14, FLS99, HT15, IKOS09, KvM02, KLN15, LTW18, Mic00, Nor09, RRR21, SV10, Wil16, vMS05, GHY89]. Proper [DHH96, HSS01, KST99b, KS96a]. Properties [Aar03, AS08a, AFNS09, ALN^{+}12, Aur87, BV06, BSHR05, Ben77, CKS02, CK12, CC83, CP05, CSS09, FN07, FR24, GPSS06, HR93b, HL14, HHN04, HJ90, KSS20, Lie76, MWY78, McI74, TZ22, Tri94, Val11, Wad76, Yao88, Zim98, BORW88, CGH^{+}88, FL83b, Gal84a, Rom82]. Property [AS08b, Bis78, BFNR08, CS05, DRTV18, FPS16, GR11a, HK07, HW04, JK82, NS13, CRGM81]. Property-Testing [HK07]. Prophet [DFKL20]. Proportion [Che01]. Propositional [ABSRW02, ABSRW04, CZ06, Lad77, LTW18]. Prospects [Mis77]. Protecting [FRR^{+}14]. Protection [Kao97]. Protein [FW07]. Protocol [FM97, HT17, Hav04, MW99a, MS17]. Protocols [BTY94, CRV10, CK96b, DIM97, FK00, GJKP04, HLR96, HIK^{+}11, HHRS15, HNO^{+}20, HHL18, KLL^{+}15, KLR10, PR08b, PTW11, SV00a, CR86, JKS84, Wil86]. Prouhet [GGG18]. Provability [Lad77]. Provably [AGKM16, Bak86, CK01, KW00, SC79b]. Prover [FK00, VV17]. Provers [BOHP14, KRT10]. Proving [AW76, Bra75, CW04, CLD73, GS00, GKM83, MP87, Pet83]. Proximity [BSGH^{+}06, Bol76, CGM^{+}22, GR11b, HPK14, HPK15, LPT98, Rie99, Rie02, Ros75b]. Proximity-Oblivious [GR11b]. Pseudo [AS05a, BBS86b, BCG20, LS16]. Pseudo-Approximation [LS16]. Pseudo-distributions [BCG20]. Pseudo-Line [AS05a]. Pseudo-Random [BBS86b]. Pseudodisks [CHW13, MRR15]. PseudoPolynomial [BGK20]. Pseudorandom [ABSRW04, App13, BV10, BRRY14, BCG20, GKL93, GMRZ13, HILL99, HRV13, HU22, Lut93, MZ13, NRR02, RT88, Vio07, Zuc19, BM84, LR88b]. Pseudorandomness [GKM18]. Pseudotriangulations [BKPS06]. PSPACE [HP10, MHSR98, Mei13]. PSPACE-Completeness [HP10]. PSPACE-Hard [MHSR98]. PTAS [CHJ18, CK13, FRS19, GR16, Kho06]. Public [BBR88, CS03b, GS88, NS12, PTW11, SZ00]. Public-Coin [PTW11]. Public-Key [CS03b, GS88, NS12]. Puiseux [Mur90]. Puiseux-Series [Mur90]. Pumping [EPR81, Var97]. Pure [BPP89, JS20, PB85]. Purely [Fri90a]. push [CM89]. Pushdown [BGS72, CC83, LLS84, Ram84, Rei72, Car79, RH78]. Putnam [BKPS02]. Puzzles [LPS20, Maz76]. Pyramid [MS87b].

Q [SJ05]. Q-Heaps [SJ05]. QMA [BJSW20, BG22, GK12a, GN16]. QMA-Complete [GK12a]. QMA-Hardness [BG22]. QoS [KLM^{+}04]. Quadratic [ET93, NS09, Sag83, TW14, Lic87, Lic89, PST88]. Quadtrees [CH03a, LM12, DL90]. Quality [BP10, CD03, EW12, MV00]. Quantified [BC10, Che08, CD16]. Quantile [GM09].

Quantitative [BLS84, FKKN11]. Quantum [Aar06, AD14, AA18, Aar20, ADH97, AvDK^{+}07, ATS07, ABO08, AE15, ACG^{+}16, Amb07, ACR^{+}10, Amb16, AdW22, AL10, ABP19, AGS21, Bac21, BYJK08, BBD^{+}97, BOHP14, BBBV97, BV97, BJKP05, BNT^{+}19, BGK^{+}18, BG22, BP02, BJ99, BDH^{+}05, BFNR08, BCF^{+}14, CK17, CK12, CKS17, DFSS08, DPVR12, Dru15, DS02, DHHM06, FIM^{+}14, GKRdW09, GKK^{+}09, GL23, GN16, HHKL21, HRTS03, KY10, Kla07a, KŠdW07, KLN15, KLN19, Kup05, LM08a, MSS07, MNRS11, Mah20, Mah22, MS08a, MN01, MRRS07, MRŚ10, Reg04, RT10, Rei14, RV01, SG04, She12, SZ08, Sho97, Sim97a, TS11, Val02, Vaz97, Wat09, Zha09, dW03, vDHI06, vDMMS07, vT04]. Quartet [SY12]. Quasi [ACOH^{+}10, BS00, BY98, CJK^{+}22, CPPT24, FGT21, Fre78, IM20, JR20, MW20, MRR15]. Quasi-Convex [BY98]. Quasi-Gray [Fre78]. Quasi-NC [FGT21]. Quasi-Polynomial [CPPT24, JR20, MRR15, CJK^{+}22]. Quasi-polytime [MW20]. Quasi-Randomness [ACOH^{+}10]. Quasi-Uniform [IM20]. Quasi-Varieties [BS00]. Quasilinear [Fu95]. Quasirandom [FGS12]. Quaternion [KV10, KV12]. Qubit [KMP^{+}20]. Queries [AAK^{+}06, AAK^{+}12, AWZ00, ABOK20, AdFM18, BLR14, BL00, BV96, BN10, CvdM00, Cha92a, CGL97, CKT00, CH03a, CH06, Cla88, CHKZ03, CH05, DS07, Eri00, FH11, Fu95, GKV06, GS17b, HPK13, HHH05, KKPV07, LPT98, Liu22, Mer01, PVV98, Sch04a, Ull16, Vai89b, Wil85a, YDEP89, Bol81, CBH84, JK83]. Query [Aar03, ABP19, AGM13, BSS08, CFK^{+}21, Cha86, CVY19, CGS22, CK12, DW82, DH13, DHHM06, Efr12, GPW20, Gop08, GHRU13, HHW98, HY20, IKW12, KSS20, LM08a, MRS18, Pag01, Rei14, She12, Wag90, dW03].

Query-Answering [Cha86]. Query-Efficient [Gop08]. Query-to-Communication [CFK^{+}21, GPW20]. Querying [CNVW10]. Question [BGS75]. questions [Wey79]. Queue [ABD^{+}07, BFG^{+}19, Bro78, DFP13, GMR98, HPT99, HP99b, LLV92, PS91, PR06]. Queue-Read [GMR98]. Queue-Write [GMR98]. Queued [HSV05]. Queueing [BGL04, Gam03, Lav73, LS75, MSV87, CRS87]. Queues [AB94, BST95, CGS99, HR92, MS99, Tho00, HR87]. Queuing [ÀBS04, Kon75, KR78, LPSR04]. Quick [Tho05]. Quickest [FS07a]. Quickselect [MR01a]. Quicksort [HN02, MR01a, Sed77]. Quiescent [ACT00]. Quittable [GHKT12]. Quorum [MRW00, NW98]. Quotas [RRW01].

Rabbit [HPK14]. Rabin [ACGS88].

Radicals [Lan92]. Radii

[GKK^{+}12, GHK91, VVYZ07]. Radio [BYII93, CGK07, CD18, De 10, KP04, KM98]. Radius [Mic04, ST23]. Radon [Bra98]. RAM [BCG^{+}18, KR88, Mor06, Tho00]. Ramanujan [BSS12, BATS11, MSS18, MOP22]. Ramsey [Coh21]. Random [AM06, ART09, ABL09, ABS07, ABK^{+}06, AK97a, ACRT99, App13, AL18, BES80, Bac88, BKM19, BAG01, BG81a, BLR^{+}15, BBS86b, Boo94, BMS13, BFSU98, Bro79, BLTV99, CRRZ24, Cha00, CM13b, CGV13, CH03a, CH06, CO10, COF14, CFLS97, DF18, DR95c, DR95b, Dev98, Dev99, Dev02, DN04, DKM09, DKR96, Fei06, FF93, FPV18, FSW87, Fri90b, FGK05, FMM11, FGRV14, GM01, GGGY21, GS17a, GJW24, GW17, Gil98, GKP04, GGN10, GM09, GHM^{+}11, HTC13, JS05, KL15, KRTV18, KHNS05, KS13, KMZ18, KSS20, LL92, Mac97, Mer01, Mol03, MS07b, MS04, NS93, PY91, RT89c, Rhe90, RT93, Ros14, SV08, Sch87, Sch90, Sie04, Sot13, SZ99, SZ76, Tam98, Ver09, Wal79, Wei80, Wor87, CDR86, DL90, Fri87b, Fus88, GG86, HC83, IP87]. random [KvLP88, Lue81, Meh82, RT89d, RT89e, SV84, VV89, Yao79]. Random-Access [DKR96, IP87]. Random-Assignment [Sot13]. Random-Order [GM09, KRTV18]. Random-Self-Reducibility [FF93]. Random-Walk [FMM11]. Randomization [AT98, AH08, CJST07, DD13, MMS17]. Randomized [Alb98, ANIM96, AW96, ACH10, BBN12, BK15, BCG^{+}18, BKRS00, CI98, CR10, CKP19, CKL^{+}09, CDO^{+}22, CH95, Che97, CCH^{+}16, Cla88, CGG^{+}97, DvM06, Doe04, Dot10, DR14, Gaz91b, GKKS93, GRSS98, HHR11, Kar99a, Kur87, KMRZ98, LM08a, LMSPR01, Mac97, Man95, MRS18, NSV94, PS97, PR02, RS92, RS94b, SW99, Wes94, Zha09, AAG^{+}89, AA88, KRR94, RR89, RS94a, Zha95]. Randomizing [BG93a]. Randomly [AH11, GJ99]. Randomness [ACOH^{+}10, BIW06, BEG17, BT00, CRS95, CK00, DHN02, DSY10, GR05, GS06, KS01, MS17, NS93, PU90, RSW06, SU09, Wan98, CG88, Huỳ86b]. Randomness-Optimal [CRS95]. Range [AMS13, AS10a, BCM99, Cha00, FH11, KRSV08, Lar14, Mor06, NS08, PT07, SS11b, Vai89b, Wil85a, YDEP89]. Range-Efficient [PT07]. Rank [Bsh90, CM22, DKM06b, HST11, HiO08, Lan14, Meh18, Moi16, RS10b, SS11a, BC87, ERV81, dG83]. Rank-Decompositions [HiO08]. Rank-Pairing [HST11]. Ranking [AH89, BN99a, CCWZ18, CV88, FW76, FJ84, GS91, GLW82, Li86, Tro78, Wil76, FJ90b]. Ranks [AL83, PRS97]. Rapid [CLV23, Cop82, Ros80, Bak78b]. Rapidly [CDG^{+}06]. Rate [ABO08, AIKW15, FMS05, HRZW20, LS75]. Rate-based [FMS05]. Rates [BGL04, CK93, FK00, SS89b]. Ratio [BD97, BLN01, BHZ23, FKK^{+}08, GW93a, HKK^{+}07, Hum89, Mes10, TV20, CCG^{+}97]. Rational [BCGW93, Che89, GKS94, KW93, KK77, PW04, RS91c, Str74, WP03, BS86, BCN81, Gal84a, Gal84b, HH79, HH80, Kal85a, Loo83, vzG86]. Ratios [FP78]. Ray [Aga92, AM93, AS96, CPT96, ES22, MS97, dBG08]. Rays [IST12, ES88, MD73]. Razor [Nat99]. Reachability [AB24, BCR18, CHKZ03, Fin20, RZ08, RZ16, May84]. Reaching [Hie10]. Read [ABK^{+}98, BCG20, BHH95, BC98, DZ97, FW90, GMR98, GKS93, GM91c, Pon98]. Read- [ABK^{+}98]. Read-Once [BCG20, BHH95, BC98, DZ97, GKS93, GM91c, Pon98]. Readies [BKO88]. Ready [DD81]. Real [BNGNS01, BP98, Bra08, CG97, CK94, EGH07, FR75, HHKL21, HJLS14, HH72, Hoo90, KS95, KDA00, PR05, Ren92, Ris85, BOFKT88, HJLS89]. Real-Time [BNGNS01, HH72, KS95, KDA00]. Real-Weighted [PR05]. Real-World [EGH07]. realistic [AHMP87]. Realization [Gel75a, Gel75b]. Realizations [Hen92]. Realize [Che76]. Reals [KM13, KHNS05, PVV98]. rear [OH87]. Rearrangeability [Çam03, LDHX99]. Rearrangeable [DGHK98, NV03, Ngo05]. Rearrangement [AHK^{+}10a, PB83b]. Rearrangements [BP96]. Reasoning [HvdMV04, IISI05]. Reassignment [DDMM05]. Recognition [BHIK97, CCKM13, CPS85, FT88, Iba74, Kos75, MSZ11, VTL82, Hsu95]. Recognizable [WS78]. Recognize [aBC08]. Recognizing [CFSZ02, FS86a, FGK05, HSS01, HM99, JO92, KMMS06, MMN14, KM89]. Recombinant [XLXJ09]. Recombinations [CCC^{+}09]. Recommendation [DFG13]. Reconfiguring [CMS97, DFK^{+}19]. Reconstructing [ALRS98, FHK^{+}88, GKS10, SY12, WW06, XLXJ09]. Reconstruction [AM01, BDMI14, BMS13, EMS03, JR13, KPS13, KW95, PW04, SS10, WP03]. Record [ER80]. Recourse [MSVW16]. Recoverability [GLS^{+}24]. Recoverable [KLR19]. Recovery [Ara86, CK01, GLPS12, HRZW20]. Rectangle [BJK^{+}94, CDL86, KMR88b]. Rectangles [AGMV00, AES10, GLM^{+}16, KMR22, Tót08, KMR88b]. Rectangular [AB98, BHPT24, Cop82, EMSV12, He93, Kun93]. Rectilinear [CKT00, GT01, GAOPL^{+}21, SS05b, VW85a, YLW95, YS93]. Recurrence [GJMM00, HN02, LR89]. Recurrences [BGS07, Fid85, KLS^{+}24, LR88a]. Recursion [HR86, SMSR75]. Recursions [BN99a]. Recursive [BV93, Car84, DS76, KS01, Muc76, Pro81, SP82, Ver97, AM80, HR86, RD81, RS83]. Recursively [Kum96]. Red [AS90]. Red-Blue [AS90]. Reduce [AH08, TY84, TY85]. Reduces [BHL95]. Reducibilities [GH92, Mer02, Yes83]. Reducibility [AHOW92, CSV84, FF93, HU72, HS95, KKMO00, KPR^{+}13, OW91, Sel78]. Reducible [BK88, Boo94, Fu95, Sel88, Sol76, Sha79]. Reducing [CK00, CKR72, DH92, HW86, TVZ22]. Reduction [AGU72, BKS92, BADTS22, BGA23, CP19b, EWS05, HT98, KRS12, KS15, Wad78]. Reductions [BG93a, BT06, DR14, Hir23, LZ00, LL78, MR07a, Moi13, Kal85b, Ukk83]. Reducts [BMPP19]. Redundancy [CDK^{+}20, Pag01]. Redundant [BSGK^{+}10, XLXJ09]. Reed [CW07, GGG18, GLS^{+}24, KRZSW23, ST23]. Reference [FP78, ST72, IKP96]. Refined [GSVY16]. Refinement [CD03, MvV92, PT87a]. Refinements [BEGJ00, KL72]. Refining [KF80]. Reflexive [Vik03]. Regime [BKM19, GLLZ19, HL14]. Region [EHS^{+}19, Lar77, SY92, Bol81]. Regions [BH13, CNVW08, CNVW10, CK95, ELRS03, MS92, HJW85, KM96]. Register [CW94, FP86, RS85a, Set75]. Registers [BOC92, SWPL11]. Regression [CDMI^{+}16, DDH^{+}09]. Regular [BKM19, BN99b, BRRY14, CT99, DK14, EPR81, GW17, GGM12, GGM11, GKK13, Goe93, GJ13, HY87, HR78, Iwa93, KT15, Lev76, MNS10, MW95, Urq11, Vad01, Var97, aBC08, Bab80, PCHM85, SH85]. Regularity [AFNS09, ACOH^{+}10, CR00, FPS16, GGM11, HNR08, KRT03]. Reingold [Car79]. Rejection [Per74, Wil73]. Related [AGvM^{+}18, AALM90, AHP08, BGNV18, BCH86, BD93, CKS02, CW75, CRS95, CEG^{+}93, CS18, CK13, DS00, EG95, ES22, GSVY16, GPST92, HLCW91, HR03, MNR90, Mit99, MS97, MS01, RS92, Sah74, CSY87, FL83a, LW86, Meg83a, RS94a, Rom82, Yao82, ZS89]. Relating [AHOW92, Bor77, RI89]. Relation [GMWW17, Wad76]. Relational [ASU79, AGM13, BMSU81, BPSV04, SU77, ASSU81, HITK79, Hul86, IIK85, Sag83]. Relations [BG84, Dob05, EWS05, FRW88, GRS93, HJLS14, VL00, HJLS89, SW79]. Relationship [CR05, FS07b, MW99a]. Relative [BG81a, BEGJ00, FFK96, Hul86, Wra78]. Relativizations [BGS75, Bal85, BLS84, GG86, SB84, SXB83, Kur85, TW89]. Relativized [Ko89, Hel84, KF80]. Relativizing [BWX82]. Relaxation [BGWZ20, CVZ14, Mur90, Mur95]. Relaxations [AM13, ESV20, FK03, KMR22, Li20, TZ17]. Relaxed [CDG^{+}09, CGS22, Fil19, GPT00, GL21, CS89]. Relaxed-Voronoi [Fil19]. Release [FR94, GJST81, SC79a, Sim83]. Releasing [GHRU13]. Reliability [GJ19, JR94, Kar99a, SSS91, Val79, Pro86, SW85, SWCP96]. Reliable [ACT00, PY87].

Remainder [BG95a, BG96]. Remarks [Fab77, McK79]. Rematch [CCWZ18]. Removal [Fil19]. Removals [Sch13]. Remove [KKN15]. Removing [EC95]. Renaming [AF01, ACHP19]. Renewal [OC75]. Reordering [ACER22, EÖW14]. Repacking [BBGR08]. Repairman [Sit21]. Repeated [RSW06]. Repeats [Sch98a]. Repetition [BVY22, Hai13, Rao11, Raz98, Raz11]. Replacement [FRT93, MVW04, MO74, WE80]. Replication [CK04b]. Reporting [AKS14, Cha00, Mor06, SS11b]. Representability [YR92]. Representation [ADS86, DL91, DHH96, Esw75, MR01b, Gus88b, HH79, HH80, Kun93, RI89]. Representations [BBC^{+}22, BGG22, Bür00b, HRTS03, Iba73, KKLV11, NN14, ST94, vzG86, Har80]. representative [IIK85]. Represented [AO95]. Representing [HSS01, BT80c]. Require [BSGK^{+}10, SV10]. Required [BD76, BD78a]. Requirements [CG94, KK77, Ord89]. Requires [CKPR02]. Resampling [HV20]. Reset [Epp90]. Resettable [CPS16]. Residual [KL15, SSS91]. Resilience [BL16, BN00, LRS17]. Resilient [CHJT04, DRTV18, KZ07, NS12, SR97]. Resistance [AH11, CGP^{+}20]. Resistances [SS11c]. Resistant [GKW20, GHM^{+}11, Pio04]. Resolution [AR08, BKPS02, BBI16, BS09, BKPY18, BN20, BGT16, BEGJ00, CM13b, CVZ14, CadHS00, DK15, FSZ21, FHH^{+}93, Gal77, Goe93, KL72, MS07b, Nor09, RU98, SBI04, Urq11, Sta80, Wil86]. Resolved [JLL12]. Resolving [ABOK20]. Resource [AST99, BvMR^{+}00, BL02, BFL02, DIM97, GG75, GJ75, Hit07, JR94, KK06, WL95]. Resource-Bounded [BvMR^{+}00, BFL02, Hit07]. Response [Ber76, GLP75, SLW^{+}98, JS89a]. Restarts [CJST07]. Restricted [ACHAH16, BBGR08, CGV13, Gar74, GMP23, GKM83, JR20, KT94, Lar87, Man08, ML00, CMST85, CR80, KMZ87, Sag83, Wey79]. Restricted-Use [ACHAH16]. restricting [Lon85, Lon88]. Restrictions [SBI04]. Restructuring [VL00]. Resultant [Pan94]. Results [AMS11, AST76, BGS98, BCR16, Bod21, CP09, CKT99, CJ92, CKN22, EGM03, EGK^{+}14, FKS93, GSVY16, GJ75, GKV06, IJTW95, KTW99, KKMO07, KZ20, LPSR04, MT83, Mil72, PST00, Sch04a, SWK09, TS81, Tri94, DF95, IM83, IIK85, Mit85, NHL82, Pet83, Ukk83]. Resynthesis [LFS00]. Retraction [Vik04]. Retrieval [CCFM04, KV86, Kou77, NN17, Riv76, SW94, Wil82, WY07, Bol81, CW83]. Returns [CKLW01]. Reuse [Kal95]. Revenue [CHK13, DFL^{+}22, JLTX20, JJLZ22]. Reversal [BNP74, BT05, CBGWW13, eCY91, IJ91, KLL90]. Reversal-Bounded [BNP74]. Reversals [BP96, GMP98, KST99a]. Reversible [Ben89, LS90]. Revisited [AHS92, BN99a, BFL02, CHP08, FFTC20, FS86b, HK87, KPS94, LMSPR01, LT82, NW93, Spr02, YLW95, AG86]. Revisiting [SW19, VP03]. Reward [BCK^{+}07]. rewrite [DMT88]. Rewriting [Ges02, GKM83, YH90, Oya93]. RH [IL04]. Ribbon [AKK^{+}09]. Rich [PT09]. Riemannian [BRLW19]. Right [BEM10, Gra74, Nor89]. right- [Nor89]. Rigid [BHPT24, GJ00]. Rigorous [FN99]. Ring [CKL98, FS92, MV75, PRS94, AAG^{+}89]. Ring-Sum-Expansions [FS92]. Rings [HM91, HP98, Tun91, HS87]. Risch [Dav86]. Rises [CGW16]. River [LP83a, SD88]. RL [HZ20]. RMRs [GHW10]. RNA [BGSW19]. RNC [RV98]. ROABP [AGKS15]. Robbers [AGG^{+}19]. Robin [DMV04]. Robot [AWZ00, CP05, DRW98, GMR97, KMMT09, AK88, HJW85]. Robots [AP06, BDL00, CFPS12, CP08, DFK^{+}19, FYO^{+}15, ISK^{+}12, SY99, SY06]. Robust [BSGH^{+}06, BP00, BDMS13, CDK^{+}20, CDMI^{+}16, CTX10, DKK^{+}19a, DKK^{+}19b, HPK14, KMW10, LPT98, LH00, RS96]. Robustly [BK16]. Robustness [BD92, BHT98, Rub99]. roommate [Gus88b]. Root [BCR16, GK98b, KS94, MvV92]. Root-Finding [KS94]. Rooted [BH80, GH97, JNS06]. Roots [CJL03, BOFKT88]. Rosser [Ott86]. rotations [Wil89]. Round [BSS21, BGK^{+}18, CPV22, CLM^{+}20, CDP21, HHRS15, LPS20, RRR21, SV08, AB79]. round-off [AB79]. Rounding [AKOT03, Doe04, FNR15, GG12]. Roundoff [Mil76]. Rounds [ABT21, CKPR02, GM98a, LPSPP05, NW93, Pip87]. Routing [AFHB^{+}00, AIK^{+}20, AAE13, Bak86, BFL06, BJK^{+}94, BM04b, BFvRV15, BHV99, BT05, CM99, CKR01, CBGWW13, Chu16, CKN22, CG94, DKKP99, FHK78, FJ89, Gam03, GP01, GJM98, GKR05, JC91, KM94a, KPST94, KNPS24, LLRS01, LP83a, LSZ11, MS99, PU90, SV00a, SD88, Szk99, WLB^{+}99, FJ90a, GL87, Jaf85]. Row [CCK^{+}22, Sib98]. Row-Invariant [CCK^{+}22]. Row-Major [Sib98]. Rows [CDG^{+}06]. RSA [ACGS88]. Rudimentary [Wra78]. Rule [BPP89, DF18, PB85]. Rules [DS76, Ges02, JK86, Pla80]. Ruling [BBO22, Kho06]. Rumor [CHHKM17]. Run [LV06]. Runs [BII^{+}17].

Safe [BL00, Spr02, Yan82]. Salesman [ACK^{+}99, AM01, BGK16, BHZ23, CKT99, FKR95, FS07b, FSS13, MPP22, PS77, Pap92, SS95, TV22, CP84, Fri87b, HT82a, Kar79, RSL77, SRP83]. Salesman-Based [AM01]. Salesmen [AABV98]. Salesperson [ACM^{+}99]. Sample [DGR99, FX15]. Sampler [KM08]. Samplers [BFKV11, DHK^{+}21]. Samples [BMS13, CRS15, CFGM16, HMR18, HTC13, Nat99, SY12]. Sampling [AJ22, AK23, BHMO22, BC22, BCM99, CGR18, Cha00, Che92, CDRR07, CDK^{+}11, CDG^{+}06, DDH^{+}09, DLM22, Eft16, FVY21, FRV10, GJ99, GPRS11, IM20, JW21, KL15, MR01a, MRRS07, MRS08, MS04, PRS23, SF11, SS12b, Vio20, Vis91, Wat14, WE80]. Sampling-Based [SS12b, SF11]. Santa [Sve12]. Santha [BEG17]. Sat [AGH17, AM06, CO10, FGK05, GS17a, GN16, Her14]. Satisfaction [BYGNR98, BK16, BP20, BMPP19, BGWZ20, BG21, BGG22, BM14, CRZ22, CM13b, Che08, CVY19, DR13, FV98, FH06, GJK^{+}22, GS17b, KOWZ23, MS07a, Mol03, MS07b, Vik04]. Satisfiability [BCHS94, FPV18, FMdR10, Fra91, GKMP09, IK22a, KS98, KST93, PSWvZ17, CF86, Iwa89]. Satisfy [ABK^{+}98]. satisfying [FHK^{+}88]. Satterthwaite [FKKN11]. Saving [GS80]. Savitch [Gef93]. scalar [BEW80a]. Scale [AAZ16]. Scaled [Hit04, CC86]. Scaling [ABL09, BPS10, GT89, Gol95, Iwa03, KLR21]. Scan [CKT93]. Scan-first [CKT93]. Scanning [CKR72, CH82a, STU97]. Scatter [ACM^{+}99]. Schedule [CKP14, HS85b, KK86, LY89]. Scheduler [CB74]. Schedulers [FMS05]. Schedules [CS80, GIS77, LY05, Sve12]. Scheduling [AC82, Alb99, Ann19, BCS08, BP10, BP14, BNGNS01, BNL03, BYHN^{+}06, Bas73, Bas76, Ber76, BJR89, BL98, BS03, BH75, CI98, CLT05, CLL08, CMNS01, Che93, CM01, CH87, CK13, CJST07, CKMP95, CGJ78, CGJL85, CV88, DGBL00, Dob84, DW85, DH04, EÖW14, FR94, FL83c, Gab88, GW93a, GG75, GJ75, GJ77, GJST81, GRS05, GL95, HT21, HLL95, HM87, HL05b, HCAL89, IMP14, IKKP19, IM20, Jaf80, JP05, JT10, KS77, KP05, KS95, KDA00, LS77, LR21, LS98, Li20, LS97, MSS04, MRT07, MS90, MRSG05, PY79, PT87b, QS06, Rou04, RC03, SLW^{+}98, Sch04b, Set76, SLC91, SL96, SSW94, SWW95, SU05, SSY90, Sve11, Tsa92, VCJS22, WC92, WL95, Whi90, BD81, BW79, BD86, CCG^{+}97, Fri84, Fri87a, HS88, Kun81, LS84, NHL82, SC79a]. scheduling [Sim83, SW89, VV89]. Schema [GMN09]. Schemas [AM75, Cha76, MNS10, SI87, Wey79]. Schemata [CD72, Mil72, Muc76, Hul86]. Scheme [AAK^{+}06, BGK16, CK05, CM01, CAdVdM21, DKKP99, GMY12, Han96, HL04a, Jan10, Kar99a, Kle08, KS00, KR07, Mit99, MRR15, SVV12, Val82, WJG00, WLB^{+}99, YHC87, GMR88, HS88, HCS80, Mur87, Pro88, Yao81]. Schemes [AK03, BMSU81, BK15, BGS72, CM87b, CVZ14, CPPT24, CAKM19, CS03b, DDDR11, EJS05, FSZ21, FH11, FKLS10, HW93, JKLS05, KKKP04, Sag88, Sit21, TA99, AM80, HR86]. Schnorr [KHNS05]. Schrijver [BPS07, FK03, GMPT10, Hav24, PS12]. Science [AEM^{+}11, AK07, ANR09, BBS13, BSO14, BCAKT23, CR17, DT08, DMK19, HIMW16, KRS20, LMOT14, LU10, MV18, Rou15]. score [KM96]. Scoring [Sch98a]. Scott [Wad76]. SDD [KMP14]. SDDM [DPPR20]. SDPs [GMPT10]. Search [Aar06, AM93, AFB96, AAK90, Ann19, Ara86, AGK^{+}04, BCNW09, BG94, BAFN99, BST85, BG93a, Cha86, CAKM19, CGM^{+}22, DR95c, DR95b, Dev02, DN04, DKM09, Dye86, DS00, FG98, FN93, FW14, FS86b, FKK^{+}08, FRS19, Gar74, GR93, GMS81, Got81, HPK13, HPK14, HPK15, HW87, HY20, Kao95, KBP03, Kir83, KPS94, KS00, Kni88, KR05, KOR00, LW98b, LZ22a, LSV13, LS22, LM94a, LAB01, LST99, MNRS11, McC85, Meh77, Meh79, MMR95, NS08, NR73, OPS04, PS77, Pol95, Pro92, RS94b, SY91, ST94, Tar72, Wil13, Zha09, CKT93, Hag90, IA86, LW86, MR96, PY88, PB83b, TN82, Wil89, Zha95]. Searchable [ANSS21]. Searches [MM93, Smi86]. Searching [AMS13, AM85, AHHP00, AF03, Bei90, CR10, Cha98b, Che92, DL76, FLST86, HS97, Hir80, LMP02, Lar14, McD88, MS76b, RS91b, Riv77, SS11b, SSW03, Sni85, SY92, Wil85b, AG86, Cha88, GO80, HW86, JS89a, RS91a, Ryt80]. Searchlight [SSY90]. Second [CBCG05, LSSV08, TZ12, BSO14, CCRZWN22, PKS^{+}13]. Second-Order [CBCG05, LSSV08]. Secret [IQ19]. Secretary [DK14, FZ18, FSZ22, Sot13]. Secrets [AARV21]. Section [AEM^{+}11, AK07, AKS09, ACHCP21, ANR09, AIM16, BBS13, BSO14, BCAKT23, CCRZWN22, CDG11, CR17, DKI18, DMK19, DKSS22, EvMP^{+}12, Gol98, HIMW16, IKM^{+}12, KRS20, LMOT14, LU10, MS16, MV18, PKS^{+}13, Rou15, SK18, Vaz97, VNA20]. Secure [ABT21, CPV22, CS03b, HIK^{+}11, IKOS09, IKK^{+}11, KLR10, GMR88]. Security [AFRV19, BCHK07, CLP16, CPS16, HHT97, IJ94, KLR10, MS17, RS10c, Gus88a, MRS88]. Seed [GRS06, TS11]. Seeds [CRS14]. Seek [CKR72]. Seen [CRZ22]. Segment [mCL04, Mye85]. Segments [AMS98, BP00, BDD^{+}07, CEG^{+}93, Goo91, LSH05, PS91, Tót03, Rap89]. Selecting [CEG^{+}94, DZ99, JM78, Pip87, YY82]. Selection [CSSS89, DKM^{+}11, FSZ21, FK15, FJ84, GR17, Hya76, Maz76, Pip91, SV08, TB84, Yao80, ALM96, FJ90b, JM96b, Rei85, RD81, RS83, Wil86]. selectively [TY84, TY85]. Selector [HHN04]. Self [AKK^{+}09, ACG^{+}05, ACC^{+}11, BGKP99, CD17, CMS14, DIM97, DHM12, Dot10, EKS00, FF93, Fre84, GMS81, HS95, Hen76, KS20, LM08b, MOOY02, Rub91, SW10, Sel88, SW07, Tam98, Var00, vDMMS07, TN82, ST86]. Self-Adjusting [KS20, ST86]. Self-Assembled [SW07]. Self-Assembly [AKK^{+}09, ACG^{+}05, CD17, Dot10, SW10]. Self-Embedding [Tam98]. Self-Improving [ACC^{+}11, CMS14]. Self-Organizing [GMS81, Hen76, Fre84, TN82]. Self-P-Printability [Rub91]. Self-Reducibility [HS95]. Self-Reducible [Sel88]. Self-Similar [LM08b]. Self-Stabilization [Var00]. Self-Stabilizing [BGKP99, DIM97, DHM12, MOOY02]. Self-Testing [EKS00, vDMMS07]. Selfish [BFG^{+}07, CKV10]. Semantic [BRV16, Rut90, Coh83]. Semantics [May81, dB84, BT83a, Don79, FL83b, Hen80]. Semi [Cha03]. Semi-Online [Cha03]. Semialgebraic [AMS13, FPS16, SS11b]. Semidefinite [MR99]. Semidirect [TT01]. Semigroup [AAK^{+}12]. Semigroups [Boo78]. Semijoins [BG81b]. Semilinear [BC03]. Semiring [IS90]. Semisorted [AM85]. Sense [BV99, FRS03]. Sensitive [Boo72b, BCM99, CHPJ20, CKPT19, ES05, Ezr16, GM91b, Goo91, KS05, SCY00]. Sensitivity [DRST14, DRT92, GR02, GP18, VY23, AMS84]. Sensor [CTX10]. Sensors [CP08]. Sentences [Tun91, Gra84]. Separable [CM87b, GMSV15, KM97b, Vég16]. Separated [AOS06, GZ05]. Separates [AA18]. Separating [AKSW22, BW05, Blu94, Boo91, BFvMT00, Leu98, Nor09, She09, RS81]. Separation [BYJK08, BG98, CS99, CKP19, GKR21, GKK^{+}09, MRS18, PS02, Urq11]. Separations [AARV21, GKRdW09, SG04]. Separator [EGIS98, GSV99, LT80]. Separator-Based [EGIS98]. Separators [BT01, FHL08, FW07, KK98, MZ08, Pla90, Ord89]. Sequence [BG95a, BH13, GKL13, Jac72, Kun74, MYSH77, NN14, NSH^{+}08, SS17, BG96]. Sequences [BNBK^{+}89, CMSS00, DLS81, ES81, Epp90, Gel75b, IISI05, Jac72, KW94, Sch16, SSW03, ST94, Tom92, Alb85, BM84, Pro78, Sut91, YI83]. Sequencing [Ado77, DD81, ELMS^{+}12, BD78b].

Sequential [Bea91, GS17a, IKM85, MS90, BC82, Gur82, TN82]. Sequentiality [EHS12, Oya93]. sequentially [BC84]. Serial [BB95]. serializability [RSL84]. Seriation [ABH98]. Series [BT80b, CDTT95, Kal93, Mur90, VTL82, EMC85, FLST86, SW85, SWCP96]. Series-Parallel [CDTT95, EMC85, FLST86, SW85, SWCP96]. Server [AHKL07, BP10, BNFN01, BKRS00, CKV10, Sit14, AKPW95, HR87, KRR94]. Servers [KM08, CL91]. Service [BP10, BNMSS98, EW12, Gav72a, CRS87]. Session [RW03]. Sessions [AZ04]. Set [AGKS15, AAA^{+}09, AK97b, AIGT94, AKL21, BCS10, BFK^{+}14, BBFT23, BYGNR98, BGGS20, BHNW23, BHK09, BS93, CCK^{+}20, CL19, CPPT24, EHM^{+}12, Esw75, ENZ00, Ezr16, FK03, GGP11, Gav72b, GGM11, GM82, GKKS93, GJW18, GGL^{+}13, HL05a, Hir80, HU73, HPZZ21, Joh88, KB76, Luc90, MRRT08, MRR15, RV98, RLDL96, Rub91, SZ00, She16, TT77, Blu86, CNS82, FT83, GS89, Hoc82, HT82b, JK86, Kin86, LP83b, Lub86, MN88, Pre79, Rap89, Sha85, WT89]. set-covering [LP83b]. Set-Maxima [GKKS93]. Set-Multilinear [AGKS15]. Set-Partitioning [Joh88]. SETH [BI18]. Sets [AOS06, AMS13, AGKS15, ASSS16, ASU75, AHOW92, ACRT99, BBS86a, BBO22, BN82, BGG18, BCG^{+}21, BFS06, BH77, BB80, BK88, BJ82, BHT98, BCL75, BC03, Cha92a, CG01, CPY89, Cus98, DY10, DW11, DKSS13b, DFJ02, EPR81, Esw75, FT88, FF93, FW98, FT06, FGLM98, For79, FL94, Fu95, GPSS06, GPSZ08, GHJY91, HOS00, HH91, HR97, Hit07, HZ20, HL72, Kap00, KRS10b, Kum96, Lev76, Nat91, NS22, Ogi95, OW91, Per74, Rie99, Rie02, Rot17, Sel88, Sol76, ST94, TIAS77, VVYZ07, Var97, vDMMS07, AR88, BORW88, JKS84, Kur85, LLK80, ORS86, Ukk83, Yes83]. Setting [ABOK20, Epp03]. Settings [vMP19]. Setup [Che93, FGS85, BD78b, HR87]. Seventh [DKI18, DMK19, FGKO08]. several [FJ81, RSL77, adH86]. Shadow [Aar20]. Shallow [AES99, AGS21, ES15]. Shallow-Light [ES15]. Shape [HPW04]. Shaped [HOS92]. Shapes [Dot10, GMRZ13, SW07]. Shared [ACFW98, BGV00, CadHS00, GK97, GMR99, MPS92, Nis94, PPS00, SWPL11, BN89, GJ86, MM83]. shared-memory [BN89]. shared-processor [MM83]. Sharing [FKLS10, GPRS11, KLSvZ08, NR95, GJ86]. Sharp [AM06, Aus10, CRRZ24]. Shatter [Ezr16]. Shattering [CLP20]. Shelf [BS83]. Shellsort [Cyp93a]. Sherali [AM13, TZ17]. Shift [RS85a, vDHI06, Nor89]. Shifted [BGKS12]. Shifts [MRRS07]. Shiloach [EiC17]. Shooting [Aga92, AM93, AS96, CPT96, ES22, IST12, MS97, dBG08]. Shop [AC82, SSW94]. Short [ALS22, BSS08, BGRV98, CM22, CGP^{+}20, CRS14, HT15, MW88, NP22, RU98, TS11]. Short-Term [MW88]. Shortcuts [DHP13]. Shorter [BGH^{+}15, BSGH^{+}06]. Shortest [AFGN22, AFGN23, AW01, ABL^{+}02, ACIM99, BK10b, BFKL21, Ber16, BH19, BCM^{+}15, CCE13, CL77, Cha10, CKT00, CHW13, CNVW08, CNVW10, CJ14, CPT96, DY10, DHZ00, EN19, Epp98, Fre76, GS21b, GKL13, Gol95, Gol08, HP99a, HC99, HKN16, HKN21, HS97, HS99, HSY22, HY87, Jac90, JL95a, KKP93, MS92, Mou90, PR05, Rod10, RZ12, SCY00, SS86a, Sha87, Spi73, SP75, Swe99, TY97, VA00, Wil18, WZ15, Blo83, Fed87, Hor87, MT87, Pro81]. Shortest-Path [CL77, HS97, Blo83]. Shrinkage [Hås98]. Shuffle [Çam03, EL86, Mor08]. Shuffle-Exchange [Çam03, EL86]. SIAM [BX92, BG96, CCG^{+}97, Col93, Coo81, FJ90b, HH80, HN79, Kad91, Lan91, Lon88, RS94a, SWCP96, Vai90]. Side [AdBS10, CRZ22, DPVR12, DLL^{+}03]. Side-Effects [DLL^{+}03]. Sided [AGM^{+}11, AS08a, KSS20]. Sieve [MRŚ10, PST88]. Sign [CM22, GJ14, KMP^{+}20, RS10b]. Sign-Rank [CM22, RS10b]. Signature [O’R86, GMR88]. Signatures [CGW16, PP97]. Signed [KST99a]. Signs [GGJT10]. Similar [LM08b, Lub81]. Similarities [ACKS15]. Simple [AFRV19, BCN^{+}20, BBS86b, BM02, BD06, CRRZ24, Che76, CW98, CDP21, Eft16, ELRS03, Fre93, GLPW16, GM98b, Gus83, Gus90, HB75, HU75, HN94, HZ20, HM99, HS90, HL72, ILPR81, JLTX20, JDU^{+}74, KOT00, KKMS11, KM00, LV91, MWY78, MS99, MW95, Mit99, Ned99, PSWvZ17, Rod10, SY91, TY84, TY85, ZS89, CI83, Hul86, IL82a, Lub86, Rap89, TV88b, TV88a]. Simpler [ACG^{+}16, CLNS15, CH02, GK07, Her14, KST99a, Kei85]. Simplex [BNS18, DH20, KMSZ05, Ver09, ZS02]. Simplices [Pel94, Pel96]. Simplicity [Bal85]. Simplification [Lan92, SMSR75, SV88]. Simplified [IJKW10, KTZ13]. Simulated [BŠVV08, Val02]. Simulating [Dev12, GK98a, HHKL21, MS08a, RT10]. Simulation [And99, BL04, BP15, CPS16, EWS05, GMR99, HU22, Mac98, RV01, SV84, Vit85a, AHMP87, KvLP88, Kin86, Vit85b]. Simulations [ACRT99, CL94, CadHS00, HB94, LL90, LL92, PR06, IP87, adH86]. Simultaneous [BGKL03, Hås88, PY91, CDR86, Lag85]. simultaneously [HM81]. Sine [PM03]. Single [Ado77, Ala14, AKL21, BCR18, BGK20, BDMT98, DDDR11, GMM09, HKN21, Juk06, KLM^{+}17, KTW99, KM08, KS02, MV13, WZ15, Blu86, CK87, HL96, IM83, Ked85]. single-chip [Ked85]. single-operation [Blu86]. Single-Parameter [DDDR11]. Single-Pass [AKL21]. Single-Source [BCR18, BDMT98, HKN21, KS02, WZ15, HL96]. single-tape [IM83]. single-valued [CK87]. Singleton [ST23]. Singular [GL23, WM97]. Sink [BKM^{+}17, Che08, GMM09]. Sinks [BKN09, MN95]. SINR [ABOK20]. Sixth [Bab06, MV18, SK18]. Size [AHS92, AGvM^{+}18, ACR^{+}10, AES10, ALS22, AGM13, BYCDM92, BH12, BS09, BP90, Bod21, Bor77, BCE95, BM14, CFKX07, CT00, CHM13, Cyp93a, Dev99, Dev12, EM03, Ezr16, FSS20, Gal77, GJ85, GJMM00, HHKW07, Kav14, KMV91, Kla07b, KvM02, Koc92, KRT17, Lar77, LM99, LSS^{+}21, MR14, MR08, PS91, PRS23, Pur78, RRSS09, RSY08, RT90, RT89c, Rhe90, Fri86a, Fri87b, HT82b, Kal85a, KS85, Lon85, Lon88, IPS97]. Size-constrained [Dev12]. Size-Depth [BCE95]. Size-Hierarchy [LSS^{+}21]. Size-Popularity [Kav14]. Size-Sensitive [Ezr16]. Size-Space [BS09]. Sized [AN93, EL08, FL86, LS18, SvSE03, Mur87]. Sizes [EGH07, MSS18, RT93, RT89d, RT89e]. Skeletons [AGG^{+}19, OL74]. Sketch [LWW21]. Sketches [CGP^{+}20, CTX10, HT21]. Sketching [AKR18, GLPW16, MNS11]. Skew [HKP14, Jon87]. Skylines [HTC13]. Slice [Val86]. Sliding [BO10, DGIM02]. Slightly [LMS18]. Slope [CSSS89]. Slowdown [GM20]. Small [ACG^{+}16, ABS07, AES10, BFK^{+}14, BPU92, Bie90, BD93, Bod96, BCH97, BDGJ99, CLS19, Coh95, DLM22, DY10, FR24, GW93b, GH97, GNS20, GNW23, Hås14a, HSS13, Hit04, HZ20, KW00, KRY96, KW98, Lov21, Mar99, Mar08, NN93, New02, NTZ16, Pla90, RS10a, RS22, Rao09, Ros18, SBI04, Bol81, KS96a]. Small-Bias [NN93]. Small-Depth [BPU92, CLS19, Hås14a]. Small-Size [AES10]. Small-Success [HZ20]. Smaller [CRSW13]. Smallest [Fre97, GG12, HR93c, RG77, Tre12]. Smart [SLW^{+}98, SLW^{+}98]. Smarter [LH00]. Smith [Ili89a, KB79]. Smooth [KS20]. Smoothed [AV09, BCM^{+}15, DH20, MO12, Ver09]. Snapshot [Hav04]. Snapshots [AR98a, DHPW99]. SNP [BMM21, FV98]. Social [ACKS15, CG18, FRS^{+}20, MR10]. Soft [KTZ13]. Software [Bak97]. Solomon [CW07, GGG18, GLS^{+}24, KRZSW23, ST23]. Solovay [AL82]. Solution [BM14, Kon75, LR89, MPR03, Nic88, NS22, PR93, SP82, Bin84, CC82, Fri87b, LW88a]. Solutions [ABL09, BEW80a, BEG^{+}02, FPV18, GGGY21, GK12b, HNOS96, JP05, KMW75, Kre90, MO12, MS04, Mur90, Nat95, Ren92, SVV12, WZ06, AK88, BLR80, GI80]. Solvability [CN99, Jay98]. Solvable [AR02, BK16, Vin04]. Solve [AT98, GHKT12, MS92, MRRT08, SY91]. Solvers [LMM99]. Solving [AdW22, BPSV04, Cha89, Hås88, Hal75, HS85a, JPZ08, KM13, KMP14, Koz21, LT91, AS80, Ili89b, Zel88]. Some [AW21, BC91, BSHR05, Ben77, BS00, Boo91, BG82, BMS13, Cus98, DFVW99, DHHM06, Fab77, FHK78, Hon82, Huỳ86b, IM83, Iga77, IIK85, LMS98b, LS98, Lub81, McK79, Mil72, NR13, Pla78, RR04, Rom82, SD88, Sti90, TS81, Tri94, WWW87, vDHI06, AK88, Fre81, Hol81b, LVW84, MS84, SJ81, SS89b, Fat74]. Sometimes [LL99]. son [LW86]. Sort [Che01, Col88, DY75, Col93]. Sorted [AHHP00, FJ84, SSW03, BT80c, FJ90b]. Sorters [SS89a]. Sorting [ABK00, ALS22, BP96, CMSS00, Cyp93a, DKM^{+}11, DY75, ES81, FT88, GKKS93, GP83, Goo99, HH81, HS02, JB90, KST99a, LP98, LM99, LSH05, Liu72, MSV06, MS76b, Par77, Pio04, Pip87, Sib98, Tan78, YHC87, YY85, Zav77, adHW87, AA88, AV87, BN89, BC82, RR89, Rei85]. Sound [Mic00]. Soundness [Coo78, Coo81, KLX10]. Source [BCR18, BADTS22, BSRZ15, BDMT98, BKM^{+}17, CCE13, Coh16, Coh21, HKN21, KS02, WZ15, HL96]. Sources [ACRT99, BJS97, BIW06, BKN09, BEG17, GRS06, KZ07, MN95, Rao09, SZ99, Vio14, CG88]. Space [AGMV00, AHPY13, AKS14, ABSRW02, AS97, ANSS21, Bak78a, BGNV18, BBRS98, BE98, Bea91, BTY94, BBR^{+}99, BBI16, BS09, Ben89, BN20, BL98, BGT16, Boo79, BWX82, BL93, Bor77, BFadH^{+}87, BM99, BHV99, CR80, DvM06, DH92, Edm98, Eri00, EWS05, ES22, FN99, FLN^{+}15, FH11, FJ90a, GS80, Gef91, GMP98, GLT80, HPK14, HST05, HSS09, Hon82, Iba74, JTT00, KŠdW07, KOR98, Lad89, LS90, LL90, LR96, LP13, Mac98, MOOY02, MRSV21, NW06, Nor09, PP08, Par77, Pel94, Pfa83, SCY00, Tót03, Tót08, Tri08, Vai88, Vai89b, YDEP89, Yao94, Zhu21, vMP19, BC82, BKRU94, CI83, Gef93, IM83, Imm88, Sys81, Tom82]. Space-Bounded [BL93, LL90, vMP19]. Space-Efficient [BGNV18, BL98, FH11, Mac98, BHV99, FJ90a]. Space-Time [Eri00, Vai89b]. Space-Width [BN20]. Spaces [ALN^{+}12, BM04a, BCG^{+}21, BCM99, Che09, KM97b, KOR00, NN93, SS86a, Yao82]. Spacious [Nor09]. Span [BW05, Rei14]. Spanner [CHD21, EP04, FS20]. Spanners [AdBS19, ABP18, BGJ^{+}12, BCC^{+}09, BDMS13, CLNS15, CLPR10, GLN02, KPX10, KMRS18, KP98, Coh98]. Spanning [ABN15, AN19, Aga92, CRT05, CT76, CT00, CLL05, CT90, CEF^{+}05, CS09, DRT92, DGGW20, Elk06, EEST08, ES15, EP09, FL12, Fre97, FGRV14, Gab77, GM78, GG12, GKP98, GH97, HKRT95, HL04a, HK01, HLCW91, Hu74, KR95, KRY96, KR02, Kun74, LPSPP05, MMS17, NSV94, NSH^{+}08, OL74, PR00, PR02, STU97, SP75, Vai88, WLB^{+}99, EMC85, Fre85, HY88, KS93, KIM81, KMZ87, Yao82]. Spanoids [DGGW20]. Spans [Hit04]. Sparse [AHOW92, ABCP98, BBS86a, BGG18, BCR16, BK88, CG01, DG93, DFJ02, FT97, FMS02, For79, Fu95, GLPS12, GKS90, GKS94, GLN02, HR97, JO92, KW93, KPST94, Kur85, LS95, Lov21, Man95, Nat95, OW91, PR93, RB91, Sem09, Sol76, SU77, Tho05, Vad01, CKT93, PA79, PA80, Ukk83, Yes83, BHL95]. Sparsest [AMS11, AHK10b]. Sparsification [AdW22, CX18, CGP^{+}20, DPPR20, EGIS98, FHHP19, KLM^{+}17, LS18, Moi13, SS11c, ST11]. Sparsifiers [BSS12, EGK^{+}14]. Sparsity [GOS^{+}11]. Spatial [AJ22, BGG^{+}19, GKV06, GMP05, KKPV07, LSH05, PT92, SS90]. Special [AGK^{+}09, AEM^{+}11, AK07, AKS09, ACHCP21, ANR09, AIM16, Bab06, BBS13, BSO14, BCAKT23, CCRZWN22, CDG11, CR17, DKI18, DT08, DMK19, DKSS22, EvMP^{+}12, FGKO08, GS21b, GS06, Gol98, HIMW16, IKM^{+}12, KRS20, LMOT14, LU10, MS16, MV18, PKS^{+}13, Rou15, SK18, Vaz97, VNA20, Che86]. Specialization [CLD73]. Species [MLZ00]. Specific [BC76, GR17, DL80]. Specification [Pig91]. Specifications [AFL08, RH93, YR92, BT83a]. Specified [MHSR98]. Spectra [Fri91, HN77, KT15, Gra84, HN79, TN82]. Spectral [AK97a, ABH98, Cha98b, CGP^{+}20, KSV08, KLM^{+}17, Kel06, KLR21, LZ22b, LS18, PSZ17, SB13, ST11, BS92]. Spectrum [KM93]. Speed [BPS10, CLL08, DKKP99, FT06, Kos75, MS87a, PP95, RS85b]. Speed-Up [FT06, PP95, MS87a]. Speeding [AU73]. Speedup [AdW22, BH98]. Speedy [CLL08]. Spheres [MS97]. spilling [BJR89]. Spin [AJ22]. Spira [CL77]. Spirality [DLV98]. Splay [CMSS00, Col00]. Splicing [Gat92, Kim97]. Split [BYHN^{+}06, MNA88]. Splitting [GPSZ08, LY13, San75, vKLM10]. Splitting-Off [LY13]. Splittings [BHT98]. Spot [MS17]. Spot-Checking [MS17]. Spreading [CHS12, CHHKM17]. Squad [CDDS89, DHM12, GK12b]. Square [GJ00, Gia95, HL04b, JV75, KV86, LM12, SRP83]. squarefree [WT79]. Squares [AF23, BHK^{+}19, AB96, CL89]. Squaring [LM12]. Squarish [DJZC00]. squashed [CS89]. SRT [MP05]. ST [CDTT95]. ST-Digraphs [CDTT95]. Stabbing [Aga92, AAK^{+}12]. Stabbing-Semigroup [AAK^{+}12]. Stability [ÀBS04, ADK^{+}08, AKK08, BNS^{+}21, CGGS19, Chv77, CRS87, DDR07, Gam03, KPR^{+}13, LPSR04, Mil75]. Stabilization [BBD^{+}97, Var00]. Stabilizing [BGKP99, DIM97, DHM12, MOOY02]. Stable [BFG03, IL86, Kav14, NH90, Par77, Sub94, Gus87b, Gus88b]. Stack [HPT99, HP99b, LLS84, MO74, Rei72]. Stackelberg [Rou04]. Stacks [IJ91, Yao81]. Stage [Çam03, CG07b, DGV05]. Stamping [DS97]. Standard [AvDK^{+}07, BV14, CLP16, CV83]. Star [AAOS97, BV93, NSH^{+}08]. Star-Tree [BV93]. Start [GJ77]. Start-Times [GJ77]. Stash [KMW10]. State [ATS07, BBP00, CGG01, DS90, EWS05, GKRdW09, Iba73, San75, FG82, GR85, PS81]. State-Splitting [San75]. Stateful [FKT17]. Stateless [AK09]. Statements [Rus77a]. States [Aar20, AFB94, CHPW98, Hie10, QY21]. Static [FH11, LL95, MR01b, Pag01, SV00a, Yu22]. Statistical [BCH^{+}20, DO04, GHRU13, HNO^{+}09, IJ94, Gus88a]. Statistically [HNO^{+}09, HHRS15]. Statistics [DGIM02, GM09, YHC87]. Stealing [BFG03]. Steiner [AKR95, ABHK11, BHL18, BD97, BGA23, CHJ18, CFHM20, DM88, DH92, ES15, FR06, Fil19, FKLS10, GGK16, HLP17, HKK^{+}07, JMP11, KKN15, KLSvZ08, MR07b, Nut10, Pro88, SS05b, Sny92, Tre00]. Step [ACHP19, YI83]. Stewart [CS04]. STOC [AGK^{+}09, ACHCP21, AIM16, Bab06, CCRZWN22, CDG11, DKI18, DKSS22, EvMP^{+}12, FGKO08, IKM^{+}12, MS16, PKS^{+}13, SK18, VNA20, AKS09]. Stochastic [AIK19, AKU05, CFW93a, CFW93b, CFLS97, Dob05, DFKL20, GPRS11, GKR12, JV75, LS97, PT87b, RU98, San75, SU05, ST00, SS12b, CKS85, OH87]. Stochasticity [LM94b]. Stop [PP97]. Stoquastic [BT10]. Storage [BKK^{+}04, CKS11, CW75, CPW82, DFSS08, Lav73, Ros75a, Sch80, TS11, BC84, BW79, CL79, CKS85, LW80]. Store [SV00a]. Store-and-Forward [SV00a]. Stores [BGS72]. Storing [Mai79]. Straight [AU72, IL82b, MRK88]. Straight-Line [IL82b, MRK88]. Strassen [AL82, Sch81a]. Strategies [ADK04, ACFW98, AIS08, BKK^{+}10, CHK13, CKLW01, MR01a, Rou04, AG86, OH87]. Strategy [IKL^{+}04, MPR22]. Strategyproof [KLSvZ08]. Stream [BNL04, CCM^{+}19, CDK^{+}11, DGIM02, FKM^{+}09, GM09, JW21, PT07]. Streaming [AKL21, BFKL21, CCKM13, GG10, MMN14]. Streams [FKSV02, GM09, KLM^{+}17, LNW19]. Streets [IKL^{+}04]. Strength [Gus91]. Strengths [BBBV97]. Stretch [AN19, App13, BEGL24, DFK^{+}19, EEST08, EP09, FGG08, MRSG05, NS00, Xia13, Coh98]. Strict [BL04, FKLS10, HH72]. Strictly [HHW05, RW05]. String [Abr87, AALM90, BGP^{+}22, BG92b, Col94, CHPZ95, CH97, CH02, CGG^{+}97, FG98, GS80, GP90, GG91, GG92, Ges02, GV05, LMS98a, MS10, MM93, Riv77, TU93, AG86, AB96, Bak78b, BG90, FL83b, GO80, Liu81, Ryt80, Yao79]. String-Matching [GS80]. String-Searching [Riv77, Ryt80]. Strings [ABK^{+}06, AHK^{+}10a, AHHP00, Bak97, BLR^{+}15, KMP77, KM99, Sch98a, ST72, HC83]. Strip [BHI^{+}13, Ste97]. Strip-Packing [Ste97]. Strong [AJ22, AHLM07, BGG^{+}19, CR22, DORS08, GH92, GGM12, GIP20, GMP05, Kao95, KŠdW07, MRRS07, MRS08, RRSS09, Raz11, Saf06, She12, BM84]. Stronger [KLN15]. Strongly [BI18, CFHM20, IKP96, Kao93, KST99b, Vég16, Bab80]. Structural [BST95, CGH^{+}88, DGL23, GKMP09, JL77]. Structure [AAK^{+}12, AHS92, AS86, AR02, BH18, BV93, BW99, BDV21, BG21, BHT98, Cha92a, CKST99, DV00, Eps79, FV98, FW98, Gan95, GM15, HHL18, LWY20, LKS97, LL78, LSV07, Mac99, MM98, SMSR75, TV99, BT80c, Dub90, Gus88b, Ili89a, Ili89b, RC79, dG83]. Structured [ACBG^{+}17, BGP^{+}23, KZ20]. Structures [AFG^{+}14, And99, ADG08, BHI18, Bit79, CGdW13, Che12, Dur10, DGGW20, EFN18, Fre97, GRSS98, Hen00, Hos75, JT73, KV86, KMV91, KCH82, KS20, LP13, LRS01, PVV98, TV91, Vio12a, Wil85a, Cha88, Fre84, Fre85, FS89, Fre81]. Structuring [KMR88a]. Stubborn [CPPW12]. Study [BK07, FV98, Vad06, Coh83, JS89a]. Sub [CLLS96, MR08]. Sub-Bus [CLLS96]. Sub-Constant [MR08]. Subadditive [Fei09]. subclasses [Huỳ86a]. Subcubic [BGSW19]. Subdeterminant [ESV20]. Subdivision [EGS86, LP77, Pre79, PT89]. Subdivisions [HST05, Kir83, Mit99]. Subexponential [CG07a, Efr12, FV13, FLM^{+}22, Gär95, JPZ08, KvM02, Kup05, MPP22]. Subexponential-Time [Kup05]. Subgraph [BCR18, CKS02, Cha06, CVV03, CN85, CFHM20, CGM^{+}22, FL12, GG12, Kho06, KSS93, KR23, LRR17, CHT93]. Subgraphs [AFN07, AK23, Bar18, CT00, CV14, DLR95, FR24, FTV15, GM15, HKRT95, KN05, WW13]. Subgroup [DISW13, HRTS03, Kup05, MvV92]. Subject [CCPV11, DP20, LY89]. Sublinear [ABP18, BBRS98, BCR16, BPP23, CRT05, CLM05, CEF^{+}05, CS09, CS24, ELRS17, ERS20, FG98, GKP98, HPK14, KPS13, KMR87, Tom82]. sublinearity [Sch88]. Sublinearly [EGS10]. Sublogarithmic [CP09, Gef91, GMP98, LR96, Gef93, RR89]. Submodular [BH18, BFNS15, BFG23, CCPV11, CG10, CVZ14, Dob21, FMV11, FV16, FZ18, FW14, Iwa03, KMZ18, PR99, SF11, Von13]. Submodularity [MR10, PRR03]. Subpolynomial [Gro03]. Subpowers [IMM^{+}10]. Subquadratic [AAKS14, BI18, CL94]. subregion [Kun93]. subregion-tree [Kun93]. subroutines [RD81, RS83]. Subsequence [GG10, HSSS22, HL87]. Subsequences [JL95a, Fus88]. Subset [BGNV18, CDK^{+}11, CH03c, DV00, ENZ00, GM91a, MPP22, Fri86b, FK89]. Subset-Sum [GM91a]. Subsets [GAOPL^{+}21, KY09]. Subspace [BSK12, LWW21]. Substitutability [CGV11]. Substring [MS10, Mar08]. Subtree [AK97b, CFCH^{+}00, Rey77, VR89]. Subtrees [Dev02, Rus81]. Success [HZ20]. Successive [BCM^{+}15]. Succinct [BCG^{+}18, FK89, MR01b, Pra75a, Vio12a, Yu22]. Succinctness [SS77a, Har80, RI89]. Suffice [AM06]. Suffix [CH03b, DSR92, Gia95, GV05, KLP96, MM93, Szp93]. Suffix-Prefix-Matching [KLP96]. Sum [AGKS15, BGNV18, BBFT23, BHK^{+}19, Che05, DDG^{+}17, FS92, GM91a, GKK^{+}12, IS90, JT86, Fri86b, FK89]. Sum-of-Digits [Che05]. Sum-of-Squares [BHK^{+}19]. Summation [KW00]. Sums [AF23, BMS86, Bjö14, CDK^{+}11, Dev02, DLWZ14, Fei06, HF98, HR03, Yao85a, Rom82]. Sums-of-Squares [AF23]. Super [GHH^{+}17]. Super-Polylogarithmic [GHH^{+}17]. Superconcentrators [Pip77]. Supercritical [BN20]. Superexponential [LMS18]. Superlinear [Amb16, BBI16]. Superlogarithmic [BH03]. Superpolylogarithmic [Gab07]. Superpolynomial [AM05, BBI16, Wil13]. Supersequences [JL95a]. Superstring [Swe99]. Superstrings [TY97]. Supertree [JLL12]. Supertrees [FBGSV15]. Support [RRSS09]. Supported [AH11, DFM23]. Suppressed [EW12]. Sure [Sch87]. Surface [AS98, Bie90, CEFN23, CDRR07, Mou90]. Surfaces [AHPY13, HO02, dVE10]. Survivable [GKR12, LNSS09, LS13]. Swapping [AGL99]. Swarm [DFK^{+}19]. Sweep [BP00]. Sweeping [PS91]. Switches [AH08, EW12, KLM^{+}04]. Switching [BMIM07, FY96, KB76, Ngo05, SBI04, TU81, TS81, Yao78]. Symbolic [GL80, HS75, RT82, Wir79, Hal75]. Symmetric [ANSS21, Aza92, BWY15a, BG21, GL80, Gro03, GLV13, MRS08, Val11, Vio07, SL80, Zwi91]. Symmetries [FW76, GJK^{+}22, Ros72]. Symmetrization [BBD^{+}97]. Symmetry [CIRR16, DISW13, MOOY02, Von13]. Symposium [AGK^{+}09, AKS09, ACHCP21, AIM16, Bab06, BBS13, BSO14, BCAKT23, CCRZWN22, CDG11, CR17, DKI18, DMK19, DKSS22, EvMP^{+}12, FGKO08, HIMW16, IKM^{+}12, KRS20, LMOT14, MS16, MV18, PKS^{+}13, Rou15, SK18, VNA20]. Synchronization [AAT97, AHR96, Bar84, GK05, GK12b, LS10, Ord89]. Synchronizer [PU89a]. Synchronizing [HZ77]. Synchronous [AG91, FM97, HS79, JKLAR90, KK02, Rei84]. Syntactic [KMSV98]. Syntactically [RSY08]. Syntax [Bak78a]. Synthesis [Liu72, RS85a, Gus83, Mor80]. Synthesizing [Ber76]. System [AW76, BEG^{+}02, Coo78, CPW82, Gav72a, HT84b, KS77, Kon75, Lav73, LMM99, YH90, CBH84, Coo81, CRS87, Hal75, KMST87, LS84, MM83, SC79a]. Systems [AJ22, AKK08, BNL03, BPS07, BKRS00, BPR00, BEGJ00, BJSW20, BCLL21, CKM^{+}14, CC07, Cha15, CKS17, CP05, CSW98, DTA94, EKL10, Ezr16, FM03, FRS03, Gat92, GS73, Ger76, GK96, Gol98, GS12, GKM83, Hen76, Hie10, Jaf80, JKLAR90, Kim97, KR78, KS95, KDA00, KMP14, KZ20, Lad77, LL95, LL00a, MRW00, MP05, MVW04, MPR03, MPR22, MRSV21, NW98, Nat95, PR93, RS03b, Rus77b, AM80, AS80, BV84, Bin84, CC82, DMT88, ERV81, FN80, Gab88, GMR89, GI80, HCAL89, Ili89b, KN85, Kun81, Ott86, Oya93, RS85b, Ren89, RSL84, Zel88]. systolic [IP87, MR84]. Szelepcsényi [Gef93]. Szilard [Iga77].

Table

[AHM^{+}08, BK88, OW91, Ukk83, Yes83]. Tables [AM85, BHV99, CDG^{+}06, CDR10, DO04, GM79, Kao97]. Tabulated [Kao97]. Tabulation [TZ12]. Tabulation-Based [TZ12]. Taking [HP00b]. Tally [Gef93, Rub91, TW89, BHL95]. Tampered [CGL20]. Tangent [BDD^{+}07]. Tape [Bis78, HW87, Iga77, LL90, FG82, HM81, IM83, MS87a]. tapes [ĎG84]. Targeted [HU22]. Tarry [GGG18]. Task [BKRS00, BCLL21, FM03, Jaf80, JP05, KS77, BD78b]. Tasks [AR02, BNMSS98, CIRR16, CN99, Dob84, GK99, GL95, HL05b, MRT07, PY79, WC92, WL95, GJST81, NHL82]. Taxonomy [ABF^{+}99]. Technique [AU73, AK97a, CV88, Dye86, GW95, KS03b, Mac97, Ram94, Vis91, ACG89, Bak78b, FN80, Pet83]. Techniques [BDP02, BKV11, CMNS01, EGK^{+}14, HR03, KS96b, MS87b, RR04, SJ05, SW76, Ver97]. Telephone [EK05]. Template [FMUY83]. Temporal [CRRZ24, HLL95, IISI05, TT01]. Temporary [AZ04]. Tensor [AL10, BC87, GGR11, HRZW20, MS08a, PRS97, SZ08]. Tensors [AL83, Bsh90, Rom82]. Term [DKS88, MW88, Oya93]. term-rewriting [Oya93]. Terminal [AK11, BKS92, Boo72b, GJ19, Kar99a, Shi79b, SW85, SWCP96]. Terminals [CFHM20, CAdVdM21, FR06, IY23]. Terminating [GKV06]. Termination [Ges02, GMP23, GKM83, Lie76, CR86]. Terms [GGM11, Che86, PCHM85, SSC85].

Terrain [AdBS19, BMKM07, BRS97, KK11]. Terrains [AS96]. Test [BSHR05, Bie90, GM15, GNS20, KLX10, LPPS17, MR08, SS77b, SS78, Tuc80, AL82, Iwa89, RD81, RS83, TY84, TY85, Pig91]. Testability [CVY19]. Testable [AE15, AS08a, AS08b, AFNS09, BSGK^{+}10, BWY15a, GHJY91, KM94b, Mei09, NS13]. Tester [CS16, KS11, KSS20]. Testers [BSGK^{+}10, CS05, DR06, IKW12]. Testing [AFN07, ABC^{+}13, BB21, BDMT98, Boo78, BFNR08, CRS15, CFGM16, CSS09, DGL23, DDG^{+}17, DGV05, DT96, DRTV18, DP94b, DXSC94, DS07, EGH07, EKS00, ET75, Fis05, FN07, FM08, FPS16, GT01, GK97, GR11a, GR11b, GOS^{+}11, GNW23, HK07, HSS13, HL14, HPZZ21, IQ19, JR13, KMSV13, KKR04, KR06, KMS18, KS13, KT88, LS19, MORS10, New02, PRR03, RS96, SS12a, SW06b, ST94, TZ22, Val11, VV17, vDMMS07, FS86a, JS82, KS91]. Tests [HH91, Leh82, vzG91, Pla80]. Text [GV05, HSS09]. th [Fri86a, JM78, MTZC81, YY82]. Their [AL18, BC91, CHPJ20, CGL20, Che09, DN07, EIM01, FW07, GVY96, HPM06, IQ19, LC89, Lic82, LMPP15, PW11, PM03, RS03b, RRW01, Sch98a, SOAD09, VW85a, Web93, BG09, RI89, dG83]. Theorem [Ald75, ABI97, AR98b, BBP00, BKRS00, Bra08, CCL13, CM13b, CLD73, CP19a, CHZ07, DGL23, DR06, Dob05, ESS93, Eps79, FKKN11, Fri90a, GR02, GS00, GM15, Hai13, Iba74, KLST01, LSS^{+}21, LT80, Ram84, Raz98, Wil73, Car79, MY91, Pet83, RC79, SSC85, BII^{+}17, Bar18]. Theorems [Bra75, Dur10, GVY96, Hal08, IJKW10, KMS18, KŠdW07, MW93, She12, Sky76, TW14, BT83a, Gef93]. Theoretic [HZ77, Lin84, LMPP15, WY07, AKPW95, Gus88a, SP82, TC84]. Theoretical [Cyp93b, KY10, MS76a, CK80]. Theoretically [HKL00, KLR10, Lu14]. Theories [BEW80b, BÉ88, EIM01]. Theory [AGK^{+}09, AKS09, ACHCP21, AC78, AIM16, Bab06, BV97, BT80a, BFKV11, BLMW11, Cha89, CCRZWN22, CDG11, CG07a, Che97, CCC^{+}13, DKI18, DFVW99, DKSS22, DG93, EvMP^{+}12, FGKO08, FV98, FR75, Gel75a, Ger76, HP00a, IKM^{+}12, KB76, LV95, MS16, MS01, MS08b, PKS^{+}13, Rév́85, Rot16, SK18, VNA20, JS82, SS86b]. There [CG01]. Theta [BKM19, BGG18, ABSD^{+}98]. Theta}-Maxima [ABSD^{+}98]. Thinking [Pra75b]. Third [EvMP^{+}12, Rou15]. Thirty [AKS09, AGK^{+}09, Bab06, FGKO08]. Thirty-Eighth [AGK^{+}09]. Thirty-Ninth [AKS09]. Thirty-Seventh [FGKO08]. Thirty-Sixth [Bab06]. Thorp [Mor08]. Thorup [PR14]. threadability [CR80]. Threatening [AGG^{+}19]. Three [AS96, AST97, BDD^{+}07, Cha00, Cha92b, Che08, Coh16, DO04, FKKN11, GK99, Gus87b, HT17, HP99a, IJ94, JKK06, KMSZ05, Kun74, LS02, LR86, LC90, LS93, MW99b, MS97, PT87a, RS92, Vid16, ZS02, Dye84, ES91, Mil88, RS94a, Vid20]. Three-Chain [LS93]. Three-Dimensional [BDD^{+}07, Cha92b, LR86, LC90, MW99b, RS92, IJ94, ES91, RS94a]. Three-Element [Che08]. Three-Party [HT17]. Three-Player [Vid16, Vid20]. Three-Processor [GK99]. Three-Source [Coh16]. Three-valued [JKK06]. three-variable [Dye84]. Three-Way [DO04]. Threshold [AM06, Aza92, BH12, DRST14, Eft16, GK98a, HHT97, HTC13, IPS97, MZ13, RS10a, RS22, RT92, She13, BS92, FSW87, Fri86a, Ord89]. Thresholds [BL21, CRRZ24]. Throughput [BNGNS01, BKPY18]. Thue [KN85, Ott86]. Tight [ACER22, ABC^{+}13, AO95, AHHP00, AFRV19, AKL21, AV87, BRS^{+}21, BHK^{+}19, BGS98, BGLL99, BFNS15, BKS15, Cha94, CFHM20, CKL98, Col94, CSW98, DP94a, EPA99, FKR95, GLM^{+}99, HHRS15, HKN21, IKKP19, JLTX20, JJLZ22, KKR04, KMMT09, KP00b, LM99, LWW21, Mao93, PR00, Ram94, VZ13, Zha09, Zhu21, dBBKB^{+}20, dBBKBK23]. Tightening [dVE10]. Tighter [BHMO22, CHPZ95, CH97, Fri84, Fri87a]. Tile [CGR12]. Time [AC82, AT77, AG91, AAKS14, AFB94, AOT89, AM96, AHLT99, AM01, AAT97, ACR^{+}10, AO12, AI77, AHK10b, ABCP98, AALR02, BI18, BPS10, BSS21, BNGNS01, BL04, BEK14, BBRS98, BE98, BGK16, BGS15, BGS18, BGK20, Bea91, BBR^{+}99, BBI16, Ben89, BH80, Bie90, Bis78, BH19, Bod96, Boo72a, BWX82, Bor77, BFadH^{+}87, BHIK97, BKM^{+}17, BRS98, BDK00, BG82, BM99, BP81, BFNS15, BGK^{+}08, BPP89, BP93b, BY98, CCK^{+}17, CKM^{+}14, CS99, CW04, CJK^{+}22, CCK^{+}20, CLL08, CCC^{+}09, CP09, CL13, CP19a, CRT05, CMNS01, CK05, CM01, CHH02, CD17, CHM96, CT90, CW98, CPPT24, CFW93a, CV88, CSSS89, CH03c, Col74, CTX10, COS99, CGG^{+}97, CGH^{+}98, CGG01, CEF^{+}05, CS09, CS24, DF18, DFM23, DHH96, DD81, DvM06, DKR96, DRT92, DS97, DGP07, DSS90, DS90, DHPW99]. Time [ET93, ELRS17, ERS20, Edm98, Elk06, EGS10, Eri00, EJS05, ESY17, EHJM00, FBL03, FG98, FN99, FS07a, FPS13, FR94, Fu95, GM91a, GKP98, GKM^{+}21, GLP75, GLPS12, GKK13, GLLON18, Gol08, GJ13, GK12b, GMY12, GL13, Gra94, GS02, GJ19, GKP22, HHKL21, HJLS14, Had75, HH81, HSSS22, HvdMV04, HKRT95, HH72, HL04a, HW93, HSS09, IKS10, Jai07, JO92, JP05, JKLS05, JTT00, JS93, KW00, KKP93, Kar99a, KU99, KM94a, KTW99, KŠdW07, Kle08, KvM02, Ko91, KR07, KS95, KDA00, KP04, KCV92, KM75, Kup05, Kur87, Kut91, LS18, LS02, LL95, LP03, LS90, Li20, LPS20, LL90, Lu14, LMR90, MV13, Mit99, Mor08, MRR15, NN17, NPSW23, NW06, OS07, Ogi95, OW91, Pag01, PP08, PSW90, PRS23, PU87, PY91, Par77]. Time [PR17, PU90, PR00, PR02, PB85, Raj96, Reg96, RT82, RLDL96, RZ16, Rub91, SS17, Saw01, Sch13, Sch98b, SLW^{+}98, Sch04b, Sei77, Sel78, Sho97, Sie04, Sit21, ST13, Spi73, SVV12, SZ76, Tod91, TO92, Tze92, Vai89b, Val02, Wal98, Wan92, Wan98, WZ21, Wat14, WROM86, WLB^{+}99, Yao94, YYI12, dBBKB^{+}20, AB96, AS80, Blo83, Blu86, BORW88, BC82, BG90, BKRU94, CHT93, CI83, CMST85, CB81, CDR86, Dye84, ES91, FG82, GJST81, GJ86, GG86, GS87, HJLS89, Hag90, Hor87, IM83, Iwa89, Kad88, Kad91, Kal85b, KK86, KF80, KR88, Ko89, KM89, LLK80, LY89, Liu81, Meg83a, MM83, MT87, Mye85, RR89, Rei83, RY87, SW85, SWCP96, Sch78, Sha79, Sim83, TY84, TY85, TV88b, TV88a, Tom82, Ukk83, Wil85b, Yes83]. Time-Adaptive [AAT97]. Time-and-Space [HSS09]. Time-Approximation [Elk06]. Time-Bounded [Ko91, LL90, SZ76]. Time-Complexity [DGP07]. Time-decaying [CTX10]. Time-Indexed [Li20]. Time-Lapse [DHPW99]. Time-Lock [LPS20]. Time-Optimal [DKR96, NN17]. Time-Randomness [PU90]. time-sharing [GJ86]. Time-Space [BE98, Bea91, BBR^{+}99, BBI16, BWX82, BFadH^{+}87, DvM06, Edm98, KŠdW07, LS90, Yao94, BC82, IM83]. Time-Stamping [DS97]. Time-Work [PR02]. Time/Space [Ben89, FN99]. Timed [GL13]. Times [Che93, CKR72, CKL98, DD81, FR94, GM01, GJ77, Tet99, BD78b, GJST81, HR87, HCAL89, PY88, SC79a, Sim83]. Timetable [EIS76]. Timing [GM01, RW03, SLC91]. Timing-Based [GM01]. Tiny [FLS04, FSS20]. TMs [IM83]. Toeplitz [Bin84, BC87, LCC90, Pan00]. Toeplitz-Like [Pan00]. Token [PU89b]. Tolerance [AL07, DPPU88, KPS13, LMS98b, MSZ11, Sch87]. Tolerant [AP06, ABO08, BCR18, BCH97, CLPR10, KP05, LM99, MMSA93, YY85, vDMMS07, GGL98]. Toll [HN02]. Tomography [Aar20]. Tools [FMUY83]. Top [BJK^{+}94, KMSV13, KS15, MGHK09, NN17, SS12a, SS05a]. Top- [NN17]. Top-Bottom [BJK^{+}94]. Top-Down [MGHK09, SS05a]. Top-Fanin [SS12a]. Topological [BAG01, Dob89, GM15, HS06, NSW11, Mil88]. Topologies [AKU05]. Topology [BNFN01, BP20, CDRR07, KOWZ23, SZ00]. Tori [CC07, ECM03]. Torus [RS08]. Torus-Based [RS08]. Total [BGT16, DHK11, FS81, GRS93, HMR98, Kao97, KTW99, LP03, Li20, MMSA93, Opa79, Sch81b]. Totally [GHSZ08, GTV21]. Tour [AST76]. Tournaments [CHZ07, SSW03]. Tours [ABD^{+}05, BY98, Jac90, TV20]. Towers [CS04]. Trace [Lem75, SL80]. Trace-Orthogonal [Lem75, SL80]. Tracing [GKW20]. Tractability [BMM21, CHM13, FG01, IMM^{+}10, KST99b, MR14, DF95]. Tractable [AR08, CLX17, CCK^{+}20, CLP^{+}19, FR06, LPPS17, VHPT09, BS86]. Trade [AKW00, BBI16, Ben89, BN20, Elk06, FN99, PU90, Vai89b, VW85b]. Trade-Off [AKW00, PU90, Elk06]. Trade-Offs [Ben89, Vai89b, BBI16, BN20, FN99, VW85b]. Tradeoff [Bea91, BBR^{+}99, BFadH^{+}87, Kav14, LS90, PU87, Yao94, BC82, IM83]. Tradeoffs [BTY94, BS09, BG82, BCE95, CCKM13, CMM10, DSY10, Edm98, Eri00, IPS97, KŠdW07, SU09]. Trading [BKRU94]. Traffic [CKV10, GJ86]. Trails [CHH02]. Traitor [GKW20]. Transaction [HT84b]. Transaction-Oriented [HT84b]. Transcendence [ASSS16]. Transdichotomous [CP09]. Transducers [Bak78a, CC83, MGHK09, Web93, CK87, Gur82]. Transfer [Whi90, JKS84]. Transfers [CH87, CGJL85, GLP75]. transfinite [RS91a]. Transform [AC09, Bra98, GKM18, CSS90, Sut91, CDMI^{+}16]. Transformation [BG95b, BCC^{+}13, GL23, LC05, SJ05, YI83].

Transformations

[EL86, JL77, Mil76, SV00a, Ye90, Mor80]. Transforming [KCH82]. Transforms [BC91, Cla89, DHR97, FA77, Pro92, PM03]. Transition [Jac72]. Transitive [AGU72, BGJ^{+}12, BB88, BB95, Boo79, DHR97, GKV06, TV91, UY91, Sch78, Tom82]. Transitive-Closure [BGJ^{+}12, UY91]. Translatability [Wey79]. Translating [AM75, FIM^{+}14]. Translation [Bak78a, FIM^{+}14, Rei72, SPH84b]. Translational [AS97]. Translations [EM03, HB75]. Transmission [KMRS18, PRS94]. Transportation [AT77, Fre93, TN95, AK88]. Transposition [Cha73]. Transshipment [BFKL21]. Transversals [EG95, EGM03, KRS10a]. Trapdoor [PW11]. Travel [DRW98, AK88]. Traveling [ACK^{+}99, AM01, ACM^{+}99, BGK16, BHZ23, CKT99, FKR95, FS07b, FSS13, GKR05, MPP22, PS77, Pap92, Sit21, SS95, TV22, CP84, HT82a, Kar79, RSL77]. Traveling-Salesman [FKR95, Kar79]. travelling [Fri87b, SRP83]. Traversal [BBR^{+}99, FFK^{+}98, Tom92, FL83a]. Treatment [FA77]. Tree [AN19, ABF^{+}99, AC11, And99, ABHK11, Bak78a, BV93, BGLL99, BGP^{+}23, Bod96, BGA23, CCK^{+}22, CRT05, Che92, CT99, CH03b, CH03c, CEF^{+}05, DMW05, EHS^{+}19, Elk06, EJ99, EM03, FT97, FL83a, FG92, Gia95, GadHW96, GGK16, GLS^{+}24, HL04a, Hos75, KW95, Kao98, KM95, KS00, KLS^{+}24, KM77, KR23, LNRS92, LAB01, LPSPP05, LST99, LL92, MGHK09, MTZC81, MMS17, MR91, MR07b, NP97, PSW90, PT87b, PR00, PS12, Pur78, Raj96, Rus81, Sei90, SI87, ST94, Szp93, Tre00, WJG00, Wil00, aBC08, ASSU81, ERV81, Kun93]. Tree-Adjoining [Raj96]. Tree-Decompositions [Bod96]. Tree-Depth [CCK^{+}22, KR23]. Tree-Like [PS12]. tree-rank [ERV81]. Tree-Structured [BGP^{+}23].

Tree-Walking [aBC08]. Tree-Width [DMW05]. Trees [AdBS10, ABN15, Aga92, AHU76, AK97b, AH89, AABV98, BYCDM92, BR79, BEGL24, BAFN99, Ben95, BDFC05, BST85, BH80, BCLR92, BLR^{+}15, BT80a, BB95, BPWY99, Bro79, BGKP99, BM02, BCLL21, CLWZ04, CT76, CH06, CLL05, CT90, CH87, CDTT95, CK04b, CKM10, CFCH^{+}00, CMSS00, Col00, CH05, CKP86, CGG01, CK01, CLY06, CS09, DP94a, DSR92, DR95c, DR95b, Dev98, Dev99, DJZC00, Dev02, DN04, DKM09, Dev12, DRT92, EEST08, ES15, EP09, EiC17, EP98, FS86b, FP86, Fre97, FGRV14, Gab77, GM78, GKP98, Gar74, GW77, GH97, GHT76, GT98, Got81, GV05, GLW82, HKK^{+}07, HF98, HLCW91, HR82, Hu74, Ita76, JS05, JMP11, KWY98, Kao98, KLST00, KR95, KRY96, KK85, KPS94, KR02, Kun74, KC94, KM93, LW98b, Lar87, LP94, LP95, LM94a, Li86, LSS13, Lou83]. Trees [MLZ00, Mai79, McC85, Meh77, MT86, MPRS79, MYSH77, MR01b, NSV94, NR73, Nis91, NRS99, OS07, OL74, OPR^{+}84, PRS23, PR99, Pro92, RH77, Rus78, STU97, SCH81c, SY12, SP75, Tro78, Vai88, Vai87, WG05, WC90, WZ21, Wil00, WROM86, WLB^{+}99, Yao91, AMS84, AKR95, BH81b, CW83, CL91, CMST85, CB81, DM88, EMC85, FLST86, Fre85, HY88, KS93, KIM81, KMZ87, LW86, MR80, Meh82, Mil88, Pro88, RS78, Shi79a, Vai90, Wil89, Yao82, ZR79, ZS89]. Treewidth [BST23, Bod96, BH98, BDD^{+}16, BT01, CR05, FFKP18, FKTV08, FLM^{+}22, LPPS17, OR21]. trellis [IKM85]. Trevisan [DPVR12]. Trials [Hol73, Hol74]. Triangle [CL09, GRS16, LM12, MSS07]. Triangle-Dense [GRS16]. Triangles [AWY18, BFvRV15, ELRS17, ES05, ES22, MPS^{+}94, PT02]. Triangular [He93, HL04b]. Triangularization [HM91]. Triangulated [BX91, EHSS99, HO02, BX92]. Triangulating [LM12, TV88b, TV88a].

Triangulation

[CW98, ET93, Ram94, Xia13, YY14]. Triangulations [BRLW19, BM04b, BFvRV15, FTV15, LM12, ZH05, vKLM10]. Triconnected [FRT93, HT73]. tridiagonal [Zel88]. Trie [LSH05]. Trie-Based [LSH05]. Tries [NS08, PHNS09, Sch04a]. Triple [COS99]. Triple-free [COS99]. Triplets [CK01, JNS06]. Trivial [KST93]. Truly [BGSW19]. Truncated [MS04, FHK^{+}88]. Truth [AHM^{+}08, BK88, OW91, Ukk83, Yes83]. Truth-Table [BK88, OW91, Ukk83, Yes83]. Truthful [AKSW22, BSS14, CK13, DSS18, DDDR11, Dob21]. TSP [ABHK11, BCK^{+}07, CG12, Kle08, MSVW16, Mit99, TVZ22, Tre00, dBBKBK23]. Tucker [VP03]. tuple [BV84]. Turing [Bis78, CG97, Fu95, Gil77, HR97, KvLP88, Ko91, KLL90, LL90, MS87a, Pri79, Rub91]. Turing-Complete [HR97]. Turn [ABD^{+}05]. Turning [FSS20]. Turtle [Mar99]. Tutte [GJ14, Ver05]. Twice [BSS12, GSV99, SV95]. Twice-Ramanujan [BSS12]. Two [AM06, AGM^{+}11, ABF94, AB98, ABT21, ANSS21, BCR80, Bak86, BKS92, BADTS22, BSRZ15, BH19, BDFP86, BCD^{+}89a, BCD^{+}89b, CK95, CM94, Coh21, CGH^{+}98, CGG01, DDMM05, DGV05, ĎG84, DS07, DS90, FK00, FKS93, Gab77, GP96, GJ77, GLM^{+}99, HNO^{+}20, Har99, HM87, HN94, HL72, IJTW95, IISI05, ISK^{+}12, KMR87, KU99, Ken76, KMV92, KMPHT14, KMP^{+}20, KT15, LS77, LPS20, LMR90, Meh18, MMP08, MO74, PST00, SV08, Set76, She13, Shp09, TN82, Tom82, Ukk83, AS80, BS83, BT83a, CF86, CGJT80, Coh83, CK87, Dye84, EW86, Gab88, Gol81, Gur82, HM81, HR87, Huỳ86a, MS87a, NHL82, SW79, VV89]. Two- [KU99, Dye84]. Two-Dimensional [ABF94, AB98, ANSS21, CK95, CGH^{+}98, GP96, Har99, BS83, CGJT80, EW86, Gol81]. Two-Level [MO74]. Two-Local [KMP^{+}20].

Two-Party [HNO^{+}20, SV08]. Two-Player [Meh18]. Two-Processor [GJ77, VV89]. Two-Prover [FK00]. Two-Round [LPS20]. Two-Sided [AGM^{+}11]. Two-Source [BADTS22, BSRZ15, Coh21]. Two-Stage [DGV05]. Two-State [CGG01]. Two-Terminal [BKS92]. Two-Variable [KMPHT14, KT15, PST00]. Two-Way [DDMM05, DS90, IJTW95, CK87, Gur82, HM81, MS87a]. Type [Hal08, Hav04, JV08, KC96, BT83a, Don79, KMS18, RI89]. Type- [KC96]. Typed [Jaf80]. Types [DCV90, Jay98, LM82, Sco76, Dob89, NHL82]. Typical [BV06].

UET [Gab88]. Ullman [APS74]. Ultimate [JS89a, KS86]. Ultrafast [CLP20]. Ultrametrics [ABN15]. Unambiguity [ST09b]. Unambiguous [HR97, RA00, SS77a, GG86, SH85]. Unanimity [DKK^{+}19a]. Unapproximable [Zuc96]. Unary [FP86, GHK96, KT88, Iba78]. unate [Zwi91]. Unbalanced [AK23]. Unbiased [CG88]. Unbounded [AFG^{+}14, Bei90, Fei06, RS85b, RS91b, RS91a]. Unbounded-Length [AFG^{+}14]. Uncapacitated [BA10, CS03a]. Uncertainty [CS19, FMP^{+}03, MMS17]. Uncolorability [CM87a]. Unconditional [Elk06, Vad06]. Unconstrained [BFNS15]. Uncoordinated [AGM^{+}11]. Uncrowded [BKL99]. Undecidability [AKK^{+}09, Mil72]. Undecidable [BJKP05, GK99, Mar99, CV85]. Understanding [FW90]. Undirected [AZ07, BK10b, BCCK19, BBR^{+}99, Bjö14, Chu16, Edm98, GM78, HS85a, KR95, Kle08, MRSV21, PR05, RZ12, Seb97, STU97, Tri08, BKRU94, HJ85, Rei83, Shi79a]. Unequal [DD81, GY96, FK88]. Unfair [BCLL21, FM03]. Unfamiliar [BRS97]. Unfolding [AAOS97]. Unidirectional [AG94a]. Unification [LSSV08]. Unified [CDW21, CPT96, FA77]. Uniform [AG94b, BTW00, BJ99, CK09, CS80, Dob84, FL83c, GW17, GJMM00, GK18, GMP23, GIS77, GKM83, HC83, IM20, IJK09, IJKW10, JS05, KW93, KS17a, LS98, MSW81, PP08, SU09, Yun79, Fri87a, Fus88, Gab88, HS88, SC79a]. Uniformity [MTV10, ST09a]. Uniformly [GKP04, GM92]. Unifying [GPV94, Păt11]. Unimodal [GR93]. Unimodular [GTV21, KM97b]. Unimodularity [DHK11]. unindexed [Wil85b]. uninteresting [SPH84b]. Union [AIGT94, AST97, AdBES14, BS93, Efr05, ES05, Luc90, NSV94, PT02, Blu86, MNA88, WT89]. union-split-find [MNA88]. Uniprocessor [KS95]. Unique [Bea91, BE17, CRS95, DFHS08, GL17, Hen92, Her14, KRT10, LR88a, ST94]. Unique-SAT [Her14]. Uniquely [AO95, HNOS96]. Uniqueness [AMT07, CLV23, Eft16, EHS^{+}19, DMT88]. Unit [CLL08, CDO^{+}22, FR94, GZ05, KPST94, MMR95, GJST81, Sim83, SW89, SRP83]. Unit-Demand [CDO^{+}22]. Unit-Disk [GZ05]. unit-length [SW89]. Unit-Speed [CLL08]. Unit-Time [FR94, GJST81, Sim83]. Unitary [DS02, FA77]. univariate [Kal85b]. Universal [ÀBS04, BNBK^{+}89, BG09, BMIM07, CF22, CH85, Dev98, EMSV12, ELMS^{+}12, GMWW17, MS17, Sie04, Tom92, vDMMS07, Pri79]. Universalization [ADK04]. Universally [BN14, GRS12]. Unknown [AH00, AW21, De 10, SW99]. Unlabelled [Wor87]. Unless [AR08, BI18, KvM02]. Unlimited [AIGT94]. Unordered [KM95]. Unoriented [ABSD^{+}98]. Unpredictable [BBS86b]. Unranking [GLW82, Li86]. Unrecognizable [TT13]. Unrelated [BSS21, Sve12]. Unreliable [ELMS^{+}12, ISK^{+}12]. Unrestricted [BTY94, Goe93]. Unrooted [KLST00]. unsatisfiability [Pla80]. Unsatisfiable [FGK05]. Unsolvability [HJ90, Iba78]. Unsplittable [AAE13, BSW14, KS02]. Unsupervised [NSW11]. unsymmetrical [Sch79]. untangle [MY91]. Untangling [LDRT24]. Until [TT01]. Untrusted [CKS11]. untyped [JK83]. Unweighted [EP09, GS21a]. Update [Alb98, BGS15, BGS18, FG98, RZ16]. Updates [TA99]. Updating [SP75, Fre85]. Upper [ABS07, BC91, CEEP16, CFKX07, Chi76, CH97, CDR86, DP94a, DKM^{+}94, GG92, Gaz91a, GadHW96, GM18, PY91]. Upward [BDMT98, GT01, HL96]. Use [ACHAH16, SW21, Cha88]. user [BD81]. Uses [Lic82]. Using [AN19, AST99, BIW06, BATS11, BOC92, BJ99, BFvMT00, BJK05, CFK^{+}21, CCH^{+}16, CIL94, DKSS13a, DLM22, DH13, ES81, Gan95, GJMM00, GKL13, GPST92, GKS93, GSY95, HRTS03, HL04a, HVV06, HR92, Koz21, KM93, LM08a, LR21, Mei13, SJ05, Sho92, Tet99, Yac98, CRS15, Fil19, HS88, KLL17, OH87, VSBR83]. Utilitarian [BKV11, GKLV14]. Utilities [GMSV15, Jai07]. Utility [Fei09, GRS12]. Utility-maximizing [GRS12]. Utilization [GS73, BH81b].

Vacillation [Cas99]. Valiant [ABW18]. Validity [JT86]. Value [Alb09, Cus98, FK03, GL23, RY96, Wal79, Hen80, IR96]. Valued [CRZ22, HKP14, KTZ15, KKR17, TZ17, Web93, CK87, HITK79, JKK06, RS83]. Values [CCL13, FHOS97]. Vanishing [CGW16]. Vapnik [DSS18, DHV22]. variability [RS85b]. Variable [EL08, FL86, GLP75, IL82b, KKMO07, KMPHT14, KT15, LS97, NS22, PST00, SvSE03, Dye84, Kni88, Mur87]. Variable-Sized [EL08, SvSE03, Mur87]. Variables [ART09, CM94, Fei06, Fra91, HN94, KST93, LMR90, ST09a, AS80, Fri86a, FHK^{+}88, JK83]. Variance [CDK^{+}11, DR95b, Fei06, Kou82]. Variance-Optimal [CDK^{+}11]. Variants [NH90, LVW84]. Variational [MMR95]. Varietal [Koz09]. Varieties [BS00, BT83b, BÉ88, Gal84b, MS08b, dG83]. Various [Sky76, CK80]. Varying [HN02]. Vazirani [BEG17]. VC [KW93]. Vector [BDL14, BEW80b, DGY11, FLS04, IKKP19, KPV14, Wir79, RS83, Sys81]. vector-valued [RS83]. Vectors [Hir80, MV75]. Vegas [Yu22]. Vehicle [CKR01]. Verifiable [CCM^{+}19]. Verification [Coo78, DGL23, DRT92, Ger76, Mah22, RS00, SHK^{+}12, SPH84a, Coo81, MR84]. Verifying [AFL08, CC97, Kre90, dOS94]. Version [Bar18, BBP00, BGS18, FKKN11, GR93, HNR08]. Versions [Zuc96, Gef93]. Versus [ALN^{+}12, BG94, BDV21, BKK^{+}04, Gas87, Goe93, LL95, FMS05, FN07, GGM12, GP04, HHH05, KMSV98, Ros18, SU09, Wil16]. Vertex [BX91, BBFT23, BYGNR98, BCG^{+}21, BHI18, CVV03, CL09, DV00, DGKR05, DP20, EGK^{+}14, EGIS98, ENZ00, FL12, FHL08, Gal80, GMPT10, GKPS08, GTV21, Hal02, KKL04, KLL17, Moi13, Oht76, RTL76, BX92, CKT93, Hoc82, MT84, RLWW97]. vertex-coloring [MT84]. Vertex-Connectivity [KKL04]. vertex-disjoint [RLWW97]. Vertex-Weighted [CL09]. Vertical [AES99, PS91, dBG08]. vertices [BM88]. Very [BL93, Ebe89, Gus90, HS90, SZ99]. Via [AJ22, BG95a, DLWZ14, Mur95, AFGN22, AFGN23, AN06, ANSS21, Bar18, BEGL24, BCGL17, BHMO22, BSRZ15, BGP^{+}22, BGG^{+}19, BG96, BHK09, BRLW19, BGG22, BGSW19, BCLL21, BNS18, CCWZ18, CLP20, CVZ14, CK00, CLV23, CT00, CV14, CGP^{+}20, CHKZ03, DSS18, DL91, ESV20, FX15, FW14, FTV15, FLM^{+}22, FGRV14, GP18, GKM18, GQ17, GLS^{+}24, GTV21, GHH^{+}17, HLP17, HT21, HV20, IM20, JV08, KLL^{+}15, KN05, KLS^{+}24, LS22, LS16, Li20, LSS^{+}21, MNRS11, MR12, Mur90, Nat99, NPSW23, PS97, Pro92, RSW06, RS10c, Rot16, Sto79, Wil18]. View [KS20, NSW11]. Views [KMSV98, SI87]. Violations [Cha05]. Virtual [AIK^{+}20, EGOS07, SV00b]. Virtually [KK02]. Virtue [AMRU22]. Visibility [GM91b, HM95, KM00, PS91]. visits [CP84]. VLSI [AKW00, BG82, Cyp93b, EGS03, Ham83, HS85a]. Voltage [LY05]. Volume [CV18, DF88, ZS02]. Volumes [DGH98, LS22]. Voronoi [AdBMS98, BRLW19, CHLX17, CW98, Fil19, GKM^{+}21, KS03a, MV13, IIM85, LW80, LD81]. Voting [CK14]. vs [BvMR^{+}00, DF92, GPW18, MS01].

w [Ngo05]. Wait [AR02, CHJT04, CIL94, Hav04, LH00, SZ00]. Wait-Free [AR02, CHJT04, CIL94, Hav04, LH00, SZ00]. Waiting [CFW93a, CKL98, MM83]. Waiting-Time [CFW93a]. Wake [CGK07]. Wake-Up [CGK07]. Wakeup [FMRT96]. Walk [ABS07, Amb07, FMM11, HP00b, MNRS11]. Walking [BBR^{+}99, IKL^{+}04, LM15, aBC08]. Walks [DF18, Gil98, GH73, GH74, KSS20, LV22b, MS04, Ver09]. Walksat [COF14]. Walrasian [FGL16]. WANs [KK02]. Want [BDP23]. Wardrop [FRV10]. WARM [ML00]. WARM-UP [ML00]. Watermarking [CHN^{+}18]. Watson [Dev12]. Way [App13, BYJK08, BCPR16, BP02, CPS16, DDMM05, DO04, DS90, Fab77, GKK^{+}09, Got81, HNO^{+}09, HRV13, HO14, IJ87, IJTW95, KYN07, Kla07b, KMN^{+}22, CK87, Gur82, HILL99, HM81, MS87a]. WDM [Ngo05, RW05]. Weak [ACG^{+}16, ACRT99, ACH10, BHMO22, CHS12, GGM12, Koz21, Pac74, Spr02, SZ99, CG88]. Weaker [BOBP^{+}04]. Weakest [GHKT12]. Weakly [EHSS99, KMR22, Lut95, LZ00]. Weaknesses [BBBV97]. Web [CKL^{+}13]. Weight [AABV98, BX91, CRT05, CT90, CPPT24, CEF^{+}05, CS09, FHL08, GKP98, HC99, HL87, KLST01, KRY96, LPSPP05, PR00, RLDL96, YY14, BX92, Hsu85, KMZ87]. Weighted [AFB96, BX91, BPS10, BSS21, BGK20, Ber16, BGW00, CCL16, Cha10, CKT00, CD03, CGGS19, CL09, DGJ09, Gab77, GS21a, GS21b, HP98, IK22b, KR95, Li20, MM02, Meh77, MR07b, MRR15, PR05, Rod10, Sch98a, SLW^{+}98, Sch04b, Vai87, Vai90, WW13, Zim98, BX92, BHL18, GMG86, HLP17, KK86, SR83, WE80]. Weights [Kle08, LW18, Nut10, Lue81]. Weil [BF03b]. Weintraub [BéT89]. Welfare [CG10, CG18, Fei09, KMZ18]. Well [ADG08, DFM23, GZ05, MMN14, PSZ17]. Well-Behaved [ADG08]. Well-Clustered [PSZ17]. Well-Parenthesized [MMN14]. Well-Separated [GZ05]. Weyl [ES81]. Where [BP98]. Whether [Eve75, FT83]. Which [Str74, Tom82]. While [AM75]. While-Schemas [AM75]. White [HP10]. Whitney [MY91]. Whole [AP07, ACGS88]. whose [Fus88]. Width [BK10a, BN20, BDFP86, CR05, DMW05, FT06, FGLS10, FGLS14, GP04, Hli05, New02, Nor09, VW85b]. Williamson [Kar99b]. Windows [BNL03, BO10, DGIM02, GKP22]. Winners [BV06]. Winnow [Hit07, Ser02]. Wire [Szk99]. Wireless [HHMW17, HT21]. Wiring [SW99]. Within [Hir23, BG93b, GSV99, HHH98, SV95]. Without [ACIM99, AALR02, DLL^{+}03, DKK^{+}19b, AK23, AHKL07, CCC^{+}09, CDR86, EKS00, GM20, HY87, Mul99, PY91, VL00, WE80]. Witness [GP96, MW20]. Witnesses [DLM22]. Wobegon [BKK^{+}10]. Woeginger [CCG^{+}97]. Wolfe [DHR20, KY15]. Wood [SWCP96]. Word [BMPT97, Loh06, NS22, RS91c, Wan99]. Words [GY96, Loh06]. Work [BFG03, CBGWW13, DHW98, Fin20, GRS05, LS75, Mal05, MPS92, PP95, PR02, Rei99, Sit14, MS87a]. Work-Competitive [GRS05]. Work-Efficient [Fin20]. Work-Optimal [Mal05, MPS92]. Work-Preserving [PP95]. Work-Stealing [BFG03]. working [HT82b]. Works [PSZ17]. World [EGH07, LR96]. Worlds [IKK^{+}11]. Worst [AV09, BCGL17, BT06, CW75, CMS97, DMV04, GW93a, Hir23, Ili89a, Ili89b, JDU^{+}74, KK86, KRT17, KV98, LS77, LL00b, MR07a, Pap81, Riv77, Sta80, SS89b, Blu86, CCG^{+}97, Cha84, LP83b, Mao93, Ren89]. Worst-Case [AV09, BCGL17, BT06, CW75, CMS97, DMV04, Hir23, KRT17, MR07a, Riv77, Ili89a, Ili89b, JDU^{+}74, Pap81, SS89b, Blu86, Cha84, Mao93, Ren89]. Wrap [WG05]. Wrap-Around [WG05]. Wright [ST00]. Write [AW97, DKR96, GMR98, HS02, Mal05, FRW88, KR88]. Write-All [AW97, Mal05]. Writes [PY91, CDR86]. Writing [AW76, HSV05].

XML [AFL08, CKM10, GMN09, MNS10]. XOR [Vid20, HHL18, Vid16]. XOS [FV16].

Yao [CPV22]. Yields [CAKM19, FRS19, Hit07].

Zag [BATS11]. Zaps [DN07]. Zero [BDMP91, BCH^{+}20, BJSW20, BG22, CKPR02, DDPY08, FLS99, Fri90a, GK96, HNO^{+}09, IKOS09, IS90, Jac90, KLL^{+}15, PTW11, Vad06, Wat09, XLXJ09]. Zero-Communication [KLL^{+}15]. Zero-Knowledge [BDMP91, BJSW20, BG22, CKPR02, DDPY08, GK96, HNO^{+}09, IKOS09, PTW11, Wat09]. Zero-Recombinant [XLXJ09]. Zero-Sum [IS90]. Zeros [BG92a, BP98, Fri90b, Ris85, Loo83, Ren89]. Zig [BATS11]. Zig-Zag [BATS11]. Ziv [FNV13, KS00, KM99, LST99]. Zone [AMT07, ESS93]. Zwick [PR14].

References

Alon:1988:ACD

[AA88] N. Alon and Y. Azar. The average complexity of deterministic and randomized parallel comparison-sorting algorithms. SIAM Journal on Computing, 17(6):1178–1192, December 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:1989:FAM

[AA89] N. Alon and Y. Azar. Finding an approximate maximum. SIAM Journal on Computing, 18(2):258–267, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aaronson:2018:FPO

[AA18] Scott Aaronson and Andris Ambainis. Forrelation: a problem that optimally separates quantum from classical computing. SIAM Journal on Computing, 47(3):982– 1038, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:2009:OSC

[AAA^{+}09] Noga Alon, Baruch Awerbuch, Yossi Azar, Niv Buchbinder, and Joseph (Seffi) Naor. The online set cover problem. SIAM Journal on Computing, 39(2):361–370, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alamdari:2017:HMP

[AABC^{+}17] Soroush Alamdari, Patrizio Angelini, Fidel Barrera-Cruz, Timothy M. Chan, Giordano Da Lozzo, Giuseppe Di Battista, Fabrizio Frati, Penny Haxell, Anna Lubiw, Maurizio Patrignani, Vincenzo Roselli, Sahil Singla, and Bryan T. Wilkinson. How to morph planar graph drawings. SIAM Journal on Computing, 46(2): 824–852, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Awerbuch:1998:NAG

[AABV98] Baruch Awerbuch, Yossi Azar, Avrim Blum, and Santosh Vempala. New approximation guarantees for minimumweight k-trees and prizecollecting salesmen. SIAM Journal on Computing, 28 (1):254–262, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28826.

Awerbuch:2013:PRU

[AAE13] Baruch Awerbuch, Yossi Azar, and Amir Epstein. The price of routing unsplittable flow. SIAM Journal on Computing, 42(1):160–177, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alistarh:2023:WEB

[AAE^{+}23] Dan Alistarh, James Aspnes, Faith Ellen, Rati Gelashvili, and Leqi Zhu. Why extensionbased proofs fail. SIAM Journal on Computing, 52 (4):913–944, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1375851.

Ailon:2012:IAA

[AAELvZ12] Nir Ailon, Noa Avigdor-Elgrabli, Edo Liberty, and Anke van Zuylen. Improved approximation algorithms for bipartite correlation clustering. SIAM Journal on Computing, 41(5):1110–1121, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:2021:EAG

[AAEZ21] Pankaj K. Agarwal, Boris Aronov, Esther Ezra, and Joshua Zahl. Efficient algorithm for generalized polynomial partitioning and its applications. SIAM Journal on Computing, 50(2):760–787, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abrahamson:1989:BCR

[AAG^{+}89] Karl Abrahamson, Andrew Adler, Rachel Gelbart, Lisa Higham, and David Kirkpatrick. The bit complexity of randomized leader election on a ring. SIAM Journal on Computing, 18(1):12–29, February 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aggarwal:1990:PDF

[AAK90] Alok Aggarwal, Richard J. Anderson, and Ming-Yang Kao. Parallel depth-first search in general directed graphs. SIAM Journal on Computing, 19(2): 397–409, April 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abiteboul:2006:CLS

[AAK^{+}06] Serge Abiteboul, Stephen Alstrup, Haim Kaplan, Tova Milo, and Theis Rauhe. Compact labeling scheme for ancestor queries. SIAM Journal on Computing, 35(6):1295– 1309, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:2012:ODD

[AAK^{+}12] Pankaj K. Agarwal, Lars Arge, Haim Kaplan, Eyal Molad, Robert E. Tarjan, and Ke Yi. An optimal dynamic data structure for stabbingsemigroup queries. SIAM Journal on Computing, 41(1):

104–127, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v41/i1/p104_s1.

Agarwal:2014:CDF

[AAKS14] Pankaj K. Agarwal, Rinat Ben Avraham, Haim Kaplan, and Micha Sharir. Computing the discrete Fréchet distance in subquadratic time. SIAM Journal on Computing, 43(2): 429–449, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Apostolico:1990:EPA

[AALM90] Alberto Apostolico, Mikhail J. Atallah, Lawrence L. Larmore, and Scott McFaddin. Efficient parallel algorithms for string editing and related problems. SIAM Journal on Computing, 19(5):968–988, October 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Amir:2006:FM

[AALP06] Amihood Amir, Yonatan Aumann, Moshe Lewenstein, and Ely Porat. Function matching. SIAM Journal on Computing, 35(5):1007–1022, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Awerbuch:2002:MFT

[AALR02] Baruch Awerbuch, Yossi Azar, Stefano Leonardi, and Oded Regev. Minimizing the flow time without migration. SIAM Journal on Computing, 31 (5):1370–1382, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37446.

Agarwal:1997:SUP

[AAOS97] Pankaj K. Agarwal, Boris Aronov, Joseph O’Rourke, and Catherine A. Schevon. Star unfolding of a polytope with applications. SIAM Journal on Computing, 26(6): 1689–1713, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25337.

Aaronson:2003:ABF

[Aar03] Scott Aaronson. Algorithms for Boolean function query properties. SIAM Journal on Computing, 32(5):1140– 1157, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37964.

Aaronson:2006:LBL

[Aar06] Scott Aaronson. Lower bounds for local search by quantum arguments. SIAM Journal on Computing, 35(4):804–824, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aaronson:2020:STQ

[Aar20] Scott Aaronson. Shadow tomography of quantum states. SIAM Journal on Computing, 49(5):STOC18–368–STOC18– 394, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Applebaum:2021:CDS

[AARV21] Benny Applebaum, Barak Arkis, Pavel Raykov, and Prashant Nalini Vasudevan. Conditional disclosure of secrets: Amplification, closure, amortization, lower-bounds, and separations. SIAM Journal on Computing, 50(1):32– 67, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:1997:CEF

[AAS97] Pankaj K. Agarwal, Boris Aronov, and Micha Sharir. Computing envelopes in four dimensions with applications. SIAM Journal on Computing, 26(6):1714–1732, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26572.

Alur:1997:TAA

[AAT97] Rajeev Alur, Hagit Attiya, and Gadi Taubenfeld. Timeadaptive algorithms for synchronization. SIAM Journal on Computing, 26(2):539–556, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26524.

Agarwal:2013:CCB

[AAvK^{+}13] Pankaj K. Agarwal, Boris Aronov, Marc van Kreveld, Maarten Löffler, and Rodrigo I. Silveira. Computing correlation between piecewiselinear functions. SIAM Journal on Computing, 42(5):1867– 1887, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Andrews:2016:MCN

[AAZ16] Matthew Andrews, Spyridon Antonakopoulos, and Lisa Zhang. Minimum-cost network design with (dis)economies of scale. SIAM Journal on Computing, 45(1):49–66, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aggarwal:1979:REM

[AB79] Vijay B. Aggarwal and James W.

Burgmeier. A round-off error model with applications to arithmetic expressions. SIAM Journal on Computing, 8(1): 60–72, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Atkinson:1994:PQP

[AB94] M. D. Atkinson and Robert Beals. Priority queues and permutations. SIAM Journal on Computing, 23(6): 1225–1230, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24089.

Apostolico:1996:OTP

[AB96] Alberto Apostolico and Dany Breslauer. An optimal O(log log n)-time parallel algorithm for detecting all squares in a string. SIAM Journal on Computing, 25(6): 1318–1331, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26040.

Amir:1998:TDP

[AB98] Amihood Amir and Gary Benson. Two-dimensional periodicity in rectangular arrays. SIAM Journal on Computing, 27(1):90–106, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29832.

Abboud:2024:RPN

[AB24] Amir Abboud and Greg Bodwin. Reachability preservers: New extremal bounds and approximation algorithms. SIAM Journal on Computing, 53(2): 221–246, March 2024. CODEN SMJCAT. ISSN 1095-7111.

ajBojanczyk:2008:TWA

[aBC08] Miko aj Bojańczyk and Thomas Colcombet. Tree-walking automata do not recognize all regular languages. SIAM Journal on Computing, 38(2):658– 701, ???? 2008. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:2013:NTB

[ABC^{+}13] Noga Alon, Eric Blais, Sourav Chakraborty, David García-Soriano, and Arie Matsliah. Nearly tight bounds for testing function isomorphism. SIAM Journal on Computing, 42(2): 459–493, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Awerbuch:1998:NLT

[ABCP98] Baruch Awerbuch, Bonnie Berger, Lenore Cowen, and David Peleg. Near-linear time construction of sparse neighborhood covers. SIAM Journal on Computing, 28 (1):263–277, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27189.

Arkin:2005:OCT

[ABD^{+}05] Esther M. Arkin, Michael A.

Bender, Erik D. Demaine, Sándor P. Fekete, Joseph S. B. Mitchell, and Saurabh Sethia. Optimal covering tours with turn costs. SIAM Journal on Computing, 35(3):531–566, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arge:2007:OCO

[ABD^{+}07] Lars Arge, Michael A. Bender, Erik D. Demaine, Bryan Holland-Minkley, and J. Ian Munro. An optimal cacheoblivious priority queue and its application to graph algorithms. SIAM Journal on Computing, 36(6):1672–1695, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Amir:1994:AIA

[ABF94] Amihood Amir, Gary Benson, and Martin Farach. An alphabet independent approach to two-dimensional pattern matching. SIAM Journal on Computing, 23(2):313–323, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22632.

Agarwala:1999:ANT

[ABF^{+}99] Richa Agarwala, Vineet Bafna, Martin Farach, Mike Paterson, and Mikkel Thorup. On the approximability of numerical taxonomy (fitting distances by tree metrics). SIAM Journal on Computing, 28(3):1073– 1085, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
29633.

Atkins:1998:SAS

[ABH98] Jonathan E. Atkins, Erik G.

Boman, and Bruce Hendrickson. A spectral algorithm for seriation and the consecutive ones problem. SIAM Journal on Computing, 28 (1):297–310, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28577.

Archer:2011:IAA

[ABHK11] Aaron Archer, Mohammad-Hossein Bateni, Mohammad-Taghi Hajiaghayi, and Howard Karloff. Improved approximation algorithms for prizecollecting Steiner tree and TSP. SIAM Journal on Computing, 40(2):309–332, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i2/
p309_s1.

Allender:1997:FOI

[ABI97] Eric Allender, José Balcázar, and Neil Immerman. A first-order isomorphism theorem. SIAM Journal on Computing, 26(2):557–567, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27023.

Aizenstein:1998:LRS

[ABK^{+}98] Howard Aizenstein, Avrim Blum, Roni Khardon, Eyal Kushilevitz, Leonard Pitt, and Dan Roth. On learning read-k-satisfy-j DNF. SIAM Journal on Computing, 27(6): 1515–1530, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27439.

Adler:2000:PSL

[ABK00] Micah Adler, John W. Byers, and Richard M. Karp. Parallel sorting with limited bandwidth. SIAM Journal on Computing, 29(6):1997– 2015, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        31588.

Alon:2004:LHM

[ABK^{+}04] Noga Alon, Richard Beigel, Simon Kasif, Steven Rudich, and Benny Sudakov. Learning a hidden matching. SIAM Journal on Computing, 33(2): 487–501, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.

org/sam-bin/dbq/article/
42013.

Allender:2006:PRS

[ABK^{+}06] Eric Allender, Harry Buhrman, Michal Koucký, Dieter van Melkebeek, and Detlef Ronneburger. Power from random strings. SIAM Journal on Computing, 35(6):1467–1493, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Allender:2009:CNA

[ABKPM09] Eric Allender, Peter Bürgisser, Johan Kjeldgaard-Pedersen, and Peter Bro Miltersen. On the complexity of numerical analysis. SIAM Journal on Computing, 38(5):1987–2006, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Azar:1999:BA

[ABKU99] Yossi Azar, Andrei Z. Broder, Anna R. Karlin, and Eli Upfal. Balanced allocations. SIAM Journal on Computing, 29(1):180–200, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28849.

Arvind:2000:CMG

[ABL00] V. Arvind, R. Beigel, and A. Lozano. The complexity of modular graph automorphism. SIAM Journal on Computing, 30(4):1299–1320, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35822.

Agarwal:2002:CCS

[ABL^{+}02] Pankaj K. Agarwal, Therese Biedl, Sylvain Lazard, Steve Robbins, Subhash Suri, and Sue Whitesides. Curvatureconstrained shortest paths in a convex polygon. SIAM Journal on Computing, 31(6): 1814–1851, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37455.

Aldous:2009:DPO

[ABL09] David J. Aldous, Charles Bordenave, and Marc Lelarge. Dynamic programming optimization over random data: The scaling exponent for nearoptimal solutions. SIAM Journal on Computing, 38(6):2382– 2410, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abraham:2015:EMU

[ABN15] Ittai Abraham, Yair Bartal, and Ofer Neiman. Embedding metrics into ultrametrics and graphs into spanning trees with constant average distortion. SIAM Journal on Computing, 44(1):160–192, ???? 2015. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Aharonov:2008:FTQ

[ABO08] Dorit Aharonov and Michael Ben-Or. Fault-tolerant quantum computation with constant error rate. SIAM Journal on Computing, 38(4):1207– 1282, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aronov:2020:RSQ

[ABOK20] Boris Aronov, Gali Bar-On, and Matthew J. Katz. Resolving SINR queries in a dynamic setting. SIAM Journal on Computing, 49(6):1271– 1290, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abboud:2018:HLB

[ABP18] Amir Abboud, Greg Bodwin, and Seth Pettie. A hierarchy of lower bounds for sublinear additive spanners. SIAM Journal on Computing, 47(6): 2203–2236, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arunachalam:2019:QQA

[ABP19] Srinivasan Arunachalam, Jop Briët, and Carlos Palazuelos. Quantum query algorithms are completely bounded forms. SIAM Journal on Computing, 48(3):903–925, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abrahamson:1987:GSM

[Abr87] Karl Abrahamson. Generalized string matching. SIAM Journal on Computing, 16(6):1039– 1051, December 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Annexstein:1990:GAG

[ABR90] Fred Annexstein, Marc Baumslag, and Arnold L. Rosenberg. Group action graphs and parallel architectures. SIAM Journal on Computing, 19(3):544– 569, June 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alvarez:2004:CUS

[ÀBS04] Carme Àlvarez, Maria Blesa, and Maria Serna. A characterization of universal stability in the adversarial queuing model. SIAM Journal on Computing, 34(1):41–66, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43552.

Alekhnovich:2007:LUB

[ABS07] Mikhail Alekhnovich and Eli Ben-Sasson. Linear upper bounds for random walk on small density random 3-CNFs. SIAM Journal on Computing, 36(5):1248–1263, ???? 2007.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Avis:1998:UMP

[ABSD^{+}98] David Avis, Bryan Beresford-Smith, Luc Devroye, Hossam Elgindy, Eric Guévremont, Ferran Hurtado, and Binhai Zhu. Unoriented thetamaxima in the plane: Complexity and algorithms. SIAM Journal on Computing, 28 (1):278–296, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27787.

Alekhnovich:2002:SCP

[ABSRW02] Michael Alekhnovich, Eli Ben-Sasson, Alexander A. Razborov, and Avi Wigderson. Space complexity in propositional calculus. SIAM Journal on Computing, 31 (4):1184–1211, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36673.

Alekhnovich:2004:PGP

[ABSRW04] Michael Alekhnovich, Eli Ben-Sasson, Alexander A. Razborov, and Avi Wigderson. Pseudorandom generators in propositional proof complexity. SIAM Journal on Computing, 34(1):67–88, February 2004. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38994.

Applebaum:2021:PSC

[ABT21] Benny Applebaum, Zvika Brakerski, and Rotem Tsabary. Perfect secure computation in two rounds. SIAM Journal on Computing, 50(1):68–97, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abboud:2018:ICC

[ABW18] Amir Abboud, Arturs Backurs, and Virginia Vassilevska Williams. If the current clique algorithms are optimal, so is Valiant’s parser. SIAM Journal on Computing, 47(6):2527– 2555, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arjomandi:1978:PCG

[AC78] Eshrat Reghbati (Arjomandi) and D. G. Corneil. Parallel computations in graph theory. SIAM Journal on Computing, 7(2):230–237, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Achugbue:1982:SOS

[AC82] James O. Achugbue and Francis Y. Chin. Scheduling the open shop to minimize mean flow time. SIAM Journal on Computing, 11(4):709–720, ???? 1982. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Ailon:2009:FJL

[AC09] Nir Ailon and Bernard Chazelle.

The fast Johnson–Lindenstrauss transform and approximate nearest neighbors. SIAM Journal on Computing, 39(1):302– 322, ???? 2009. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ailon:2011:FTM

[AC11] Nir Ailon and Moses Charikar.

Fitting tree metrics: Hierarchical clustering and phylogeny. SIAM Journal on Computing, 40(5):1275–1291, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i5/p1275_s1.

Auer:2002:NMB

[ACBFS02] Peter Auer, Nicolò Cesa-Bianchi, Yoav Freund, and Robert E. Schapire. The nonstochastic multiarmed bandit problem. SIAM Journal on Computing, 32(1):48– 77, February 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39837.

Alon:2017:NMA

[ACBG^{+}17] Noga Alon, Nicolò Cesa-Bianchi, Claudio Gentile, Shie Mannor, Yishay Mansour, and Ohad Shamir. Nonstochastic multi-armed bandits with graph-structured feedback. SIAM Journal on Computing, 46(6):1785–1826, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ailon:2011:SIA

[ACC^{+}11] Nir Ailon, Bernard Chazelle, Kenneth L. Clarkson, Ding Liu, Wolfgang Mulzer, and C. Seshadhri. Self-improving algorithms. SIAM Journal on Computing, 40(2):350–375, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i2/
p350_s1.

Aceto:1994:AFC

[Ace94] Luca Aceto. On “Axiomatising Finite Concurrent Processes”. SIAM Journal on Computing, 23(4):852–863, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24360.  See

[Hen88].

Adamaszek:2022:ATB

[ACER22] Anna Adamaszek, Artur Czumaj, Matthias Englert, and Harald Räcke. Almost tight bounds for reordering buffer management. SIAM Journal on Computing, 51(3):701–722, ???? 2022. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1326167.

Attiya:1998:SMC

[ACFW98] Hagit Attiya, Soma Chaudhuri, Roy Friedman, and Jennifer L. Welch. Shared memory consistency conditions for nonsequential execution: Definitions and programming strategies. SIAM Journal on Computing, 27(1):65–89, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27839.

Atallah:1989:CDC

[ACG89] Mikhail J. Atallah, Richard Cole, and Michael T. Goodrich. Cascading divide-and-conquer: a technique for designing parallel algorithms. SIAM Journal on Computing, 18(3):499–532, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aggarwal:2005:CGM

[ACG^{+}05] Gagan Aggarwal, Qi Cheng, Michael H. Goldwasser, Ming-Yang Kao, Pablo Moisset de Espanes, and Robert T. Schweller. Complexities for generalized models of selfassembly. SIAM Journal on Computing, 34(6):1493–1515, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        44520.

Aharonov:2016:SPE

[ACG^{+}16] Dorit Aharonov, André Chailloux, Maor Ganz, Iordanis Kerenidis, and Loïck Magnin. A simpler proof of the existence of quantum weak coin flipping with arbitrarily small bias. SIAM Journal on Computing, 45(3):633–679, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alexi:1988:RRF

[ACGS88] Werner Alexi, Benny Chor, Oded Goldreich, and Claus-P. Schnorr. RSA and Rabin functions: certain parts are as hard as the whole. SIAM Journal on Computing, 17(2):194–209, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Attiya:2010:LBR

[ACH10] Hagit Attiya and Keren Censor-Hillel. Lower bounds for randomized consensus under a weak adversary. SIAM Journal on Computing, 39(8): 3885–3904, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v39/i8/p3885_s1.

Aspnes:2016:LBR

[ACHAH16] James Aspnes, Keren Censor-Hillel, Hagit Attiya, and Danny Hendler. Lower bounds for restricted-use objects. SIAM Journal on Computing, 45(3):734–762, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Andoni:2021:SSA

[ACHCP21] Alexandr Andoni, Keren Censor-Hillel, Jing Chen, and Debmalya Panigrahi. Special section on the 48th Annual ACM Symposium on Theory of Computing (STOC 2016). SIAM Journal on Computing, 50(3):STOC16–i–STOC16–iii, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Attiya:2019:BSN

[ACHP19] Hagit Attiya, Armando Castañeda, Maurice Herlihy, and Ami Paz. Bounds on the step and namespace complexity of renaming. SIAM Journal on Computing, 48(1):1–32, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aingworth:1999:FED

[ACIM99] D. Aingworth, C. Chekuri, P. Indyk, and R. Motwani. Fast estimation of diameter and shortest paths (without matrix multiplication). SIAM Journal on Computing, 28 (4):1167–1181, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30342.

Awerbuch:1998:OBP

[ACK^{+}98] Baruch Awerbuch, Israel Cidon, Shay Kutten, Yishay Mansour, and David Peleg. Optimal broadcast with partial knowledge. SIAM Journal on Computing, 28(2):511–524, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27993.

Aggarwal:1999:AMT

[ACK^{+}99] Alok Aggarwal, Don Coppersmith, Sanjeev Khanna, Rajeev Motwani, and Baruch Schieber. The angular-metric traveling salesman problem. SIAM Journal on Computing, 29(3):697–711, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31272.

Abraham:2015:LDI

[ACKS15] Ittai Abraham, Shiri Chechik, David Kempe, and Aleksandrs Slivkins. Low-distortion inference of latent similarities from a multiplex social network. SIAM Journal on Computing, 44(3):617–668, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arnon:1984:CADa

[ACM84a] Dennis S. Arnon, George E.

Collins, and Scott McCallum. Cylindrical algebraic decomposition. I. the basic algorithm. SIAM Journal on Computing, 13(4):865–877, November 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arnon:1984:CADb

[ACM84b] Dennis S. Arnon, George E.

Collins, and Scott McCallum. Cylindrical algebraic decomposition. II. an adjacency algorithm for the plane. SIAM Journal on Computing, 13(4): 878–889, November 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arkin:1999:MST

[ACM^{+}99] Esther M. Arkin, Yi-Jen Chiang, Joseph S. B. Mitchell, Steven S. Skiena, and Tae-Cheon Yang. On the maximum scatter traveling salesperson problem. SIAM Journal on Computing, 29(2):515–544, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32028.

Alon:2010:QRA

[ACOH^{+}10] Noga Alon, Amin Coja-Oghlan, Hiêp Hàn, Mihyun Kang, Vojtĕch Rödl, and Mathias Schacht. Quasirandomness and algorithmic regularity for graphs with general degree distributions. SIAM Journal on Computing, 39(6):2336–2362, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ambainis:2010:FSC

[ACR^{+}10] A. Ambainis, A. M. Childs, B. W. Reichardt, R. Špalek, and S. Zhang. Any AND-OR formula of size N can be evaluated in time N^{1/2+o(1)} on a quantum computer. SIAM Journal on Computing, 39(6): 2513–2530, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Andreev:1999:WRS

[ACRT99] Alexander E. Andreev, Andrea E. F. Clementi, José D. P. Rolim, and Luca Trevisan. Weak random sources, hitting sets, and BPP simulations. SIAM Journal on Computing, 28(6):2103–2116, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32563.

Aguilera:2000:QRC

[ACT00] Marcos Kawazoe Aguilera, Wei Chen, and Sam Toueg. On quiescent reliable communication. SIAM Journal on Computing, 29(6):2040–2073, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34129.

Aaronson:2014:FCQ

[AD14] Scott Aaronson and Andrew Drucker. A full characterization of quantum advice. SIAM Journal on Computing, 43(3): 1131–1183, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aronov:2014:IBU

[AdBES14] Boris Aronov, Mark de Berg, Esther Ezra, and Micha Sharir. Improved bounds for the union of locally fat objects in the plane. SIAM Journal on Computing, 43(2):543–572, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:1998:CLA

[AdBMS98] Pankaj K. Agarwal, Mark de Berg, Jiří Matousek, and Otfried Schwarzkopf. Constructing levels in arrangements and higher order Voronoi diagrams. SIAM Journal on Computing, 27(3):654– 667, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28184.

Abam:2010:KTL

[AdBS10] Mohammad Ali Abam, Mark de Berg, and Bettina Speckmann. Kinetic kd-trees and longest-side kd-trees. SIAM Journal on Computing, 39(4): 1219–1232, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abam:2019:GSP

[AdBS19] Mohammad Ali Abam, Mark de Berg, and Mohammad Javad Rezaei Seraji. Geodesic spanners for points on a polyhedral terrain. SIAM Journal on Computing, 48(6):1796– 1810, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arya:2018:APM

[AdFM18] Sunil Arya, Guilherme D. da Fonseca, and David M. Mount. Approximate polytope membership queries. SIAM Journal on Computing, 47(1): 1–51, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Atserias:2008:PUE

[ADG08] Albert Atserias, Anuj Dawar, and Martin Grohe. Preservation under extensions on well-behaved finite structures. SIAM Journal on Computing, 38(4):1364–1381, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aland:2011:EPA

[ADG^{+}11] Sebastian Aland, Dominic Dumrauf, Martin Gairing, Burkhard Monien, and Florian Schoppmann. Exact price of anarchy for polynomial congestion games. SIAM Journal on Computing, 40(5): 1211–1233, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i5/p1211_s1.

aufderHeide:1986:ESA

[adH86] Friedhelm Meyer auf der Heide. Efficient simulations among several models of parallel computers. SIAM Journal on Computing, 15(1):106–119, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Adleman:1997:QC

[ADH97] Leonard M. Adleman, Jonathan Demarrais, and Ming-Deh A. Huang. Quantum computability. SIAM Journal on Computing, 26(5):1524–1540, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29363.

aufderHeide:1987:CPS

[adHW87] Friedhelm Meyer auf der Heide and Avi Wigderson. The complexity of parallel sorting. SIAM Journal on Computing, 16(1):100–107, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Akcoglu:2004:FUI

[ADK04] Karhan Akcoglu, Petros Drineas, and Ming-Yang Kao. Fast universalization of investment strategies. SIAM Journal on Computing, 34(1):1–22, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40561.

Anshelevich:2008:PSN

[ADK^{+}08] Elliot Anshelevich, Anirban Dasgupta, Jon Kleinberg, Éva Tardos, Tom Wexler, and Tim Roughgarden. The price of stability for network design with fair cost allocation. SIAM Journal on Computing, 38(4): 1602–1623, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aggarwal:2018:NMC

[ADL18] Divesh Aggarwal, Yevgeniy Dodis, and Shachar Lovett. Non-malleable codes from additive combinatorics. SIAM Journal on Computing, 47(2): 524–546, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Adolphson:1977:SMJ

[Ado77] Donald L. Adolphson. Single machine job sequencing with precedence constraints. SIAM Journal on Computing, 6(1): 40–54, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ausiello:1986:MRD

[ADS86] G. Ausiello, A. D’Atri, and D. Saccà. Minimal representation of directed hypergraphs. SIAM Journal on Computing, 15(2):418–431, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Apers:2022:QSG

[AdW22] Simon Apers and Ronald de Wolf. Quantum speedup for graph sparsification, cut approximation, and Laplacian solving. SIAM Journal on Computing, 51(6):1703–1742, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1391018.

Aharonov:2015:QLT

[AE15] Dorit Aharonov and Lior Eldar. Quantum locally testable codes. SIAM Journal on Computing, 44(5):1230–1262, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aaronson:2011:SSF

[AEM^{+}11] Scott Aaronson, Jeff Erickson, Mohammad Mahdian, R. Ravi, and Emanuele Viola. Special section on foundations of computer science. SIAM Journal on Computing, 40(3):770, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i3/p770_s1.

Agarwal:1999:VDS

[AES99] Pankaj K. Agarwal, Alon Efrat, and Micha Sharir. Vertical decomposition of shallow levels in 3-dimensional arrangements and its applications. SIAM Journal on Computing, 29(3):912–953, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29593.

Aronov:2010:SSN

[AES10] Boris Aronov, Esther Ezra, and Micha Sharir. Small-size ϵ-nets for axis-parallel rectangles and boxes. SIAM Journal on Computing, 39(7):3248– 3282, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aronov:2020:CPP

[AEZ20] Boris Aronov, Esther Ezra, and Joshua Zahl. Constructive polynomial partitioning for algebraic curves in R^{3} with applications. SIAM Journal on Computing, 49(6):1109–1127, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Attiya:1998:CCH

[AF98] Hagit Attiya and Roy Friedman. A correctness condition for high-performance multiprocessors. SIAM Journal on Computing, 27(6): 1637–1670, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28921.

Attiya:2001:AEA

[AF01] Hagit Attiya and Arie Fouren.

Adaptive and efficient algorithms for lattice agreement and renaming. SIAM Journal on Computing, 31(2):642–664, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36600.

Arya:2003:ECC

[AF03] Sunil Arya and Ho-Yam Addy Fu. Expected-case complexity of approximate nearest neighbor searching. SIAM Journal on Computing, 32(3):793–815, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36634.

Atserias:2023:DEM

[AF23] Albert Atserias and Joanna Fijalkow. Definable ellipsoid method, sums-of-squares proofs, and the graph isomorphism problem. SIAM Journal on Computing, 52 (5):1193–1229, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1338435.

Agarwala:1994:PTA

[AFB94] Richa Agarwala and David Fernández-Baca. A polynomialtime algorithm for the perfect phylogeny problem when the number of character states is fixed. SIAM Journal on Computing, 23(6):1216–1224, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        24458.

Agarwala:1996:WMS

[AFB96] Richa Agarwala and David Fernández-Baca. Weighted multidimensional search and its application to convex optimization. SIAM Journal on Computing, 25(1):83–99, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        24192.

Amir:2014:MUL

[AFG^{+}14] Amihood Amir, Gianni Franceschini, Roberto Grossi, Tsvi Kopelowitz, Moshe Lewenstein, and Noa Lewenstein. Managing unbounded-length keys in comparison-driven data structures with applications to online indexing. SIAM Journal on Computing, 43(4):1396– 1416, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abraham:2022:MES

[AFGN22] Ittai Abraham, Arnold Filtser, Anupam Gupta, and Ofer Neiman. Metric embedding via shortest path decompositions. SIAM Journal on Computing, 51(2):290–314, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1296021. See corri-

gendum [AFGN23].

Abraham:2023:CME

[AFGN23] Ittai Abraham, Arnold Filtser, Anupam Gupta, and Ofer Neiman. Corrigendum: Metric embedding via shortest path decompositions. SIAM Journal on Computing, 52(5):1319– 1320, October 2023. CODEN SMJCAT. ISSN 1095-7111. See [AFGN22].

Andrews:2000:GDR

[AFHB^{+}00] Matthew Andrews, Antonio Fernández, Mor Harchol-Balter, Tom Leighton, and Lisa Zhang. General dynamic routing with per-packet delay guarantees of O(Distance+ 1/SessionRate). SIAM Journal on Computing, 30(5):1594– 1623, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35061.

Arenas:2008:CVC

[AFL08] Marcelo Arenas, Wenfei Fan, and Leonid Libkin. On the complexity of verifying consistency of XML specifications. SIAM Journal on Computing, 38(3):841–880, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:2007:ETB

[AFN07] Noga Alon, Eldar Fischer, and Ilan Newman. Efficient testing of bipartite graphs for forbidden induced subgraphs. SIAM Journal on Computing, 37(3): 959–976, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:2009:CCT

[AFNS09] Noga Alon, Eldar Fischer, Ilan Newman, and Asaf Shapira. A combinatorial characterization of the testable graph properties: It’s all about regularity. SIAM Journal on Computing, 39(1):143–167, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arnon-Friedman:2019:STD

[AFRV19] Rotem Arnon-Friedman, Renato Renner, and Thomas Vidick. Simple and tight device-independent security proofs. SIAM Journal on Computing, 48(1):181–225, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Apostolico:1986:BMG

[AG86] Alberto Apostolico and Raffaele Giancarlo. The Boyer– Moore–Galil string searching strategies revisited. SIAM Journal on Computing, 15(1): 98–105, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Afek:1991:TMB

[AG91] Yehuda Afek and Eli Gafni.

Time and message bounds for election in synchronous and asynchronous complete networks. SIAM Journal on Computing, 20(2):376–394, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Afek:1994:DAU

[AG94a] Yehuda Afek and Eli Gafni.

Distributed algorithms for unidirectional networks. SIAM Journal on Computing, 23(6): 1152–1178, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/23277.

Allender:1994:UCL

[AG94b] Eric Allender and Vivek Gore. A uniform circuit lower bound for the permanent. SIAM Journal on Computing, 23 (5):1026–1049, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/23390.

Agarwal:1992:RSO

[Aga92] Pankaj K. Agarwal. Ray shooting and other applications of spanning trees with low stabbing number. SIAM Journal on Computing, 21(3):540–570, June 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abraham:2019:CRT

[AGG^{+}19] Ittai Abraham, Cyril Gavoille, Anupam Gupta, Ofer Neiman, and Kunal Talwar. Cops, robbers, and threatening skeletons: Padded decomposition for minor-free graphs. SIAM Journal on Computing, 48(3): 1120–1145, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Austrin:2017:SNH

[AGH17] Per Austrin, Venkatesan Guruswami, and Johan Håstad.

(2 + ε)-sat is NP-hard. SIAM Journal on Computing, 46(5): 1554–1573, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arya:2004:LSH

[AGK^{+}04] Vijay Arya, Naveen Garg, Rohit Khandekar, Adam Meyerson, Kamesh Munagala, and Vinayaka Pandit. Local search heuristics for kmedian and facility location problems. SIAM Journal on Computing, 33(3):544–562, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41640.

Aaronson:2009:SIT

[AGK^{+}09] Scott Aaronson, Sudipto Guha, Jon Kleinberg, Frank McSherry, Dieter van Melkebeek, and Amit Sahai. Special issue on the Thirty-Eighth Annual ACM Symposium On Theory Of Computing (STOC 2006). SIAM Journal on Computing, 39(1):vii, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arora:2016:CNM

[AGKM16] Sanjeev Arora, Rong Ge, Ravi Kannan, and Ankur Moitra. Computing a nonnegative matrix factorization — provably. SIAM Journal on Computing, 45(4):1582–1611, ???? 2016.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agrawal:2015:HSR

[AGKS15] Manindra Agrawal, Rohit Gurjar, Arpita Korwar, and Nitin Saxena. Hitting-sets for ROABP and sum of setmultilinear circuits. SIAM Journal on Computing, 44(3): 669–697, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Anily:1999:SPL

[AGL99] Shoshana Anily, Michel Gendreau, and Gilbert Laporte. The swapping problem on a line. SIAM Journal on Computing, 29(1):327–335, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32310.

Ackermann:2011:UTS

[AGM^{+}11] Heiner Ackermann, Paul W. Goldberg, Vahab S. Mirrokni, Heiko Röglin, and Berthold Vöcking. Uncoordinated twosided matching markets. SIAM Journal on Computing, 40(1): 92–106, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i1/p92_s1.

Atserias:2013:SBQ

[AGM13] Albert Atserias, Martin Grohe, and Dániel Marx. Size bounds and query plans for relational joins. SIAM Journal on Computing, 42(4):1737–1767, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:2000:BSP

[AGMV00] Pankaj K. Agarwal, Edward F.

Grove, T. M. Murali, and Jeffrey Scott Vitter. Binary space partitions for fat rectangles. SIAM Journal on Computing, 29(5):1422–1448, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32057.

Arunachalam:2021:QHL

[AGS21] Srinivasan Arunachalam, Alex Bredariol Grilo, and Aarthi Sundaram. Quantum hardness of learning shallow classical circuits. SIAM Journal on Computing, 50(3):972–1013, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aho:1972:TRD

[AGU72] A. V. Aho, M. R. Garey, and J. D. Ullman. The transitive reduction of a directed graph. SIAM Journal on Computing, 1(2):131–137, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Allender:2018:MCS

[AGvM^{+}18] Eric Allender, Joshua A. Grochow, Dieter van Melkebeek, Cristopher Moore, and Andrew Morgan. Minimum circuit size, graph isomorphism, and related problems. SIAM Journal on Computing, 47(4): 1339–1372, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Anily:1989:RBB

[AH89] S. Anily and R. Hassin. Ranking the best binary trees. SIAM Journal on Computing, 18(5):882–892, October 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Albers:2000:EUE

[AH00] Susanne Albers and Monika R. Henzinger. Exploring unknown environments. SIAM Journal on Computing, 29 (4):1164–1188, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32428.

Attiya:2008:RDR

[AH08] Hagit Attiya and David Hay. Randomization does not reduce the average delay in parallel packet switches. SIAM Journal on Computing, 37(5): 1613–1636, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Austrin:2011:RSI

[AH11] Per Austrin and Johan Håstad.

Randomly supported independence and resistance. SIAM Journal on Computing, 40 (1):1–27, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i1/p1_s1.

Andersson:2000:TBS

[AHHP00] Arne Andersson, Torben Hagerup, Johan Håstad, and Ola Petersson. Tight bounds for searching a sorted array of strings. SIAM Journal on Computing, 30(5):1552– 1578, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32988.

Amir:2010:CIR

[AHK^{+}10a] Amihood Amir, Tzvika Hartman, Oren Kapah, Avivit Levy, and Ely Porat. On the cost of interchange rearrangement in strings. SIAM Journal on Computing, 39(4): 1444–1461, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arora:2010:ASC

[AHK10b] Sanjeev Arora, Elad Hazan, and Satyen Kale. O(\sqrt{log n}) approximation to SPARSEST CUT in Õ(n^{2}) time. SIAM Journal on Computing, 39(5): 1748–1771, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Awerbuch:2007:LCS

[AHKL07] Baruch Awerbuch, Mohammad T. Hajiaghayi, Robert Kleinberg, and Tom Leighton. Localized client-server load balancing without global information. SIAM Journal on Computing, 37(4):1259–1279, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Athreya:2007:ESD

[AHLM07] Krishna B. Athreya, John M. Hitchcock, Jack H. Lutz, and Elvira Mayordomo. Effective strong dimension in algorithmic information and computational complexity. SIAM Journal on Computing, 37(3):671– 705, ???? 2007. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alstrup:1999:DLT

[AHLT99] Stephen Alstrup, Dov Harel, Peter W. Lauridsen, and Mikkel Thorup. Dominators in linear time. SIAM Journal on Computing, 28(6): 2117–2132, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31726.

Allender:2008:MDN

[AHM^{+}08] Eric Allender, Lisa Hellerstein, Paul McCabe, Toniann Pitassi, and Michael Saks. Minimizing disjunctive normal form formulas and AC^{0} circuits given a truth table. SIAM Journal on Computing, 38(1):63–84, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alt:1987:DSI

[AHMP87] Helmut Alt, Torben Hagerup, Kurt Mehlhorn, and Franco P. Preparata. Deterministic simulation of idealized parallel computers on more realistic ones. SIAM Journal on Computing, 16(5):808–835, October 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Allender:1992:RER

[AHOW92] Eric Allender, Lane A. Hemachandra, Mitsunori Ogiwara, and Osamu Watanabe. Relating equivalence and reducibility to sparse sets. SIAM Journal on Computing, 21(3):521–539, June 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aronov:2008:ADR

[AHP08] Boris Aronov and Sariel Har-Peled. On approximating the depth and related problems. SIAM Journal on Computing, 38(3):899–921, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:2013:ESC

[AHPY13] Pankaj K. Agarwal, Sariel Har-Peled, and Hai Yu. Embeddings of surfaces, curves, and moving points in Euclidean space. SIAM Journal on Computing, 42(2):442–458, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Attiya:1996:OCS

[AHR96] Hagit Attiya, Amir Herzberg, and Sergio Rajsbaum. Optimal clock synchronization under different delay assumptions. SIAM Journal on Computing, 25(2):369–389, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26632.

Aldous:1992:MSD

[AHS92] David Aldous, Micha Hofri, and Wojciech Szpankowski. Maximum size of a dynamic data structure: Hashing with lazy deletion revisited. SIAM Journal on Computing, 21(4): 713–732, August 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aho:1976:FLC

[AHU76] A. V. Aho, J. E. Hopcroft, and J. D. Ullman. On finding lowest common ancestors in trees. SIAM Journal on Computing, 5(1):115–132, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aoshima:1977:CFH

[AI77] K. Aoshima and M. Iri. Comments on F. Hadlock’s paper: “Finding a Maximum Cut of a Planar Graph in Polynomial Time”. SIAM Journal on Computing, 6(1):86–87, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Apostolico:1994:SUP

[AIGT94] Alberto Apostolico, Giuseppe F.

Italiano, Giorgio Gambosi, and Maurizio Talamo. The set union problem with unlimited backtracking. SIAM Journal on Computing, 23(1):50– 70, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
17359.

Applebaum:2006:C

[AIK06] Benny Applebaum, Yuval Ishai, and Eyal Kushilevitz. Cryptography in NC^{0}. SIAM Journal on Computing, 36(4): 845–888, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Applebaum:2014:HGA

[AIK14] Benny Applebaum, Yuval Ishai, and Eyal Kushilevitz. How to garble arithmetic circuits. SIAM Journal on Computing, 43(2):905–929, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Achlioptas:2019:LLF

[AIK19] Dimitris Achlioptas, Fotis Iliopoulos, and Vladimir Kolmogorov. A local lemma for focused stochastic algorithms. SIAM Journal on Computing, 48(5):1583–1602, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Antoniadis:2020:HHE

[AIK^{+}20] Antonios Antoniadis, Sungjin Im, Ravishankar Krishnaswamy, Benjamin Moseley, Viswanath Nagarajan, Kirk Pruhs, and Clifford Stein. Hallucination helps: Energy efficient virtual circuit routing. SIAM Journal on Computing, 49(1):37– 66, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abraham:2007:PM

[AIKM07] David J. Abraham, Robert W. Irving, Telikepalli Kavitha, and Kurt Mehlhorn. Popular matchings. SIAM Journal on Computing, 37(4):1030–1045, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Applebaum:2015:EFC

[AIKW15] Benny Applebaum, Yuval Ishai, Eyal Kushilevitz, and Brent Waters. Encoding functions with constant online rate, or how to compress garbled circuit keys. SIAM Journal on Computing, 44(2):433–466, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aspnes:2016:SSF

[AIM16] James Aspnes, Yuval Ishai, and Peter Bro Miltersen. Special section on the Forty-Fifth Annual ACM Symposium on the Theory of Computing (STOC 2013). SIAM Journal on Computing, 45(2): 459–460, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Augustine:2008:OPS

[AIS08] John Augustine, Sandy Irani, and Chaitanya Swamy. Optimal power-down strategies. SIAM Journal on Computing, 37(5):1499–1516, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Anand:2022:PSI

[AJ22] Konrad Anand and Mark Jerrum. Perfect sampling in infinite spin systems via strong spatial mixing. SIAM Journal on Computing, 51 (4):1280–1295, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1437433.

Atallah:1988:ESS

[AK88] Mikhail J. Atallah and S. Rao Kosaraju. Efficient solutions to some transportation problems with applications to minimizing robot arm travel. SIAM Journal on Computing, 17(5): 849–869, October 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:1997:STC

[AK97a] Noga Alon and Nabil Kahale. A spectral technique for coloring random 3-colorable graphs. SIAM Journal on Computing, 26(6):1733–1748, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27024.

Amir:1997:MAS

[AK97b] Amihood Amir and Dmitry Keselman. Maximum agreement subtree in a set of evolutionary trees: Metrics and efficient algorithms. SIAM Journal on Computing, 26(6): 1656–1669, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26946.

Arora:2003:ASM

[AK03] Sanjeev Arora and George Karakostas. Approximation schemes for minimum latency problems. SIAM Journal on Computing, 32(5):1317– 1337, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39965.

Achlioptas:2007:SSF

[AK07] Dimitris Achlioptas and Vladlen Koltun. Special section on foundations of computer science. SIAM Journal on Computing, 37(1):165, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Awerbuch:2009:SDG

[AK09] Baruch Awerbuch and Rohit Khandekar. Stateless distributed gradient descent for positive linear programs. SIAM Journal on Computing, 38(6):2468–2486, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Andoni:2010:CHE

[AK10] Alexandr Andoni and Robert Krauthgamer. The computational hardness of estimating edit distance. SIAM Journal on Computing, 39(6):2398– 2429, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Anshelevich:2011:TBM

[AK11] Elliot Anshelevich and Adriana Karagiozova. Terminal backup, 3D matching, and covering cubic graphs. SIAM Journal on Computing, 40(3): 678–708, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i3/p678_s1.

Applebaum:2023:SGF

[AK23] Benny Applebaum and Eliran Kachlon. Sampling graphs without forbidden subgraphs and unbalanced expanders with negligible error. SIAM Journal on Computing, 52(6): 1321–1368, November 2023. CODEN SMJCAT. ISSN 1095-7111.

Akkoyunlu:1973:EMC

[Akk73] E. A. Akkoyunlu. The enumeration of maximal cliques of large graphs. SIAM Journal on Computing, 2(1):1–6, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Anshelevich:2008:SLB

[AKK08] Elliot Anshelevich, David Kempe, and Jon Kleinberg. Stability of load balancing algorithms in dynamic adversarial systems. SIAM Journal on Computing, 37(5):1656–1673, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Adleman:2009:UIR

[AKK^{+}09] Leonard Adleman, Jarkko Kari, Lila Kari, Dustin Reishus, and Petr Sosik. The undecidability of the infinite ribbon problem: Implications for computing by selfassembly. SIAM Journal on Computing, 38(6):2356–2381, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Assadi:2021:TBS

[AKL21] Sepehr Assadi, Sanjeev Khanna, and Yang Li. Tight bounds for single-pass streaming complexity of the set cover problem. SIAM Journal on Computing, 50(3):STOC16–341–STOC16– 376, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Asano:2003:MRU

[AKOT03] Tetsuo Asano, Naoki Katoh, Koji Obokata, and Takeshi Tokuyama. Matrix rounding under the L_{p}-discrepancy measure and its application to digital halftoning. SIAM Journal on Computing, 32(6): 1423–1435, December 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41751.

Alon:1995:GTG

[AKPW95] Noga Alon, Richard M. Karp, David Peleg, and Douglas West. A graph-theoretic game and its application to the kserver problem. SIAM Journal on Computing, 24(1):78– 100, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        22447.

Agrawal:1995:WTC

[AKR95] Ajit Agrawal, Philip Klein, and R. Ravi. When trees collide: an approximation algorithm for the generalized Steiner problem on networks. SIAM Journal on Computing, 24(3):440–456, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/23623.

Andoni:2018:SEE

[AKR18] Alexandr Andoni, Robert Krauthgamer, and Ilya Razenshteyn. Sketching and embedding are equivalent for norms. SIAM Journal on Computing, 47(3):890–916, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Allender:2009:SST

[AKS09] Eric Allender, Vladlen Koltun, and Maxim Sviridenko. Special section on the thirty-ninth annual ACM symposium on theory of computing (STOC 2007). SIAM Journal on Computing, 39(3):978, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aiger:2014:RNH

[AKS14] Dror Aiger, Haim Kaplan, and Micha Sharir. Reporting neighbors in high-dimensional Euclidean space. SIAM Journal on Computing, 43(4):1363– 1395, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Assadi:2022:SCC

[AKSW22] Sepehr Assadi, Hrishikesh Khandeparkar, Raghuvansh R. Saxena, and S. Matthew Weinberg. Separating the communication complexity of truthful and nontruthful algorithms for combinatorial auctions. SIAM Journal on Computing, 51(3):STOC20– 75–STOC20–114, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1370021.

Anagnostopoulos:2005:LBA

[AKU05] Aris Anagnostopoulos, Adam Kirsch, and Eli Upfal. Load balancing in arbitrary network topologies with stochastic adversarial input. SIAM Journal on Computing, 34(3):616–639, June 2005. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43783.

Alon:1990:LC

[AKW90] Noga Alon, Mauricio Karchmer, and Avi Wigderson. Linear circuits over GF(2). SIAM Journal on Computing, 19(6): 1064–1067, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aggarwal:2000:NDP

[AKW00] Alok Aggarwal, Jon Kleinberg, and David P. Williamson. Node-disjoint paths on the mesh and a new trade-off in VLSI layout. SIAM Journal on Computing, 29(4):1321– 1333, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        31273.

Atkin:1982:PTS

[AL82] A. O. L. Atkin and R. G. Larson. On a primality test of Solovay and Strassen. SIAM Journal on Computing, 11(4): 789–791, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Atkinson:1983:RT

[AL83] M. D. Atkinson and S. Lloyd. The ranks of m × n × (mn − 2) tensors. SIAM Journal on Computing, 12(4):611–615, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Avis:1988:PAH

[AL88] David Avis and C. W. Lai.

The probabilistic analysis of a heuristic for the assignment problem. SIAM Journal on Computing, 17(4):732–741, August 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aiello:2007:HCH

[AL07] William Aiello and F. T.

Leighton. Hamming codes, hypercube embeddings, and fault tolerance. SIAM Journal on Computing, 37(3):783–803, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arad:2010:QCE

[AL10] Itai Arad and Zeph Landau.

Quantum computation and the evaluation of tensor networks. SIAM Journal on Computing, 39(7):3089–3121, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Applebaum:2018:AAA

[AL18] Benny Applebaum and Shachar Lovett. Algebraic attacks against random local functions and their countermeasures. SIAM Journal on Computing, 47(1):52–79, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alaei:2014:BCA

[Ala14] Saeed Alaei. Bayesian combinatorial auctions: Expanding single buyer mechanisms to many buyers. SIAM Journal on Computing, 43(2):930–972, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alblas:1985:FMP

[Alb85] Henk Alblas. Finding minimal pass sequences for attribute grammars. SIAM Journal on Computing, 14(4):889– 914, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Albers:1998:IRL

[Alb98] Susanne Albers. Improved randomized on-line algorithms for the list update problem. SIAM Journal on Computing, 27(3):682–693, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27785.

Albers:1999:BBO

[Alb99] Susanne Albers. Better bounds for online scheduling. SIAM Journal on Computing, 29(2): 459–473, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.

org/sam-bin/dbq/article/
32487.

Albers:2009:VCN

[Alb09] Susanne Albers. On the value of coordination in network design. SIAM Journal on Computing, 38(6):2273–2302, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alder:1975:CTH

[Ald75] M. D. Alder. A convergence theorem for hierarchies of model neurones. SIAM Journal on Computing, 4(4): 491–506, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arora:1996:LAP

[ALM96] Sanjeev Arora, F. T. Leighton, and Bruce M. Maggs. Online algorithms for path selection in a nonblocking network. SIAM Journal on Computing, 25(3):600–625, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22149.

Andrews:1999:AMH

[ALMZ99] Matthew Andrews, Tom Leighton, P. Takis Metaxas, and Lisa Zhang. Automatic methods for hiding latency in parallel and distributed computation. SIAM Journal on Computing, 29(2):615–647, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32650.

Arora:2012:LVG

[ALN^{+}12] Sanjeev Arora, László Lovász, Ilan Newman, Yuval Rabani, Yuri Rabinovich, and Santosh Vempala. Local versus global properties of metric spaces. SIAM Journal on Computing, 41(1):250–271, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ar:1998:RAF

[ALRS98] Sigal Ar, Richard J. Lipton, Ronitt Rubinfeld, and Madhu Sudan. Reconstructing algebraic functions from mixed data. SIAM Journal on Computing, 28(2):487–510, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29757.

Asharov:2022:SSK

[ALS22] Gilad Asharov, Wei-Kai Lin, and Elaine Shi. Sorting short keys in circuits of size o(n log n). SIAM Journal on Computing, 51(3):424–466, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1380983.

Archer:2008:FBA

[ALW08] Aaron Archer, Asaf Levin, and David P. Williamson. A faster, better approximation algorithm for the minimum latency problem. SIAM Journal on Computing, 37(5):1472– 1498, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ashcroft:1975:TPS

[AM75] Edward Ashcroft and Zohar Manna. Translating program schemas to while-schemas. SIAM Journal on Computing, 4(2):125–146, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Apt:1980:CFS

[AM80] Krzysztof R. Apt and Lambert G. L. T. Meertens. Completeness with finite systems of intermediate assertions for recursive program schemes. SIAM Journal on Computing, 9(4): 665–671, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alt:1985:SST

[AM85] Helmut Alt and Kurt Mehlhorn.

Searching semisorted tables. SIAM Journal on Computing, 14(4):840–848, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:1993:RSP

[AM93] Pankaj K. Agarwal and Jiří Matoušek. Ray shooting and parametric search. SIAM Journal on Computing, 22(4): 794–806, August 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ajtai:1996:DTP

[AM96] Miklos Ajtai and Nimrod Megiddo. A deterministic Poly(log log N)-time Nprocessor algorithm for linear programming in fixed dimension. SIAM Journal on Computing, 25(6):1171–1195, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23485.

Althaus:2001:TSB

[AM01] Ernst Althaus and Kurt Mehlhorn. Traveling salesmanbased curve reconstruction in polynomial time. SIAM Journal on Computing, 31(1):27– 66, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36611.

Amano:2004:PAM

[AM04] Kazuyuki Amano and Akira Maruoka. The potential of the approximation method.

SIAM Journal on Computing, 33(2):433–447, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38445.

Amano:2005:SLB

[AM05] Kazuyuki Amano and Akira Maruoka. A superpolynomial lower bound for a circuit computing the clique function with at most (1/6)log log n negation gates. SIAM Journal on Computing, 35(1):201–216, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Achlioptas:2006:RST

[AM06] Dimitris Achlioptas and Cristopher Moore. Random k-SAT: Two moments suffice to cross a sharp threshold. SIAM Journal on Computing, 36(3):740– 762, ???? 2006. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Atserias:2013:SAR

[AM13] Albert Atserias and Elitza Maneva. Sherali–Adams relaxations and indistinguishability in counting logics. SIAM Journal on Computing, 42(1):112– 137, ???? 2013. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ambainis:2007:QWA

[Amb07] Andris Ambainis. Quantum walk algorithm for element distinctness. SIAM Journal on Computing, 37(1):210–239, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ambainis:2016:SAE

[Amb16] Andris Ambainis. Superlinear advantage for exact quantum algorithms. SIAM Journal on Computing, 45(2):617–631, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arya:2007:OEC

[AMMW07] Sunil Arya, Theocharis Malamatos, David M. Mount, and Ka Chun Wong. Optimal expected-case planar point location. SIAM Journal on Computing, 37(2):584–610, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Augustine:2022:BAP

[AMRU22] John Augustine, William K. Moses, Amanda Redlich, and Eli Upfal. Balanced allocation: Patience is not a virtue. SIAM Journal on Computing, 51(6):1743–1768, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/17M1155375.

Agarwal:1984:COC

[AMS84] Sunita Agarwal, A. K. Mittal, and P. Sharma. Constrained optimum communication trees and sensitivity analysis. SIAM Journal on Computing, 13(2): 315–328, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:1998:CMF

[AMS98] Pankaj K. Agarwal, Jiří Matousek, and Otfried Schwarzkopf. Computing many faces in arrangements of lines and segments. SIAM Journal on Computing, 27(2):491–505, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26616.

Ambuhl:2011:IRM

[AMS11] Christoph Ambühl, Monaldo Mastrolilli, and Ola Svensson. Inapproximability results for maximum edge biclique, minimum linear arrangement, and sparsest cut. SIAM Journal on Computing, 40(2):567–596, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i2/p567_s1.

Agarwal:2013:RSS

[AMS13] Pankaj K. Agarwal, Jirí Matousek, and Micha Sharir. On range searching with semialgebraic sets. II. SIAM Journal on Computing, 42(6):2039– 2062, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Asano:2007:ZDE

[AMT07] Tetsuo Asano, JiřÍ Matoušek, and Takeshi Tokuyama. Zone diagrams: Existence, uniqueness, and algorithmic challenge. SIAM Journal on Computing, 37(4):1182–1198, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:1993:CFG

[AN93] Noga Alon and Moni Naor. Coin-flipping games immune against linear-sized coalitions. SIAM Journal on Computing, 22(2):403–417, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:2006:ACN

[AN06] Noga Alon and Assaf Naor. Approximating the cut-norm via Grothendieck’s inequality. SIAM Journal on Computing, 35(4):787–803, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Abraham:2019:UPD

[AN19] Ittai Abraham and Ofer Neiman. Using petal-decompositions to build a low stretch spanning tree. SIAM Journal on Computing, 48(2):227–248, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Anderson:1999:TDS

[And99] Richard J. Anderson. Tree data structures for N-body simulation. SIAM Journal on Computing, 28(6):1923–1940, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32630.

Ahmadian:2020:BGM

[ANFSW20] Sara Ahmadian, Ashkan Norouzi-Fard, Ola Svensson, and Justin Ward. Better guarantees for k-means and Euclidean kmedian by primal-dual algorithms. SIAM Journal on Computing, 49(4):FOCS17– 97–FOCS17–156, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Allenberg-Navony:1996:ARC

[ANIM96] Nechama Allenberg-Navony, Alon Itai, and Shlomo Moran. Average and randomized complexity of distributed problems. SIAM Journal on Computing, 25(6):1254–1267, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
25243.

Annamalai:2019:LLS

[Ann19]  Chidambaram  Annamalai.

Lazy local search meets machine scheduling. SIAM Journal on Computing, 48(5):1503– 1543, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Anonymous:1995:MEL

[Ano95] Anonymous. In memoriam: Eugene L. Lawler. SIAM Journal on Computing, 24(1):1–2, February 1995. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Andrew:2009:SSF

[ANR09] Matthew Andrew, Ashwin Nayak, and Rajmohan Rajaraman. Special section on foundations of computer science. SIAM Journal on Computing, 39(2):545, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Asharov:2021:SSE

[ANSS21] Gilad Asharov, Moni Naor, Gil Segev, and Ido Shahaf. Searchable symmetric encryption: Optimal locality in linear space via two-dimensional balanced allocations. SIAM Journal on Computing, 50(5): 1501–1536, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Andersson:1995:NTB

[AO95] Arne Andersson and Thomas Ottmann. New tight bounds on uniquely represented dictionaries. SIAM Journal on Computing, 24(5):1091–1103, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24110.

Andoni:2012:AED

[AO12] Alexandr Andoni and Krzysztof Onak. Approximating edit distance in near-linear time. SIAM Journal on Computing, 41(6):1635–1648, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:2006:CMS

[AOS06] Pankaj K. Agarwal, Mark Overmars, and Micha Sharir. Computing maximally separated sets in the plane. SIAM Journal on Computing, 36(3): 815–834, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ahuja:1994:IAB

[AOST94] Ravindra K. Ahuja, James B.

Orlin, Clifford Stein, and Robert E. Tarjan. Improved algorithms for bipartite network flow. SIAM Journal on Computing, 23(5):906–933, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19933.

Ahuja:1989:ITB

[AOT89] Ravindra K. Ahuja, James B.

Orlin, and Robert E. Tarjan.

Improved time bounds for the maximum flow problem. SIAM Journal on Computing, 18(5): 939–954, October 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aho:1972:MDE

[AP72] Alfred V. Aho and Thomas G. Peterson. A minimum distance error-correcting parser for context-free languages. SIAM Journal on Computing, 1(4):305–312, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agmon:2006:FTG

[AP06] Noa Agmon and David Peleg. Fault-tolerant gathering algorithms for autonomous mobile robots. SIAM Journal on Computing, 36(1):56–82, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alekseyev:2007:WGD

[AP07] Max A. Alekseyev and Pavel A. Pevzner. Whole genome duplications and contracted breakpoint graphs. SIAM Journal on Computing, 36(6):1748– 1763, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Applebaum:2013:PGL

[App13] Benny Applebaum. Pseudorandom generators with long stretch and low locality from random local one-way functions. SIAM Journal on Computing, 42(5):2008–2037, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Adams:1974:NHU

[APS74] J. M. Adams, J. M. Phelan, and R. H. Stark. A note on the hecht-ullman characterization of nonreducible flow graphs. SIAM Journal on Computing, 3(3):222–223, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:1993:CCA

[APS93] Pankaj K. Agarwal, Marco Pellegrini, and Micha Sharir. Counting circular arc intersections. SIAM Journal on Computing, 22(4):778–793, August 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Allender:1988:PS

[AR88] Eric W. Allender and Roy S.

Rubinstein. P-printable sets. SIAM Journal on Computing, 17(6):1193–1202, December 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Attiya:1998:ASO

[AR98a] Hagit Attiya and Ophir Rachman. Atomic snapshots in O(n log n) operations. SIAM Journal on Computing, 27(2): 319–340, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
27946.

Aumann:1998:AMC

[AR98b] Yonatan Aumann and Yuval Rabani. An O(log k) approximate Min-Cut Max-Flow theorem and approximation algorithm. SIAM Journal on Computing, 27(1):291–301, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28598.

Attiya:2002:CSW

[AR02] Hagit Attiya and Sergio Rajsbaum. The combinatorial structure of wait-free solvable tasks. SIAM Journal on Computing, 31(4):1286–1313, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33068.

Alekhnovich:2008:RAU

[AR08] Michael Alekhnovich and Alexander A. Razborov. Resolution is not automatizable unless W [P] is tractable. SIAM Journal on Computing, 38(4): 1347–1363, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arazi:1986:BSP

[Ara86] Benjamin Arazi. A binary search with a parallel recovery of the bits. SIAM Journal on Computing, 15(3):851–855, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alonso:1997:ACC

[ARS97] Laurent Alonso, Edward M.

Reingold, and René Schott. The average-case complexity of determining the majority. SIAM Journal on Computing, 26(1):1–14, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27591.

Achlioptas:2009:RFF

[ART09] Dimitris Achlioptas and Federico Ricci-Tersenghi. Random formulas have frozen variables. SIAM Journal on Computing, 39(1):260–280, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aspvall:1980:PTA

[AS80] Bengt Aspvall and Yossi Shiloach. A polynomial time algorithm for solving systems of linear inequalities with two variables per inequality. SIAM Journal on Computing, 9(4): 827–845, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ambos-Spies:1986:ISK

[AS86] Klaus Ambos-Spies. An inhomogeneity in the structure of Karp degrees. SIAM Journal on Computing, 15(4):958–963, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:1990:RBI

[AS90] Pankaj K. Agarwal and Micha Sharir. Red-blue intersection detection algorithms, with applications to motion planning and collision detection. SIAM Journal on Computing, 19(2): 297–321, April 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agarwal:1996:RSA

[AS96] Pankaj K. Agarwal and Micha Sharir. Ray shooting amidst convex polyhedra and polyhedral terrains in three dimensions. SIAM Journal on Computing, 25(1):100–116, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24436.

Aronov:1997:TMP

[AS97] Boris Aronov and Micha Sharir. On translational motion planning of a convex polyhedron in 3-space. SIAM Journal on Computing, 26(6): 1785–1803, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26660.

Agarwal:1998:SAG

[AS98] Pankaj K. Agarwal and Subhash Suri. Surface approximation and geometric partitions. SIAM Journal on Computing, 27(4):1016–1035, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26980.

Agarwal:2005:PLA

[AS05a] Pankaj K. Agarwal and Micha Sharir. Pseudo-line arrangements: Duality, algorithms, and applications. SIAM Journal on Computing, 34(3):526– 552, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43390.

Albers:2005:PGA

[AS05b] Susanne Albers and Markus Schmidt. On the performance of greedy algorithms in packet buffering. SIAM Journal on Computing, 35(2):278– 304, ???? 2005. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:2008:CNG

[AS08a] Noga Alon and Asaf Shapira.

A characterization of the (natural) graph properties testable with one-sided error. SIAM Journal on Computing, 37(6): 1703–1727, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Alon:2008:EMG

[AS08b] Noga Alon and Asaf Shapira. Every monotone graph property is testable. SIAM Journal on Computing, 38(2):505–522, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aronov:2010:AHR

[AS10a] Boris Aronov and Micha Sharir. Approximate halfspace range counting. SIAM Journal on Computing, 39(7): 2704–2725, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Asadpour:2010:AAM

[AS10b] Arash Asadpour and Amin Saberi. An approximation algorithm for max-min fair allocation of indivisible goods. SIAM Journal on Computing, 39(7):2970–2989, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Asharov:2016:LPI

[AS16] Gilad Asharov and Gil Segev. Limits on the power of indistinguishability obfuscation and functional encryption. SIAM Journal on Computing, 45(6):

2117–2176, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Asano:1987:ADE

[Asa87] Takao Asano. An application of duality to edge-deletion problems. SIAM Journal on Computing, 16(2):312–331, April 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

An:2017:LBA

[ASS17] Hyung-Chan An, Mohit Singh, and Ola Svensson. LP-based algorithms for capacitated facility location. SIAM Journal on Computing, 46(1):272–306, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Agrawal:2016:JHC

[ASSS16] Manindra Agrawal, Chandan Saha, Ramprasad Saptharishi, and Nitin Saxena. Jacobian hits circuits: Hitting sets, lower bounds for depth-D occur-k formulas and depth-3 transcendence degree-k circuits. SIAM Journal on Computing, 45(4):1533–1562, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aho:1981:ITL

[ASSU81] A. V. Aho, Y. Sagiv, T. G. Szymanski, and J. D. Ullman. Inferring a tree from lowest common ancestors with an application to the optimization of relational expressions. SIAM Journal on Computing, 10(3):405– 421, ???? 1981. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Asano:1976:GRT

[AST76] Takao Asano, Michiro Shibui, and Itsuo Takanami. General results on tour lengths in machines and digraphs. SIAM Journal on Computing, 5(4): 629–645, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aronov:1997:UCP

[AST97] Boris Aronov, Micha Sharir, and Boaz Tagansky. The union of convex polyhedra in three dimensions. SIAM Journal on Computing, 26(6): 1670–1688, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25075.

Attiya:1999:LLR

[AST99] Hagit Attiya, Hadas Shachnai, and Tami Tamir. Local labeling and resource allocation using preprocessing. SIAM Journal on Computing, 28 (4):1397–1414, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28564.

Aho:1975:EPF

[ASU75] A. V. Aho, K. Steiglitz, and J. D. Ullman. Evaluating polynomials at fixed sets of points. SIAM Journal on Computing, 4(4):533–539, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aho:1979:EAR

[ASU79] A. V. Aho, Y. Sagiv, and J. D.

Ullman. Equivalences among relational expressions. SIAM Journal on Computing, 8(2): 218–246, May 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Adolphson:1977:LTA

[AT77] D. L. Adolphson and G. N.

Thomas. A linear time algorithm for a 2 × n transportation problem. SIAM Journal on Computing, 6(3):481–486, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aguilera:1998:FDR

[AT98] Marcos Kawazoe Aguilera and Sam Toueg. Failure detection and randomization: a hybrid approach to solve consensus. SIAM Journal on Computing, 28(3):890–903, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31291.

Agrawal:2000:FIP

[AT00] Manindra Agrawal and Thomas Thierauf. The formula isomorphism problem. SIAM Journal on Computing, 30(3):990– 1009, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        34364.

Atallah:1982:FCI

[Ata82] Mikhail J. Atallah. Finding the cyclic index of an irreducible, nonnegative matrix. SIAM Journal on Computing, 11(3): 567–570, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aharonov:2007:AQS

[ATS07] Dorit Aharonov and Amnon Ta-Shma. Adiabatic quantum state generation. SIAM Journal on Computing, 37(1):47– 82, ???? 2007. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aho:1972:OSL

[AU72] Alfred V. Aho and Jeffrey D. Ullman. Optimization of straight line programs. SIAM Journal on Computing, 1(1):1– 19, ???? 1972. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aho:1973:TSP

[AU73] A. V. Aho and J. D. Ullman. A technique for speeding up lr(k) parsers. SIAM Journal on Computing, 2(2):106–127, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aurenhammer:1987:PDP

[Aur87] F. Aurenhammer. Power diagrams: Properties, algorithms, and applications. SIAM Journal on Computing, 16(1):78– 96, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Austrin:2010:TSI

[Aus10] Per Austrin. Towards sharp inapproximability for any 2-CSP. SIAM Journal on Computing, 39(6):2430–2463, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Azar:1987:TCB

[AV87] Yossi Azar and Uzi Vishkin.

Tight comparison bounds on the complexity of parallel sorting. SIAM Journal on Computing, 16(3):458–464, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Arge:2003:OEM

[AV03] Lars Arge and Jeffrey Scott Vitter. Optimal external memory interval management. SIAM Journal on Computing, 32(6):1488–1508, December 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40481.

Arthur:2009:WCS

[AV09] David Arthur and Sergei Vassilvitskii. Worst-case and smoothed analysis of the ICP algorithm, with an application to the k-means method. SIAM Journal on Computing, 39(2): 766–782, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aharonov:2007:AQC

[AvDK^{+}07] Dorit Aharonov, Wim van Dam, Julia Kempe, Zeph Landau, Seth Lloyd, and Oded Regev. Adiabatic quantum computation is equivalent to standard quantum computation. SIAM Journal on Computing, 37(1):166–194, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Anderson:2012:LCL

[AvMSS12] Matthew Anderson, Dieter van Melkebeek, Nicole Schweikardt, and Luc Segoufin. Locality from circuit lower bounds. SIAM Journal on Computing, 41(6):1481–1523, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ashcroft:1976:LFS

[AW76] E. A. Ashcroft and W. W. Wadge. Lucid — a formal system for writing and proving programs. SIAM Journal on Computing, 5(3):336– 354, September 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Aspnes:1996:RCE

[AW96] James Aspnes and Orli Waarts.

Randomized consensus in expected O(n log^{2} n) operations per processor. SIAM Journal on Computing, 25(5):1024– 1044, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
24088.

Anderson:1997:ACW

[AW97] Richard J. Anderson and Heather Woll. Algorithms for the Certified Write-All Problem. SIAM Journal on Computing, 26(5):1277–1283, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31912.

Agarwal:2001:AAC

[AW01] Pankaj K. Agarwal and Hongyan Wang. Approximation algorithms for curvatureconstrained shortest paths. SIAM Journal on Computing, 30(6):1739–1772, December 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30779.

Alman:2021:LAK

[AW21] Josh Alman and Virginia Vassilevska Williams. Limits on all known (and some unknown) approaches to matrix multiplication. SIAM Journal on Computing, 52(6):FOCS18– 285–FOCS18–315, November 2021. CODEN SMJCAT. ISSN 1095-7111.

Abboud:2018:MTB

[AWY18] Amir Abboud, Virginia Vassilevska Williams, and Huacheng Yu. Matching triangles and basing hardness on an extremely popular conjecture. SIAM Journal on Computing, 47(3):1098–1122, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Angluin:2000:RND

[AWZ00] Dana Angluin, Jeffery Westbrook, and Wenhong Zhu. Robot navigation with distance queries. SIAM Journal on Computing, 30(1):110– 144, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        33005.

Andrews:2004:ETS

[AZ04] Matthew Andrews and Lisa Zhang. The effects of temporary sessions on network performance. SIAM Journal on Computing, 33(3):659–673, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41722.

Andrews:2007:HUC

[AZ07] Matthew Andrews and Lisa Zhang. Hardness of the undirected congestion minimization problem. SIAM Journal on Computing, 37(1):112–131, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Azar:1992:LBT

[Aza92] Yossi Azar. Lower bounds for threshold and symmetric functions in parallel computation. SIAM Journal on Computing, 21(2):329–338, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Byrka:2010:OBA

[BA10] Jaroslaw Byrka and Karen Aardal. An optimal bifactor approximation algorithm for the metric uncapacitated facility location problem. SIAM Journal on Computing, 39(6): 2212–2231, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Babai:1980:CCL

[Bab80] László Babai. On the complexity of canonical labeling of strongly regular graphs. SIAM Journal on Computing, 9(1): 212–216, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Babai:2006:SID

[Bab06] László Babai. Special issue dedicated to the Thirty-Sixth Annual ACM Symposium on Theory of Computing (STOC 2004). SIAM Journal on Computing, 35(4):ix–x, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bach:1988:HGF

[Bac88] Eric Bach. How to generate factored random numbers. SIAM Journal on Computing, 17(2):179–193, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Backens:2021:FDI

[Bac21] Miriam Backens. A full dichotomy for hol^{c}, inspired by quantum computation. SIAM Journal on Computing, 50(6): 1739–1799, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Aroya:2022:ERT

[BADTS22] Avraham Ben-Aroya, Dean Doron, and Amnon Ta-Shma. An efficient reduction from two-source to nonmalleable extractors: Achieving near-logarithmic minentropy. SIAM Journal on Computing, 51(2):STOC17– 31–STOC17–49, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1133245.

Ben-Asher:1999:OST

[BAFN99] Yosi Ben-Asher, Eitan Farchi, and Ilan Newman. Optimal search in trees. SIAM Journal on Computing, 28(6): 2090–2102, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31858.

Ben-Amram:2001:TLB

[BAG01] Amir M. Ben-Amram and Zvi Galil. Topological lower bounds on algebraic random access machines. SIAM Journal on Computing, 31(3):722– 761, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32939.

Baker:1978:GSD

[Bak78a] Brenda S. Baker. Generalized syntax directed translation, tree transducers, and linear space. SIAM Journal on Computing, 7(3):376–391, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baker:1978:TER

[Bak78b] Theodore P. Baker. A technique for extending rapid exact-match string matching to arrays of more than one dimension. SIAM Journal on Computing, 7(4):533–541, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baker:1986:PGA

[Bak86] Brenda S. Baker. A provably good algorithm for the two module routing problem. SIAM Journal on Computing, 15(1):162–188, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baker:1997:PDS

[Bak97] Brenda S. Baker. Parameterized duplication in strings: Algorithms and an application to software maintenance. SIAM Journal on Computing, 26(5):1343–1362, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24670.

Balcazar:1985:SRN

[Bal85] José L. Balcázar. Simplicity, relativizations and nondeterminism. SIAM Journal on Computing, 14(1):148–157, ???? 1985. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Barz:1984:PSM

[Bar84] H. W. Barz. The power of synchronization mechanisms. SIAM Journal on Computing, 13(4):726–749, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barman:2018:ANE

[Bar18]  Siddharth Barman.  Ap-

proximating Nash equilibria and dense subgraphs via an approximate version of Carathéodory’s Theorem. SIAM Journal on Computing, 47(3):960–981, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bass:1972:NIC

[Bas72] Leonard J. Bass. A note on the intersection of complexity classes of functions. SIAM Journal on Computing, 1(4): 288–289, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bass:1973:OPS

[Bas73] L. J. Bass. On optimal processor scheduling for multiprogramming. SIAM Journal on Computing, 2(4):273–280, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [Bas76].

Bass:1976:EOP

[Bas76] L. J. Bass. Errata: On optimal processor scheduling for multiprogramming. SIAM Journal on Computing, 5(4):758, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Bas73].

Ben-Aroya:2011:CCA

[BATS11] Avraham Ben-Aroya and Amnon Ta-Shma. A combinatorial construction of almost-Ramanujan graphs using the zig-zag product. SIAM Journal on Computing, 40(2):267–290, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i2/
                                        p267_s1.

Bazzi:2009:PIC

[Baz09] Louay M. J. Bazzi. Polylogarithmic independence can fool DNF formulas. SIAM Journal on Computing, 38(6):2220– 2272, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1980:ESC

[BB80] Ronald V. Book and Franz-Josef Brandenburg. Equality sets and complexity classes. SIAM Journal on Computing, 9(4):729–743, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bollobas:1988:TOG

[BB88] Béla Bollobás and Graham Brightwell. Transitive orientations of graphs. SIAM Journal on Computing, 17(6):1119– 1133, December 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bonet:1995:STC

[BB95] Maria Luisa Bonet and Samuel R.

Buss. The serial transitive closure problem for trees. SIAM Journal on Computing, 24(1):109–122, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22530.

Belovs:2021:PLB

[BB21] Aleksandrs Belovs and Eric Blais. A polynomial lower bound for testing monotonicity. SIAM Journal on Computing, 50(3):STOC16–406– STOC16–433, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bennett:1997:SWQ

[BBBV97] Charles H. Bennett, Ethan Bernstein, Gilles Brassard, and Umesh Vazirani. Strengths and weaknesses of quantum computing. SIAM Journal on Computing, 26(5):1510– 1523, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
30093.

Berczi:2022:AMR

[BBC^{+}22] Kristóf Bérczi, Endre Boros, Ondrej Cepek, Petr Kucera, and Kazuhisa Makino. Approximating minimum representations of key Horn functions. SIAM Journal on Computing, 51(1):116–138, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1275681.

Barak:2013:HCI

[BBCR13] Boaz Barak, Mark Braverman, Xi Chen, and Anup Rao. How to compress interactive communication. SIAM Journal on Computing, 42(3):1327–1363, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Berman:1986:CFP

[BBD^{+}86] Francine Berman, Mary Ellen Bock, Eric Dittert, Michael J. O’Donnell, and Darrell Plank. Collections of functions for perfect hashing. SIAM Journal on Computing, 15(2):604– 618, ???? 1986. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barenco:1997:SQC

[BBD^{+}97] Adriano Barenco, Andre Berthiaume, David Deutsch, Artur Ekert, Richard Jozsa, and Chiara Macchiavello. Stabilization of quantum computations by symmetrization. SIAM Journal on Computing, 26(5):1541–1557, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30245.

Bansal:2023:MSV

[BBFT23] Nikhil Bansal, Jatin Batra, Majid Farhadi, and Prasad Tetali. On min sum vertex cover and generalized min sum set cover. SIAM Journal on Computing, 52(2):327–357, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1434052.

Balogh:2008:LBO

[BBGR08] János Balogh, József Békési, Gábor Galambos, and Gerhard Reinelt. Lower bound for the online bin packing problem with restricted repacking. SIAM Journal on Computing, 38(1):398–410, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beame:2016:TST

[BBI16] Paul Beame, Chris Beck, and Russell Impagliazzo. Timespace trade-offs in resolution: Superpolynomial lower bounds for superlinear space. SIAM Journal on Computing, 45(4):

1612–1645, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Balcan:2013:CPA

[BBM13] Maria-Florina Balcan, Avrim Blum, and Yishay Mansour. Circumventing the price of anarchy: Leading dynamics to good behavior. SIAM Journal on Computing, 42(1):230–264, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bansal:2012:RCA

[BBN12] Nikhil Bansal, Niv Buchbinder, and Joseph (Seffi) Naor. Randomized competitive algorithms for generalized caching. SIAM Journal on Computing, 41(2):391–414, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Balliu:2022:DLB

[BBO22] Alkida Balliu, Sebastian Brandt, and Dennis Olivetti. Distributed lower bounds for ruling sets. SIAM Journal on Computing, 51(1):70–115, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1381770.

Bassino:2000:FSV

[BBP00] Frédérique Bassino, Marie-Pierre Béal, and Dominique Perrin. A finite state version of the Kraft–McMillan theorem. SIAM Journal on Computing, 30(4):1211–1230, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34390.

Bennett:1988:PAP

[BBR88] Charles H. Bennett, Gilles Brassard, and Jean-Marc Robert. Privacy amplification by public discussion. SIAM Journal on Computing, 17(2): 210–229, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Beame:1999:TST

[BBR^{+}99] Paul Beame, Allan Borodin, Prabhakar Raghavan, Walter L. Ruzzo, and Martin Tompa. A time-space tradeoff for undirected graph traversal by walking automata. SIAM Journal on Computing, 28(3):1051–1072, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28294.

Bartal:2004:FDA

[BBR04] Yair Bartal, John W. Byers, and Danny Raz. Fast, distributed approximation algorithms for positive linear programming with applications to flow control. SIAM Journal on Computing, 33(6):

1261–1279, December 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37938.

Barnes:1998:SSP

[BBRS98] Greg Barnes, Jonathan F. Buss, Walter L. Ruzzo, and Baruch Schieber. A sublinear space, polynomial time algorithm for directed s-t connectivity. SIAM Journal on Computing, 27(5):1273–1282, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28315.

Balcazar:1986:SSL

[BBS86a] José L. Balcázar, Ronald V. Book, and Uwe Schöning. Sparse sets, lowness and highness. SIAM Journal on Computing, 15(3):739–747, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blum:1986:SUP

[BBS86b] L. Blum, M. Blum, and M. Shub. A simple unpredictable pseudo-random number generator. SIAM Journal on Computing, 15(2):364–383, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Balcan:2013:SSF

[BBS13] Maria Florina Balcan, Mark Braverman, and Daniel A. Spielman. Special section on the Fiftieth Annual IEEE Symposium on Foundations of Computer Science (FOCS 2009). SIAM Journal on Computing, 42(6):2286, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Borodin:1976:NAC

[BC76] Allan Borodin and Stephen Cook. On the number of additions to compute specific polynomials. SIAM Journal on Computing, 5(1):146–157, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Borodin:1982:TST

[BC82] A. Borodin and S. Cook. A time-space tradeoff for sorting on a general sequential model of computation. SIAM Journal on Computing, 11(2):287– 297, ???? 1982. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baker:1984:ICA

[BC84] B. S. Baker and E. G. Coffman, Jr. Insertion and compaction algorithms in sequentially allocated storage. SIAM Journal on Computing, 13(3):600–609, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bini:1987:TRB

[BC87] D. Bini and M. Capovani. Tensor rank and border rank of band Toeplitz matrices. SIAM Journal on Computing, 16(2): 252–258, April 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baum:1991:SLU

[BC91] Ulrich Baum and Michael Clausen. Some lower and upper complexity bounds for generalized Fourier transforms and their inverses. SIAM Journal on Computing, 20(3):451– 459, June 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bshouty:1998:IAR

[BC98] Nader H. Bshouty and Richard Cleve. Interpolating arithmetic read-once formulas in parallel. SIAM Journal on Computing, 27(2):401–413, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28812.

Blum:2000:OAI

[BC00] Avrim Blum and Prasad Chalasani. An online algorithm for improving performance in navigation. SIAM Journal on Computing, 29(6): 1907–1938, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29059.

Burgisser:2003:CCC

[BC03] Peter Bürgisser and Felipe Cucker. Counting complexity classes for numeric computations I: Semilinear sets. SIAM Journal on Computing, 33(1):227–260, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41595.

Bodirsky:2010:QEC

[BC10] Manuel Bodirsky and Hubie Chen. Quantified equality constraints. SIAM Journal on Computing, 39(8):3682–3699, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bhandari:2022:IBP

[BC22] Siddharth Bhandari and Sayantan Chakraborty. Improved bounds for perfect sampling of k-colorings in graphs. SIAM Journal on Computing, 51 (3):STOC20–54–STOC20–74, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1366666.

Boyle:2023:SSF

[BCAKT23] Elette Boyle, Vincent Cohen-Addad, Alexandra Kolla, and Mikkel Thorup. Special section on the Fifty-Ninth Annual IEEE Symposium on Foundations of Computer Science (2018). SIAM Journal on Computing, 52(6):FOCS18–i– FOCS18–i, December 2023. CODEN SMJCAT. ISSN 1095-7111.

Bose:2009:SCP

[BCC^{+}09] Prosenjit Bose, Paz Carmi, Mathieu Couture, Anil Maheshwari, Pat Morin, and Michiel Smid. Spanners of complete k-partite geometric graphs. SIAM Journal on Computing, 38(5):1803–1820, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Butman:2013:PMU

[BCC^{+}13] Ayelet Butman, Peter Clifford, Raphaël Clifford, Markus Jalsenius, Noa Lewenstein, Benny Porat, Ely Porat, and Benjamin Sach. Pattern matching under polynomial transformation. SIAM Journal on Computing, 42(2):611–633, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baswana:2019:DDU

[BCCK19] Surender Baswana, Shreejit Ray Chaudhury, Keerti Choudhary, and Shahbaz Khan. Dynamic DFS in undirected graphs: Breaking the O(m) barrier. SIAM Journal on Computing, 48(4):1335– 1363, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Borodin:1989:ETA

[BCD^{+}89a] Allan Borodin, Stephen A.

Cook, Patrick W. Dymond, Walter L. Ruzzo, and Martin Tompa. Erratum: “Two applications of inductive counting for complementation problems”. SIAM Journal on Computing, 18(6):1283, December 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [BCD^{+}89b].

Borodin:1989:TAI

[BCD^{+}89b] Allan Borodin, Stephen A.

Cook, Patrick W. Dymond, Walter L. Ruzzo, and Martin Tompa. Two applications of inductive counting for complementation problems. SIAM Journal on Computing, 18(3): 559–578, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [BCD^{+}89a].

Bshouty:1995:SDT

[BCE95] Nader H. Bshouty, Richard Cleve, and Wayne Eberly. Size-depth tradeoffs for algebraic formulas. SIAM Journal on Computing, 24(4):682– 705, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23258.

Buhrman:2014:PBQ

[BCF^{+}14] Harry Buhrman, Nishanth Chandran, Serge Fehr, Ran Gelles, Vipul Goyal, Rafail Ostrovsky, and Christian Schaffner. Position-based quantum cryptography: Impossibility and constructions. SIAM Journal on Computing, 43(1):150–178, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bitansky:2018:IOR

[BCG^{+}18] Nir Bitansky, Ran Canetti, Sanjam Garg, Justin Holmgren, Abhishek Jain, Huijia Lin, Rafael Pass, Sidharth Telang, and Vinod Vaikuntanathan. Indistinguishability obfuscation for RAM programs and succinct randomized encodings. SIAM Journal on Computing, 47(3):1123– 1210, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Braverman:2020:PPD

[BCG20] Mark Braverman, Gil Cohen, and Sumegha Garg. Pseudorandom pseudo-distributions with near-optimal error for read-once branching programs. SIAM Journal on Computing, 49(5):STOC18–242–STOC18– 299, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bei:2021:ISV

[BCG^{+}21] Xiaohui Bei, Shiteng Chen, Ji Guan, Youming Qiao, and Xiaoming Sun. From independent sets and vertex colorings to isotropic spaces and isotropic decompositions: Another bridge between graphs and alternating matrix spaces. SIAM Journal on Computing, 50(3):924–971, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bei:2017:WCM

[BCGL17] Xiaohui Bei, Ning Chen, Nick Gravin, and Pinyan Lu. Worstcase mechanism design via Bayesian analysis. SIAM Journal on Computing, 46(4):1428– 1448, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buss:1992:OPA

[BCGR92] S. Buss, S. Cook, A. Gupta, and V. Ramachandran. An optimal parallel algorithm for formula evaluation. SIAM Journal on Computing, 21(4):755– 780, August 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bajaj:1993:FRP

[BCGW93] Chanderjit Bajaj, John Canny, Thomas Garrity, and Joe Warren. Factoring rational polynomials over the complex numbers. SIAM Journal on Computing, 22(2):318–331, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beame:1986:LDC

[BCH86] Paul W. Beame, Stephen A. Cook, and H. James Hoover. Log depth circuits for division and related problems. SIAM Journal on Computing, 15(4): 994–1003, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bruck:1997:FTM

[BCH97] Jehoshua Bruck, Robert Cypher, and Ching-Tien Ho. Faulttolerant meshes with small degree. SIAM Journal on Computing, 26(6):1764–1784, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        27499.

Bouland:2020:PSZ

[BCH^{+}20] Adam Bouland, Lijie Chen, Dhiraj Holden, Justin Thaler, and Prashant Nalini Vasudevan. On the power of statistical zero knowledge. SIAM Journal on Computing, 49(4):FOCS17– 1–FOCS17–58, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boneh:2007:CCS

[BCHK07] Dan Boneh, Ran Canetti, Shai Halevi, and Jonathan Katz. Chosen-ciphertext security from identity-based encryption. SIAM Journal on Computing, 36(5):1301–1328, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boros:1994:CIS

[BCHS94] E. Boros, Y. Crama, P. L.

Hammer, and M. Saks. A complexity index for satisfiability problems. SIAM Journal on Computing, 23(1):45– 49, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
22862.

Blum:2007:AAO

[BCK^{+}07] Avrim Blum, Shuchi Chawla, David R. Karger, Terran Lane, Adam Meyerson, and Maria Minkoff. Approximation algorithms for orienteering and discounted-reward TSP. SIAM Journal on Computing, 37(2): 653–670, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bumby:1975:ICH

[BCL75] R. Bumby, E. Cooper, and D. Latch. Interactive computation of homology of finite partially ordered sets. SIAM Journal on Computing, 4(3):321– 325, ???? 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bubeck:2021:MTS

[BCLL21] Sébastien Bubeck, Michael B. Cohen, James R. Lee, and Yin Tat Lee. Metrical task systems on trees via mirror descent and unfair gluing. SIAM Journal on Computing, 50(3): 909–923, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bhatt:1992:EET

[BCLR92] Sandeep N. Bhatt, Fan R. K. Chung, F. Thomson Leighton, and Arnold L. Rosenberg. Efficient embeddings of trees in hypercubes. SIAM Journal on Computing, 21(1):151–162, February 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bronnimann:1999:PRS

[BCM99] Hervé Brönnimann, Bernard Chazelle, and Jiri Matousek. Product range spaces, sensitive sampling, and derandomization. SIAM Journal on Computing, 28(5):1552–1575, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26032.

Brunsch:2015:SAS

[BCM^{+}15] Tobias Brunsch, Kamiel Cornelissen, Bodo Manthey, Heiko Röglin, and Clemens Rösner. Smoothed analysis of the successive shortest path algorithm. SIAM Journal on Computing, 44(6):1798–1819, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Burgisser:2017:MMP

[BCMW17] Peter Bürgisser, Matthias Christandl, Ketan D. Mulmuley, and Michael Walter. Membership in moment polytopes is in NP and coNP. SIAM Journal on Computing, 46(3):972– 991, ???? 2017. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boasson:1981:RIC

[BCN81] Luc Boasson, Bruno Courcelle, and Maurice Nivat. The rational index: a complexity measure for languages. SIAM Journal on Computing, 10(2): 284–296, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Becchetti:2020:FYP

[BCN^{+}20] Luca Becchetti, Andrea E.

Clementi, Emanuele Natale, Francesco Pasquale, and Luca Trevisan. Find your place: Simple distributed algorithms for community detection. SIAM Journal on Computing, 49(4):821–864, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beimel:2009:PAS

[BCNW09] Amos Beimel, Paz Carmi, Kobbi Nissim, and Enav Weinreb. Private approximation of search problems. SIAM Journal on Computing, 38(5):1728– 1760, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bitansky:2016:EEO

[BCPR16] Nir Bitansky, Ran Canetti, Omer Paneth, and Alon Rosen. On the existence of extractable one-way functions. SIAM Journal on Computing, 45(5): 1910–1952, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baker:1980:OPT

[BCR80] Brenda S. Baker, E. G. Coffman, Jr., and Ronald L. Rivest. Orthogonal packings in two dimensions. SIAM Journal on Computing, 9(4):846– 855, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bi:2016:SRD

[BCR16] Jingguo Bi, Qi Cheng, and J. Maurice Rojas. Sublinear root detection and new hardness results for sparse polynomials over finite fields. SIAM Journal on Computing, 45(4): 1433–1447, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baswana:2018:FTS

[BCR18] Surender Baswana, Keerti Choudhary, and Liam Roditty. Fault-tolerant subgraph for single-source reachability: General and optimal. SIAM Journal on Computing, 47(1):80– 95, ???? 2018. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bansal:2008:IAA

[BCS08] Nikhil Bansal, Don Coppersmith, and Maxim Sviridenko. Improved approximation algorithms for broadcast scheduling. SIAM Journal on Computing, 38(3):1157–1174, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bansal:2010:NAM

[BCS10] Nikhil Bansal, Alberto Caprara, and Maxim Sviridenko. A new approximation method for set covering problems, with applications to multidimensional bin packing. SIAM Journal on Computing, 39(4):1256–1278, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Berenbrink:2006:BAH

[BCSV06] Petra Berenbrink, Artur Czumaj, Angelika Steger, and Berthold Vöcking. Balanced allocations: The heavily loaded case. SIAM Journal on Computing, 35(6):1350–1385, ???? 2006. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Brockett:1976:NMR

[BD76] Roger W. Brockett and David Dobkin. On the number of multiplications required for matrix multiplication. SIAM Journal on Computing, 5(4): 624–628, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [BD78a].

Brockett:1978:ENM

[BD78a] R. W. Brockett and D. Dobkin. Errata: On the number of multiplications required for matrix multiplication. SIAM Journal on Computing, 7(2):238, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [BD76].

Bruno:1978:CTS

[BD78b] John Bruno and Peter Downey. Complexity of task sequencing with deadlines, setup times and changeover costs. SIAM Journal on Computing, 7(4): 393–404, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barak:1981:DPS

[BD81] Amnon B. Barak and Peter J. Downey. Distributed processor scheduling and user countermeasures. SIAM Journal on Computing, 10(4):809–819, ???? 1981. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Bruno:1986:PBP

[BD86] John L. Bruno and Peter J.

Downey. Probabilistic bounds on the performance of list scheduling. SIAM Journal on Computing, 15(2):409–417, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bajaj:1992:CDP

[BD92] Chanderjit L. Bajaj and Tamal K. Dey. Convex decomposition of polyhedra and robustness. SIAM Journal on Computing, 21(2):339–364, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bienstock:1993:BSC

[BD93] Daniel Bienstock and Nicole Diaz. Blocking small cuts in a network, and related problems. SIAM Journal on Computing, 22(3):482–499, June 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Borchers:1997:SRG

[BD97] Al Borchers and Ding-Zhu Du. The k-Steiner ratio in graphs. SIAM Journal on Computing, 26(3):857–869, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28108.

Bulatov:2006:SAM

[BD06] Andrei Bulatov and Víctor Dalmau. A simple algorithm for Mal’tsev constraints. SIAM Journal on Computing, 36(1): 16–27, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bronnimann:2007:LFL

[BDD^{+}07] Hervé Brönnimann, Olivier Devillers, Vida Dujmović, Hazel Everett, Marc Glisse, Xavier Goaoc, Sylvain Lazard, Hyeon-Suk Na, and Sue Whitesides. Lines and free line segments tangent to arbitrary three-dimensional convex polyhedra. SIAM Journal on Computing, 37(2):522–551, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bodlaender:2016:AAT

[BDD^{+}16] Hans L. Bodlaender, Pål Grønås Drange, Markus S. Dregi, Fedor V. Fomin, Daniel Lokshtanov, and Michal Pilipczuk. A c^{k}n 5-approximation algorithm for treewidth. SIAM Journal on Computing, 45(2):317–378, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bender:2005:COB

[BDFC05] Michael A. Bender, Erik D. Demaine, and Martin Farach-Colton. Cache-oblivious Btrees. SIAM Journal on Computing, 35(2):341–358, ????

2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Borodin:1986:BWT

[BDFP86] Allan Borodin, Danny Dolev, Faith E. Fich, and Wolfgang Paul. Bounds for width two branching programs. SIAM Journal on Computing, 15(2): 549–560, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bansal:2019:AKC

[BDG19] Nikhil Bansal, Daniel Dadush, and Shashwat Garg. An algorithm for Komlós conjecture matching Banaszczyk’s bound. SIAM Journal on Computing, 48(2):534–553, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bubley:1999:ACC

[BDGJ99] Russ Bubley, Martin Dyer, Catherine Greenhill, and Mark Jerrum. On approximately counting colorings of small degree graphs. SIAM Journal on Computing, 29(2):387–400, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33817.

Buhrman:2005:QAE

[BDH^{+}05] Harry Buhrman, Christoph Dürr, Mark Heiligman, Peter Høyer, Frédéric Magniez, Miklos Santha, and Ronald de Wolf. Quantum algorithms for element distinctness. SIAM Journal on Computing, 34(6): 1324–1330, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40278.

Boyce:1985:FEP

[BDIG85] James E. Boyce, David P. Dobkin, Robert L. (Scot) Drysdale III, and Leo J. Guibas. Finding extremal polygons. SIAM Journal on Computing, 14(1):134–147, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brandstadt:2000:LTA

[BDK00] Andreas Brandstädt, Feodor F. Dragan, and Ekkehard Köhler. Linear time algorithms for Hamiltonian problems on (claw,net)-free graphs. SIAM Journal on Computing, 30 (5):1662–1677, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35777.

Batu:2005:CAE

[BDKR05] Tugkan Batu, Sanjoy Dasgupta, Ravi Kumar, and Ronitt Rubinfeld. The complexity of approximating the entropy. SIAM Journal on Computing, 35(1):132–150, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boissonnat:2000:MPL

[BDL00] Jean-Daniel Boissonnat, Olivier Devillers, and Sylvain Lazard. Motion planning of legged robots. SIAM Journal on Computing, 30(1):218–246, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32628.

Bhowmick:2014:NBM

[BDL14] Abhishek Bhowmick, Zeev Dvir, and Shachar Lovett. New bounds for matching vector families. SIAM Journal on Computing, 43(5):1654–1683, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boutsidis:2014:NOC

[BDMI14] Christos Boutsidis, Petros Drineas, and Malik Magdon-Ismail. Near-optimal columnbased matrix reconstruction. SIAM Journal on Computing, 43(2):687–717, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blum:1991:NZK

[BDMP91] Manuel Blum, Alfredo De Santis, Silvio Micali, and Giuseppe Persiano. Noninteractive zeroknowledge. SIAM Journal on Computing, 20(6):1084–1118, December 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bose:2013:RGS

[BDMS13] Prosenjit Bose, Vida Dujmović, Pat Morin, and Michiel Smid. Robust geometric spanners. SIAM Journal on Computing, 42(4):1720–1736, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bertolazzi:1998:OUP

[BDMT98] Paola Bertolazzi, Giuseppe Di Battista, Carlo Mannino, and Roberto Tamassia. Optimal upward planarity testing of single-source digraphs. SIAM Journal on Computing, 27(1):132–169, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27962.

Boreale:2002:PTC

[BDP02] Michele Boreale, Rocco De Nicola, and Rosario Pugliese. Proof techniques for cryptographic processes. SIAM Journal on Computing, 31(3):947– 986, June 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37786.

Bouchard:2023:WGN

[BDP23] Sébastien Bouchard, Yoann Dieudonné, and Andrzej Pelc. Want to gather? No need to chatter! SIAM Journal on Computing, 52(2):358–411, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1362899.

Bousquet:2018:MF

[BDT18] Nicolas Bousquet, Jean Daligault, and Stéphan Thomassé. Multicut is FPT. SIAM Journal on Computing, 47(1):166– 207, ???? 2018. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bitansky:2021:SVH

[BDV21] Nir Bitansky, Akshay Degwekar, and Vinod Vaikuntanathan. Structure versus hardness through the obfuscation lens. SIAM Journal on Computing, 50(1):98–144, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bloom:1988:VIT

[BÉ88] Stephen L. Bloom and Zoltán Ésik. Varieties of iteration theories. SIAM Journal on Computing, 17(5):939–966, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barnes:1998:TSL

[BE98] Greg Barnes and Jeff A. Edmonds. Time-space lower bounds for directed st-connectivity on graph automata models. SIAM Journal on Computing, 27(4):1190–1202, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29440.

Braverman:2017:LUC

[BE17] Mark Braverman and Klim Efremenko. List and unique coding for interactive communication in the presence of adversarial noise. SIAM Journal on Computing, 46(1):388–428, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beame:1991:GST

[Bea91] Paul Beame. A general sequential time-space tradeoff for finding unique elements. SIAM Journal on Computing, 20(2): 270–277, April 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boros:2002:DBG

[BEG^{+}02] E. Boros, K. Elbassioni, V. Gurvich, L. Khachiyan, and K. Makino. Dual-bounded generating problems: All minimal integer solutions for a monotone system of linear inequalities. SIAM Journal on Computing, 31(5):1624–1643, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38876.

Beigi:2017:DRE

[BEG17] Salman Beigi, Omid Etesami, and Amin Gohari. Deterministic randomness extraction from generalized and distributed Santha–Vazirani sources. SIAM Journal on Computing, 46(1):1–36, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bonet:2000:RCR

[BEGJ00] Maria Luisa Bonet, Juan Luis Esteban, Nicola Galesi, and Jan Johannsen. On the relative complexity of resolution refinements and cutting planes proof systems. SIAM Journal on Computing, 30(5):1462–1484, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35247.

Becker:2024:DLS

[BEGL24] Ruben Becker, Yuval Emek, Mohsen Ghaffari, and Christoph Lenzen. Decentralized lowstretch trees via low diameter graph decompositions. SIAM Journal on Computing, 53(2): 247–286, March 2024. CODEN SMJCAT. ISSN 1095-7111.

Beigel:1990:USA

[Bei90] Richard Beigel. Unbounded searching algorithms. SIAM Journal on Computing, 19(3): 522–537, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barenboim:2014:DCL

[BEK14] Leonid Barenboim, Michael Elkin, and Fabian Kuhn. Distributed (Δ+1)-coloring in linear (in Δ) time. SIAM Journal on Computing, 43(1):72– 95, ???? 2014. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boros:2010:LRM

[BEM10] Endre Boros, Khaled Elbassioni, and Kazuhisa Makino. Left-to-right multiplication for monotone Boolean dualization. SIAM Journal on Computing, 39(7):3424–3439, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Benson:1977:SPP

[Ben77] David B. Benson. Some preservation properties of normal form grammars. SIAM Journal on Computing, 6(2):381– 402, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bennett:1989:TST

[Ben89] Charles H. Bennett. Time/ space trade-offs for reversible computation. SIAM Journal on Computing, 18(4):766–776, August 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Benczur:1995:CDN

[Ben95] András A. Benczúr. Counterexamples for directed and node capacitated cut-trees. SIAM Journal on Computing, 24(3):505–510, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23673.

Bernstein:1976:SRF

[Ber76] Arthur Bernstein. Synthesizing a response function with a feedback scheduling algorithm. SIAM Journal on Computing, 5(2):309–317, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Berger:1997:FMM

[Ber97] Bonnie Berger. The fourth moment method. SIAM Journal on Computing, 26(4):1188– 1207, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
24000.

Bernstein:2016:MSP

[Ber16] Aaron Bernstein. Maintaining shortest paths under deletions in weighted directed graphs. SIAM Journal on Computing, 45(2):548–574, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Babai:1980:RGI

[BES80] László Babai, Paul Erdős, and Stanley M. Selkow. Random graph isomorphism. SIAM Journal on Computing, 9(3): 628–635, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barahona:1989:NWM

[BéT89] Francisco Barahona and éva Tardos. Note on Weintraub’s minimum-cost circulation algorithm. SIAM Journal on Computing, 18(3):579–583, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bloom:1980:SIE

[BEW80a] Stephen L. Bloom, Calvin C. Elgot, and Jesse B. Wright. Solutions of the iteration equation and extensions of the scalar iteration operation. SIAM Journal on Computing, 9(1):25–45, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bloom:1980:VIP

[BEW80b] Stephen L. Bloom, Calvin C. Elgot, and Jesse B. Wright. Vector iteration in pointed iterative theories. SIAM Journal on Computing, 9(3):525– 540, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bermond:1994:BGB

[BF94] Jean-Claude Bermond and Pierre Fraigniaud. Broadcasting and gossiping in de Bruijn networks. SIAM Journal on Computing, 23(1):212– 225, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
19785.

Bohman:2003:ADP

[BF03a] Tom Bohman and Alan Frieze.

Arc-disjoint paths in expander digraphs. SIAM Journal on Computing, 32(2):326–344, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39858.

Boneh:2003:IBE

[BF03b] Dan Boneh and Matthew Franklin. Identity-based encryption from the Weil pairing. SIAM Journal on Computing, 32(3):586–615, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39852.

Borodin:1987:TST

[BFadH^{+}87] A. Borodin, F. Fich, F. Meyer auf der Heide, E. Upfal, and A. Wigderson. A time-space tradeoff for element distinctness. SIAM Journal on Computing, 16(1):97–99, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Berenbrink:2003:NWS

[BFG03] Petra Berenbrink, Tom Friedetzky, and Leslie Ann Goldberg. The natural workstealing algorithm is stable. SIAM Journal on Computing, 32(5):1260–1279, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39955.

Berenbrink:2007:DSL

[BFG^{+}07] Petra Berenbrink, Tom Friedetzky, Leslie Ann Goldberg, Paul W. Goldberg, Zengjian Hu, and Russell Martin. Distributed selfish load balancing. SIAM Journal on Computing, 37(4):1163–1181, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bekos:2019:PGB

[BFG^{+}19] Michael A. Bekos, Henry Förster, Martin Gronemann, Tamara Mchedlidze, Fabrizio Montecchiani, Chrysanthi Raftopoulou, and Torsten Ueckerdt. Planar graphs of bounded degree have bounded queue number. SIAM Journal on Computing, 48(5):1487– 1502, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buchbinder:2023:DAS

[BFG23] Niv Buchbinder, Moran Feldman, and Mohit Garg. Deterministic 1/2 (1/2 + ϵ)- approximation for submodular maximization over a matroid. SIAM Journal on Computing, 52(4):945–967, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M125515X.

Bansal:2014:MMG

[BFK^{+}14] Nikhil Bansal, Uriel Feige, Robert Krauthgamer, Konstantin Makarychev, Viswanath Nagarajan, Joseph Seffi Naor, and Roy Schwartz. Min-max graph partitioning and small set expansion. SIAM Journal on Computing, 43(2):872–904, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Becker:2021:NOA

[BFKL21] Ruben Becker, Sebastian Forster, Andreas Karrenbauer, and Christoph Lenzen. Nearoptimal approximate shortest paths and transshipment in distributed and streaming models. SIAM Journal on Computing, 50(3):815–856, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bodirsky:2011:BSP

[BFKV11] Manuel Bodirsky, Éric Fusy, Mihyun Kang, and Stefan Vigerske. Boltzmann samplers, Pólya theory, and cycle pointing. SIAM Journal on Computing, 40(3):721–769, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i3/
                                        p721_s1.

Buhrman:2002:RBK

[BFL02] Harry Buhrman, Lance Fortnow, and Sophie Laplante. Resource-bounded Kolmogorov complexity revisited. SIAM Journal on Computing, 31(3): 887–905, June 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        34388.

Bartal:2006:LBL

[BFL06] Yair Bartal, Amos Fiat, and Stefano Leonardi. Lower bounds for on-line graph problems with application to online circuit and optical routing. SIAM Journal on Computing, 36(2):354–393, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buhrman:2008:QPT

[BFNR08] Harry Buhrman, Lance Fortnow, Ilan Newman, and Hein Röhrig. Quantum property testing. SIAM Journal on Computing, 37(5):1387–1400, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buchbinder:2015:TLT

[BFNS15] Niv Buchbinder, Moran Feldman, Joseph Seffi Naor, and Roy Schwartz. A tight linear time (1/2)-approximation for unconstrained submodular maximization. SIAM Journal on Computing, 44(5):1384– 1402, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brown:1989:NBC

[BFP89] Cynthia A. Brown, Larry Finkelstein, and Paul Walton Purdom, Jr. A new base change algorithm for permutation groups. SIAM Journal on Computing, 18(5):1037–1047, October 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beigel:2006:IOA

[BFS06] Richard Beigel, Lance Fortnow, and Frank Stephan. Infinitely-often autoreducible sets. SIAM Journal on Computing, 36(3):595–608, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Broder:1998:OCE

[BFSU98] Andrei Z. Broder, Alan M.

Frieze, Stephen Suen, and Eli Upfal. Optimal construction of edge-disjoint paths in random graphs. SIAM Journal on Computing, 28(2):541–573, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29080. See cor-

rigendum [Fri02].

Broder:1994:ECE

[BFU94] Andrei Z. Broder, Alan M. Frieze, and Eli Upfal. Existence and construction of edge-disjoint paths on expander graphs. SIAM Journal on Computing, 23(5):976– 989, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23202.

Buhrman:2000:SCC

[BFvMT00] Harry Buhrman, Lance Fortnow, Dieter van Melkebeek, and Leen Torenvliet. Separating complexity classes using autoreducibility. SIAM Journal on Computing, 29 (5):1497–1520, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33473.

Bose:2015:OLR

[BFvRV15] Prosenjit Bose, Rolf Fagerberg, André van Renssen, and Sander Verdonschot. Optimal local routing on Delaunay triangulations defined by empty equilateral triangles. SIAM Journal on Computing, 44(6): 1626–1649, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bennett:1981:RRO

[BG81a] Charles H. Bennett and John Gill. Relative to a random oracle A, P^{A} ≠ NP^{A} ≠ co−NP^{A} with probability 1. SIAM Journal on Computing, 10(1):96–113, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bernstein:1981:PNS

[BG81b] Philip A. Bernstein and Nathan Goodman. The power of natural semijoins. SIAM Journal on Computing, 10(4): 751–771, February 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brent:1982:SAT

[BG82] Richard P. Brent and Leslie M.

Goldschlager. Some area-time tradeoffs for VLSI. SIAM Journal on Computing, 11(4): 737–747, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blass:1984:ERI

[BG84] Andreas Blass and Yuri Gurevich. Equivalence relations, invariants, and normal forms. SIAM Journal on Computing, 13(4):682–689, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Breslauer:1990:OTP

[BG90] Dany Breslauer and Zvi Galil. An optimal O(log log n) time parallel string matching algorithm. SIAM Journal on Computing, 19(6):1051–1058, December 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bini:1992:CPZ

[BG92a] Dario Bini and Luca Gemignani. On the complexity of polynomial zeros. SIAM Journal on Computing, 21(4):781–799, August 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Breslauer:1992:LBP

[BG92b] Dany Breslauer and Zvi Galil. A lower bound for parallel string matching. SIAM Journal on Computing, 21(5):856– 862, October 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blass:1993:RRS

[BG93a] Andreas Blass and Yuri Gurevich. Randomizing reductions of search problems. SIAM Journal on Computing, 22(5): 949–975, October 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buss:1993:NW

[BG93b] Jonathan F. Buss and Judy Goldsmith. Nondeterminism within P^{∗}. SIAM Journal on Computing, 22(3):560–572, June 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bellare:1994:CDV

[BG94] Mihir Bellare and Shafi Goldwasser. The complexity of decision versus search. SIAM Journal on Computing, 23 (1):97–119, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22828.

Bini:1995:FPC

[BG95a] Dario Bini and Luca Gemignani.

Fast parallel computation of the polynomial remainder sequence via Bézout and Hankel matrices. SIAM Journal on Computing, 24(1):63– 77, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/

20190. See erratum [BG96].

Blass:1995:MTC

[BG95b] Andreas Blass and Yuri Gurevich. Matrix transformation is complete for the average case. SIAM Journal on Computing, 24(1):3–29, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23207.

Bini:1996:EFP

[BG96] Dario Bini and Luca Gemignani. Erratum: “Fast parallel computation of the polynomial remainder sequence via Bézout and Hankel matrices” [SIAM J. Comput. 24 (1995), no. 1, 63–77, MR 95j:65048]. SIAM Journal on Computing, 25(6): 1358, December 1996. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [BG95a].

Beigel:1998:DSF

[BG98] R. Beigel and J. Goldsmith. Downward separation fails catastrophically for limited nondeterminism classes. SIAM Journal on Computing, 27 (5):1420–1429, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27742.

Barak:2009:UAT

[BG09] Boaz Barak and Oded Goldreich. Universal arguments and their applications. SIAM Journal on Computing, 38(5): 1661–1694, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brakensiek:2021:PCS

[BG21] Joshua Brakensiek and Venkatesan Guruswami. Promise constraint satisfaction: Algebraic structure and a symmetric Boolean dichotomy. SIAM Journal on Computing, 50(6): 1663–1700, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Broadbent:2022:QHC

[BG22] Anne Broadbent and Alex Bredariol Grilo. QMA-hardness of consistency of local density matrices with applications to quantum zero-knowledge. SIAM Journal on Computing, 51(4):1400–1450, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M140729X.

Byrka:2023:BAB

[BGA23] Jarosław Byrka, Fabrizio Grandoni, and Afrouz Jabal Ameli. Breaching the 2-approximation barrier for connectivity augmentation: a reduction to Steiner tree. SIAM Journal on Computing, 52(3):718–739, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421143.

Bansal:2018:LTF

[BGG18] Nikhil Bansal, Anupam Gupta, and Guru Guruganesh. On the Lovász theta function for independent sets in sparse graphs. SIAM Journal on Computing, 47(3):1039–1055, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bezakova:2019:ACD

[BGG^{+}19] Ivona Bezáková, Andreas Galanis, Leslie Ann Goldberg, Heng Guo, and Daniel Stefankovic. Approximation via correlation decay when strong spatial mixing fails. SIAM Journal on Computing, 48(2): 279–349, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brakensiek:2022:CSP

[BGG22] Joshua Brakensiek, Sivakanth Gopi, and Venkatesan Guruswami. Constraint satisfaction problems with global modular constraints: Algorithms and hardness via polynomial representations. SIAM Journal on Computing, 51(3):577–626, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/19M1291054.

Bhattiprolu:2023:IMN

[BGG^{+}23] Vijay Bhattiprolu, Mrinal Kanti Ghosh, Venkatesan Guruswami, Euiwoong Lee, and Madhur Tulsiani. Inapproximability of matrix p → q norms. SIAM Journal on Computing, 52(1):132–155, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/18M1233418.

Bshouty:1998:ELD

[BGGM98] Nader H. Bshouty, Paul W.

Goldberg, Sally A. Goldman, and H. David Mathias. Exact learning of discretized geometric concepts. SIAM Journal on Computing, 28(2):674–699, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27424.

Bezakova:2020:IIS

[BGGS20] Ivona Bezáková, Andreas Galanis, Leslie Ann Goldberg, and Daniel Stefankovic. Inapproximability of the independent set polynomial in the complex plane. SIAM Journal on Computing, 49(5): STOC18–395–STOC18–448, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barak:2015:MLC

[BGH^{+}15] Boaz Barak, Parikshit Gopalan, Johan Håstad, Raghu Meka, Prasad Raghavendra, and David Steurer. Making the long code shorter. SIAM Journal on Computing, 44(5):1287– 1324, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bhattacharyya:2012:TCS

[BGJ^{+}12] Arnab Bhattacharyya, Elena Grigorescu, Kyomin Jung, Sofya Raskhodnikova, and David P. Woodruff. Transitiveclosure spanners. SIAM Journal on Computing, 41(6):1380– 1425, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buchsbaum:2008:LTA

[BGK^{+}08] Adam L. Buchsbaum, Loukas Georgiadis, Haim Kaplan, Anne Rogers, Robert E. Tarjan, and Jeffery R. Westbrook. Linear-time algorithms for dominators and other pathevaluation problems. SIAM Journal on Computing, 38(4): 1533–1573, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bartal:2016:TSP

[BGK16] Yair Bartal, Lee-Ad Gottlieb, and Robert Krauthgamer. The traveling salesman problem: Low-dimensionality implies a polynomial time approximation scheme. SIAM Journal on Computing, 45(4):1563– 1581, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Braverman:2018:NOB

[BGK^{+}18] Mark Braverman, Ankit Garg, Young Kun Ko, Jieming Mao, and Dave Touchette. Near-optimal bounds on the bounded-round quantum communication complexity of disjointness. SIAM Journal on Computing, 47(6):2277–2314, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Batra:2020:CFA

[BGK20] Jatin Batra, Naveen Garg, and Amit Kumar. Constant factor approximation algorithm for weighted flow-time on a single machine in PseudoPolynomial time. SIAM Journal on Computing, 52(6):FOCS18– 158–FOCS18–188, September 2020. CODEN SMJCAT. ISSN 1095-7111.

Babai:2003:CCS

[BGKL03] László Babai, Anna Gál, Peter G. Kimmel, and Satyanarayana V. Lokam. Communication complexity of simultaneous messages. SIAM Journal on Computing, 33 (1):137–166, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37594.

Bruell:1999:SSA

[BGKP99] Steven C. Bruell, Sukumar Ghosh, Mehmet Hakan Karaata, and Sriram V. Pemmaraju. Self-stabilizing algorithms for finding centers and medians of trees. SIAM Journal on Computing, 29(2): 600–614, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
27156.

Bourgain:2012:HSP

[BGKS12] Jean Bourgain, Moubariz Z. Garaev, Sergei V. Konyagin, and Igor E. Shparlinski. On the hidden shifted power problem. SIAM Journal on Computing, 41(6):1524–1557, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bhattacharjee:2004:IFA

[BGL04] Rajat Bhattacharjee, Ashish Goel, and Zvi Lotker. Instability of FIFO at arbitrarily low rates in the adversarial queueing model. SIAM Journal on Computing, 34(2):318–332, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42680.

Bhatt:1999:TBL

[BGLL99] Sandeep Bhatt, David Greenberg, Tom Leighton, and Pangfeng Liu. Tight bounds for on-line tree embeddings. SIAM Journal on Computing, 29(2):474–491, April 1999.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30871.

Bansal:2018:FSE

[BGNV18] Nikhil Bansal, Shashwat Garg, Jesper Nederlof, and Nikhil Vyas. Faster space-efficient algorithms for subset sum, ksum, and related problems. SIAM Journal on Computing, 47(5):1755–1777, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bernardini:2022:EDS

[BGP^{+}22] Giulia Bernardini, Paweł Gawrychowski, Nadia Pisanti, Solon P. Pissis, and Giovanna Rosone. Elastic-degenerate string matching via fast matrix multiplication. SIAM Journal on Computing, 51(3):549–576, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1368033.

Bhattacharyya:2023:NOL

[BGP^{+}23] Arnab Bhattacharyya, Sutanu Gayen, Eric Price, Vincent Y. F. Tan, and N. V. Vinodchandran. Near-optimal learning of tree-structured distributions by Chow and Liu. SIAM Journal on Computing, 52(3):761–793, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1489678.

Bermond:1998:FGS

[BGRV98] Jean-Claude Bermond, Luisa Gargano, Adele A. Rescigno, and Ugo Vaccaro. Fast gossiping by short messages. SIAM Journal on Computing, 27(4):917–941, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28361.

Brown:1972:PSP

[BGS72] Steven Brown, David Gries, and Thomas Szymanski. Program schemes with pushdown stores. SIAM Journal on Computing, 1(3):242–268, ???? 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baker:1975:RQ

[BGS75] Theodore Baker, John Gill, and Robert Solovay. Relativizations of the P =?N P question. SIAM Journal on Computing, 4(4):431–442, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bellare:1998:FBP

[BGS98] Mihir Bellare, Oded Goldreich, and Madhu Sudan. Free bits, PCPs, and nonapproximability — towards tight results. SIAM Journal on Computing, 27(3):804–915, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30253.

Bostan:2007:LRP

[BGS07] Alin Bostan, Pierrick Gaudry, and Éric Schost. Linear recurrences with polynomial coefficients and application to integer factorization and Cartier– Manin operator. SIAM Journal on Computing, 36(6):1777– 1806, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baswana:2015:FDM

[BGS15] Surender Baswana, Manoj Gupta, and Sandeep Sen. Fully dynamic maximal matching in O(log n) update time. SIAM Journal on Computing, 44(1): 88–113, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See corrected version [BGS18].

Baswana:2018:FDM

[BGS18] Surender Baswana, Manoj Gupta, and Sandeep Sen. Fully dynamic maximal matching in O(log n) update time (corrected version). SIAM Journal on Computing, 47(3):617–650, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [BGS15].

Bringmann:2019:TSA

[BGSW19] Karl Bringmann, Fabrizio Grandoni, Barna Saha, and Virginia Vassilevska Williams. Truly subcubic algorithms for language edit distance and RNA folding via fast boundeddifference min-plus product. SIAM Journal on Computing, 48(2):481–512, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bonacina:2016:TSR

[BGT16] Ilario Bonacina, Nicola Galesi, and Neil Thapen. Total space in resolution. SIAM Journal on Computing, 45(5):1894– 1909, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barve:2000:ACP

[BGV00] Rakesh D. Barve, Edward F. Grove, and Jeffrey Scott Vitter. Application-controlled paging for a shared cache. SIAM Journal on Computing, 29(4):1290–1303, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32427.

Buchsbaum:2000:DWF

[BGW00] Adam L. Buchsbaum, Raffaele Giancarlo, and Jeffery R. Westbrook. On the determinization of weighted finite automata. SIAM Journal on Computing, 30(5):1502– 1531, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34667.

Brakensiek:2020:PCB

[BGWZ20] Joshua Brakensiek, Venkatesan Guruswami, Marcin Wrochna, and Stanislav Zivný. The power of the combined basic linear programming and affine relaxation for promise constraint satisfaction problems. SIAM Journal on Computing, 49(6):1232–1248, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bruno:1975:SCJ

[BH75] John Bruno and Micha Hofri.

On scheduling chains of jobs on one processor with limited preemption. SIAM Journal on Computing, 4(4):478–490, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Berman:1977:IDO

[BH77] L. Berman and J. Hartmanis.

On isomorphisms and density of NP and other complete sets. SIAM Journal on Computing, 6(2):305–322, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beyer:1980:CTG

[BH80] Terry Beyer and Sandra Mitchell Hedetniemi. Constant time generation of rooted trees. SIAM Journal on Computing, 9(4):706–712, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beeri:1981:PFD

[BH81a] C. Beeri and P. Honeyman. Preserving functional dependencies. SIAM Journal on Computing, 10(3):647–656, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bitner:1981:KCO

[BH81b] James R. Bitner and Shou Hsuan Huang. Key comparison optimal 2 − 3 trees with maximum utilization. SIAM Journal on Computing, 10(3):558– 570, ???? 1981. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Becker:1987:OLP

[BH87] B. Becker and G. Hotz. On the optimal layout of planar graphs with fixed boundary. SIAM Journal on Computing, 16(5):946–972, October 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bodlaender:1998:PAO

[BH98] Hans L. Bodlaender and Torben Hagerup. Parallel algorithms with optimal speedup for bounded treewidth. SIAM Journal on Computing, 27(6): 1725–1746, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28985.

Bjorklund:2003:FPS

[BH03] Andreas Björklund and Thore Husfeldt. Finding a path of superlogarithmic length. SIAM Journal on Computing, 32(6):1395–1402, December 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41676.

Beame:2012:MCC

[BH12] Paul Beame and Trinh Huynh.

Multiparty communication complexity and threshold circuit size of oAC^{0}. SIAM Journal on Computing, 41(3):484– 518, ???? 2012. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Burton:2013:LRS

[BH13] Benjamin A. Burton and Mathias Hiron. Locating regions in a sequence under density constraints. SIAM Journal on Computing, 42(3):1201– 1215, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Balcan:2018:SFL

[BH18] Maria-Florina Balcan and Nicholas J. A. Harvey. Submodular functions: Learnability, structure, and optimization. SIAM Journal on Computing, 47(3):703–754, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bjorklund:2019:STD

[BH19] Andreas Björklund and Thore Husfeldt. Shortest two disjoint paths in polynomial time. SIAM Journal on Computing, 48(6):1698–1710, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bshouty:1995:LAR

[BHH95] Nader H. Bshouty, Thomas R. Hancock, and Lisa Hellerstein. Learning arithmetic read-once formulas. SIAM Journal on Computing, 24 (4):706–735, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/23664.

Bansal:2013:HAS

[BHI^{+}13] Nikhil Bansal, Xin Han, Kazuo Iwama, Maxim Sviridenko, and Guochuan Zhang. A harmonic algorithm for the 3D strip packing problem. SIAM Journal on Computing, 42(2): 579–592, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bhattacharya:2018:DFD

[BHI18] Sayan Bhattacharya, Monika Henzinger, and Giuseppe F. Italiano. Deterministic fully dynamic data structures for vertex cover and matching. SIAM Journal on Computing, 47(3):859–887, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boros:1997:PTR

[BHIK97] Endre Boros, Peter L. Hammer, Toshihide Ibaraki, and Kazuhiko Kawakami. Polynomialtime recognition of 2-monotonic positive Boolean functions given by an oracle. SIAM Journal on Computing, 26 (1):93–109, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26908.

Bjorklund:2009:SPI

[BHK09] Andreas Björklund, Thore Husfeldt, and Mikko Koivisto. Set partitioning via inclusionexclusion. SIAM Journal on Computing, 39(2):546–563, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barak:2019:NTS

[BHK^{+}19] Boaz Barak, Samuel Hopkins, Jonathan Kelner, Pravesh K. Kothari, Ankur Moitra, and Aaron Potechin. A nearly tight sum-of-squares lower bound for the planted clique problem. SIAM Journal on Computing, 48(2):687–735, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buhrman:1995:SRC

[BHL95] Harry Buhrman, Edith Hemaspaandra, and Luc Longpré. SPARSE reduces conjunctively to TALLY. SIAM Journal on Computing, 24(4):673–681, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24672.

Bateni:2018:IAA

[BHL18] Mohammad Hossein Bateni, Mohammad Taghi Hajiaghayi, and Vahid Liaghat. Improved approximation algorithms for (budgeted) nodeweighted Steiner problems. SIAM Journal on Computing, 47(4):1275–1293, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beimel:2022:TBM

[BHMO22] Amos Beimel, Iftach Haitner, Nikolaos Makriyannis, and Eran Omri. Tighter bounds on MultiParty coin flipping via augmented weak martingales and differentially private sampling. SIAM Journal on Computing, 51(4):1126–1171, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/18M1210782.

Bhattacharya:2023:DNO

[BHNW23] Sayan Bhattacharya, Monika Henzinger, Danupon Nanongkai, and Xiaowei Wu. Deterministic near-optimal approximation algorithms for dynamic set cover. SIAM Journal on Computing, 52(5):1132–1192, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1428649.

Bhangale:2024:RMR

[BHPT24] Amey Bhangale, Prahladh Harsha, Orr Paradise, and Avishay Tal. Rigid matrices from rectangular PCPs. SIAM Journal on Computing, 53(2): 480–523, April 2024. CODEN SMJCAT. ISSN 1095-7111.

Buhrman:1998:SRS

[BHT98] Harry Buhrman, Albrecht Hoene, and Leen Torenvliet. Splittings, robustness, and structure of complete sets. SIAM Journal on Computing, 27(3):637–653, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27972.

Buhrman:1999:SER

[BHV99] Harry Buhrman, Jaap-Henk Hoepman, and Paul Vitányi. Space-efficient routing tables for almost all networks and the incompressibility method. SIAM Journal on Computing, 28(4):1414–1432, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30848.

Brodowsky:2023:ARK

[BHZ23] Ulrich A. Brodowsky, Stefan Hougardy, and Xianghui Zhong. The approximation ratio of the k -opt heuristic for the Euclidean traveling salesman problem. SIAM Journal on Computing, 52(4):841–864, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M146199X.

Bienstock:2006:AFP

[BI06] D. Bienstock and G. Iyengar. Approximating fractional packings and coverings in O(1/ϵ) iterations. SIAM Journal on Computing, 35(4): 825–854, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Backurs:2018:EDC

[BI18] Arturs Backurs and Piotr Indyk. Edit distance cannot be computed in strongly subquadratic time (unless SETH is false). SIAM Journal on Computing, 47(3):1087–1097, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bienstock:1990:LTT

[Bie90] D. Bienstock. Linear-time test for small face covers in any fixed surface. SIAM Journal on Computing, 19(5):907–911, October 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bannai:2017:RT

[BII^{+}17] Hideo Bannai, Tomohiro I, Shunsuke Inenaga, Yuto Nakashima, Masayuki Takeda, and Kazuya Tsuruta. The “Runs” Theorem. SIAM Journal on Computing, 46(5):1501– 1514, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bini:1984:PSC

[Bin84] Dario Bini. Parallel solution of certain Toeplitz linear systems. SIAM Journal on Computing, 13(2):268–276, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Biskup:1978:TMO

[Bis78] Joachim Biskup. The time measure of one-tape Turing machines does not have the parallel computation property. SIAM Journal on Computing, 7(1):115–117, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bitner:1979:HDO

[Bit79] James R. Bitner. Heuristics that dynamically organize data structures. SIAM Journal on Computing, 8(1):82–110, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bitner:1982:AOA

[Bit82] James R. Bitner. An asymptotically optimal algorithm for the Dutch National Flag problem. SIAM Journal on Computing, 11(2):243–262, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barak:2006:ERU

[BIW06] Boaz Barak, Russell Impagliazzo, and Avi Wigderson. Extracting randomness using few independent sources. SIAM Journal on Computing, 36(4): 1095–1118, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Booth:1982:DSC

[BJ82] Kellogg S. Booth and J. Howard Johnson. Dominating sets in chordal graphs. SIAM Journal on Computing, 11(1):191–199, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bshouty:1999:LDU

[BJ99] Nader H. Bshouty and Jeffrey C. Jackson. Learning DNF over the uniform distribution using a quantum example oracle. SIAM Journal on Computing, 28(3):1136–1153, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29312.

Berkman:1994:TBR

[BJK^{+}94] Omer Berkman, Joseph Ja’Ja’, Sridhar Krishnamurthy, Ramakrishna Thurimella, and Uzi Vishkin. Top-bottom routing around a rectangle is as easy as computing prefix minima. SIAM Journal on Computing, 23(3):449–465, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/21827.

Bulatov:2005:CCC

[BJK05] Andrei Bulatov, Peter Jeavons, and Andrei Krokhin. Classifying the complexity of constraints using finite algebras. SIAM Journal on Computing, 34(3):720–742, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37667.

Blondel:2005:DUP

[BJKP05] Vincent D. Blondel, Emmanuel Jeandel, Pascal Koiran, and Natacha Portier. Decidable and undecidable problems about quantum automata. SIAM Journal on Computing, 34(6):1464–1473, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42586.

Barrett:2000:FLC

[BJM00] Chris Barrett, Riko Jacob, and Madhav Marathe. Formal-language-constrained path problems. SIAM Journal on Computing, 30(3):809–837, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33771.

Bjorklund:2014:DSU

[Bjö14] Andreas Björklund. Determinant sums for undirected Hamiltonicity. SIAM Journal on Computing, 43(1):280–299, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bernstein:1989:SAL

[BJR89] David Bernstein, Jeffrey M. Jaffe, and Michael Rodeh. Scheduling arithmetic and load operations in parallel with no spilling. SIAM Journal on Computing, 18(6):1098–1127, December 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baliga:1997:LMS

[BJS97] Ganesh Baliga, Sanjay Jain, and Arun Sharma. Learning from multiple sources of inaccurate data. SIAM Journal on Computing, 26(4):961– 990, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23946.

Broadbent:2020:ZKP

[BJSW20] Anne Broadbent, Zhengfeng Ji, Fang Song, and John Watrous. Zero-knowledge proof systems for QMA. SIAM Journal on Computing, 49(2):245– 283, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1988:STT

[BK88] Ronald V. Book and Ker-I Ko.

On sets truth-table reducible to sparse sets. SIAM Journal on Computing, 17(5):903–919, October 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bshouty:1990:MPF

[BK90] Nader H. Bshouty and Michael Kaminski. Multiplication of polynomials over finite fields. SIAM Journal on Computing, 19(3):452–456, June 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bradford:2007:PSC

[BK07] Phillip G. Bradford and Michael N. Katehakis. A probabilistic study on combinatorial expanders and hashing. SIAM Journal on Computing, 37(1):83–111, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barto:2010:CDI

[BK10a] Libor Barto and Marcin Kozik. Congruence distributivity implies bounded width. SIAM Journal on Computing, 39(4): 1531–1542, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baswana:2010:FAA

[BK10b] Surender Baswana and Telikepalli Kavitha. Faster algorithms for all-pairs approximate shortest paths in undirected graphs. SIAM Journal on Computing, 39(7):2865–2896, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Briest:2011:BCE

[BK11] Patrick Briest and Piotr Krysta. Buying cheap is expensive: Approximability of combinatorial pricing problems. SIAM Journal on Computing, 40(6):1554–1586, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i6/p1554_s1.

Benczur:2015:RAS

[BK15] András A. Benczúr and David R. Karger. Randomized approximation schemes for cuts and flows in capacitated graphs. SIAM Journal on Computing, 44(2):290–319, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barto:2016:RSC

[BK16] Libor Barto and Marcin Kozik.

Robustly solvable constraint satisfaction problems. SIAM Journal on Computing, 45(4): 1646–1669, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bertram–Kretzberg:2000:AHP

[BKHL00] Claudia Bertram-Kretzberg, Thomas Hofmeister, and Hanno Lefmann. An algorithm for Heilbronn’s problem. SIAM Journal on Computing, 30(2):383–390, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34887.

Buchsbaum:2004:OVL

[BKK^{+}04] Adam L. Buchsbaum, Howard Karloff, Claire Kenyon, Nick Reingold, and Mikkel Thorup. OPT versus LOAD in dynamic storage allocation. SIAM Journal on Computing, 33(3):632–646, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42394.

Broder:2010:HPL

[BKK^{+}10] Andrei Z. Broder, Adam Kirsch, Ravi Kumar, Michael Mitzenmacher, Eli Upfal, and Sergei Vassilvitskii. The hiring problem and Lake Wobegon strategies. SIAM Journal on Computing, 39(4):1233–1255, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bertram-Kretzberg:1999:AAU

[BKL99] Claudia Bertram-Kretzberg and Hanno Lefmann. The algorithmic aspects of uncrowded hypergraphs. SIAM Journal on Computing, 29 (1):201–230, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32371.

Borradaile:2017:MSM

[BKM^{+}17] Glencora Borradaile, Philip N. Klein, Shay Mozes, Yahav Nussbaum, and Christian Wulff-Nilsen. Multiple-source multiple-sink maximum flow in directed planar graphs in nearlinear time. SIAM Journal on Computing, 46(4):1280–1303, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Banks:2019:LTF

[BKM19] Jess Banks, Robert Kleinberg, and Cristopher Moore. The Lovász theta function for random regular graphs and community detection in the hard regime. SIAM Journal on Computing, 48(3):1098–1119, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barto:2009:CDH

[BKN09] Libor Barto, Marcin Kozik, and Todd Niven. The CSP dichotomy holds for digraphs with no sources and no sinks (a positive answer to a conjecture of Bang-Jensen and Hell). SIAM Journal on Computing, 38(5):1782–1802, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bansal:2010:AGD

[BKN10] Nikhil Bansal, Rohit Khandekar, and Viswanath Nagarajan. Additive guarantees for degree-bounded directed network design. SIAM Journal on Computing, 39(4):1413– 1431, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bergstra:1988:RFA

[BKO88] J. A. Bergstra, J. W. Klop, and E. R. Olderog. Readies and failures in the algebra of communicating processes. SIAM Journal on Computing, 17(6):1134–1177, December 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beame:2002:ERD

[BKPS02] Paul Beame, Richard Karp, Toniann Pitassi, and Michael Saks. The efficiency of resolution and Davis–Putnam procedures. SIAM Journal on Computing, 31(4):1048–1075, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36915.

Bronnimann:2006:CEP

[BKPS06] Hervé Brönnimann, Lutz Kettner, Michel Pocchiola, and Jack Snoeyink. Counting and enumerating pointed pseudotriangulations with the Greedy Flip Algorithm. SIAM Journal on Computing, 36(3):721–739, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bender:2018:CRC

[BKPY18] Michael A. Bender, Tsvi Kopelowitz, Seth Pettie, and Maxwell Young. Contention resolution with constant throughput and log-logstar channel accesses. SIAM Journal on Computing, 47(5):1735–1754, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blum:2000:DTT

[BKRS00] Avrim Blum, Howard Karloff, Yuval Rabani, and Michael Saks. A decomposition theorem for task systems and bounds for randomized server problems. SIAM Journal on Computing, 30(5):1624– 1661, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35188.

Broder:1994:TST

[BKRU94] Andrei Z. Broder, Anna R.

Karlin, Prabhakar Raghavan, and Eli Upfal. Trading space for time in undirected s-t connectivity. SIAM Journal on Computing, 23(2):324–334, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19014.

Bein:1992:ORT

[BKS92] Wolfgang W. Bein, Jerzy Kamburowski, and Matthias F. M. Stallmann. Optimal reduction of two-terminal directed acyclic graphs. SIAM Journal on Computing, 21(6):1112– 1129, December 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bulanek:2015:TLB

[BKS15] Jan Bulánek, Michal Koucký, and Michael Saks. Tight lower bounds for the online labeling problem. SIAM Journal on Computing, 44(6):1765– 1797, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Briest:2011:ATU

[BKV11] Patrick Briest, Piotr Krysta, and Berthold Vöcking. Approximation techniques for utilitarian mechanism design. SIAM Journal on Computing, 40(6):1587–1622, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i6/p1587_s1.

Book:1993:LVH

[BL93] Ronald V. Book and Jack H. Lutz. On languages with very high space-bounded Kolmogorov complexity. SIAM Journal on Computing, 22(2): 395–402, April 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blumofe:1998:SES

[BL98] Robert D. Blumofe and Charles E. Leiserson. Spaceefficient scheduling of multithreaded computations. SIAM Journal on Computing, 27 (1):202–229, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25947.

Breutzmann:1999:EMC

[BL99] Josef M. Breutzmann and Jack H. Lutz. Equivalence of measures of complexity classes. SIAM Journal on Computing, 29(1):302–326, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30226.

Benedikt:2000:SCQ

[BL00] Michael Benedikt and Leonid Libkin. Safe constraint queries. SIAM Journal on Computing, 29(5):1652–1682, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34248.

Buhrman:2002:CRB

[BL02] Harry Buhrman and Luc Longpré. Compressibility and resource bounded measure. SIAM Journal on Computing, 31(3):876–886, June 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31712.

Barak:2004:SPT

[BL04] Boaz Barak and Yehuda Lindell. Strict polynomial-time in simulation and extraction. SIAM Journal on Computing, 33(4):783–818, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42797.

Balcan:2016:CUP

[BL16] Maria Florina Balcan and Yingyu Liang. Clustering under perturbation resilience. SIAM Journal on Computing, 45(1):102–155, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bausch:2021:ETA

[BL21] Johannes Bausch and Felix Leditzky. Error thresholds for arbitrary Pauli noise. SIAM Journal on Computing, 50(4): 1410–1460, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blaser:2004:CCA

[Blä04]  Markus Bläser.  A com-

plete characterization of the algebras of minimal bilinear complexity. SIAM Journal on Computing, 34(2):277–298, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43827.

Burgisser:2011:OMI

[BLMW11] Peter Bürgisser, J. M. Landsberg, Laurent Manivel, and Jerzy Weyman. An overview of mathematical issues arising in the geometric complexity theory approach to VP ≠ VNP. SIAM Journal on Computing, 40(4):1179–1209, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i4/p1179_s1.

Boyar:2001:AFG

[BLN01] Joan Boyar, Kim S. Larsen, and Morten N. Nielsen. The accommodating function: a generalization of the competitive ratio. SIAM Journal on Computing, 31(1):233– 258, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36178.

Bansal:2015:MCM

[BLNZ15] Nikhil Bansal, Kang-Won Lee, Viswanath Nagarajan, and Murtaza Zafer. Minimum congestion mapping in a cloud. SIAM Journal on Computing, 44(3):819–843, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bloniarz:1983:SPA

[Blo83] Peter A. Bloniarz. A shortestpath algorithm with expected time O(n^{2} log n log^{∗} n). SIAM Journal on Computing, 12(3): 588–600, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bini:1980:ASB

[BLR80] Dario Bini, Grazia Lotti, and Francesco Romani. Approximate solutions for the bilinear form computational problem. SIAM Journal on Computing, 9(4):692–697, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barcelo:2014:EAC

[BLR14] Pablo Barceló, Leonid Libkin, and Miguel Romero. Efficient approximations of conjunctive queries. SIAM Journal on Computing, 43(3):1085– 1130, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bille:2015:RAG

[BLR^{+}15] Philip Bille, Gad M. Landau, Rajeev Raman, Kunihiko Sadakane, Srinivasa Rao Satti, and Oren Weimann. Random access to grammar-compressed strings and trees. SIAM Journal on Computing, 44(3):513– 539, ???? 2015. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1984:QRC

[BLS84] Ronald V. Book, Timothy J. Long, and Alan L. Selman.

Quantitative relativizations of complexity classes. SIAM Journal on Computing, 13(3): 461–487, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Babai:1997:FMP

[BLS97] László Babai, Eugene M. Luks, and Ákos Seress. Fast management of permutation groups. I. SIAM Journal on Computing, 26(5):1310–1342, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22941.

Buhrman:1999:KRG

[BLTV99] Harry Buhrman, Ming Li, John Tromp, and Paul Vitányi. Kolmogorov random graphs and the incompressibility method. SIAM Journal on Computing, 29(2):590–599, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32780.

Blum:1986:SOW

[Blu86] Norbert Blum. On the singleoperation worst-case time complexity of the disjoint set union problem. SIAM Journal on Computing, 15(4):1021–1024, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blum:1994:SDF

[Blu94] Avrim L. Blum. Separating distribution-free and mistakebound learning models over the Boolean domain. SIAM Journal on Computing, 23 (5):990–1000, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/23455.

Blum:1984:HGC

[BM84] Manuel Blum and Silvio Micali. How to generate cryptographically strong sequences of pseudorandom bits. SIAM Journal on Computing, 13(4): 850–864, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bienstock:1988:CCV

[BM88] Daniel Bienstock and Clyde L. Monma. On the complexity of covering vertices by faces in a planar graph. SIAM Journal on Computing, 17(1):53– 76, February 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brodnik:1999:MCT

[BM99] Andrej Brodnik and J. Ian Munro. Membership in constant time and almostminimum space. SIAM Journal on Computing, 28(5):1627– 1640, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
29416.

Bshouty:2002:SLA

[BM02] Nader H. Bshouty and Yishay Mansour. Simple learning algorithms for decision trees and multivariate polynomials. SIAM Journal on Computing, 31(6):1909–1925, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32058.

Bartal:2004:MMS

[BM04a] Yair Bartal and Manor Mendel.

Multiembedding of metric spaces. SIAM Journal on Computing, 34(1):248–259, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43312.

Bose:2004:ORT

[BM04b] Prosenjit Bose and Pat Morin.

Online routing in triangulations. SIAM Journal on Computing, 33(4):937–951, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36938.

Bulatov:2014:CSP

[BM14] Andrei A. Bulatov and Dániel Marx. Constraint satisfaction parameterized by solution size. SIAM Journal on Computing, 43(2):573–616, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Busch:2007:UBP

[BMIM07] Costas Busch, Malik Magdon-Ismail, and Marios Mavronicolas. Universal bufferless packet switching. SIAM Journal on Computing, 37(4):1139– 1162, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Moshe:2007:CFA

[BMKM07] Boaz Ben-Moshe, Matthew J. Katz, and Joseph S. B. Mitchell. A constant-factor approximation algorithm for optimal 1.5D terrain guarding. SIAM Journal on Computing, 36(6):1631–1647, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bodirsky:2021:PAT

[BMM21] Manuel Bodirsky, Florent Madelaine, and Antoine Mottet. A proof of the algebraic tractability conjecture for monotone monadic SNP. SIAM Journal on Computing, 50(4):1359–1409, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bodirsky:2019:CSP

[BMPP19] Manuel Bodirsky, Barnaby Martin, Michael Pinsker, and András Pongrácz. Constraint satisfaction problems for reducts of homogeneous graphs. SIAM Journal on Computing, 48(4):1224–1264, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beaudry:1997:FMW

[BMPT97] Martin Beaudry, Pierre McKenzie, Pierre Peladeau, and Denis Therien. Finite monoids: From word to circuit evaluation. SIAM Journal on Computing, 26(1):138–152, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24953.

Buhrman:2002:BO

[BMRV02] H. Buhrman, P. B. Miltersen, J. Radhakrishnan, and S. Venkatesh. Are bitvectors optimal? SIAM Journal on Computing, 31(6): 1723–1744, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40529.

Bach:1986:SDP

[BMS86] Eric Bach, Gary Lee Miller, and Jeffrey O. Shallit. Sums of divisors, perfect numbers and factoring. SIAM Journal on Computing, 15(4):1143– 1154, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bresler:2013:RMR

[BMS13] Guy Bresler, Elchanan Mossel, and Allan Sly. Reconstruction of Markov random fields from samples: Some observations and algorithms. SIAM Journal on Computing, 42(2): 563–578, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beeri:1981:ERD

[BMSU81] Catriel Beeri, Alberto O. Mendelzon, Yehoshua Sagiv, and Jeffrey D. Ullman. Equivalence of relational database schemes. SIAM Journal on Computing, 10(2):352–370, May 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1978:LLI

[BN78] Ronald V. Book and Maurice Nivat. Linear languages and the intersection closures of classes of languages. SIAM Journal on Computing, 7(2): 167–177, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Banaschewski:1982:CPO

[BN82] Bernhard Banaschewski and Evelyn Nelson. Completions of partially ordered sets. SIAM Journal on Computing, 11(3): 521–528, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bilardi:1989:ABS

[BN89] Gianfranco Bilardi and Alexandru Nicolau. Adaptive bitonic sorting: an optimal parallel algorithm for shared-memory machines. SIAM Journal on Computing, 18(2):216–228, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bellantoni:1999:RPR

[BN99a] Stephen J. Bellantoni and Karl-Heinz Niggl. Ranking primitive recursions: The low Grzegorczyk classes revisited. SIAM Journal on Computing, 29(2):401–415, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28175.

Bertsch:1999:RCD

[BN99b] Eberhard Bertsch and Mark-Jan Nederhof. Regular closure of deterministic languages. SIAM Journal on Computing, 29(1):81–102, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28682.

Boppana:2000:PIL

[BN00] Ravi B. Boppana and Babu O. Narayanan. Perfect-information leader election with optimal resilience. SIAM Journal on Computing, 29(4):1304– 1320, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        30718.

Blumrosen:2010:CPD

[BN10] Liad Blumrosen and Noam Nisan. On the computational power of demand queries. SIAM Journal on Computing, 39(4):1372–1391, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brenner:2014:IDP

[BN14] Hai Brenner and Kobbi Nissim. Impossibility of differentially private universally optimal mechanisms. SIAM Journal on Computing, 43(5):1513– 1540, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Berkholz:2020:SSW

[BN20] Christoph Berkholz and Jakob Nordström. Supercritical space-width trade-offs for resolution. SIAM Journal on Computing, 49(1):98–118, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bar-Noy:1989:BUS

[BNBK^{+}89] Amotz Bar-Noy, Allan Borodin, Mauricio Karchmer, Nathan Linial, and Michael Werman. Bounds on universal sequences. SIAM Journal on Computing, 18(2):268–277, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bar-Noy:1999:BAP

[BNCK^{+}99] Amotz Bar-Noy, Ran Canetti, Shay Kutten, Yishay Mansour, and Baruch Schieber. Bandwidth allocation with preemption. SIAM Journal on Computing, 28(5):1806–1828, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32123.

Bar-Noy:2001:LLB

[BNFN01] Amotz Bar-Noy, Ari Freund, and Joseph (Seffi) Naor. On-line load balancing in a hierarchical server topology. SIAM Journal on Computing, 31(2):527–549, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34613.

Bar-Noy:2000:MMH

[BNGNS00] Amotz Bar-Noy, Sudipto Guha, Joseph (Seffi) Naor, and Baruch Schieber. Message multicasting in heterogeneous networks. SIAM Journal on Computing, 30(2):347–358, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34790.

Bar-Noy:2001:ATM

[BNGNS01] Amotz Bar-Noy, Sudipto Guha, Joseph (Seffi) Naor, and Baruch Schieber. Approximating the throughput of multiple machines in real-time scheduling. SIAM Journal on Computing, 31(2):331–352, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35413.

Bar-Noy:2003:WSP

[BNL03] Amotz Bar-Noy and Richard E. Ladner. Windows scheduling problems for broadcast systems. SIAM Journal on Computing, 32(4):1091–1113, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40447.

Bar-Noy:2004:EAO

[BNL04] Amotz Bar-Noy and Richard E. Ladner. Efficient algorithms for optimal stream merging for media-on-demand. SIAM Journal on Computing, 33 (5):1011–1034, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38924.

Bar-Noy:1998:GFS

[BNMSS98] Amotz Bar-Noy, Alain Mayer, Baruch Schieber, and Madhu Sudan. Guaranteeing fair service to persistent dependent tasks. SIAM Journal on Computing, 27(4):1168–1189, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28209.

Book:1974:RBA

[BNP74] Ronald Book, Maurice Nivat, and Michael Paterson. Reversal-bounded acceptors and intersections of linear languages. SIAM Journal on Computing, 3(4):283–295, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buchbinder:2018:SPE

[BNS18] Niv Buchbinder, Joseph Seffi Naor, and Roy Schwartz. Simplex partitioning via exponential clocks and the multiwaycut problem. SIAM Journal on Computing, 47(4):1463– 1482, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bassily:2021:ASA

[BNS^{+}21] Raef Bassily, Kobbi Nissim, Adam Smith, Thomas Steinke, Uri Stemmer, and Jonathan Ullman. Algorithmic stability for adaptive data analysis. SIAM Journal on Computing, 50(3):STOC16–377–STOC16– 405, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beals:1998:CNL

[BNT98] Robert Beals, Tetsuro Nishino, and Keisuke Tanaka. On the complexity of negation-limited Boolean networks. SIAM Journal on Computing, 27 (5):1334–1347, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27513.

Brassard:2019:NIQ

[BNT^{+}19] Gilles Brassard, Ashwin Nayak, Alain Tapp, Dave Touchette, and Falk Unger. Noisy interactive quantum communication. SIAM Journal on Computing, 48(4):1147–1195, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Braverman:2010:ECS

[BO10] Vladimir Braverman and Rafail Ostrovsky. Effective computations on sliding windows. SIAM Journal on Computing, 39(6):2113–2131, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buresh-Oppenheim:2004:BDF

[BOBP^{+}04] Joshua Buresh-Oppenheim, Paul Beame, Toniann Pitassi, Ran Raz, and Ashish Sabharwal. Bounded-depth Frege lower bounds for weaker pigeonhole principles. SIAM Journal on Computing, 34(2): 261–276, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43314.

Ben-Or:1992:CAF

[BOC92] Michael Ben-Or and Richard Cleve. Computing algebraic formulas using a constant number of registers. SIAM Journal on Computing, 21(1):54– 58, February 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bodlaender:1996:LTA

[Bod96] Hans L. Bodlaender. A lineartime algorithm for finding tree-decompositions of small treewidth. SIAM Journal on Computing, 25(6):1305–1317, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
25121.

Bodwin:2021:NRL

[Bod21] Greg Bodwin. New results on linear size distance preservers. SIAM Journal on Computing, 50(2):662–673, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Or:1988:FPA

[BOFKT88] Michael Ben-Or, Ephraim Feig, Dexter Kozen, and Prasoon Tiwari. A fast parallel algorithm for determining all roots of a polynomial with real roots. SIAM Journal on Computing, 17(6):1081–1092, December 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Or:2014:QMI

[BOHP14] Michael Ben-Or, Avinatan Hassidim, and Haran Pilpel. Quantum multiprover interactive proofs with communicating provers. SIAM Journal on Computing, 43(3):987– 1011, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bollman:1976:PPE

[Bol76] D. Bollman. On preserving proximity in extendible arrays. SIAM Journal on Computing, 5(2):318–323, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bolour:1981:ORA

[Bol81] Azad Bolour. Optimal retrieval algorithms for small region queries. SIAM Journal on Computing, 10(4):721–741, ???? 1981. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1972:LAP

[Boo72a] Ronald V. Book. On languages accepted in polynomial time. SIAM Journal on Computing, 1(4):281–287, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1972:TCC

[Boo72b]  Ronald V. Book.  Termi-

nal context in context-sensitive grammars. SIAM Journal on Computing, 1(1):20–30, ???? 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Booth:1978:ITG

[Boo78] Kellogg S. Booth. Isomorphism testing for graphs, semigroups, and finite automata are polynomially equivalent problems. SIAM Journal on Computing, 7(3):273–279, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1979:PST

[Boo79] Ronald V. Book. Polynomial space and transitive closure. SIAM Journal on Computing, 8(3):434–439, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1991:SOS

[Boo91] Ronald V. Book. Some observations on separating complexity classes. SIAM Journal on Computing, 20(2):246–258, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1994:LRA

[Boo94] Ronald V. Book. On languages reducible to algorithmically random languages. SIAM Journal on Computing, 23(6):1275–1282, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23814.

Borodin:1977:RTS

[Bor77] Allan Borodin. On relating time and space to size and depth. SIAM Journal on Computing, 6(4):733–744, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1988:LPS

[BORW88] R. Book, P. Orponen, D. Russo, and O. Watanabe. Lowness properties of sets in the exponential-time hierarchy. SIAM Journal on Computing, 17(3):504–516, June 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barak:2007:DC

[BOV07] Boaz Barak, Shien Jin Ong, and Salil Vadhan. Derandomization in cryptography. SIAM Journal on Computing, 37(2): 380–400, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brown:1981:ATA

[BP81] Cynthia A. Brown and Paul Walton Purdom, Jr. An average time analysis of backtracking. SIAM Journal on Computing, 10(3):583–593, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bilardi:1990:CAO

[BP90] G. Bilardi and F. P. Preparata.

Characterization of associative operations with prefix circuits of constant depth and linear size. SIAM Journal on Computing, 19(2):246–255, April 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bui:1992:PPG

[BP92] Thang Nguyen Bui and Andrew Peck. Partitioning planar graphs. SIAM Journal on Computing, 21(2):203–215, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bini:1993:IPP

[BP93a] Dario Bini and Victor Pan. Improved parallel polynomial division. SIAM Journal on Computing, 22(3):617–626, June 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bugrara:1993:ATA

[BP93b] Khaled M. Bugrara and Paul Walton Purdom, Jr. Average time analysis of clause order backtracking. SIAM Journal on Computing, 22(2): 303–317, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bafna:1996:GRS

[BP96] Vineet Bafna and Pavel A. Pevzner. Genome rearrangements and sorting by reversals. SIAM Journal on Computing, 25(2):272–289, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25062.

Bini:1998:CME

[BP98] Dario Bini and Victor Y. Pan. Computing matrix eigenvalues and polynomial zeros where the output is real. SIAM Journal on Computing, 27(4):1099–1115, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18248.

Boissonnat:2000:RPS

[BP00] Jean-Daniel Boissonnat and Franco P. Preparata. Robust plane sweep for intersecting segments. SIAM Journal on Computing, 29(5):1401– 1421, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32937.

Brodsky:2002:CWQ

[BP02] Alex Brodsky and Nicholas Pippenger. Characterizations of 1-way quantum finite automata. SIAM Journal on Computing, 31(5):1456– 1478, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35344.

Bansal:2010:SSB

[BP10] Nikhil Bansal and Kirk R.

Pruhs. Server scheduling to balance priorities, fairness, and average quality of service. SIAM Journal on Computing, 39(7):3311–3335, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bansal:2014:GS

[BP14] Nikhil Bansal and Kirk Pruhs.

The geometry of scheduling. SIAM Journal on Computing, 43(5):1684–1698, ???? 2014.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bitansky:2015:NBB

[BP15] Nir Bitansky and Omer Paneth. On non-black-box simulation and the impossibility of approximate obfuscation. SIAM Journal on Computing, 44(5):1325–1383, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barto:2020:TID

[BP20] Libor Barto and Michael Pinsker. Topology is irrelevant (in a dichotomy conjecture for infinite domain constraint satisfaction problems). SIAM Journal on Computing, 49(2):365–393, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bugrara:1989:EAT

[BPP89] Khaled M. Bugrara, Youfang Pan, and Paul Walton Purdom, Jr. Exponential average time for the pure literal rule. SIAM Journal on Computing, 18(2):409–418, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bressan:2023:SAL

[BPP23] Marco Bressan, Enoch Peserico, and Luca Pretto. Sublinear algorithms for local graph-centrality estimation.

SIAM Journal on Computing, 52(4):968–1008, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1266976.

Bonet:2000:IAF

[BPR00] Maria Luisa Bonet, Toniann Pitassi, and Ran Raz. On interpolation and automatization for Frege systems. SIAM Journal on Computing, 29(6):1939–1967, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35323.

Beame:2007:LBL

[BPS07] Paul Beame, Toniann Pitassi, and Nathan Segerlind. Lower bounds for Lovász–Schrijver systems and beyond follow from multiparty communication complexity. SIAM Journal on Computing, 37(3):845–869, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bansal:2010:SSW

[BPS10] Nikhil Bansal, Kirk Pruhs, and Cliff Stein. Speed scaling for weighted flow time. SIAM Journal on Computing, 39(4): 1294–1308, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Biskup:2004:SER

[BPSV04] Joachim Biskup, Jan Paredaens, Thomas Schwentick, and Jan Van den Bussche. Solving equations in the relational algebra. SIAM Journal on Computing, 33(5):1052–1066, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39085.

Bellantoni:1992:ASD

[BPU92] Stephen Bellantoni, Toniann Pitassi, and Alasdair Urquhart. Approximation and small-depth Frege proofs. SIAM Journal on Computing, 21(6):1161–1179, December 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bonet:1999:CET

[BPWY99] Maria Bonet, Cynthia Phillips, Tandy Warnow, and Shibu Yooseph. Constructing evolutionary trees in the presence of polymorphic characters. SIAM Journal on Computing, 29(1):103–131, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32463.

Bagchi:1979:OT

[BR79] A. Bagchi and J. K. Roy. On V -optimal trees. SIAM Journal on Computing, 8(4):524– 541, ???? 1979. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brand:1975:PTM

[Bra75] D. Brand. Proving theorems with the modification method. SIAM Journal on Computing, 4(4):412–430, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brady:1998:FDA

[Bra98] Martin L. Brady. A fast discrete approximation algorithm for the Radon transform. SIAM Journal on Computing, 27(1):107–119, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25667.

Brattka:2008:PRN

[Bra08] Vasco Brattka. Plottable real number functions and the computable graph theorem. SIAM Journal on Computing, 38(1): 303–328, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Braverman:2015:IIC

[Bra15] Mark Braverman. Interactive information complexity. SIAM Journal on Computing, 44(6): 1698–1739, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boissonnat:2019:ATD

[BRLW19] Jean-Daniel Boissonnat, Mael Rouxel-Labbé, and Mathijs H. M. J. Wintraecken. Anisotropic triangulations via discrete Riemannian Voronoi diagrams. SIAM Journal on Computing, 48(3):1046–1097, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brown:1978:IAB

[Bro78] Mark R. Brown. Implementation and analysis of binomial queue algorithms. SIAM Journal on Computing, 7(3):298– 319, ???? 1978. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brown:1979:PAR

[Bro79] Mark R. Brown. A partial analysis of random heightbalanced trees. SIAM Journal on Computing, 8(1):33–41, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brouwer:1999:ABD

[Bro99] A. E. Brouwer. An associative block design ABD(8, 5). SIAM Journal on Computing, 28(6):1970–1971, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31662.

Braverman:2014:PGR

[BRRY14] Mark Braverman, Anup Rao, Ran Raz, and Amir Yehudayoff. Pseudorandom generators for regular branching programs. SIAM Journal on Computing, 43(3):973–986, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blum:1997:NUG

[BRS97] Avrim Blum, Prabhakar Raghavan, and Baruch Schieber. Navigating in unfamiliar geometric terrain. SIAM Journal on Computing, 26(1):110– 137, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
19493.

Bradford:1998:EMC

[BRS98] Phillip G. Bradford, Gregory J. E. Rawlins, and Gregory E. Shannon. Efficient matrix chain ordering in polylog time. SIAM Journal on Computing, 27(2):466–490, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27069.

Baev:2008:AAD

[BRS08] Ivan Baev, Rajmohan Rajaraman, and Chaitanya Swamy. Approximation algorithms for data placement problems. SIAM Journal on Computing, 38(4):1411–1429, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Backurs:2021:TTA

[BRS^{+}21] Arturs Backurs, Liam Roditty, Gilad Segal, Virginia Vassilevska Williams, and Nicole Wein. Toward tight approximation bounds for graph diameter and eccentricities. SIAM Journal on Computing, 50(4): 1155–1199, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barcelo:2016:SAG

[BRV16] Pablo Barceló, Miguel Romero, and Moshe Y. Vardi. Semantic acyclicity on graph databases. SIAM Journal on Computing, 45(4):1339–1376, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Belkin:2018:EOD

[BRV18] Mikhail Belkin, Luis Rademacher, and James Voss. Eigenvectors of orthogonally decomposable functions. SIAM Journal on Computing, 47(2):547–615, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Barak:1976:PEB

[BS76] Amnon Barak and Eliahu Shamir. On the parallel evaluation of Boolean expressions. SIAM Journal on Computing, 5(4):678–681, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Burton:1982:HLA

[BS82] Robert P. Burton and David R.

Smith. A hidden-line algorithm for hyperspace. SIAM Journal on Computing, 11(1): 71–80, February 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Baker:1983:SAT

[BS83] Brenda S. Baker and Jerald S. Schwarz. Shelf algorithms for two-dimensional packing problems. SIAM Journal on Computing, 12(3):508– 525, ???? 1983. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blum:1986:ERF

[BS86] Lenore Blum and Michael Shub. Evaluating rational functions: infinite precision is finite cost and tractable on average. SIAM Journal on Computing, 15(2):384–398, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bruck:1992:PTF

[BS92] Jehoshua Bruck and Roman Smolensky. Polynomial threshold functions, AC^{0} functions, and spectral norms. SIAM Journal on Computing, 21(1):

33–42, February 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bollobas:1993:PAD

[BS93] Béla Bollobás and Istvan Simon. Probabilistic analysis of disjoint set union algorithms. SIAM Journal on Computing, 22(5):1053–1074, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bergman:2000:CSP

[BS00] Clifford Bergman and Giora Slutzki. Complexity of some problems concerning varieties and quasi-varieties of algebras. SIAM Journal on Computing, 30(2):359–382, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34594.

Braun:2003:PPS

[BS03] Oliver Braun and Günter Schmidt. Parallel processor scheduling with limited number of preemptions. SIAM Journal on Computing, 32(3): 671–680, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41069.

Ben-Sasson:2009:SST

[BS09] Eli Ben-Sasson. Size-space tradeoffs for resolution. SIAM Journal on Computing, 38(6): 2511–2525, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Belkin:2015:PLD

[BS15] Mikhail Belkin and Kaushik Sinha. Polynomial learning of distribution families. SIAM Journal on Computing, 44(4): 889–911, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Sasson:2006:RPP

[BSGH^{+}06] Eli Ben-Sasson, Oded Goldreich, Prahladh Harsha, Madhu Sudan, and Salil Vadhan. Robust PCPs of proximity, shorter PCPs, and applications to coding. SIAM Journal on Computing, 36(4):889–974, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Sasson:2010:LTC

[BSGK^{+}10] Eli Ben-Sasson, Venkatesan Guruswami, Tali Kaufman, Madhu Sudan, and Michael Viderman. Locally testable codes require redundant testers. SIAM Journal on Computing, 39(7):3230–3247, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bshouty:1989:LBM

[Bsh89] Nader H. Bshouty. A lower bound for matrix multiplication. SIAM Journal on Computing, 18(4):759–765, August 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bshouty:1990:MRT

[Bsh90] Nader H. Bshouty. Maximal rank of m × n × (mn − k) tensors. SIAM Journal on Computing, 19(3):467–471, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bagchi:1994:PID

[BSH94a] A. Bagchi, E. F. Schmeichel, and S. L. Hakimi. Parallel information dissemination by packets. SIAM Journal on Computing, 23(2):355–372, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/19267.

Bshouty:1994:CBF

[Bsh94b]  Nader H. Bshouty.  On

the complexity of bilinear forms over associative algebras. SIAM Journal on Computing, 23(4):815–833, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/19269.

Ben-Sasson:2005:SPH

[BSHR05] Eli Ben-Sasson, Prahladh Harsha, and Sofya Raskhodnikova. Some 3CNF properties are hard to test. SIAM Journal on Computing, 35(1):1–21, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Sasson:2012:ADS

[BSK12] Eli Ben-Sasson and Swastik Kopparty. Affine dispersers from subspace polynomials. SIAM Journal on Computing, 41(4):880–914, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Sasson:2014:SSF

[BSO14] Eli Ben-Sasson and Rafail Ostrovsky. Special section on the Fifty-Second IEEE Annual Symposium on Foundations of Computer Science (FOCS 2011). SIAM Journal on Computing, 43(2):654, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Sasson:2015:ATS

[BSRZ15] Eli Ben-Sasson and Noga Ron-Zewi. From affine to twosource extractors via approximate duality. SIAM Journal on Computing, 44(6):1670– 1697, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ben-Sasson:2008:SPP

[BSS08] Eli Ben-Sasson and Madhu Sudan. Short PCPs with polylog query complexity. SIAM Journal on Computing, 38(2):551– 607, ???? 2008. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Batson:2012:TRS

[BSS12] Joshua Batson, Daniel A. Spielman, and Nikhil Srivastava. Twice-Ramanujan sparsifiers. SIAM Journal on Computing, 41(6):1704–1721, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Babaioff:2014:CTM

[BSS14] Moshe Babaioff, Yogeshwer Sharma, and Aleksandrs Slivkins. Characterizing truthful multi-armed bandit mechanisms. SIAM Journal on Computing, 43(1):194–230, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bansal:2021:LRI

[BSS21] Nikhil Bansal, Aravind Srinivasan, and Ola Svensson. Liftand-round to improve weighted completion time on unrelated machines. SIAM Journal on Computing, 50(3):STOC16– 138–STOC16–159, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bent:1985:BST

[BST85] Samuel W. Bent, Daniel D. Sleator, and Robert E. Tarjan. Biased search trees. SIAM Journal on Computing, 14(3):

545–568, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buchsbaum:1995:DSB

[BST95] Adam L. Buchsbaum, Rajamani Sundar, and Robert E. Tarjan. Data-structural bootstrapping, linear path compression, and catenable heapordered double-ended queues. SIAM Journal on Computing, 24(6):1190–1206, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24214.

Baste:2023:HMB

[BST23] Julien Baste, Ignasi Sau, and Dimitrios M. Thilikos. Hitting minors on bounded treewidth graphs. IV. An optimal algorithm. SIAM Journal on Computing, 52(4):865–912, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M140482X.

Bezakova:2008:ASA

[BŠVV08] Ivona Bezáková, Daniel Štefankovič, Vijay V. Vazirani, and Eric Vigoda. Accelerating simulated annealing for the permanent and combinatorial counting problems. SIAM Journal on Computing, 37(5):1429– 1454, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bonsma:2014:CFA

[BSW14] Paul Bonsma, Jens Schulz, and Andreas Wiese. A constantfactor approximation algorithm for unsplittable flow on paths. SIAM Journal on Computing, 43(2):767–799, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bloom:1980:COM

[BT80a] Stephen L. Bloom and Ralph Tindell. Compatible orderings on the metric theory of trees. SIAM Journal on Computing, 9(4):683–691, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brent:1980:CCG

[BT80b] Richard P. Brent and Joseph F. Traub. On the complexity of composition and generalized composition of power series. SIAM Journal on Computing, 9(1):54–66, February 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brown:1980:DAD

[BT80c] Mark R. Brown and Robert E. Tarjan. Design and analysis of a data structure for representing sorted lists. SIAM Journal on Computing, 9(3):594– 614, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bergstra:1983:IFA

[BT83a] J. A. Bergstra and J. V.

Tucker. Initial and final algebra semantics for data type specifications: two characterization theorems. SIAM Journal on Computing, 12(2):366– 387, ???? 1983. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bloom:1983:VIT

[BT83b] Stephen L. Bloom and Ralph Tindell. Varieties of “If– Then–Else”. SIAM Journal on Computing, 12(4):677–707, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Braschi:1994:NIC

[BT94] Bertrand Braschi and Denis Trystram. A new insight into the Coffman–Graham algorithm. SIAM Journal on Computing, 23(3):662–669, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18188.

Buhrman:2000:RH

[BT00] Harry Buhrman and Leen Torenvliet. Randomness is hard. SIAM Journal on Computing, 30(5):1485–1501, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36014.

Bouchitte:2001:TMF

[BT01] Vincent Bouchitté and Ioan Todinca. Treewidth and minimum fill-in: Grouping the minimal separators. SIAM Journal on Computing, 31 (1):212–232, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35968.

Busch:2005:ALR

[BT05] Costas Busch and Srikanta Tirthapura. Analysis of link reversal routing algorithms. SIAM Journal on Computing, 35(2):305–326, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bogdanov:2006:WCA

[BT06] Andrej Bogdanov and Luca Trevisan. On worst-case to average-case reductions for NP problems. SIAM Journal on Computing, 36(4):1119–1159, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bravyi:2010:CSF

[BT10] Sergey Bravyi and Barbara Terhal. Complexity of stoquastic frustration-free Hamiltonians. SIAM Journal on Computing, 39(4):1462–1485, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blais:2015:ABF

[BT15] Eric Blais and Li-Yang Tan.

Approximating Boolean functions with depth-2 circuits. SIAM Journal on Computing, 44(6):1583–1600, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bun:2020:NOL

[BT20] Mark Bun and Justin Thaler.

A nearly optimal lower bound on the approximate degree of AC^{0}. SIAM Journal on Computing, 49(4):FOCS17– 59–FOCS17–96, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brazil:2000:MNU

[BTW00] M. Brazil, D. A. Thomas, and J. F. Weng. Minimum networks in uniform orientation metrics. SIAM Journal on Computing, 30(5):1579– 1593, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34719.

Beame:1994:CST

[BTY94] Paul Beame, Martin Tompa, and Peiyuan Yan. Communicationspace tradeoffs for unrestricted protocols. SIAM Journal on Computing, 23(3):652–661, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/19688.

Burgisser:2000:CCI

[Bür00a] Peter Bürgisser. The computational complexity of immanants. SIAM Journal on Computing, 30(3):1023–1040, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36788.

Burgisser:2000:CCE

[Bür00b] Peter Bürgisser. The computational complexity to evaluate representations of general linear groups. SIAM Journal on Computing, 30(3):1010–1022, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36789.

Bun:2018:FCP

[BUV18] Mark Bun, Jonathan Ullman, and Salil Vadhan. Fingerprinting codes and the price of approximate differential privacy. SIAM Journal on Computing, 47(5):1888–1938, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Beeri:1984:FST

[BV84] C. Beeri and M. Y. Vardi. Formal systems for tuple and equality generating dependencies. SIAM Journal on Computing, 13(1):76–98, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Berkman:1993:RST

[BV93] Omer Berkman and Uzi Vishkin. Recursive star-tree parallel data structure. SIAM Journal on Computing, 22(2): 221–242, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bergadano:1996:LBA

[BV96] Francesco Bergadano and Stefano Varricchio. Learning behaviors of automata from multiplicity and equivalence queries. SIAM Journal on Computing, 25(6):1268–1280, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26091.

Bernstein:1997:QCT

[BV97] Ethan Bernstein and Umesh Vazirani. Quantum complexity theory. SIAM Journal on Computing, 26(5):1411– 1473, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
30092.

Boldi:1999:CDS

[BV99] Paolo Boldi and Sebastiano Vigna. Complexity of deciding sense of direction. SIAM Journal on Computing, 29(3): 779–789, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        31080.

Beier:2006:TPW

[BV06] Rene Beier and Berthold Vöcking. Typical properties of winners and losers [0.2ex] in discrete optimization. SIAM Journal on Computing, 35(4): 855–881, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bogdanov:2010:PBP

[BV10] Andrej Bogdanov and Emanuele Viola. Pseudorandom bits for polynomials. SIAM Journal on Computing, 39(6):2464–2486, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Brakerski:2014:EFH

[BV14] Zvika Brakerski and Vinod Vaikuntanathan. Efficient fully homomorphic encryption from (standard) LWE. SIAM Journal on Computing, 43(2):831– 871, ???? 2014. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buhrman:2000:GRB

[BvMR^{+}00] Harry Buhrman, Dieter van Melkebeek, Kenneth W. Regan, D. Sivakumar, and Martin Strauss. A generalization of resource-bounded measure, with application to the BPP vs. EXP problem. SIAM Journal on Computing, 30(2):576– 601, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34389.

Bavarian:2022:APR

[BVY22] Mohammad Bavarian, Thomas Vidick, and Henry Yuen. Anchored parallel repetition for nonlocal games. SIAM Journal on Computing, 51(2):214–253, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1405927.

Bitner:1979:ONO

[BW79] J. R. Bitner and C. K.

Wong. Optimal and nearoptimal scheduling algorithms for batched processing in linear storage. SIAM Journal on Computing, 8(4):479–498, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bik:1999:ANS

[BW99] Aart J. C. Bik and Harry A. G. Wijshoff. Automatic nonzero structure analysis.

SIAM Journal on Computing, 28(5):1576–1587, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29595.

Beimel:2005:SPM

[BW05] Amos Beimel and Enav Weinreb. Separating the power of monotone span programs over different fields. SIAM Journal on Computing, 34 (5):1196–1215, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44403.

Boutsidis:2017:OCM

[BW17] Christos Boutsidis and David P. Woodruff. Optimal CUR matrix decompositions. SIAM Journal on Computing, 46(2): 543–589, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Book:1982:RTS

[BWX82] Ronald V. Book, Christopher B. Wilson, and Mei-Rui Xu. Relativizing time, space, and time-space. SIAM Journal on Computing, 11(3):571–581, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Blais:2015:PSF

[BWY15a] Eric Blais, Amit Weinstein, and Yuichi Yoshida. Partially symmetric functions are efficiently isomorphism testable. SIAM Journal on Computing, 44(2):411–432, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Boissonnat:2015:ADM

[BWY15b] Jean-Daniel Boissonnat, Camille Wormser, and Mariette Yvinec. Anisotropic Delaunay mesh generation. SIAM Journal on Computing, 44(2):467–512, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Balas:1991:MWC

[BX91] Egon Balas and Jue Xue. Minimum weighted coloring of triangulated graphs, with application to maximum weight vertex packing and clique finding in arbitrary graphs. SIAM Journal on Computing, 20(2): 209–221, April 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See addendum [BX92].

Balas:1992:AMW

[BX92] Egon Balas and Jue Xue. Addendum: “Minimum weighted coloring of triangulated graphs, with application to maximum weight vertex packing and clique finding in arbitrary graphs” [SIAM J. Comput. 20 (1991), no. 2, 209–221, MR 92h:68069]. SIAM Journal on Computing, 21(5):1000, October 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [BX91].

Balas:1986:FMC

[BY86] Egon Balas and Chang Sung Yu. Finding a maximum clique in an arbitrary graph. SIAM Journal on Computing, 15(4): 1054–1068, November 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Buss:1998:LTM

[BY98] Samuel R. Buss and Peter N. Yianilos. Linear and O(n log n) time minimum-cost matching algorithms for quasi-convex tours. SIAM Journal on Computing, 27(1):170–201, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26724.

Baeza-Yates:1992:ASI

[BYCDM92] R. Baeza-Yates, R. Casas, J. Díaz, and C. Martínez. On the average size of the intersection of binary trees. SIAM Journal on Computing, 21(1): 24–32, February 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bar-Yehuda:1998:AAF

[BYGNR98] Reuven Bar-Yehuda, Dan Geiger, Joseph (Seffi) Naor, and Ron M. Roth. Approximation algorithms for the feedback vertex set problem with applications to constraint satisfaction and Bayesian inference. SIAM Journal on Computing, 27(4):942–959, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30510.

Bar-Yehuda:2006:SSI

[BYHN^{+}06] R. Bar-Yehuda, M. M. Halldórsson, J. (S. ) Naor, H. Shachnai, and I. Shapira. Scheduling split intervals. SIAM Journal on Computing, 36(1):1–15, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bar-Yehuda:1993:MCM

[BYII93] Reuven Bar-Yehuda, Amos Israeli, and Alon Itai. Multiple communication in multihop radio networks. SIAM Journal on Computing, 22(4):875–887, August 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Bar-Yossef:2008:ESQ

[BYJK08] Ziv Bar-Yossef, T. S. Jayram, and Iordanis Kerenidis. Exponential separation of quantum and classical one-way communication complexity. SIAM Journal on Computing, 38(1): 366–384, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Czumaj:2000:CRH

[CadHS00] Artur Czumaj, Friedhelm Meyer auf der Heide, and Volker Stemann. Contention resolution in hashing based shared memory simulations. SIAM Journal on Computing, 29 (5):1703–1739, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29564.

Cohen-Addad:2021:NLA

[CAdVdM21] Vincent Cohen-Addad, Éric Colin de Verdière, and Arnaud de Mesmay. A near-linear approximation scheme for multicuts of embedded graphs with a fixed number of terminals. SIAM Journal on Computing, 50(1):1–31, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen-Addad:2019:LSY

[CAKM19] Vincent Cohen-Addad, Philip N. Klein, and Claire Mathieu. Local search yields approximation schemes for k-means and k-median in Euclidean and minor-free metrics. SIAM Journal on Computing, 48(2): 644–667, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cam:2003:RSS

[Çam03] Hasan Çam. Rearrangeability of (2n − 1)-stage shuffleexchange networks. SIAM Journal on Computing, 32(3): 557–585, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34484.

Carlson:1979:CRP

[Car79] Carl R. Carlson. A counterexample to Reingold’s pushdown permuter characterization theorem. SIAM Journal on Computing, 8(2):199–201, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cartwright:1984:RPD

[Car84] Robert Cartwright. Recursive programs as definitions in firstorder logic. SIAM Journal on Computing, 13(2):374–408, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Case:1999:PVL

[Cas99] John Case. The power of vacillation in language learning. SIAM Journal on Computing, 28(6):1941–1969, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24969.

Chua:1974:AFS

[CB74] Y. S. Chua and A. J. Bernstein. Analysis of a feedback scheduler. SIAM Journal on Computing, 3(3):159–176, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Colbourn:1981:LTA

[CB81] Charles J. Colbourn and Kellogg S. Booth. Linear time automorphism algorithms for trees, interval graphs, and planar graphs. SIAM Journal on Computing, 10(1):203–225, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cesa-Bianchi:2005:SOP

[CBCG05] Nicolò Cesa-Bianchi, Alex Conconi, and Claudio Gentile. A second-order perceptron algorithm. SIAM Journal on Computing, 34(3):640–668, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43254.

Charron-Bost:2013:LRR

[CBGWW13] Bernadette Charron-Bost, Antoine Gaillard, Jennifer L. Welch, and Josef Widder. Link reversal routing with binary link labels: Work complexity. SIAM Journal on Computing, 42(2):634–661, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chiu:1984:OCQ

[CBH84] Dah Ming Chiu, Philip A.

Bernstein, and Yu Chi Ho. Optimizing chain queries in a distributed database system. SIAM Journal on Computing, 13(1):116–134, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chou:1982:ASS

[CC82] Tsu Wu J. Chou and George E.

Collins. Algorithms for the solution of systems of linear Diophantine equations. SIAM Journal on Computing, 11(4): 687–708, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Choffrut:1983:PFP

[CC83] Christian Choffrut and Karel Culik II. Properties of finite and pushdown transducers. SIAM Journal on Computing, 12(2):300–315, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cabay:1986:ACS

[CC86] Stanley Cabay and Dong-Koo Choi. Algebraic computations of scaled Padé fractions. SIAM Journal on Computing, 15(1): 243–270, February 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:1997:ANP

[CC97] Liming Cai and Jianer Chen. On the amount of nondeterminism and the power of verifying. SIAM Journal on Computing, 26(3):733–750, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25829.

Chen:1999:MPM

[CC99] Chi-Chang Chen and Jianer Chen. The maximum partition matching problem with applications. SIAM Journal on Computing, 28(3):935–954, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30012.

Chang:2007:DMS

[CC07] Guey-Yun Chang and Gen-Huey Chen. (t,k)-diagnosability of multiprocessor systems with applications to grids and tori. SIAM Journal on Computing, 37(4):1280–1298, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coy:2023:DMP

[CC23] Sam Coy and Artur Czumaj. Deterministic massively parallel connectivity. SIAM Journal on Computing, 52(5):1269– 1318, October 2023. CODEN SMJCAT. ISSN 1095-7111.

Chan:2009:LTH

[CCC^{+}09] Mee Yee Chan, Wun-Tat Chan, Francis Y. L. Chin, Stanley P. Y. Fung, and Ming-Yang Kao. Linear-time haplotype inference on pedigrees without recombinations and mating loops. SIAM Journal on Computing, 38(6):2179– 2197, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:2013:ATC

[CCC^{+}13] David A. Cohen, Martin C.

Cooper, Páidí Creed, Peter G. Jeavons, and Stanislav Zivný. An algebraic theory of complexity for discrete optimization. SIAM Journal on Computing, 42(5):1915–1939, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cabello:2013:MSS

[CCE13] Sergio Cabello, Erin W.

Chambers, and Jeff Erickson. Multiple-source shortest paths in embedded graphs. SIAM Journal on Computing, 42(4): 1542–1571, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Charikar:2004:ICD

[CCFM04] Moses Charikar, Chandra Chekuri, Tomas Feder, and Rajeev Motwani. Incremental clustering and dynamic information retrieval. SIAM Journal on Computing, 33(6): 1417–1440, December 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41849.

Chandrasekaran:1997:NLS

[CCG^{+}97] R. Chandrasekaran, Bo Chen, Gábor Galambos, P. R. Narayanan, André Van Vliet, and Gerhard J. Woeginger. A note on: “An on-line scheduling heuristic with better worst case ratio than Graham’s list scheduling” [SIAM J. Comput. 22 (1993), no. 2, 349–355; MR 94b:90031] by Galambos and Woeginger. SIAM Journal on Computing, 26(3):870–872, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25877.  See

[GW93a].

Cai:1998:CBF

[CCH98] Liming Cai, Jianer Chen, and Johan Håstad. Circuit bottom fan-in and computational power. SIAM Journal on Computing, 27(2):341–355, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28243.

Chitnis:2016:DFA

[CCH^{+}16] Rajesh Chitnis, Marek Cygan, MohammadTaghi Hajiaghayi, Marcin Pilipczuk, and Michal Pilipczuk. Designing FPT algorithms for cut problems using randomized contractions. SIAM Journal on Computing, 45(4):1171–1229, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cabello:2017:PPN

[CCK^{+}17] Sergio Cabello, Josef Cibulka, Jan Kyncl, Maria Saumell, and Pavel Valtr. Peeling potatoes near-optimally in near-linear time. SIAM Journal on Computing, 46(5):1574–1602, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chalermsook:2020:GET

[CCK^{+}20] Parinya Chalermsook, Marek Cygan, Guy Kortsarz, Bundit Laekhanukit, Pasin Manurangsi, Danupon Nanongkai, and Luca Trevisan. From gapexponential time hypothesis to fixed parameter tractable inapproximability: Clique, dominating set, and more. SIAM Journal on Computing, 49(4): 772–810, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2022:MOT

[CCK^{+}22] Timothy F. Chan, Jacob W.

Cooper, Martin Koutecký, Daniel Král, and Kristýna Pekárková. Matrices of optimal tree-depth and a rowinvariant parameterized algorithm for integer programming. SIAM Journal on Computing, 51(3):664–700, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1353502.

Chakrabarti:2013:ICT

[CCKM13] Amit Chakrabarti, Graham Cormode, Ranganath Kondapally, and Andrew McGregor. Information cost tradeoffs for augmented index and streaming language recognition. SIAM Journal on Computing, 42(1):61–83, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:2013:GHC

[CCL13] Jin-Yi Cai, Xi Chen, and Pinyan Lu. Graph homomorphisms with complex values: a dichotomy theorem. SIAM Journal on Computing, 42(3): 924–1029, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:2016:NWC

[CCL16] Jin-Yi Cai, Xi Chen, and Pinyan Lu. Nonnegative weighted #CSP: an effective complexity dichotomy. SIAM Journal on Computing, 45(6): 2177–2198, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chakrabarti:2019:VSC

[CCM^{+}19] Amit Chakrabarti, Graham Cormode, Andrew McGregor, Justin Thaler, and Suresh Venkatasubramanian. Verifiable stream computation and Arthur–Merlin communication. SIAM Journal on Computing, 48(4):1265–1299, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Calinescu:2011:MMS

[CCPV11] Gruia Calinescu, Chandra Chekuri, Martin Pál, and Jan Vondrák. Maximizing a monotone submodular function subject to a matroid constraint. SIAM Journal on Computing, 40(6):1740–1766, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i6/p1740_s1.

Citrini:1986:DMP

[CCRM86] Claudio Citrini, Stefano Crespi-Reghizzi, and Dino Mandrioli. On deterministic multipass analysis. SIAM Journal on Computing, 15(3):668–693, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chattopadhyay:2022:SSF

[CCRZWN22] Arkadev Chattopadhyay, Marek Cygan, Noga Ron-Zewi, and Christian Wulff-Nilsen. Special section on the Fifty-Second Annual ACM Symposium on the Theory of Computing (STOC 2020). SIAM Journal on Computing, 51(3): STOC20–i–STOC20–i, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975494.

Chan:2018:RAG

[CCWZ18] T-H. Hubert Chan, Fei Chen, Xiaowei Wu, and Zhichao Zhao. Ranking on arbitrary graphs: Rematch via continuous linear programming. SIAM Journal on Computing, 47(4): 1529–1546, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Constable:1972:CPS

[CD72] Robert L. Constable and Gries David. On classes of program schemata. SIAM Journal on Computing, 1(1):66–118, ???? 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheng:2003:QMW

[CD03] Siu-Wing Cheng and Tamal K. Dey. Quality meshing with weighted Delaunay refinement. SIAM Journal on Computing, 33(1):69–93, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41880.

Chen:2016:DQC

[CD16] Hubie Chen and Víctor Dalmau. Decomposing quantified conjunctive (or disjunctive) formulas. SIAM Journal on Computing, 45(6):2066– 2086, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2017:PTH

[CD17] Ho-Lin Chen and David Doty.

Parallelism and time in hierarchical self-assembly. SIAM Journal on Computing, 46(2): 661–709, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Czumaj:2018:DCR

[CD18] Artur Czumaj and Peter Davies. Deterministic communication in radio networks. SIAM Journal on Computing, 47(1):218–240, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coan:1989:DFS

[CDDS89] Brian A. Coan, Danny Dolev, Cynthia Dwork, and Larry Stockmeyer. The distributed firing squad problem. SIAM Journal on Computing, 18(5): 990–1012, October 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cryan:2006:RMM

[CDG^{+}06] Mary Cryan, Martin Dyer, Leslie Ann Goldberg, Mark Jerrum, and Russell Martin. Rapidly mixing Markov chains for sampling contingency tables with a constant number of rows. SIAM Journal on Computing, 36(1):247–278, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2009:MER

[CDG^{+}09] T.-H. Hubert Chan, Kedar Dhamdhere, Anupam Gupta, Jon Kleinberg, and Aleksandrs Slivkins. Metric embeddings with relaxed guarantees. SIAM Journal on Computing, 38(6): 2303–2329, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chawla:2011:SSF

[CDG11] Shuchi Chawla, Cynthia Dwork, and Venkat Guruswami. Special section on the Fortieth Annual ACM Symposium On Theory Of Computing (STOC 2008). SIAM Journal on Computing, 40(6):1738, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i6/p1738_s1.

Corneil:2013:LBC

[CDH13] Derek G. Corneil, Barnaby Dalton, and Michel Habib.

LDFS-based certifying algorithm for the minimum path cover problem on cocomparability graphs. SIAM Journal on Computing, 42(3):792–807, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See corrigendum [DCH21].

Cohen:2011:ESS

[CDK^{+}11] Edith Cohen, Nick Duffield, Haim Kaplan, Carsten Lund, and Mikkel Thorup. Efficient stream sampling for varianceoptimal estimation of subset sums. SIAM Journal on Computing, 40(5):1402–1431, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i5/
p1402_s1.

Chee:2020:RPP

[CDK^{+}20] Yeow Meng Chee, Duc Tu Dao, Han Mao Kiah, San Ling, and Hengjia Wei. Robust positioning patterns with low redundancy. SIAM Journal on Computing, 49(2):284–317, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chazelle:1986:CLE

[CDL86] B. Chazelle, R. L. Drysdale, III, and D. T. Lee. Computing the largest empty rectangle. SIAM Journal on Computing, 15(1):300–315, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Clarkson:2016:FCT

[CDMI^{+}16] Kenneth L. Clarkson, Petros Drineas, Malik Magdon-Ismail, Michael W. Mahoney, Xiangrui Meng, and David P. Woodruff. The Fast Cauchy Transform and faster robust linear regression. SIAM Journal on Computing, 45(3):763– 810, ???? 2016. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2022:COL

[CDO^{+}22] Xi Chen, Ilias Diakonikolas, Anthi Orfanou, Dimitris Paparas, Xiaorui Sun, and Mihalis Yannakakis. On the complexity of optimal lottery pricing and randomized mechanisms for a unit-demand buyer. SIAM Journal on Computing, 51(3):492–548, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/17M1136481.

Czumaj:2021:SDC

[CDP21] Artur Czumaj, Peter Davies, and Merav Parter. Simple, deterministic, constant-round coloring in congested clique and MPC. SIAM Journal on Computing, 50(5):1603–1626, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cook:1986:ULT

[CDR86] Stephen Cook, Cynthia Dwork, and Rüdiger Reischuk. Upper and lower time bounds for parallel random access machines without simultaneous writes. SIAM Journal on Computing, 15(1):87–97, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cryan:2010:ACI

[CDR10] Mary Cryan, Martin Dyer, and Dana Randall. Approximately counting integral flows and cell-bounded contingency tables. SIAM Journal on Computing, 39(7):2683–2703, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheng:2007:SMS

[CDRR07] Siu-Wing Cheng, Tamal K.

Dey, Edgar A. Ramos, and Tathagata Ray. Sampling and meshing a surface with guaranteed topology and geometry. SIAM Journal on Computing, 37(4):1199–1227, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:1995:DGD

[CDTT95] Robert F. Cohen, Giuseppe Di Battista, Roberto Tamassia, and Ioannis G. Tollis. Dynamic graph drawings: Trees, series-parallel digraphs, and planar ST-Digraphs. SIAM Journal on Computing, 24 (5):970–1001, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23572.

Cai:2021:DBU

[CDW21] Yang Cai, Nikhil R. Devanur, and S. Matthew Weinberg. A duality-based unified approach to Bayesian mechanism design. SIAM Journal on Computing, 50(3):STOC16–160–STOC16– 200, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chattopadhyay:2016:ULB

[CEEP16] Arkadev Chattopadhyay, Jeff Edmonds, Faith Ellen, and Toniann Pitassi. Upper and lower bounds on the power of advice. SIAM Journal on Computing, 45(4):1412–1432, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Czumaj:2005:AWE

[CEF^{+}05] Artur Czumaj, Funda Ergün, Lance Fortnow, Avner Magen, Ilan Newman, Ronitt Rubinfeld, and Christian Sohler. Approximating the weight of the Euclidean minimum spanning tree in sublinear time. SIAM Journal on Computing, 35(1): 91–109, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chambers:2023:MCS

[CEFN23] Erin W. Chambers, Jeff Erickson, Kyle Fox, and Amir Nayyeri. Minimum cuts in surface graphs. SIAM Journal on Computing, 52(1):156–195, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1291820.

Chazelle:1993:CFA

[CEG^{+}93] Bernard Chazelle, Herbert Edelsbrunner, Leonidas Guibas, Micha Sharir, and Jack Snoeyink. Computing a face in an arrangement of line segments and related problems. SIAM Journal on Computing, 22(6):1286–1302, December 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chazelle:1994:SHC

[CEG^{+}94] Bernard Chazelle, Herbert Edelsbrunner, Leonidas J. Guibas, John E. Hershberger, Raimund Seidel, and Micha Sharir. Selecting heavily covered points. SIAM Journal on Computing, 23(6): 1138–1151, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/17991.

Chakrabarty:2018:OBB

[CEKP18] Deeparnab Chakrabarty, Alina Ene, Ravishankar Krishnaswamy, and Debmalya Panigrahi. Online buy-at-bulk network design. SIAM Journal on Computing, 47(4):1505–1528, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chambers:2012:HFC

[CEN12] Erin W. Chambers, Jeff Erickson, and Amir Nayyeri. Homology flows, cohomology cuts. SIAM Journal on Computing, 41(6):1605–1634, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chao:1986:PAT

[CF86] Ming-Te Chao and John Franco. Probabilistic analysis of two heuristics for the 3satisfiability problem. SIAM Journal on Computing, 15(4): 1106–1118, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:2022:HAM

[CF22] Jin-Yi Cai and Zhiguo Fu. Holographic algorithm with matchgates is universal for planar #CSP over Boolean domain. SIAM Journal on Computing, 51(2):STOC17– 50–STOC17–151, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1131672.

Cole:2000:AMA

[CFCH^{+}00] Richard Cole, Martin Farach-Colton, Ramesh Hariharan, Teresa Przytycka, and Mikkel Thorup. An O(n log n) algorithm for the maximum agreement subtree problem for binary trees. SIAM Journal on Computing, 30(5):1385– 1404, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
31347.

Cohen:2012:EFM

[CFF^{+}12] Edith Cohen, Michal Feldman, Amos Fiat, Haim Kaplan, and Svetlana Olonetsky. Envyfree makespan approximation. SIAM Journal on Computing, 41(1):12–25, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v41/i1/p12_s1.

Chakraborty:2016:PCS

[CFGM16] Sourav Chakraborty, Eldar Fischer, Yonatan Goldhirsh, and Arie Matsliah. On the power of conditional samples in distribution testing. SIAM Journal on Computing, 45(4): 1261–1296, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chitnis:2020:TBP

[CFHM20] Rajesh H. Chitnis, Andreas E.

Feldmann, MohammadTaghi HajiAghayi, and Daniel Marx. Tight bounds for planar strongly connected Steiner subgraph with fixed number of terminals (and extensions). SIAM Journal on Computing, 49(2):318–364, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cardinal:2010:EAP

[CFJ^{+}10] Jean Cardinal, Samuel Fiorini, Gwenaël Joret, Raphaël M. Jungers, and J. Ian Munro. An efficient algorithm for partial order production. SIAM Journal on Computing, 39(7): 2927–2940, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2007:OCF

[CFK^{+}07] Ke Chen, Amos Fiat, Haim Kaplan, Meital Levy, Ji í Matoušek, Elchanan Mossel, János Pach, Micha Sharir, Shakhar Smorodinsky, Uli Wagner, and Emo Welzl. Online conflict-free coloring for intervals. SIAM Journal on Computing, 36(5):1342–1359, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chattopadhyay:2021:QCL

[CFK^{+}21] Arkadev Chattopadhyay, Yuval Filmus, Sajin Koroth, Or Meir, and Toniann Pitassi. Query-to-communication lifting using low-discrepancy gadgets. SIAM Journal on Computing, 50(1):171–210, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2007:PDK

[CFKX07] Jianer Chen, Henning Fernau, Iyad A. Kanj, and Ge Xia. Parametric duality and kernelization: Lower bounds and upper bounds on kernel size. SIAM Journal on Computing, 37(4):1077–1106, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Condon:1997:RDH

[CFLS97] Anne Condon, Joan Feigenbaum, Carsten Lund, and Peter Shor. Random debaters and the hardness of approximating stochastic functions. SIAM Journal on Computing, 26(2):369–400, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26073.

Chan:2011:DBN

[CFLY11] Yuk Hei Chan, Wai Shing Fung, Lap Chi Lau, and Chun Kong Yung. Degree bounded network design with metric costs. SIAM Journal on Computing, 40(4):953–980, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i4/
                                        p953_s1.

Cieliebak:2012:DCM

[CFPS12] Mark Cieliebak, Paola Flocchini, Giuseppe Prencipe, and Nicola Santoro. Distributed computing by mobile robots: Gathering. SIAM Journal on Computing, 41(4):829–879, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chvatal:2002:RDF

[CFSZ02] V. Chvátal, J. Fonlupt, L. Sun, and A. Zemirline. Recognizing dart-free perfect graphs. SIAM Journal on Computing, 31(5):1315–1338, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35477.

Coffman:1993:OSA

[CFW93a] E. G. Coffman, Jr., Leopold Flatto, and Paul E. Wright. Optimal stochastic allocation of machines under waitingtime constraints. SIAM Journal on Computing, 22(2):332– 348, April 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1993:SCO

[CFW93b] E. G. Coffman, Jr., Leopold Flatto, and Paul E. Wright. A stochastic checkpoint optimization problem. SIAM Journal on Computing, 22(3):650– 659, June 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Corneil:1973:ADC

[CG73] D. G. Corneil and B. Graham. An algorithm for determining the chromatic number of a graph. SIAM Journal on Computing, 2(4):311–318, December 1973. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chor:1988:UBS

[CG88] Benny Chor and Oded Goldreich. Unbiased bits from sources of weak randomness and probabilistic communication complexity. SIAM Journal on Computing, 17(2):230– 261, April 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Cypher:1994:RDF

[CG94] Robert Cypher and Luis Gravano. Requirements for deadlock-free, adaptive packet routing. SIAM Journal on Computing, 23(6):1266–1274, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23183.

Cucker:1997:PRT

[CG97] Felipe Cucker and Dima Grigoriev. On the power of real Turing machines over binary inputs. SIAM Journal on Computing, 26(1):243– 254, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
27034.

Cucker:2001:TNS

[CG01] Felipe Cucker and Dima Grigoriev. There are no sparse NP_{w}-hard sets. SIAM Journal on Computing, 31(1):193– 198, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37934.

Charikar:2005:ICA

[CG05] Moses Charikar and Sudipto Guha. Improved combinatorial algorithms for facility location problems. SIAM Journal on Computing, 34(4):803– 824, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39859.

Chen:2007:IBS

[CG07a] Yijia Chen and Martin Grohe. An isomorphism between subexponential and parameterized complexity theory. SIAM Journal on Computing, 37(4):1228–1258, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Correa:2007:IBN

[CG07b] José R. Correa and Michel X.

Goemans. Improved bounds on nonblocking 3-stage Clos networks. SIAM Journal on Computing, 37(3):870–894, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chakrabarty:2010:ABA

[CG10] Deeparnab Chakrabarty and Gagan Goel. On the approximability of budgeted allocations and improved lower bounds for submodular welfare maximization and GAP. SIAM Journal on Computing, 39(6): 2189–2211, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2012:ATM

[CG12] T.-H. Hubert Chan and Anupam Gupta. Approximating TSP on metrics with bounded global growth. SIAM Journal on Computing, 41(3):587–617, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cole:2018:ANS

[CG18] Richard Cole and Vasilis Gkatzelis. Approximating the Nash social welfare with indivisible items. SIAM Journal on Computing, 47(3):1211–1236, ???? 2018. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2013:ECD

[CGdW13] Victor Chen, Elena Grigorescu, and Ronald de Wolf. Error-correcting data structures. SIAM Journal on Computing, 42(1):84–111, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Crochemore:1997:CTR

[CGG^{+}97] Maxime Crochemore, Zvi Galil, Leszek Ga̧sieniec, Kunsoo Park, and Wojciech Rytter. Constant-time randomized parallel string matching. SIAM Journal on Computing, 26(4):950–960, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28007.

Cryan:2001:ETC

[CGG01] Mary Cryan, Leslie Ann Goldberg, and Paul W. Goldberg. Evolutionary trees can be learned in polynomial time in the two-state general Markov model. SIAM Journal on Computing, 31(2):375–397, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34249.

Chor:1995:PCI

[CGGK95] Benny Chor, Mihály Geréb-Graus, and Eyal Kushilevitz. Private computations over the integers. SIAM Journal on Computing, 24(2):376–386, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19499.

Creus:2016:HPE

[CGGR16] Carles Creus, Adrià Gascón, Guillem Godoy, and Lander Ramos. The HOM problem is EXPTIME-complete. SIAM Journal on Computing, 45(4): 1230–1260, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Christodoulou:2019:PSW

[CGGS19] George Christodoulou, Martin Gairing, Yiannis Giannakopoulos, and Paul G. Spirakis. The price of stability of weighted congestion games. SIAM Journal on Computing, 48(5):1544–1582, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:1988:BHS

[CGH^{+}88] Jin-Yi Cai, Thomas Gundermann, Juris Hartmanis, Lane A. Hemachandra, Vivian Sewelson, Klaus Wagner, and Gerd Wechsung. The Boolean hierarchy. I. structural properties. SIAM Journal on Computing, 17(6):1232–1252, December 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:1989:BHI

[CGH^{+}89] Jin-Yi Cai, Thomas Gundermann, Juris Hartmanis, Lane A. Hemachandra, Vivian Sewelson, Klaus Wagner, and Gerd Wechsung. The Boolean hierarchy. II. applications. SIAM Journal on Computing, 18(1):95–111, February 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Crochemore:1998:CTO

[CGH^{+}98] Maxime Crochemore, Leszek Gasieniec, Ramesh Hariharan, S. Muthukrishnan, and Wojciech Rytter. A constant time optimal parallel algorithm for two-dimensional pattern matching. SIAM Journal on Computing, 27(3):668–681, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28006.

Chandrasekaran:2013:DAL

[CGH13] Karthekeyan Chandrasekaran, Navin Goyal, and Bernhard Haeupler. Deterministic algorithms for the Lovász local lemma. SIAM Journal on Computing, 42(6):2132–2155, ???? 2013. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1978:APM

[CGJ78] E. G. Coffman, Jr., M. R.

Garey, and D. S. Johnson. An application of bin-packing to multiprocessor scheduling. SIAM Journal on Computing, 7(1):1–17, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1983:DP

[CGJ83] E. G. Coffman, Jr., M. R.

Garey, and D. S. Johnson. Dynamic bin packing. SIAM Journal on Computing, 12(2): 227–258, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1985:SFT

[CGJL85] E. G. Coffman, Jr., M. R.

Garey, D. S. Johnson, and A. S. Lapaugh. Scheduling file transfers. SIAM Journal on Computing, 14(3):744–780, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1980:PBL

[CGJT80] E. G. Coffman, Jr., M. R.

Garey, D. S. Johnson, and R. E. Tarjan. Performance bounds for level-oriented twodimensional packing algorithms. SIAM Journal on Computing, 9(4):808–826, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chrobak:2007:WPM

[CGK07] Marek Chrobak, Leszek Gasieniec, and Dariusz R. Kowalski. The wake-up problem in multihop radio networks. SIAM Journal on Computing, 36(5): 1453–1471, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:1997:BQA

[CGL97] Richard Chang, William I. Gasarch, and Carsten Lund. On bounded queries and approximation. SIAM Journal on Computing, 26(1):188– 209, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        26648.

Chattopadhyay:2020:NEC

[CGL20] Eshan Chattopadhyay, Vipul Goyal, and Xin Li. Nonmalleable extractors and codes, with their many tampered extensions. SIAM Journal on Computing, 49(5):999–1040, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Conte:2022:PSM

[CGM^{+}22] Alessio Conte, Roberto Grossi, Andrea Marino, Takeaki Uno, and Luca Versari. Proximity search for maximal subgraph enumeration. SIAM Journal on Computing, 51 (5):1580–1625, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1375048.

Chandran:2014:PBC

[CGMO14] Nishanth Chandran, Vipul Goyal, Ryan Moriarty, and Rafail Ostrovsky. Positionbased cryptography. SIAM Journal on Computing, 43(4): 1291–1341, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chu:2020:GSS

[CGP^{+}20] Timothy Chu, Yu Gao, Richard Peng, Sushant Sachdeva, Saurabh Sawlani, and Junxing Wang. Graph sparsification, spectral sketches, and faster resistance computation via short cycle decompositions. SIAM Journal on Computing, 52 (6):FOCS18–85–FOCS18–157, June 2020. CODEN SMJCAT. ISSN 1095-7111.

Chandran:2012:TCL

[CGR12] Harish Chandran, Nikhil Gopalkrishnan, and John Reif. Tile complexity of linear assemblies. SIAM Journal on Computing, 41(4):1051–1073, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Canonne:2018:SC

[CGR18] Clément L. Canonne, Themis Gouleakis, and Ronitt Rubinfeld. Sampling correctors. SIAM Journal on Computing, 47(4):1373–1423, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cherkassky:1999:BHL

[CGS99] Boris V. Cherkassky, Andrew V. Goldberg, and Craig Silverstein. Buckets, heaps, lists, and monotone priority queues. SIAM Journal on Computing, 28(4):1326– 1346, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        31349.

Chiesa:2022:RLC

[CGS22] Alessandro Chiesa, Tom Gur, and Igor Shinkar. Relaxed locally correctable codes with nearly-linear block length and constant query complexity. SIAM Journal on Computing, 51(6):1839–1865, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M135515X.

Chen:2011:OEF

[CGV11] Ning Chen, Arpita Ghosh, and Sergei Vassilvitskii. Optimal envy-free pricing with metric substitutability. SIAM Journal on Computing, 40(3): 623–645, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i3/p623_s1.

Cheraghchi:2013:RIF

[CGV13] Mahdi Cheraghchi, Venkatesan Guruswami, and Ameya Velingker. Restricted isometry of Fourier matrices and list decodability of random linear codes. SIAM Journal on Computing, 42(5):1888–1914, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:2016:CDR

[CGW16] Jin-Yi Cai, Heng Guo, and Tyson Williams. A complete dichotomy rises from the capture of vanishing signatures. SIAM Journal on Computing, 45(5):1671–1728, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2006:ALC

[CGYZ06] Guantao Chen, Zhicheng Gao, Xingxing Yu, and Wenan Zang. Approximating longest cycles in graphs with bounded degrees. SIAM Journal on Computing, 36(3):635–656, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1982:EPS

[CH82a] E. G. Coffman, Jr. and Micha Hofri. On the expected performance of scanning disks. SIAM Journal on Computing, 11(1): 60–70, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cole:1982:ECB

[CH82b] Richard Cole and John Hopcroft. On edge coloring bipartite graphs. SIAM Journal on Computing, 11(3):540–546, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cook:1985:DUC

[CH85] Stephen A. Cook and H. James Hoover. A depth-universal circuit. SIAM Journal on Computing, 14(4):833–839, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Choi:1987:SFT

[CH87] Hyeong-Ah Choi and S. Louis Hakimi. Scheduling file transfers for trees and odd cycles. SIAM Journal on Computing, 16(1):162–168, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheriyan:1995:RMF

[CH95] Joseph Cheriyan and Torben Hagerup. A randomized maximum-flow algorithm.

SIAM Journal on Computing, 24(2):203–226, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22152.

Cole:1997:TUB

[CH97] Richard Cole and Ramesh Hariharan. Tighter upper bounds on the exact complexity of string matching. SIAM Journal on Computing, 26(3):803–856, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24694.

Cole:2002:ASM

[CH02] Richard Cole and Ramesh Hariharan. Approximate string matching: a simpler faster algorithm. SIAM Journal on Computing, 31(6):1761–1782, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37052.

Chern:2003:PMQ

[CH03a] Hua-Huai Chern and Hsien-Kuei Hwang. Partial match queries in random quadtrees. SIAM Journal on Computing, 32(4):904–915, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41213.

Cole:2003:FST

[CH03b] Richard Cole and Ramesh Hariharan. Faster suffix tree construction with missing suffix links. SIAM Journal on Computing, 33(1):26–42, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42446.

Cole:2003:TPM

[CH03c] Richard Cole and Ramesh Hariharan. Tree pattern matching to subset matching in linear time. SIAM Journal on Computing, 32(4):1056–1066, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38270.

Cole:2005:DLQ

[CH05] Richard Cole and Ramesh Hariharan. Dynamic LCA queries on trees. SIAM Journal on Computing, 34(4):894– 923, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        37053.

Chern:2006:PMQ

[CH06] Hua-Huai Chern and Hsien-Kuei Hwang. Partial match queries in random k-d trees.

SIAM Journal on Computing, 35(6):1440–1466, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chase:1973:TG

[Cha73] Phillip J. Chase. Transposition graphs. SIAM Journal on Computing, 2(2):128–133, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chandra:1976:GPS

[Cha76] Ashok K. Chandra. Generalized program schemas. SIAM Journal on Computing, 5(3): 402–413, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chazelle:1984:CPP

[Cha84] Bernard Chazelle. Convex partitions of polyhedra: a lower bound and worst-case optimal algorithm. SIAM Journal on Computing, 13(3):488–507, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chazelle:1986:FSN

[Cha86] Bernard Chazelle. Filtering search: a new approach to query-answering. SIAM Journal on Computing, 15(3):703– 724, ???? 1986. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chazelle:1988:FAD

[Cha88] Bernard Chazelle. A functional approach to data structures and its use in multidimensional searching. SIAM Journal on Computing, 17(3): 427–462, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:1989:DTS

[Cha89] Edward P. F. Chan. A design theory for solving the anomalies problem. SIAM Journal on Computing, 18(3):429–448, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:1991:EGO

[Cha91] Mee Yee Chan. Embedding of grids into optimal hypercubes. SIAM Journal on Computing, 20(5):834–864, October 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:1992:SBQ

[Cha92a] Richard Chang. On the structure of bounded queries to arbitrary NP sets. SIAM Journal on Computing, 21(4):743–754, August 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chazelle:1992:OAI

[Cha92b] Bernard Chazelle. An optimal algorithm for intersecting three-dimensional convex polyhedra. SIAM Journal on Computing, 21(4):671–696, August 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chaudhuri:1994:TBO

[Cha94]  Shiva Chaudhuri.  Tight

bounds on oblivious chaining. SIAM Journal on Computing, 23(6):1253–1265, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23027.

Chang:1998:EAD

[Cha98a]  Maw-Shang Chang.  Effi-

cient algorithms for the domination problems on interval and circular-arc graphs. SIAM Journal on Computing, 27(6):1671–1694, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23843.

Chazelle:1998:SAL

[Cha98b] Bernard Chazelle. A spectral approach to lower bounds with applications to geometric searching. SIAM Journal on Computing, 27(2):545–556, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27566.

Chan:2000:RSH

[Cha00] Timothy M. Chan. Random sampling, halfspace range reporting, and construction of (≤ k)-levels in three dimensions. SIAM Journal on Computing, 30(2):561–575, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34918.

Chan:2003:SOM

[Cha03] Timothy M. Chan. Semionline maintenance of geometric optima and measures. SIAM Journal on Computing, 32(3):700–716, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40438.

Chan:2005:LDL

[Cha05] Timothy M. Chan. Lowdimensional linear programming with violations. SIAM Journal on Computing, 34 (4):879–893, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43940.

Chan:2006:DSC

[Cha06] Timothy M. Chan. Dynamic subgraph connectivity with geometric applications. SIAM Journal on Computing, 36(3): 681–694, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2010:MAA

[Cha10] Timothy M. Chan. More algorithms for all-pairs shortest paths in weighted graphs. SIAM Journal on Computing, 39(5):2075–2089, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chazelle:2015:DIS

[Cha15] Bernard Chazelle. Diffusive influence systems. SIAM Journal on Computing, 44(5):1403– 1442, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Censor-Hillel:2021:DSA

[CHD21] Keren Censor-Hillel and Michal Dory. Distributed spanner approximation. SIAM Journal on Computing, 50(3):1103–1147, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cherniavsky:1976:SPR

[Che76] John C. Cherniavsky. Simple programs realize exactly Presburger formulas. SIAM Journal on Computing, 5(4):666– 677, ???? 1976. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cherry:1986:IFT

[Che86] G. W. Cherry. Integration in finite terms with special functions: the logarithmic integral. SIAM Journal on Computing, 15(1):1–21, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cherry:1989:ARE

[Che89] G. W. Cherry. An analysis of the rational exponential integral. SIAM Journal on Computing, 18(5):893–905, October 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:1992:HSM

[Che92] Pang C. Chen. Heuristic sampling: a method for predicting the performance of tree searching programs. SIAM Journal on Computing, 21(2):295–315, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:1993:BHP

[Che93] Bo Chen. A better heuristic for preemptive parallel machine scheduling with batch setup times. SIAM Journal on Computing, 22(6):1303– 1318, December 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheriyan:1997:RAP

[Che97] Joseph Cheriyan. Randomized Õ(M(|V |)) algorithms for problems in matching theory. SIAM Journal on Computing, 26(6):1635–1655, December 1997. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25622.

Chen:2001:PES

[Che01]  Jing-Chao Chen.  Propor-

tion extend sort. SIAM Journal on Computing, 31 (1):323–330, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34290.

Cheng:2005:BSD

[Che05] Qi Cheng. On the bounded sum-of-digits discrete logarithm problem in finite fields. SIAM Journal on Computing, 34(6):1432–1442, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44603.

Chen:2008:CQC

[Che08] Hubie Chen. The complexity of quantified constraint satisfaction: Collapsibility, sink algebras, and the three-element case. SIAM Journal on Computing, 37(5):1674–1701, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2009:CMM

[Che09] Ke Chen. On coresets for kmedian and k-means clustering in metric and Euclidean spaces and their applications. SIAM Journal on Computing, 39(3):923–947, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chepoi:2012:NLP

[Che12] Victor Chepoi. Nice labeling problem for event structures: a counterexample. SIAM Journal on Computing, 41(4):715– 727, ???? 2012. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2002:FDE

[CHH02] Zhi-Zhong Chen, Xin He, and Chun-Hsi Huang. Finding double Euler trails of planar graphs in linear time. SIAM Journal on Computing, 31(4):1255–1285, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35432.

Censor-Hillel:2017:RSN

[CHHKM17] Keren Censor-Hillel, Bernhard Haeupler, Jonathan Kelner, and Petar Maymounkov. Rumor spreading with no dependence on conductance. SIAM Journal on Computing, 46(1): 58–79, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chin:1976:GAU

[Chi76] Francis Y. Chin. A generalized asymptotic upper bound on fast polynomial evaluation and interpolation. SIAM Journal on Computing, 5(4):682– 690, ???? 1976. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chin:1977:PFE

[Chi77] Francis Y. Chin. The partial fraction expansion problem and its inverse. SIAM Journal on Computing, 6(3):554– 562, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2018:PSF

[CHJ18] T-H. Hubert Chan, Shuguang Hu, and Shaofeng H.-C. Jiang. A PTAS for the Steiner forest problem in doubling metrics. SIAM Journal on Computing, 47(4):1705–1734, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chandra:2004:GIC

[CHJT04] Tushar Chandra, Vassos Hadzilacos, Prasad Jayanti, and Sam Toueg. Generalized irreducibility of consensus and the equivalence of t-resilient and wait-free implementations of consensus. SIAM Journal on Computing, 34(2):333–357, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34436.

Chen:2003:CFE

[CHK03] Zhi-Zhong Chen, Xin He, and Ming-Yang Kao. Common-face embeddings of planar graphs. SIAM Journal on Computing, 32(2):408–434, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37775.

Chakraborty:2013:DNP

[CHK13] Tanmoy Chakraborty, Zhiyi Huang, and Sanjeev Khanna. Dynamic and nonuniform pricing strategies for revenue maximization. SIAM Journal on Computing, 42(6):2424–2451, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chekuri:2010:AAN

[CHKS10] C. Chekuri, M. T. Hajiaghayi, G. Kortsarz, and M. R. Salavatipour. Approximation algorithms for nonuniform buy-at-bulk network design. SIAM Journal on Computing, 39(5):1772–1798, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:2003:RDQ

[CHKZ03] Edith Cohen, Eran Halperin, Haim Kaplan, and Uri Zwick. Reachability and distance queries via 2-hop labels. SIAM Journal on Computing, 32 (5):1338–1355, October 2003.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40309.

Chen:2017:CIV

[CHLX17] Danny Z. Chen, Ziyun Huang, Yangwei Liu, and Jinhui Xu. On clustering induced Voronoi diagrams. SIAM Journal on Computing, 46(6):1679–1711, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheriyan:1996:TMF

[CHM96] Joseph Cheriyan, Torben Hagerup, and Kurt Mehlhorn. An o(n^{3})-time maximum-flow algorithm. SIAM Journal on Computing, 25(6):1144–1170, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
27837.

Chitnis:2013:FPT

[CHM13] Rajesh Chitnis, Mohammad-Taghi Hajiaghayi, and Dániel Marx. Fixed-parameter tractability of directed multiway cut parameterized by the size of the cutset. SIAM Journal on Computing, 42(4):1674– 1696, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:2018:WCC

[CHN^{+}18] Aloni Cohen, Justin Holmgren, Ryo Nishimaki, Vinod Vaikuntanathan, and Daniel Wichs. Watermarking cryptographic capabilities. SIAM Journal on Computing, 47(6): 2157–2202, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2008:EOP

[CHP08] Ke Chen and Sariel Har-Peled. The Euclidean orienteering problem revisited. SIAM Journal on Computing, 38(1):385–397, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2020:LSO

[CHPJ20] Timothy M. Chan, Sariel Har-Peled, and Mitchell Jones. On locality-sensitive orderings and their applications. SIAM Journal on Computing, 49(3):583– 600, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2022:OAG

[CHPJ22] Timothy M. Chan, Sariel Har-Peled, and Mitchell Jones. Optimal algorithms for geometric centers and depth. SIAM Journal on Computing, 51(3):627–663, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1423324.

Condon:1998:PFA

[CHPW98] Anne Condon, Lisa Hellerstein, Samuel Pottle, and Avi Wigderson. On the power of finite automata with both nondeterministic and probabilistic states. SIAM Journal on Computing, 27(3):739–762, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26557.

Cole:1995:TLB

[CHPZ95] Richard Cole, Ramesh Hariharan, Mike Paterson, and Uri Zwick. Tighter lower bounds on the exact complexity of string matching. SIAM Journal on Computing, 24(1):30–45, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24582.

Censor-Hillel:2012:FIS

[CHS12] Keren Censor-Hillel and Hadas Shachnai. Fast information spreading in graphs with large weak conductance. SIAM Journal on Computing, 41(6): 1451–1465, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:1993:TAM

[CHT93] Jiazhen Cai, Xiaofeng Han, and Robert E. Tarjan. An O(m log n)-time algorithm for the maximal planar subgraph problem. SIAM Journal on Computing, 22(6):1142–1162, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chuzhoy:2016:RUG

[Chu16] Julia Chuzhoy. Routing in undirected graphs with constant congestion. SIAM Journal on Computing, 45(4):1490– 1532, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chvatal:1977:DSN

[Chv77] V. Chvátal. Determining the stability number of a graph. SIAM Journal on Computing, 6(4):643–662, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2013:CSP

[CHW13] Danny Z. Chen, John Hershberger, and Haitao Wang. Computing shortest paths amid convex pseudodisks. SIAM Journal on Computing, 42(3):1158–1184, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2007:MMT

[CHZ07] Xujin Chen, Xiaodong Hu, and Wenan Zang. A min-max theorem on tournaments. SIAM Journal on Computing, 37(3):

923–937, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:1983:STC

[CI83] Tat Hung Chan and Oscar H.

Ibarra. On the space and time complexity of functions computable by simple programs. SIAM Journal on Computing, 12(4):708–716, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Canetti:1998:BPP

[CI98] Ran Canetti and Sandy Irani. Bounding the power of preemption in randomized scheduling. SIAM Journal on Computing, 27(4):993– 1015, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28329.

Chor:1994:WFC

[CIL94] Benny Chor, Amos Israeli, and Ming Li. Wait-free consensus using asynchronous hardware. SIAM Journal on Computing, 23(4):701–712, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19263.

Castanneda:2016:GSB

[CIRR16] Armando Castanñeda, Damien Imbs, Sergio Rajsbaum, and Michel Raynal. Generalized symmetry breaking tasks and nondeterminism in concurrent objects. SIAM Journal on Computing, 45(2):379–414, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheng:1992:NRD

[CJ92] Siu Wing Cheng and Ravi Janardan. New results on dynamic planar point location. SIAM Journal on Computing, 21(5):972–999, October 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheng:2014:ASD

[CJ14] Siu-Wing Cheng and Jiongxin Jin. Approximate shortest descending paths. SIAM Journal on Computing, 43(2):410–428, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Calude:2022:DPG

[CJK^{+}22] Cristian S. Calude, Sanjay Jain, Bakhadyr Khoussainov, Wei Li, and Frank Stephan. Deciding parity games in quasi-polynomial time. SIAM Journal on Computing, 51(2): STOC17–152–STOC17–188, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/17M1145288.

Chen:2003:CPR

[CJL03] Zhi-Zhong Chen, Tao Jiang, and Guohui Lin. Computing phylogenetic roots with bounded degrees and errors. SIAM Journal on Computing, 32(4):864–879, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38915.

Case:2006:ICP

[CJM^{+}06] John Case, Sanjay Jain, Eric Martin, Arun Sharma, and Frank Stephan. Identifying clusters from positive data. SIAM Journal on Computing, 36(1):28–55, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chrobak:2007:OSE

[CJST07] Marek Chrobak, Wojciech Jawor, Jiří Sgall, and Tomáš Tichý. Online scheduling of equal-length jobs: Randomization and restarts help. SIAM Journal on Computing, 36(6): 1709–1728, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Corneil:1980:TAV

[CK80] D. G. Corneil and D. G. Kirkpatrick. A theoretical analysis of various heuristics for the graph isomorphism problem. SIAM Journal on Computing, 9(2):281–297, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Culik:1987:EPS

[CK87] Karel Culik, II and Juhani Karhumäki. The equivalence problem for single-valued twoway transducers (on NPDTOL languages) is decidable. SIAM Journal on Computing, 16(2): 221–230, April 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Colte:1991:BFI

[CK91] Peter Colte and Evangelos Kranakis. Boolean functions, invariance groups, and parallel complexity. SIAM Journal on Computing, 20(3):553–590, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chu:1993:NMC

[CK93] Jiang-Hsing Chu and Gary D. Knott. A new method for computing page-fault rates. SIAM Journal on Computing, 22(6): 1319–1330, December 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Culik:1994:FAC

[CK94] Karel Culik, II and Juhani Karhumäki. Finite automata computing real functions. SIAM Journal on Computing, 23(4):789–814, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22489.

Chou:1995:CCT

[CK95] Arthur W. Chou and Ker-I Ko. Computational complexity of two-dimensional regions. SIAM Journal on Computing, 24(5):923–947, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25456.

Chang:1996:BHP

[CK96a] Richard Chang and Jim Kadin.

The Boolean hierarchy and the polynomial hierarchy: a closer connection. SIAM Journal on Computing, 25(2):340–354, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/17806.

Cypher:1996:BEM

[CK96b] Robert Cypher and Smaragda Konstantinidou. Bounds on the efficiency of messagepassing protocols for parallel computers. SIAM Journal on Computing, 25(5):1082– 1104, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26216.

Chen:2000:RRI

[CK00] Zhi-Zhong Chen and Ming-Yang Kao. Reducing randomness via irrational numbers. SIAM Journal on Computing, 29(4):1247–1256, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34160.

Csuros:2001:PFA

[CK01] Miklós Csurös and Ming-Yang Kao. Provably fast and accurate recovery of evolutionary trees through harmonic greedy triplets. SIAM Journal on Computing, 31(1):306– 322, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        37905.

Chekuri:2004:MPP

[CK04a] Chandra Chekuri and Sanjeev Khanna. On multidimensional packing problems. SIAM Journal on Computing, 33(4):837–851, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35626.

Cohen:2004:BRA

[CK04b] Edith Cohen and Haim Kaplan. Balanced-replication algorithms for distribution trees.

SIAM Journal on Computing, 34(1):227–247, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41214.

Chekuri:2005:PTA

[CK05] Chandra Chekuri and Sanjeev Khanna. A polynomial time approximation scheme for the multiple knapsack problem. SIAM Journal on Computing, 35(3):713–728, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:2009:PEA

[CK09] Kevin L. Chang and Ravi Kannan. Pass-efficient algorithms for learning mixtures of uniform distributions. SIAM Journal on Computing, 39(3): 783–812, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Childs:2012:QQC

[CK12] Andrew M. Childs and Robin Kothari. Quantum query complexity of minor-closed graph properties. SIAM Journal on Computing, 41(6):1426–1450, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Christodoulou:2013:DTP

[CK13] George Christodoulou and Annamária Kovács. A deterministic truthful PTAS for scheduling related machines. SIAM Journal on Computing, 42(4): 1572–1595, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chierichetti:2014:VLI

[CK14] Flavio Chierichetti and Jon Kleinberg. Voting with limited information and many alternatives. SIAM Journal on Computing, 43(5):1615–1653, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chailloux:2017:PLQ

[CK17] André Chailloux and Iordanis Kerenidis. Physical limitations of quantum cryptographic primitives or optimal bounds for quantum coin flipping and bit commitment. SIAM Journal on Computing, 46(5):1647–1677, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1998:PRC

[CKL98] E. G. Coffman, Jr., Nabil Kahale, and F. T. Leighton. Processor-ring communication: a tight asymptotic bound on packet waiting times. SIAM Journal on Computing, 27 (5):1221–1236, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26863.

Chen:2009:RDC

[CKL^{+}09] Jianer Chen, Joachim Kneis, Songjian Lu, Daniel Mölle, Stefan Richter, Peter Rossmanith, Sing-Hoi Sze, and Fenghui Zhang. Randomized divide-and-conquer: Improved path, matching, and packing algorithms. SIAM Journal on Computing, 38(6):2526–2547, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chierichetti:2013:MCW

[CKL^{+}13] Flavio Chierichetti, Ravi Kumar, Silvio Lattanzi, Alessandro Panconesi, and Prabhakar Raghavan. Models for the compressible Web. SIAM Journal on Computing, 42(5):1777– 1802, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2001:OBH

[CKLW01] Gen-Huey Chen, Ming-Yang Kao, Yuh-Dauh Lyuu, and Hsing-Kuo Wong. Optimal buy-and-hold strategies for financial markets with bounded daily returns. SIAM Journal on Computing, 31(2):447–459, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35884.

Cohen:2010:LDX

[CKM10] Edith Cohen, Haim Kaplan, and Tova Milo. Labeling dynamic XML trees. SIAM Journal on Computing, 39(5):2048– 2074, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cadek:2014:PTC

[CKM^{+}14] Martin Cadek, Marek Krcál, Jirí Matousek, Lukás Vokrínek, and Uli Wagner. Polynomialtime computation of homotopy groups and Postnikov systems in fixed dimension. SIAM Journal on Computing, 43(5): 1728–1780, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cidon:1995:GPS

[CKMP95] Israel Cidon, Shay Kutten, Yishay Mansour, and David Peleg. Greedy packet scheduling. SIAM Journal on Computing, 24(1):148–157, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/21769.

Chaudhury:2021:LCG

[CKMS21] Bhaskar Ray Chaudhury, Telikepalli Kavitha, Kurt Mehlhorn, and Alkmini Sgouritsa. A little charity guarantees almost envy-freeness. SIAM Journal on Computing, 50(4): 1336–1358, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chuzhoy:2022:NHR

[CKN22] Julia Chuzhoy, David H. K.

Kim, and Rachit Nimavat. New hardness results for routing on disjoint paths. SIAM Journal on Computing, 51(2): STOC17–189–STOC17–237, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1146580.

Coppersmith:1986:AMT

[CKP86] D. Coppersmith, M. M. Klawe, and N. J. Pippenger. Alphabetic minimax trees of degree at most t. SIAM Journal on Computing, 15(1):189–192, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chierichetti:2014:HSC

[CKP14] Flavio Chierichetti, Jon Kleinberg, and Alessandro Panconesi. How to schedule a cascade in an arbitrary graph. SIAM Journal on Computing, 43(6):1906–1920, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:2019:ESB

[CKP19] Yi-Jun Chang, Tsvi Kopelowitz, and Seth Pettie. An exponential separation between randomized and deterministic complexity in the LOCAL model. SIAM Journal on Computing, 48(1):122–143, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Canetti:2002:BBC

[CKPR02] Ran Canetti, Joe Kilian, Erez Petrank, and Alon Rosen. Black-box concurrent zeroknowledge requires (almost) logarithmically many rounds. SIAM Journal on Computing, 32(1):1–47, February 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39294.

Chierichetti:2019:DLS

[CKPT19] Flavio Chierichetti, Ravi Kumar, Alessandro Panconesi, and Erisa Terolli. On the distortion of locality sensitive hashing. SIAM Journal on Computing, 48(2):350–372, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1972:ASP

[CKR72] E. G. Coffman, L. A. Klimko, and Barbara Ryan. Analysis of scanning policies for reducing disk seek times. SIAM Journal on Computing, 1(3):269– 279, ???? 1972. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Charikar:2001:ACV

[CKR01] Moses Charikar, Samir Khuller, and Balaji Raghavachari. Algorithms for capacitated vehicle routing. SIAM Journal on Computing, 31(3):665–682, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39205.

Calinescu:2004:AAE

[CKR04] Gruia Calinescu, Howard Karloff, and Yuval Rabani. Approximation algorithms for the 0-extension problem. SIAM Journal on Computing, 34(2): 358–372, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39597.

Chekuri:2015:MFC

[CKRV15] Chandra Chekuri, Sreeram Kannan, Adnan Raja, and Pramod Viswanath. Multicommodity flows and cuts in polymatroidal networks. SIAM Journal on Computing, 44(4): 912–943, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1985:SMF

[CKS85] E. G. Coffman, Jr., T. T.

Kadota, and L. A. Shepp. A stochastic model of fragmentation in dynamic storage allocation. SIAM Journal on Computing, 14(2):416–425, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chakrabarti:2002:ESC

[CKS02] Amit Chakrabarti, Subhash Khot, and Yaoyun Shi. Evasiveness of subgraph containment and related properties. SIAM Journal on Computing, 31(3):866–875, June 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38200.

Chekuri:2009:EDP

[CKS09] Chandra Chekuri, Sanjeev Khanna, and F. Bruce Shepherd. Edge-disjoint paths in planar graphs with constant congestion. SIAM Journal on Computing, 39(1):281–301, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cachin:2011:FAU

[CKS11] Christian Cachin, Idit Keidar, and Alexander Shraer. Fail-aware untrusted storage. SIAM Journal on Computing, 40(2):493–533, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i2/p493_s1.

Chekuri:2013:ANM

[CKS13] Chandra Chekuri, Sanjeev Khanna, and F. Bruce Shepherd. The all-or-nothing multicommodity flow problem. SIAM Journal on Computing, 42(4):1467–1493, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Childs:2017:QAS

[CKS17] Andrew M. Childs, Robin Kothari, and Rolando D. Somma. Quantum algorithm for systems of linear equations with exponentially improved dependence on precision. SIAM Journal on Computing, 46(6):1920–1950, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Crescenzi:1999:SAC

[CKST99] Pierluigi Crescenzi, Viggo Kann, Riccardo Silvestri, and Luca Trevisan. Structure in approximation classes. SIAM Journal on Computing, 28 (5):1759–1782, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30422.

Cheriyan:1993:SFS

[CKT93] Joseph Cheriyan, Ming-Yang Kao, and Ramakrishna Thurimella. Scan-first search and sparse certificates: an improved parallel algorithm for k-vertex connectivity. SIAM Journal on Computing, 22(1):157–174, February 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chandra:1999:NRO

[CKT99] Barun Chandra, Howard Karloff, and Craig Tovey. New results on the old kopt algorithm for the traveling salesman problem. SIAM Journal on Computing, 28(6): 1998–2029, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25124.

Chen:2000:SPQ

[CKT00] Danny Z. Chen, Kevin S. Klenk, and Hung-Yi T. Tu. Shortest path queries among weighted obstacles in the rectilinear plane. SIAM Journal on Computing, 29(4):1223– 1246, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
30719.

Czumaj:2010:STA

[CKV10] Artur Czumaj, Piotr Krysta, and Berthold Vöcking. Selfish traffic allocation for server farms. SIAM Journal on Computing, 39(5):1957–1987, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Carson:1977:NSA

[CL77] John S. Carson and Averill M.

Law. A note on Spira’s algorithm for the all-pairs shortestpath problem. SIAM Journal on Computing, 6(4):696– 699, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1979:CAE

[CL79] E. G. Coffman, Jr. and Joseph Y.-T. Leung. Combinatorial analysis of an efficient algorithm for processor and storage allocation. SIAM Journal on Computing, 8(2):202– 217, ???? 1979. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coffman:1989:APS

[CL89] E. G. Coffman, Jr. and J. C.

Lagarias. Algorithms for packing squares: a probabilistic analysis. SIAM Journal on Computing, 18(1):166–185, February 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chrobak:1991:OLA

[CL91] Marek Chrobak and Lawrence L.

Larmore. An optimal online algorithm for k servers on trees. SIAM Journal on Computing, 20(1):144–148, February 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:1994:SSB

[CL94] Jin-Yi Cai and Richard J. Lipton. Subquadratic simulations of balanced formulae by branching programs. SIAM Journal on Computing, 23(3):563–572, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/18133.

Czumaj:2009:FHV

[CL09] Artur Czumaj and Andrzej Lingas. Finding a heaviest vertex-weighted triangle is not harder than matrix multiplication. SIAM Journal on Computing, 39(2):431–444, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:2013:CGP

[CL13] Hsien-Chih Chang and Hsueh-I Lu. Computing the girth of a planar graph in linear time. SIAM Journal on Computing, 42(3):1077–1094, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2019:CIP

[CL19] Yijia Chen and Bingkai Lin. The constant inapproximability of the parameterized dominating set problem. SIAM Journal on Computing, 48(2): 513–533, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheng:2021:APP

[CL21] Siu-Wing Cheng and Man-Kit Lau. Adaptive planar point location. SIAM Journal on Computing, 50(4):1200–1247, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Clarkson:1988:RAC

[Cla88] Kenneth L. Clarkson. A randomized algorithm for closestpoint queries. SIAM Journal on Computing, 17(4):830–847, August 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Clausen:1989:FFT

[Cla89] Michael Clausen. Fast Fourier transforms for metabelian groups. SIAM Journal on Computing, 18(3):584–593, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:1973:SPT

[CLD73] C. L. Chang, R. C. T. Lee, and J. K. Dixon. The specialization of programs by theorem proving. SIAM Journal on Computing, 2(1):7–15, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chiang:2005:OST

[CLL05] Yi-Ting Chiang, Ching-Chi Lin, and Hsueh-I Lu. Orderly spanning trees with applications. SIAM Journal on Computing, 34(4):924–945, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41138.

Chan:2008:EUS

[CLL08] Ho-Leung Chan, Tak-Wah Lam, and Kin-Shing Liu. Extra unit-speed machines are almost as powerful as speedy machines for flow time scheduling. SIAM Journal on Computing, 37(5):1595–1612, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheung:2013:GCN

[CLL13] Ho Yee Cheung, Lap Chi Lau, and Kai Man Leung. Graph connectivities, network coding, and expander graphs. SIAM Journal on Computing, 42(3): 733–751, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2019:EAO

[CLL19] Jing Chen, Bo Li, and Yingkai Li. Efficient approximations for the online dispersion problem. SIAM Journal on Computing, 48(2):373–416, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Condon:1996:CSB

[CLLS96] Anne Condon, Richard Ladner, Jordan Lampe, and Rakesh Sinha. Complexity of sub-bus mesh computations. SIAM Journal on Computing, 25(3):520–539, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25687.

Chazelle:2005:SGA

[CLM05] Bernard Chazelle, Ding Liu, and Avner Magen. Sublinear geometric algorithms. SIAM Journal on Computing, 35(3): 627–646, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Czumaj:2020:RCP

[CLM^{+}20] Artur Czumaj, Jakub Lacki, Aleksander Madry, Slobodan Mitrović, Krzysztof Onak, and Piotr Sankowski. Round compression for parallel matching algorithms. SIAM Journal on Computing, 49(5):STOC18–1– STOC18–44, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2015:NDS

[CLNS15] T.-H. Hubert Chan, Mingfei Li, Li Ning, and Shay Solomon. New doubling spanners: Better and simpler. SIAM Journal on Computing, 44(1):37– 53, ???? 2015. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Canetti:2016:AHC

[CLP16] Ran Canetti, Huijia Lin, and Rafael Pass. Adaptive hardness and composable security in the plain model from standard assumptions. SIAM Journal on Computing, 45(5):1793– 1834, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cygan:2019:MBF

[CLP^{+}19] Marek Cygan, Daniel Lokshtanov, Marcin Pilipczuk, Michał Pilipczuk, and Saket Saurabh. Minimum bisection is fixed-parameter tractable. SIAM Journal on Computing, 48(2):417–450, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:2020:DCU

[CLP20] Yi-Jun Chang, Wenzheng Li, and Seth Pettie. Distributed (Δ + 1)-coloring via ultrafast graph shattering. SIAM Journal on Computing, 49(3):497– 539, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chechik:2010:FTS

[CLPR10] S. Chechik, M. Langberg, D. Peleg, and L. Roditty. Fault tolerant spanners for general graphs. SIAM Journal on Computing, 39(7):3403–3423, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chillara:2019:SDM

[CLS19] Suryajith Chillara, Nutan Limaye, and Srikanth Srinivasan. Small-depth multilinear formula lower bounds for iterated matrix multiplication with applications. SIAM Journal on Computing, 48(1):70–92, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2005:NOD

[CLT05] Ho-Leung Chan, Tak-Wah Lam, and Kar-Keung To. Nonmigratory online deadline scheduling on multiprocessors. SIAM Journal on Computing, 34(3):669–682, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43576.

Chen:2023:RMG

[CLV23] Zongchen Chen, Kuikui Liu, and Eric Vigoda. Rapid mixing of Glauber dynamics up to uniqueness via contraction. SIAM Journal on Computing, 52(1):196–237, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M136685X.

Chen:2004:MAO

[CLWZ04] William Y. C. Chen, Xueliang Li, Chao Wang, and Xiaoyan Zhang. The minimum all-ones problem for trees. SIAM Journal on Computing, 33(2):379–392, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42162.

Cai:2011:CCH

[CLX11] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. Computational complexity of Holant problems. SIAM Journal on Computing, 40(4):1101–1132, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i4/
                                        p1101_s1.

Cai:2017:HAM

[CLX17] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. Holographic algorithms with matchgates capture precisely tractable planar #CSP. SIAM Journal on Computing, 46(3):853–889, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Curran:2006:FFI

[CLY06] Sean Curran, Orlando Lee, and Xingxing Yu. Finding four independent trees. SIAM Journal on Computing, 35(5):1023– 1058, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:1987:GMU

[CM87a] Jin-Yi Cai and Gabriele E.

Meyer. Graph minimal uncolorability is D^{p}-complete. SIAM Journal on Computing, 16(2):259–277, April 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:1987:ISD

[CM87b] Edward P. F. Chan and Alberto O. Mendelzon. Independent and separable database schemes. SIAM Journal on Computing, 16(5):841–851, October 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheriyan:1989:APP

[CM89] J. Cheriyan and S. N. Maheshwari. Analysis of preflow push algorithms for maximum network flow. SIAM Journal on Computing, 18(6):1057– 1086, December 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:1994:IAL

[CM94] Edith Cohen and Nimrod Megiddo. Improved algorithms for linear inequalities with two variables per inequality. SIAM Journal on Computing, 23(6):1313–1347, December 1994. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25632.

Chalasani:1999:ACR

[CM99] Prasad Chalasani and Rajeev Motwani. Approximating capacitated routing and delivery problems. SIAM Journal on Computing, 28(6): 2133–2149, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29546.

Chen:2001:PTA

[CM01] Jianer Chen and Antonio Miranda. A polynomial time approximation scheme for general multiprocessor job scheduling. SIAM Journal on Computing, 31(1):1–17, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34811.

Cabello:2013:AOE

[CM13a] Sergio Cabello and Bojan Mohar. Adding one edge to planar graphs makes crossing number and 1-planarity hard. SIAM Journal on Computing, 42(5): 1803–1829, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2013:DTR

[CM13b] Siu On Chan and Michael Molloy. A dichotomy theorem for the resolution complexity of random constraint satisfaction problems. SIAM Journal on Computing, 42(1):27– 60, ???? 2013. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cubitt:2016:CCL

[CM16] Toby Cubitt and Ashley Montanaro. Complexity classification of local Hamiltonian problems. SIAM Journal on Computing, 45(2):268–316, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chattopadhyay:2022:SLE

[CM22] Arkadev Chattopadhyay and Nikhil S. Mande. A short list of equalities induces large signrank. SIAM Journal on Computing, 51(3):820–848, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1271348.

Charikar:2010:LGT

[CMM10] Moses Charikar, Konstantin Makarychev, and Yury Makarychev. Local global tradeoffs in metric embeddings. SIAM Journal on Computing, 39(6):2487–2512, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chekuri:2001:ATA

[CMNS01] C. Chekuri, R. Motwani, B. Natarajan, and C. Stein. Approximation techniques for average completion time scheduling. SIAM Journal on Computing, 31(1):146–166, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32718.

Cole:1997:RAF

[CMS97] Richard J. Cole, Bruce M. Maggs, and Ramesh K. Sitaraman. Reconfiguring arrays with faults. Part I: Worstcase faults. SIAM Journal on Computing, 26(6): 1581–1611, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25501.

Clarkson:2014:SIA

[CMS14] Kenneth L. Clarkson, Wolfgang Mulzer, and C. Seshadhri. Self-improving algorithms for coordinatewise maxima and convex hulls. SIAM Journal on Computing, 43(2): 617–653, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2018:CAO

[CMS18] Lin Chen, Nicole Megow, and Kevin Schewior. An O(log ⇕)- competitive algorithm for online machine minimization. SIAM Journal on Computing, 47(6):2057–2077, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cole:2000:DFCa

[CMSS00] Richard Cole, Bud Mishra, Jeanette Schmidt, and Alan Siegel. On the dynamic finger conjecture for splay trees. Part I: Splay sorting log n-block sequences. SIAM Journal on Computing, 30(1):1–43, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32698.

Chung:1985:PTA

[CMST85] Moon Jung Chung, Fillia Makedon, Ivan Hal Sudborough, and Jonathan Turner. Polynomial time algorithms for the MIN CUT problem on degree restricted trees. SIAM Journal on Computing, 14(1): 158–177, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chiba:1985:ASL

[CN85] Norishige Chiba and Takao Nishizeki. Arboricity and subgraph listing algorithms. SIAM Journal on Computing, 14(1): 210–223, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chor:1999:SAE

[CN99] Benny Chor and Lee-Bath Nelson. Solvability in asynchronous environments II: Finite interactive tasks. SIAM Journal on Computing, 29(2): 351–377, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        29497.

Chuzhoy:2006:CPH

[CN06] Julia Chuzhoy and Joseph (Seffi) Naor. Covering problems with hard capacities. SIAM Journal on Computing, 36(2):498–515, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chuzhoy:2007:HML

[CN07] Julia Chuzhoy and Joseph (Seffi) Naor. The hardness of metric labeling. SIAM Journal on Computing, 36(5):1376–1386, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2018:TOM

[CN18] Timothy M. Chan and Yakov Nekrich. Towards an optimal method for dynamic planar point location. SIAM Journal on Computing, 47(6): 2337–2361, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chiba:1982:AAM

[CNS82] Norishige Chiba, Takao Nishizeki, and Nobuji Saito. An approximation algorithm for the maximum independent set problem on planar graphs. SIAM Journal on Computing, 11(4):663– 675, ???? 1982. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheng:2008:ASP

[CNVW08] Siu-Wing Cheng, Hyeon-Suk Na, Antoine Vigneron, and Yajun Wang. Approximate shortest paths in anisotropic regions. SIAM Journal on Computing, 38(3):802–824, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheng:2010:QAS

[CNVW10] Siu-Wing Cheng, Hyeon-Suk Na, Antoine Vigneron, and Yajun Wang. Querying approximate shortest paths in anisotropic regions. SIAM Journal on Computing, 39(5): 1888–1918, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Comer:1982:GPA

[CO82] Douglas Comer and Michael J.

O’Donnell. Geometric problems with application to hashing. SIAM Journal on Computing, 11(2):217–226, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coja-Oghlan:2010:BAR

[CO10] Amin Coja-Oghlan. A better algorithm for random k-SAT. SIAM Journal on Computing, 39(7):2823–2864, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coja-Oghlan:2014:AWR

[COF14] Amin Coja-Oghlan and Alan Frieze. Analyzing Walksat on random formulas. SIAM Journal on Computing, 43(4):1456– 1485, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohn:1983:ETS

[Coh83] Avra Cohn. The equivalence of two semantic definitions: a case study in LCF. SIAM Journal on Computing, 12(2): 267–285, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:1995:AMF

[Coh95] Edith Cohen. Approximate max-flow on small depth networks. SIAM Journal on Computing, 24(3):579–597, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23671.

Cohen:1998:FAC

[Coh98] Edith Cohen. Fast algorithms for constructing t-spanners and paths with stretch t. SIAM Journal on Computing, 28 (1):210–236, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26129.

Cohen:2016:LCB

[Coh16] Gil Cohen. Local correlation breakers and applications to three-source extractors and mergers. SIAM Journal on Computing, 45(4):1297–1338, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:2021:TSD

[Coh21] Gil Cohen. Two-source dispersers for polylogarithmic entropy and improved Ramsey graphs. SIAM Journal on Computing, 50(3):STOC16– 30–STOC16–67, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Collins:1974:CTE

[Col74] George E. Collins. The computing time of the Euclidean algorithm. SIAM Journal on Computing, 3(1):1–10, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cole:1988:PMS

[Col88] Richard Cole. Parallel merge sort. SIAM Journal on Computing, 17(4):770–785, August 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cole:1993:CPM

[Col93] Richard Cole. Correction: “Parallel merge sort” [SIAM J. Comput. 17 (1988), no. 4, 770– 785, MR 89m:68015]. SIAM Journal on Computing, 22(6): 1349, December 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cole:1994:TBC

[Col94] Richard Cole. Tight bounds on the complexity of the Boyer– Moore string matching algorithm. SIAM Journal on Computing, 23(5):1075–1091, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/19554.

Cole:2000:DFCb

[Col00] Richard Cole. On the dynamic finger conjecture for splay trees. Part II: The proof. SIAM Journal on Computing, 30(1):44–85, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32699.

Cook:1974:FOG

[Coo74] Curtis R. Cook. First order graph grammars. SIAM Journal on Computing, 3(1):90–99, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cook:1978:SCA

[Coo78] Stephen A. Cook. Soundness and completeness of an axiom system for program verification. SIAM Journal on Computing, 7(1):70–90, February 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See corrigendum [Coo81].

Cook:1981:CSC

[Coo81] Stephen A. Cook. Corrigendum: “Soundness and completeness of an axiom system for program verification” [SIAM J. Comput. 7 (1978), no. 1, 70–90, MR 58 #13843]. SIAM Journal on Computing, 10(3):612, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Coo78].

Coppersmith:1982:RMR

[Cop82] D. Coppersmith. Rapid multiplication of rectangular matrices. SIAM Journal on Computing, 11(3):467–471, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Corneil:1999:LTA

[COS99] Derek G. Corneil, Stephan Olariu, and Lorna Stewart. Linear time algorithms for dominating pairs in asteroidal triple-free graphs. SIAM Journal on Computing, 28 (4):1284–1297, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28237.

Constable:1976:P

[CP76] Robert L. Constable and David Park. Preface. SIAM Journal on Computing, 5(3): 333–335, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cosmadakis:1984:TSP

[CP84] Stavros S. Cosmadakis and Christos H. Papadimitriou. The traveling salesman problem with many visits to few cities. SIAM Journal on Computing, 13(1):99–108, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cohen:2005:CPG

[CP05] Reuven Cohen and David Peleg. Convergence properties of the gravitational algorithm in asynchronous robot systems. SIAM Journal on Computing, 34(6):1516–1528, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/44647.

Cohen:2008:CAM

[CP08] Reuven Cohen and David Peleg. Convergence of autonomous mobile robots with inaccurate sensors and movements. SIAM Journal on Computing, 38(1):276–302, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2009:TRC

[CP09] Timothy M. Chan and Mihai Pătraşcu. Transdichotomous results in computational geometry, I: Point location in sublogarithmic time. SIAM Journal on Computing, 39(2): 703–729, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2012:AOG

[CP12] Jen-Yeu Chen and Gopal Pandurangan. Almost-optimal gossip-based aggregate computation. SIAM Journal on Computing, 41(3):455–483, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chang:2019:THT

[CP19a] Yi-Jun Chang and Seth Pettie. A time hierarchy theorem for the LOCAL model. SIAM Journal on Computing, 48(1):33–69, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2019:DRC

[CP19b] Shiteng Chen and Periklis A. Papakonstantinou. Depth reduction for composites. SIAM Journal on Computing, 48(2): 668–686, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cygan:2016:KAE

[CPP16] Marek Cygan, Marcin Pilipczuk, and Michal Pilipczuk. Known algorithms for edge clique cover are probably optimal. SIAM Journal on Computing, 45(1):67–83, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chudnovsky:2024:QPT

[CPPT24] Maria Chudnovsky, Marcin Pilipczuk, Micha Pilipczuk, and Stéphan Thomassé. Quasipolynomial time approximation schemes for the maximum weight independent set problem in H-free graphs. SIAM Journal on Computing, 53(1): 47–86, February 2024. CO-DEN SMJCAT. ISSN 1095-7111.

Cygan:2012:PAC

[CPPW12] Marek Cygan, Marcin Pilipczuk, Michal Pilipczuk, and Jakub Onufry Wojtaszczyk. A polynomial algorithm for 3-compatible coloring and the stubborn list partition problem (the stubborn problem is stubborn no more). SIAM Journal on Computing, 41(4):815–828, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2011:DCC

[CPR11] Timothy M. Chan, Mihai Pǎtrascu, and Liam Roditty. Dynamic connectivity: Connecting to networks and geometry. SIAM Journal on Computing, 40(2):333–349, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i2/
p333_s1.

Corneil:1985:LRA

[CPS85] D. G. Corneil, Y. Perl, and L. K. Stewart. A linear recognition algorithm for cographs. SIAM Journal on Computing, 14(4):926–934, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chung:2016:NBB

[CPS16] Kai-Min Chung, Rafael Pass, and Karn Seth. Non-black-box simulation from one-way functions and applications to resettable security. SIAM Journal on Computing, 45(2):415–458, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chiang:1996:UAD

[CPT96] Yi-Jen Chiang, Franco P.

Preparata, and Roberto Tamassia. A unified approach to dynamic point location, ray shooting, and shortest paths in planar maps. SIAM Journal on Computing, 25(1):207– 233, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
22451.

Canetti:2022:EYC

[CPV22] Ran Canetti, Oxana Poburinnaya, and Muthuramakrishnan Venkitasubramaniam. Equivocating Yao: Constant-round adaptively secure multiparty computation in the plain model. SIAM Journal on Computing, 51(2):STOC17– 333–STOC17–399, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1151602.

Coppersmith:1982:AGM

[CPW82] D. Coppersmith, D. S. Parker, and C. K. Wong. Analysis of a general mass storage system. SIAM Journal on Computing, 11(1):94–116, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Culik:1989:LSC

[CPY89] Karel Culik, II, Jan Pachl, and Sheng Yu. On the limit sets of cellular automata. SIAM Journal on Computing, 18(4):

831–842, August 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cook:1980:SLB

[CR80] Stephen A. Cook and Charles W.

Rackoff. Space lower bounds for maze threadability on restricted machines. SIAM Journal on Computing, 9(3):636– 652, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chin:1986:OTP

[CR86] Francis Chin and K. V. S.

Ramarao. Optimal termination protocols for network partitioning. SIAM Journal on Computing, 15(1):131–144, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chung:1991:NMI

[CR91] Shun-Ping Chung and Keith W.

Ross. On nonblocking multirate interconnection networks. SIAM Journal on Computing, 20(4):726–736, August 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Czygrinow:2000:ARL

[CR00] Andrzej Czygrinow and Vojtech Rödl. An algorithmic regularity lemma for hypergraphs. SIAM Journal on Computing, 30(4):1041–1066, August 2000. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35172.

Corneil:2005:RBC

[CR05] Derek G. Corneil and Udi Rotics. On the relationship between clique-width and treewidth. SIAM Journal on Computing, 34(4):825– 847, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38535.

Chakrabarti:2010:ORC

[CR10] Amit Chakrabarti and Oded Regev. An optimal randomized cell probe lower bound for approximate nearest neighbor searching. SIAM Journal on Computing, 39(5):1919– 1940, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chakrabarti:2012:OLB

[CR12] Amit Chakrabarti and Oded Regev. An optimal lower bound on the communication complexity of gap-Hamming-distance. SIAM Journal on Computing, 41(5):1299–1317, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chuzhoy:2017:SSF

[CR17] Julia Chuzhoy and Alexander Russell. Special section on the Fifty-Fifth Annual ACM Symposium on Foundations of Computer Science (FOCS 2014). SIAM Journal on Computing, 46(1):160, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2022:SAC

[CR22] Lijie Chen and Hanlin Ren.

Strong average-case circuit lower bounds from nontrivial derandomization. SIAM Journal on Computing, 51(3): STOC20–115–STOC20–173, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1364886.

Crespi-Reghizzi:1981:OPG

[CRGM81] Stefano Crespi-Reghizzi, Giovanni Guida, and Dino Mandrioli. Operator precedence grammars and the noncounting property. SIAM Journal on Computing, 10(1):174–191, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Casteigts:2024:STR

[CRRZ24] Arnaud Casteigts, Michael Raskin, Malte Renken, and Viktor Zamaraev. Sharp thresholds in random simple temporal graphs. SIAM Journal on Computing, 53(2):346– 388, March 2024. CODEN SMJCAT. ISSN 1095-7111.

Courcoubetis:1987:SQS

[CRS87] C. A. Courcoubetis, M. I. Reiman, and B. Simon. Stability of a queueing system with concurrent service and locking. SIAM Journal on Computing, 16(1):169–178, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chari:1995:ROU

[CRS95] Suresh Chari, Pankaj Rohatgi, and Aravind Srinivasan. Randomness-optimal unique element isolation with applications to perfect matching and related problems. SIAM Journal on Computing, 24 (5):1036–1050, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25033.

Cohen:2014:NES

[CRS14] Gil Cohen, Ran Raz, and Gil Segev. Nonmalleable extractors with short seeds and applications to privacy amplification. SIAM Journal on Computing, 43(2):450–476, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Canonne:2015:TPD

[CRS15] Clément L. Canonne, Dana Ron, and Rocco A. Servedio. Testing probability distributions using conditional samples. SIAM Journal on Computing, 44(3):540–616, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Celis:2013:BBS

[CRSW13] L. Elisa Celis, Omer Reingold, Gil Segev, and Udi Wieder. Balls and bins: Smaller hash families and faster evaluation. SIAM Journal on Computing, 42(3):1030–1050, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chazelle:2005:AMS

[CRT05] Bernard Chazelle, Ronitt Rubinfeld, and Luca Trevisan. Approximating the minimum spanning tree weight in sublinear time. SIAM Journal on Computing, 34(6): 1370–1379, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40324.

Chen:2010:DNP

[CRV10] Ho-Lin Chen, Tim Roughgarden, and Gregory Valiant. Designing network protocols for good equilibria. SIAM Journal on Computing, 39(5):1799– 1832, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Carbonnel:2022:CGV

[CRZ22] Clément Carbonnel, Miguel Romero, and Stanislav Zivný. The complexity of generalvalued constraint satisfaction problems seen from the other side. SIAM Journal on Computing, 51(1):19–69, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/19M1250121.

Cho:1980:BLS

[CS80] Yookun Cho and Sartaj Sahni. Bounds for list schedules on uniform processors. SIAM Journal on Computing, 9(1): 91–103, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:1989:RSE

[CS89] Ming-Syan Chen and Kang G. Shin. On relaxed squashed embedding of graphs into a hypercube. SIAM Journal on Computing, 18(6):1226–1244, December 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cai:1999:FSA

[CS99] Jin-Yi Cai and Alan L. Selman. Fine separation of averagetime complexity classes. SIAM Journal on Computing, 28 (4):1310–1325, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31171.

Chudak:2003:IAA

[CS03a] Fabián A. Chudak and David B.

Shmoys. Improved approximation algorithms for the uncapacitated facility location problem. SIAM Journal on Computing, 33(1):1–25, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40575.

Cramer:2003:DAP

[CS03b] Ronald Cramer and Victor Shoup. Design and analysis of practical public-key encryption schemes secure against adaptive chosen ciphertext attack. SIAM Journal on Computing, 33(1):167–226, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40377.

Chen:2004:FSC

[CS04] Xiao Chen and Jian Shen.

On the Frame–Stewart conjecture about the Towers of Hanoi. SIAM Journal on Computing, 33(3):584–589, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43101.

Czumaj:2005:ACP

[CS05] Artur Czumaj and Christian Sohler. Abstract combinatorial programs and efficient property testers. SIAM Journal on Computing, 34(3):580–615, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/44199.

Chien:2007:API

[CS07] Steve Chien and Alistair Sinclair. Algebras with polynomial identities and computing the determinant. SIAM Journal on Computing, 37(1):252– 266, ???? 2007. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Czumaj:2009:EWM

[CS09] Artur Czumaj and Christian Sohler. Estimating the weight of metric minimum spanning trees in sublinear time. SIAM Journal on Computing, 39(3): 904–922, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chakrabarty:2016:MTB

[CS16] D. Chakrabarty and C. Seshadhri. An o(n) monotonicity tester for Boolean functions over the hypercube. SIAM Journal on Computing, 45(2): 461–472, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chekuri:2018:AAE

[CS18] Chandra Chekuri and Anastasios Sidiropoulos. Approximation algorithms for Euler genus and related problems. SIAM Journal on Computing, 47(4): 1610–1643, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Christodoulou:2019:DNG

[CS19] George Christodoulou and Alkmini Sgouritsa. Designing networks with good equilibria under uncertainty. SIAM Journal on Computing, 48(4): 1364–1396, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Czumaj:2024:STA

[CS24] Artur Czumaj and Christian Sohler. Sublinear time approximation of the cost of a metric k-nearest neighbor graph. SIAM Journal on Computing, 53(2):524–571, April 2024. CODEN SMJCAT. ISSN 1095-7111.

Csanky:1976:FPM

[Csa76] L. Csanky. Fast parallel matrix inversion algorithms. SIAM Journal on Computing, 5(4): 618–623, December 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cypher:1990:HTC

[CSS90] R. E. Cypher, J. L. C. Sanz, and L. Snyder. The Hough transform has O(N) complexity on N × N mesh connected computers. SIAM Journal on Computing, 19(5):805–820, October 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Czumaj:2009:THP

[CSS09] Artur Czumaj, Asaf Shapira, and Christian Sohler. Testing hereditary properties of nonexpanding bounded-degree graphs. SIAM Journal on Computing, 38(6):2499–2510, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cole:1989:OTA

[CSSS89] Richard Cole, Jeffrey S. Salowe, W. L. Steiger, and Endre Szemerédi. An optimaltime algorithm for slope selection. SIAM Journal on Computing, 18(4):792–810, August 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chandra:1984:CDR

[CSV84] Ashok K. Chandra, Larry Stockmeyer, and Uzi Vishkin. Constant depth reducibility. SIAM Journal on Computing, 13(2):423–439, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cormen:1998:ATB

[CSW98] Thomas H. Cormen, Thomas Sundquist, and Leonard F. Wisniewski. Asymptotically tight bounds for performing BMMC permutations on parallel disk systems. SIAM Journal on Computing, 28 (1):105–136, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28368.

Cole:1987:HRP

[CSY87] Richard Cole, Micha Sharir, and Chee K. Yap. On k-hulls and related problems. SIAM Journal on Computing, 16(1): 61–77, February 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chudnovsky:2024:FCF

[CSZ24a] Maria Chudnovsky, Sophie Spirkl, and Mingxian Zhong. Four-coloring P_{6}-free graphs. I. Extending an excellent precoloring. SIAM Journal on Computing, 53(1):111–145, February 2024. CODEN SMJCAT. ISSN 1095-7111.

Chudnovsky:2024:FCB

[CSZ24b] Maria Chudnovsky, Sophie Spirkl, and Mingxian Zhong. Four-coloring P_{6}-free graphs. II. Finding an excellent precoloring. SIAM Journal on Computing, 53(1):146–187, February 2024. CODEN SMJCAT. ISSN 1095-7111.

Cheriton:1976:FMS

[CT76] David Cheriton and Robert Endre Tarjan. Finding minimum spanning trees. SIAM Journal on Computing, 5(4):724–742, December 1976. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chin:1990:ITC

[CT90] F. Chin and H. F. Ting. Improving the time complexity of message-optimal distributed algorithms for minimumweight spanning trees. SIAM Journal on Computing, 19(4): 612–626, August 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:1999:RTE

[CT99] Weimin Chen and Volker Turau. On regular tree embeddings. SIAM Journal on Computing, 29(1):288–301, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32430.

Cheriyan:2000:AMS

[CT00] Joseph Cheriyan and Ramakrishna Thurimella. Approximating minimum-size kconnected spanning subgraphs via matching. SIAM Journal on Computing, 30(2):528–560, April 2000. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33920.

Cormode:2010:TDS

[CTX10] Graham Cormode, Srikanta Tirthapura, and Bojian Xu. Time-decaying sketches for robust aggregation of sensor data. SIAM Journal on Computing, 39(4):1309–1339, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cunningham:1986:IBM

[Cun86] William H. Cunningham. Improved bounds for matroid partition and intersection algorithms. SIAM Journal on Computing, 15(4):948–957, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cusick:1998:VSS

[Cus98] Thomas W. Cusick. Value sets of some polynomials over finite fields GF(22m). SIAM Journal on Computing, 27 (1):120–131, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27035.

Chen:1983:AEI

[CV83] Wen Chin Chen and Jeffrey Scott Vitter. Analysis of early-insertion standard coalesced hashing. SIAM Journal on Computing, 12(4):667– 676, November 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chandra:1985:IPF

[CV85] Ashok K. Chandra and Moshe Y. Vardi. The implication problem for functional and inclusion dependencies is undecidable. SIAM Journal on Computing, 14(3):671–677, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cole:1988:APS

[CV88] Richard Cole and Uzi Vishkin. Approximate parallel scheduling. Part I: The basic technique with applications to optimal parallel list ranking in logarithmic time. SIAM Journal on Computing, 17(1):128–142, February 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chierichetti:2010:LNL

[CV10] Flavio Chierichetti and Andrea Vattani. The local nature of list colorings for graphs of high girth. SIAM Journal on Computing, 39(6):2232–2250, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cheriyan:2014:AMC

[CV14] Joseph Cheriyan and László A. Végh. Approximating minimumcost k-node connected subgraphs via independence-free graphs. SIAM Journal on Computing, 43(4):1342–1362, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cousins:2018:GCA

[CV18] Ben Cousins and Santosh Vempala. Gaussian cooling and O^{∗}(n^{3}) algorithms for volume and Gaussian volume. SIAM Journal on Computing, 47(3): 1237–1273, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chan:2000:COO

[CvdM00] Edward P. F. Chan and Ron van der Meyden. Containment and optimization of objectpreserving conjunctive queries. SIAM Journal on Computing, 29(4):1371–1400, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26244.

Cheriyan:2003:AAM

[CVV03] Joseph Cheriyan, Santosh Vempala, and Adrian Vetta. An approximation algorithm for the minimum-cost k-vertex connected subgraph. SIAM Journal on Computing, 32 (4):1050–1055, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39228.

Chen:2019:CQT

[CVY19] Hubie Chen, Matt Valeriote, and Yuichi Yoshida. Constantquery testability of assignments to constraint satisfaction problems. SIAM Journal on Computing, 48(3):1022– 1045, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chekuri:2014:SFM

[CVZ14] Chandra Chekuri, Jan Vondrák, and Rico Zenklusen. Submodular function maximization via the multilinear relaxation and contention resolution schemes. SIAM Journal on Computing, 43(6):1831–1879, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chandra:1975:WCA

[CW75] Ashok K. Chandra and C. K. Wong. Worst-case analysis of a placement algorithm related to storage allocation. SIAM Journal on Computing, 4(3):249– 263, ???? 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Coppersmith:1982:ACM

[CW82] D. Coppersmith and S. Winograd. On the asymptotic complexity of matrix multiplication. SIAM Journal on Computing, 11(3):472–492, August 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Choy:1983:COL

[CW83] David M. Choy and C. K.

Wong. Construction of optimal α-β leaf trees with applications to prefix code and information retrieval. SIAM Journal on Computing, 12(3):426– 446, ???? 1983. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chaudhuri:1994:BCM

[CW94] Soma Chaudhuri and Jennifer L. Welch. Bounds on the costs of multivalued register implementations. SIAM Journal on Computing, 23(2): 335–354, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
19097.

Chin:1998:FCD

[CW98] Francis Chin and Cao An Wang. Finding the constrained Delaunay triangulation and constrained Voronoi diagram of a simple polygon in linear time. SIAM Journal on Computing, 28(2):471–486, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28591.

Cai:2004:PCL

[CW04] Jin-Yi Cai and Osamu Watanabe. On proving circuit lower bounds against the polynomial-time hierarchy. SIAM Journal on Computing, 33(4):984–1009, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42271.

Cheng:2007:LBD

[CW07] Qi Cheng and Daqing Wan. On the list and bounded distance decodability of Reed– Solomon codes. SIAM Journal on Computing, 37(1):195–209, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chekuri:2018:MCS

[CX18] Chandra Chekuri and Chao Xu. Minimum cuts and sparsification in hypergraphs. SIAM Journal on Computing, 47(6): 2118–2156, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chen:2004:CGB

[CXY04] Guantao Chen, Jun Xu, and Xingxing Yu. Circumference of graphs with bounded degree. SIAM Journal on Computing, 33(5):1136–1170, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43647.

Cypher:1993:LBS

[Cyp93a] Robert Cypher. A lower bound on the size of shellsort sorting networks. SIAM Journal on Computing, 22(1):62– 71, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Cypher:1993:TAV

[Cyp93b] Robert Cypher. Theoretical aspects of VLSI pin limitations. SIAM Journal on Computing, 22(2):356–378, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Creignou:2006:CCC

[CZ06] Nadia Creignou and Bruno Zanuttini. A complete classification of the complexity of propositional abduction. SIAM Journal on Computing, 36(1):207–229, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ciardo:2023:CNA

[CZ23] Lorenzo Ciardo and Stanislav Zivný. CLAP: A new algorithm for promise CSPs. SIAM Journal on Computing, 52(1):1–37, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1476435.

Davenport:1986:RDE

[Dav86] J. H. Davenport. The Risch differential equation problem. SIAM Journal on Computing, 15(4):903–918, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

deBruin:1984:ECS

[dB84] Arie de Bruin. On the existence of Cook semantics. SIAM Journal on Computing, 13(1): 1–13, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

deBerg:2020:FET

[dBBKB^{+}20] Mark de Berg, Hans L. Bodlaender, Sándor Kisfaludi-Bak, Dániel Marx, and Tom C. van der Zanden. A framework for exponential-time-hypothesis–tight algorithms and lower bounds in geometric intersection graphs. SIAM Journal on Computing, 49(6): 1291–1331, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

deBerg:2023:ETE

[dBBKBK23] Mark de Berg, Hans L. Bodlaender, Sándor Kisfaludi-Bak, and Sudeshna Kolay. An ETHtight exact algorithm for Euclidean TSP. SIAM Journal on Computing, 52(3):740–760, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1469122.

deBerg:2008:VRS

[dBG08] Mark de Berg and Chris Gray.

Vertical ray shooting and computing depth orders for fat objects. SIAM Journal on Computing, 38(1):257–275, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dezani-Ciancaglini:1998:FMC

[DCdP98] Mariangiola Dezani-Ciancaglini, Ugo de’Liguoro, and Adolfo Piperno. A filter model for concurrent λ-calculus. SIAM Journal on Computing, 27 (5):1376–1419, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27586.

Dusart:2021:CLB

[DCH21] Jérémie Dusart, Derek G.

Corneil, and Michel Habib. Corrigendum: LDFS-based certifying algorithm for the minimum path cover problem on cocomparability graphs. SIAM Journal on Computing, 50(3):1148–1153, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [CDH13].

Dezani-Ciancaglini:1990:PTI

[DCV90] M. Dezani-Ciancaglini and B. Venneri. Partial types and intervals. SIAM Journal on Computing, 19(4):644–658, August 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dessouky:1981:SJU

[DD81] M. I. Dessouky and J. S. Deogun. Sequencing jobs with unequal ready times to minimize mean flow time. SIAM Journal on Computing, 10(1):192–202, February 1981. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dobzinski:2013:PRA

[DD13] Shahar Dobzinski and Shaddin Dughmi. On the power of randomization in algorithmic mechanism design. SIAM Journal on Computing, 42(6): 2287–2304, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dhangwatnotai:2011:TAS

[DDDR11] Peerapong Dhangwatnotai, Shahar Dobzinski, Shaddin Dughmi, and Tim Roughgarden. Truthful approximation schemes for singleparameter agents. SIAM Journal on Computing, 40(3): 915–933, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i3/p915_s1.

David:2017:DST

[DDG^{+}17] Roee David, Irit Dinur, Elazar Goldenberg, Guy Kindler, and Igor Shinkar. Direct sum testing. SIAM Journal on Computing, 46(4):1336–1369, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dasgupta:2009:SAC

[DDH^{+}09] Anirban Dasgupta, Petros Drineas, Boulos Harb, Ravi Kumar, and Michael W. Mahoney. Sampling algorithms and coresets for ℓ_{p} regression. SIAM Journal on Computing, 38(5):2060–2078, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dalal:2005:TWC

[DDMM05] Ketan Dalal, Luc Devroye, Ebrahim Malalla, and Erin McLeish. Two-way chaining with reassignment. SIAM Journal on Computing, 35(2): 327–340, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dolev:2000:NC

[DDN00] Danny Dolev, Cynthia Dwork, and Moni Naor. Nonmalleable cryptography. SIAM Journal on Computing, 30(2):391–437, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29156.

DeSantis:2008:MFC

[DDPY08] Alfredo De Santis, Giovanni Di Crescenzo, Giuseppe Persiano, and Moti Yung. On monotone formula composition of perfect zero-knowledge languages. SIAM Journal on Computing, 38(4):1300–1329, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Davidovitch:2007:SMC

[DDR07] Lior Davidovitch, Shlomi Dolev, and Sergio Rajsbaum. Stability of multivalued continuous consensus. SIAM Journal on Computing, 37(4):1057– 1076, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Daskalakis:2016:HGC

[DDY16] Constantinos Daskalakis, Ilias Diakonikolas, and Mihalis Yannakakis. How good is the chord algorithm? SIAM Journal on Computing, 45(3):811– 858, ???? 2016. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

DeMarco:2010:DBU

[De 10] Gianluca De Marco. Distributed broadcast in unknown radio networks. SIAM Journal on Computing, 39(6):2162– 2175, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Devroye:1998:ULL

[Dev98] Luc Devroye. Universal limit laws for depths in random trees. SIAM Journal on Computing, 28(2):409–432, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28395.

Devroye:1999:HSR

[Dev99] Luc Devroye. The height and size of random hash trees and random pebbled hash trees. SIAM Journal on Computing, 28(4):1215–1224, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32617.

Devroye:2002:LLS

[Dev02] Luc Devroye. Limit laws for sums of functions of subtrees of random binary search trees. SIAM Journal on Computing, 32(1):152–171, February 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38392.

Devroye:2012:SSC

[Dev12] Luc Devroye. Simulating sizeconstrained Galton–Watson trees. SIAM Journal on Computing, 41(1):1–11, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v41/i1/p1_s1.

Dyer:1988:CCV

[DF88] M. E. Dyer and A. M. Frieze. On the complexity of computing the volume of a polyhedron. SIAM Journal on Computing, 17(5):967–974, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dolev:1992:DVN

[DF92] Danny Dolev and Tomás Feder. Determinism vs. nondeterminism in multiparty communication complexity. SIAM Journal on Computing, 21(5): 889–895, October 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Downey:1995:FPT

[DF95] Rod G. Downey and Michael R. Fellows. Fixed-parameter tractability and completeness. I. basic results. SIAM Journal on Computing, 24(4):873– 921, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        22822.

David:2018:RWM

[DF18] Roee David and Uriel Feige. Random walks with the minimum degree local rule have O(n^{2}) cover time. SIAM Journal on Computing, 47(3):755– 768, ???? 2018. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Deng:2013:CPR

[DFG13] Ting Deng, Wenfei Fan, and Floris Geerts. On the complexity of package recommendation problems. SIAM Journal on Computing, 42(5):1940– 1986, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Demaine:2008:CCH

[DFHS08] Erik D. Demaine, Uriel Feige, MohammadTaghi Hajiaghayi, and Mohammad R. Salavatipour. Combination can be hard: Approximability of the unique coverage problem. SIAM Journal on Computing, 38(4):1464–1483, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dyer:1998:ACH

[DFJ98] Martin Dyer, Alan Frieze, and Mark Jerrum. Approximately counting Hamilton paths and cycles in dense graphs. SIAM Journal on Computing, 27 (5):1262–1272, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26112.

Dyer:2002:CIS

[DFJ02] Martin Dyer, Alan Frieze, and Mark Jerrum. On counting independent sets in sparse graphs. SIAM Journal on Computing, 31(5):1527– 1541, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38384.

Demaine:2019:CMP

[DFK^{+}19] Erik D. Demaine, Sándor P. Fekete, Phillip Keldenich, Henk Meijer, and Christian Scheffer. Coordinated motion planning: Reconfiguring a swarm of labeled robots with bounded stretch. SIAM Journal on Computing, 48(6):1727– 1762, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dutting:2020:PIM

[DFKL20] Paul Dütting, Michal Feldman, Thomas Kesselheim, and Brendan Lucier. Prophet inequalities made easy: Stochastic optimization by pricing nonstochastic inputs. SIAM Journal on Computing, 49(3):540– 582, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Daskalakis:2022:MIN

[DFL^{+}22] Constantinos Daskalakis, Maxwell Fishelson, Brendan Lucier, Vasilis Syrgkanis, and Santhoshini Velusamy. Multiitem nontruthful auctions achieve good revenue. SIAM Journal on Computing, 51 (6):1796–1838, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1471742.

Deligkas:2023:PTA

[DFM23] Argyrios Deligkas, Michail Fasoulakis, and Evangelos Markakis. A polynomialtime algorithm for 1/2-well-supported Nash equilibria in bimatrix games. SIAM Journal on Computing, 52 (5):1083–1096, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1549237.

DiBattista:2013:QNP

[DFP13] Giuseppe Di Battista, Fabrizio Frati, and János Pach. On the queue number of planar graphs. SIAM Journal on Computing, 42(6):2243–2285, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Damgaard:2008:CBQ

[DFSS08] Ivan B. Damgård, Serge Fehr, Louis Salvail, and Christian Schaffner. Cryptography in the bounded-quantum-storage model. SIAM Journal on Computing, 37(6):1865–1890, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Downey:1999:PCS

[DFVW99] Rod G. Downey, Michael R. Fellows, Alexander Vardy, and Geoff Whittle. The parametrized complexity of some fundamental problems in coding theory. SIAM Journal on Computing, 29(2):545–570, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32357.

deGroote:1983:CDA

[dG83] Hans F. de Groote. Characterization of division algebras of minimal rank and the structure of their algorithm varieties. SIAM Journal on Computing, 12(1):101–117, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Duris:1984:TTB

[ĎG84] Pavol Ďuriš and Zvi Galil. Two tapes are better than one for nondeterministic machines. SIAM Journal on Computing, 13(2):219–227, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dur:1993:ACT

[DG93] A. Dür and J. Grabmeier. Applying coding theory to sparse interpolation. SIAM Journal on Computing, 22(4):695–704, August 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Deng:2000:PSP

[DGBL00] Xiaotie Deng, Nian Gu, Tim Brecht, and KaiCheng Lu. Preemptive scheduling of parallel jobs on multiprocessors. SIAM Journal on Computing, 30(1):145–160, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31559.

Dvir:2020:SAS

[DGGW20] Zeev Dvir, Sivakanth Gopi, Yuzhou Gu, and Avi Wigderson. Spanoids — an abstraction of spanning structures, and a barrier for LCCs. SIAM Journal on Computing, 49(3): 465–496, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dyer:1998:CCM

[DGH98] Martin Dyer, Peter Gritzmann, and Alexander Hufnagel. On the complexity of computing mixed volumes. SIAM Journal on Computing, 27(2):356–400, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27838.

Du:1998:MRC

[DGHK98] D. Z. Du, B. Gao, F. K. Hwang, and J. H. Kim. On multirate rearrangeable Clos networks. SIAM Journal on Computing, 28(2):463–470, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28471.

Datar:2002:MSS

[DGIM02] Mayur Datar, Aristides Gionis, Piotr Indyk, and Rajeev Motwani. Maintaining stream statistics over sliding windows. SIAM Journal on Computing, 31(6):1794–1813, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39836.

Dyer:2009:CWB

[DGJ09] Martin Dyer, Leslie Ann Goldberg, and Mark Jerrum. The complexity of weighted Boolean CSP. SIAM Journal on Computing, 38(5):1970– 1986, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Diakonikolas:2010:BIF

[DGJ^{+}10] Ilias Diakonikolas, Parikshit Gopalan, Ragesh Jaiswal, Rocco A. Servedio, and Emanuele Viola. Bounded independence fools halfspaces. SIAM Journal on Computing, 39(8):3441–3462, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dinur:2005:NMP

[DGKR05] Irit Dinur, Venkatesan Guruswami, Subhash Khot, and Oded Regev. A new multilayered PCP and the hardness of hypergraph vertex cover. SIAM Journal on Computing, 34(5):1129–1146, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44305.

DallAgnol:2023:STL

[DGL23] Marcel Dall Agnol, Tom Gur, and Oded Lachish. A structural theorem for local algorithms with applications to coding, testing, and verification. SIAM Journal on Computing, 52(6):1413–1463, December 2023. CODEN SMJ-CAT. ISSN 1095-7111.

Dutta:2010:FAD

[DGLV10] Partha Dutta, Rachid Guerraoui, Ron R. Levy, and Marko Vukolić. Fast access to distributed atomic memory. SIAM Journal on Computing, 39(8):3752–3783, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v39/i8/p3752_s1.

Dutta:2007:TCL

[DGP07] Partha Dutta, Rachid Guerraoui, and Bastian Pochon. The time-complexity of local decision in distributed agreement. SIAM Journal on Computing, 37(3):722–756, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Daskalakis:2009:CCN

[DGP09] Constantinos Daskalakis, Paul W. Goldberg, and Christos H. Papadimitriou. The complexity of computing a Nash equilibrium. SIAM Journal on Computing, 39(1):195–259, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Decatur:1999:CSC

[DGR99] Scott E. Decatur, Oded Goldreich, and Dana Ron. Computational sample complexity. SIAM Journal on Computing, 29(3):854–879, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32564.

DeBonis:2005:OTS

[DGV05] Annalisa De Bonis, Leszek Gasieniec, and Ugo Vaccaro. Optimal two-stage algorithms for group testing problems. SIAM Journal on Computing, 34(5):1253–1270, October 2005. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42800.

Dvir:2011:MVC

[DGY11] Zeev Dvir, Parikshit Gopalan, and Sergey Yekhanin. Matching vector codes. SIAM Journal on Computing, 40(4):1154– 1178, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i4/p1154_s1.

Du:1992:RSP

[DH92] D. Z. Du and F. K. Hwang. Reducing the Steiner problem in a normed space. SIAM Journal on Computing, 21(6):1001– 1007, December 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dosa:2004:BOA

[DH04] György Dósa and Yong He.

Better online algorithms for scheduling with machine cost. SIAM Journal on Computing, 33(5):1035–1051, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43395.

Dinur:2013:CLE

[DH13] Irit Dinur and Prahladh Harsha. Composition of low-error 2-query PCPs using decodable PCPs. SIAM Journal on Computing, 42(6):2452–2486, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dadush:2020:FSA

[DH20] Daniel Dadush and Sophie Huiberts. A friendly smoothed analysis of the simplex method. SIAM Journal on Computing, 49(5):STOC18–449–STOC18– 499, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Deng:1996:LTR

[DHH96] Xiaotie Deng, Pavol Hell, and Jing Huang. Linear-time representation algorithms for proper circular-arc graphs and proper interval graphs. SIAM Journal on Computing, 25(2):390–403, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26909.

Durr:2006:QQC

[DHHM06] Christoph Dürr, Mark Heiligman, Peter HOyer, and Mehdi Mhalla. Quantum query complexity of some graph problems. SIAM Journal on Computing, 35(6):1310–1328, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Demaine:2007:DOA

[DHIP07] Erik D. Demaine, Dion Harmon, John Iacono, and Mihai Pătraşcu. Dynamic optimality — almost. SIAM Journal on Computing, 37(1):240–251, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dey:2011:OHC

[DHK11] Tamal K. Dey, Anil N. Hirani, and Bala Krishnamoorthy. Optimal homologous cycles, total unimodularity, and linear programming. SIAM Journal on Computing, 40(4): 1026–1044, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i4/p1026_s1.

Dinur:2021:LDD

[DHK^{+}21] Irit Dinur, Prahladh Harsha, Tali Kaufman, Inbal Livni Navon, and Amnon Ta-Shma. List-decoding with double samplers. SIAM Journal on Computing, 50(2):301–349, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dolev:2012:OSS

[DHM12] Danny Dolev, Ezra N. Hoch, and Yoram Moses. An optimal self-stabilizing firing squad. SIAM Journal on Computing, 41(2):415–435, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Downey:2002:RCD

[DHN02] Rod G. Downey, Denis R. Hirschfeldt, and André Nies. Randomness, computability, and density. SIAM Journal on Computing, 31(4):1169– 1183, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        37693.

Driemel:2013:JYD

[DHP13] Anne Driemel and Sariel Har-Peled. Jaywalking your dog: Computing the Fréchet distance with shortcuts. SIAM Journal on Computing, 42(5): 1830–1866, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dwork:1999:TLS

[DHPW99] Cynthia Dwork, Maurice Herlihy, Serge Plotkin, and Orli Waarts. Time-lapse snapshots. SIAM Journal on Computing, 28(5):1848–1874, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24368.

Driscoll:1997:FDP

[DHR97] J. R. Driscoll, D. M. Healy, Jr., and D. N. Rockmore. Fast discrete polynomial transforms with applications to data analysis for distance transitive graphs. SIAM Journal on Computing, 26(4):1066– 1099, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
24012.

DeLoera:2020:MEN

[DHR20] Jesús A. De Loera, Jamie Haddock, and Luis Rademacher. The minimum Euclidean-norm point in a convex polytope: Wolfe’s combinatorial algorithm is exponential. SIAM Journal on Computing, 49(1): 138–169, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ducoffe:2022:DED

[DHV22] Guillaume Ducoffe, Michel Habib, and Laurent Viennot. Diameter, eccentricities and distance oracle computations on H-minor free graphs and graphs of bounded (distance) Vapnik–Chervonenkis dimension. SIAM Journal on Computing, 51(5):1506–1534, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M136551X.

Dwork:1998:PWE

[DHW98] Cynthia Dwork, Joseph Y.

Halpern, and Orli Waarts. Performing work efficiently in the presence of faults. SIAM Journal on Computing, 27(5):1457–1491, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25552.

Dor:2000:APA

[DHZ00] Dorit Dor, Shay Halperin, and Uri Zwick. All-pairs almost shortest paths. SIAM Journal on Computing, 29 (5):1740–1759, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32790.

Dolev:1997:RBS

[DIM97] Shlomi Dolev, Amos Israeli, and Shlomo Moran. Resource bounds for selfstabilizing message-driven protocols. SIAM Journal on Computing, 26(1):273–290, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23507.

Decker:2013:HSS

[DISW13] Thomas Decker, Gábor Ivanyos, Miklos Santha, and Pawel Wocjan. Hidden symmetry subgroup problems. SIAM Journal on Computing, 42(5): 1987–2007, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dahlhaus:1994:CMC

[DJP^{+}94] E. Dahlhaus, D. S. Johnson, C. H. Papadimitriou, P. D. Seymour, and M. Yannakakis. The complexity of multiterminal cuts. SIAM Journal on Computing, 23(4):864– 894, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
22529.

Devroye:2000:ST

[DJZC00] Luc Devroye, Jean Jabbour, and Carlos Zamora-Cura. Squarish k–d trees. SIAM Journal on Computing, 30(5):1678–1700, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35892.

Dinitz:2014:MSR

[DK14] Michael Dinitz and Guy Kortsarz. Matroid secretary for regular and decomposable matroids. SIAM Journal on Computing, 43(5):1807–1830, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

DeMarco:2015:FND

[DK15] Gianluca De Marco and Dariusz R. Kowalski. Fast nonadaptive deterministic algorithm for conflict resolution in a dynamic multipleaccess channel. SIAM Journal on Computing, 44(3):868–888, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Daskalakis:2018:SSF

[DKI18] Costis Daskalakis, Yael Kalai, and Sandy Irani. Special section on the Forty-Seventh Annual ACM Symposium on Theory of Computing (STOC 2015). SIAM Journal on Computing, 47(3):888–889, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dalmau:2019:RAP

[DKK^{+}19a] Víctor Dalmau, Marcin Kozik, Andrei Krokhin, Konstantin Makarychev, Yury Makarychev, and Jakub Oprsal. Robust algorithms with polynomial loss for near-unanimity CSPs. SIAM Journal on Computing, 48(6):1763–1795, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Diakonikolas:2019:REH

[DKK^{+}19b] Ilias Diakonikolas, Gautam Kamath, Daniel Kane, Jerry Li, Ankur Moitra, and Alistair Stewart. Robust estimators in high-dimensions without the computational intractability. SIAM Journal on Computing, 48(2):742–864, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dolev:1999:BAR

[DKKP99] Shlomi Dolev, Evangelos Kranakis, Danny Krizanc, and David Peleg. Bubbles: Adaptive routing scheme for high-speed dynamic networks. SIAM Journal on Computing, 29(3):804–833, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31661.

Dagum:2000:OAM

[DKLR00] Paul Dagum, Richard Karp, Michael Luby, and Sheldon Ross. An optimal algorithm for Monte Carlo estimation. SIAM Journal on Computing, 29(5):1484–1496, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31530.

Dietzfelbinger:1994:DPH

[DKM^{+}94] Martin Dietzfelbinger, Anna Karlin, Kurt Mehlhorn, Friedhelm Meyer auf der Heide, Hans Rohnert, and Robert E. Tarjan. Dynamic perfect hashing: Upper and lower bounds. SIAM Journal on Computing, 23(4):738–761, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19409.

Drineas:2006:FMCa

[DKM06a] Petros Drineas, Ravi Kannan, and Michael W. Mahoney. Fast Monte Carlo algorithms for matrices I: Approximating matrix multiplication. SIAM Journal on Computing, 36(1): 132–157, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Drineas:2006:FMCb

[DKM06b] Petros Drineas, Ravi Kannan, and Michael W. Mahoney. Fast Monte Carlo algorithms for matrices II: Computing a low-rank approximation to a matrix. SIAM Journal on Computing, 36(1):158– 183, ???? 2006. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Drineas:2006:FMCc

[DKM06c] Petros Drineas, Ravi Kannan, and Michael W. Mahoney. Fast Monte Carlo algorithms for matrices III: Computing a compressed approximate matrix decomposition. SIAM Journal on Computing, 36(1): 184–206, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Devroye:2009:RHS

[DKM09] Luc Devroye, James King, and Colin McDiarmid. Random hyperplane search trees. SIAM Journal on Computing, 38(6):

2411–2425, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Daskalakis:2011:SSP

[DKM^{+}11] Constantinos Daskalakis, Richard M.

Karp, Elchanan Mossel, Samantha J. Riesenfeld, and Elad Verbin. Sorting and selection in posets. SIAM Journal on Computing, 40(3):597–622, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i3/
p597_s1.

Dietzfelbinger:1996:FTO

[DKR96] Martin Dietzfelbinger, Mirosław Kutyłowski, and Rüdiger Reischuk. Feasible time-optimal algorithms for Boolean functions on exclusive-write parallel random-access machines. SIAM Journal on Computing, 25(6):1196–1230, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22428.

Dwork:1988:PAT

[DKS88] Cynthia Dwork, Paris C.

Kanellakis, and Larry Stockmeyer. Parallel algorithms for term matching. SIAM Journal on Computing, 17(4):711–731, August 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

De:2013:FIM

[DKSS13a] Anindya De, Piyush P. Kurur, Chandan Saha, and Ramprasad Saptharishi. Fast integer multiplication using modular arithmetic. SIAM Journal on Computing, 42(2):685–699, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dvir:2013:EMM

[DKSS13b] Zeev Dvir, Swastik Kopparty, Shubhangi Saraf, and Madhu Sudan. Extensions to the method of multiplicities, with applications to Kakeya sets and mergers. SIAM Journal on Computing, 42(6):2305– 2328, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Drucker:2022:SSF

[DKSS22] Andy Drucker, Ravi Kumar, Amit Sahai, and Mohit Singh. Special section on the Forty-Ninth Annual ACM Symposium on the Theory of Computing (STOC 2017). SIAM Journal on Computing, 51(2): STOC17–i–STOC17–ii, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22N975482.

Dobkin:1976:MSP

[DL76] David P. Dobkin and Richard J. Lipton. Multidimensional searching problems. SIAM Journal on Computing, 5(2): 181–186, June 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dobkin:1980:ACM

[DL80] David Dobkin and Richard J.

Lipton. Addition chain methods for the evaluation of specific polynomials. SIAM Journal on Computing, 9(1):121– 125, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Devroye:1990:ARD

[DL90] Luc Devroye and Louise Laforest. An analysis of random ddimensional quadtrees. SIAM Journal on Computing, 19(5): 821–832, October 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Davida:1991:FPA

[DL91] George I. Davida and Bruce Litow. Fast parallel arithmetic via modular representation. SIAM Journal on Computing, 20(4):756–765, August 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Deng:2003:GDD

[DLL^{+}03] Xiaotie Deng, Guojun Li, Zimao Li, Bin Ma, and Lusheng Wang. Genetic design of drugs without side-effects. SIAM Journal on Computing, 32 (4):1073–1090, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39782.

Dell:2022:ACS

[DLM22] Holger Dell, John Lapinskas, and Kitty Meeks. Approximately counting and sampling small witnesses using a colorful decision oracle. SIAM Journal on Computing, 51(4):849–899, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M130604X.

Duke:1995:FAA

[DLR95] Richard A. Duke, Hanno Lefmann, and Vojtěch Rödl. A fast approximation algorithm for computing the frequencies of subgraphs in a given graph. SIAM Journal on Computing, 24(3):598–620, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24763.

Downey:1981:CSA

[DLS81] Peter Downey, Benton Leong, and Ravi Sethi. Computing sequences with addition chains. SIAM Journal on Computing, 10(3):638–646, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

DiBattista:1998:SOO

[DLV98] Giuseppe Di Battista, Giuseppe Liotta, and Francesco Vargiu. Spirality and optimal orthogonal drawings. SIAM Journal on Computing, 27(6): 1764–1811, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26284.

Dodis:2014:PAN

[DLWZ14] Yevgeniy Dodis, Xin Li, Trevor D. Wooley, and David Zuckerman. Privacy amplification and nonmalleable extractors via character sums. SIAM Journal on Computing, 43(2): 800–830, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

DAtri:1988:DHG

[DM88] Alessandro D’Atri and Marina Moscarini. Distancehereditary graphs, Steiner trees, and connected domination. SIAM Journal on Computing, 17(3):521–538, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Deng:1997:CDC

[DM97] Xiaotie Deng and Sanjeev Mahajan. The cost of derandomization: Computability or competitiveness. SIAM Journal on Computing, 26(3):786– 802, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        20230.

Dinur:2019:SSF

[DMK19] Irit Dinur, Or Meir, and Swastik Kopparty. Special section on the Fifty-Seventh Annual IEEE Symposium on Foundations of Computer Science (FOCS 2016). SIAM Journal on Computing, 48(2): 451, ???? 2019. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dinur:2009:CHA

[DMR09] Irit Dinur, Elchanan Mossel, and Oded Regev. Conditional hardness for approximate coloring. SIAM Journal on Computing, 39(3):843–873, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dirksen:2024:FME

[DMS24] Sjoerd Dirksen, Shahar Mendelson, and Alexander Stollenwerk. Fast metric embedding into the Hamming cube. SIAM Journal on Computing, 53(2): 315–345, March 2024. CODEN SMJCAT. ISSN 1095-7111.

Dershowitz:1988:EUC

[DMT88] Nachum Dershowitz, Leo Marcus, and Andrzej Tarlecki. Existence, uniqueness, and construction of rewrite systems. SIAM Journal on Computing, 17(4):629–639, August 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Devroye:2004:WCR

[DMV04] Luc Devroye, Pat Morin, and Alfredo Viola. On worstcase Robin Hood hashing. SIAM Journal on Computing, 33(4):923–936, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40337.

Dujmovic:2005:LGB

[DMW05] Vida Dujmovic, Pat Morin, and David R. Wood. Layout of graphs with bounded tree-width. SIAM Journal on Computing, 34(3):553–579, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41614.

Devroye:2004:DFS

[DN04] Luc Devroye and Ralph Neininger. Distances and finger search in random binary search trees. SIAM Journal on Computing, 33(3):647–658, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42452.

Dwork:2007:ZTA

[DN07] Cynthia Dwork and Moni Naor. Zaps and their applications. SIAM Journal on Computing, 36(6):1513–1543, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dekel:1981:PMG

[DNS81] Eliezer Dekel, David Nassimi, and Sartaj Sahni. Parallel matrix and graph algorithms. SIAM Journal on Computing, 10(4):657–675, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

DeLoera:2004:CTW

[DO04] Jesus De Loera and Shmuel Onn. The complexity of three-way statistical tables. SIAM Journal on Computing, 33(4):819–836, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40380.

Dobson:1984:SIT

[Dob84] Gregory Dobson. Scheduling independent tasks on uniform processors. SIAM Journal on Computing, 13(4):705– 716, ???? 1984. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Doberkat:1989:TCI

[Dob89] Ernst-Erich Doberkat. Topological completeness in an ideal model for polymorphic types. SIAM Journal on Computing, 18(5):977–989, October 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Doberkat:2005:SRC

[Dob05] Ernst-Erich Doberkat. Stochastic relations: Congruences, bisimulations and the Hennessy– Milner theorem. SIAM Journal on Computing, 35(3):590–626, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dobzinski:2021:BLB

[Dob21] Shahar Dobzinski. Breaking the logarithmic barrier for truthful combinatorial auctions with submodular bidders. SIAM Journal on Computing, 50(3):STOC16–1– STOC16–17, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Doerr:2004:NRR

[Doe04] Benjamin Doerr. Nonindependent randomized rounding and an application to digital halftoning. SIAM Journal on Computing, 34(2):299–317, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43015.

Donahue:1979:SDT

[Don79] James Donahue. On the semantics of “data type”. SIAM Journal on Computing, 8(4): 546–560, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dodis:2008:FEH

[DORS08] Yevgeniy Dodis, Rafail Ostrovsky, Leonid Reyzin, and Adam Smith. Fuzzy extractors: How to generate strong keys from biometrics and other noisy data. SIAM Journal on Computing, 38(1):97–139, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

deBerg:1994:CVD

[dOS94] Mark de Berg, Mark Overmars, and Otfried Schwarzkopf. Computing and verifying depth orders. SIAM Journal on Computing, 23(2):437–446, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22374.

Doty:2010:RSA

[Dot10]  David Doty.  Randomized

self-assembly for exact shapes. SIAM Journal on Computing, 39(8):3521–3552, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dowling:1990:FPH

[Dow90]  Michael L. Dowling.  A

fast parallel Horner algorithm. SIAM Journal on Computing, 19(1):133–142, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

DeSantis:1994:TUL

[DP94a] Alfredo De Santis and Giuseppe Persiano. Tight upper and lower bounds on the path length of binary trees. SIAM Journal on Computing, 23 (1):12–23, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/19732.

Du:1994:CGT

[DP94b] Ding Zhu Du and Haesun Park. On competitive group testing. SIAM Journal on Computing, 23(5):1019–1025, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
24669.

Duan:2020:COG

[DP20] Ran Duan and Seth Pettie. Connectivity oracles for graphs subject to vertex failures. SIAM Journal on Computing, 49(6):1363–1396, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Durfee:2020:DPS

[DPPR20] David Durfee, John Peebles, Richard Peng, and Anup B. Rao. Determinant-preserving sparsification of SDDM matrices. SIAM Journal on Computing, 49(4):FOCS17–350– FOCS17–408, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dwork:1988:FTN

[DPPU88] Cynthia Dwork, David Peleg, Nicholas Pippenger, and Eli Upfal. Fault tolerance in networks of bounded degree. SIAM Journal on Computing, 17(5):975–988, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dieudonne:2015:HMA

[DPV15] Yoann Dieudonné, Andrzej Pelc, and Vincent Villain. How to meet asynchronously at polynomial cost. SIAM Journal on Computing, 44(3):844– 867, ???? 2015. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

De:2012:TEP

[DPVR12] Anindya De, Christopher Portmann, Thomas Vidick, and Renato Renner. Trevisan’s extractor in the presence of quantum side information. SIAM Journal on Computing, 41(4): 915–940, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Delcher:1995:AEM

[DR95a] A. L. Delcher and S. Rao Kosaraju. An N C algorithm for evaluating monotone planar circuits. SIAM Journal on Computing, 24(2):369–375, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22627.

Devroye:1995:VHR

[DR95b] Luc Devroye and Bruce Reed.

On the variance of the height of random binary search trees. SIAM Journal on Computing, 24(6):1157–1162, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23754.

Devroye:1995:GRB

[DR95c] Luc Devroye and John Michael Robson. On the generation of random binary search trees. SIAM Journal on Computing, 24(6):1141–1156, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22495.

Dinur:2006:ATT

[DR06] Irit Dinur and Omer Reingold.

Assignment testers: Towards a combinatorial proof of the PCP theorem. SIAM Journal on Computing, 36(4):975– 1024, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dyer:2013:EDC

[DR13] Martin Dyer and David Richerby. An effective dichotomy for the counting constraint satisfaction problem. SIAM Journal on Computing, 42(3):1245–1274, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dughmi:2014:BBR

[DR14] Shaddin Dughmi and Tim Roughgarden. Black-box randomized reductions in algorithmic mechanism design. SIAM Journal on Computing, 43(1): 312–336, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Diakonikolas:2014:ASN

[DRST14] Ilias Diakonikolas, Prasad Raghavendra, Rocco A. Servedio, and Li-Yang Tan. Average sensitivity and noise sensitivity of polynomial threshold functions. SIAM Journal on Computing, 43(1):231–253, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dixon:1992:VSA

[DRT92] Brandon Dixon, Monika Rauch, and Robert E. Tarjan. Verification and sensitivity analysis of minimum spanning trees in linear time. SIAM Journal on Computing, 21(6):1184–1192, December 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dutting:2021:CC

[DRTC21] P. Dütting, T. Roughgarden, and I. Talgam-Cohen. The complexity of contracts. SIAM Journal on Computing, 50(1): 211–254, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dixit:2018:ERP

[DRTV18] Kashyap Dixit, Sofya Raskhodnikova, Abhradeep Thakurta, and Nithin Varma. Erasureresilient property testing. SIAM Journal on Computing, 47(2):295–329, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Drucker:2015:NLC

[Dru15] Andrew Drucker. New limits to classical and quantum instance compression. SIAM Journal on Computing, 44(5): 1443–1479, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dudek:1998:LRM

[DRW98] Gregory Dudek, Kathleen Romanik, and Sue Whitesides.

Localizing a robot with minimum travel. SIAM Journal on Computing, 27(2):583–604, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27920.

Downey:1976:CCR

[DS76] Peter J. Downey and Ravi Sethi. Correct computation rules for recursive languages. SIAM Journal on Computing, 5(3):378–401, September 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dolev:1983:AAB

[DS83] D. Dolev and H. R. Strong. Authenticated algorithms for Byzantine agreement. SIAM Journal on Computing, 12(4): 656–666, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dwork:1990:TCG

[DS90] Cynthia Dwork and Larry Stockmeyer. A time complexity gap for two-way probabilistic finite-state automata. SIAM Journal on Computing, 19(6):1011–1023, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Deogun:1994:PAH

[DS94] Jitender S. Deogun and George Steiner. Polynomial algorithms for Hamiltonian cycle in cocomparability graphs. SIAM Journal on Computing, 23(3):520–552, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/20037.

Dolev:1997:BCT

[DS97] Danny Dolev and Nir Shavit.

Bounded concurrent timestamping. SIAM Journal on Computing, 26(2):418–455, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19264.

Dyer:2000:FOP

[DS00] Martin E. Dyer and Sandeep Sen. Fast and optimal parallel multidimensional search in PRAMs with applications to linear programming and related problems. SIAM Journal on Computing, 30(5):1443– 1461, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32572.

Durr:2002:DPU

[DS02] Christoph Dürr and Miklos Santha. A decision procedure for unitary linear quantum cellular automata. SIAM Journal on Computing, 31 (4):1076–1089, August 2002.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32770.

Dvir:2007:LDC

[DS07] Zeev Dvir and Amir Shpilka. Locally decodable codes with two queries and polynomial identity testing for depth 3 circuits. SIAM Journal on Computing, 36(5):1404–1434, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Devroye:1992:NHS

[DSR92] Luc Devroye, Wojciech Szpankowski, and Bonita Rais. A note on the height of suffix trees. SIAM Journal on Computing, 21(1):48–53, February 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dwork:1990:FPC

[DSS90] Cynthia Dwork, David Shmoys, and Larry Stockmeyer. Flipping persuasively in constant time. SIAM Journal on Computing, 19(3):472–499, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Daniely:2018:ITM

[DSS18] Amit Daniely, Michael Schapira, and Gal Shahaf. Inapproximability of truthful mechanisms via generalizations of the Vapnik–Chervonenkis dimension. SIAM Journal on Computing, 47(1):96–120, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dekel:2008:FKB

[DSSS08] Ofer Dekel, Shai Shalev-Shwartz, and Yoram Singer. The Forgetron: a kernel-based perceptron on a budget. SIAM Journal on Computing, 37(5): 1342–1372, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dvir:2010:HRT

[DSY10] Zeev Dvir, Amir Shpilka, and Amir Yehudayoff. Hardnessrandomness tradeoffs for bounded depth arithmetic circuits. SIAM Journal on Computing, 39(4):1279–1293, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

DiBattista:1996:LPT

[DT96] Giuseppe Di Battista and Roberto Tamassia. On-line planarity testing. SIAM Journal on Computing, 25(5):956– 997, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28073.

Dor:1997:GDN

[DT97] Dorit Dor and Michael Tarsi.

Graph decomposition is NP-Complete: a complete proof of Holyer’s conjecture. SIAM Journal on Computing, 26 (4):1166–1187, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22950.

Dinur:2008:SIF

[DT08] Irit Dinur and Eva Tardos. Special issue on foundations of computer science. SIAM Journal on Computing, 37(6):vii, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Das:1994:DDS

[DTA94] A. Das, K. Thulasiraman, and V. K. Agarwal. Diagnosis of t/(t + 1)-diagnosable systems. SIAM Journal on Computing, 23(5):895–905, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18753.

Demetrescu:2008:ODA

[DTCR08] Camil Demetrescu, Mikkel Thorup, Rezaul Alam Chowdhury, and Vijaya Ramachandran. Oracles for distances avoiding a failed node or link. SIAM Journal on Computing, 37(5):1299–1318, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dube:1990:SPI

[Dub90] Thomas W. Dubé. The structure of polynomial ideals and gröbner bases. SIAM Journal on Computing, 19(4):750–773, August 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dunn:1973:CGA

[Dun73] J. C. Dunn. Continuous group averaging and pattern classification problems. SIAM Journal on Computing, 2(4):253– 272, ???? 1973. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Duris:2010:EPT

[Dur10] David Duris. Extension preservation theorems on classes of acyclic finite structures. SIAM Journal on Computing, 39(8): 3670–3681, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Duval:1991:AFP

[Duv91] Dominique Duval. Absolute factorization of polynomials: a geometric approach. SIAM Journal on Computing, 20(1): 1–21, February 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dinitz:2000:GSE

[DV00] Yefim Dinitz and Alek Vainshtein. The general structure of edge-connectivity of a vertex subset in a graph and its incremental maintenance. odd case. SIAM Journal on Computing, 30(3):753–808, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33004.

deVerdiere:2010:TNP

[dVE10] Éric Colin de Verdière and Jeff Erickson. Tightening nonsimple paths and cycles on surfaces. SIAM Journal on Computing, 39(8):3784–3813, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v39/i8/
p3784_s1.

Diehl:2006:TSL

[DvM06] Scott Diehl and Dieter van Melkebeek. Time-space lower bounds for the polynomialtime hierarchy on randomized machines. SIAM Journal on Computing, 36(3):563–594, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dalirrooyfard:2021:GPD

[DVW21] Mina Dalirrooyfard, Thuy Duong Vuong, and Virginia Vassilevska Williams. Graph pattern detection: Hardness for all induced patterns and faster noninduced cycles. SIAM Journal on Computing, 50(5): 1627–1662, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Davis:1982:CPQ

[DW82] Henry W. Davis and Leon E.

Winslow. Computational power in query languages. SIAM Journal on Computing, 11(3):547–554, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dolev:1985:SFG

[DW85] Danny Dolev and Manfred Warmuth. Scheduling flat graphs. SIAM Journal on Computing, 14(3):638–657, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

deWolf:2003:NQQ

[dW03] Ronald de Wolf. Nondeterministic quantum query and communication complexities. SIAM Journal on Computing, 32(3):681–699, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40734.

Dvir:2011:KSN

[DW11] Zeev Dvir and Avi Wigderson. Kakeya sets, new mergers, and old extractors. SIAM Journal on Computing, 40(3): 778–792, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i3/p778_s1.

Dughmi:2021:ABP

[DX21] Shaddin Dughmi and Haifeng Xu. Algorithmic Bayesian persuasion. SIAM Journal on Computing, 50(3):STOC16– 68–STOC16–97, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Du:1994:MCG

[DXSC94] D.-Z. Du, G.-L. Xue, S.-Z. Sun, and S.-W. Cheng. Modifications of competitive group testing. SIAM Journal on Computing, 23(1):82–96, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22761.

Drysdale:1975:IDS

[DY75] R. L. (Scot) Drysdale, III and Frank H. Young. Improved divide/sort/merge sorting networks. SIAM Journal on Computing, 4(3):264–270, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Diakonikolas:2010:SAP

[DY10] Ilias Diakonikolas and Mihalis Yannakakis. Small approximate Pareto sets for biobjective shortest paths and other problems. SIAM Journal on Computing, 39(4):1340–1371, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dyer:1984:LTA

[Dye84] M. E. Dyer. Linear time algorithms for two- and three-variable linear programs. SIAM Journal on Computing, 13(1):31–45, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dyer:1986:MST

[Dye86] M. E. Dyer. On a multidimensional search technique and its application to the Euclidean one-centre problem. SIAM Journal on Computing, 15(3): 725–738, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dyer:1991:CLP

[Dye91] Martin Dyer. On counting lattice points in polyhedra. SIAM Journal on Computing, 20(4): 695–707, August 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Dubiner:1997:ARO

[DZ97] Moshe Dubiner and Uri Zwick.

Amplification by read-once formulas. SIAM Journal on Computing, 26(1):15–38, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23633.

Dor:1999:SM

[DZ99] Dorit Dor and Uri Zwick. Selecting the median. SIAM Journal on Computing, 28 (5):1722–1758, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28861.

Eberly:1989:VFP

[Ebe89] Wayne Eberly. Very fast parallel polynomial arithmetic. SIAM Journal on Computing, 18(5):955–976, October 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Emiris:1995:GAR

[EC95] Ioannis Z. Emiris and John F. Canny. A general approach to removing degeneracies. SIAM Journal on Computing, 24(3): 650–664, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23591.

Ellis:2003:MOE

[ECM03] John Ellis, Stirling Chow, and Dennis Manke. Many to one embeddings from grids into cylinders, tori, and hypercubes. SIAM Journal on Computing, 32(2):386–407, April 2003. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37765.

Chen:1991:RC

[eCY91] Jian er Chen and Chee-Keng Yap. Reversal complexity. SIAM Journal on Computing, 20(4):622–638, August 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Edmonds:1998:TST

[Edm98]  Jeff A. Edmonds.  Time-

space tradeoffs for undirected st-connectivity on a graph automata. SIAM Journal on Computing, 27(5):1492– 1513, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
27713.

Elkin:2008:LSS

[EEST08] Michael Elkin, Yuval Emek, Daniel A. Spielman, and Shang-Hua Teng. Lowerstretch spanning trees. SIAM Journal on Computing, 38(2): 608–628, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Elkin:2018:PMS

[EFN18] Michael Elkin, Arnold Filtser, and Ofer Neiman. Prioritized metric structures and embedding. SIAM Journal on Computing, 47(3):829–858, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Efrat:2005:CUC

[Efr05] Alon Efrat. The complexity of the union of (α, β)-covered objects. SIAM Journal on Computing, 34(4):775–787, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40751.

Efremenko:2012:QLD

[Efr12] Klim Efremenko. 3-query locally decodable codes of subexponential length. SIAM Journal on Computing, 41(6):1694– 1703, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Efthymiou:2016:SAS

[Eft16] Charilaos Efthymiou. A simple algorithm for sampling colorings of G(n, d/n) up to the Gibbs uniqueness threshold. SIAM Journal on Computing, 45(6):2087–2116, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eiter:1995:IMT

[EG95] Thomas Eiter and Georg Gottlob. Identifying the minimal transversals of a hypergraph and related problems. SIAM Journal on Computing, 24(6):1278–1304, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25029.

Eppstein:2007:ICG

[EGH07] David Eppstein, Michael T.

Goodrich, and Daniel S. Hirschberg. Improved combinatorial group testing algorithms for real-world problem sizes. SIAM Journal on Computing, 36(5):1360–1375, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eppstein:1998:SBS

[EGIS98] David Eppstein, Zvi Galil, Giuseppe F. Italiano, and Thomas H. Spencer. Separatorbased sparsification II: Edge and vertex connectivity. SIAM Journal on Computing, 28 (1):341–381, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26907.

Englert:2014:VSN

[EGK^{+}14] Matthias Englert, Anupam Gupta, Robert Krauthgamer, Harald Räcke, Inbal Talgam-Cohen, and Kunal Talwar. Vertex sparsifiers: New results from old techniques. SIAM Journal on Computing, 43(4): 1239–1262, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eiter:2003:NRM

[EGM03] Thomas Eiter, Georg Gottlob, and Kazuhisa Makino. New results on Monotone dualization and generating hypergraph transversals. SIAM Journal on Computing, 32(2):514–537, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40639.

Eisenbrand:2007:NAV

[EGOS07] Friedrich Eisenbrand, Fabrizio Grandoni, Gianpaolo Oriolo, and Martin Skutella. New approaches for virtual private network design. SIAM Journal on Computing, 37(3):706–721, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Edelsbrunner:1986:OPL

[EGS86] Herbert Edelsbrunner, Leonidas J. Guibas, and Jorge Stolfi. Optimal point location in a monotone subdivision. SIAM Journal on Computing, 15(2):317– 340, ???? 1986. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Even:2003:IAC

[EGS03] Guy Even, Sudipto Guha, and Baruch Schieber. Improved approximations of crossings in graph drawings and VLSI layout areas. SIAM Journal on Computing, 32(1):231– 252, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37352.

Eppstein:2010:LTA

[EGS10] David Eppstein, Michael T.

Goodrich, and Darren Strash. Linear-time algorithms for geometric graphs with sublinearly many edge crossings. SIAM Journal on Computing, 39(8): 3814–3829, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v39/i8/p3814_s1.

Etzioni:2000:OIG

[EHJM00] Oren Etzioni, Steve Hanks, Tao Jiang, and Omid Madani. Optimal information gathering on the Internet with time and cost constraints. SIAM Journal on Computing, 29 (5):1596–1620, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31446.

Emek:2012:OSP

[EHM^{+}12] Yuval Emek, Magnús M.

Halldórsson, Yishay Mansour, Boaz Patt-Shamir, Jaikumar Radhakrishnan, and Dror Rawitz. Online set packing.

SIAM Journal on Computing, 41(4):728–746, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ene:2017:GPU

[EHPR17] Alina Ene, Sariel Har-Peled, and Benjamin Raichel. Geometric packing under nonuniform constraints. SIAM Journal on Computing, 46(6):1745– 1784, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ellen:2012:ISC

[EHS12] Faith Ellen, Danny Hendler, and Nir Shavit. On the inherent sequentiality of concurrent objects. SIAM Journal on Computing, 41(3):519–536, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Efthymiou:2019:CML

[EHS^{+}19] Charilaos Efthymiou, Thomas P. Hayes, Daniel Stefankovic, Eric Vigoda, and Yitong Yin. Convergence of MCMC and loopy BP in the tree uniqueness region for the hard-core model. SIAM Journal on Computing, 48(2):581–643, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eschen:1999:WTC

[EHSS99] Elaine Eschen, Ryan B. Hayward, Jeremy Spinrad, and R. Sritharan. Weakly triangulated comparability graphs. SIAM Journal on Computing, 29(2):378–386, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32233.

Esteban:2017:CSA

[EiC17] Juan Luis Esteban and Ramon Ferrer i Cancho. A correction on Shiloach’s algorithm for minimum linear arrangement of trees. SIAM Journal on Computing, 46(3):1146– 1151, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Shi79a].

Eiter:2001:DHT

[EIM01] Thomas Eiter, Toshihide Ibaraki, and Kazuhisa Makino. Disjunctions of Horn theories and their cores. SIAM Journal on Computing, 31 (1):269–288, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35084.

Even:1976:CTM

[EIS76] S. Even, A. Itai, and A. Shamir.

On the complexity of timetable and multicommodity flow problems. SIAM Journal on Computing, 5(4):691–703, December 1976. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Emerson:1999:CTA

[EJ99] E. Allen Emerson and Charanjit S. Jutla. The complexity of tree automata and logics of programs. SIAM Journal on Computing, 29 (1):132–158, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30474.

Erlebach:2005:PTA

[EJS05] Thomas Erlebach, Klaus Jansen, and Eike Seidel. Polynomial-time approximation schemes for geometric intersection graphs. SIAM Journal on Computing, 34(6): 1302–1323, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40267.

Elkin:2005:CLA

[EK05] Michael Elkin and Guy Kortsarz. A combinatorial logarithmic approximation algorithm for the directed telephone broadcast problem. SIAM Journal on Computing, 35(3): 672–689, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Esparza:2010:CLF

[EKL10] Javier Esparza, Stefan Kiefer, and Michael Luttenberger. Computing the least fixed point of positive polynomial systems. SIAM Journal on Computing, 39(6):2282–2335, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Evans:2016:MCL

[EKLS16] William Evans, David Kirkpatrick, Maarten Löffler, and Frank Staals. Minimizing colocation potential of moving entities. SIAM Journal on Computing, 45(5):1870–1893, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ergun:2001:CAC

[EKR01] Funda Ergün, S. Ravi Kumar, and Ronitt Rubinfeld. Checking approximate computations of polynomials and functional equations. SIAM Journal on Computing, 31(2):550–576, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33761.

Ergun:2000:STG

[EKS00] Funda Ergün, S. Ravi Kumar, and D. Sivakumar. Self-testing without the generator bottleneck. SIAM Journal on Computing, 29(5):1630–1651, October 2000. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31116.

Etzion:1986:EAG

[EL86] T. Etzion and A. Lempel. An efficient algorithm for generating linear transformations in a shuffle-exchange network. SIAM Journal on Computing, 15(1):216–221, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Epstein:2008:AGC

[EL08] Leah Epstein and Asaf Levin. An APTAS for generalized cost variable-sized bin packing. SIAM Journal on Computing, 38(1):411–428, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Elkin:2006:ULB

[Elk06] Michael Elkin. An unconditional lower bound on the timeapproximation trade-off for the distributed minimum spanning tree problem. SIAM Journal on Computing, 36(2):433–456, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Epstein:2012:USU

[ELMS^{+}12] Leah Epstein, Asaf Levin, Alberto Marchetti-Spaccamela, Nicole Megow, Julián Mestre, Martin Skutella, and Leen Stougie. Universal sequencing on an unreliable machine. SIAM Journal on Computing, 41(3):565–586, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Even:2003:CFC

[ELRS03] Guy Even, Zvi Lotker, Dana Ron, and Shakhar Smorodinsky. Conflict-free colorings of simple geometric regions with applications to frequency assignment in cellular networks. SIAM Journal on Computing, 33(1):94–136, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43184.

Eden:2017:ACT

[ELRS17] Talya Eden, Amit Levi, Dana Ron, and C. Seshadhri. Approximately counting triangles in sublinear time. SIAM Journal on Computing, 46(5):1603– 1646, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Engelfriet:2003:MTT

[EM03] Joost Engelfriet and Sebastian Maneth. Macro tree translations of linear size increase are MSO definable. SIAM Journal on Computing, 32(4):950–1006, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39451.

El-Mallah:1985:OCS

[EMC85] Ehab S. El-Mallah and Charles J. Colbourn. Optimum communication spanning trees in seriesparallel networks. SIAM Journal on Computing, 14(4):915– 925, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

El-Mabrouk:2003:RDG

[EMS03] Nadia El-Mabrouk and David Sankoff. The reconstruction of doubled genomes. SIAM Journal on Computing, 32(3): 754–792, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        37717.

Eppstein:2012:AUC

[EMSV12] David Eppstein, Elena Mumford, Bettina Speckmann, and Kevin Verbeek. Area-universal and constrained rectangular layouts. SIAM Journal on Computing, 41(3):537–564, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Engelberg:2016:EOG

[EN16] Roee Engelberg and Joseph Seffi Naor. Equilibria in online games. SIAM Journal on Computing, 45(2):232–267, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Elkin:2019:HCH

[EN19] Michael Elkin and Ofer Neiman.

Hopsets with constant hopbound, and applications to approximate shortest paths. SIAM Journal on Computing, 48(4):1436–1480, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Engelfriet:1986:CLG

[Eng86] Joost Engelfriet. The complexity of languages generated by attribute grammars. SIAM Journal on Computing, 15(1): 70–86, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Even:1999:FAG

[ENRS99] Guy Even, Joseph (Seffi)

Naor, Satish Rao, and Baruch Schieber. Fast approximate graph partitioning algorithms. SIAM Journal on Computing, 28(6):2187–2214, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30821.

Even:2000:AAS

[ENZ00] Guy Even, Joseph (Seffi) Naor, and Leonid Zosin. An 8approximation algorithm for the subset feedback vertex set problem. SIAM Journal on Computing, 30(4):1231– 1252, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        34004.

Edelsbrunner:1986:CAL

[EOS86] H. Edelsbrunner, J. O’Rourke, and R. Seidel. Constructing Arrangements of Lines and Hyperplanes with Applications. SIAM Journal on Computing, 15(2):341–363, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Englert:2014:PRO

[EÖW14] Matthias Englert, Deniz Özmen, and Matthias Westermann. The power of reordering for online minimum makespan scheduling. SIAM Journal on Computing, 43(3):1220–1237, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Evans:1998:ACL

[EP98] William Evans and Nicholas Pippenger. Average-case lower bounds for noisy Boolean decision trees. SIAM Journal on Computing, 28(2):433–446, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31010.

Elkin:2004:SCG

[EP04] Michael Elkin and David Peleg. (1 + ϵ, β)-spanner constructions for general graphs.

SIAM Journal on Computing, 33(3):608–631, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39338.

Emek:2009:AMM

[EP09] Yuval Emek and David Peleg.

Approximating minimum maxstretch spanning trees on unweighted graphs. SIAM Journal on Computing, 38(5):1761– 1781, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Edmonds:1999:TLB

[EPA99] Jeff Edmonds, Chung Keung Poon, and Dimitris Achlioptas. Tight lower bounds for st-connectivity on the NN-JAG model. SIAM Journal on Computing, 28(6): 2257–2284, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29594.

Eppstein:1990:RSM

[Epp90] David Eppstein. Reset sequences for monotonic automata. SIAM Journal on Computing, 19(3):500–510, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eppstein:1998:FSP

[Epp98] David Eppstein. Finding the k shortest paths. SIAM Journal on Computing, 28(2):652–673, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29047.

Eppstein:2003:SPE

[Epp03] David Eppstein. Setting parameters by example. SIAM Journal on Computing, 32(3): 643–653, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        37008.

Ehrenfeucht:1981:PLR

[EPR81] A. Ehrenfeucht, R. Parikh, and G. Rozenberg. Pumping lemmas for regular sets. SIAM Journal on Computing, 10(3): 536–541, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Epstein:1979:NST

[Eps79] H. I. Epstein. A natural structure theorem for complex fields. SIAM Journal on Computing, 8(3):320–325, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ehrig:1980:MRH

[ER80] Hartmut Ehrig and Barry K. Rosen. The mathematics of record handling. SIAM Journal on Computing, 9(3):441– 469, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Erickson:1999:NLB

[Eri99] Jeff Erickson. New lower bounds for convex hull problems in odd dimensions. SIAM Journal on Computing, 28 (4):1198–1214, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31541.

Erickson:2000:STT

[Eri00]  Jeff Erickson.  Space-time

tradeoffs for emptiness queries. SIAM Journal on Computing, 29(6):1968–1996, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33721.

Eden:2020:ANC

[ERS20] Talya Eden, Dana Ron, and C. Seshadhri. On approximating the number of k-cliques in sublinear time. SIAM Journal on Computing, 49(4):747–771, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ehrenfeucht:1981:ESF

[ERV81] A. Ehrenfeucht, G. Rozenberg, and D. Vermeir. On ETOL systems with finite tree-rank.

SIAM Journal on Computing, 10(1):40–58, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ellis:1981:FSW

[ES81] Martin H. Ellis and J. Michael Steele. Fast sorting of Weyl sequences using comparisons. SIAM Journal on Computing, 10(1):88–95, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Edelsbrunner:1988:PCP

[ES88] Herbert Edelsbrunner and Steven S. Skiena. Probing convex polygons with x-rays. SIAM Journal on Computing, 17(5):870–882, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Edelsbrunner:1991:TAT

[ES91] Herbert Edelsbrunner and Weiping Shi. An O(n log^{2} h) time algorithm for the threedimensional convex hull problem. SIAM Journal on Computing, 20(2):259–269, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ezra:2005:OSC

[ES05] Esther Ezra and Micha Sharir. Output-sensitive construction of the union of triangles. SIAM Journal on Computing, 34(6):1331–1351, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44424.

Elkin:2015:SSL

[ES15] Michael Elkin and Shay Solomon. Steiner shallow-light trees are exponentially lighter than spanning ones. SIAM Journal on Computing, 44(4): 996–1025, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ezra:2022:RST

[ES22] Esther Ezra and Micha Sharir.

On ray shooting for triangles in 3-space and related problems. SIAM Journal on Computing, 51(4):1065–1095, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1408245.

Edelsbrunner:1993:ZTH

[ESS93] Herbert Edelsbrunner, Raimund Seidel, and Micha Sharir. On the zone theorem for hyperplane arrangements. SIAM Journal on Computing, 22(2): 418–429, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ebrahimi:2020:SMN

[ESV20] Javad Ebrahimi, Damian Straszak, and Nisheeth Vishnoi. Subdeterminant maximization via nonconvex relaxations and anti-concentration. SIAM Journal on Computing, 49(6):1249–1270, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eswaran:1975:FRF

[Esw75] Kapali P. Eswaran. Faithful representation of a family of sets by a set of intervals. SIAM Journal on Computing, 4(1):56–68, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Etessami:2017:PTA

[ESY17] Kousha Etessami, Alistair Stewart, and Mihalis Yannakakis. A polynomial time algorithm for computing extinction probabilities of multitype branching processes. SIAM Journal on Computing, 46(5): 1515–1553, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Even:1975:NFT

[ET75] Shimon Even and R. Endre Tarjan. Network flow and testing graph connectivity. SIAM Journal on Computing, 4(4): 507–518, December 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eswaran:1976:AP

[ET76] Kapali P. Eswaran and R. Endre Tarjan. Augmentation problems. SIAM Journal on Computing, 5(4):653–665, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Edelsbrunner:1993:QTA

[ET93] Herbert Edelsbrunner and Tiow Seng Tan. A quadratic time algorithm for the minmax length triangulation. SIAM Journal on Computing, 22(3): 527–551, June 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eppstein:2021:NAC

[EV21] David Eppstein and Vijay V.

Vazirani. NC algorithms for computing a perfect matching and a maximum flow in onecrossing-minor-free graphs. SIAM Journal on Computing, 50(3):1014–1033, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Even:1975:ADW

[Eve75] Shimon Even. An algorithm for determining whether the connectivity of a graph is at least k. SIAM Journal on Computing, 4(3):393–396, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Etessami:2012:SSF

[EvMP^{+}12] Kousha Etessami, Dieter van Melkebeek, Seth Pettie, John Watrous, and Salil Vadhan. Special section on the Forty-Third Annual ACM Symposium on Theory of Computing (STOC 2011). SIAM Journal on Computing, 41(5):1233– 1234, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Epstein:2005:OOA

[EvS05] Leah Epstein and Rob van Stee. Optimal online algorithms for multidimensional packing problems. SIAM Journal on Computing, 35(2):431– 448, ???? 2005. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Edelsbrunner:1986:CBT

[EW86] H. Edelsbrunner and E. Welzl. Constructing belts in twodimensional arrangements with applications. SIAM Journal on Computing, 15(1):271–284, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Eidenbenz:2003:AAM

[EW03] Stephan J. Eidenbenz and Peter Widmayer. An approximation algorithm for minimum convex cover with logarithmic performance guarantee. SIAM Journal on Computing, 32(3):654–670, June 2003. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40513.

Englert:2012:CSP

[EW12] Matthias Englert and Matthias Westermann. Considering suppressed packets improves buffer management in quality of service switches. SIAM Journal on Computing, 41(5): 1166–1192, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Etessami:2005:FSR

[EWS05] Kousha Etessami, Thomas Wilke, and Rebecca A. Schuller. Fair simulation relations, parity games, and state space reduction for Büchi automata. SIAM Journal on Computing, 34(5):1159–1175, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42067.

Etessami:2010:CNE

[EY10] Kousha Etessami and Mihalis Yannakakis. On the complexity of Nash equilibria and other fixed points. SIAM Journal on Computing, 39(6):2531– 2597, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ezra:2016:SSD

[Ezr16] Esther Ezra. A size-sensitive discrepancy bound for set systems of bounded primal shatter dimension. SIAM Journal on Computing, 45(1):84– 101, ???? 2016. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fino:1977:UTD

[FA77] Bernard J. Fino and V. Ralph Algazi. A unified treatment of discrete fast unitary transforms. SIAM Journal on Computing, 6(4):700–717, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fabri:1977:SRW

[Fab77] Janet Fabri. Some remarks on p-way merging. SIAM Journal on Computing, 6(2):268– 271, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fateman:1974:PMP

[Fat74] Richard J. Fateman. Polynomial multiplication, powers, and asymptotic analysis: Some comments. SIAM Journal on Computing, 3(3):196–213, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fernandez-Baca:2015:FPA

[FBGSV15] David Fernández-Baca, Sylvain Guillemot, Brad Shutters, and Sudheer Vakati. Fixedparameter algorithms for finding agreement supertrees. SIAM Journal on Computing, 44(2):384–410, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fernandez-Baca:2003:PTA

[FBL03] David Fernández-Baca and Jens Lagergren. A polynomialtime algorithm for near-perfect phylogeny. SIAM Journal on Computing, 32(5):1115– 1127, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35083.

Federickson:1987:FAS

[Fed87] Greg N. Federickson. Fast algorithms for shortest paths in planar graphs, with applications. SIAM Journal on Computing, 16(6):1004–1022, December 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feige:2006:SIR

[Fei06] Uriel Feige. On sums of independent random variables with unbounded variance and estimating the average degree in a graph. SIAM Journal on Computing, 35(4):964–984, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feige:2009:MWW

[Fei09] Uriel Feige. On maximizing welfare when utility functions are subadditive. SIAM Journal on Computing, 39(1):122– 142, ???? 2009. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feigenbaum:1993:RSR

[FF93] Joan Feigenbaum and Lance Fortnow. Random-self-reducibility of complete sets. SIAM Journal on Computing, 22(5):994– 1005, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Farruggia:2019:BDC

[FFFV19] Andrea Farruggia, Paolo Ferragina, Antonio Frangioni, and Rossano Venturini. Bicriteria data compression. SIAM Journal on Computing, 48(5): 1603–1642, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fenner:1996:ICH

[FFK96] Stephen Fenner, Lance Fortnow, and Stuart A. Kurtz. The isomorphism conjecture holds relative to an oracle. SIAM Journal on Computing, 25(1):193–206, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24830.

Fiat:1998:CAL

[FFK^{+}98] Amos Fiat, Dean P. Foster, Howard Karloff, Yuval Rabani, Yiftach Ravid, and Sundar Vishwanathan. Competitive algorithms for layered graph traversal. SIAM Journal on Computing, 28(2):447–462, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27994.

Feldmann:2018:ELH

[FFKP18] Andreas Emil Feldmann, Wai Shing Fung, Jochen Könemann, and Ian Post. A (1 + ε)-embedding of low highway dimension graphs into bounded treewidth graphs. SIAM Journal on Computing, 47(4):1667–1704, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feige:2020:AMR

[FFTC20] Uriel Feige, Michal Feldman, and Inbal Talgam-Cohen. Approximate modularity revisited. SIAM Journal on Computing, 49(1):67–97, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedman:1982:PTA

[FG82] E. P. Friedman and S. A.

Greibach. A polynomial time algorithm for deciding the equivalence problem for 2tape deterministic finite state acceptors. SIAM Journal on Computing, 11(1):166–183, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1992:PEM

[FG92] Greg N. Frederickson and D. J. Guan. Preemptive ensemble motion planning on a tree. SIAM Journal on Computing, 21(6):1130–1152, December 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ferragina:1998:OLS

[FG98] Paolo Ferragina and Roberto Grossi. Optimal on-line search and sublinear time update in string matching. SIAM Journal on Computing, 27(3):713– 736, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        28611.

Flum:2001:FPT

[FG01] Jörg Flum and Martin Grohe. Fixed-parameter tractability, definability, and modelchecking. SIAM Journal on Computing, 31(1):113–145, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        36076.

Flum:2004:PCC

[FG04] Jörg Flum and Martin Grohe.

The parameterized complexity of counting problems. SIAM Journal on Computing, 33(4):892–922, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42720.

Farshi:2008:ISF

[FGG08] Mohammad Farshi, Panos Giannopoulos, and Joachim Gudmundsson. Improving the stretch factor of a geometric network by edge augmentation. SIAM Journal on Computing, 38(1):226–240, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedman:2005:RMU

[FGK05] Joel Friedman, Andreas Goerdt, and Michael Krivelevich. Recognizing more unsatisfiable random k-SAT instances efficiently. SIAM Journal on Computing, 35(2):408– 430, ???? 2005. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fagin:2008:SID

[FGKO08] Ronald Fagin, Anupam Gupta, Ravi Kumar, and Ryan O’Donnell. Special issue dedicated to the Thirty-Seventh Annual ACM Symposium on Theory of Computing (STOC 2005). SIAM Journal on Computing, 38(2):vii, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2009:ALP

[FGKP09] Vitaly Feldman, Parikshit Gopalan, Subhash Khot, and Ashok Kumar Ponnuswami. On agnostic learning of parities, monomials, and halfspaces. SIAM Journal on Computing, 39(2):606–645, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2016:CWE

[FGL16] Michal Feldman, Nick Gravin, and Brendan Lucier. Combinatorial Walrasian equilibrium. SIAM Journal on Computing, 45(1):29–48, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fortnow:1998:PS

[FGLM98] Lance Fortnow, Judy Goldsmith, Matthew A. Levy, and Stephen Mahaney. L-printable sets. SIAM Journal on Computing, 28(1):137–151, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30044.

Fomin:2010:ICW

[FGLS10] Fedor V. Fomin, Petr A. Golovach, Daniel Lokshtanov, and Saket Saurabh. Intractability of clique-width parameterizations. SIAM Journal on Computing, 39(5):1941–1956, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fomin:2014:AOL

[FGLS14] Fedor V. Fomin, Petr A. Golovach, Daniel Lokshtanov, and Saket Saurabh. Almost optimal lower bounds for problems parameterized by clique-width. SIAM Journal on Computing, 43(5):1541–1563, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frieze:2014:ERS

[FGRV14] Alan Frieze, Navin Goyal, Luis Rademacher, and Santosh Vempala. Expanders via random spanning trees. SIAM Journal on Computing, 43(2): 497–513, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2012:ALM

[FGRW12] Vitaly Feldman, Venkatesan Guruswami, Prasad Raghavendra, and Yi Wu. Agnostic learning of monomials by halfspaces is hard. SIAM Journal on Computing, 41(6):1558– 1590, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Faigle:1985:AAS

[FGS85] Ulrich Faigle, Gerhard Gierz, and Rainer Schrader. Algorithmic approaches to setup minimization. SIAM Journal on Computing, 14(4):954–965, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedrich:2012:QLB

[FGS12] Tobias Friedrich, Martin Gairing, and Thomas Sauerwald. Quasirandom load balancing. SIAM Journal on Computing, 41(4):747–771, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fenner:2021:BPM

[FGT21] Stephen Fenner, Rohit Gurjar, and Thomas Thierauf. Bipartite perfect matching is in quasi-NC. SIAM Journal on Computing, 50(3):STOC16– 218–STOC16–235, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feder:2006:FCS

[FH06] Tomás Feder and Pavol Hell. Full constraint satisfaction problems. SIAM Journal on Computing, 36(1):230–246, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fischer:2011:SEP

[FH11] Johannes Fischer and Volker Heun. Space-efficient preprocessing schemes for range minimum queries on static arrays. SIAM Journal on Computing, 40(2):465–492, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i2/p465_s1.

Formann:1993:DGP

[FHH^{+}93] M. Formann, T. Hagerup, J. Haralambides, M. Kaufmann, F. T. Leighton, A. Symvonis, E. Welzl, and G. Woeginger. Drawing graphs in the plane with high resolution. SIAM Journal on Computing, 22(5):1035–1052, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fung:2019:GFG

[FHHP19] Wai-Shing Fung, Ramesh Hariharan, Nicholas J. A. Harvey, and Debmalya Panigrahi. A general framework for graph sparsification. SIAM Journal on Computing, 48(4):1196– 1223, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1978:AAS

[FHK78] Greg N. Frederickson, Matthew S.

Hecht, and Chul E. Kim. Approximation algorithms for some routing problems. SIAM Journal on Computing, 7(2): 178–193, May 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frieze:1988:RTI

[FHK^{+}88] Alan M. Frieze, Johan Håstad, Ravi Kannan, Jeffrey C. Lagarias, and Adi Shamir. Reconstructing truncated integer variables satisfying linear congruences. SIAM Journal on Computing, 17(2):262–280, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Feige:2002:ADN

[FHKS02] Uriel Feige, Magnús M. Halldórsson, Guy Kortsarz, and Aravind Srinivasan. Approximating the domatic number. SIAM Journal on Computing, 32(1):172–195, February 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38075.

Feige:2008:IAA

[FHL08] Uriel Feige, MohammadTaghi Hajiaghayi, and James R. Lee. Improved approximation algorithms for minimum weight vertex separators. SIAM Journal on Computing, 38(2):629– 657, ???? 2008. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fenner:1997:OCV

[FHOS97] Stephen Fenner, Steven Homer, Mitsunori Ogihara, and Alan Selman. Oracles that compute values. SIAM Journal on Computing, 26(4):1043–1065, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24743.

Friedman:1996:EPC

[FHS96] J. Friedman, J. Hershberger, and J. Snoeyink. Efficiently planning compliant motion in the plane. SIAM Journal on Computing, 25(3):562–599, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26319.

Fiduccia:1985:EFL

[Fid85] Charles M. Fiduccia. An efficient formula for linear recurrences. SIAM Journal on Computing, 14(1):106–112, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Filtser:2019:SPR

[Fil19]  Arnold Filtser.  Steiner

point removal with distortion O(log k) using the relaxed-Voronoi algorithm. SIAM Journal on Computing, 48(2): 249–278, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedl:2014:HTT

[FIM^{+}14] Katalin Friedl, Gábor Ivanyos, Frédéric Magniez, Miklos Santha, and Pranab Sen. Hidden translation and translating coset in quantum computing. SIAM Journal on Computing, 43(1):1–24, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fineman:2020:NWE

[Fin20] Jeremy T. Fineman. Nearly work-efficient parallel algorithm for digraph reachability. SIAM Journal on Computing, 49(5):STOC18–500–STOC18– 539, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fischer:2005:DTI

[Fis05]  Eldar Fischer.  The diffi-

culty of testing for isomorphism against a graph that is given in advance. SIAM Journal on Computing, 34 (5):1147–1158, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44560.

Frederickson:1981:AAS

[FJ81] Greg N. Frederickson and Joseph Ja’Ja’. Approximation algorithms for several graph augmentation problems. SIAM Journal on Computing, 10(2): 270–283, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1984:GSR

[FJ84] Greg N. Frederickson and Donald B. Johnson. Generalized selection and ranking: Sorted matrices. SIAM Journal on Computing, 13(1):14–30, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [FJ90b].

Frederickson:1989:EMR

[FJ89] Greg N. Frederickson and Ravi Janardan. Efficient message routing in planar networks. SIAM Journal on Computing, 18(4):843–857, August 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1990:SEM

[FJ90a] Greg N. Frederickson and Ravi Janardan. Space-efficient message routing in c-decomposable networks. SIAM Journal on Computing, 19(1):164–181, February 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1990:EGS

[FJ90b] Greg N. Frederickson and Donald B. Johnson. Erratum: “Generalized selection and ranking: sorted matrices” [SIAM J. Comput. 13 (1984), no. 1, 14–30, MR 85c:68025]. SIAM Journal on Computing, 19(1):205–206, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [FJ84].

Friesen:1988:AHA

[FK88] D. K. Friesen and F. S. Kuhl. Analysis of a hybrid algorithm for packing unequal bins. SIAM Journal on Computing, 17(1):23–40, February 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Furst:1989:SCA

[FK89] Merrick L. Furst and Ravi Kannan. Succinct certificates for almost all subset sum problems. SIAM Journal on Computing, 18(3):550–558, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Floyd:1990:AM

[FK90] Robert W. Floyd and Donald E. Knuth. Addition machines. SIAM Journal on Computing, 19(2):329–340, April 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feige:2000:TPP

[FK00] Uriel Feige and Joe Kilian. Two-prover protocols — low error at affordable rates. SIAM Journal on Computing, 30 (1):324–346, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32537.

Feige:2002:PAM

[FK02] Uriel Feige and Robert Krauthgamer.

A polylogarithmic approximation of the minimum bisection. SIAM Journal on Computing, 31(4):1090–1118, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38766.

Feige:2003:PVL

[FK03] Uriel Feige and Robert Krauthgamer.

The probable value of the Lovász–Schrijver relaxations for maximum independent set. SIAM Journal on Computing, 32(2):345–370, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40118.

Fischer:2015:OIS

[FK15] Felix Fischer and Max Klimm.

Optimal impartial selection. SIAM Journal on Computing, 44(5):1263–1285, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fleischer:2008:COA

[FKK^{+}08] Rudolf Fleischer, Tom Kamphans, Rolf Klein, Elmar Langetepe, and Gerhard Trippen. Competitive online approximation of the optimal search ratio. SIAM Journal on Computing, 38(3):881–898, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedgut:2011:QVG

[FKKN11] Ehud Friedgut, Gil Kalai, Nathan Keller, and Noam Nisan. A quantitative version of the Gibbard–Satterthwaite theorem for three alternatives. SIAM Journal on Computing, 40(3):934–952, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i3/p934_s1.

Fleischer:2010:SCS

[FKLS10] Lisa Fleischer, Jochen Könemann, Stefano Leonardi, and Guido Schäfer. Strict cost sharing schemes for Steiner forest. SIAM Journal on Computing, 39(8):3616–3632, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feigenbaum:2009:GDD

[FKM^{+}09] Joan Feigenbaum, Sampath Kannan, Andrew McGregor, Siddharth Suri, and Jian Zhang. Graph distances in the data-stream model. SIAM Journal on Computing, 38(5): 1709–1727, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feder:1995:ACC

[FKNN95] Tomás Feder, Eyal Kushilevitz, Moni Naor, and Noam Nisan. Amortized communication complexity. SIAM Journal on Computing, 24(4):736– 750, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23586.

Frieze:1995:WAB

[FKR95] Alan Frieze, Richard M. Karp, and Bruce Reed. When is the assignment bound tight for the asymmetric traveling-salesman problem? SIAM Journal on Computing, 24(3):484–493, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23538.

FernandezdelaVega:1993:TPR

[FKS93] Wenceslas Fernandez de la Vega, Sampath Kannan, and Miklós Sántha. Two probabilistic results on merging. SIAM Journal on Computing, 22(2):261–271, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feigenbaum:2002:ADA

[FKSV02] Joan Feigenbaum, Sampath Kannan, Martin J. Strauss, and Mahesh Viswanathan. An approximate L^{1}-difference algorithm for massive data streams. SIAM Journal on Computing, 32(1):131–151, February 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        36170.

Feige:2017:CGC

[FKT17] Uriel Feige, Tomer Koren, and Moshe Tennenholtz. Chasing ghosts: Competing with stateful policies. SIAM Journal on Computing, 46(1):190–223, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fomin:2008:EAT

[FKTV08] Fedor V. Fomin, Dieter Kratsch, Ioan Todinca, and Yngve Villanger. Exact algorithms for treewidth and minimum fill-in. SIAM Journal on Computing, 38(3):1058–1079, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fenner:1983:TTR

[FL83a] Trevor I. Fenner and Georghios Loizou. Tree traversal related algorithms for generating integer partitions. SIAM Journal on Computing, 12(3):551–564, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fleck:1983:FSA

[FL83b] A. C. Fleck and R. S. Limaye. Formal semantics and abstract properties of string pattern operations and extended formal language description mechanisms. SIAM Journal on Computing, 12(1): 166–188, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friesen:1983:BMS

[FL83c] D. K. Friesen and M. A.

Langston. Bounds for multifit scheduling on uniform processors. SIAM Journal on Computing, 12(1):60–70, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friesen:1986:VSP

[FL86] D. K. Friesen and M. A.

Langston. Variable sized bin packing. SIAM Journal on Computing, 15(1):222–230, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fu:1994:CNH

[FL94] Bin Fu and Hong Zhou Li.

Closeness of NP-Hard sets to other complexity classes. SIAM Journal on Computing, 23(2):255–260, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22593.

Fakcharoenphol:2012:AAV

[FL12] Jittat Fakcharoenphol and Bundit Laekhanukit. An O(log^{2} k)-approximation algorithm for the k-vertex connected spanning subgraph problem. SIAM Journal on Computing, 41(5):1095–1109, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fomin:2022:SPA

[FLM^{+}22] Fedor V. Fomin, Daniel Lokshtanov, Dániel Marx, Marcin Pilipczuk, Michał Pilipczuk, and Saket Saurabh. Subexponential parameterized algorithms for planar and apexminor-free graphs via low treewidth pattern covering. SIAM Journal on Computing, 51(6):1866–1930, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1262504.

Fournier:2015:LBD

[FLMS15] Hervé Fournier, Nutan Limaye, Guillaume Malod, and Srikanth Srinivasan. Lower bounds for depth-4 formulas computing iterated matrix multiplication. SIAM Journal on Computing, 44(5):1173– 1201, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Filmus:2015:SCP

[FLN^{+}15] Yuval Filmus, Massimo Lauria, Jakob Nordström, Noga Ron-Zewi, and Neil Thapen. Space complexity in polynomial calculus. SIAM Journal on Computing, 44(4):1119–1153, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feige:1999:MNZ

[FLS99] Uriel Feige, Dror Lapidot, and Adi Shamir. Multiple noninteractive zero knowledge proofs under general assumptions. SIAM Journal on Computing, 29(1):1–28, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23001.

Feige:2004:GTV

[FLS04] Uriel Feige, Michael Langberg, and Gideon Schechtman. Graphs with tiny vector chromatic numbers and huge chromatic numbers. SIAM Journal on Computing, 33(6): 1338–1368, December 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43139.

Faigle:1986:STS

[FLST86] U. Faigle, L. Lovász, R. Schrader, and Gy. Turán. Searching in trees, series-parallel and interval orders. SIAM Journal on Computing, 15(4):1075–1084, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fomin:2020:BK

[FLST20] Fedor V. Fomin, Daniel Lokshtanov, Saket Saurabh, and Dimitrios M. Thilikos. Bidimensionality and kernels. SIAM Journal on Computing, 49(6):1397–1422, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:1997:OPP

[FM97] Pesech Feldman and Silvio Micali. An optimal probabilistic protocol for synchronous Byzantine agreement. SIAM Journal on Computing, 26 (4):873–933, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/18708.

Fiat:2003:BAU

[FM03] Amos Fiat and Manor Mendel. Better algorithms for unfair metrical task systems and applications. SIAM Journal on Computing, 32(6):1403–1422, December 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        37615.

Fischer:2008:TGI

[FM08] Eldar Fischer and Arie Matsliah. Testing graph isomorphism. SIAM Journal on Computing, 38(1):207–225, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fischer:2010:ASE

[FMdR10] Eldar Fischer, Frédéric Magniez, and Michel de Rougemont. Approximate satisfiability and equivalence. SIAM Journal on Computing, 39(6): 2251–2281, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frieze:2011:ARW

[FMM11] Alan Frieze, Páll Melsted, and Michael Mitzenmacher. An analysis of random-walk cuckoo hashing. SIAM Journal on Computing, 40(2):291–308, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i2/
p291_s1.

Fischer:1982:LBL

[FMP82] Michael J. Fischer, Albert R.

Meyer, and Michael S. Paterson. Ω(n log n) lower bounds on length of Boolean formulas. SIAM Journal on Computing, 11(3):416–427, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fich:1995:PP

[FMP95] Faith E. Fich, J. Ian Munro, and Patricio V. Poblete. Permuting in place. SIAM Journal on Computing, 24(2):266–278, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23864.

Feder:2003:CMU

[FMP^{+}03] Tomás Feder, Rajeev Motwani, Rina Panigrahy, Chris Olston, and Jennifer Widom. Computing the median with uncertainty. SIAM Journal on Computing, 32(2):538–547, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39566.

Frieze:1990:GML

[FMR90] Alan Frieze, Colin McDiarmid, and Bruce Reed. Greedy matching on the line. SIAM Journal on Computing, 19(4): 666–672, August 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fischer:1996:WP

[FMRT96] Michael J. Fischer, Shlomo Moran, Steven Rudich, and Gadi Taubenfeld. The wakeup problem. SIAM Journal on Computing, 25(6):1332–1357, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
25495.

Feder:2002:ALC

[FMS02] Tomás Feder, Rajeev Motwani, and Carlos Subi. Approximating the longest cycle problem in sparse graphs. SIAM Journal on Computing, 31(5):1596–1607, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39548.

Fatourou:2005:EOV

[FMS05] Panagiota Fatourou, Marios Mavronicolas, and Paul Spirakis. Efficiency of oblivious versus nonoblivious schedulers for optimistic, ratebased flow control. SIAM Journal on Computing, 34 (5):1216–1252, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43275.

Fischer:2010:AHP

[FMS10] Eldar Fischer, Arie Matsliah, and Asaf Shapira. Approximate hypergraph partitioning and applications. SIAM Journal on Computing, 39(7):3155– 3185, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fagin:1983:TTD

[FMUY83] Ronald Fagin, David Maier, Jeffrey D. Ullman, and Mihalis Yannakakis. Tools for template dependencies. SIAM Journal on Computing, 12(1):36– 59, ???? 1983. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feige:2011:MNM

[FMV11] Uriel Feige, Vahab S. Mirrokni, and Jan Vondrák. Maximizing non-monotone submodular functions. SIAM Journal on Computing, 40(4):1133–1153, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i4/
                                        p1133_s1.

Feldman:1980:MPT

[FN80] Jerome A. Feldman and Anil Nigam. A model and proof technique for message-based systems. SIAM Journal on Computing, 9(4):768–784, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fiat:1993:IPS

[FN93] Amos Fiat and Moni Naor. Implicit O(1) probe search. SIAM Journal on Computing, 22(1): 1–10, February 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fiat:1999:RTS

[FN99] Amos Fiat and Moni Naor.

Rigorous time/space tradeoffs for inverting functions. SIAM Journal on Computing, 29(3):790–803, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28051.

Fischer:2007:TVE

[FN07] Eldar Fischer and Ilan Newman. Testing versus estimation of graph properties. SIAM Journal on Computing, 37(2): 482–501, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fukunaga:2015:IRA

[FNR15] Takuro Fukunaga, Zeev Nutov, and R. Ravi. Iterative rounding approximation algorithms for degree-bounded node-connectivity network design. SIAM Journal on Computing, 44(5):1202–1229, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ferragina:2013:BCL

[FNV13] Paolo Ferragina, Igor Nitto, and Rossano Venturini. On the bit-complexity of Lempel– Ziv compression. SIAM Journal on Computing, 42(4):1521– 1541, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fortune:1979:NSC

[For79] Steven Fortune. A note on sparse complete sets. SIAM Journal on Computing, 8(3): 431–433, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2008:LMP

[FOS08] Jon Feldman, Ryan O’Donnell, and Rocco A. Servedio. Learning mixtures of product distributions over discrete domains. SIAM Journal on Computing, 37(5):1536–1564, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fagin:1978:ECE

[FP78] Ronald Fagin and Thomas G. Price. Efficient calculation of expected miss ratios in the independent reference model. SIAM Journal on Computing, 7(3):288–297, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Flajolet:1986:RAU

[FP86] P. Flajolet and H. Prodinger. Register allocation for unary– binary trees. SIAM Journal on Computing, 15(3):629–640, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ferns:2011:BMC

[FPP11] Norm Ferns, Prakash Panangaden, and Doina Precup.

Bisimulation metrics for continuous Markov decision processes. SIAM Journal on Computing, 40(6):1662–1714, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i6/
p1662_s1.

Fountoulakis:2013:ITC

[FPS13] Nikolaos Fountoulakis, Konstantinos Panagiotou, and Angelika Steger. On the insertion time of cuckoo hashing. SIAM Journal on Computing, 42(6): 2156–2181, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fox:2016:PRL

[FPS16] Jacob Fox, János Pach, and Andrew Suk. A polynomial regularity lemma for semialgebraic hypergraphs and its applications in geometry and property testing. SIAM Journal on Computing, 45(6):2199– 2223, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2018:CRS

[FPV18] Vitaly Feldman, Will Perkins, and Santosh Vempala. On the complexity of random satisfiability problems with planted solutions. SIAM Journal on Computing, 47(4):1294–1338, ???? 2018. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Ferrante:1975:DPF

[FR75] Jeanne Ferrante and Charles Rackoff. A decision procedure for the first order theory of real addition with order. SIAM Journal on Computing, 4(1):69–76, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Flajolet:1980:NGC

[FR80] P. Flajolet and Lyle Ramshaw. A note on Gray code and oddeven merge. SIAM Journal on Computing, 9(1):142–158, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1994:NAS

[FR94] Greg N. Frederickson and Susan H. Rodger. An NC algorithm for scheduling unittime jobs with arbitrary release times and deadlines. SIAM Journal on Computing, 23(1):185–211, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/18671.

Feldman:2006:DSN

[FR06] Jon Feldman and Matthias Ruhl. The directed Steiner network problem is tractable for a constant number of terminals. SIAM Journal on Computing, 36(2):543–561, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Focke:2024:CSI

[FR24] Jacob Focke and Marc Roth.

Counting small induced subgraphs with hereditary properties. SIAM Journal on Computing, 53(2):189–220, March 2024. CODEN SMJCAT. ISSN 1095-7111.

Franco:1991:EIV

[Fra91] John Franco. Elimination of infrequent variables improves average case performance of satisfiability algorithms. SIAM Journal on Computing, 20(6): 1119–1127, December 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fredman:1976:NBC

[Fre76] Michael L. Fredman. New bounds on the complexity of the shortest path problem. SIAM Journal on Computing, 5(1):83–89, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fredman:1978:OCG

[Fre78] Michael L. Fredman. Observations on the complexity of generating quasi-gray codes. SIAM Journal on Computing, 7(2):134–146, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fredman:1981:LBC

[Fre81] Michael L. Fredman. Lower bounds on the complexity of some optimal data structures. SIAM Journal on Computing, 10(1):1–10, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1984:SOH

[Fre84] Greg N. Frederickson. Selforganizing heuristics for implicit data structures. SIAM Journal on Computing, 13(2): 277–291, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1985:DSL

[Fre85] Greg N. Frederickson. Data structures for on-line updating of minimum spanning trees, with applications. SIAM Journal on Computing, 14(4):781– 798, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1993:NCS

[Fre93] Greg N. Frederickson. A note on the complexity of a simple transportation problem. SIAM Journal on Computing, 22(1): 57–61, February 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1997:ADS

[Fre97] Greg N. Frederickson. Ambivalent data structures for dynamic 2-edge-connectivity and k smallest spanning trees. SIAM Journal on Computing, 26(2):484–538, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22682.

Filos-Ratsikas:2023:CEC

[FRGH^{+}23] Aris Filos-Ratsikas, Yiannis Giannakopoulos, Alexandros Hollender, Philip Lazos, and Diogo Poças. On the complexity of equilibrium computation in first-price auctions. SIAM Journal on Computing, 52(1):80–131, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1435823.

Friesen:1984:TBM

[Fri84] Donald K. Friesen. Tighter bounds for the multifit processor scheduling algorithm. SIAM Journal on Computing, 13(1):170–181, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedman:1986:CSM

[Fri86a] Joel Friedman. Constructing O(n log n) size monotone formulae for the k th threshold function of n Boolean variables. SIAM Journal on Computing, 15(3):641–654, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frieze:1986:LOA

[Fri86b] A. M. Frieze. On the Lagarias– Odlyzko algorithm for the subset sum problem. SIAM Journal on Computing, 15(2):536– 539, ???? 1986. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friesen:1987:TBL

[Fri87a] Donald K. Friesen. Tighter bounds for LPT scheduling on uniform processors. SIAM Journal on Computing, 16(3): 554–560, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frieze:1987:ESR

[Fri87b] A. M. Frieze. On the exact solution of random travelling salesman problems with medium size integer coefficients. SIAM Journal on Computing, 16(6):1052–1072, December 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedman:1990:DTP

[Fri90a] Joel Friedman. A density theorem for purely iterative zero finding methods. SIAM Journal on Computing, 19(1):124– 132, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedman:1990:RPA

[Fri90b] Joel Friedman. Random polynomials and approximate zeros of Newton’s method. SIAM Journal on Computing, 19(6): 1068–1099, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedman:1991:SIH

[Fri91] Joel Friedman. The spectra of infinite hypertrees. SIAM Journal on Computing, 20(5): 951–961, October 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friedman:1993:NPG

[Fri93] Joel Friedman. A note on poset geometries. SIAM Journal on Computing, 22(1):72– 78, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frieze:2002:CED

[Fri02] Alan Frieze. Corrigendum:

Edge-disjoint paths in expander graphs. SIAM Journal on Computing, 31(3):988, June 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39509. See [?].

Feige:1994:CNI

[FRPU94] Uriel Feige, Prabhakar Raghavan, David Peleg, and Eli Upfal. Computing with noisy information. SIAM Journal on Computing, 23(5):1001– 1018, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
19587.

Faust:2014:PCC

[FRR^{+}14] Sebastian Faust, Tal Rabin, Leonid Reyzin, Eran Tromer, and Vinod Vaikuntanathan. Protecting circuits from computationally bounded and noisy leakage. SIAM Journal on Computing, 43(5):1564– 1614, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Flocchini:2003:BCS

[FRS03] P. Flocchini, A. Roncato, and N. Santoro. Backward consistency and sense of direction in advanced distributed systems. SIAM Journal on Computing, 32(2):281–306, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37729.

Friggstad:2019:LSY

[FRS19] Zachary Friggstad, Mohsen Rezapour, and Mohammad R. Salavatipour. Local search yields a PTAS for k-means in doubling metrics. SIAM Journal on Computing, 48(2):452– 480, ???? 2019. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fox:2020:FCS

[FRS^{+}20] Jacob Fox, Tim Roughgarden, C. Seshadhri, Fan Wei, and Nicole Wein. Finding cliques in social networks: a new distribution-free model. SIAM Journal on Computing, 49(2): 448–464, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fussell:1993:FTC

[FRT93] Donald Fussell, Vijaya Ramachandran, and Ramakrishna Thurimella. Finding triconnected components by local replacement. SIAM Journal on Computing, 22(3):587– 616, June 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fischer:2010:FCW

[FRV10] Simon Fischer, Harald Räcke, and Berthold Vöcking. Fast convergence to Wardrop equilibria by adaptive sampling methods. SIAM Journal on Computing, 39(8):3700–3735, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v39/i8/
p3700_s1.

Fich:1988:RBC

[FRW88] Faith E. Fich, Prabhakar Ragde, and Avi Wigderson. Relations between concurrentwrite models of parallel computation. SIAM Journal on Computing, 17(3):606–627, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Flon:1981:TCP

[FS81] Lawrence Flon and Norihisa Suzuki. The total correctness of parallel programs. SIAM Journal on Computing, 10(2): 227–246, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feigenbaum:1986:RCG

[FS86a] Joan Feigenbaum and Alejandro A. Schäffer. Recognizing composite graphs is equivalent to testing graph isomorphism. SIAM Journal on Computing, 15(2):619–627, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Flajolet:1986:DST

[FS86b] Philippe Flajolet and Robert Sedgewick. Digital search trees revisited. SIAM Journal on Computing, 15(3):748–767, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frederickson:1989:ADS

[FS89] Greg N. Frederickson and Mandayam A. Srinivas. Algorithms and data structures for an expanded family of matroid intersection problems. SIAM Journal on Computing, 18(1):

112–138, February 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fischer:1992:LRS

[FS92] Paul Fischer and Hans Ulrich Simon. On learning ring-sum-expansions. SIAM Journal on Computing, 21(1):181–192, February 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fleischer:2007:QFT

[FS07a] Lisa Fleischer and Martin Skutella. Quickest flows over time. SIAM Journal on Computing, 36(6):1600–1630, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Frieze:2007:PRB

[FS07b] Alan Frieze and Gregory B.

Sorkin. The probabilistic relationship between the assignment and asymmetric traveling salesman problems. SIAM Journal on Computing, 36(5): 1435–1452, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Filtser:2020:GSE

[FS20] Arnold Filtser and Shay Solomon. The greedy spanner is existentially optimal. SIAM Journal on Computing, 49(2): 429–447, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Friggstad:2013:ATS

[FSS13] Zachary Friggstad, Mohammad R. Salavatipour, and Zoya Svitkina. Asymmetric traveling salesman path and directed latency problems. SIAM Journal on Computing, 42(4): 1596–1619, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2020:TBD

[FSS20] Dan Feldman, Melanie Schmidt, and Christian Sohler. Turning big data into tiny data: Constant-size coresets for kmeans, PCA, and projective clustering. SIAM Journal on Computing, 49(3):601–657, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fogelman-Soulie:1987:RIT

[FSW87] Françoise Fogelman-Soulié and Gérard Weisbuch. Random iterations of threshold networks and associative memory. SIAM Journal on Computing, 16(1): 203–220, February 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2021:OCR

[FSZ21] Moran Feldman, Ola Svensson, and Rico Zenklusen. Online contention resolution schemes with applications to Bayesian selection problems. SIAM Journal on Computing, 50(2): 255–300, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2022:FSP

[FSZ22] Moran Feldman, Ola Svensson, and Rico Zenklusen. A framework for the secretary problem on the intersection of matroids. SIAM Journal on Computing, 51(3):766–819, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1411822.

Fischer:1983:WSM

[FT83] Patrick C. Fischer and Don Min Tsou. Whether a set of multivalued dependencies implies a join dependency is NP-hard. SIAM Journal on Computing, 12(2):259–266, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Faigle:1988:SRP

[FT88] U. Faigle and Gy. Turán. Sorting and recognition problems for ordered sets. SIAM Journal on Computing, 17(1):100–113, February 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Farach:1997:SDP

[FT97] Martin Farach and Mikkel Thorup. Sparse dynamic programming for evolutionarytree comparison. SIAM Journal on Computing, 26(1):210– 230, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26242.

Fomin:2006:DSP

[FT06] Fedor V. Fomin and Dimitrios M. Thilikos. Dominating sets in planar graphs: Branchwidth and exponential speedup. SIAM Journal on Computing, 36(2):281–309, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fomin:2015:LIS

[FTV15] Fedor V. Fomin, Ioan Todinca, and Yngve Villanger. Large induced subgraphs via triangulations and CMSO. SIAM Journal on Computing, 44(1):54– 87, ???? 2015. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fu:1995:QQE

[Fu95] Bin Fu. With quasilinear queries EXP is not polynomial time Turing reducible to sparse sets. SIAM Journal on Computing, 24(5):1082–1090, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23718.

Furer:2009:FIM

[Für09] Martin Fürer. Faster integer multiplication. SIAM Journal on Computing, 39(3):979– 1005, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fushimi:1988:DUR

[Fus88] Masanori Fushimi. Designing a uniform random number generator whose subsequences are k-distributed. SIAM Journal on Computing, 17(1):89– 99, February 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feder:1998:CSM

[FV98] Tomás Feder and Moshe Y.

Vardi. The computational structure of monotone monadic SNP and constraint satisfaction: a study through datalog and group theory. SIAM Journal on Computing, 28 (1):57–104, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26676.

Fomin:2013:SPA

[FV13] Fedor V. Fomin and Yngve Villanger. Subexponential parameterized algorithm for minimum fill-in. SIAM Journal on Computing, 42(6):2197– 2216, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2016:OBA

[FV16] Vitaly Feldman and Jan Vondrák. Optimal bounds on approximation of submodular and XOS functions by juntas. SIAM Journal on Computing, 45(3):1129–1170, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feng:2021:DSG

[FVY21] Weiming Feng, Nisheeth K. Vishnoi, and Yitong Yin. Dynamic sampling from graphical models. SIAM Journal on Computing, 50(2):350–381, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fillmore:1974:BCD

[FW74] Jay P. Fillmore and S. G. Williamson. On backtracking: a combinatorial description of the algorithm. SIAM Journal on Computing, 3(1):41–55, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fillmore:1976:RAS

[FW76] Jay P. Fillmore and S. G. Williamson. Ranking algorithms: The symmetries and colorations of the n-cube. SIAM Journal on Computing, 5(2):297–304, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fich:1990:TUE

[FW90] Faith E. Fich and Avi Wigderson. Toward understanding exclusive read. SIAM Journal on Computing, 19(4):718–727, August 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Felsner:1998:MCP

[FW98] Stefan Felsner and Lorenz Wernisch. Maximum k-chains in planar point sets: Combinatorial structure and algorithms. SIAM Journal on Computing, 28(1):192–209, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26617.

Fu:2007:GST

[FW07] Bin Fu and Wei Wang. Geometric separators and their applications to protein folding in the HP-model. SIAM Journal on Computing, 37(4): 1014–1029, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Filmus:2014:MSM

[FW14] Yuval Filmus and Justin Ward.

Monotone submodular maximization over a matroid via non-oblivious local search. SIAM Journal on Computing, 43(2):514–542, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2015:SCB

[FX15] Vitaly Feldman and David Xiao. Sample complexity bounds on differentially private learning via communication complexity. SIAM Journal on Computing, 44(6):1740– 1764, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Fujita:1996:OGG

[FY96] Satoshi Fujita and Masafumi Yamashita. Optimal group gossiping in hypercubes under a circuit- switching model. SIAM Journal on Computing, 25(5):1045–1060, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25287.

Fujinaga:2015:PFO

[FYO^{+}15] Nao Fujinaga, Yukiko Yamauchi, Hirotaka Ono, Shuji Kijima, and Masafumi Yamashita. Pattern formation by oblivious asynchronous mobile robots. SIAM Journal on Computing, 44(3):740–785, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Feldman:2018:SSP

[FZ18] Moran Feldman and Rico Zenklusen. The submodular secretary problem goes linear.

SIAM Journal on Computing, 47(2):330–366, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gabow:1977:TAG

[Gab77] Harold N. Gabow. Two algorithms for generating weighted spanning trees in order. SIAM Journal on Computing, 6(1): 139–150, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gabow:1988:SUS

[Gab88] Harold N. Gabow. Scheduling UET systems on two uniform processors and length two pipelines. SIAM Journal on Computing, 17(4):810–829, August 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gabow:2007:FPC

[Gab07] Harold N. Gabow. Finding paths and cycles of superpolylogarithmic length. SIAM Journal on Computing, 36(6): 1648–1671, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gil:1996:TMH

[GadHW96] Joseph Gil, Friedhelm Meyer auf der Heide, and Avi Wigderson. The tree model for hashing: Lower and upper bounds. SIAM Journal on Computing, 25(5):936–955, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25572.

Galil:1977:RCB

[Gal77] Zvi Galil. On resolution with clauses of bounded size. SIAM Journal on Computing, 6(3): 444–459, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1980:FVC

[Gal80] Zvi Galil. Finding the vertex connectivity of graphs. SIAM Journal on Computing, 9(1): 197–199, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gallier:1984:RAB

[Gal84a] Jean H. Gallier. n-rational algebras. I. basic properties and free algebras. SIAM Journal on Computing, 13(4):750–775, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gallier:1984:RAIb

[Gal84b] Jean H. Gallier. n-rational algebras. II. varieties and logic of inequalities. SIAM Journal on Computing, 13(4):776–794, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1988:P

[Gal88] Zvi Galil. Preface. SIAM Journal on Computing, 17(2):v–vii, ???? 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gamarnik:2003:SAN

[Gam03]  David Gamarnik.  Stabil-

ity of adaptive and nonadaptive packet routing policies in adversarial queueing networks. SIAM Journal on Computing, 32(2):371–385, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36916.

Ganz:1995:EPU

[Gan95] Jürg Ganz. Evaluation of polynomials using the structure of the coefficients. SIAM Journal on Computing, 24(3):473–483, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22558.

Gonzalez-Aguilar:2021:MRC

[GAOPL^{+}21] Hernán González-Aguilar, David Orden, Pablo Pérez-Lantero, David Rappaport, Carlos Seara, Javier Tejel, and Jorge Urrutia. Maximum rectilinear convex subsets. SIAM Journal on Computing, 50(1):145–170, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garey:1974:OBS

[Gar74] M. R. Garey. Optimal binary search trees with restricted maximal depth. SIAM Journal on Computing, 3(2):101– 110, ???? 1974. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gartner:1995:SAA

[Gär95] Bernd Gärtner. A subexponential algorithm for abstract optimization problems. SIAM Journal on Computing, 24(5):1018–1035, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25028.

Gasarch:1987:ODV

[Gas87] William Gasarch. Oracles for deterministic versus alternating classes. SIAM Journal on Computing, 16(4):613–627, August 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gatterdam:1992:ASS

[Gat92] R. W. Gatterdam. Algorithms for splicing systems. SIAM Journal on Computing, 21(3): 507–520, June 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gaver:1972:ASS

[Gav72a] Donald P. Gaver. Asymptotic service system output, with application to multiprogramming. SIAM Journal on Computing, 1(2):138–145, ???? 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gavril:1972:AMC

[Gav72b] Fănică Gavril. Algorithms for minimum coloring, maximum clique, minimum covering by cliques, and maximum independent set of a chordal graph. SIAM Journal on Computing, 1(2):180–187, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gazit:1991:NUB

[Gaz91a]  Hillel Gazit.  New upper

bounds in Klee’s measure problem. SIAM Journal on Computing, 20(6):1034–1045, December 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gazit:1991:ORP

[Gaz91b] Hillel Gazit. An optimal randomized parallel algorithm for finding connected components in a graph. SIAM Journal on Computing, 20(6):1046– 1067, December 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gouyou-Beauchamps:1982:HCP

[GB82] D. Gouyou-Beauchamps. The Hamiltonian circuit problem is polynomial for 4-connected planar graphs. SIAM Journal on Computing, 11(3):529–539, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Geffert:1991:NCS

[Gef91] Viliam Geffert. Nondeterministic, computations in sublogarithmic space and space constructibility. SIAM Journal on Computing, 20(3):484–498, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Geffert:1993:TVS

[Gef93] Viliam Geffert. Tally versions of the Savitch and Immerman–Szelepcsényi theorems for sublogarithmic space. SIAM Journal on Computing, 22(1):102–113, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Geller:1975:RFEa

[Gel75a] Dennis P. Geller. Realization with feedback encoding. I: Analogues of the classical theory. SIAM Journal on Computing, 4(1):12–33, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Geller:1975:RFEb

[Gel75b] Dennis P. Geller. Realization with feedback encoding. II: Applications to distinguishing sequences. SIAM Journal on Computing, 4(1):34–48, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gerhart:1976:PTP

[Ger76] Susan L. Gerhart. Proof theory of partial correctness verification systems. SIAM Journal on Computing, 5(3):355– 377, ???? 1976. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Geser:2002:DTG

[Ges02] Alfons Geser. Decidability of termination of grid string rewriting rules. SIAM Journal on Computing, 31(4):1156– 1168, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
33297.

Garey:1975:BMS

[GG75] M. R. Garey and R. L. Graham. Bounds for multiprocessor scheduling with resource constraints. SIAM Journal on Computing, 4(2):187–200, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Geske:1986:RUR

[GG86] John Geske and Joachim Grollmann. Relativizations of unambiguous and random polynomial time classes. SIAM Journal on Computing, 15(2): 511–519, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1991:ECS

[GG91] Zvi Galil and Raffaele Giancarlo. On the exact complexity of string matching: Lower bounds. SIAM Journal on Computing, 20(6):1008–1020, December 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1992:ECS

[GG92] Zvi Galil and Raffaele Giancarlo. On the exact complexity of string matching: Upper bounds. SIAM Journal on Computing, 21(3):407–437, June 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gal:2010:LBS

[GG10] Anna Gál and Parikshit Gopalan. Lower bounds on streaming algorithms for approximating the length of the longest increasing subsequence. SIAM Journal on Computing, 39(8):3463–3479, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gabow:2012:IRA

[GG12] Harold N. Gabow and Suzanne R. Gallagher. Iterated rounding algorithms for the smallest k-edge connected spanning subgraph. SIAM Journal on Computing, 41(1):61–103, ???? 2012. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v41/i1/
p61_s1.

Gandikota:2018:NHR

[GGG18] Venkata Gandikota, Badih Ghazi, and Elena Grigorescu. NP-hardness of Reed–Solomon decoding, and the Prouhet– Tarry–Escott problem. SIAM Journal on Computing, 47(4): 1547–1584, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galanis:2021:CSR

[GGGY21] Andreas Galanis, Leslie Ann Goldberg, Heng Guo, and Kuan Yang. Counting solutions to random CNF formulas. SIAM Journal on Computing, 50(6):1701–1738, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garg:2016:CIO

[GGH^{+}16] Sanjam Garg, Craig Gentry, Shai Halevi, Mariana Raykova, Amit Sahai, and Brent Waters. Candidate indistinguishability obfuscation and functional encryption for all circuits. SIAM Journal on Computing, 45(3): 882–929, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galanis:2016:ACC

[GGJ16] Andreas Galanis, Leslie Ann Goldberg, and Mark Jerrum. Approximately counting Hcolorings is #BIS-hard. SIAM Journal on Computing, 45(3): 680–711, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldberg:2010:CDP

[GGJT10] Leslie Ann Goldberg, Martin Grohe, Mark Jerrum, and Marc Thurley. A complexity dichotomy for partition functions with mixed signs. SIAM Journal on Computing, 39(7): 3336–3402, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gereb-Graus:1992:ACP

[GGK92] Mihály Geréb-Graus and Danny Krizanc. The average complexity of parallel comparison merging. SIAM Journal on Computing, 21(1):43–47, February 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gu:2016:PDM

[GGK16] Albert Gu, Anupam Gupta, and Amit Kumar. The power of deferral: Maintaining a constant-competitive Steiner tree online. SIAM Journal on Computing, 45(1):1–28, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gennaro:2005:BEG

[GGKT05] Rosario Gennaro, Yael Gertner, Jonathan Katz, and Luca Trevisan. Bounds on the efficiency of generic cryptographic constructions. SIAM Journal on Computing, 35(1):217–246, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldreich:1998:FTC

[GGL98] Oded Goldreich, Shafi Goldwasser, and Nathan Linial. Fault-tolerant computation in the full information model. SIAM Journal on Computing, 27(2):506–544, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24668.

Grandoni:2013:SCO

[GGL^{+}13] Fabrizio Grandoni, Anupam Gupta, Stefano Leonardi, Pauli Miettinen, Piotr Sankowski, and Mohit Singh. Set covering with our eyes closed. SIAM Journal on Computing, 42(3): 808–830, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gimenez:2011:DRS

[GGM11] Omer Giménez, Guillem Godoy, and Sebastian Maneth. Deciding regularity of the set of instances of a set of terms with regular constraints is EXPTIME-complete. SIAM Journal on Computing, 40(2): 446–464, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i2/p446_s1.

Gelade:2012:REC

[GGM12] Wouter Gelade, Marc Gyssens, and Wim Martens. Regular expressions with counting: Weak versus strong determinism. SIAM Journal on Computing, 41(1):160–190, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v41/i1/p160_s1.

Gobel:2015:CLM

[GGM^{+}15] Andreas Göbel, Leslie Ann Goldberg, Colin McQuillan, David Richerby, and Tomoyuki Yamakami. Counting list matrix partitions of graphs. SIAM Journal on Computing, 44(4): 1089–1118, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldreich:2010:IHR

[GGN10] Oded Goldreich, Shafi Goldwasser, and Asaf Nussboim. On the implementation of huge random objects. SIAM Journal on Computing, 39(7):2761– 2822, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gafni:2011:CED

[GGP11] Eli Gafni, Rachid Guerraoui, and Bastian Pochon. The complexity of early deciding set agreement. SIAM Journal on Computing, 40(1):63–78, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i1/
p63_s1.

Gopalan:2011:LDT

[GGR11] Parikshit Gopalan, Venkatesan Guruswami, and Prasad Raghavendra. List decoding tensor products and interleaved codes. SIAM Journal on Computing, 40(5):1432–1462, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i5/
p1432_s1.

Gallo:1989:FPM

[GGT89] Giorgio Gallo, Michael D.

Grigoriadis, and Robert E. Tarjan. A fast parametric maximum flow algorithm and applications. SIAM Journal on Computing, 18(1):30– 55, February 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goodman:1973:EWG

[GH73] S. Goodman and S. Hedetniemi. Eulerian walks in graphs. SIAM Journal on Computing, 2(1):16–27, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goodman:1974:HWG

[GH74] S. E. Goodman and S. T. Hedetniemi. On Hamiltonian walks in graphs. SIAM Journal on Computing, 3(3):214– 221, ???? 1974. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ganesan:1992:CPS

[GH92] K. Ganesan and Steven Homer. Complete problems and strong polynomial reducibilities. SIAM Journal on Computing, 21(4):733–742, August 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ge:1997:DRS

[GH97] Zhengyu Ge and S. Louis Hakimi. Disjoint rooted spanning trees with small depths in deBruijn and Kautz graphs. SIAM Journal on Computing, 26(1):79–92, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24419.

Goldmann:1998:MCC

[GH98] Mikael Goldmann and Johan Håstad. Monotone circuits for connectivity have depth (log n)^{2−o(1)}. SIAM Journal on Computing, 27 (5):1283–1294, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28563.

Gelles:2017:CIC

[GH17] Ran Gelles and Bernhard Haeupler. Capacity of interactive communication over erasure channels and channels with feedback. SIAM Journal on Computing, 46(4):1449– 1472, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guruswami:2017:SPH

[GHH^{+}17] Venkatesan Guruswami, Prahladh Harsha, Johan Håstad, Srikanth Srinivasan, and Girish Varma. Super-polylogarithmic hypergraph coloring hardness via low-degree long codes. SIAM Journal on Computing, 46(1): 132–159, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goos:2024:FCB

[GHJ^{+}24] Mika Göös, Alexandros Hollender, Siddhartha Jain, Gilbert Maystre, William Pires, Robert Robere, and Ran Tao. Further collapses in tfnp. SIAM Journal on Computing, 53(3): 573–587, May 2024. CODEN SMJCAT. ISSN 1095-7111.

Goldsmith:1991:NTS

[GHJY91] Judy Goldsmith, Lane A. Hemachandra, Deborah Joseph, and Paul Young. Near-testable sets. SIAM Journal on Computing, 20(3):506–523, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gritzmann:1991:GBR

[GHK91] Peter Gritzmann, Laurent Habsieger, and Victor Klee. Good and bad radii of convex polygons. SIAM Journal on Computing, 20(2):395–403, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grove:1996:ACP

[GHK96] Adam J. Grove, Joseph Y. Halpern, and Daphne Koller. Asymptotic conditional probabilities: The unary case. SIAM Journal on Computing, 25(1):1–51, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25703.

Guerraoui:2012:WFD

[GHKT12] Rachid Guerraoui, Vassos Hadzilacos, Petr Kuznetsov, and Sam Toueg. The weakest failure detectors to solve quittable consensus and nonblocking atomic commit. SIAM Journal on Computing, 41(6): 1343–1379, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guruswami:2011:BRO

[GHM^{+}11] Venkatesan Guruswami, Johan Håstad, Rajsekar Manokaran, Prasad Raghavendra, and Moses Charikar. Beating the random ordering is hard: Every ordering CSP is approximation resistant. SIAM Journal on Computing, 40(3): 878–914, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i3/p878_s1.

Gupta:2013:PRC

[GHRU13] Anupam Gupta, Moritz Hardt, Aaron Roth, and Jonathan Ullman. Privately releasing conjunctions and the statistical query barrier. SIAM Journal on Computing, 42(4):1494– 1520, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guruswami:2002:HAH

[GHS02] Venkatesan Guruswami, Johan Håstad, and Madhu Sudan. Hardness of approximate hypergraph coloring. SIAM Journal on Computing, 31(6):1663–1686, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37716.

Gil:2008:EAD

[GHSZ08] Ángel J. Gil, Miki Hermann, Gernot Salzer, and Bruno Zanuttini. Efficient algorithms for description problems over finite totally ordered domains. SIAM Journal on Computing, 38(3):922–945, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goodman:1976:MT

[GHT76] S. Goodman, S. Hedetniemi, and R. E. Tarjan. b-matchings in trees. SIAM Journal on Computing, 5(1):104–108, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Golab:2010:RLE

[GHW10] Wojciech Golab, Danny Hendler, and Philipp Woelfel. An O(1) RMRs leader election algorithm. SIAM Journal on Computing, 39(7):2726–2760, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1989:MKI

[GHY89] Zvi Galil, Stuart Haber, and Moti Yung. Minimumknowledge interactive proofs for decision problems. SIAM Journal on Computing, 18(4): 711–739, August 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gurari:1980:PSC

[GI80] Eitan M. Gurari and Oscar H.

Ibarra. Path systems: constructions, solutions and applications. SIAM Journal on Computing, 9(2):348–374, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1992:FDA

[GI92] Zvi Galil and Giuseppe F. Italiano. Fully dynamic algorithms for 2-edge connectivity. SIAM Journal on Computing, 21(6):1047–1069, December 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1993:MEC

[GI93] Zvi Galil and Giuseppe F.

Italiano. Maintaining the 3edge-connected components of a graph on-line. SIAM Journal on Computing, 22(1):11– 28, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Giancarlo:1995:GST

[Gia95] Raffaele Giancarlo. A generalization of the suffix tree to square matrices, with applications. SIAM Journal on Computing, 24(3):520–562, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23198.

Giesbrecht:1995:NOA

[Gie95] Mark Giesbrecht. Nearly optimal algorithms for canonical matrix forms. SIAM Journal on Computing, 24 (5):948–969, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25268.

Gill:1977:CCP

[Gil77] John Gill. Computational complexity of probabilistic Turing machines. SIAM Journal on Computing, 6(4):675– 695, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gillman:1998:CBR

[Gil98]  David Gillman.  A Cher-

noff bound for random walks on expander graphs. SIAM Journal on Computing, 27 (4):1203–1220, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26876.

Georgiadis:2020:SCD

[GIP20] Loukas Georgiadis, Giuseppe F. Italiano, and Nikos Parotsidis. Strong connectivity in directed graphs under failures, with applications. SIAM Journal on Computing, 49(5):865–926, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gonzalez:1977:BLS

[GIS77] Teofilo Gonzalez, Oscar H.

Ibarra, and Sartaj Sahni. Bounds for LPT schedules on uniform processors. SIAM Journal on Computing, 6(1): 155–166, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garey:1975:CRM

[GJ75] M. R. Garey and D. S. Johnson. Complexity results for multiprocessor scheduling under resource constraints. SIAM Journal on Computing, 4(4): 397–411, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garey:1977:TPS

[GJ77] M. R. Garey and D. S. Johnson. Two-processor scheduling with start-times and deadlines*. SIAM Journal on Computing, 6(3):416–426, September 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gonzalez:1982:EAE

[GJ82] Teofilo Gonzalez and Joseph Ja’Ja’. Evaluation of arithmetic expressions with algebraic identities. SIAM Journal on Computing, 11(4):633–662, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garey:1985:CFM

[GJ85] M. R. Garey and D. S. Johnson. Composing functions to minimize image size. SIAM Journal on Computing, 14(2): 500–503, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gaver:1986:PST

[GJ86] Donald P. Gaver and Patricia A. Jacobs. Processorshared time-sharing models in heavy traffic. SIAM Journal on Computing, 15(4):1085– 1100, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldberg:1999:RSM

[GJ99] Leslie Ann Goldberg and Mark Jerrum. Randomly sampling molecules. SIAM Journal on Computing, 29(3):834–853, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31886.

Gabow:2000:HMS

[GJ00] Harold N. Gabow and Tibor Jordán. How to make a square grid framework with cables rigid. SIAM Journal on Computing, 30(2):649–680, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34718.

Goldberg:2013:PTA

[GJ13] Leslie Ann Goldberg and Mark Jerrum. A polynomial-time algorithm for estimating the partition function of the ferromagnetic Ising model on a regular matroid. SIAM Journal on Computing, 42(3):1132– 1157, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldberg:2014:CCS

[GJ14] Leslie Ann Goldberg and Mark Jerrum. The complexity of computing the sign of the Tutte polynomial. SIAM Journal on Computing, 43(6):1921– 1952, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guo:2019:PTA

[GJ19] Heng Guo and Mark Jerrum.

A polynomial-time approximation algorithm for all-terminal network reliability. SIAM Journal on Computing, 48(3): 964–978, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gillibert:2022:WSE

[GJK^{+}22] Pierre Gillibert, Julius Jonusas, Michael Kompatscher, Antoine Mottet, and Michael Pinsker. When symmetries are not enough: a hierarchy of hard constraint satisfaction problems. SIAM Journal on Computing, 51(2):175–213, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1383471.

Goldberg:2004:BCB

[GJKP04] Leslie Ann Goldberg, Mark Jerrum, Sampath Kannan, and Mike Paterson. A bound on the capacity of backoff and acknowledgment-based protocols. SIAM Journal on Computing, 33(2):313–331, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38185.

Goldberg:1997:DLC

[GJLR97] Leslie Ann Goldberg, Mark Jerrum, Tom Leighton, and Satish Rao. Doubly logarithmic communication algorithms for optical-communication parallel computers. SIAM Journal on Computing, 26(4):1100– 1119, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
25948.

Goldberg:1998:LBR

[GJM98] Leslie Ann Goldberg, Mark Jerrum, and Philip D. MacKenzie. An Ω(\sqrt{log log n}) lower bound for routing in optical networks. SIAM Journal on Computing, 27(4):1083– 1098, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
27256.

Gaujal:2000:CUR

[GJMM00] Bruno Gaujal, Alain Jean-Marie, and Jean Mairesse. Computations of uniform recurrence equations using minimal memory size. SIAM Journal on Computing, 30 (5):1701–1738, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29035.

Garey:1981:SUT

[GJST81] M. R. Garey, D. S. Johnson, B. B. Simons, and R. E. Tarjan. Scheduling unit-time tasks with arbitrary release times and deadlines. SIAM Journal on Computing, 10(2):256–269, May 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garey:1976:PHC

[GJT76] M. R. Garey, D. S. Johnson, and R. Endre Tarjan. The planar Hamiltonian circuit problem is NP-complete. SIAM Journal on Computing, 5(4): 704–714, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goos:2018:ECI

[GJW18] Mika Göös, Rahul Jain, and Thomas Watson. Extension complexity of independent set polytopes. SIAM Journal on Computing, 47(1):241–269, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gamarnik:2024:HRO

[GJW24] David Gamarnik, Aukosh Jagannath, and Alexander S. Wein. Hardness of random optimization problems for Boolean circuits, low-degree polynomials, and Langevin dynamics. SIAM Journal on Computing, 53(1):1–46, February 2024. CODEN SMJCAT. ISSN 1095-7111.

Gabow:1982:AEC

[GK82] Harold N. Gabow and Oded Kariv. Algorithms for edge coloring bipartite graphs and multigraphs. SIAM Journal on Computing, 11(1):117–129, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldreich:1996:CZK

[GK96] Oded Goldreich and Hugo Krawczyk. On the composition of zero-knowledge proof systems. SIAM Journal on Computing, 25(1):169–192, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22068.

Gibbons:1997:TSM

[GK97] Phillip B. Gibbons and Ephraim Korach. Testing shared memories. SIAM Journal on Computing, 26(4):1208–1244, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27961.

Goldmann:1998:STC

[GK98a] Mikael Goldmann and Marek Karpinski. Simulating threshold circuits by majority circuits. SIAM Journal on Computing, 27(1):230–246, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27451.

Grigoriev:1998:CAC

[GK98b] Dima Grigoriev and Marek Karpinski. Computing the additive complexity of algebraic circuits with root extracting. SIAM Journal on Computing, 27(3):694–701, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25831.

Gafni:1999:TPT

[GK99] Eli Gafni and Elias Koutsoupias. Three-processor tasks are undecidable. SIAM Journal on Computing, 28(3):970–983, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30576.

Goldstein:2005:CNS

[GK05] Darin Goldstein and Kojiro Kobayashi. On the complexity of network synchronization. SIAM Journal on Computing, 35(3):567–589, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garg:2007:FSA

[GK07] Naveen Garg and Jochen Könemann. Faster and simpler algorithms for multicommodity flow and other fractional packing problems. SIAM Journal on Computing, 37(2):630– 652, ???? 2007. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gurevich:2010:CGB

[GK10] Maxim Gurevich and Idit Keidar. Correctness of gossipbased membership under message loss. SIAM Journal on Computing, 39(8):3830–3859, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v39/i8/
                                        p3830_s1.

Gharibian:2012:AAQ

[GK12a] Sevag Gharibian and Julia Kempe. Approximation algorithms for QMA-complete problems. SIAM Journal on Computing, 41(4):1028–1050, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldstein:2012:MTS

[GK12b] Darin Goldstein and Kojiro Kobayashi. On minimal-time solutions of firing squad synchronization problems for networks. SIAM Journal on Computing, 41(3):618–669, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Giannakopoulos:2018:DOA

[GK18] Yiannis Giannakopoulos and Elias Koutsoupias. Duality and optimality of auctions for uniform distributions. SIAM Journal on Computing, 47(1): 121–165, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gavinsky:2009:ESO

[GKK^{+}09] Dmitry Gavinsky, Julia Kempe, Iordanis Kerenidis, Ran Raz, and Ronald de Wolf. Exponential separation for one-way quantum communication complexity, with applications to cryptography. SIAM Journal on Computing, 38(5):1695– 1708, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gibson:2012:CMS

[GKK^{+}12] Matt Gibson, Gaurav Kanade, Erik Krohn, Imran A. Pirwani, and Kasturi Varadarajan. On clustering to minimize the sum of radii. SIAM Journal on Computing, 41(1):

47–60, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v41/i1/p47_s1.

Goel:2013:PMT

[GKK13] Ashish Goel, Michael Kapralov, and Sanjeev Khanna. Perfect matchings in O(n log n) time in regular bipartite graphs. SIAM Journal on Computing, 42(3): 1392–1404, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ghaffari:2018:NOD

[GKK^{+}18] Mohsen Ghaffari, Andreas Karrenbauer, Fabian Kuhn, Christoph Lenzen, and Boaz Patt-Shamir. Near-optimal distributed maximum flow. SIAM Journal on Computing, 47(6):2078–2117, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goddard:1993:ORA

[GKKS93] Wayne Goddard, Claire Kenyon, Valerie King, and Leonard J. Schulman. Optimal randomized algorithms for local sorting and set-maxima. SIAM Journal on Computing, 22(2): 272–283, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gupta:2016:ACC

[GKKS16] Ankit Gupta, Pritish Kamath, Neeraj Kayal, and Ramprasad Saptharishi. Arithmetic circuits: a chasm at depth 3. SIAM Journal on Computing, 45(3):1064–1079, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldreich:1993:EPG

[GKL93] Oded Goldreich, Hugo Krawczyk, and Michael Luby. On the existence of pseudorandom generators. SIAM Journal on Computing, 22(6):1163–1175, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goaoc:2013:BCS

[GKL13] Xavier Goaoc, Hyo-Sil Kim, and Sylvain Lazard. Boundedcurvature shortest paths through a sequence of points using convex optimization. SIAM Journal on Computing, 42(2):662– 684, ???? 2013. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grandoni:2014:UMD

[GKLV14] Fabrizio Grandoni, Piotr Krysta, Stefano Leonardi, and Carmine Ventre. Utilitarian mechanism design for multiobjective optimization. SIAM Journal on Computing, 43(4): 1263–1290, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guttag:1983:PUT

[GKM83] J. V. Guttag, D. Kapur, and D. R. Musser. On proving uniform termination and restricted termination of rewriting systems. SIAM Journal on Computing, 12(1):189–214, February 1983. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gopalan:2018:PDF

[GKM18] Parikshit Gopalan, Daniel M. Kane, and Raghu Meka. Pseudorandomness via the discrete Fourier transform. SIAM Journal on Computing, 47(6):2451– 2487, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gawrychowski:2021:VDP

[GKM^{+}21] Paweł Gawrychowski, Haim Kaplan, Shay Mozes, Micha Sharir, and Oren Weimann. Voronoi diagrams on planar graphs, and computing the diameter in deterministic Õ(n^{5/3}) time. SIAM Journal on Computing, 50(2):509–554, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gopalan:2009:CBS

[GKMP09] Parikshit Gopalan, Phokion G. Kolaitis, Elitza Maneva, and Christos H. Papadimitriou. The connectivity of Boolean satisfiability: Computational and structural dichotomies. SIAM Journal on Computing, 38(6):2330–2355, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garay:1998:STD

[GKP98] Juan A. Garay, Shay Kutten, and David Peleg. A sublinear time distributed algorithm for minimum-weight spanning trees. SIAM Journal on Computing, 27(1):302–316, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26111.

Goldberg:2004:CCC

[GKP04] Leslie Ann Goldberg, Steven Kelk, and Mike Paterson. The complexity of choosing an Hcoloring (nearly) uniformly at random. SIAM Journal on Computing, 33(2):416–432, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40836.

Gupta:2022:CTW

[GKP22] Anupam Gupta, Amit Kumar, and Debmalya Panigrahi. Caching with time Windows and delays. SIAM Journal on Computing, 51(4):975–1017, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1346286.

Grandoni:2008:PDB

[GKPS08] F. Grandoni, J. Könemann, A. Panconesi, and M. Sozio. A primal-dual bicriteria distributed algorithm for capacitated vertex cover. SIAM Journal on Computing, 38(3): 825–840, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gupta:2005:TPD

[GKR05] Anupam Gupta, Amit Kumar, and Rajeev Rastogi. Traveling with a Pez dispenser (or, routing issues in MPLS). SIAM Journal on Computing, 34(2):453–474, April 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40992.

Gupta:2012:OSS

[GKR12] Anupam Gupta, Ravishankar Krishnaswamy, and R. Ravi. Online and stochastic survivable network design. SIAM Journal on Computing, 41(6): 1649–1672, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ganor:2021:ESC

[GKR21] Anat Ganor, Gillat Kol, and Ran Raz. Exponential separation of communication and external information. SIAM Journal on Computing, 50(3):STOC16–236–STOC16– 254, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gavinsky:2009:BEQ

[GKRdW09] Dmitry Gavinsky, Julia Kempe, Oded Regev, and Ronald de Wolf. Bounded-error quantum state identification and exponential separations in communication complexity. SIAM Journal on Computing, 39(1):1–24, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grigoriev:1990:FPA

[GKS90] Dima Yu. Grigoriev, Marek Karpiński, and Michael F. Singer. Fast parallel algorithms for sparse multivariate polynomial interpolation over finite fields. SIAM Journal on Computing, 19(6):1059– 1063, December 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldman:1993:EIR

[GKS93] Sally A. Goldman, Michael J.

Kearns, and Robert E. Schapire. Exact identification of readonce formulas using fixed points of amplification functions. SIAM Journal on Computing, 22(4):705–726, August 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grigoriev:1994:CCS

[GKS94] Dima Grigoriev, Marek Karpinski, and Michael F. Singer. Computational complexity of sparse rational interpolation. SIAM Journal on Computing, 23(1):1–11, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/19406.

Goyal:2008:LBN

[GKS08] Navin Goyal, Guy Kindler, and Michael Saks. Lower bounds for the noisy broadcast problem. SIAM Journal on Computing, 37(6):1806–1841, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gopalan:2010:HRM

[GKS10] Parikshit Gopalan, Subhash Khot, and Rishi Saket. Hardness of reconstructing multivariate polynomials over finite fields. SIAM Journal on Computing, 39(6):2598–2621, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guo:2022:DAH

[GKSS22] Zeyu Guo, Mrinal Kumar, Ramprasad Saptharishi, and Noam Solomon. Derandomization from algebraic hardness. SIAM Journal on Computing, 51(2):315–335, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1347395.

Geerts:2006:LCR

[GKV06] Floris Geerts, Bart Kuijpers, and Jan Van den Bussche. Linearization and completeness results for terminating transitive closure queries on spatial databases. SIAM Journal on Computing, 35(6):1386– 1439, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goyal:2020:CRT

[GKW20] Rishab Goyal, Venkata Koppula, and Brent Waters. Collusion resistant traitor tracing from learning with errors. SIAM Journal on Computing, 49(5):STOC18–94–STOC18– 241, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ginsburg:1977:DCC

[GL77] Seymour Ginsburg and Nancy Lynch. Derivation complexity in context-free grammar forms. SIAM Journal on Computing, 6(1):123–138, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

George:1980:OAS

[GL80] Alan George and Joseph W. H.

Liu. An optimal algorithm for symbolic factorization of symmetric matrices. SIAM Journal on Computing, 9(3):583– 593, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gonzalez:1987:AAR

[GL87] Teofilo F. Gonzalez and Sing-Ling Lee. A 1.6 approximation algorithm for routing multiterminal nets. SIAM Journal on Computing, 16(4):669–704, August 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gillies:1995:STA

[GL95] Donald W. Gillies and Jane W.-S. Liu. Scheduling tasks with AND/OR precedence constraints. SIAM Journal on Computing, 24(4):797– 810, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        21866.

Goller:2013:BTM

[GL13] Stefan Göller and Markus Lohrey. Branching-time model checking of one-counter processes and timed automata. SIAM Journal on Computing, 42(3):884–923, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guruswami:2017:NON

[GL17] Venkatesan Guruswami and Euiwoong Lee. Nearly optimal NP-hardness of unique coverage. SIAM Journal on Computing, 46(3):1018–1028, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gur:2021:PRL

[GL21] Tom Gur and Oded Lachish.

On the power of relaxed local decoding algorithms. SIAM Journal on Computing, 50(2): 788–813, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gharibian:2023:DQS

[GL23] Sevag Gharibian and François Le Gall. Dequantizing the quantum singular value transformation: Hardness and applications to quantum chemistry and the quantum PCP conjecture. SIAM Journal on Computing, 52(4):1009–1038, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1513721.

Glier:2003:KCD

[Gli03]  Oliver Glier.  Kolmogorov

complexity and deterministic context-free languages. SIAM Journal on Computing, 32 (5):1389–1394, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41775.

Golab:2018:CAP

[GLLON18] Wojciech Golab, Xiaozhou Steve Li, Alejandro López-Ortiz, and Naomi Nishimura. Computing k-atomicity in polynomial time. SIAM Journal on Computing, 47(2):420–455, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guo:2019:CHC

[GLLZ19] Heng Guo, Chao Liao, Pinyan Lu, and Chihao Zhang. Counting hypergraph colorings in the local lemma regime. SIAM Journal on Computing, 48(4): 1397–1424, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ghosh:1999:TAT

[GLM^{+}99] Bhaskar Ghosh, F. T. Leighton, Bruce M. Maggs, S. Muthukrishnan, C. Greg Plaxton, R. Rajaraman, Andréa W. Richa, newline Robert E. Tarjan, and David Zuckerman. Tight analyses of two local load balancing algorithms. SIAM Journal on Computing, 29(1):29–64, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29220.

Goos:2016:RNJ

[GLM^{+}16] Mika Göös, Shachar Lovett, Raghu Meka, Thomas Watson, and David Zuckerman. Rectangles are nonnegative juntas. SIAM Journal on Computing, 45(5):1835–1869, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gudmundsson:2002:FGA

[GLN02] Joachim Gudmundsson, Christos Levcopoulos, and Giri Narasimhan. Fast greedy algorithms for constructing sparse geometric spanners. SIAM Journal on Computing, 31 (5):1479–1500, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38294.

Gelenbe:1975:RTF

[GLP75] Erol Gelenbe, Jacques Lenfant, and Dominique Potier. Response time of a fixed-head disk to transfers of variable length. SIAM Journal on Computing, 4(4):461–473, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gilbert:2012:ASR

[GLPS12] Anna C. Gilbert, Yi Li, Ely Porat, and Martin J. Strauss. Approximate sparse recovery: Optimizing time and measurements. SIAM Journal on Computing, 41(2):436–453, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ghashami:2016:FDS

[GLPW16] Mina Ghashami, Edo Liberty, Jeff M. Phillips, and David P. Woodruff. Frequent directions: Simple and deterministic matrix sketching. SIAM Journal on Computing, 45(5):1762– 1792, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guo:2024:ILD

[GLS^{+}24] Zeyu Guo, Ray Li, Chong Shangguan, Itzhak Tamo, and Mary Wootters. Improved list-decodability and list-recoverability of Reed– Solomon codes via tree packings. SIAM Journal on Computing, 53(2):389–430, March 2024. CODEN SMJCAT. ISSN 1095-7111.

Gilbert:1980:PPC

[GLT80] John R. Gilbert, Thomas Lengauer, and Robert Endre Tarjan. The pebbling problem is complete in polynomial space. SIAM Journal on Computing, 9(3):513–524, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guo:2013:CSB

[GLV13] Heng Guo, Pinyan Lu, and Leslie G. Valiant. The complexity of symmetric Boolean parity Holant problems. SIAM Journal on Computing, 42(1): 324–356, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gupta:1982:RUT

[GLW82] Udai Gupta, D. T. Lee, and C. K. Wong. Ranking and unranking of 2-3 trees. SIAM Journal on Computing, 11(3): 582–590, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gabow:1978:FAS

[GM78] Harold N. Gabow and Eugene W. Myers. Finding all spanning trees of directed and undirected graphs. SIAM Journal on Computing, 7(3): 280–287, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gonnet:1979:EOH

[GM79] Gaston H. Gonnet and J. Ian Munro. Efficient ordering of hash tables. SIAM Journal on Computing, 8(3):463–478, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gligor:1982:FAS

[GM82] Virgil Gligor and David Maier.

Finding augmented-set bases. SIAM Journal on Computing, 11(3):602–609, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gonnet:1986:HH

[GM86] Gaston H. Gonnet and J. Ian Munro. Heaps on heaps. SIAM Journal on Computing, 15(4): 964–971, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guessarian:1987:AIT

[GM87] Irène Guessarian and José Meseguer. On the axiomatization of “If–Then–Else”. SIAM Journal on Computing, 16(2): 332–357, April 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1991:ALT

[GM91a] Zvi Galil and Oded Margalit. An almost linear-time algorithm for the dense subsetsum problem. SIAM Journal on Computing, 20(6):1157– 1189, December 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ghosh:1991:OSA

[GM91b] Subir Kumar Ghosh and David M. Mount. An outputsensitive algorithm for computing visibility graphs. SIAM Journal on Computing, 20(5): 888–910, October 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gu:1991:ABD

[GM91c] Qian Ping Gu and Akira Maruoka. Amplification of bounded depth monotone read-once Boolean formulae. SIAM Journal on Computing, 20(1):41–55, February 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gu:1992:LMB

[GM92] Qian Ping Gu and Akira Maruoka. Learning monotone Boolean functions by uniformly distributed examples. SIAM Journal on Computing, 21(3):587–599, June 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garay:1998:FPB

[GM98a] Juan A. Garay and Yoram Moses. Fully polynomial Byzantine agreement for n > 3t processors in t + 1 rounds. SIAM Journal on Computing, 27(1):247–290, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26523.

Gil:1998:SFP

[GM98b] Joseph Gil and Yossi Matias. Simple fast parallel hashing by oblivious execution. SIAM Journal on Computing, 27(5):1348–1375, October 1998. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29158.

Greenhalgh:2000:CCG

[GM00] David Greenhalgh and Stephen Marshall. Convergence criteria for genetic algorithms. SIAM Journal on Computing, 30(1):269–282, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32565.

Gafni:2001:ATB

[GM01] Eli Gafni and Michael Mitzenmacher. Analysis of timingbased mutual exclusion with random times. SIAM Journal on Computing, 31(3):816–837, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36491.

Guha:2009:SOO

[GM09] Sudipto Guha and Andrew McGregor. Stream order and order statistics: Quantile estimation in randomorder streams. SIAM Journal on Computing, 38(5):2044– 2059, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grohe:2015:STI

[GM15] Martin Grohe and Dániel Marx. Structure theorem and isomorphism test for graphs with excluded topological subgraphs. SIAM Journal on Computing, 44(1):114–159, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gottlob:2018:ANU

[GM18] Georg Gottlob and Enrico Malizia. Achieving new upper bounds for the hypergraph duality problem through logic. SIAM Journal on Computing, 47(2):456–492, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grossman:2020:ADS

[GM20] Ofer Grossman and Dana Moshkovitz. Amplification and derandomization without slowdown. SIAM Journal on Computing, 49(5):959–998, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gusfield:1987:FAB

[GMFB87] Dan Gusfield, Charles Martel, and David Fernández-Baca. Fast algorithms for bipartite network flow. SIAM Journal on Computing, 16(2): 237–251, April 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1986:AFM

[GMG86] Zvi Galil, Silvio Micali, and Harold Gabow. An O(EV log V) algorithm for finding a maximal weighted matching in general graphs. SIAM Journal on Computing, 15(1):120–130, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guha:2009:CFA

[GMM09] Sudipto Guha, Adam Meyerson, and Kamesh Munagala. A constant factor approximation for the single sink edge installation problem. SIAM Journal on Computing, 38(6): 2426–2442, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gelade:2009:OSL

[GMN09] Wouter Gelade, Wim Martens, and Frank Neven. Optimizing schema languages for XML: Numerical constraints and interleaving. SIAM Journal on Computing, 38(5):2021–2043, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Geffert:1998:SBS

[GMP98] Viliam Geffert, Carlo Mereghetti, and Giovanni Pighizzini. Sublogarithmic bounds on space and reversals. SIAM Journal on Computing, 28(1):325– 340, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.

org/sam-bin/dbq/article/
30130.

Grigni:2000:DDG

[GMP00] Michelangelo Grigni, Vincent Mirelli, and Christos H. Papadimitriou. On the difficulty of designing good classifiers. SIAM Journal on Computing, 30(1):318–323, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30814.

Goldberg:2005:SSM

[GMP05] Leslie Ann Goldberg, Russell Martin, and Mike Paterson. Strong spatial mixing with fewer colors for lattice graphs. SIAM Journal on Computing, 35(2):486–517, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gogacz:2023:URC

[GMP23] Tomasz Gogacz, Jerzy Marcinkowski, and Andreas Pieris. Uniform restricted chase termination. SIAM Journal on Computing, 52(3):641–683, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1377035.

Georgiou:2010:IGV

[GMPT10] Konstantinos Georgiou, Avner Magen, Toniann Pitassi, and Iannis Tourlakis. Integrality gaps of 2 − o(1) for vertex cover SDPs in the Lovász– Schrijver hierarchy. SIAM Journal on Computing, 39(8): 3553–3570, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldwasser:1988:DSS

[GMR88] Shafi Goldwasser, Silvio Micali, and Ronald L. Rivest. A digital signature scheme secure against adaptive chosenmessage attacks. SIAM Journal on Computing, 17(2):281– 308, April 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Goldwasser:1989:KCI

[GMR89] Shafi Goldwasser, Silvio Micali, and Charles Rackoff. The knowledge complexity of interactive proof systems. SIAM Journal on Computing, 18(1): 186–208, February 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Guibas:1997:RLP

[GMR97] Leonidas J. Guibas, Rajeev Motwani, and Prabhakar Raghavan. The robot localization problem. SIAM Journal on Computing, 26 (4):1120–1138, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23325.

Gibbons:1998:QRQ

[GMR98] Phillip B. Gibbons, Yossi Matias, and Vijaya Ramachandran. The queue-read queuewrite PRAM model: Accounting for contention in parallel algorithms. SIAM Journal on Computing, 28(2):733–769, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27491.

Goldberg:1999:OSS

[GMR99] Leslie Ann Goldberg, Yossi Matias, and Satish Rao. An optical simulation of shared memory. SIAM Journal on Computing, 28(5):1829– 1847, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
29050.

Gopalan:2013:PGC

[GMRZ13] Parikshit Gopalan, Raghu Meka, Omer Reingold, and David Zuckerman. Pseudorandom generators for combinatorial shapes. SIAM Journal on Computing, 42(3):1051–1076, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gonnet:1981:ESO

[GMS81] Gaston H. Gonnet, J. Ian Munro, and Hendra Suwanda.

Exegesis of self-organizing linear search. SIAM Journal on Computing, 10(3):613–637, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garg:2015:CPA

[GMSV15] Jugal Garg, Ruta Mehta, Milind Sohoni, and Vijay V. Vazirani. A complementary pivot algorithm for market equilibrium under separable, piecewise-linear concave utilities. SIAM Journal on Computing, 44(6):1820–1847, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gavinsky:2017:TBF

[GMWW17] Dmitry Gavinsky, Or Meir, Omri Weinstein, and Avi Wigderson. Toward better formula lower bounds: The composition of a function and a universal relation. SIAM Journal on Computing, 46(1):114– 131, ???? 2017. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Golin:2012:HCL

[GMY12] Mordecai J. Golin, Claire Mathieu, and Neal E. Young. Huffman coding with letter costs: a linear-time approximation scheme. SIAM Journal on Computing, 41(3):684–713, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gosset:2016:QSQ

[GN16] David Gosset and Daniel Nagaj. Quantum 3-SAT is QMA_{1}-Complete. SIAM Journal on Computing, 45(3):1080– 1128, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grohe:2020:FIT

[GNS20] Martin Grohe, Daniel Neuen, and Pascal Schweitzer. A faster isomorphism test for graphs of small degree. SIAM Journal on Computing, 52(6):FOCS18– 1–FOCS18–36, February 2020. CODEN SMJCAT. ISSN 1095-7111.

Grohe:2023:ITG

[GNW23] Martin Grohe, Daniel Neuen, and Daniel Wiebking. Isomorphism testing for graphs excluding small minors. SIAM Journal on Computing, 52 (1):238–272, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1401930.

Guibas:1980:NPL

[GO80] Leo J. Guibas and Andrew M.

Odlyzko. A new proof of the linearity of the Boyer– Moore string searching algorithm. SIAM Journal on Computing, 9(4):672–682, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goerdt:1993:RRV

[Goe93] Andreas Goerdt. Regular resolution versus unrestricted resolution. SIAM Journal on Computing, 22(4):661–683, August 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gold:1978:DPE

[Gol78] E. Mark Gold. Deadlock prediction: Easy and difficult cases. SIAM Journal on Computing, 7(3):320–336, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Golan:1981:PBO

[Gol81]  Igal Golan.  Performance

bounds for orthogonal oriented two-dimensional packing algorithms. SIAM Journal on Computing, 10(3):571–582, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldberg:1995:SAS

[Gol95] Andrew V. Goldberg. Scaling algorithms for the shortest paths problem. SIAM Journal on Computing, 24(3):494–504, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23117.

Goldwasser:1998:ISS

[Gol98] Shafi Goldwasser. Introduction to special section on probabilistic proof systems. SIAM Journal on Computing, 27(3):737–738, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97405.

Goldberg:2008:PSP

[Gol08]  Andrew V. Goldberg.  A

practical shortest path algorithm with linear expected time. SIAM Journal on Computing, 37(5):1637–1655, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goodrich:1991:ILS

[Goo91] Michael T. Goodrich. Intersecting line segments in parallel with an output- sensitive number of processors. SIAM Journal on Computing, 20(4): 737–755, August 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goodrich:1999:CEP

[Goo99] Michael T. Goodrich. Communicationefficient parallel sorting. SIAM Journal on Computing, 29(2): 416–432, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
29414.

Goldreich:1998:CCK

[GOP98] Oded Goldreich, Rafail Ostrovsky, and Erez Petrank.

Computational complexity and knowledge complexity. SIAM Journal on Computing, 27 (4):1116–1141, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28052.

Gopalan:2008:QEA

[Gop08] Parikshit Gopalan. Queryefficient algorithms for polynomial interpolation over composites. SIAM Journal on Computing, 38(3):1033–1057, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gopalan:2011:TFD

[GOS^{+}11] Parikshit Gopalan, Ryan O’Donnell, Rocco A. Servedio, Amir Shpilka, and Karl Wimmer. Testing Fourier dimensionality and sparsity. SIAM Journal on Computing, 40(4): 1075–1100, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i4/p1075_s1.

Gotlieb:1981:OMW

[Got81] L. Gotlieb. Optimal multi-way search trees. SIAM Journal on Computing, 10(3):422–433, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goodman:1983:MS

[GP83] Jacob E. Goodman and Richard Pollack. Multidimensional sorting. SIAM Journal on Computing, 12(3):484–507, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1990:IAA

[GP90] Zvi Galil and Kunsoo Park.

An improved algorithm for approximate string matching. SIAM Journal on Computing, 19(6):989–999, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1996:AIT

[GP96] Zvi Galil and Kunsoo Park.

Alphabet-independent twodimensional witness computation. SIAM Journal on Computing, 25(5):907–935, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24194.

Gavoille:2001:CIR

[GP01] Cyril Gavoille and David Peleg. The compactness of interval routing for almost all graphs. SIAM Journal on Computing, 31(3):706–721, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35171.

Gottlob:2004:HMC

[GP04] Georg Gottlob and Reinhard Pichler. Hypergraphs in model checking: Acyclicity and hypertree-width versus clique-width. SIAM Journal on Computing, 33(2):351–378, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39680.

Goos:2018:CLB

[GP18] Mika Göös and Toniann Pitassi. Communication lower bounds via critical block sensitivity. SIAM Journal on Computing, 47(5):1778–1806, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gupta:2011:SCS

[GPRS11] Anupam Gupta, Martin Pál, R. Ravi, and Amitabh Sinha. Sampling and cost-sharing: Approximation algorithms for stochastic optimization problems. SIAM Journal on Computing, 40(5):1361–1401, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i5/
p1361_s1.

Glasser:2006:PNC

[GPSS06] Christian Glaßer, A. Pavan, Alan L. Selman, and Samik Sengupta. Properties of NPcomplete sets. SIAM Journal on Computing, 36(2):516–542, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldberg:1992:UIP

[GPST92] Andrew V. Goldberg, Serge A.

Plotkin, David B. Shmoys, and Éva Tardos. Using interiorpoint methods for fast parallel algorithms for bipartite matching and related problems. SIAM Journal on Computing, 21(1):140–150, February 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Glasser:2008:SNC

[GPSZ08] Christian Glaßer, A. Pavan, Alan L. Selman, and Liyu Zhang. Splitting NP-complete sets. SIAM Journal on Computing, 37(5):1517–1535, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gambosi:2000:ARO

[GPT00] Giorgio Gambosi, Alberto Postiglione, and Maurizio Talamo. Algorithms for the relaxed online bin-packing model. SIAM Journal on Computing, 30(5):1532–1551, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18040.

Gyssens:1994:GBA

[GPV94] Marc Gyssens, Jan Paredaens, and Dirk Van Gucht. A grammar-based approach towards unifying hierarchical data models. SIAM Journal on Computing, 23(6): 1093–1137, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/18816.

Goos:2018:DCV

[GPW18] Mika Göös, Toniann Pitassi, and Thomas Watson. Deterministic communication vs. partition number. SIAM Journal on Computing, 47(6):2435– 2450, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goos:2020:QCL

[GPW20] Mika Göös, Toniann Pitassi, and Thomas Watson. Queryto-communication lifting for BPP. SIAM Journal on Computing, 49(4):FOCS17–441– FOCS17–461, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grochow:2017:AGI

[GQ17] Joshua A. Grochow and Youming Qiao. Algorithms for group isomorphism via group extensions and cohomology. SIAM Journal on Computing, 46(4): 1153–1216, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gouda:1985:PNC

[GR85] Mohamed G. Gouda and Louis E. Rosier. Priority networks of communicating finite state machines. SIAM Journal on Computing, 14(3):569–584, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldstein:1993:FVK

[GR93] Arthur S. Goldstein and Edward M. Reingold. A Fibonacci version of Kraft’s inequality applied to discrete unimodal search. SIAM Journal on Computing, 22(4):751– 777, August 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gal:2002:TSA

[GR02] Anna Gál and Adi Rosén. A theorem on sensitivity and applications in private computation. SIAM Journal on Computing, 31(5):1424–1437, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38529.

Gal:2005:LBA

[GR05] Anna Gál and Adi Rosén.

Ω(log n) lower bounds on the amount of randomness in 2private computation. SIAM Journal on Computing, 34 (4):946–959, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43278.

Guruswami:2009:HLH

[GR09] Venkatesan Guruswami and Prasad Raghavendra. Hardness of learning halfspaces with noise. SIAM Journal on Computing, 39(2):742–765, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldreich:2011:AAP

[GR11a] Oded Goldreich and Dana Ron. Algorithmic aspects of property testing in the dense graphs model. SIAM Journal on Computing, 40(2):376–445, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i2/
p376_s1.

Goldreich:2011:POT

[GR11b] Oded Goldreich and Dana Ron. On proximity-oblivious testing. SIAM Journal on Computing, 40(2):534–566, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i2/
p534_s1.

Goldwasser:2015:HCP

[GR15] Shafi Goldwasser and Guy N.

Rothblum. How to compute in the presence of leakage. SIAM Journal on Computing, 44(5): 1480–1549, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grandoni:2016:PPP

[GR16] Fabrizio Grandoni and Thomas Rothvoß. Pricing on paths: a PTAS for the highway problem. SIAM Journal on Computing, 45(2):216–231, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gupta:2017:PAA

[GR17] Rishi Gupta and Tim Roughgarden. A PAC approach to application-specific algorithm selection. SIAM Journal on Computing, 46(3):992–1017, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goos:2021:NOC

[GR21] Mika Göös and Aviad Rubinstein. Near-optimal communication lower bounds for approximate Nash equilibria. SIAM Journal on Computing, 52(6):FOCS18–316–FOCS18– 348, November 2021. CODEN SMJCAT. ISSN 1095-7111.

Graham:1974:BRC

[Gra74] Susan L. Graham. On bounded right context languages and grammars. SIAM Journal on Computing, 3(3):224–254, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grandjean:1984:SFO

[Gra84] Étienne Grandjean. The spectra of first-order sentences and computational complexity. SIAM Journal on Computing, 13(2):356–373, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grandjean:1988:NNC

[Gra88] Étienne Grandjean. A natural NP-complete problem with a nontrivial lower bound. SIAM Journal on Computing, 17(4): 786–809, August 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gradel:1990:DGC

[Grä90a] Erich Grädel. Domino games and complexity. SIAM Journal on Computing, 19(5):787–804, October 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grandjean:1990:NLB

[Gra90b] Étienne Grandjean. A nontrivial lower bound for an NP problem on automata. SIAM Journal on Computing, 19(3): 438–451, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grandjean:1994:LTA

[Gra94]  Étienne Grandjean.  Lin-

ear time algorithms and NP-Complete problems. SIAM Journal on Computing, 23(3): 573–597, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
22320.

Greibach:1973:HCF

[Gre73]  Sheila A. Greibach.  The

hardest context-free language. SIAM Journal on Computing, 2(4):304–310, ???? 1973. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Greibach:1974:JPH

[Gre74] S. A. Greibach. Jump PDA’s and hierarchies of deterministic context-free languages. SIAM Journal on Computing, 3(2):111–127, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grolmusz:2003:CES

[Gro03] Vince Grolmusz. Computing elementary symmetric polynomials with a subpolynomial number of multiplications. SIAM Journal on Computing, 32(6):1475–1487, December 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42465.

Goyal:2021:AAN

[GRRV21] Vipul Goyal, Silas Richelson, Alon Rosen, and Margarita Vald. An algebraic approach to nonmalleability. SIAM Journal on Computing, 50(5):1537– 1579, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldman:1993:LBR

[GRS93] Sally A. Goldman, Ronald L. Rivest, and Robert E. Schapire. Learning binary relations and total orders. SIAM Journal on Computing, 22(5):1006–1034, October 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Georgiou:2005:WCS

[GRS05] Chryssis Georgiou, Alexander Russell, and Alexander A. Shvartsman. Workcompetitive scheduling for cooperative computing with dynamic groups. SIAM Journal on Computing, 34(4):848– 862, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        44044.

Gabizon:2006:DEB

[GRS06] Ariel Gabizon, Ran Raz, and Ronen Shaltiel. Deterministic extractors for bit-fixing sources by obtaining an independent seed. SIAM Journal on Computing, 36(4):1072–1094, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ghosh:2012:UUM

[GRS12] Arpita Ghosh, Tim Roughgarden, and Mukund Sundararajan. Universally utilitymaximizing privacy mechanisms. SIAM Journal on Computing, 41(6):1673–1693, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gupta:2016:DTD

[GRS16] Rishi Gupta, Tim Roughgarden, and C. Seshadhri. Decompositions of triangledense graphs. SIAM Journal on Computing, 45(2):197–215, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Golin:1998:RDS

[GRSS98] Mordecai Golin, Rajeev Raman, Christian Schwarz, and Michiel Smid. Randomized data structures for the dynamic closest-pair problem. SIAM Journal on Computing, 27(4):1036–1072, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27771.

Gaver:1973:AMP

[GS73] D. P. Gaver and G. S. Shedler.

Approximate models for processor utilization in multiprogrammed computer systems.

SIAM Journal on Computing, 2(3):183–192, ???? 1973. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galil:1980:SSF

[GS80] Zvi Galil and Joel Seiferas. Saving space in fast stringmatching. SIAM Journal on Computing, 9(2):417–438, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gurevich:1987:ECT

[GS87] Yuri Gurevich and Saharon Shelah. Expected computation time for Hamiltonian path problem. SIAM Journal on Computing, 16(3):486–502, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grollmann:1988:CMP

[GS88] Joachim Grollmann and Alan L. Selman. Complexity measures for public-key cryptosystems. SIAM Journal on Computing, 17(2):309–335, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Goldberg:1989:NPA

[GS89] Mark Goldberg and Thomas Spencer. A new parallel algorithm for the maximal independent set problem. SIAM Journal on Computing, 18(2):

419–427, April 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldberg:1991:CR

[GS91] Andrew V. Goldberg and Michael Sipser. Compression and ranking. SIAM Journal on Computing, 20(3):524–536, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goldreich:2000:CCL

[GS00] Oded Goldreich and Shmuel Safra. A combinatorial consistency lemma with application to proving the PCP theorem. SIAM Journal on Computing, 29(4):1132–1154, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31574.

Grandjean:2002:MIC

[GS02] Etienne Grandjean and Thomas Schwentick. Machine-independent characterizations and complete problems for deterministic linear time. SIAM Journal on Computing, 32 (1):196–230, February 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36024.

Goldreich:2006:SIR

[GS06] Oded Goldreich and Madhu Sudan. Special issue on randomness and complexity. SIAM Journal on Computing, 36(4):ix–xi, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gurevich:2007:MPM

[GS07] Yuri Gurevich and Paul Schupp. Membership problem for the modular group. SIAM Journal on Computing, 37(2): 425–459, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Groth:2012:ENP

[GS12] Jens Groth and Amit Sahai. Efficient noninteractive proof systems for bilinear groups. SIAM Journal on Computing, 41(5):1193–1232, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gamarnik:2017:PSL

[GS17a] David Gamarnik and Madhu Sudan. Performance of sequential local algorithms for the random NAE-K-SAT problem. SIAM Journal on Computing, 46(2):590–619, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Greco:2017:PLC

[GS17b] Gianluigi Greco and Francesco Scarcello. The power of local consistency in conjunctive queries and constraint satisfaction problems. SIAM Journal on Computing, 46(3):1111– 1145, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gabow:2021:AWMa

[GS21a] Harold N. Gabow and Piotr Sankowski. Algorithms for weighted matching generalizations I: Bipartite graphs, b-matching, and unweighted f-factors. SIAM Journal on Computing, 50(2):440–486, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gabow:2021:AWMb

[GS21b] Harold N. Gabow and Piotr Sankowski. Algorithms for weighted matching generalizations II: f -factors and the special case of shortest paths. SIAM Journal on Computing, 50(2):555–601, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Glasser:2004:DNP

[GSSZ04] Christian Glasser, Alan L. Selman, Samik Sengupta, and Liyu Zhang. Disjoint NP-pairs. SIAM Journal on Computing, 33(6):1369–1416, December 2004. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42584.

Garg:1999:FSC

[GSV99] Naveen Garg, Huzur Saran, and Vijay V. Vazirani. Finding separator cuts in planar graphs within twice the optimal. SIAM Journal on Computing, 29(1):159–179, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27169.

Galanis:2016:FPM

[GSVY16] Andreas Galanis, Daniel Stefankovic, Eric Vigoda, and Linji Yang. Ferromagnetic Potts model: Refined #BIShardness and related results. SIAM Journal on Computing, 45(6):2004–2065, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grigoriev:1995:CAF

[GSY95] Dima Grigoriev, Michael Singer, and Andrew Yao. On computing algebraic functions using logarithms and exponentials. SIAM Journal on Computing, 24(2):242–246, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24501.

Gabow:1989:FSA

[GT89] Harold N. Gabow and Robert E.

Tarjan. Faster scaling algorithms for network problems. SIAM Journal on Computing, 18(5):1013–1036, October 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goodrich:1998:DTD

[GT98] Michael T. Goodrich and Roberto Tamassia. Dynamic trees and dynamic point location. SIAM Journal on Computing, 28(2):612–636, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25437.

Grolmusz:2000:LBC

[GT00] Vince Grolmusz and Gábor Tardos. Lower bounds for (MODp − MODm) circuits. SIAM Journal on Computing, 29(4):1209–1222, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34085.

Garg:2001:CCU

[GT01] Ashim Garg and Roberto Tamassia. On the computational complexity of upward and rectilinear planarity testing. SIAM Journal on Computing, 31(2):601–625, April 2001. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27712.

Gurjar:2021:IVL

[GTV21] Rohit Gurjar, Thomas Thierauf, and Nisheeth K. Vishnoi. Isolating a vertex via lattices: Polytopes with totally unimodular faces. SIAM Journal on Computing, 50(2):636– 661, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gurari:1982:EPD

[Gur82] Eitan M. Gurari. The equivalence problem for deterministic two-way sequential transducers is decidable. SIAM Journal on Computing, 11(3):448–452, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gusfield:1983:SCM

[Gus83] Dan Gusfield. Simple constructions for multiterminal network flow synthesis. SIAM Journal on Computing, 12(1): 157–165, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gusfield:1987:OMG

[Gus87a] Dan Gusfield. Optimal mixed graph augmentation. SIAM Journal on Computing, 16(4): 599–612, August 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gusfield:1987:TFA

[Gus87b] Dan Gusfield. Three fast algorithms for four problems in stable marriage. SIAM Journal on Computing, 16(1):111–128, February 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gusfield:1988:GTA

[Gus88a] Dan Gusfield. A graph theoretic approach to statistical data security. SIAM Journal on Computing, 17(3):552–571, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gusfield:1988:SSR

[Gus88b] Dan Gusfield. The structure of the stable roommate problem: efficient representation and enumeration of all stable assignments. SIAM Journal on Computing, 17(4):742–769, August 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gusfield:1990:VSM

[Gus90] Dan Gusfield. Very simple methods for all pairs network flow analysis. SIAM Journal on Computing, 19(1):143–155, February 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gusfield:1991:CSG

[Gus91] Dan Gusfield. Computing the strength of a graph. SIAM Journal on Computing, 20(4): 639–654, August 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Grossi:2005:CSA

[GV05] Roberto Grossi and Jeffrey Scott Vitter. Compressed suffix arrays and suffix trees with applications to text indexing and string matching. SIAM Journal on Computing, 35(2):378–407, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Gowers:2019:IGP

[GV19] W. T. Gowers and Emanuele Viola. Interleaved group products. SIAM Journal on Computing, 48(2):554–580, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garg:1996:AMF

[GVY96] Naveen Garg, Vijay V. Vazirani, and Mihalis Yannakakis. Approximate max-flow min- (multi)cut theorems and their applications. SIAM Journal on Computing, 25(2):235–251, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24301.

Garsia:1977:NAM

[GW77] Adriano M. Garsia and Michelle L.

Wachs. A new algorithm for minimum cost binary trees. SIAM Journal on Computing, 6(4):622–642, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Galambos:1993:LSH

[GW93a] Gábor Galambos and Gerhard J. Woeginger. An on-line scheduling heuristic with better worst case ratio than Graham’s list scheduling. SIAM Journal on Computing, 22(2): 349–355, April 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See note [CCG^{+}97].

Gavalda:1993:CCS

[GW93b] Ricard Gavaldà and Osamu Watanabe. On the computational complexity of small descriptions. SIAM Journal on Computing, 22(6):1257–1275, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Goemans:1995:GAT

[GW95] Michel X. Goemans and David P. Williamson. A general approximation technique for constrained forest problems. SIAM Journal on Computing, 24(2):296–317, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24261.

Gao:2017:UGR

[GW17] Pu Gao and Nicholas Wormald. Uniform generation of random regular graphs. SIAM Journal on Computing, 46(4):1395– 1427, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Golin:1996:PCE

[GY96] Mordecai J. Golin and Neal Young. Prefix codes: Equiprobable words, unequal letter costs. SIAM Journal on Computing, 25(6):1281–1292, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26838.

Gao:2005:WSP

[GZ05] Jie Gao and Li Zhang. Wellseparated pair decomposition for the unit-disk graph metric and its applications. SIAM Journal on Computing, 35(1): 151–169, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hadlock:1975:FMC

[Had75] F. Hadlock. Finding a maximum cut of a planar graph in polynomial time. SIAM Journal on Computing, 4(3):221– 225, ???? 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hagerup:1990:PDF

[Hag90]  Torben Hagerup.  Planar

depth-first search in O(log n) parallel time. SIAM Journal on Computing, 19(4):678–704, August 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haitner:2013:PRT

[Hai13] Iftach Haitner. A parallel repetition theorem for any interactive argument. SIAM Journal on Computing, 42(6):2487– 2501, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hall:1975:SPE

[Hal75] Andrew D. Hall, Jr. Solving a problem in eigenvalue approximation with a symbolic algebra system. SIAM Journal on Computing, 4(2):163– 174, ???? 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Halperin:2002:IAA

[Hal02]  Eran Halperin.  Improved

approximation algorithms for the vertex cover problem in graphs and hypergraphs. SIAM Journal on Computing, 31(5):1608–1623, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38109.

Halman:2008:AAD

[Hal08] Nir Halman. On the algorithmic aspects of discrete and lexicographic helly-type theorems and the discrete LP-type model. SIAM Journal on Computing, 38(1):1–45, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hambrusch:1983:VAC

[Ham83] Susanne E. Hambrusch. VLSI algorithms for the connected component problem. SIAM Journal on Computing, 12(2): 354–365, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Han:1996:FDS

[Han96] Yijie Han. A fast derandomization scheme and its applications. SIAM Journal on Computing, 25(1):52–82, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19270.

Hartmanis:1980:SDR

[Har80] J. Hartmanis. On the succinctness of different representations of languages. SIAM Journal on Computing, 9(1): 114–120, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hartmann:1985:MCM

[Har85] W. Hartmann. On the multiplicative complexity of modules over associative algebras. SIAM Journal on Computing, 14(2):383–395, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Harvey:1999:CTD

[Har99] Warwick Harvey. Computing two-dimensional integer hulls. SIAM Journal on Computing, 28(6):2285–2299, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28977.

Harvey:2009:AAM

[Har09] Nicholas J. A. Harvey. Algebraic algorithms for matching and matroid problems. SIAM Journal on Computing, 39(2): 679–702, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Harris:2020:DLA

[Har20] David G. Harris. Distributed local approximation algorithms for maximum matching in graphs and hypergraphs. SIAM Journal on Computing, 49(4):711–746, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haastad:1988:SSM

[Hås88] Johan Håstad. Solving simultaneous modular equations of low degree. SIAM Journal on Computing, 17(2):336–341, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Haastad:1998:SEM

[Hås98] Johan Håstad. The shrinkage exponent of de Morgan formulas is 2. SIAM Journal on Computing, 27(1):48– 64, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        26155.

Haastad:2014:CPS

[Hås14a] Johan Håstad. On the correlation of parity and smalldepth circuits. SIAM Journal on Computing, 43(5):1699– 1708, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haastad:2014:NHM

[Hås14b] Johan Håstad. On the NPhardness of max-not-2. SIAM Journal on Computing, 43(1): 179–193, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Havlicek:2004:NHT

[Hav04] John Havlicek. A note on the homotopy type of wait-free atomic snapshot protocol complexes. SIAM Journal on Computing, 33(5):1215–1222, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33722.

Haviv:2024:FPA

[Hav24] Ishay Haviv. Fixed-parameter algorithms for the Kneser and Schrijver problems. SIAM Journal on Computing, 53(2): 287–314, March 2024. CODEN SMJCAT. ISSN 1095-7111.

Hartmanis:1975:SGN

[HB75] J. Hartmanis and T. P. Baker.

On simple Goedel numberings and translations. SIAM Journal on Computing, 4(1):1–11, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Herley:1994:DSP

[HB94] Kieran T. Herley and Gianfranco Bilardi. Deterministic simulations of PRAMs on bounded degree networks. SIAM Journal on Computing, 23(2):276–292, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/16476.

Hickey:1983:URG

[HC83] Timothy Hickey and Jacques Cohen. Uniform random generation of strings in a contextfree language. SIAM Journal on Computing, 12(4):645–655, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

He:1999:ASP

[HC99] Xin He and Zhi-Zhong Chen. An algorithm for shortest paths in bipartite digraphs with concave weight matrices and its applications. SIAM Journal on Computing, 29(1):65–80, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32225.

Hwang:1989:SPG

[HCAL89] Jing Jang Hwang, Yuan-Chieh Chow, Frank D. Anger, and Chung-Yee Lee. Scheduling precedence graphs in systems with interprocessor communication times. SIAM Journal on Computing, 18(2):244–257, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hagerup:1989:OPC

[HCD89] Torben Hagerup, Marek Chrobak, and Krzysztof Diks. Optimal parallel 5-colouring of planar graphs. SIAM Journal on Computing, 18(2):288–300, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hunt:1980:CCP

[HCS80] H. B. Hunt, III, R. L. Constable, and S. Sahni. On the computational complexity of program scheme equivalence. SIAM Journal on Computing, 9(2):396–416, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

He:1993:FRD

[He93] Xin He. On finding the rectangular duals of planar triangular graphs. SIAM Journal on Computing, 22(6):1218– 1226, December 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

He:1999:FPP

[He99] Xin He. On floor-plan of plane graphs. SIAM Journal on Computing, 28(6): 2150–2167, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30887.

Heller:1984:RPE

[Hel84] Hans Heller. On relativized polynomial and exponential computations. SIAM Journal on Computing, 13(4):717–725, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hendricks:1976:ASO

[Hen76] W. J. Hendricks. An account of self-organizing systems. SIAM Journal on Computing, 5(4): 715–723, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hennessy:1980:SCV

[Hen80] M. C. B. Hennessy. The semantics of call-by-value and call-by-name in a nondeterministic environment. SIAM Journal on Computing, 9(1): 67–84, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hennessy:1988:AFC

[Hen88] Matthew Hennessy. Axiomatising finite concurrent processes. SIAM Journal on Computing, 17(5):997–1017, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Ace94].

Hendrickson:1992:CUG

[Hen92] Bruce Hendrickson. Conditions for unique graph realizations. SIAM Journal on Computing, 21(1):65–84, February 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Henzinger:2000:IDS

[Hen00] Monika R. Henzinger. Improved data structures for fully dynamic biconnectivity.

SIAM Journal on Computing, 29(6):1761–1815, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26390.

Hertli:2014:SFS

[Her14] Timon Hertli. 3-SAT faster and simpler — unique-SAT bounds for PPSZ hold in general. SIAM Journal on Computing, 43(2):718–729, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hampapuram:1998:OBB

[HF98] Haripriyan Hampapuram and Michael L. Fredman. Optimal biweighted binary trees and the complexity of maintaining partial sums. SIAM Journal on Computing, 28(1):1–9, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29159.

Harrison:1972:RTS

[HH72] Michael A. Harrison and Ivan M. Havel. Real-time strict deterministic languages. SIAM Journal on Computing, 1(4): 333–349, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hehner:1979:NRR

[HH79] E. C. R. Hehner and R. N. S. Horspool. A new representation of the rational numbers for fast easy arithmetic. SIAM Journal on Computing, 8(2): 124–134, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See corrigendum [HH80].

Hehner:1980:CNR

[HH80] E. C. R. Hehner and R. N. S. Horspool. Corrigendum: “A new representation of the rational numbers for fast easy arithmetic” [SIAM J. Comput. 8 (1979), no. 2, 124–134, MR 80h:68027]. SIAM Journal on Computing, 9(1):217, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [HH79].

Haggkvist:1981:PSC

[HH81] Roland Häggkvist and Pavol Hell. Parallel sorting with constant time for comparisons. SIAM Journal on Computing, 10(3):465–472, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hemachandra:1991:SEI

[HH91] Lane A. Hemachandra and Albrecht Hoene. On sets with efficient implicit membership tests. SIAM Journal on Computing, 20(6):1148–1156, December 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hemaspaandra:1998:DCW

[HHH98] Edith Hemaspaandra, Lane A.

Hemaspaandra, and Harald Hempel. A downward collapse within the polynomial hierarchy. SIAM Journal on Computing, 28(2):383–393, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30647.

Hemaspaandra:2005:EDC

[HHH05] Edith Hemaspaandra, Lane A.

Hemaspaandra, and Harald Hempel. Extending downward collapse from 1-versus-2 queries to m-versus-m + 1 queries. SIAM Journal on Computing, 34(6):1352–1369, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39100.

Haah:2021:QAS

[HHKL21] Jeongwan Haah, Matthew B.

Hastings, Robin Kothari, and Guang Hao Low. Quantum algorithm for simulating real time evolution of lattice Hamiltonians. SIAM Journal on Computing, 52(6):FOCS18– 250–FOCS18–284, January 2021. CODEN SMJCAT. ISSN 1095-7111.

Hemaspaandra:2007:CCS

[HHKW07] Lane A. Hemaspaandra, Christopher M. Homan, Sven Kosub, and Klaus W. Wagner. The complexity of computing the size of an interval. SIAM Journal on Computing, 36(5):1264– 1300, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hatami:2018:SPX

[HHL18] Hamed Hatami, Kaave Hosseini, and Shachar Lovett. Structure of protocols for XOR functions. SIAM Journal on Computing, 47(1):208–217, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Halldorsson:2017:POW

[HHMW17] Magnús M. Halldórsson, Stephan Holzer, Pradipta Mitra, and Roger Wattenhofer. The power of oblivious wireless power. SIAM Journal on Computing, 46(3):1062–1086, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hemaspaandra:2004:APS

[HHN04] Lane A. Hemaspaandra, Harald Hempel, and Arfst Nickelsen. Algebraic properties for selector functions. SIAM Journal on Computing, 33(6): 1309–1337, December 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42755.

Haitner:2011:PRI

[HHR11] Iftach Haitner, Danny Harnik, and Omer Reingold. On the power of the randomized iterate. SIAM Journal on Computing, 40(6):1486–1528, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i6/
p1486_s1.

Hsu:2016:PMA

[HHR^{+}16] Justin Hsu, Zhiyi Huang, Aaron Roth, Tim Roughgarden, and Zhiwei Steven Wu. Private matchings and allocations. SIAM Journal on Computing, 45(6):1953–1984, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haitner:2015:FCI

[HHRS15] Iftach Haitner, Jonathan J.

Hoch, Omer Reingold, and Gil Segev. Finding collisions in interactive protocols — tight lower bounds on the round and communication complexities of statistically hiding commitments. SIAM Journal on Computing, 44(1):193–242, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Han:1997:TCC

[HHT97] Yenjo Han, Lane A. Hemaspaandra, and Thomas Thierauf. Threshold computation and cryptographic security. SIAM Journal on Computing, 26(1):59–78, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24046.

Hemaspaandra:1998:QO

[HHW98] Lane A. Hemaspaandra, Harald Hempel, and Gerd Wechsung. Query order. SIAM Journal on Computing, 28(2): 637–651, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
29763.

Hwang:2005:SNM

[HHW05] Frank K. Hwang, Yong He, and Yang Wang. Strictly nonblocking multirate log_{d}(N, m, p) networks. SIAM Journal on Computing, 34(5):1271– 1278, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43315.

Hierons:2010:RDS

[Hie10] Robert M. Hierons. Reaching and distinguishing states of distributed systems. SIAM Journal on Computing, 39(8): 3480–3500, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haitner:2011:BBC

[HIK^{+}11] Iftach Haitner, Yuval Ishai, Eyal Kushilevitz, Yehuda Lindell, and Erez Petrank. Blackbox constructions of protocols for secure computation. SIAM Journal on Computing, 40(2): 225–266, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i2/p225_s1.

Hoffmann:2001:PEP

[HIKK01] Frank Hoffmann, Christian Icking, Rolf Klein, and Klaus Kriegel. The polygon exploration problem. SIAM Journal on Computing, 31(2):577–600, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34867.

Haastad:1999:PGO

[HILL99] Johan Håstad, Russell Impagliazzo, Leonid A. Levin, and Michael Luby. A pseudorandom generator from any one-way function. SIAM Journal on Computing, 28 (4):1364–1396, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24470.

Hardt:2016:SSF

[HIMW16] Moritz Hardt, Yuval Ishai, Raghu Meka, and Virginia Vassilevska Williams. Special section on the Fifty-Fourth Annual IEEE Symposium on Foundations of Computer Science (FOCS 2013). SIAM Journal on Computing, 45(3): 881, ???? 2016. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hlineny:2008:FBD

[HiO08] Petr Hlinĕný and Sang il Oum. Finding branch-decompositions and rank-decompositions. SIAM Journal on Computing, 38(3): 1012–1032, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hirschberg:1980:CSS

[Hir80] D. S. Hirschberg. On the complexity of searching a set of vectors. SIAM Journal on Computing, 9(1):126–129, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hirahara:2023:NBB

[Hir23] Shuichi Hirahara. Non-black-box worst-case to averagecase reductions within NP. SIAM Journal on Computing, 52(6):FOCS18–349–FOCS18– 382, December 2023. CODEN SMJCAT. ISSN 1095-7111.

Hitchcock:2004:SSS

[Hit04] John M. Hitchcock. Small spans in scaled dimension. SIAM Journal on Computing, 34(1):170–194, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42641.

Hitchcock:2007:OLR

[Hit07] John M. Hitchcock. Online learning and resource-bounded dimension: Winnow yields new lower bounds for hard sets. SIAM Journal on Computing, 36(6):1696–1708, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hagihara:1979:DPM

[HITK79] Kenichi Hagihara, Minoru Ito, Kenichi Taniguchi, and Tadao Kasami. Decision problems for multi-valued dependencies in relational databases. SIAM Journal on Computing, 8(2): 247–264, May 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hassin:1985:AMF

[HJ85] Refael Hassin and Donald B.

Johnson. An O(n log^{2} n) algorithm for maximum flow in undirected planar networks. SIAM Journal on Computing, 14(3):612–624, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hentzel:1990:CUP

[HJ90] I. R. Hentzel and D. Pokrass Jacobs. Complexity and unsolvability properties of nilpotency. SIAM Journal on Computing, 19(1):32–43, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haastad:1989:PTA

[HJLS89] J. Håstad, B. Just, J. C. Lagarias, and C.-P. Schnorr. Polynomial time algorithms for finding integer relations among real numbers. SIAM Journal on Computing, 18(5):859–881, October 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [HJLS14].

Haastad:2014:EPT

[HJLS14] J. Håstad, B. Just, J. C. Lagarias, and C. P. Schnorr. Erratum: Polynomial Time Algorithms for Finding Integer Relations Among Real Numbers. SIAM Journal on Computing, 43(1):254, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [HJLS89].

Hopcroft:1984:MPD

[HJW84] John E. Hopcroft, Deborah A. Joseph, and Sue H. Whitesides. Movement problems for 2-dimensional linkages. SIAM Journal on Computing, 13(3): 610–629, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hopcroft:1985:MRA

[HJW85] John E. Hopcroft, Deborah A.

Joseph, and Sue H. Whitesides. On the movement of robot arms in 2-dimensional bounded regions. SIAM Journal on Computing, 14(2):315– 333, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hopcroft:1973:AMM

[HK73] John E. Hopcroft and Richard M.

Karp. An n^{5/2} algorithm for maximum matchings in bipartite graphs. SIAM Journal on Computing, 2(4):225–231, December 1973. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hofri:1987:PLR

[HK87] Micha Hofri and Alan G. Konheim. Padded lists revisited. SIAM Journal on Computing, 16(6):1073–1114, December 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Herzberg:2000:EDM

[HK00] Amir Herzberg and Shay Kutten. Early detection of message forwarding faults. SIAM Journal on Computing, 30 (4):1169–1196, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31274.

Henzinger:2001:MMS

[HK01] Monika R. Henzinger and Valerie King. Maintaining minimum spanning forests in dynamic graphs. SIAM Journal on Computing, 31(2):364–374, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32720.

Halevy:2007:DFP

[HK07] Shirley Halevy and Eyal Kushilevitz. Distribution-free property-testing. SIAM Journal on Computing, 37(4):1107– 1138, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hazan:2011:HHI

[HK11] Elad Hazan and Robert Krauthgamer. How hard is it to approximate the best Nash equilibrium? SIAM Journal on Computing, 40(1): 79–91, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i1/p79_s1.

Halperin:2007:IRG

[HKK^{+}07] Eran Halperin, Guy Kortsarz, Robert Krauthgamer, Aravind Srinivasan, and Nan Wang. Integrality ratio for group Steiner trees and directed Steiner trees. SIAM Journal on Computing, 36(5):

1494–1511, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

He:2000:FGM

[HKL00] Xin He, Ming-Yang Kao, and Hsueh-I Lu. A fast general methodology for informationtheoretically optimal encodings of graphs. SIAM Journal on Computing, 30(3):838–846, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35911.

Hariharan:2008:FAM

[HKM08] Ramesh Hariharan, Telikepalli Kavitha, and Kurt Mehlhorn. Faster algorithms for minimum cycle basis in directed graphs. SIAM Journal on Computing, 38(4):1430–1447, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Henzinger:2016:DAA

[HKN16] Monika Henzinger, Sebastian Krinninger, and Danupon Nanongkai. Dynamic approximate all-pairs shortest paths: Breaking the O(mn) barrier and derandomization. SIAM Journal on Computing, 45(3): 947–1006, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Henzinger:2021:DAT

[HKN21] Monika Henzinger, Sebastian Krinninger, and Danupon Nanongkai. A deterministic almost-tight distributed algorithm for approximating single-source shortest paths. SIAM Journal on Computing, 50(3):STOC16–98–STOC16– 137, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Huber:2014:SBV

[HKP14] Anna Huber, Andrei Krokhin, and Robert Powell. Skew bisubmodularity and valued CSPs. SIAM Journal on Computing, 43(3):1064–1084, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Han:1995:CMS

[HKRT95] Xiaofeng Han, Pierre Kelsen, Vijaya Ramachandran, and Robert Tarjan. Computing minimal spanning subgraphs in linear time. SIAM Journal on Computing, 24(6): 1332–1358, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/22450.

Hazan:2017:NOA

[HKSS17] Elad Hazan, Satyen Kale, and Shai Shalev-Shwartz. Nearoptimal algorithms for online matrix prediction. SIAM Journal on Computing, 46(2):744– 773, ???? 2017. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hwang:1972:SAM

[HL72] F. K. Hwang and S. Lin. A simple algorithm for merging two disjoint linearly ordered sets. SIAM Journal on Computing, 1(1):31–39, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hirschberg:1986:ACA

[HL86] D. S. Hirschberg and L. L. Larmore. Average case analysis of marking algorithms. SIAM Journal on Computing, 15(4): 1069–1074, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hirschberg:1987:LWS

[HL87] D. S. Hirschberg and L. L.

Larmore. The least weight subsequence problem. SIAM Journal on Computing, 16(4): 628–638, August 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hutton:1996:UPD

[HL96] Michael D. Hutton and Anna Lubiw. Upward planar drawing of single-source acyclic digraphs. SIAM Journal on Computing, 25(2):291–311, April 1996. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23590.

Hassin:2004:EPT

[HL04a] Refael Hassin and Asaf Levin. An efficient polynomial time approximation scheme for the constrained minimum spanning tree problem using matroid intersection. SIAM Journal on Computing, 33(2):261– 268, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
42677.

Honkala:2004:ICT

[HL04b] Iiro Honkala and Tero Laihonen. On identifying codes in the triangular and square grids. SIAM Journal on Computing, 33(2):304–312, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43311.

Hassin:2005:BTG

[HL05a] Refael Hassin and Asaf Levin. A better-than-greedy approximation algorithm for the minimum set cover problem. SIAM Journal on Computing, 35(1): 189–200, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Huo:2005:OSP

[HL05b] Yumei Huo and Joseph Y.- T. Leung. Online scheduling of precedence constrained tasks. SIAM Journal on Computing, 34(3):743–762, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44444.

Hatami:2014:CTA

[HL14] Hamed Hatami and Shachar Lovett. Correlation testing for affine invariant properties on F_{p}^{n} in the high error regime. SIAM Journal on Computing, 43(4):1417–1455, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ho:1991:MDS

[HLCW91] Jan-Ming Ho, D. T. Lee, Chia-Hsiang Chang, and C. K. Wong. Minimum diameter spanning trees and related problems. SIAM Journal on Computing, 20(5):987–997, October 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hlineny:2005:PAM

[Hli05] Petr Hlinený. A parametrized algorithm for matroid branchwidth. SIAM Journal on Computing, 35(2):259–277, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Han:1995:SJT

[HLL95] Ching-Chih Han, Kwei-Jay Lin, and Jane W.-S. Liu. Scheduling jobs with temporal distance constraints. SIAM Journal on Computing, 24 (5):1104–1121, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/21808.

Hajiaghayi:2017:ONW

[HLP17] MohammadTaghi Hajiaghayi, Vahid Liaghat, and Debmalya Panigrahi. Online nodeweighted Steiner forest and extensions via disk paintings. SIAM Journal on Computing, 46(3):911–935, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haastad:1996:ABP

[HLR96] Johan Håstad, Tom Leighton, and Brian Rogoff. Analysis of backoff protocols for multiple access channels. SIAM Journal on Computing, 25(4):740– 774, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23382.

Haramaty:2018:BIP

[HLV18] Elad Haramaty, Chin Ho Lee, and Emanuele Viola. Bounded independence plus noise fools products. SIAM Journal on Computing, 47(2):493–523, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hopcroft:1973:DAC

[HM73] J. E. Hopcroft and J. Musinski.

Duality applied to the complexity of matrix multiplication and other bilinear forms. SIAM Journal on Computing, 2(3):159–173, April 1973. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hartmanis:1981:LSC

[HM81] J. Hartmanis and S. Mahaney. Languages simultaneously complete for one-way and two-way log-tape automata. SIAM Journal on Computing, 10(2):383–390, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Helmbold:1987:TPS

[HM87] David Helmbold and Ernst Mayr. Two processor scheduling is in N C. SIAM Journal on Computing, 16(4):747–759, August 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hafner:1991:AFT

[HM91] James L. Hafner and Kevin S.

McCurley. Asymptotically fast triangularization of matrices over rings. SIAM Journal on Computing, 20(6):1068–1083, December 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Heffernan:1995:OAC

[HM95] Paul J. Heffernan and Joseph S. B. Mitchell. An optimal algorithm for computing visibility in the plane. SIAM Journal on Computing, 24(1):184–201, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22150.

Hsu:1999:FSA

[HM99] Wen-Lian Hsu and Tze-Heng Ma. Fast and simple algorithms for recognizing chordal comparability graphs and interval graphs. SIAM Journal on Computing, 28(3):1004– 1020, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
22481.

Hind:1998:TCC

[HMR98] Hugh Hind, Michael Molloy, and Bruce Reed. Total coloring with Δ + poly(log Δ) colors. SIAM Journal on Computing, 28(3):816–821, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29457.

Huang:2018:MMY

[HMR18] Zhiyi Huang, Yishay Mansour, and Tim Roughgarden. Making the most of your samples. SIAM Journal on Computing, 47(3):651–674, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hunt:1998:CPC

[HMRS98] Harry B. Hunt, III, Madhav V.

Marathe, Venkatesh Radhakrishnan, and Richard E. Stearns. The complexity of planar counting problems. SIAM Journal on Computing, 27(4):1142–1167, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30460.

Halpern:2001:CEB

[HMW01] Joseph Y. Halpern, Yoram Moses, and Orli Waarts. A characterization of eventual Byzantine agreement. SIAM Journal on Computing, 31(3): 838–865, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34021.

Hikita:1977:CCS

[HN77] T. Hikita and A. Nozaki.

A completeness criterion for spectra. SIAM Journal on Computing, 6(2):285–297, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See corrigenda [HN79].

Hikita:1979:CCC

[HN79] T. Hikita and A. Nozaki. Corrigenda: “A completeness criterion for spectra” [SIAM J. Comput. 6 (1977), no. 2, 285– 297, MR 57 #2759]. SIAM Journal on Computing, 8(4): 656, ???? 1979. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [HN77].

Hochbaum:1994:SFA

[HN94] Dorit S. Hochbaum and Joseph (Seffi) Naor. Simple and fast algorithms for linear and integer programs with two variables per inequality. SIAM Journal on Computing, 23(6):1179–1192, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25187.

Hwang:2002:PCL

[HN02] Hsien-Kuei Hwang and Ralph Neininger. Phase change of limit laws in the quicksort recurrence under varying toll functions. SIAM Journal on Computing, 31(6): 1687–1722, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38390.

Harnik:2010:CIC

[HN10] Danny Harnik and Moni Naor.

On the compressibility of N P instances and cryptographic applications. SIAM Journal on Computing, 39(5):1667–1713, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haitner:2009:SHC

[HNO^{+}09] Iftach Haitner, Minh-Huyen Nguyen, Shien Jin Ong, Omer Reingold, and Salil Vadhan. Statistically hiding commitments and statistical zeroknowledge arguments from any one-way function. SIAM Journal on Computing, 39(3):1153– 1218, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haitner:2020:CTP

[HNO^{+}20] Iftach Haitner, Kobbi Nissim, Eran Omri, Ronen Shaltiel, and Jad Silbak. Computational two-party correlation: a dichotomy for key-agreement protocols. SIAM Journal on Computing, 49(6):1041–1082, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hemaspaandra:1996:CSU

[HNOS96] Lane A. Hemaspaandra, Ashish V.

Naik, Mitsunori Ogihara, and Alan L. Selman. Computing solutions uniquely collapses the polynomial hierarchy. SIAM Journal on Computing, 25(4):697–708, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26831.

Haxell:2008:AVH

[HNR08] P. E. Haxell, B. Nagle, and V. Rödl. An algorithmic version of the hypergraph regularity method. SIAM Journal on Computing, 37(6):1728– 1776, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hiptmair:2002:GTS

[HO02] R. Hiptmair and J. Ostrowski. Generators of H_{1}(Γ_{h}, Z) for triangulated surfaces: Construction and classification. SIAM Journal on Computing, 31(5):1405–1423, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38652.

Haitner:2014:CFC

[HO14] Iftach Haitner and Eran Omri. Coin flipping with constant bias implies one-way functions. SIAM Journal on Computing, 43(2):389–409, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hochbaum:1982:AAS

[Hoc82] Dorit S. Hochbaum. Approximation algorithms for the set covering and vertex cover problems. SIAM Journal on Computing, 11(3):555–556, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Holland:1973:GAO

[Hol73] John H. Holland. Genetic algorithms and the optimal allocation of trials. SIAM Journal on Computing, 2(2):88– 105, April 1973. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [Hol74].

Holland:1974:EGA

[Hol74] John H. Holland. Erratum:

Genetic algorithms and the optimal allocation of trials. SIAM Journal on Computing, 3(4): 326, ???? 1974. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Hol73].

Holyer:1981:NCE

[Hol81a]  Ian Holyer.  The NP-

completeness of edge-coloring. SIAM Journal on Computing, 10(4):718–720, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Holyer:1981:NCS

[Hol81b]  Ian Holyer.  The NP-

completeness of some edgepartition problems. SIAM Journal on Computing, 10(4): 713–717, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hong:1982:SDS

[Hon82] Jia Wei Hong. On some deterministic space complexity problems. SIAM Journal on Computing, 11(3):591–601, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hoover:1990:FRF

[Hoo90] H. James Hoover. Feasible real functions and arithmetic circuits. SIAM Journal on Computing, 19(1):182–204, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Horton:1987:PTA

[Hor87] J. D. Horton. A polynomialtime algorithm to find the shortest cycle basis of a graph. SIAM Journal on Computing, 16(2):358–366, April 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hosken:1975:OPT

[Hos75] W. H. Hosken. Optimum partitions of tree addressing structures. SIAM Journal on Computing, 4(3):341–347, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Halperin:1992:EMP

[HOS92] Dan Halperin, Mark H. Overmars, and Micha Sharir. Efficient motion planning for an L-shaped object. SIAM Journal on Computing, 21(1):1–23, February 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Han:2000:FSC

[HOS00] Xiaoxu Han, Suely Oliveira, and David Stewart. Finding sets covering a point with application to mesh-free Galerkin methods. SIAM Journal on Computing, 30 (4):1368–1383, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35936.

Higham:1998:AOE

[HP98] Lisa Higham and Teresa Przytycka. Asymptotically optimal election on weighted rings. SIAM Journal on Computing, 28(2):720–732, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28879.

Har-Peled:1999:CAS

[HP99a] Sariel Har-Peled. Constructing approximate shortest path maps in three dimensions. SIAM Journal on Computing, 28(4):1182–1197, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32522.

Heath:1999:SQLb

[HP99b] Lenwood S. Heath and Sriram V. Pemmaraju. Stack and queue layouts of directed acyclic graphs: Part II. SIAM Journal on Computing, 28(5):1588–1626, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29155.

Har-Peled:2000:CPC

[HP00a] Sariel Har-Peled. Constructing planar cuttings in theory and practice. SIAM Journal on Computing, 29(6): 2016–2039, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35023.

Har-Peled:2000:TWP

[HP00b] Sariel Har-Peled. Taking a walk in a planar arrangement. SIAM Journal on Computing, 30(4):1341–1367, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36262.

Hertel:2010:PCB

[HP10] Philipp Hertel and Toniann Pitassi. The PSPACEcompleteness of black-white pebbling. SIAM Journal on Computing, 39(6):2622–2682, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Har-Peled:2013:ANN

[HPK13] Sariel Har-Peled and Nirman Kumar. Approximate nearest neighbor search for lowdimensional queries. SIAM Journal on Computing, 42(1): 138–159, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Har-Peled:2014:RHR

[HPK14] Sariel Har-Peled and Nirman Kumar. Down the rabbit hole: Robust proximity search and density estimation in sublinear space. SIAM Journal on Computing, 43(4):1486–1511, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Har-Peled:2015:AMD

[HPK15] Sariel Har-Peled and Nirman Kumar. Approximating minimization diagrams and generalized proximity search. SIAM Journal on Computing, 44(4): 944–974, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Har-Peled:2006:FCN

[HPM06] Sariel Har-Peled and Manor Mendel. Fast construction of nets in low-dimensional metrics and their applications. SIAM Journal on Computing, 35(5):1148–1184, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Har-Peled:2017:AAP

[HPQ17] Sariel Har-Peled and Kent Quanrud. Approximation algorithms for polynomialexpansion and low-density graphs. SIAM Journal on Computing, 46(6):1712–1744, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Heath:1999:SQLa

[HPT99] Lenwood S. Heath, Sriram V. Pemmaraju, and Ann N. Trenk. Stack and queue layouts of directed acyclic graphs: Part I. SIAM Journal on Computing, 28(4):1510–1539, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28028.

Har-Peled:2004:SFO

[HPW04] Sariel Har-Peled and Yusu Wang. Shape fitting with outliers. SIAM Journal on Computing, 33(2):269–285, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42796.

Huang:2021:CCS

[HPZZ21] Dawei Huang, Seth Pettie, Yixiang Zhang, and Zhijun Zhang. The communication complexity of set intersection and multiple equality testing. SIAM Journal on Computing, 50(2):674–717, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haskins:1973:TCP

[HR73] Loren Haskins and Donald J.

Rose. Toward characterization of perfect elimination digraphs. SIAM Journal on Computing, 2(4):217–224, ???? 1973. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hunt:1978:CPB

[HR78] H. B. Hunt, III and D. J.

Rosenkrantz. Computational parallels between the regular and context-free languages. SIAM Journal on Computing, 7(1):99–114, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hong:1982:GAB

[HR82] Jia Wei Hong and Arnold L.

Rosenberg. Graphs that are almost binary trees. SIAM Journal on Computing, 11(2):227– 242, ???? 1982. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hunt:1986:RSR

[HR86] H. B. Hunt, III and D. J. Rosenkrantz. Recursion schemes and recursive programs are exponentially hard to analyze. SIAM Journal on Computing, 15(3):831–850, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hofri:1987:OCT

[HR87] Micha Hofri and Keith W. Ross. On the optimal control of two queues with server setup times and its analysis. SIAM Journal on Computing, 16(2): 399–420, April 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Heath:1992:LGU

[HR92] Lenwood S. Heath and Arnold L. Rosenberg. Laying out graphs using queues. SIAM Journal on Computing, 21(5):927–958, October 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Halstenberg:1993:DMC

[HR93a] Bernd Halstenberg and Rüdiger Reischuk. Different modes of communication. SIAM Journal on Computing, 22(5):913–934, October 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Harel:1993:DPN

[HR93b] David Harel and Danny Raz. Deciding properties of nonregular programs. SIAM Journal on Computing, 22(4):857–874, August 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hsu:1993:FSA

[HR93c] Tsan-Sheng Hsu and Vijaya Ramachandran. Finding a smallest augmentation to biconnect a graph. SIAM Journal on Computing, 22(5):889– 912, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hemaspaandra:1997:UCB

[HR97] Lane A. Hemaspaandra and Jörg Rothe. Unambiguous computation: Boolean hierarchies and sparse Turingcomplete sets. SIAM Journal on Computing, 26(3):634–653, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26197.

Husfeldt:2003:NLB

[HR03] Thore Husfeldt and Theis Rauhe. New lower bound techniques for dynamic partial sums and related problems. SIAM Journal on Computing, 32(3):736–753, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39159.

Hunt:1987:CCA

[HRB87] H. B. Hunt, III, D. J. Rosenkrantz, and P. A. Bloniarz. On the computational complexity of algebra on lattices. SIAM Journal on Computing, 16(1):129–148, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hallgren:2003:HSP

[HRTS03] Sean Hallgren, Alexander Russell, and Amnon Ta-Shma. The hidden subgroup problem and quantum computation using group representations. SIAM Journal on Computing, 32(4):916–934, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39450.

Haitner:2013:EIC

[HRV13] Iftach Haitner, Omer Reingold, and Salil Vadhan. Efficiency improvements in constructing pseudorandom generators from one-way functions. SIAM Journal on Computing, 42(3):1405–1430, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Henzinger:2020:LFP

[HRW20] Monika Henzinger, Satish Rao, and Di Wang. Local flow partitioning for faster edge connectivity. SIAM Journal on Computing, 49(1):1–36, ???? 2020.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hemenway:2020:LLR

[HRZW20] Brett Hemenway, Noga Ron-Zewi, and Mary Wootters. Local list recovery of highrate tensor codes and applications. SIAM Journal on Computing, 49(4):FOCS17– 157–FOCS17–195, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Horowitz:1975:CPS

[HS75] Ellis Horowitz and Sartaj Sahni. The computation of powers of symbolic polynomials. SIAM Journal on Computing, 4(2):201–208, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Harper:1979:LBS

[HS79] L. H. Harper and J. E. Savage. Lower bounds on synchronous combinational complexity. SIAM Journal on Computing, 8(2):115–119, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hu:1982:CMC

[HS82] T. C. Hu and M. T. Shing.

Computation of matrix chain products. Part I. SIAM Journal on Computing, 11(2):362– 373, ???? 1982. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hu:1984:CMC

[HS84] T. C. Hu and M. T. Shing. Computation of matrix chain products. Part II. SIAM Journal on Computing, 13(2):228– 251, ???? 1984. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See comments [SW19].

Hambrusch:1985:SUG

[HS85a] Susanne E. Hambrusch and Janos Simon. Solving undirected graph problems on VLSI. SIAM Journal on Computing, 14(3):527–544, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hart:1985:CPP

[HS85b] Sergiu Hart and Micha Sharir. Concurrent probabilistic programs, or: how to schedule if you must. SIAM Journal on Computing, 14(4):991– 1012, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hunt:1987:NAO

[HS87] H. B. Hunt, III and R. E. Stearns. Nonlinear algebra and optimization on rings are “hard”. SIAM Journal on Computing, 16(5):910–929, October 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hochbaum:1988:PAS

[HS88] Dorit S. Hochbaum and David B. Shmoys. A polynomial approximation scheme for scheduling on uniform processors: using the dual approximation approach. SIAM Journal on Computing, 17(3):539– 551, June 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hunt:1990:CVS

[HS90] H. B. Hunt, III and R. E.

Stearns. The complexity of very simple Boolean formulas with applications. SIAM Journal on Computing, 19(1):44– 70, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hemaspaandra:1995:ECG

[HS95] Lane A. Hemaspaandra and Riccardo Silvestri. Easily checked generalized selfreducibility. SIAM Journal on Computing, 24(4):840– 858, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23490.

Hershberger:1997:MSS

[HS97] John Hershberger and Subhash Suri. Matrix searching with the shortest-path metric. SIAM Journal on Computing, 26(6):1612–1634, December 1997. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25357.

Hershberger:1999:OAE

[HS99] John Hershberger and Subhash Suri. An optimal algorithm for Euclidean shortest paths in the plane. SIAM Journal on Computing, 28(6): 2215–2256, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28960.

Han:2002:PIS

[HS02] Yijie Han and Xiaojun Shen. Parallel integer sorting is more efficient than parallel comparison sorting on exclusive write PRAMs. SIAM Journal on Computing, 31(6): 1852–1878, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35244.

Hromkovic:2003:NCL

[HS03] Juraj Hromkovic and Georg Schnitger. Nondeterministic communication with a limited number of advice bits. SIAM Journal on Computing, 33(1):43–68, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41462.

Hoest:2006:TTC

[HS06] Gunnar Hoest and Nir Shavit.

Toward a topological characterization of asynchronous complexity. SIAM Journal on Computing, 36(2):457–497, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hell:2001:FDA

[HSS01] Pavol Hell, Ron Shamir, and Roded Sharan. A fully dynamic algorithm for recognizing and representing proper interval graphs. SIAM Journal on Computing, 31(1):289– 305, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37221.

Hon:2009:BTS

[HSS09] Wing-Kai Hon, Kunihiko Sadakane, and Wing-Kin Sung. Breaking a time-and-space barrier in constructing full-text indices. SIAM Journal on Computing, 38(6):2162– 2178, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haramaty:2013:OTM

[HSS13] Elad Haramaty, Amir Shpilka, and Madhu Sudan. Optimal testing of multivariate polynomials over small prime fields.

SIAM Journal on Computing, 42(2):536–562, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

HajiAghayi:2022:ALC

[HSSS22] MohammadTaghi HajiAghayi, Masoud Seddighin, Saeedreza Seddighin, and Xiaorui Sun. Approximating longest common subsequence in linear time: Beating the \sqrt{n} barrier. SIAM Journal on Computing, 51(4):1341–1367, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1272068.

Hershberger:2005:BSP

[HST05] John Hershberger, Subhash Suri, and Csaba D. Tóth. Binary space partitions of orthogonal subdivisions. SIAM Journal on Computing, 34(6): 1380–1397, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44570.

Haeupler:2011:RPH

[HST11] Bernhard Haeupler, Siddhartha Sen, and Robert E. Tarjan. Rank-pairing heaps. SIAM Journal on Computing, 40(6): 1463–1485, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.

siam.org/sicomp/resource/
1/smjcat/v40/i6/p1463_s1.

Hsu:1985:MWC

[Hsu85] Wen-Lian Hsu. Maximum weight clique algorithms for circular-arc graphs and circle graphs. SIAM Journal on Computing, 14(1):224–231, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hsu:1995:ARI

[Hsu95] Wen Lian Hsu. O(M · N)

algorithms for the recognition and isomorphism problems on circular-arc graphs. SIAM Journal on Computing, 24(3):411–439, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26072.

Hutchinson:2005:DBP

[HSV05] David A. Hutchinson, Peter Sanders, and Jeffrey Scott Vitter. Duality between prefetching and queued writing with parallel disks. SIAM Journal on Computing, 34(6): 1443–1463, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43157.

Harary:2007:CCF

[HSV07] Frank Harary, Wolfgang Slany, and Oleg Verbitsky. On the computational complexity of the forcing chromatic number. SIAM Journal on Computing, 37(1):1–19, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Helmbold:1992:LIL

[HSW92] David Helmbold, Robert Sloan, and Manfred K. Warmuth. Learning integer lattices. SIAM Journal on Computing, 21(2):240–266, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hershberger:2022:NOA

[HSY22] John Hershberger, Subhash Suri, and Hakan Yildiz. A near-optimal algorithm for shortest paths among curved obstacles in the plane. SIAM Journal on Computing, 51 (4):1296–1340, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1428248.

Hopcroft:1973:DGT

[HT73] J. E. Hopcroft and R. E. Tarjan. Dividing a graph into triconnected components. SIAM Journal on Computing, 2(3): 135–158, August 1973. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Halton:1982:FAE

[HT82a] J. H. Halton and R. Terada. A fast algorithm for the Euclidean traveling salesman problem, optimal with probability one. SIAM Journal on Computing, 11(1):28–46, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hofri:1982:WSS

[HT82b] M. Hofri and P. Tzelnic. The working set size distribution for the Markov chain model of program behavior. SIAM Journal on Computing, 11(3): 453–466, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Harel:1984:FAF

[HT84a] Dov Harel and Robert Endre Tarjan. Fast algorithms for finding nearest common ancestors. SIAM Journal on Computing, 13(2):338–355, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Heyman:1984:DPT

[HT84b] D. P. Heyman and S. Tsur.

Disk performance in a transactionoriented system. SIAM Journal on Computing, 13(4):669– 681, ???? 1984. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Heckler:1998:CAP

[HT98] C. Heckler and L. Thiele. Complexity analysis of a parallel lattice basis reduction algorithm. SIAM Journal on Computing, 27(5):1295–1302, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29562.

Hrubes:2015:SPD

[HT15] Pavel Hrubes and Iddo Tzameret. Short proofs for the determinant identities. SIAM Journal on Computing, 44(2): 340–383, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Haitner:2017:AOF

[HT17] Iftach Haitner and Eliad Tsfadia. An almost-optimally fair three-party coin-flipping protocol. SIAM Journal on Computing, 46(2):479–542, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Halldorsson:2021:EWS

[HT21] Magnús M. Halldórsson and Tigran Tonoyan. Effective wireless scheduling via hypergraph sketches. SIAM Journal on Computing, 50(2):718–759, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hwang:2013:TPD

[HTC13] Hsien-Kuei Hwang, Tsung-Hsi Tsai, and Wei-Mei Chen. Threshold phenomena in kdominant skylines of random samples. SIAM Journal on Computing, 42(2):405–441, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hecht:1972:FGR

[HU72] Matthew S. Hecht and Jeffrey D. Ullman. Flow graph reducibility. SIAM Journal on Computing, 1(2):188–202, June 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hopcroft:1973:SMA

[HU73] J. E. Hopcroft and J. D. Ullman. Set merging algorithms. SIAM Journal on Computing, 2(4):294–303, December 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hu:1974:OCS

[Hu74] T. C. Hu. Optimum communication spanning trees. SIAM Journal on Computing, 3(3): 188–195, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hecht:1975:SAG

[HU75] Matthew S. Hecht and Jeffrey D. Ullman. A simple algorithm for global data flow analysis problems. SIAM Journal on Computing, 4(4):519–532, December 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hoza:2022:TPG

[HU22] William M. Hoza and Chris Umans. Targeted pseudorandom generators, simulation advice generators, and derandomizing logspace. SIAM Journal on Computing, 51(2): STOC17–281–STOC17–304, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1145707.

Hull:1986:RIC

[Hul86] Richard Hull. Relative information capacity of simple relational database schemata. SIAM Journal on Computing, 15(3):856–886, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Humenik:1989:REM

[Hum89] Keith Humenik. Ratio estimators are maximum-likelihood estimators for non-context-free grammars. SIAM Journal on Computing, 18(5):1048–1055, October 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Huynh:1986:CMP

[Huỳ86a] Dũng T. Huỳnh. The complexity of the membership problem for two subclasses of polynomial ideals. SIAM Journal on Computing, 15(2):581–594, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Huynh:1986:SOA

[Huỳ86b] Dũng T. Huỳnh. Some observations about the randomness of hard problems. SIAM Journal on Computing, 15(4): 1101–1105, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Harvey:2020:APL

[HV20] Nicholas J. A. Harvey and Jan Vondrák. An algorithmic proof of the Lovász local lemma via resampling oracles. SIAM Journal on Computing, 49(2):394–428, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Halpern:2004:CAR

[HvdMV04] Joseph Y. Halpern, Ron van der Meyden, and Moshe Y. Vardi. Complete axiomatizations for reasoning about knowledge and time. SIAM Journal on Computing, 33(3): 674–703, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32090.

Healy:2006:UNA

[HVV06] Alexander Healy, Salil Vadhan, and Emanuele Viola. Using nondeterminism to amplify hardness. SIAM Journal on Computing, 35(4):903–931, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hopcroft:1986:RMO

[HW86] J. E. Hopcroft and G. T. Wilfong. Reducing multiple object motion planning to graph searching. SIAM Journal on Computing, 15(3):768– 785, ???? 1986. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hu:1987:BST

[HW87] T. C. Hu and Michelle L. Wachs. Binary search on a tape. SIAM Journal on Computing, 16(3):573–590, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hernandez:1993:BCT

[HW93] Héctor J. Hernández and Ke Wang. On the boundedness of constant-time-maintainable database schemes. SIAM Journal on Computing, 22(1):29– 45, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hofting:1995:MCP

[HW95] Franz Höfting and Egon Wanke. Minimum cost paths in periodic graphs. SIAM Journal on Computing, 24 (5):1051–1067, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23437.

Hemaspaandra:2002:MPB

[HW02] Edith Hemaspaandra and Gerd Wechsung. The minimization problem for Boolean formulas. SIAM Journal on Computing, 31(6):1948–1958, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36263.

Hougardy:2004:PEG

[HW04] Stefan Hougardy and Annegret Wagler. Perfectness is an elusive graph property. SIAM Journal on Computing, 34(1):109–117, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42679.

Hwang:1980:OME

[Hwa80] F. K. Hwang. Optimal merging of 3 elements with n elements. SIAM Journal on Computing, 9(2):298–320, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Huang:1987:DRS

[HY87] Wen Qi Huang and Xiang Dong Yu. A DNF without regular shortest consensus path. SIAM Journal on Computing, 16(5):836–840, October 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

He:1988:NOP

[HY88] Xin He and Yaacov Yesha. A nearly optimal parallel algorithm for constructing depth first spanning trees in planar graphs. SIAM Journal on Computing, 17(3):486–491, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Huang:2017:CCD

[HY17] Zengfeng Huang and Ke Yi. The communication complexity of distributed epsilonapproximations. SIAM Journal on Computing, 46(4):1370– 1394, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hubacek:2020:HCL

[HY20] Pavel Hubácek and Eylon Yogev. Hardness of continuous local search: Query complexity and cryptographic lower bounds. SIAM Journal on Computing, 49(6):1128–1172, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hyafil:1976:BS

[Hya76] Laurent Hyafil. Bounds for selection. SIAM Journal on Computing, 5(1):109–114, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hyafil:1979:PEM

[Hya79] Laurent Hyafil. On the parallel evaluation of multivariate polynomials. SIAM Journal on Computing, 8(2):120– 123, ???? 1979. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Henderson:1977:GTC

[HZ77] Peter B. Henderson and Yechezkel Zalcstein. A graphtheoretic characterization of the PV_{chunk} class of synchronizing primitives. SIAM Journal on Computing, 6(1):88– 108, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hoza:2020:SOH

[HZ20] William M. Hoza and David Zuckerman. Simple optimal hitting sets for small-success RL. SIAM Journal on Computing, 49(4):811–820, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Imai:1986:EAG

[IA86] Hiroshi Imai and Takao Asano.

Efficient algorithms for geometric graph search problems. SIAM Journal on Computing, 15(2):478–494, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibaraki:1973:FSR

[Iba73] Toshihide Ibaraki. Finite state representations of discrete optimization problems. SIAM Journal on Computing, 2(3): 193–210, ???? 1973. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1974:HTP

[Iba74] Oscar H. Ibarra. A hierarchy theorem for polynomialspace recognition. SIAM Journal on Computing, 3(3):184– 187, ???? 1974. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1978:UEP

[Iba78] Oscar H. Ibarra. The unsolvability of the equivalence problem for ε-free NGSM’s with unary input (output) alphabet and applications. SIAM Journal on Computing, 7(4):524– 532, ???? 1978. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Igarashi:1977:TCS

[Iga77] Y. Igarashi. The tape comilexity of some classes of Szilard languages. SIAM Journal on Computing, 6(3):460–466, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ito:1985:SRR

[IIK85] Minoru Ito, Motoaki Iwasaki, and Tadao Kasami. Some results on the representative instance in relational databases. SIAM Journal on Computing, 14(2):334–354, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Imai:1985:VDL

[IIM85] Hiroshi Imai, Masao Iri, and Kazuo Murota. Voronoĭ diagrams in the Laguerre geometry and its applications. SIAM Journal on Computing, 14(1): 93–105, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ishihara:2005:TRA

[IISI05] Yasunori Ishihara, Shin Ishii, Hiroyuki Seki, and Minoru Ito. Temporal reasoning about two concurrent sequences of events. SIAM Journal on Computing, 34(2):498–513, April 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36217.

Ibarra:1987:OWC

[IJ87] Oscar H. Ibarra and Tao Jiang.

On one-way cellular arrays. SIAM Journal on Computing, 16(6):1135–1154, December 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1991:PAO

[IJ91] Oscar H. Ibarra and Tao Jiang. The power of alternating onereversal counters and stacks. SIAM Journal on Computing, 20(2):278–290, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Irving:1994:TDS

[IJ94] Robert W. Irving and Mark R. Jerrum. Three-dimensional statistical data security problems. SIAM Journal on Computing, 23(1):170–184, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/19101.

Impagliazzo:2009:ALD

[IJK09] Russell Impagliazzo, Ragesh Jaiswal, and Valentine Kabanets. Approximate listdecoding of direct product codes and uniform hardness amplification. SIAM Journal on Computing, 39(2):564–605, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Impagliazzo:2010:UDP

[IJKW10] Russell Impagliazzo, Ragesh Jaiswal, Valentine Kabanets, and Avi Wigderson. Uniform direct product theorems: Simplified, optimized, and derandomized. SIAM Journal on Computing, 39(4):1637–1665, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1995:NDR

[IJTW95] Oscar H. Ibarra, Tao Jiang, Nicholas Tran, and Hui Wang. New decidability results concerning two-way counter machines. SIAM Journal on Computing, 24(1):123–137, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24062.

Idziak:2022:SMC

[IK22a] Paweł M. Idziak and Jacek Krzaczkowski. Satisfiability in MultiValued circuits. SIAM Journal on Computing, 51(3):337–378, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/18M1220194.

Iwata:2022:WLM

[IK22b] Satoru Iwata and Yusuke Kobayashi. A weighted linear matroid parity algorithm. SIAM Journal on Computing, 51(2):STOC17– 238–STOC17–280, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1141709.

Ishai:2011:ABB

[IKK^{+}11] Yuval Ishai, Jonathan Katz, Eyal Kushilevitz, Yehuda Lindell, and Erez Petrank. On achieving the “best of both worlds” in secure multiparty computation. SIAM Journal on Computing, 40(1):122–141, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i1/
                                        p122_s1.

Kawarabayashi:2018:ANM

[iKK18] Ken ichi Kawarabayashi and Yusuke Kobayashi. All-or-nothing multicommodity flow problem with bounded fractionality in planar graphs. SIAM Journal on Computing, 47(4):1483–1504, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Im:2019:TBO

[IKKP19] Sungjin Im, Nathaniel Kell, Janardhan Kulkarni, and Debmalya Panigrahi. Tight bounds for online vector scheduling. SIAM Journal on Computing, 48(1):93–121, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Icking:2004:OCS

[IKL^{+}04] Christian Icking, Rolf Klein, Elmar Langetepe, Sven Schuierer, and Ines Semrau. An optimal competitive strategy for walking in streets. SIAM Journal on Computing, 33(2):462–486, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41935.

Ibarra:1985:SMC

[IKM85] Oscar H. Ibarra, Sam M. Kim, and Shlomo Moran. Sequential machine characterizations of trellis and cellular automata and applications. SIAM Journal on Computing, 14(2):426– 447, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Immorlica:2012:SSF

[IKM^{+}12] Nicole Immorlica, Jonathan Katz, Michael Mitzenmacher, Rocco Servedio, and Chris Umans. Special section on the Forty-First Annual ACM Symposium on Theory of Computing (STOC 2009). SIAM Journal on Computing, 41(6): 1591–1592, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ishai:2009:ZKP

[IKOS09] Yuval Ishai, Eyal Kushilevitz, Rafail Ostrovsky, and Amit Sahai. Zero-knowledge proofs from secure multiparty computation. SIAM Journal on Computing, 39(3):1121–1152, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Irani:1996:SCA

[IKP96] Sandy Irani, Anna R. Karlin, and Steven Phillips. Strongly competitive algorithms for paging with locality of reference. SIAM Journal on Computing, 25(3):477–497, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23635.

Ivanyos:2010:DPT

[IKS10] Gábor Ivanyos, Marek Karpinski, and Nitin Saxena. Deterministic polynomial time algorithms for matrix completion problems. SIAM Journal on Computing, 39(8):3736–3751, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v39/i8/
                                        p3736_s1.

Impagliazzo:2012:NDP

[IKW12] Russell Impagliazzo, Valentine Kabanets, and Avi Wigderson. New direct-product testers and 2-query PCPs. SIAM Journal on Computing, 41(6):1722– 1768, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1981:CPF

[IL81] Oscar H. Ibarra and Brian S. Leininger. Characterizations of Presburger functions. SIAM Journal on Computing, 10(1): 22–39, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1982:CEP

[IL82a] Oscar H. Ibarra and Brian S.

Leininger. The complexity of the equivalence problem for simple loop-free programs. SIAM Journal on Computing, 11(1):15–27, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1982:SLP

[IL82b] Oscar H. Ibarra and Brian S.

Leininger. Straight-line programs with one input variable. SIAM Journal on Computing, 11(1):1–14, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Irving:1986:CCS

[IL86] Robert W. Irving and Paul Leather. The complexity of counting stable marriages. SIAM Journal on Computing, 15(3):655–667, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ivkovic:1998:FDA

[IL98] Zoran Ivkovic and Errol L.

Lloyd. Fully dynamic algorithms for bin packing: Being (mostly) myopic helps.

SIAM Journal on Computing, 28(2):574–611, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27674.

Ivansson:2004:ARM

[IL04] Lars Ivansson and Jens Lagergren. Algorithms for RH mapping: New ideas and improved analysis. SIAM Journal on Computing, 34(1):89– 108, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38835.

Iliopoulos:1989:WCCa

[Ili89a] Costas S. Iliopoulos. Worstcase complexity bounds on algorithms for computing the canonical structure of finite Abelian groups and the Hermite and Smith normal forms of an integer matrix. SIAM Journal on Computing, 18(4): 658–669, August 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Iliopoulos:1989:WCCb

[Ili89b] Costas S. Iliopoulos. Worstcase complexity bounds on algorithms for computing the canonical structure of infinite Abelian groups and solving systems of linear Diophantine equations. SIAM Journal on Computing, 18(4):670–678, August 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Itai:1981:CGS

[ILPR81] Alon Itai, Richard J. Lipton, Christos H. Papadimitriou, and M. Rodeh. Covering graphs by simple circuits. SIAM Journal on Computing, 10(4):746–750, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1983:STS

[IM83] Oscar H. Ibarra and Shlomo Moran. Some time-space tradeoff results concerning single-tape and offline TMs. SIAM Journal on Computing, 12(2):388–394, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Im:2020:FSI

[IM20] Sungjin Im and Benjamin Moseley. Fair scheduling via iterative quasi-uniform sampling. SIAM Journal on Computing, 49(3):658–680, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Immerman:1987:LCC

[Imm87] Neil Immerman. Languages that capture complexity classes. SIAM Journal on Computing, 16(4):760–778, August 1987.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Immerman:1988:NSC

[Imm88] Neil Immerman. Nondeterministic space is closed under complementation. SIAM Journal on Computing, 17(5):935–938, October 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Immerman:1989:EPC

[Imm89] Neil Immerman. Expressibility and parallel complexity. SIAM Journal on Computing, 18(3): 625–638, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Idziak:2010:TLA

[IMM^{+}10] Pawe Idziak, Petar Marković, Ralph McKenzie, Matthew Valeriote, and Ross Willard. Tractability and learnability arising from algebras with few subpowers. SIAM Journal on Computing, 39(7):3023–3037, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Im:2014:OSG

[IMP14] Sungjin Im, Benjamin Moseley, and Kirk Pruhs. Online scheduling with general cost functions. SIAM Journal on Computing, 43(1):126–143, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ibarra:1987:ESS

[IP87] Oscar H. Ibarra and Michael A.

Palis. On efficient simulations of systolic arrays by randomaccess machines. SIAM Journal on Computing, 16(2):367– 377, April 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Itai:1982:HPG

[IPS82] Alon Itai, Christos H. Papadimitriou, and Jayme Luiz Szwarcfiter. Hamilton paths in grid graphs. SIAM Journal on Computing, 11(4):676–686, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Impagliazzo:1997:SDT

[IPS97] Russell Impagliazzo, Ramamohan Paturi, and Michael E. Saks. Size–Depth tradeoffs for threshold circuits. SIAM Journal on Computing, 26(3): 693–707, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28296.

Ivanyos:2019:ABA

[IQ19] Gábor Ivanyos and Youming Qiao. Algorithms based on *-algebras, and their applications to isomorphism of polynomials with one secret, group isomorphism, and polynomial identity testing. SIAM Journal on Computing, 48(3):926– 963, ???? 2019. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Itai:1978:FMC

[IR78] Alon Itai and Michael Rodeh. Finding a minimum circuit in a graph. SIAM Journal on Computing, 7(4):413–423, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Irani:1996:VCD

[IR96] Sandy Irani and Yuval Rabani. On the value of coordination in distributed decision making. SIAM Journal on Computing, 25(3):498–519, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26142.

Itai:1979:MFP

[IS79] Alon Itai and Yossi Shiloach. Maximum flow in planar networks. SIAM Journal on Computing, 8(2):135–150, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Iwano:1990:SCP

[IS90] Kazuo Iwano and Kenneth Steiglitz. A semiring on convex polygons and zero-sum cycle problems. SIAM Journal on Computing, 19(5):883–901, October 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Izumi:2012:GPT

[ISK^{+}12] Taisuke Izumi, Samia Souissi, Yoshiaki Katayama, Nobuhiro Inuzuka, Xavier Défago, Koichi Wada, and Masafumi Yamashita. The gathering problem for two oblivious robots with unreliable compasses. SIAM Journal on Computing, 41(1):26–46, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v41/i1/p26_s1.

Ishaque:2012:SPR

[IST12] Mashhood Ishaque, Bettina Speckmann, and Csaba D. Tóth. Shooting permanent rays among disjoint polygons in the plane. SIAM Journal on Computing, 41(4):1005– 1027, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Itai:1976:OAT

[Ita76] Alon Itai. Optimal alphabetic trees. SIAM Journal on Computing, 5(1):9–18, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Indyk:2022:OEM

[IW22] Piotr Indyk and Tal Wagner.

Optimal (Euclidean) metric compression. SIAM Journal on Computing, 51(3):467–491, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1371324.

Iwama:1989:CST

[Iwa89] Kazuo Iwama. CNF-satisfiability test by counting and polynomial average time. SIAM Journal on Computing, 18(2):385– 391, April 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Iwama:1993:R

[Iwa93] Kazuo Iwama. ASPACE(o(log log n)) is regular. SIAM Journal on Computing, 22(1):136–146, February 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Iwata:2003:FSA

[Iwa03]  Satoru Iwata.  A faster

scaling algorithm for minimizing submodular functions. SIAM Journal on Computing, 32(4):833–840, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39781.

Iwata:2016:HIL

[IWY16] Yoichi Iwata, Magnus Wahlström, and Yuichi Yoshida. Halfintegrality, LP-branching, and FPT algorithms. SIAM Journal on Computing, 45(4):1377– 1411, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Iwata:2023:FME

[IY23] Satoru Iwata and Yu Yokoi.

Finding maximum edge-disjoint paths between multiple terminals. SIAM Journal on Computing, 52(5):1230–1268, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1494804.

Jackson:1972:EFS

[Jac72] D. M. Jackson. Expectations of functions of sequences over finite alphabets with given transition probabilities by methods independent of sequence length. SIAM Journal on Computing, 1(3):203–217, ???? 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jackson:1990:SCC

[Jac90] Bill Jackson. Shortest circuit covers and postman tours in graphs with a nowhere zero 4flow. SIAM Journal on Computing, 19(4):659–665, August 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jaffe:1980:BST

[Jaf80] Jeffrey M. Jaffe. Bounds on the scheduling of typed task systems. SIAM Journal on Computing, 9(3):541–551, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jaffe:1985:DMR

[Jaf85] Jeffrey M. Jaffe. Distributed multidestination routing: the constraints of local information. SIAM Journal on Computing, 14(4):875–888, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jain:2007:PTA

[Jai07] Kamal Jain. A polynomial time algorithm for computing an Arrow–Debreu market equilibrium for linear utilities. SIAM Journal on Computing, 37(1):303–318, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

JaJa:1979:OEP

[JáJ79] Joseph JáJá. Optimal evaluation of pairs of bilinear forms. SIAM Journal on Computing, 8(3):443–462, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

JaJa:1980:CBF

[JáJ80] Joseph JáJá. On the complexity of bilinear forms with commutativity. SIAM Journal on Computing, 9(4):713– 728, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jansen:2010:PAS

[Jan10] Klaus Jansen. Parameterized approximation scheme for the multiple knapsack problem. SIAM Journal on Computing, 39(4):1392–1412, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jayanti:1998:SCC

[Jay98] Prasad Jayanti. Solvability of consensus: Composition breaks down for nondeterministic types. SIAM Journal on Computing, 28(3):782–797, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28008.

Johnstone:1990:SPA

[JB90] John K. Johnstone and Chanderjit L. Bajaj. Sorting points along an algebraic curve. SIAM Journal on Computing, 19(5):925–967, October 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

JaJa:1991:PAC

[JC91] Joseph JáJá and Shing-Chong Chang. Parallel algorithms for channel routing in the knockknee model. SIAM Journal on Computing, 20(2):228–245, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Johnson:1974:WCP

[JDU^{+}74] D. S. Johnson, A. Demers, J. D. Ullman, M. R. Garey, and R. L. Graham. Worstcase performance bounds for simple one-dimensional packing algorithms. SIAM Journal on Computing, 3(4):299–325, December 1974. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jin:2022:TRG

[JJLZ22] Yaonan Jin, Shunhua Jiang, Pinyan Lu, and Hengjie Zhang. Tight revenue gaps among multiunit mechanisms. SIAM Journal on Computing, 51 (5):1535–1579, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1456364.

Jensen:1982:CMP

[JK82] Per M. Jensen and Bernhard Korte. Complexity of matroid property algorithms. SIAM Journal on Computing, 11(1): 184–190, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Johnson:1983:OCQ

[JK83] D. S. Johnson and A. Klug. Optimizing conjunctive queries that contain untyped variables. SIAM Journal on Computing, 12(4):616–640, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jouannaud:1986:CSR

[JK86] Jean-Pierre Jouannaud and Hélène Kirchner. Completion of a set of rules modulo a set of equations. SIAM Journal on Computing, 15(4):1155–1194, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jonsson:2006:ATV

[JKK06] Peter Jonsson, Mikael Klasson, and Andrei Krokhin. The approximability of three-valued MAX CSP. SIAM Journal on Computing, 35(6):1329–1349, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jover:1990:ASC

[JKLAR90] J. M. Jover, T. Kailath, H. Lev-Ari, and S. K. Rao. On the analysis of synchronous computing systems. SIAM Journal on Computing, 19(4): 627–643, August 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jansen:2005:PTA

[JKLS05] Klaus Jansen, Marek Karpinski, Andrzej Lingas, and Eike Seidel. Polynomial time approximation schemes for MAX-BISECTION on planar and geometric graphs. SIAM Journal on Computing, 35(1): 110–119, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jonsson:2008:MOG

[JKN08] Peter Jonsson, Fredrik Kuivinen, and Gustav Nordh. MAX ONES generalized to larger domains. SIAM Journal on Computing, 38(1):329–365, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jenssen:2020:ABH

[JKP20] Matthew Jenssen, Peter Keevash, and Will Perkins. Algorithms for #BIS-hard problems on expander graphs. SIAM Journal on Computing, 49(4):681–710, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

JaJa:1984:ITU

[JKS84] J. Ja’Ja’, V. K. Prasanna Kumar, and J. Simon. Information transfer under different sets of protocols. SIAM Journal on Computing, 13(4):840– 849, ???? 1984. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Joshi:1977:CSD

[JL77] Aravind K. Joshi and Leon S. Levy. Constraints on structural descriptions: Local transformations. SIAM Journal on Computing, 6(2):272–284, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jiang:1995:ASC

[JL95a] Tao Jiang and Ming Li. On the approximation of shortest common supersequences and longest common subsequences. SIAM Journal on Computing, 24(5):1122–1139, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23842.

Juedes:1995:CDH

[JL95b] David W. Juedes and Jack H.

Lutz. The complexity and distribution of hard problems. SIAM Journal on Computing, 24(2):279–295, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23813.

Jansson:2012:CIM

[JLL12] Jesper Jansson, Richard S.

Lemence, and Andrzej Lingas. The complexity of inferring a minimally resolved phylogenetic supertree. SIAM Journal on Computing, 41(1):272–291, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jin:2020:TRG

[JLTX20] Yaonan Jin, Pinyan Lu, Zhihao Gavin Tang, and Tao Xiao. Tight revenue gaps among simple mechanisms. SIAM Journal on Computing, 49(5):927– 958, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Johnson:1978:STE

[JM78] Donald B. Johnson and Tetsuo Mizoguchi. Selecting the K th element in X + Y and X_{1} + X_{2} + ⋯ + X_{m}. SIAM Journal on Computing, 7(2):147– 153, ???? 1978. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Johnson:1986:NBA

[JM86] Rodney W. Johnson and Aileen M. McLoughlin. Noncommutative bilinear algorithms for 3 × 3 matrix multiplication. SIAM Journal on Computing, 15(2):595–603, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jim:1996:FAC

[JM96a] Trevor Jim and Albert R. Meyer. Full abstraction and the context lemma. SIAM Journal on Computing, 25(3): 663–696, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        22421.

Jimbo:1996:MCS

[JM96b] S. Jimbo and A. Maruoka. A method of constructing selection networks with O(log n) depth. SIAM Journal on Computing, 25(4):709–739, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24832.

Jiang:2011:NOB

[JMP11] Tao Jiang, Zevi Miller, and Dan Pritikin. Near optimal bounds for Steiner trees in the hypercube. SIAM Journal on Computing, 40(5):1340–1360, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i5/
p1340_s1.

Jansson:2006:ACR

[JNS06] Jesper Jansson, Nguyen Bao Nguyen, and Wing-Kin Sung. Algorithms for combining rooted triplets into a galled phylogenetic network. SIAM Journal on Computing, 35(5): 1098–1121, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jamison:1992:RSG

[JO92] Beverly Jamison and Stephan Olariu. Recognizing P_{4}-sparse graphs in linear time. SIAM Journal on Computing, 21(2): 381–406, April 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Johnson:1975:FAE

[Joh75] Donald B. Johnson. Finding all the elementary circuits of a directed graph. SIAM Journal on Computing, 4(1):77–84, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

John:1988:NLB

[Joh88] John Welliaveetil John. A new lower bound for the setpartitioning problem. SIAM Journal on Computing, 17(4): 640–647, August 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jones:1987:NBS

[Jon87] Douglas W. Jones. A note on bottom-up skew heaps. SIAM Journal on Computing, 16(1): 108–110, February 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jansen:2005:GMT

[JP05] Klaus Jansen and Lorant Porkolab. General multiprocessor task scheduling: Approximate solutions in linear time. SIAM Journal on Computing, 35(3):519–530, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jurdzinski:2008:DSA

[JPZ08] Marcin Jurdziński, Mike Paterson, and Uri Zwick. A deterministic subexponential algorithm for solving parity games. SIAM Journal on Computing, 38(4):1519–1532, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jiang:1993:MNP

[JR93] Tao Jiang and B. Ravikumar.

Minimal NFA problems are hard. SIAM Journal on Computing, 22(6):1117–1141, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Johnson:1994:CNR

[JR94] Donald B. Johnson and Larry Raab. Complexity of network reliability and optimal resource placement problems. SIAM Journal on Computing, 23(3):510–519, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23114.

Jha:2013:TRL

[JR13] Madhav Jha and Sofya Raskhodnikova. Testing and reconstruction of Lipschitz functions with applications to data privacy. SIAM Journal on Computing, 42(2):700–731, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jansen:2020:QPA

[JR20] Klaus Jansen and Lars Rohwedder. A quasi-polynomial approximation for the restricted assignment problem. SIAM Journal on Computing, 49(6):1083–1108, ???? 2020.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jones:1981:NCG

[JS81] Neil D. Jones and Sven Skyum. A note on the complexity of general D0L membership. SIAM Journal on Computing, 10(1):114–117, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

JaJa:1982:PAG

[JS82] Joseph Ja’Ja’ and Janos Simon. Parallel algorithms in graph theory: planarity testing. SIAM Journal on Computing, 11(2):314–328, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jacquet:1989:UCB

[JS89a] Philippe Jacquet and Wojciech Szpankowski. Ultimate characterizations of the burst response of an interval searching algorithm: a study of a functional equation. SIAM Journal on Computing, 18(4):777–791, August 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jerrum:1989:AP

[JS89b] Mark Jerrum and Alistair Sinclair. Approximating the permanent. SIAM Journal on Computing, 18(6):1149–1178, December 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jerrum:1993:PTA

[JS93] Mark Jerrum and Alistair Sinclair. Polynomial-time approximation algorithms for the Ising model. SIAM Journal on Computing, 22(5):1087– 1116, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jackson:2005:LRL

[JS05] Jeffrey C. Jackson and Rocco A.

Servedio. Learning random log-depth decision trees under uniform distribution. SIAM Journal on Computing, 34 (5):1107–1128, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44455.

Jain:2008:MCI

[JS08] Sanjay Jain and Frank Stephan.

Mitotic classes in inductive inference. SIAM Journal on Computing, 38(4):1283–1299, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jukna:2020:ALP

[JS20] Stasys Jukna and Hannes Seiwert. Approximation limitations of pure dynamic programming. SIAM Journal on Computing, 49(1):170–205, ???? 2020. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Jump:1973:EAC

[JT73] J. Robert Jump and P. S. Thiagarajan. On the equivalence of asynchronous control structures. SIAM Journal on Computing, 2(2):67–87, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

JaJa:1986:VDS

[JT86] Joseph Ja’Ja’ and Jean Takche. On the validity of the direct sum conjecture. SIAM Journal on Computing, 15(4):1004– 1020, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jansen:2010:AAS

[JT10] Klaus Jansen and Ralf Thöle. Approximation algorithms for scheduling parallel jobs. SIAM Journal on Computing, 39(8): 3571–3615, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jayanti:2000:TSL

[JTT00] Prasad Jayanti, King Tan, and Sam Toueg. Time and space lower bounds for nonblocking implementations. SIAM Journal on Computing, 30(2):438– 456, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.

org/sam-bin/dbq/article/
31729.

Jukna:2006:DSL

[Juk06] Stasys Jukna. Disproving the single level conjecture. SIAM Journal on Computing, 36(1): 83–98, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Just:1992:GCF

[Jus92] Bettina Just. Generalizing the continued fraction algorithm to arbitrary dimensions. SIAM Journal on Computing, 21(5): 909–926, October 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jackson:1975:EGD

[JV75] D. M. Jackson and G. H. J.

Van Rees. The enumeration of generalized double stochastic nonnegative integer square matrices. SIAM Journal on Computing, 4(4):474–477, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jain:2008:ECA

[JV08] Kamal Jain and Vijay V. Vazirani. Equitable cost allocations via primal–dual-type algorithms. SIAM Journal on Computing, 38(1):241–256, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Jayaram:2021:PSD

[JW21] Rajesh Jayaram and David Woodruff. Perfect L_{p} sampling in a data stream. SIAM Journal on Computing, 50(2):382– 439, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Joichi:1980:CGC

[JWW80] J. T. Joichi, Dennis E. White, and S. G. Williamson. Combinatorial Gray codes. SIAM Journal on Computing, 9(1): 130–141, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kadin:1988:PTH

[Kad88]  Jim Kadin.  The polyno-

mial time hierarchy collapses if the Boolean hierarchy collapses. SIAM Journal on Computing, 17(6):1263–1282, December 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [Kad91].

Kadin:1991:EPT

[Kad91] Jim Kadin. Erratum: “The polynomial time hierarchy collapses if the Boolean hierarchy collapses” [SIAM J. Comput. 17 (1988), no. 6, 1263–1282, MR 90a:03060]. SIAM Journal on Computing, 20(2):404, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Kad88].

Kasai:1979:CPG

[KAI79] Takumi Kasai, Akeo Adachi, and Shigeki Iwata. Classes of pebble games and complete problems. SIAM Journal on Computing, 8(4):574–586, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kalorkoti:1985:LBF

[Kal85a] K. A. Kalorkoti. A lower bound for the formula size of rational functions. SIAM Journal on Computing, 14(3):678– 687, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaltofen:1985:PTR

[Kal85b] Erich Kaltofen. Polynomialtime reductions from multivariate to bi- and univariate integral polynomial factorization. SIAM Journal on Computing, 14(2):469–489, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kalorkoti:1993:IPF

[Kal93] K. Kalorkoti. Inverting polynomials and formal power series. SIAM Journal on Computing, 22(3):552–559, June 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kalorkoti:1995:RAM

[Kal95] K. Kalorkoti. On the reuse of additions in matrix multiplication. SIAM Journal on Computing, 24(6):1305–1312, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
25176.

Kaminski:2005:LBC

[Kam05] Michael Kaminski. A lower bound on the complexity of polynomial multiplication over finite fields. SIAM Journal on Computing, 34(4):960– 992, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
44211.

Kao:1993:LPNa

[Kao93]  Ming-Yang Kao.  Linear-

processor NC algorithms for planar directed graphs. I. strongly connected components. SIAM Journal on Computing, 22(3):431–459, June 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kao:1995:PSC

[Kao95] Ming-Yang Kao. Planar strong connectivity helps in parallel depth-first search. SIAM Journal on Computing, 24(1):46– 62, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
22707.

Kao:1997:TPA

[Kao97] Ming-Yang Kao. Total protection of analytic-invariant information in cross-tabulated tables. SIAM Journal on Computing, 26(1):231–242, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25358.

Kao:1998:TCE

[Kao98] Ming-Yang Kao. Tree contractions and evolutionary trees. SIAM Journal on Computing, 27(6):1592–1616, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28350.

Kapoor:2000:DMM

[Kap00]  Sanjiv Kapoor.  Dynamic

maintenance of maxima of 2d point sets. SIAM Journal on Computing, 29(6): 1858–1877, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34836.

Karp:1979:PAN

[Kar79] Richard M. Karp. A patching algorithm for the nonsymmetric traveling-salesman problem. SIAM Journal on Computing, 8(4):561–573, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kari:1992:NPO

[Kar92] Jarkko Kari. The nilpotency problem of one-dimensional cellular automata. SIAM Journal on Computing, 21(3):571– 586, June 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karger:1999:RFP

[Kar99a] David R. Karger. A randomized fully polynomial time approximation scheme for the all-terminal network reliability problem. SIAM Journal on Computing, 29(2):492–514, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29834.

Karloff:1999:HGG

[Kar99b] Howard Karloff. How good is the Goemans–Williamson MAX CUT algorithm? SIAM Journal on Computing, 29 (1):336–350, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32148.

Kavitha:2014:SPT

[Kav14] Telikepalli Kavitha. A sizepopularity tradeoff in the stable marriage problem. SIAM Journal on Computing, 43(1): 52–71, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kinariwala:1976:TOS

[KB76] Bharat Kinariwala and Kabekode V. S. Bhat. Theory of output set assignments and degree switching operations. SIAM Journal on Computing, 5(4): 589–601, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kannan:1979:PAC

[KB79] Ravindran Kannan and Achim Bachem. Polynomial algorithms for computing the Smith and Hermite normal forms of an integer matrix. SIAM Journal on Computing, 8(4):499–507, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khuller:2003:LSP

[KBP03] Samir Khuller, Randeep Bhatia, and Robert Pless. On local search and placement of meters in networks. SIAM Journal on Computing, 32(2):470–487, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36335.

Kuo:1994:PMP

[KC94] David Kuo and Gerard J.

Chang. The profile minimization problem in trees.

SIAM Journal on Computing, 23(1):71–81, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19607.

Kapron:1996:NCT

[KC96] B. M. Kapron and S. A. Cook. A new characterization of type-2 feasibility. SIAM Journal on Computing, 25(1):117–132, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26345.

Keohane:1982:TCS

[KCH82] John Keohane, John C. Cherniavsky, and Peter B. Henderson. On transforming control structures. SIAM Journal on Computing, 11(2):268–286, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Krumme:1992:GMT

[KCV92] David W. Krumme, George Cybenko, and K. N. Venkataraman. Gossiping in minimal time. SIAM Journal on Computing, 21(1):111–139, February 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Krishnamoorthy:1979:NDN

[KD79] M. S. Krishnamoorthy and Narsingh Deo. Node-deletion NP-complete problems. SIAM Journal on Computing, 8(4): 619–625, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Koren:2000:PMM

[KDA00] Gilad Koren, Emanuel Dar, and Amihood Amir. The power of migration in multiprocessor scheduling of realtime systems. SIAM Journal on Computing, 30(2):511–527, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32624.

Kedem:1985:OAA

[Ked85] Zvi M. Kedem. Optimal allocation of area for single-chip computations. SIAM Journal on Computing, 14(3):730–743, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Keil:1985:DPS

[Kei85] J. Mark Keil. Decomposing a polygon into simpler components. SIAM Journal on Computing, 14(4):799–817, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kelner:2006:SPE

[Kel06] Jonathan A. Kelner. Spectral partitioning, eigenvalue bounds, and circle packings for graphs of bounded genus.

SIAM Journal on Computing, 35(4):882–902, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kennedy:1976:CTA

[Ken76] Ken Kennedy. A comparison of two algorithms for global data flow analysis. SIAM Journal on Computing, 5(1):158– 180, ???? 1976. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kintala:1980:RNR

[KF80] Chandra M. R. Kintala and Patrick C. Fischer. Refining nondeterminism in relativized polynomial-time bounded computations. SIAM Journal on Computing, 9(1): 46–53, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kirkpatrick:1983:CGG

[KH83] D. G. Kirkpatrick and P. Hell. On the complexity of general graph factor problems. SIAM Journal on Computing, 12(3): 601–609, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kjos-Hanssen:2005:LCS

[KHNS05] Bjorn Kjos-Hanssen, André Nies, and Frank Stephan. Lowness for the class of Schnorr random reals. SIAM Journal on Computing, 35(3):647–657, ???? 2005. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Khot:2006:RPG

[Kho06] Subhash Khot. Ruling out PTAS for graph min-bisection, dense k-subgraph, and bipartite clique. SIAM Journal on Computing, 36(4):1025–1071, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Katoh:1981:AFM

[KIM81] N. Katoh, T. Ibaraki, and H. Mine. An algorithm for finding K minimum spanning trees. SIAM Journal on Computing, 10(2):247–255, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kim:1997:CMG

[Kim97] Sam Myo Kim. Computational modeling for genetic splicing systems. SIAM Journal on Computing, 26(5):1284– 1309, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26389.

Kingston:1986:AHA

[Kin86] Jeffrey H. Kingston. Analysis of Henriksen’s algorithm for the simulation event set. SIAM Journal on Computing, 15(3): 887–902, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kirkpatrick:1983:OSP

[Kir83] David Kirkpatrick. Optimal search in planar subdivisions. SIAM Journal on Computing, 12(1):28–35, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kirkpatrick:1977:ARR

[KK77] David G. Kirkpatrick and Zvi M. Kedem. Addition requirements for rational functions. SIAM Journal on Computing, 6(1):188–199, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kirkpatrick:1985:AMT

[KK85] David G. Kirkpatrick and Maria M. Klawe. Alphabetic minimax trees. SIAM Journal on Computing, 14(3):514–526, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kawaguchi:1986:WCB

[KK86] Tsuyoshi Kawaguchi and Seiki Kyan. Worst case bound of an LRF schedule for the mean weighted flow-time problem. SIAM Journal on Computing, 15(4):1119–1129, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kloks:1998:LAM

[KK98] T. Kloks and D. Kratsch. Listing all minimal separators of a graph. SIAM Journal on Computing, 27(3):605–613, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27087.

Kimbrel:2000:NOP

[KK00] Tracy Kimbrel and Anna R.

Karlin. Near-optimal parallel prefetching and caching. SIAM Journal on Computing, 29(4):1051–1082, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32697.

Keidar:2002:VSG

[KK02] Idit Keidar and Roger Khazan.

A virtually synchronous group multicast algorithm for WANs: Formal approach. SIAM Journal on Computing, 32(1):78– 130, February 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37875.

Kumar:2006:FMR

[KK06] Amit Kumar and Jon Kleinberg. Fairness measures for resource allocation. SIAM Journal on Computing, 36(3):657– 680, ???? 2006. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

King:2011:TGN

[KK11] James King and Erik Krohn. Terrain guarding is NP-hard. SIAM Journal on Computing, 40(5):1316–1339, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i5/p1316_s1.

Katz:2004:LSF

[KKKP04] Michal Katz, Nir A. Katz, Amos Korman, and David Peleg. Labeling schemes for flow and connectivity. SIAM Journal on Computing, 34(1):23– 40, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        43391.

Karmarkar:1993:MCA

[KKL^{+}93] N. Karmarkar, R. Karp, R. Lipton, L. Lovász, and M. Luby. A Monte-Carlo algorithm for estimating the permanent. SIAM Journal on Computing, 22(2):284–293, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kortsarz:2004:HAV

[KKL04] Guy Kortsarz, Robert Krauthgamer, and James R. Lee. Hardness of approximation for vertexconnectivity network design problems. SIAM Journal on Computing, 33(3):704–720, June 2004. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41673.

Kakade:2009:PGA

[KKL09] Sham M. Kakade, Adam Tauman Kalai, and Katrina Ligett. Playing games with approximation algorithms. SIAM Journal on Computing, 39(3): 1088–1106, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kobler:2011:IGC

[KKLV11] Johannes Köbler, Sebastian Kuhnert, Bastian Laubner, and Oleg Verbitsky. Interval graphs: Canonical representations in logspace. SIAM Journal on Computing, 40(5): 1292–1315, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i5/p1292_s1.

Keane:1984:OPP

[KKM84] M. Keane, A. G. Konheim, and I. Meilijson. The organ pipe permutation. SIAM Journal on Computing, 13(3):531–540, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kempe:2011:EGH

[KKM^{+}11] Julia Kempe, Hirotada Kobayashi, Keiji Matsumoto, Ben Toner, and Thomas Vidick. Entangled games are hard to approximate. SIAM Journal on Computing, 40(3):848–877, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i3/
p848_s1.

Kilian:2000:RCP

[KKMO00] Joe Kilian, Eyal Kushilevitz, Silvio Micali, and Rafail Ostrovsky. Reducibility and completeness in private computations. SIAM Journal on Computing, 29(4):1189–1208, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32174.

Khot:2007:OIR

[KKMO07] Subhash Khot, Guy Kindler, Elchanan Mossel, and Ryan O’Donnell. Optimal inapproximability results for MAX-CUT and other 2-variable CSPs? SIAM Journal on Computing, 37(1):319–357, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karloff:2009:EDM

[KKMR09] Howard Karloff, Subhash Khot, Aranyak Mehta, and Yuval Rabani. On earthmover distance, metric labeling, and 0-extension. SIAM Journal on Computing, 39(2):371–387, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kalai:2008:ALH

[KKMS08] Adam Tauman Kalai, Adam R.

Klivans, Yishay Mansour, and Rocco A. Servedio. Agnostically learning halfspaces. SIAM Journal on Computing, 37(6):1777–1805, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaplan:2011:OCP

[KKMS11] Haim Kaplan, Matthew J.

Katz, Gila Morgenstern, and Micha Sharir. Optimal cover of points by disks in a simple polygon. SIAM Journal on Computing, 40(6):1647–1661, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i6/
p1647_s1.

Kamma:2015:CCC

[KKN15] Lior Kamma, Robert Krauthgamer, and Huy L. Nguy˜ ên. Cutting corners cheaply, or how to remove Steiner points. SIAM Journal on Computing, 44(4): 975–995, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karger:1993:FHP

[KKP93] David R. Karger, Daphne Koller, and Steven J. Phillips.

Finding the hidden path: Time bounds for all-pairs shortest paths. SIAM Journal on Computing, 22(6):1199–1217, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kuijpers:2007:FOL

[KKPV07] Bart Kuijpers, Gabriel Kuper, Jan Paredaens, and Luc Vandeurzen. First-order languages expressing constructible spatial database queries. SIAM Journal on Computing, 36(6): 1570–1599, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaufman:2004:TBT

[KKR04] Tali Kaufman, Michael Krivelevich, and Dana Ron. Tight bounds for testing bipartiteness in general graphs. SIAM Journal on Computing, 33(6): 1441–1483, December 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43642.

Kempe:2006:CLH

[KKR06] Julia Kempe, Alexei Kitaev, and Oded Regev. The complexity of the local Hamiltonian problem. SIAM Journal on Computing, 35(5):1070– 1097, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kolmogorov:2017:CGV

[KKR17] Vladimir Kolmogorov, Andrei Krokhin, and Michal Rolínek. The complexity of generalvalued CSPs. SIAM Journal on Computing, 46(3):1087– 1110, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khuller:2004:ADM

[KKW04] Samir Khuller, Yoo-Ah Kim, and Yung-Chun (Justin) Wan. Algorithms for data migration with cloning. SIAM Journal on Computing, 33(2):448–461, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42585.

Kieburtz:1972:CCR

[KL72] Richard B. Kieburtz and David Luckham. Compatibility and complexity of refinements of the resolution principle. SIAM Journal on Computing, 1(4): 313–332, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kearns:1993:LPM

[KL93] Michael Kearns and Ming Li.

Learning in the presence of malicious errors. SIAM Journal on Computing, 22(4):807–837, August 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khanna:2000:BDP

[KL00] Sanjeev Khanna and Vincenzo Liberatore. On broadcast disk paging. SIAM Journal on Computing, 29 (5):1683–1702, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34139.

Karger:2015:FAP

[KL15] David R. Karger and Matthew S. Levine. Fast augmenting paths by random sampling from residual graphs. SIAM Journal on Computing, 44(2):320–339, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klawe:1984:LEC

[Kla84] Maria Klawe. Limitations on explicit constructions of expanding graphs. SIAM Journal on Computing, 13(1):156–166, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klauck:2007:LBQ

[Kla07a]  Hartmut Klauck.  Lower

bounds for quantum communication complexity. SIAM Journal on Computing, 37(1):20– 46, ???? 2007. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klauck:2007:OWC

[Kla07b]  Hartmut Klauck.  One-

way communication complexity and the Nečiporuk lower bound on formula size. SIAM Journal on Computing, 37(2): 552–583, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kleitman:1974:NPE

[Kle74] D. J. Kleitman. A note on perfect elimination digraphs. SIAM Journal on Computing, 3(4):280–282, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klein:1996:EPA

[Kle96]  Philip N. Klein.  Effi-

cient parallel algorithms for chordal graphs. SIAM Journal on Computing, 25(4):797– 827, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
16688.

Klein:2008:LTA

[Kle08] Philip N. Klein. A linear-time approximation scheme for TSP in undirected planar graphs with edge-weights. SIAM Journal on Computing, 37(6): 1926–1952, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klip:1979:NAP

[Kli79] Dorothea A. Klip. New algorithms for polynomial multiplication. SIAM Journal on Computing, 8(3):326–343, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kutylowski:1990:RCC

[KLL90] Mirosław Kutyłowski, Maciej Liśkiewicz, and Krzysztof Loryś. Reversal complexity classes for alternating Turing machines. SIAM Journal on Computing, 19(2):207–221, April 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kerenidis:2015:LBI

[KLL^{+}15] Iordanis Kerenidis, Sophie Laplante, Virginie Lerays, Jérémie Roland, and David Xiao. Lower bounds on information complexity via zerocommunication protocols and applications. SIAM Journal on Computing, 44(5):1550–1572, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kwok:2017:ICI

[KLL17] Tsz Chiu Kwok, Lap Chi Lau, and Yin Tat Lee. Improved Cheeger’s inequality and analysis of local graph partitioning using vertex expansion and expansion profile. SIAM Journal on Computing, 46(3):890–910, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kesselman:2004:BOM

[KLM^{+}04] Alexander Kesselman, Zvi Lotker, Yishay Mansour, Boaz Patt-Shamir, Baruch Schieber, and Maxim Sviridenko. Buffer overflow management in QoS switches. SIAM Journal on Computing, 33(3):563–583, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39966.

Kapralov:2017:SPS

[KLM^{+}17] M. Kapralov, Y. T. Lee, C. N. Musco, C. P. Musco, and A. Sidford. Single pass spectral sparsification in dynamic streams. SIAM Journal on Computing, 46(1):456–477, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kasiviswanathan:2011:WCW

[KLN^{+}11] Shiva Prasad Kasiviswanathan, Homin K. Lee, Kobbi Nissim, Sofya Raskhodnikova, and Adam Smith. What can we learn privately? SIAM Journal on Computing, 40(3): 793–826, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i3/p793_s1.

Kobayashi:2015:SMM

[KLN15] Hirotada Kobayashi, François Le Gall, and Harumichi Nishimura. Stronger methods of making quantum interactive proofs perfectly complete. SIAM Journal on Computing, 44(2):243–289, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kobayashi:2019:GQA

[KLN19] Hirotada Kobayashi, François Le Gall, and Harumichi Nishimura. Generalized quantum Arthur–Merlin games. SIAM Journal on Computing, 48(3):865–902, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kedem:1996:PSP

[KLP96] Zvi M. Kedem, Gad M. Landau, and Krishna V. Palem. Parallel suffix-prefix-matching algorithm and applications. SIAM Journal on Computing, 25(5):998–1023, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19015.

Kochol:2003:CPG

[KLR03] Martin Kochol, Vadim Lozin, and Bert Randerath. The 3colorability problem on graphs with maximum degree four. SIAM Journal on Computing, 32(5):1128–1139, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41875.

Kushilevitz:2010:ITS

[KLR10] Eyal Kushilevitz, Yehuda Lindell, and Tal Rabin. Information-theoretically secure protocols and security under composition. SIAM Journal on Computing, 39(5):2090– 2112, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kane:2019:INB

[KLR19] Daniel Kane, Shachar Lovett, and Sankeerth Rao. The independence number of the Birkhoff polytope graph, and applications to maximally recoverable codes. SIAM Journal on Computing, 48(4):1425– 1435, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kwok:2021:SAM

[KLR21] Tsz Chiu Kwok, Lap Chi Lau, and Akshay Ramachandran. Spectral analysis of matrix scaling and operator scaling. SIAM Journal on Computing, 50(3):1034–1102, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kulkarni:2024:OEC

[KLS^{+}24] Janardhan Kulkarni, Yang P.

Liu, Ashwin Sah, Mehtaab S. Sawhney, and Jakub Tarnawski. Online edge coloring via tree recurrences and correlation decay. SIAM Journal on Computing, 53(1):87–110, February 2024. CODEN SMJ-CAT. ISSN 1095-7111.

Kayal:2017:ELB

[KLSS17] Neeraj Kayal, Nutan Limaye, Chandan Saha, and Srikanth Srinivasan. An exponential lower bound for homogeneous depth four arithmetic formulas. SIAM Journal on Computing, 46(1):307–335, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kao:2000:CML

[KLST00] Ming-Yang Kao, Tak-Wah Lam, Wing-Kin Sung, and Hing-Fung Ting. Cavity matchings, label compressions, and unrooted evolutionary trees. SIAM Journal on Computing, 30(2):602–624, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33227.

Kao:2001:DTM

[KLST01] Ming-Yang Kao, Tak-Wah Lam, Wing-Kin Sung, and Hing-Fung Ting. A decomposition theorem for maximum weight bipartite matchings. SIAM Journal on Computing, 31(1):18–26, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36120.

Konemann:2008:GSC

[KLSvZ08] Jochen Könemann, Stefano Leonardi, Guido Schäfer, and Stefan H. M. van Zwam. A group-strategyproof cost sharing mechanism for the Steiner forest game. SIAM Journal on Computing, 37(5):1319– 1341, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaufman:2010:BSB

[KLX10] Tali Kaufman, Simon Litsyn, and Ning Xie. Breaking the ϵ-soundness bound of the linearity test over GF(2). SIAM Journal on Computing, 39(5): 1988–2003, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kuck:1975:TBP

[KM75] D. J. Kuck and K. Maruyama.

Time bounds on the parallel evaluation of arithmetic expressions. SIAM Journal on Computing, 4(2):147–162, June 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kundu:1977:LTP

[KM77] Sukhamay Kundu and Jayadev Misra. A linear tree partitioning algorithm. SIAM Journal on Computing, 6(1):151– 154, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ko:1981:CAD

[KM81] Ker-I Ko and Daniel Moore. Completeness, approximation and density. SIAM Journal on Computing, 10(4):787–796, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Korte:1989:ILT

[KM89] Norbert Korte and Rolf H. Möhring. An incremental linear-time algorithm for recognizing interval graphs. SIAM Journal on Computing, 18(1): 68–81, February 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kushilevitz:1993:LDT

[KM93] Eyal Kushilevitz and Yishay Mansour. Learning decision trees using the Fourier spectrum. SIAM Journal on Computing, 22(6):1331–1348, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaufmann:1994:LTA

[KM94a] Michael Kaufmann and Kurt Mehlhorn. A linear-time algorithm for the homotopic routing problem in grid graphs. SIAM Journal on Computing, 23(2):227–246, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/16210.

Kim:1994:COL

[KM94b] Sam M. Kim and Robert Mc-Naughton. Computing the order of a locally testable automaton. SIAM Journal on Computing, 23(6):1193–1215, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
21698.

Kilpelainen:1995:OUT

[KM95] Pekka Kilpeläinen and Heikki Mannila. Ordered and unordered tree inclusion. SIAM Journal on Computing, 24(2): 340–356, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
21820.

Kannan:1996:ALN

[KM96] Sampath K. Kannan and Eugene W. Myers. An algorithm for locating nonoverlapping regions of maximum alignment score. SIAM Journal on Computing, 25(3):648–662, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26267.

Karger:1997:AMC

[KM97a] David R. Karger and Rajeev Motwani. An NC algorithm for minimum cuts. SIAM Journal on Computing, 26(1):255–272, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27308.

Karzanov:1997:PMS

[KM97b] Alexander V. Karzanov and S. Thomas McCormick. Polynomial methods for separable convex optimization in unimodular linear spaces with applications. SIAM Journal on Computing, 26(4):1245– 1275, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        26369.

Kushilevitz:1998:LBB

[KM98] Eyal Kushilevitz and Yishay Mansour. An Ω(D log(N/D)) lower bound for broadcast in radio networks. SIAM Journal on Computing, 27(3):702–712, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27910.

Kosaraju:1999:CLE

[KM99] S. Rao Kosaraju and Giovanni Manzini. Compression of low entropy strings with Lempel– Ziv algorithms. SIAM Journal on Computing, 29(3):893–911, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33110.

Kapoor:2000:ECV

[KM00] Sanjiv Kapoor and S. N. Maheshwari. Efficiently constructing the visibility graph of a simple polygon with obstacles. SIAM Journal on Computing, 30(3):847–871, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25359.

Kijima:2008:AAP

[KM08] S. Kijima and T. Matsui.

Approximation algorithm and perfect sampler for closed Jackson networks with single servers. SIAM Journal on Computing, 38(4):1484–1503, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khot:2013:HAS

[KM13] Subhash Khot and Dana Moshkovitz. N P-Hardness of approximately solving linear equations over reals. SIAM Journal on Computing, 42(3): 752–791, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kratsch:2006:CAR

[KMMS06] Dieter Kratsch, Ross M. Mc-Connell, Kurt Mehlhorn, and Jeremy P. Spinrad. Certifying algorithms for recognizing interval graphs and permutation graphs. SIAM Journal on Computing, 36(2):326–353, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaplan:2022:DPL

[KMMS22] Haim Kaplan, Yishay Mansour, Yossi Matias, and Uri Stemmer. Differentially private learning of geometric concepts. SIAM Journal on Computing, 51(4):952–974, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1406428.

Koenig:2009:NTA

[KMMT09] Sven Koenig, Joseph S. B. Mitchell, Apurva Mudgal, and Craig Tovey. A near-tight approximation algorithm for the robot localization problem. SIAM Journal on Computing, 39(2):461–490, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Komargodski:2022:OWF

[KMN^{+}22] Ilan Komargodski, Tal Moran, Moni Naor, Rafael Pass, Alon Rosen, and Eylon Yogev. Oneway functions and (im)perfect obfuscation. SIAM Journal on Computing, 51(6):1769–1795, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/15M1048549.

Knuth:1977:FPM

[KMP77] Donald E. Knuth, James H.

Morris, Jr., and Vaughan R. Pratt. Fast pattern matching in strings. SIAM Journal on Computing, 6(2):323– 350, June 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Errata, see [Ryt80]. See also [?, Ryt80, ?, ?, ?].

Koutis:2014:AOS

[KMP14] Ioannis Koutis, Gary L. Miller, and Richard Peng. Approaching optimality for solving SDD linear systems. SIAM Journal on Computing, 43(1):337–354, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klassen:2020:HEC

[KMP^{+}20] Joel Klassen, Milad Marvian, Stephen Piddock, Marios Ioannou, Itay Hen, and Barbara M. Terhal. Hardness and ease of curing the sign problem for two-local qubit Hamiltonians. SIAM Journal on Computing, 49(6):1332–1362, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kieronski:2014:TVF

[KMPHT14] Emanuel Kieroński, Jakub Michaliszyn, Ian Pratt-Hartmann, and Lidia Tendera. Twovariable first-order logic with equivalence closure. SIAM Journal on Computing, 43(3): 1012–1063, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kannan:1987:SPA

[KMR87] Ravindran Kannan, Gary Miller, and Larry Rudolph. Sublinear parallel algorithm for computing the greatest common divisor of two integers. SIAM Journal on Computing, 16(1):7–16, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karp:1988:DDS

[KMR88a] Richard M. Karp, Rajeev Motwani, and Prabhakar Raghavan. Deferred data structuring. SIAM Journal on Computing, 17(5):883–902, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kong:1988:DRR

[KMR88b] T. Y. Kong, David M. Mount, and A. W. Roscoe. The decomposition of a rectangle into rectangles of minimal perimeter. SIAM Journal on Computing, 17(6):1215–1231, December 1988. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Kothari:2022:ARJ

[KMR22] Pravesh K. Kothari, Raghu Meka, and Prasad Raghavendra. Approximating rectangles by juntas and weakly exponential lower bounds for LP relaxations of CSPs. SIAM Journal on Computing, 51(2): STOC17–305–STOC17–332, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1152966.

Kaplan:2018:SDT

[KMRS18] Haim Kaplan, Wolfgang Mulzer, Liam Roditty, and Paul Seiferth. Spanners for directed transmission graphs. SIAM Journal on Computing, 47(4): 1585–1609, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kushilevitz:1998:LBR

[KMRZ98] Eyal Kushilevitz, Yishay Mansour, Michael O. Rabin, and David Zuckerman. Lower bounds for randomized mutual exclusion. SIAM Journal on Computing, 27(6): 1550–1563, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26513.

Kaltofen:1983:GCP

[KMS83] Erich Kaltofen, David R. Musser, and B. David Saunders. A generalized class of polynomials that are hard to factor. SIAM Journal on Computing, 12(3):473–483, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khot:2018:MTB

[KMS18] Subhash Khot, Dor Minzer, and Muli Safra. On monotonicity testing and Boolean isoperimetric-type theorems. SIAM Journal on Computing, 47(6):2238–2276, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Knessl:1987:AEC

[KMST87] C. Knessl, B. J. Matkowsky, Z. Schuss, and C. Tier. Asymptotic expansions for a closed multiple access system. SIAM Journal on Computing, 16(2): 378–398, April 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khanna:1998:SVC

[KMSV98] Sanjeev Khanna, Rajeev Motwani, Madhu Sudan, and Umesh Vazirani. On syntactic versus computational views of approximability. SIAM Journal on Computing, 28 (1):164–191, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28661.

Karnin:2013:DIT

[KMSV13] Zohar S. Karnin, Partha Mukhopadhyay, Amir Shpilka, and Ilya Volkovich. Deterministic identity testing of depth-4 multilinear circuits with bounded top fan-in. SIAM Journal on Computing, 42(6): 2114–2131, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaibel:2005:SAD

[KMSZ05] Volker Kaibel, Rafael Mechtel, Micha Sharir, and Günter M. Ziegler. The simplex algorithm in dimension three. SIAM Journal on Computing, 34(2):475–497, April 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43497.

Kenyon-Mathieu:1991:MSD

[KMV91] Claire M. Kenyon-Mathieu and Jeffrey Scott Vitter. The maximum size of dynamic data structures. SIAM Journal on Computing, 20(5):807–823, October 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khuller:1992:PEP

[KMV92] Samir Khuller, Stephen G. Mitchell, and Vijay V. Vazirani. Processor efficient parallel algorithms for the two disjoint paths problem and for finding a Kuratowski homeomorph. SIAM Journal on Computing, 21(3):486–506, June 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karp:1975:NOS

[KMW75] R. M. Karp, A. C. McKellar, and C. K. Wong. Near-optimal solutions to a 2-dimensional placement problem. SIAM Journal on Computing, 4(3): 271–286, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kirsch:2010:MRH

[KMW10] Adam Kirsch, Michael Mitzenmacher, and Udi Wieder. More robust hashing: Cuckoo hashing with a stash. SIAM Journal on Computing, 39(4):1543– 1561, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Korach:1987:ODC

[KMZ87] E. Korach, S. Moran, and S. Zaks. The optimality of distributive constructions of minimum weight and degree restricted spanning trees in a complete network of processors. SIAM Journal on Computing, 16(2):231–236, April 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Korula:2018:OSW

[KMZ18] Nitish Korula, Vahab Mirrokni, and Morteza Zadimoghaddam. Online submodular welfare maximization: Greedy beats 1/2 in random order. SIAM Journal on Computing, 47(3):1056–1086, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kapur:1985:KBC

[KN85] Deepak Kapur and Paliath Narendran. The Knuth– Bendix completion procedure and Thue systems. SIAM Journal on Computing, 14(4): 1052–1072, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kortsarz:2005:ANC

[KN05] Guy Kortsarz and Zeev Nutov. Approximating k-node connected subgraphs via critical graphs. SIAM Journal on Computing, 35(1):247–257, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khot:2008:LEM

[KN08] Subhash Khot and Assaf Naor.

Linear equations modulo 2 and the L_{1} diameter of convex bodies. SIAM Journal on Computing, 38(4):1448–1463, ????

2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Knight:1988:SOA

[Kni88] William J. Knight. Search in an ordered array having variable probe cost. SIAM Journal on Computing, 17(6):1203– 1214, December 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karger:1999:FCC

[KNP99] David R. Karger, Noam Nisan, and Michal Parnas. Fast connected components algorithms for the EREW PRAM. SIAM Journal on Computing, 28(3):1021–1034, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25247.

Krishnaswamy:2024:CBY

[KNPS24] Ravishankar Krishnaswamy, Viswanath Nagarajan, Kirk Pruhs, and Clifford Stein. Cluster before you hallucinate: Node-capacitated network design and energy efficient routing. SIAM Journal on Computing, 53(3):588–623, May 2024. CODEN SMJCAT. ISSN 1095-7111.

Knudsen:2019:LHA

[Knu19] Mathias Bæk Tejs Knudsen. Linear hashing is awesome. SIAM Journal on Computing, 48(2):736–741, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ko:1989:RPT

[Ko89] Ker-I Ko. Relativized polynomial time hierarchies having exactly k levels. SIAM Journal on Computing, 18(2):392– 408, April 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ko:1991:CLM

[Ko91] Ker-I Ko. On the complexity of learning minimum timebounded Turing machines. SIAM Journal on Computing, 20(5):962–986, October 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Koch:1992:ISN

[Koc92] Richard Koch. Increasing the size of a network by a constant factor can increase performance by more than a constant factor. SIAM Journal on Computing, 21(5):801–823, October 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kolmogorov:2018:CAL

[Kol18] Vladimir Kolmogorov. Commutativity in the algorithmic Lovász local lemma. SIAM Journal on Computing, 47(6): 2029–2056, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Konheim:1972:NM

[Kon72] Alan G. Konheim. A note on merging. SIAM Journal on Computing, 1(2):161–166, ???? 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Konheim:1975:ESQ

[Kon75] Alan G. Konheim. An elementary solution of the queuing system G/G/1. SIAM Journal on Computing, 4(4):540– 545, ???? 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kushilevitz:1998:LSP

[KOR98] Eyal Kushilevitz, Rafail Ostrovsky, and Adi Rosén. Logspace polynomial end-to-end communication. SIAM Journal on Computing, 27(6): 1531–1549, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29676.

Kushilevitz:2000:ESA

[KOR00] Eyal Kushilevitz, Rafail Ostrovsky, and Yuval Rabani. Efficient search for approximate nearest neighbor in high dimensional spaces. SIAM Journal on Computing, 30(2):457– 474, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34717.

Kosaraju:1975:SRC

[Kos75] S. Rao Kosaraju. Speed of recognition of context-free languages by array automata. SIAM Journal on Computing, 4(3):331–340, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaplan:2000:SCP

[KOT00] Haim Kaplan, Chris Okasaki, and Robert E. Tarjan. Simple confluently persistent catenable lists. SIAM Journal on Computing, 30(3):965–977, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33943.

Kou:1977:PCC

[Kou77] Lawrence T. Kou. Polynomial complete consecutive information retrieval problems. SIAM Journal on Computing, 6(1):67–75, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kou:1982:MVH

[Kou82] Lawrence T. Kou. Minimum variance Huffman codes. SIAM Journal on Computing, 11(1): 138–148, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Krokhin:2023:TAP

[KOWZ23] Andrei Krokhin, Jakub Oprsal, Marcin Wrochna, and Stanislav Zivný. Topology and adjunction in promise constraint satisfaction. SIAM Journal on Computing, 52(1):38–79, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1378223.

Kozik:2009:CVM

[Koz09] Marcin Kozik. A 2EXPTIME complete varietal membership problem. SIAM Journal on Computing, 38(6):2443–2467, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kozik:2021:SCU

[Koz21] Marcin Kozik. Solving CSPs using weak local consistency. SIAM Journal on Computing, 50(4):1263–1286, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karp:1982:LCC

[KP82] Richard M. Karp and Christos H. Papadimitriou. On linear characterizations of combinatorial optimization problems. SIAM Journal on Computing, 11(4):620–632, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kanellakis:1985:CDC

[KP85] Paris C. Kanellakis and Christos H. Papadimitriou. The complexity of distributed concurrency control. SIAM Journal on Computing, 14(1):52– 74, February 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kortsarz:1998:GLD

[KP98] Guy Kortsarz and David Peleg. Generating low-degree 2-spanners. SIAM Journal on Computing, 27(5):1438– 1456, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26875.

Koutsoupias:2000:BCA

[KP00a] Elias Koutsoupias and Christos H. Papadimitriou. Beyond competitive analysis. SIAM Journal on Computing, 30 (1):300–317, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29954.

Kutten:2000:TFL

[KP00b] Shay Kutten and David Peleg. Tight fault locality. SIAM Journal on Computing, 30(1):247–268, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31910.

Kowalski:2004:TDB

[KP04] Dariusz R. Kowalski and Andrzej Pelc. Time of deterministic broadcasting in radio networks with local knowledge. SIAM Journal on Computing, 33(4):870–891, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41933.

Kalyanasundaram:2005:FTS

[KP05] Bala Kalyanasundaram and Kirk R. Pruhs. Fault-tolerant scheduling. SIAM Journal on Computing, 34(3):697–719, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26179.

Karlin:2000:MP

[KPR00] Anna R. Karlin, Steven J. Phillips, and Prabhakar Raghavan. Markov paging. SIAM Journal on Computing, 30(3): 906–922, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        26804.

Kintali:2013:RAF

[KPR^{+}13] Shiva Kintali, Laura J. Poplawski, Rajmohan Rajaraman, Ravi Sundaram, and Shang-Hua Teng. Reducibility among fractional stability problems. SIAM Journal on Computing, 42(6):2063–2113, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kirschenhofer:1994:DST

[KPS94] Peter Kirschenhofer, Helmut Prodinger, and Wojciech Szpankowski. Digital search trees again revisited: The internal path length perspective. SIAM Journal on Computing, 23(3):598–616, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18936.

Kale:2013:NTE

[KPS13] Satyen Kale, Yuval Peres, and C. Seshadhri. Noise tolerance of expanders and sublinear expansion reconstruction. SIAM Journal on Computing, 42(1): 305–323, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klein:1994:FAA

[KPST94] Philip Klein, Serge Plotkin, Clifford Stein, and Éva Tardos. Faster approximation algorithms for the unit capacity concurrent flow problem with applications to routing and finding sparse cuts. SIAM Journal on Computing, 23(3):466–487, June 1994.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24117.

Khot:2014:APF

[KPV14] Subhash A. Khot, Preyas Popat, and Nisheeth K. Vishnoi. Almost polynomial factor hardness for closest vector problem with preprocessing. SIAM Journal on Computing, 43(3):1184–1205, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kanj:2010:SLS

[KPX10] Iyad A. Kanj, Ljubomir Perković, and Ge Xia. On spanners and lightweight spanners of geometric graphs. SIAM Journal on Computing, 39(6):2132–2161, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kao:1999:OBA

[KQT99] Ming-Yang Kao, Junfeng Qi, and Lei Tan. Optimal bidding algorithms against cheating in multiple-object auctions. SIAM Journal on Computing, 28(3):955–969, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30537.

Konheim:1978:FCQ

[KR78] A. G. Konheim and M. Reiser.

Finite capacity queuing systems with applications in computer modeling. SIAM Journal on Computing, 7(2):210– 229, May 1978. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klein:1988:PTA

[KR88] Philip N. Klein and John H.

Reif. Parallel time O(log n) acceptance of deterministic CFLs on an exclusive-write P-RAM. SIAM Journal on Computing, 17(3):463–485, June 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kapoor:1995:AEA

[KR95] Sanjiv Kapoor and H. Ramesh.

Algorithms for enumerating all spanning trees of undirected and weighted graphs. SIAM Journal on Computing, 24(2):247–265, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25030.

Konemann:2002:MDI

[KR02] J. Könemann and R. Ravi.

A matter of degree: Improved approximation algorithms for degree-bounded minimum spanning trees. SIAM Journal on Computing, 31(6):1783–1793, December 2002. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36917.

Konemann:2005:PDM

[KR05] J. Könemann and R. Ravi. Primal-dual meets local search: Approximating MSTs with nonuniform degree bounds. SIAM Journal on Computing, 34(3):763–773, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41804.

Kaufman:2006:TPG

[KR06] Tali Kaufman and Dana Ron. Testing polynomials over general fields. SIAM Journal on Computing, 36(3):779–802, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kolliopoulos:2007:NLT

[KR07] Stavros G. Kolliopoulos and Satish Rao. A nearly lineartime approximation scheme for the Euclidean k-median problem. SIAM Journal on Computing, 37(3):757–782, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Krauthgamer:2009:ILB

[KR09] Robert Krauthgamer and Yuval Rabani. Improved lower bounds for embeddings into L_{1}. SIAM Journal on Computing, 38(6):2487–2498, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kush:2023:TDF

[KR23] Deepanshu Kush and Benjamin Rossman. Tree-depth and the formula complexity of subgraph isomorphism. SIAM Journal on Computing, 52(1):273–325, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1372925.

Krentel:1990:FVL

[Kre90] Mark W. Krentel. On finding and verifying locally optimal solutions. SIAM Journal on Computing, 19(4):742–749, August 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karloff:1994:LBR

[KRR94] Howard Karloff, Yuval Rabani, and Yiftach Ravid. Lower bounds for randomized kserver and motion-planning algorithms. SIAM Journal on Computing, 23(2):293–312, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22483.

Kaplan:2010:LTC

[KRS10a] Haim Kaplan, Natan Rubin, and Micha Sharir. Line transversals of convex polyhedra in R^{3}. SIAM Journal on Computing, 39(7):3283– 3310, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kenyon:2010:LDM

[KRS10b] Claire Kenyon, Yuval Rabani, and Alistair Sinclair. Low distortion maps between point sets. SIAM Journal on Computing, 39(4):1617–1636, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karnin:2012:EDR

[KRS12] Zohar S. Karnin, Yuval Rabani, and Amir Shpilka. Explicit dimension reduction and its applications. SIAM Journal on Computing, 41(1):219–249, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v41/i1/
p219_s1.

Kabanets:2020:SSF

[KRS20] Valentine Kabanets, Sofya Raskhodnikova, and Chaitanya Swamy. Special section on the fifty-eighth annual IEEE Symposium on Foundations of Computer Science (FOCS 2017). SIAM Journal on Computing, 49(4):FOCS17– i–FOCS17–i, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaplan:2008:ECO

[KRSV08] Haim Kaplan, Natan Rubin, Micha Sharir, and Elad Verbin. Efficient colored orthogonal range counting. SIAM Journal on Computing, 38(3):982– 1011, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kleinberg:2000:ABB

[KRT00] Jon Kleinberg, Yuval Rabani, and Éva Tardos. Allocating bandwidth for bursty connections. SIAM Journal on Computing, 30(1):191–217, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32914.

Kohayakawa:2003:OAC

[KRT03] Y. Kohayakawa, V. Rödl, and L. Thoma. An optimal algorithm for checking regularity. SIAM Journal on Computing, 32(5):1210–1235, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40822.

Kempe:2010:UGE

[KRT10] Julia Kempe, Oded Regev, and Ben Toner. Unique games with entangled provers are easy. SIAM Journal on Computing, 39(7):3207–3229, ???? 2010.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Komargodski:2017:IAC

[KRT17] Ilan Komargodski, Ran Raz, and Avishay Tal. Improved average-case lower bounds for De Morgan formula size: Matching worst-case lower bound. SIAM Journal on Computing, 46(1):37–57, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kesselheim:2018:PBD

[KRTV18] Thomas Kesselheim, Klaus Radke, Andreas Tönnis, and Berthold Vöcking. Primal beats dual on online packing LPs in the random-order model. SIAM Journal on Computing, 47(5):1939–1964, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Krumme:1992:FGH

[Kru92] David W. Krumme. Fast gossiping for the hypercube. SIAM Journal on Computing, 21(2):365–380, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khuller:1995:AME

[KRY95] Samir Khuller, Balaji Raghavachari, and Neal Young. Approximating the minimum equivalent digraph. SIAM Journal on Computing, 24(4):859–872, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25668.

Khuller:1996:LDS

[KRY96] Samir Khuller, Balaji Raghavachari, and Neal Young. Low-degree spanning trees of small weight. SIAM Journal on Computing, 25(2):355–368, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26458.

Kopparty:2023:ILD

[KRZSW23] Swastik Kopparty, Noga Ron-Zewi, Shubhangi Saraf, and Mary Wootters. Improved list decoding of folded Reed– Solomon and multiplicity codes. SIAM Journal on Computing, 52(3):794–840, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1370215.

Kafura:1977:TSM

[KS77] D. G. Kafura and V. Y. Shen.

Task scheduling on a multiprocessor system with independent memories. SIAM Journal on Computing, 6(1):167–187, March 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ko:1985:CSC

[KS85] Ker-I Ko and Uwe Schöning. On circuit-size complexity and the low hierarchy in NP. SIAM Journal on Computing, 14(1): 41–51, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kirkpatrick:1986:UPC

[KS86] David G. Kirkpatrick and Raimund Seidel. The ultimate planar convex hull algorithm? SIAM Journal on Computing, 15(1):287–299, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khuller:1991:EPA

[KS91] Samir Khuller and Baruch Schieber. Efficient parallel algorithms for testing kconnectivity and finding disjoint s-t paths in graphs. SIAM Journal on Computing, 20(2): 352–375, April 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kao:1993:LPNb

[KS93] Ming-Yang Kao and Gregory E. Shannon. Linearprocessor NC algorithms for planar directed graphs. II. directed spanning trees. SIAM Journal on Computing, 22(3): 460–481, June 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kim:1994:PRF

[KS94] Myong-Hi Kim and Scott Sutherland. Polynomial root-finding algorithms and branched covers. SIAM Journal on Computing, 23(2):415– 436, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
20158.

Koren:1995:OLS

[KS95] Gilad Koren and Dennis Shasha. Dover: An optimal on-line scheduling algorithm for overloaded uniprocessor real-time systems. SIAM Journal on Computing, 24(2):318– 339, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23688.

Kaplan:1996:PBC

[KS96a] Haim Kaplan and Ron Shamir.

Pathwidth, bandwidth, and completion problems to proper interval graphs with small cliques. SIAM Journal on Computing, 25(3):540–561, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25814.

Kapoor:1996:NTE

[KS96b] Sanjiv Kapoor and Michiel Smid. New techniques for exact and approximate dynamic closest-point problems. SIAM Journal on Computing, 25(4):775–796, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25945.

Katz:1997:EBA

[KS97] Matthew J. Katz and Micha Sharir. An expander-based approach to geometric optimization. SIAM Journal on Computing, 26(5):1384–1408, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26864.

Kavvadias:1998:ISP

[KS98] Dimitris Kavvadias and Martha Sideri. The inverse satisfiability problem. SIAM Journal on Computing, 28(1):152– 163, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28511.

Knessl:2000:ABH

[KS00] Charles Knessl and Wojciech Szpankowski. Asymptotic behavior of the height in a digital search tree and the longest phrase of the Lempel– Ziv scheme. SIAM Journal on Computing, 30(3):923–964, June 2000. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35681.

Kucera:2001:RRE

[KS01] Antonín Kucera and T. Slaman. Randomness and recursive enumerability. SIAM Journal on Computing, 31 (1):199–211, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35744.

Kolliopoulos:2002:AAS

[KS02] Stavros G. Kolliopoulos and Clifford Stein. Approximation algorithms for single-source unsplittable flow. SIAM Journal on Computing, 31(3):919– 946, June 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35531.

Koltun:2003:DEV

[KS03a] Vladlen Koltun and Micha Sharir. 3-dimensional Euclidean Voronoi diagrams of lines with a fixed number of orientations. SIAM Journal on Computing, 32(3):616–642, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40838.

Koltun:2003:PTO

[KS03b] Vladlen Koltun and Micha Sharir. The partition technique for overlays of envelopes. SIAM Journal on Computing, 32(4):841–863, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40700.

Koltun:2005:CSC

[KS05] Vladlen Koltun and Micha Sharir. Curve-sensitive cuttings. SIAM Journal on Computing, 34(4):863–878, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43568.

Kratsch:2006:BNA

[KS06] Dieter Kratsch and Jeremy Spinrad. Between O(nm) and O(nα). SIAM Journal on Computing, 36(2):310–325, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kale:2011:ETB

[KS11] Satyen Kale and C. Seshadhri. An expansion tester for bounded degree graphs. SIAM Journal on Computing, 40(3):709–720, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i3/p709_s1.

Kopparty:2013:LLD

[KS13] Swastik Kopparty and Shubhangi Saraf. Local listdecoding and testing of random linear codes from high error. SIAM Journal on Computing, 42(3):1302–1326, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kumar:2015:LDR

[KS15] Mrinal Kumar and Shubhangi Saraf. The limits of depth reduction for arithmetic formulas: It’s all about the top fanin. SIAM Journal on Computing, 44(6):1601–1625, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Khot:2017:HCC

[KS17a] Subhash Khot and Rishi Saket. Hardness of coloring 2-colorable 12-uniform hypergraphs with 2^{(log n)^{Ω(1)}} colors. SIAM Journal on Computing, 46(1):235–271, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kumar:2017:PHD

[KS17b] Mrinal Kumar and Shubhangi Saraf. On the power of homogeneous depth 4 arithmetic circuits. SIAM Journal on Computing, 46(1):336–387, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kumar:2019:CPD

[KS19] Mrinal Kumar and Ramprasad Saptharishi. The computational power of depth five arithmetic circuits. SIAM Journal on Computing, 48(1): 144–180, ???? 2019. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kozma:2020:SHD

[KS20] László Kozma and Thatchaphol Saranurak. Smooth heaps and a dual view of selfadjusting data structures. SIAM Journal on Computing, 49(5):STOC18–45–STOC18– 93, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klauck:2007:QCS

[KŠdW07] Hartmut Klauck, Robert Špalek, and Ronald de Wolf. Quantum and classical strong direct product theorems and optimal time-space tradeoffs. SIAM Journal on Computing, 36(5):1472–1493, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kirousis:1993:PCC

[KSS93] Lefteris M. Kirousis, Maria Serna, and Paul Spirakis. Parallel complexity of the connected subgraph problem. SIAM Journal on Computing, 22(3):573–586, June 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kleinberg:2008:NFD

[KSS08] Jon Kleinberg, Mark Sandler, and Aleksandrs Slivkins. Network failure detection and graph connectivity. SIAM Journal on Computing, 38(4): 1330–1346, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kumar:2020:RWF

[KSS20] Akash Kumar, C. Seshadhri, and Andrew Stolman. Random walks and forbidden minors I: An n^{1/2+o(1)}-query onesided tester for minor closed properties on bounded degree graphs. SIAM Journal on Computing, 52(6):FOCS18– 216–FOCS18–249, December 2020. CODEN SMJCAT. ISSN 1095-7111.

Kratochvil:1993:OMO

[KST93] Jan Kratochvíl, Petr Savický, and Zsolt Tuza. One more occurrence of variables makes satisfiability jump from trivial to NP-Complete. SIAM Journal on Computing, 22(1): 203–210, February 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kaplan:1999:FSA

[KST99a] Haim Kaplan, Ron Shamir, and Robert E. Tarjan. A faster and simpler algorithm for sorting signed permutations by reversals. SIAM Journal on Computing, 29(3):880–892, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33420.

Kaplan:1999:TPC

[KST99b] Haim Kaplan, Ron Shamir, and Robert E. Tarjan. Tractability of parameterized completion problems on chordal, strongly chordal, and proper interval graphs. SIAM Journal on Computing, 28(5):1906– 1922, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
30304.

Kannan:2008:SMG

[KSV08] Ravindran Kannan, Hadi Salmasian, and Santosh Vempala. The spectral method for general mixture models. SIAM Journal on Computing, 38(3): 1141–1156, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kung:1977:FAP

[KT77] H. T. Kung and D. M. Tong. Fast algorithms for partial fraction decomposition. SIAM Journal on Computing, 6(3): 582–593, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kucera:1988:ITU

[KT88] Luděk Kučera and Věra Trnková. Isomorphism testing of unary algebras. SIAM Journal on Computing, 17(4): 673–686, August 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kobler:1994:CRA

[KT94] Johannes Köbler and Thomas Thierauf. Complexity-restricted advice functions. SIAM Journal on Computing, 23(2):261– 275, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
20035.

Kirousis:1996:LG

[KT96] Lefteris M. Kirousis and Dimitris M. Thilikos. The linkage of a graph. SIAM Journal on Computing, 25(3):626–647, June 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25570.

Kopczynski:2015:RGS

[KT15] Eryk Kopczyński and Tony Tan. Regular graphs and the spectra of two-variable logic with counting. SIAM Journal on Computing, 44(3):786–818, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kellerer:1999:ANR

[KTW99] Hans Kellerer, T. Tautenhahn, and G. Woeginger. Approximability and nonapproximability results for minimizing total flow time on a single machine. SIAM Journal on Computing, 28(4):1155–1166, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30577.

Kaplan:2013:SHS

[KTZ13] Haim Kaplan, Robert E. Tarjan, and Uri Zwick. Soft heaps simplified. SIAM Journal on Computing, 42(4):1660– 1673, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kolmogorov:2015:PLP

[KTZ15] Vladimir Kolmogorov, Johan Thapper, and Stanislav Zivný. The power of linear programming for general-valued CSPs. SIAM Journal on Computing, 44(1):1–36, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Karkkainen:1999:THD

[KU99] Juha Kärkkäinen and Esko Ukkonen. Two- and higherdimensional pattern matching in optimal expected time. SIAM Journal on Computing, 29(2):571–589, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27587.

Kedlaya:2011:FPF

[KU11] Kiran S. Kedlaya and Christopher Umans. Fast polynomial factorization and modular composition. SIAM Journal on Computing, 40(6):1767– 1802, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i6/p1767_s1.

Kummer:1996:KCI

[Kum96] Martin Kummer. Kolmogorov complexity and instance complexity of recursively enumerable sets. SIAM Journal on Computing, 25(6):1123–1143, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26878.

Kundu:1974:EGT

[Kun74] Sukhamay Kundu. Existence of graphs with three spanning trees and given degree sequence. SIAM Journal on Computing, 3(4):296–298, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kunde:1981:NLS

[Kun81] Manfred Kunde. Nonpreemptive LP-scheduling on homogeneous multiprocessor systems. SIAM Journal on Computing, 10(1):151–173, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kundu:1993:ADS

[Kun93] Sukhamay Kundu. An O(n) algorithm for determining the subregion-tree representation of a rectangular dissection. SIAM Journal on Computing, 22(1):79–101, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kuperberg:2005:STQ

[Kup05] Greg Kuperberg. A subexponentialtime quantum algorithm for the dihedral hidden subgroup problem. SIAM Journal on Computing, 35(1):170–188, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kurtz:1985:SSR

[Kur85] Stuart A. Kurtz. Sparse sets in NP − P: relativizations. SIAM Journal on Computing, 14(1): 113–119, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kurtz:1987:NRP

[Kur87] Stuart A. Kurtz. A note on randomized polynomial time. SIAM Journal on Computing, 16(5):852–853, October 1987.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kutylowski:1991:TCB

[Kut91] Mirosław Kutyłowski. Time complexity of Boolean functions on CREW PRAMs. SIAM Journal on Computing, 20(5):824–833, October 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kutz:2002:LBL

[Kut02] Martin Kutz. Lower bounds for Lucas chains. SIAM Journal on Computing, 31(6): 1896–1908, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37925.

Katz:1986:DSR

[KV86] Martin David Katz and Dennis J. Volper. Data structures for retrieval on square grids. SIAM Journal on Computing, 15(4):919–931, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Krishnan:1998:OPP

[KV98] P. Krishnan and Jeffrey Scott Vitter. Optimal prediction for prefetching in the worst case. SIAM Journal on Computing, 27(6):1617–1636, December 1998. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26181.

Kirschmer:2010:AEI

[KV10] Markus Kirschmer and John Voight. Algorithmic enumeration of ideal classes for quaternion orders. SIAM Journal on Computing, 39(5):1714– 1747, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See corrigendum [KV12].

Kirschmer:2012:CAE

[KV12] Markus Kirschmer and John Voight. Corrigendum: Algorithmic enumeration of ideal classes for quaternion orders. SIAM Journal on Computing, 41(3):714, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [KV10].

Katajainen:1988:FST

[KvLP88] Jyrki Katajainen, Jan van Leeuwen, and Martti Penttonen. Fast simulation of Turing machines by random access machines. SIAM Journal on Computing, 17(1):77– 88, February 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klivans:2002:GNS

[KvM02] Adam R. Klivans and Dieter van Melkebeek. Graph nonisomorphism has subexponential size proofs unless the polynomial-time hierarchy collapses. SIAM Journal on Computing, 31(5):1501–1526, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38965.

Karpinski:1993:VDU

[KW93] Marek Karpiński and Thorsten Werther. VC dimension and uniform learnability of sparse polynomials and rational functions. SIAM Journal on Computing, 22(6):1276–1285, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kannan:1994:IEH

[KW94] Sampath K. Kannan and Tandy J. Warnow. Inferring evolutionary history from DNA sequences. SIAM Journal on Computing, 23(4):713– 737, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
22217.

Kannan:1995:TRP

[KW95] Sampath K. Kannan and Tandy J. Warnow. Tree reconstruction from partial orders. SIAM Journal on Computing, 24(3):511–519, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25219.

Kannan:1997:FAC

[KW97] Sampath Kannan and Tandy Warnow. A fast algorithm for the computation and enumeration of perfect phylogenies. SIAM Journal on Computing, 26(6):1749–1763, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27906.

Kobler:1998:NCC

[KW98] Johannes Köbler and Osamu Watanabe. New collapse consequences of NP having small circuits. SIAM Journal on Computing, 28 (1):311–324, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29620.

Kao:2000:LTA

[KW00] Ming-Yang Kao and Jie Wang. Linear-time approximation algorithms for computing numerical summation with provably small errors. SIAM Journal on Computing, 29 (5):1568–1576, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34159.

Kannan:1998:CLC

[KWY98] Sampath Kannan, Tandy Warnow, and Shibu Yooseph.

Computing the local consensus of trees. SIAM Journal on Computing, 27(6): 1695–1724, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28764.

Kedlaya:2009:LDC

[KY09] Kiran S. Kedlaya and Sergey Yekhanin. Locally decodable codes from nice subsets of finite fields and prime factors of Mersenne numbers. SIAM Journal on Computing, 38(5): 1952–1969, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kawachi:2010:QHF

[KY10] Akinori Kawachi and Tomoyuki Yamakami. Quantum hardcore functions by complexity-theoretical quantum list decoding. SIAM Journal on Computing, 39(7):2941– 2969, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Klein:2015:NID

[KY15] Philip Klein and Neal E.

Young. On the number of iterations for Dantzig–Wolfe optimization and packingcovering approximation algorithms. SIAM Journal on Computing, 44(4):1154–1172, ???? 2015. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Kamidoi:2007:DAF

[KYN07] Yoko Kamidoi, Noriyoshi Yoshida, and Hiroshi Nagamochi. A deterministic algorithm for finding all minimum k-way cuts. SIAM Journal on Computing, 36(5):1329–1341, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kamp:2007:DEB

[KZ07] Jesse Kamp and David Zuckerman. Deterministic extractors for bit-fixing sources and exposure-resilient cryptography. SIAM Journal on Computing, 36(5):1231–1247, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Kyng:2020:HRS

[KZ20] Rasmus Kyng and Peng Zhang. Hardness results for structured linear systems. SIAM Journal on Computing, 49(4):FOCS17–280–FOCS17– 349, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

LaPoutre:2000:MEC

[La 00] Han La Poutré. Maintenance of 2- and 3-edge-connected components of graphs II. SIAM Journal on Computing, 29(5):1521–1549, October 2000. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25777.

Liu:2001:TSA

[LAB01] Pangfeng Liu, William Aiello, and Sandeep Bhatt. Tree search on an atomic model for message passing. SIAM Journal on Computing, 31(1):67– 85, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
35807.

Ladner:1977:CCP

[Lad77] Richard E. Ladner. The computational complexity of provability in systems of modal propositional logic. SIAM Journal on Computing, 6(3): 467–480, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ladner:1989:PSC

[Lad89] Richard E. Ladner. Polynomial space counting problems. SIAM Journal on Computing, 18(6):1087–1097, December 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lagarias:1985:CCS

[Lag85] J. C. Lagarias. The computational complexity of simultaneous Diophantine approximation problems. SIAM Journal on Computing, 14(1):196–209, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Landau:1985:FPA

[Lan85] Susan Landau. Factoring polynomials over algebraic number fields. SIAM Journal on Computing, 14(1):184–195, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [Lan91].

Landau:1991:EFP

[Lan91] Susan Landau. Erratum: “Factoring polynomials over algebraic number fields” [SIAM J. Comput. 14 (1985), no. 1, 184– 195, MR 86d:11102]. SIAM Journal on Computing, 20(5): 998, October 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Lan85].

Landau:1992:SNR

[Lan92] Susan Landau. Simplification of nested radicals. SIAM Journal on Computing, 21(1):85– 110, February 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Landsberg:2014:NLB

[Lan14] J. M. Landsberg. New lower bounds for the rank of matrix multiplication. SIAM Journal on Computing, 43(1):144–149, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Larson:1977:MGC

[Lar77] Richard G. Larson. Minimizing garbage collection as a function of region size. SIAM Journal on Computing, 6(4):663– 668, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Larmore:1987:HRO

[Lar87] Lawrence L. Larmore. Height restricted optimal binary trees. SIAM Journal on Computing, 16(6):1115–1123, December 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Larmore:1989:MDC

[Lar89] Lawrence L. Larmore. Minimum delay codes. SIAM Journal on Computing, 18(1):82– 94, February 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Larsen:2014:RSG

[Lar14] Kasper Green Larsen. On range searching in the group model and combinatorial discrepancy. SIAM Journal on Computing, 43(2):673–686, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lopez-Alt:2017:MFH

[LATV17] Adriana López-Alt, Eran Tromer, and Vinod Vaikuntanathan. Multikey fully homomorphic encryption and applications. SIAM Journal on Computing, 46(6):1827–1892, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lauer:1983:GAC

[Lau83] Markus Lauer. Generalized p-adic constructions. SIAM Journal on Computing, 12(2): 395–410, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lavenberg:1973:QAM

[Lav73] Stephen S. Lavenberg. Queueing analysis of a multiprogrammed computer system having a multilevel storage hierarchy. SIAM Journal on Computing, 2(4):232–252, December 1973. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Also published in/as: IBM Res.R. RJ1133, Nov.1972.

Lucier:2017:EGC

[LB17] Brendan Lucier and Allan Borodin. Equilibria of greedy combinatorial auctions. SIAM Journal on Computing, 46(2): 620–660, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Labahn:1989:MPF

[LC89] George Labahn and Stan Cabay. Matrix Padé fractions and their computation. SIAM Journal on Computing, 18(4):

639–657, August 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:1990:TDP

[LC90] Keqin Li and Kam-Hoi Cheng.

On three-dimensional packing. SIAM Journal on Computing, 19(5):847–867, October 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lee:2005:PLK

[LC05] Gyung-Ok Lee and Kwang-Moo Choe. A powerful LL(k) covering transformation. SIAM Journal on Computing, 35(2):359–377, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Labahn:1990:IBH

[LCC90] George Labahn, Dong Koo Choi, and Stan Cabay. The inverses of block Hankel and block Toeplitz matrices. SIAM Journal on Computing, 19(1): 98–123, February 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lee:1981:GVD

[LD81] D. T. Lee and R. L. Drysdale, III. Generalization of Voronoĭ diagrams in the plane. SIAM Journal on Computing, 10(1): 73–87, February 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lin:1999:RMC

[LDHX99] Guo-Hui Lin, Ding-Zhu Du, Xiao-Dong Hu, and Guoliang Xue. On rearrangeability of multirate Clos networks. SIAM Journal on Computing, 28(4):1225–1231, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31392.

Legrand-Duchesne:2024:PCU

[LDRT24] Clément Legrand-Duchesne, Ashutosh Rai, and Martin Tancer. Parameterized complexity of untangling knots. SIAM Journal on Computing, 53(2):431–479, March 2024. CODEN SMJCAT. ISSN 1095-7111.

Lehmann:1982:PT

[Leh82] Daniel J. Lehmann. On primality tests. SIAM Journal on Computing, 11(2):374–375, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lempel:1975:MFT

[Lem75]  Abraham Lempel.  Ma-

trix factorization over GF (2) and trace-orthogonal bases of GF (2^{n}). SIAM Journal on Computing, 4(2):175–186, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lenstra:1987:FMP

[Len87] Arjen K. Lenstra. Factoring multivariate polynomials over algebraic number fields. SIAM Journal on Computing, 16(3): 591–598, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leung:1998:SEA

[Leu98] Hing Leung. Separating exponentially ambiguous finite automata from polynomially ambiguous finite automata. SIAM Journal on Computing, 27(4):1073–1082, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25209.

Levy:1976:LAL

[Lev76] Leon S. Levy. Local adjunct languages and regular sets. SIAM Journal on Computing, 5(2):305–308, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Levin:1986:ACC

[Lev86] Leonid A. Levin. Average case complete problems. SIAM Journal on Computing, 15(1): 285–286, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lewis:1972:P

[Lew72] Philip M. Lewis, II. Preface.

SIAM Journal on Computing, 1(1):v–vi, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lin:2000:PLR

[LFS00] Wei-Liang Lin, Amir H. Farrahi, and M. Sarrafzadeh. On the power of logic resynthesis. SIAM Journal on Computing, 29(4):1257–1289, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33548.

Lo:2000:AUS

[LH00] Wai-Kau Lo and Vassos Hadzilacos. All of us are smarter than any of us: Nondeterministic wait-free hierarchies are not robust. SIAM Journal on Computing, 30(3):689–728, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33576.

Li:1986:RUA

[Li86] Li Wu Li. Ranking and unranking of AVL-trees. SIAM Journal on Computing, 15(4): 1025–1035, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:1989:DPE

[Li89] Shuo-Yen Robert Li. Dynamic programming by exchangeability. SIAM Journal on Computing, 18(3):463–472, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:2020:SMT

[Li20] Shi Li. Scheduling to minimize total weighted completion time via time-indexed linear programming relaxations. SIAM Journal on Computing, 49(4):FOCS17–409–FOCS17– 440, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lichtenstein:1982:PFT

[Lic82] David Lichtenstein. Planar formulae and their uses. SIAM Journal on Computing, 11(2): 329–343, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lickteig:1987:CCD

[Lic87] Thomas Lickteig. The computational complexity of division in quadratic extension fields. SIAM Journal on Computing, 16(2):278–311, April 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lickteig:1989:LBC

[Lic89] Thomas Lickteig. A lower bound on the complexity of division in finite extension fields and inversion in quadratic alternative algebras. SIAM Journal on Computing, 18(2):

209–215, April 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lien:1976:TPG

[Lie76] Y. Edmund Lien. Termination properties of generalized Petri nets. SIAM Journal on Computing, 5(2):251–265, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Liebehenschel:2003:LGG

[Lie03]  Jens Liebehenschel.  Lexi-

cographical generation of a generalized Dyck language. SIAM Journal on Computing, 32(4):880–903, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39449.

Linial:1984:ITB

[Lin84] Nathan Linial. The informationtheoretic bound is good for merging. SIAM Journal on Computing, 13(4):795–801, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Linial:1992:LDG

[Lin92] Nathan Linial. Locality in distributed graph algorithms. SIAM Journal on Computing, 21(1):193–201, February 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lipton:1978:PCH

[Lip78] Richard J. Lipton. Polynomials with 0-1 coefficients that are hard to evaluate. SIAM Journal on Computing, 7(1): 61–69, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Liu:1972:ASS

[Liu72] C. L. Liu. Analysis and synthesis of sorting algorithms. SIAM Journal on Computing, 1(4): 290–304, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Liu:1981:SPM

[Liu81] Ken Chih Liu. On string pattern matching: a new model with a polynomial time algorithm. SIAM Journal on Computing, 10(1):118–140, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Liu:2022:NOP

[Liu22] Chih-Hung Liu. Nearly optimal planar k nearest neighbors queries under general distance functions. SIAM Journal on Computing, 51(3):723–765, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1388371.

Louchard:1997:DSM

[LKS97] G. Louchard, Claire Kenyon, and R. Schott. Data structures’ maxima. SIAM Journal on Computing, 26(4):1006– 1042, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
19660.

Lynch:1978:SPR

[LL78] Nancy Lynch and Richard J. Lipton. On structure preserving reductions. SIAM Journal on Computing, 7(2):119– 126, ???? 1978. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Liskiewicz:1990:FST

[LL90] Maciej Liśkiewicz and Krzysztof Loryś. Fast simulations of time-bounded one-tape Turing machines by space-bounded ones. SIAM Journal on Computing, 19(3):511–521, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Loui:1992:OLS

[LL92] Michael C. Loui and David R. Luginbuhl. Optimal on-line simulations of tree machines by random access machines. SIAM Journal on Computing, 21(5):959–971, October 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lennerstad:1995:OET

[LL95] Håkan Lennerstad and Lars Lundberg. An optimal execution time estimate of static versus dynamic allocation in multiprocessor systems. SIAM Journal on Computing, 24 (4):751–764, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24057.

Levene:1999:NHE

[LL99] Mark Levene and George Loizou. Navigation in hypertext is easy only sometimes. SIAM Journal on Computing, 29(3):728–760, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28273.

Lennerstad:2000:OCF

[LL00a] Håkan Lennerstad and Lars Lundberg. Optimal combinatorial functions comparing multiprocess allocation performance in multiprocessor systems. SIAM Journal on Computing, 29(6):1816–1838, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
29439.

Lennerstad:2000:OWC

[LL00b] Håkan Lennerstad and Lars Lundberg. Optimal worst case formulas comparing cache memory associativity. SIAM Journal on Computing, 30(3):

872–905, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        34916.

Lawler:1980:GAM

[LLK80] E. L. Lawler, J. K. Lenstra, and A. H. G. Rinnooy Kan. Generating all maximal independent sets: NP-hardness and polynomial-time algorithms. SIAM Journal on Computing, 9(3):558–565, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leighton:2001:NAA

[LLRS01] Tom Leighton, Chi-Jen Lu, Satish Rao, and Aravind Srinivasan. New algorithmic aspects of the local lemma with applications to routing and partitioning. SIAM Journal on Computing, 31(2):626–641, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37976.

Ladner:1984:APS

[LLS84] Richard E. Ladner, Richard J. Lipton, and Larry J. Stockmeyer. Alternating pushdown and stack automata. SIAM Journal on Computing, 13(1): 135–155, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:1992:PQ

[LLV92] Ming Li, Luc Longpré, and Paul Vitányi. The power of the queue. SIAM Journal on Computing, 21(4):697–712, August 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Levy:1982:CDT

[LM82] M. R. Levy and T. S. E.

Maibaum. Continuous data types. SIAM Journal on Computing, 11(2):201–216, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lew:1994:JDE

[LM94a] William Lew and Hosam M.

Mahmoud. The joint distribution of elastic buckets in multiway search trees. SIAM Journal on Computing, 23 (5):1050–1074, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23023.

Lutz:1994:MSD

[LM94b] Jack H. Lutz and Elvira Mayordomo. Measure, stochasticity, and the density of hard languages. SIAM Journal on Computing, 23(4):762– 779, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23749.

Leighton:1999:TBS

[LM99] Tom Leighton and Yuan Ma. Tight bounds on the size of fault-tolerant merging and sorting networks with destructive faults. SIAM Journal on Computing, 29(1):258– 273, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        30529.

Laplante:2008:LBR

[LM08a] Sophie Laplante and Frédéric Magniez. Lower bounds for randomized and quantum query complexity using Kolmogorov arguments. SIAM Journal on Computing, 38(1): 46–62, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lutz:2008:DPS

[LM08b] Jack H. Lutz and Elvira Mayordomo. Dimensions of points in self-similar fractals. SIAM Journal on Computing, 38(3): 1080–1112, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Loffler:2012:TSS

[LM12] Maarten Löffler and Wolfgang Mulzer. Triangulating the square and squaring the triangle: Quadtrees and Delaunay triangulations are equivalent. SIAM Journal on Computing, 41(4):941–974, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lovett:2015:CDM

[LM15] Shachar Lovett and Raghu Meka. Constructive discrepancy minimization by walking on the edges. SIAM Journal on Computing, 44(5):1573– 1582, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leoncini:1999:PCN

[LMM99] Mauro Leoncini, Giovanni Manzini, and Luciano Margara. Parallel complexity of numerically accurate linear system solvers. SIAM Journal on Computing, 28(6): 2030–2058, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32711.

Lau:2014:SSF

[LMOT14] Lap Chi Lau, Tal Malkin, Ryan O’Donnell, and Luca Trevisan. Special section on the Fifty-First Annual IEEE Symposium on Foundations of Computer Science (FOCS 2010). SIAM Journal on Computing, 43(1):255, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Laber:2002:BSN

[LMP02] Eduardo S. Laber, Ruy L. Milidiú, and Artur A. Pessoa. On binary searching with nonuniform costs. SIAM Journal on Computing, 31 (4):1022–1047, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38199.

Lonati:2015:OPL

[LMPP15] Violetta Lonati, Dino Mandrioli, Federica Panella, and Matteo Pradella. Operator precedence languages: Their automata-theoretic and logic characterization. SIAM Journal on Computing, 44(4):1026– 1088, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lueker:1990:LPT

[LMR90] George S. Lueker, Nimrod Megiddo, and Vijaya Ramachandran. Linear programming with two variables per inequality in poly-log time. SIAM Journal on Computing, 19(6):1000–1010, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Landau:1998:ISC

[LMS98a] Gad M. Landau, Eugene W. Myers, and Jeanette P. Schmidt. Incremental string comparison. SIAM Journal on Computing, 27(2):557–582, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26481.

Leighton:1998:FTS

[LMS98b] F. Thomson Leighton, Bruce M.

Maggs, and Ramesh K. Sitaraman. On the fault tolerance of some popular boundeddegree networks. SIAM Journal on Computing, 27(5):1303– 1333, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
25516.

Lokshtanov:2018:SSP

[LMS18] Daniel Lokshtanov, Dániel Marx, and Saket Saurabh. Slightly superexponential parameterized problems. SIAM Journal on Computing, 47(3): 675–702, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leonardi:2001:LRC

[LMSPR01] Stefano Leonardi, Alberto Marchetti-Spaccamela, Alessio Presciutti, and Adi Rosén. On-line randomized call control revisited. SIAM Journal on Computing, 31(1):86– 112, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.

                                        org/sam-bin/dbq/article/
                                        34670.

Luczak:2006:PAD

[ŁN06] Tomasz Łuczak and Jaroslav Nešetřil. A probabilistic approach to the dichotomy problem. SIAM Journal on Computing, 36(3):835–843, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leighton:1992:DTE

[LNRS92] F. T. Leighton, Mark J. Newman, Abhiram G. Ranade, and Eric J. Schwabe. Dynamic tree embeddings in butterflies and hypercubes. SIAM Journal on Computing, 21(4):639–654, August 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lin:2010:GAI

[LNRW10] Guolong Lin, Chandrashekhar Nagarajan, Rajmohan Rajaraman, and David P. Williamson. A general approach for incremental approximation and hierarchical clustering. SIAM Journal on Computing, 39(8): 3633–3669, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lau:2009:SND

[LNSS09] Lap Chi Lau, Joseph (Seffi) Naor, Mohammad R. Salavatipour, and Mohit Singh. Survivable network design with degree or order constraints. SIAM Journal on Computing, 39(3):1062– 1087, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:2019:AMN

[LNW19] Yi Li, Huy L. Nguy˜ ên, and David P. Woodruff. On approximating matrix norms in data streams. SIAM Journal on Computing, 48(6):1643– 1697, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lohrey:2006:WPM

[Loh06] Markus Lohrey. Word problems and membership problems on compressed words. SIAM Journal on Computing, 35(5):1210–1240, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Long:1985:RSO

[Lon85] Timothy J. Long. On restricting the size of oracles compared with restricting access to oracles. SIAM Journal on Computing, 14(3):585–597, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [Lon88].

Long:1988:ERS

[Lon88] Timothy J. Long. Erratum:

“On restricting the size of oracles compared with restricting access to oracles” [SIAM J.

Comput. 14 (1985), no. 3, 585– 597, MR 87h:68053]. SIAM Journal on Computing, 17(3): 628, June 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Lon85].

Loos:1983:CRZ

[Loo83] Rüdiger Loos. Computing rational zeros of integral polynomials by p-adic expansion. SIAM Journal on Computing, 12(2):286–293, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Loui:1983:ODE

[Lou83] Michael C. Loui. Optimal dynamic embedding of trees into arrays. SIAM Journal on Computing, 12(3):463–472, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lovett:2021:SMM

[Lov21] Shachar Lovett. Sparse MDS matrices over small fields: a proof of the GM-MDS conjecture. SIAM Journal on Computing, 50(4):1248–1262, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lee:1977:LPP

[LP77] D. T. Lee and F. P. Preparata. Location of a point in a planar subdivision and its applications. SIAM Journal on Computing, 6(3):594–606, September 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leiserson:1983:OPR

[LP83a] Charles E. Leiserson and Ron Y. Pinter. Optimal placement for river routing. SIAM Journal on Computing, 12(3): 447–462, August 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lifschitz:1983:WMP

[LP83b] V. Lifschitz and B. Pittel’. The worst and the most probable performance of a class of setcovering algorithms. SIAM Journal on Computing, 12(2): 329–346, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Larmore:1994:FAO

[LP94] Lawrence L. Larmore and Teresa M. Przytycka. A fast algorithm for optimum height-limited alphabetic binary trees. SIAM Journal on Computing, 23(6):1283– 1312, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23116.

Larmore:1995:CHT

[LP95] Lawrence L. Larmore and Teresa M. Przytycka. Constructing Huffman trees in parallel. SIAM Journal on Computing, 24(6):1163–1169, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23324.

Leighton:1998:HSN

[LP98] Tom Leighton and C. Greg Plaxton. Hypercubic sorting networks. SIAM Journal on Computing, 27(1):1–47, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26840.

Leung:2003:MTC

[LP03] Joseph Y.-T. Leung and Michael Pinedo. Minimizing total completion time on parallel machines with deadline constraints. SIAM Journal on Computing, 32(5):1370– 1388, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
40638.

Lovett:2013:SLB

[LP13] Shachar Lovett and Ely Porat. A space lower bound for dynamic approximate membership data structures. SIAM Journal on Computing, 42(6): 2182–2196, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lokshtanov:2017:FPT

[LPPS17] Daniel Lokshtanov, Marcin Pilipczuk, Michal Pilipczuk, and Saket Saurabh. Fixedparameter tractable canonization and isomorphism test for graphs of bounded treewidth. SIAM Journal on Computing, 46(1):161–189, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lin:2020:TRN

[LPS20] Huijia Lin, Rafael Pass, and Pratik Soni. Two-round and non-interactive concurrent non-malleable commitments from time-lock puzzles. SIAM Journal on Computing, 49(4):FOCS17–196–FOCS17– 279, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lotker:2005:MWS

[LPSPP05] Zvi Lotker, Boaz Patt-Shamir, Elan Pavlov, and David Peleg. Minimum-weight spanning tree construction in O(log log n) communication rounds. SIAM Journal on Computing, 35(1):120–131, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lotker:2004:NSR

[LPSR04] Zvi Lotker, Boaz Patt-Shamir, and Adi Rosén. New stability results for adversarial queuing. SIAM Journal on Computing, 33(2):286–303, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41330.

Lotker:2009:DAM

[LPSR09] Zvi Lotker, Boaz Patt-Shamir, and Adi Rosén. Distributed approximate matching. SIAM Journal on Computing, 39(2): 445–460, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Liotta:1998:RPQ

[LPT98] Giuseppe Liotta, Franco P. Preparata, and Roberto Tamassia. Robust proximity queries: An illustration of degreedriven algorithm design. SIAM Journal on Computing, 28(3): 864–889, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        30536.

Leighton:1986:TDC

[LR86] Frank Thomson Leighton and Arnold L. Rosenberg. Threedimensional circuit layouts. SIAM Journal on Computing, 15(3):793–813, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lagarias:1988:UEP

[LR88a] Jeffrey C. Lagarias and James A. Reeds. Unique extrapolation of polynomial recurrences. SIAM Journal on Computing, 17(2): 342–362, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Luby:1988:HCP

[LR88b] Michael Luby and Charles Rackoff. How to construct pseudorandom permutations from pseudorandom functions. SIAM Journal on Computing, 17(2):373–386, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Li:1989:SDC

[LR89] Zhiyuan Li and Edward M.

Reingold. Solution of a divideand-conquer maximin recurrence. SIAM Journal on Computing, 18(6):1188–1200, December 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Liskiewicz:1996:SAS

[LR96] Maciej Liśkiewicz and Rüdiger Reischuk. The sublogarithmic alternating space world. SIAM Journal on Computing, 25(4):828–861, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25244.

Levey:2021:AMS

[LR21] Elaine Levey and Thomas Rothvo. A (1+epsilon)- approximation for makespan scheduling with precedence constraints using LP hierarchies. SIAM Journal on Computing, 50(3):STOC16–201– STOC16–217, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:2017:CSI

[LRR17] Yuan Li, Alexander Razborov, and Benjamin Rossman. On the AC^{0} complexity of subgraph isomorphism. SIAM Journal on Computing, 46(3): 936–971, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Luby:2001:MCA

[LRS01] Michael Luby, Dana Randall, and Alistair Sinclair. Markov chain algorithms for planar lattice structures. SIAM Journal on Computing, 31 (1):167–192, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36035.

Lenzen:2017:ECO

[LRS17] Christoph Lenzen, Joel Rybicki, and Jukka Suomela. Efficient counting with optimal resilience. SIAM Journal on Computing, 46(4):1473–1500, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lund:1999:CLA

[LRWY99] Carsten Lund, Nick Reingold, Jeffery Westbrook, and Dicky Yan. Competitive on-line algorithms for distributed data management. SIAM Journal on Computing, 28(3):1086– 1111, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28782.

Lavenberg:1975:DCI

[LS75] S. S. Lavenberg and G. S.

Shedler. Derivation of confidence intervals for work rate estimators in a closed queueing network. SIAM Journal on Computing, 4(2):108–124, June 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lam:1977:WCA

[LS77] Shui Lam and Ravi Sethi.

Worst case analysis of two scheduling algorithms. SIAM Journal on Computing, 6(3): 518–536, September 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lai:1984:PSM

[LS84] Ten Hwang Lai and Sartaj Sahni. Preemptive scheduling of a multiprocessor system with memories to minimize maximum lateness. SIAM Journal on Computing, 13(4): 690–704, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Levine:1990:NBT

[LS90] Robert Y. Levine and Alan T. Sherman. A note on Bennett’s time-space tradeoff for reversible computation. SIAM Journal on Computing, 19(4): 673–677, August 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lou:1993:OAM

[LS93] Ruey-Der Lou and Majid Sarrafzadeh. An optimal algorithm for the maximum threechain problem. SIAM Journal on Computing, 22(5):976–993, October 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lakshman:1995:SPI

[LS95] Y. N. Lakshman and B. David Saunders. Sparse polynomial interpolation in nonstandard bases. SIAM Journal on Computing, 24(2):387–397, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23778.

Liu:1997:SSV

[LS97] Zhen Liu and Eric Sanlaville. Stochastic scheduling with variable profile and precedence constraints. SIAM Journal on Computing, 26 (1):173–187, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/21894.

Li:1998:LAS

[LS98] Rongheng Li and Lijie Shi.

An on-line algorithm for some uniform processor scheduling. SIAM Journal on Computing, 27(2):414–422, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27969.

Lefmann:2002:DPT

[LS02] Hanno Lefmann and Niels Schmitt. A deterministic polynomial-time algorithm for Heilbronn’s problem in three dimensions. SIAM Journal on Computing, 31(6):1926– 1947, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39511.

Lipman:2010:ADC

[LS10] Julia Lipman and Quentin F.

Stout. Analysis of delays caused by local synchronization. SIAM Journal on Computing, 39(8):3860–3884, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v39/i8/
                                        p3860_s1.

Lau:2013:AAB

[LS13] Lap Chi Lau and Mohit Singh. Additive approximation for bounded degree survivable network design. SIAM Journal on Computing, 42(6):2217–2242, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:2016:AMP

[LS16] Shi Li and Ola Svensson. Approximating k-median via pseudo-approximation. SIAM Journal on Computing, 45(2): 530–547, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lee:2018:CLS

[LS18] Yin Tat Lee and He Sun. Constructing linear-sized spectral sparsification in almost-linear time. SIAM Journal on Computing, 47(6):2315–2336, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lenstra:2019:TIL

[LS19] Hendrik W. Lenstra, Jr. and Alice Silverberg. Testing isomorphism of lattices over CMorders. SIAM Journal on Computing, 48(4):1300–1334, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leme:2022:CSI

[LS22] Renato Paes Leme and Jon Schneider. Contextual search via intrinsic volumes. SIAM Journal on Computing, 51 (4):1096–1125, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1385718.

Lindenbaum:2005:PAT

[LSH05] Michael Lindenbaum, Hanan Samet, and Gisli R. Hjaltason. A probabilistic analysis of triebased sorting of large collections of line segments in spatial databases. SIAM Journal on Computing, 35(1):22– 58, ???? 2005. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Linz:2013:CTP

[LSS13] Simone Linz, Katherine St.

John, and Charles Semple. Counting trees in a phylogenetic network is #P-complete. SIAM Journal on Computing, 42(4):1768–1776, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Limaye:2021:FDS

[LSS^{+}21] Nutan Limaye, Karteek Sreenivasaiah, Srikanth Srinivasan, Utkarsh Tripathi, and S. Venkitesh. A fixed-depth size-hierarchy theorem for AC^{0}[⊕] via the coin problem. SIAM Journal on Computing, 50(4):1461– 1499, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Levy:2008:CMS

[LSSV08] Jordi Levy, Manfred Schmidt-Schauß, and Mateu Villaret. The complexity of monadic second-order unification. SIAM Journal on Computing, 38(3):1113–1140, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Louchard:1999:APG

[LST99] Guy Louchard, Wojciech Szpankowski, and Jing Tang. Average profile of the generalized digital search tree and the generalized Lempel–Ziv algorithm. SIAM Journal on Computing, 28(3):904–934, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30181.

Lynch:2007:OBS

[LSV07] Nancy Lynch, Roberto Segala, and Frits Vaandrager. Observing branching structure through probabilistic contexts.

SIAM Journal on Computing, 37(4):977–1013, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lee:2013:MMP

[LSV13] Jon Lee, Maxim Sviridenko, and Jan Vondrák. Matroid matching: The power of local search. SIAM Journal on Computing, 42(1):357–379, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lee:1990:MCC

[LSW90] D. T. Lee, M. Sarrafzadeh, and Y. F. Wu. Minimum cuts for circular-arc graphs. SIAM Journal on Computing, 19(6):1041–1050, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Levin:2011:IRG

[LSZ11] Hagay Levin, Michael Schapira, and Aviv Zohar. Interdomain routing and games. SIAM Journal on Computing, 40(6): 1892–1912, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i6/p1892_s1.

Lipton:1980:APS

[LT80] Richard J. Lipton and Robert Endre Tarjan. Applications of a planar separator theorem. SIAM Journal on Computing, 9(3):615–627, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Linial:1982:CCP

[LT82] Nathan Linial and Michael Tarsi. The counterfeit coin problem revisited. SIAM Journal on Computing, 11(3):409– 415, ???? 1982. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Luo:1991:CCS

[LT91] Zhi-Quan Luo and John N. Tsitsiklis. On the communication complexity of solving a polynomial equation. SIAM Journal on Computing, 20(5): 936–950, October 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:2018:CPP

[LTW18] Fu Li, Iddo Tzameret, and Zhengyu Wang. Characterizing propositional proofs as noncommutative formulas. SIAM Journal on Computing, 47(4):1424–1462, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lee:2010:SSF

[LU10] James R. Lee and Chris Umans. Special section on foundations of computer science. SIAM Journal on Computing, 39(6):2397, ???? 2010.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lu:2014:LTC

[Lu14]  Hsueh-I Lu.  Linear-time

compression of bounded-genus graphs into information-theoretically optimal number of bits. SIAM Journal on Computing, 43(2): 477–496, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lubiw:1981:SNC

[Lub81]  Anna Lubiw.  Some NP-

complete problems similar to graph isomorphism. SIAM Journal on Computing, 10(1): 11–21, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Luby:1986:SPA

[Lub86] Michael Luby. A simple parallel algorithm for the maximal independent set problem. SIAM Journal on Computing, 15(4):1036–1053, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lubiw:1987:DLO

[Lub87] Anna Lubiw. Doubly lexical orderings of matrices. SIAM Journal on Computing, 16(5): 854–879, October 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lucas:1990:PDS

[Luc90] Joan M. Lucas. Postorder disjoint set union is linear. SIAM Journal on Computing, 19(5): 868–882, October 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lueker:1981:OPG

[Lue81]  George S. Lueker.  Opti-

mization problems on graphs with independent random edge weights. SIAM Journal on Computing, 10(2):338–351, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lutz:1990:CMC

[Lut90] Jack H. Lutz. Category and measure in complexity classes. SIAM Journal on Computing, 19(6):1100–1131, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lutz:1993:POC

[Lut93] Jack H. Lutz. A pseudorandom oracle characterization of bpp. SIAM Journal on Computing, 22(5):1075–1086, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lutz:1995:WHP

[Lut95] Jack H. Lutz. Weakly hard problems. SIAM Journal on Computing, 24(6):1170–1189, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
24970.

Lutz:2003:DCC

[Lut03] Jack H. Lutz. Dimension in complexity classes. SIAM Journal on Computing, 32 (5):1236–1259, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41772.

Li:1991:LSC

[LV91] Ming Li and Paul M. B.

Vitányi. Learning simple concepts under simple distributions. SIAM Journal on Computing, 20(5):911–935, October 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:1995:NAF

[LV95] Ming Li and Paul Vitányi. A new approach to formal language theory by Kolmogorov complexity. SIAM Journal on Computing, 24(2):398–410, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24485.

Lovasz:2006:HRC

[LV06] László Lovász and Santosh Vempala. Hit-and-run from a corner. SIAM Journal on Computing, 35(4):985–1005, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leake:2022:CCM

[LV22a] Jonathan Leake and Nisheeth K. Vishnoi. On the computability of continuous maximum entropy distributions with applications. SIAM Journal on Computing, 51(5):1451–1505, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1440864.

Lee:2022:GWP

[LV22b] Yin Tat Lee and Santosh Vempala. Geodesic walks in polytopes. SIAM Journal on Computing, 51(2):STOC17– 400–STOC17–488, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1145999.

Leung:1984:SVB

[LVW84] Joseph Y.-T. Leung, Oliver Vornberger, and James D. Witthoff. On some variants of the bandwidth minimization problem. SIAM Journal on Computing, 13(3):650–667, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lee:1980:VDM

[LW80] D. T. Lee and C. K. Wong. Voronoĭ diagrams in L_{1} (L_{∞}) metrics with 2-dimensional storage applications. SIAM Journal on Computing, 9(1): 200–211, February 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Langenhop:1986:PRF

[LW86] Carl E. Langenhop and William E. Wright. Probabilities related to father-son distances in binary search trees. SIAM Journal on Computing, 15(2):520–530, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lengauer:1988:ESC

[LW88a] Thomas Lengauer and Egon Wanke. Efficient solution of connectivity problems on hierarchically defined graphs. SIAM Journal on Computing, 17(6):1063–1080, December 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Long:1988:DLH

[LW88b] Douglas L. Long and Avi Wigderson. The discrete logarithm hides O(log n) bits. SIAM Journal on Computing, 17(2):363–372, April 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

LaPoutre:1998:DCB

[LW98a] Johannes A. La Poutré and Jeffery Westbrook. Dynamic 2-connectivity with backtracking. SIAM Journal on Computing, 28(1):10–26, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27258.

Lai:1998:AHB

[LW98b] Tony W. Lai and Derick Wood. Adaptive heuristics for binary search trees and constant linkage cost. SIAM Journal on Computing, 27(6): 1564–1591, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25032.

Lin:2018:CBH

[LW18] Jiabao Lin and Hanpin Wang. The complexity of Boolean Holant problems with nonnegative weights. SIAM Journal on Computing, 47(3):798–828, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:2021:TBS

[LWW21] Yi Li, Ruosong Wang, and David P. Woodruff. Tight bounds for the subspace sketch problem with applications. SIAM Journal on Computing, 50(4):1287–1335, ???? 2021.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Larsen:2020:CLB

[LWY20] Kasper Green Larsen, Omri Weinstein, and Huacheng Yu. Crossing the logarithmic barrier for dynamic Boolean data structure lower bounds. SIAM Journal on Computing, 49(5):STOC18–323–STOC18– 367, ???? 2020. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Leung:1989:MSL

[LY89] Joseph Y.-T. Leung and Gilbert H. Young. Minimizing schedule length subject to minimum flow time. SIAM Journal on Computing, 18(2):314–326, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Li:2005:EAC

[LY05] Minming Li and Frances F.

Yao. An efficient algorithm for computing optimal discrete voltage schedules. SIAM Journal on Computing, 35(3):658– 671, ???? 2005. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lau:2013:EES

[LY13] Lap Chi Lau and Chun Kong Yung. Efficient edge splittingoff algorithms maintaining all-pairs edge-connectivities. SIAM Journal on Computing, 42(3):1185–1200, ???? 2013.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Lutz:2000:DWC

[LZ00] Jack H. Lutz and Yong Zhao. The density of weakly complete problems under adaptive reductions. SIAM Journal on Computing, 30(4):1197– 1210, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        32154.

Lau:2022:LSF

[LZ22a] Lap Chi Lau and Hong Zhou. A local search framework for experimental design. SIAM Journal on Computing, 51 (4):900–951, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1386542.

Lau:2022:SAN

[LZ22b] Lap Chi Lau and Hong Zhou. A spectral approach to network design. SIAM Journal on Computing, 51(4):1018–1064, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1330762.

Maass:1986:CNC

[Maa86] Wolfgang Maass. On the complexity of nonconvex covering.

SIAM Journal on Computing, 15(2):453–467, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Maass:1997:BCP

[Maa97] Wolfgang Maass. Bounds for the computational power and learning complexity of analog neural nets. SIAM Journal on Computing, 26(3):708–732, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25604.

MacKenzie:1997:RAL

[Mac97] Philip D. MacKenzie. The random adversary: a lowerbound technique for randomized parallel algorithms. SIAM Journal on Computing, 26(6): 1559–1580, December 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22403.

Macarie:1998:SED

[Mac98]  Ioan I. Macarie.  Space-

efficient deterministic simulation of probabilistic automata. SIAM Journal on Computing, 27(2):448–465, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25385.

Macarie:1999:SLP

[Mac99] Ioan I. Macarie. On the structure of logspace probabilistic complexity classes. SIAM Journal on Computing, 29(3):987–1007, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29833.

Mahadev:2020:CHE

[Mah20]  Urmila Mahadev.  Clas-

sical homomorphic encryption for quantum circuits. SIAM Journal on Computing, 52(6):FOCS18–189–FOCS18– 215, December 2020. CODEN SMJCAT. ISSN 1095-7111.

Mahadev:2022:CVQ

[Mah22] Urmila Mahadev. Classical verification of quantum computations. SIAM Journal on Computing, 51(4):1172–1229, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1371828.

Maier:1979:EMS

[Mai79] David Maier. An efficient method for storing ancestor information in trees. SIAM Journal on Computing, 8(4):599– 618, November 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mak:1997:PAH

[Mak97] Louis Mak. Parallelism always helps. SIAM Journal on Computing, 26(1):153–172, February 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26540.

Malewicz:2005:WOD

[Mal05] Grzegorz Malewicz. A workoptimal deterministic algorithm for the certified write-all problem with a nontrivial number of asynchronous processors. SIAM Journal on Computing, 34(4):993–1024, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42801.

Manber:1986:MDI

[Man86] Udi Manber. On maintaining dynamic information in a concurrent environment. SIAM Journal on Computing, 15(4): 1130–1142, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mansour:1995:RIA

[Man95] Yishay Mansour. Randomized interpolation and approximation of sparse polynomials. SIAM Journal on Computing, 24(2):357–368, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23929.

Manthey:2008:ARC

[Man08] Bodo Manthey. On approximating restricted cycle covers. SIAM Journal on Computing, 38(1):181–206, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mao:1993:TWC

[Mao93] Weizhen Mao. Tight worstcase performance bounds for next-k-fit bin packing. SIAM Journal on Computing, 22(1): 46–56, February 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Marcinkowski:1999:ATU

[Mar99] Jerzy Marcinkowski. Achilles, turtle, and undecidable boundedness problems for small DATALOG programs. SIAM Journal on Computing, 29 (1):231–257, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32214.

Marx:2008:CSP

[Mar08] Dániel Marx. Closest substring problems with small distances. SIAM Journal on Computing, 38(4):1382–1410, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Matousek:1991:ALL

[Mat91] Jiří Matoušek. Approximate levels in line arrangements. SIAM Journal on Computing, 20(2):222–227, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mayoh:1981:AGM

[May81] Brian H. Mayoh. Attribute grammars and mathematical semantics. SIAM Journal on Computing, 10(3):503– 518, August 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See also: report DAIMI PB-90, Computer Science Department, Aarhus University (August 1978).

Mayr:1984:AGP

[May84] Ernst W. Mayr. An algorithm for the general Petri net reachability problem. SIAM Journal on Computing, 13(3):441–460, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mazlack:1976:MSE

[Maz76] Lawrence J. Mazlack. Machine selection of elements in crossword puzzles: An application of computational linguistics. SIAM Journal on Computing, 5(1):51–72, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mitchell:1998:CFA

[MBCV98] Joseph S. B. Mitchell, Avrim Blum, Prasad Chalasani, and Santosh Vempala. A constantfactor approximation algorithm for the geometric k-MST problem in the plane. SIAM Journal on Computing, 28(3):771–781, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30329.

McKenzie:1987:PCA

[MC87] Pierre McKenzie and Stephen A. Cook. The parallel complexity of Abelian permutation group problems. SIAM Journal on Computing, 16(5):880– 909, October 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

McCreight:1985:PST

[McC85] Edward M. McCreight. Priority search trees. SIAM Journal on Computing, 14(2):257–276, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

McDiarmid:1979:DCN

[McD79] Colin McDiarmid. Determining the chromatic number of a graph. SIAM Journal on Computing, 8(1):1–14, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

McDiarmid:1988:ACL

[McD88] Colin McDiarmid. Averagecase lower bounds for searching. SIAM Journal on Computing, 17(5):1044–1060, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

McIlroy:1974:NBI

[McI74] M. D. McIlroy. The number of 1’s in binary integers: Bounds and extremal properties. SIAM Journal on Computing, 3(4): 255–261, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

McKay:1979:SRC

[McK79] J. McKay. Some remarks on computing Galois groups. SIAM Journal on Computing, 8(3):344–347, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Chung:2004:OAM

[mCL04] Kai min Chung and Hsueh-I Lu. An optimal algorithm for the maximum-density segment problem. SIAM Journal on Computing, 34(2):373–387, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44043.

McRae:1973:AER

[MD73] Walter B. McRae and Ernest R. Davidson. An algorithm for the extreme rays of a pointed convex polyhedral cone. SIAM Journal on Computing, 2(4): 281–293, December 1973. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mateti:1976:AEA

[MD76] Prabhaker Mateti and Narsingh Deo. On algorithms for enumerating all circuits of a graph. SIAM Journal on Computing, 5(1):90–99, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Megiddo:1983:LTA

[Meg83a] Nimrod Megiddo. Linear-time algorithms for linear programming in R^{3} and related problems. SIAM Journal on Computing, 12(4):759–776, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Megiddo:1983:TGP

[Meg83b] Nimrod Megiddo. Towards a genuinely polynomial algorithm for linear programming. SIAM Journal on Computing, 12(2):347–353, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mehlhorn:1977:BPB

[Meh77] Kurt Mehlhorn. A best possible bound for the weighted path length of binary search trees. SIAM Journal on Computing, 6(2):235–239, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mehlhorn:1979:DBS

[Meh79] Kurt Mehlhorn. Dynamic binary search. SIAM Journal on Computing, 8(2):175–198, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mehlhorn:1982:PAH

[Meh82]  Kurt Mehlhorn.  A par-

tial analysis of height-balanced trees under random insertions and deletions. SIAM Journal on Computing, 11(4):748–760, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mehta:2018:CRT

[Meh18] Ruta Mehta. Constant rank two-player games are PPADhard. SIAM Journal on Computing, 47(5):1858–1887, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Meir:2009:CCL

[Mei09] Or Meir. Combinatorial construction of locally testable codes. SIAM Journal on Computing, 39(2):491–544, ????

2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Meir:2013:IPU

[Mei13] Or Meir. IP = PSPACE using error-correcting codes. SIAM Journal on Computing, 42(1): 380–403, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Merkle:2001:GPA

[Mer01] Wolfgang Merkle. The global power of additional queries to P -random oracles. SIAM Journal on Computing, 31(2):483– 495, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
36671.

Merkle:2002:LEA

[Mer02] Wolfgang Merkle. Lattice embeddings for abstract bounded reducibilities. SIAM Journal on Computing, 31(4):1119– 1155, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38595.

Mestre:2010:ALR

[Mes10] Julián Mestre. Adaptive local ratio. SIAM Journal on Computing, 39(7):3038–3057, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Maletti:2009:PET

[MGHK09] Andreas Maletti, Jonathan Graehl, Mark Hopkins, and Kevin Knight. The power of extended top-down tree transducers. SIAM Journal on Computing, 39(2):410–430, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Marathe:1998:AAP

[MHSR98] Madhav V. Marathe, Harry B.

Hunt III, Richard E. Stearns, and Venkatesh Radhakrishnan. Approximation algorithms for PSPACE-hard hierarchically and periodically specified problems. SIAM Journal on Computing, 27 (5):1237–1261, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28525.

Makino:1997:MLI

[MI97] Kazuhisa Makino and Toshihide Ibaraki. The maximum latency and identification of positive Boolean functions. SIAM Journal on Computing, 26(5):1363–1383, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27632.

Micali:2000:CSP

[Mic00] Silvio Micali. Computationally sound proofs. SIAM Journal on Computing, 30 (4):1253–1298, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28495.

Micciancio:2004:APL

[Mic04] Daniele Micciancio. Almost perfect lattices, the covering radius problem, and applications to Ajtai’s connection factor. SIAM Journal on Computing, 34(1):118–169, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43351.

Makino:1999:HEP

[MiHI99] Kazuhisa Makino, Ken ichi Hatanaka, and Toshihide Ibaraki. Horn extensions of a partially defined Boolean function. SIAM Journal on Computing, 28(6):2168–2186, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
29795.

Miller:1972:SUR

[Mil72] Raymond E. Miller. Some undecidability results for parallel program schemata. SIAM Journal on Computing, 1(1): 119–129, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1975:CCN

[Mil75] Webb Miller. Computational complexity and numerical stability. SIAM Journal on Computing, 4(2):97–107, June 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1976:GTR

[Mil76] Webb Miller. Graph transformations for roundoff analysis. SIAM Journal on Computing, 5(2):204–216, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1988:LAT

[Mil88] Zevi Miller. A linear algorithm for topological bandwidth in degree-three trees. SIAM Journal on Computing, 17(5):1018– 1035, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miltersen:1993:CMM

[Mil93] Peter Bro Miltersen. The complexity of malign measures. SIAM Journal on Computing, 22(1):147–156, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Misra:1977:PLA

[Mis77] Jayadev Misra. Prospects and limitations of automatic assertion generation for loop programs. SIAM Journal on Computing, 6(4):718–729, ????

1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mitra:1985:PMA

[Mit85] Debasis Mitra. Probabilistic models and asymptotic results for concurrent processing with exclusive and nonexclusive locks. SIAM Journal on Computing, 14(4):1030– 1051, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mitchell:1999:GSA

[Mit99] Joseph S. B. Mitchell. Guillotine subdivisions approximate polygonal subdivisions: a simple polynomial-time approximation scheme for geometric TSP, k-MST, and related problems. SIAM Journal on Computing, 28(4):1298– 1309, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
30976.

Milidiu:2000:WAL

[ML00] Ruy Luiz Milidiú and Eduardo Sany Laber. The WARM-UP algorithm: a Lagrangian construction of length restricted Huffman codes. SIAM Journal on Computing, 30(5):1405–1426, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31981.

Ma:2000:GTS

[MLZ00] Bin Ma, Ming Li, and Louxin Zhang. From gene trees to species trees. SIAM Journal on Computing, 30(3):729–752, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34336.

Mitra:1983:AEM

[MM83] Debasis Mitra and J. A. Morrison. Asymptotic expansions of moments of the waiting time in a shared-processor of an interactive system. SIAM Journal on Computing, 12(4):789– 802, November 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Manber:1993:SAN

[MM93] Udi Manber and Gene Myers.

Suffix arrays: a new method for on-line string searches. SIAM Journal on Computing, 22(5):935–948, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Malajovich:1998:S

[MM98] Gregorio Malajovich and Klaus Meer. On the structure of N P_{C}. SIAM Journal on Computing, 28(1):27–35, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29498.

Malvestuto:2002:LAF

[MM02] F. M. Malvestuto and M. Mezzini. A linear algorithm for finding the invariant edges of an edge-weighted graph. SIAM Journal on Computing, 31 (5):1438–1455, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37606.

Madelaine:2018:CMC

[MM18] Florent R. Madelaine and Barnaby D. Martin. On the complexity of the model checking problem. SIAM Journal on Computing, 47(3):769–797, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Merino:2022:CGP

[MMM22] Arturo Merino, Ondrej Micka, and Torsten Mütze. On a combinatorial generation problem of Knuth. SIAM Journal on Computing, 51(3):379–423, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1377394.

Magniez:2014:RWP

[MMN14] Frédéric Magniez, Claire Mathieu, and Ashwin Nayak. Recognizing well-parenthesized expressions in the streaming model. SIAM Journal on Computing, 43(6):1880–1905, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mitchell:1987:DGP

[MMP87] Joseph S. B. Mitchell, David M.

Mount, and Christos H. Papadimitriou. The discrete geodesic problem. SIAM Journal on Computing, 16(4):647– 668, August 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Meyerson:2008:CDT

[MMP08] Adam Meyerson, Kamesh Munagala, and Serge Plotkin. Cost-distance: Two metric network design. SIAM Journal on Computing, 38(4):1648– 1659, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mejean:1995:VMA

[MMR95] Henri-M. Méjean, Henri Morel, and Gérard Reynaud. A variational method for analysing unit clause search. SIAM Journal on Computing, 24(3): 621–649, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23048.

Megow:2017:RHC

[MMS17] Nicole Megow, Julie Meißner, and Martin Skutella. Randomization helps computing a minimum spanning tree under uncertainty. SIAM Journal on Computing, 46(4):1217– 1240, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moser:1993:AFT

[MMSA93] Louise E. Moser, P. M. Melliar-Smith, and Vivek Agrawala. Asynchronous fault-tolerant total ordering algorithms. SIAM Journal on Computing, 22(4):727–750, August 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mekler:1987:EBI

[MN87] Alan H. Mekler and Evelyn M. Nelson. Equational bases for If–Then–Else. SIAM Journal on Computing, 16(3):465–485, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Masuda:1988:OAF

[MN88] Sumio Masuda and Kazuo Nakajima. An optimal algorithm for finding a maximum independent set of a circulararc graph. SIAM Journal on Computing, 17(1):41–52, February 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1995:FPG

[MN95] Gary L. Miller and Joseph (Seffi) Naor. Flow in planar graphs with multiple sources and sinks. SIAM Journal on Computing, 24(5):1002–1017, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/16299.

Moore:2001:PQC

[MN01] Cristopher Moore and Martin Nilsson. Parallel quantum computation and quantum codes. SIAM Journal on Computing, 31(3):799–815, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35505.

Mehlhorn:1988:LBC

[MNA88] Kurt Mehlhorn, Stefan Näher, and Helmut Alt. A lower bound on the complexity of the union-split-find problem. SIAM Journal on Computing, 17(6):1093–1102, December 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mehlhorn:1990:CGR

[MNR90] K. Mehlhorn, St. Näher, and M. Rauch. On the complexity of a game related to the dictionary problem. SIAM Journal on Computing, 19(5):902–906, October 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Magniez:2011:SQW

[MNRS11] Frédéric Magniez, Ashwin Nayak, Jérémie Roland, and Miklos Santha. Search via quantum walk. SIAM Journal on Computing, 40(1):142–164, ???? 2011. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i1/
                                        p142_s1.

Matsumoto:1985:EAF

[MNS85] Kazuhiko Matsumoto, Takao Nishizeki, and Nobuji Saito. An efficient algorithm for finding multicommodity flows in planar networks. SIAM Journal on Computing, 14(2):289– 302, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Matsumoto:1986:PMF

[MNS86] Kazuhiko Matsumoto, Takao Nishizeki, and Nobuji Saito. Planar multicommodity flows, maximum matchings and negative cycles. SIAM Journal on Computing, 15(2):495–510, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Martens:2010:CDP

[MNS10] Wim Martens, Frank Neven, and Thomas Schwentick. Complexity of decision problems for XML schemas and chain regular expressions. SIAM Journal on Computing, 39(4):1486– 1530, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mironov:2011:SAE

[MNS11] Ilya Mironov, Moni Naor, and Gil Segev. Sketching in adversarial environments. SIAM Journal on Computing, 40(6): 1845–1870, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i6/p1845_s1.

Muntz:1974:SRA

[MO74] R. R. Muntz and H. Opderbeck. Stack replacement algorithms for two-level directly addressable paged memories. SIAM Journal on Computing, 3(1):11–22, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moitra:2012:POS

[MO12] Ankur Moitra and Ryan O’Donnell. Pareto optimal solutions for smoothed analysts. SIAM Journal on Computing, 41(5):1266–1284, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moitra:2013:VSO

[Moi13] Ankur Moitra. Vertex sparsification and oblivious reductions. SIAM Journal on Computing, 42(6):2400–2423, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moitra:2016:AOA

[Moi16] Ankur Moitra. An almost optimal algorithm for computing nonnegative rank. SIAM Journal on Computing, 45(1):156– 173, ???? 2016. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Molzan:1990:ENC

[Mol90] B. Molzan. Expressibility and nonuniform complexity classes. SIAM Journal on Computing, 19(3):411–423, June 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Molloy:2003:MRC

[Mol03] Michael Molloy. Models for random constraint satisfaction problems. SIAM Journal on Computing, 32(4):935– 949, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        36866.

Molloy:2004:GDC

[Mol04] Michael Molloy. The Glauber dynamics on colorings of a graph with high girth and maximum degree. SIAM Journal on Computing, 33(3):721–737, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40178.

Mayer:2002:SSS

[MOOY02] Alain Mayer, Rafail Ostrovsky, Yoram Ofek, and Moti Yung. Self-stabilizing symmetry breaking in constant space. SIAM Journal on Computing, 31(5):1571–1595, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28599.

Mohanty:2022:ENR

[MOP22] Sidhanth Mohanty, Ryan O’Donnell, and Pedro Paredes. Explicit near-Ramanujan graphs of every degree. SIAM Journal on Computing, 51(3): STOC20–1–STOC20–23, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1342112.

Morgera:1980:ESI

[Mor80] Salvatore D. Morgera. Efficient synthesis and implementation of large discrete Fourier transformations. SIAM Journal on Computing, 9(2):251– 272, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moran:1982:ADH

[Mor82] Shlomo Moran. On the accepting density hierarchy in NP. SIAM Journal on Computing, 11(2):344–349, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mortensen:2006:FDO

[Mor06] Christian Worm Mortensen.

Fully dynamic orthogonal range reporting on RAM.

SIAM Journal on Computing, 35(6):1494–1525, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Morris:2008:MTT

[Mor08] Ben Morris. The mixing time of the Thorp shuffle. SIAM Journal on Computing, 38(2): 484–504, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Matulef:2010:TH

[MORS10] Kevin Matulef, Ryan O’Donnell, Ronitt Rubinfeld, and Rocco A. Servedio. Testing halfspaces. SIAM Journal on Computing, 39(5):2004–2047, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mount:1990:NSP

[Mou90] David M. Mount. The number of shortest paths on the surface of a polyhedron. SIAM Journal on Computing, 19(4):593–611, August 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

McColl:1977:DAB

[MP77] W. F. McColl and M. S. Patterson. The depth of all Boolean functions. SIAM Journal on Computing, 6(2): 373–380, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mosses:1987:PLC

[MP87] Peter D. Mosses and Gordon D. Plotkin. On proving limiting completeness. SIAM Journal on Computing, 16(1): 179–194, February 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mettu:2003:OMP

[MP03] Ramgopal R. Mettu and C. Greg Plaxton. The online median problem. SIAM Journal on Computing, 32(3): 816–832, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38344.

McCann:2005:SDA

[MP05] Mark McCann and Nicholas Pippenger. SRT division algorithms as dynamical systems. SIAM Journal on Computing, 34(6):1279–1301, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44106.

Marx:2022:SPA

[MPP22] Dániel Marx, Marcin Pilipczuk, and Michal Pilipczuk. A subexponential parameterized algorithm for directed subset traveling salesman problem on planar graphs. SIAM Journal on Computing, 51(2):254–289, ???? 2022. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1304088.

Mourrain:2003:ASM

[MPR03] B. Mourrain, V. Y. Pan, and O. Ruatta. Accelerated solution of multivariate polynomial systems of equations. SIAM Journal on Computing, 32(2):435–454, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38516.

Mull:2022:CBP

[MPR22] Nathan Mull, Shuo Pang, and Alexander Razborov. On CDCL-based proof systems with the ordered decision strategy. SIAM Journal on Computing, 51(4):1368–1399, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1362528.

Miller:1979:OT

[MPRS79] Raymond E. Miller, Nicholas Pippenger, Arnold L. Rosenberg, and Lawrence Snyder. Optimal 2,3-trees. SIAM Journal on Computing, 8(1):42– 59, February 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Also published in/as: IBM, TJWRC, Res.R. RC 6505, 1977.

Martel:1992:WOA

[MPS92] Charles Martel, Arvin Park, and Ramesh Subramonian. Work-optimal asynchronous algorithms for shared memory parallel computers. SIAM Journal on Computing, 21(6): 1070–1099, December 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Matousek:1994:FTD

[MPS^{+}94] Jiří Matoušek, János Pach, Micha Sharir, Shmuel Sifrony, and Emo Welzl. Fat triangles determine linearly many holes. SIAM Journal on Computing, 23(1):154–169, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18330.

Majster:1980:ELC

[MR80] Mila E. Majster and Angelika Reiser. Efficient on-line construction and correction of position trees. SIAM Journal on Computing, 9(4):785–807, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Melhem:1984:MMV

[MR84] Rami G. Melhem and Werner C.

Rheinboldt. A mathematical model for the verification of systolic networks. SIAM Journal on Computing, 13(3):541– 565, ???? 1984. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1991:PTC

[MR91] Gary L. Miller and John H. Reif. Parallel tree contraction part 2: Further applications. SIAM Journal on Computing, 20(6):1128–1147, December 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mathur:1996:GKI

[MR96] Anmol Mathur and Edward M. Reingold. Generalized Kraft’s inequality and discrete kmodal search. SIAM Journal on Computing, 25(2):420–447, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24636.

Mahajan:1999:DAA

[MR99] Sanjeev Mahajan and H. Ramesh. Derandomizing approximation algorithms based on semidefinite programming. SIAM Journal on Computing, 28 (5):1641–1663, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30932.

Martinez:2001:OSS

[MR01a] Conrado Martínez and Salvador Roura. Optimal sampling strategies in Quicksort and Quickselect. SIAM Journal on Computing, 31(3):683– 705, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
38210.

Munro:2001:SRB

[MR01b] J. Ian Munro and Venkatesh Raman. Succinct representation of balanced parentheses and static trees. SIAM Journal on Computing, 31(3):762–776, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36409.

Moses:2002:LAC

[MR02] Yoram Moses and Sergio Rajsbaum. A layered analysis of consensus. SIAM Journal on Computing, 31 (4):989–1021, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36400.

Micciancio:2007:WCA

[MR07a] Daniele Micciancio and Oded Regev. Worst-case to averagecase reductions based on Gaussian measures. SIAM Journal on Computing, 37(1):267–302, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moss:2007:AAC

[MR07b] A. Moss and Y. Rabani. Approximation algorithms for constrained node weighted Steiner tree problems. SIAM Journal on Computing, 37(2): 460–481, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moshkovitz:2008:SCE

[MR08] Dana Moshkovitz and Ran Raz. Sub-constant error low degree test of almost-linear size. SIAM Journal on Computing, 38(1):140–180, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mossel:2010:SIS

[MR10] Elchanan Mossel and Sebastien Roch. Submodularity of influence in social networks: From local to global. SIAM Journal on Computing, 39(6):2176– 2188, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moore:2012:APN

[MR12] Cristopher Moore and Alexander Russell. Approximating the permanent via nonabelian determinants. SIAM Journal on Computing, 41(2):332–355, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Marx:2014:FPT

[MR14] Dániel Marx and Igor Razgon.

Fixed-parameter tractability of multicut parameterized by the size of the cutset. SIAM Journal on Computing, 43(2): 355–388, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1988:EPE

[MRK88] Gary L. Miller, Vijaya Ramachandran, and Erich Kaltofen. Efficient parallel evaluation of straight-line code and arithmetic circuits. SIAM Journal on Computing, 17(4):687–695, August 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mustafa:2015:QPT

[MRR15] Nabil H. Mustafa, Rajiv Raman, and Saurabh Ray. Quasipolynomial time approximation scheme for weighted geometric set cover on pseudodisks and halfspaces. SIAM Journal on Computing, 44(6): 1650–1669, ???? 2015. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moore:2007:PSF

[MRRS07] Cristopher Moore, Daniel Rockmore, Alexander Russell, and Leonard J. Schulman. The power of strong Fourier sampling: Quantum algorithms for affine groups and hidden shifts.

SIAM Journal on Computing, 37(3):938–958, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mostefaoui:2008:CPC

[MRRT08] Achour Mostefaoui, Sergio Rajsbaum, Michel Raynal, and Corentin Travers. The combined power of conditions and information on failures to solve asynchronous set agreement. SIAM Journal on Computing, 38(4):1574–1601, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Micali:1988:NSP

[MRS88] Silvio Micali, Charles Rackoff, and Bob Sloan. The notion of security for probabilistic cryptosystems. SIAM Journal on Computing, 17(2):412– 426, April 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Moore:2008:SGD

[MRS08] Cristopher Moore, Alexander Russell, and Leonard J. Schulman. The symmetric group defies strong Fourier sampling. SIAM Journal on Computing, 37(6):1842–1864, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moore:2010:IQS

[MRŚ10] Cristopher Moore, Alexander Russell, and Piotr Śniady. On the impossibility of a quantum sieve algorithm for graph isomorphism. SIAM Journal on Computing, 39(6):2377– 2396, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mukhopadhyay:2018:SBD

[MRS18] Sagnik Mukhopadhyay, Jaikumar Radhakrishnan, and Swagato Sanyal. Separation between deterministic and randomized query complexity. SIAM Journal on Computing, 47(4):1644–1666, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Muthukrishnan:2005:OSM

[MRSG05] S. Muthukrishnan, Rajmohan Rajaraman, Anthony Shaheen, and Johannes E. Gehrke. Online scheduling to minimize average stretch. SIAM Journal on Computing, 34(2):433–452, April 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38754.

Murtagh:2021:DBC

[MRSV21] Jack Murtagh, Omer Reingold, Aaron Sidford, and Salil Vadhan. Derandomization beyond connectivity: Undirected Laplacian systems in nearly logarithmic space. SIAM Journal on Computing, 50(6):1892– 1922, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mounie:2007:AAS

[MRT07] Gregory Mounie, Christophe Rapine, and Denis Trystram. A \frac{3}{2}-approximation algorithm for scheduling independent monotonic malleable tasks. SIAM Journal on Computing, 37(2):401–412, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Malkhi:2000:LAB

[MRW00] Dahlia Malkhi, Michael K. Reiter, and Avishai Wool. The load and availability of Byzantine quorum systems. SIAM Journal on Computing, 29(6):1889–1906, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32523.

Manna:1976:TAO

[MS76a] Zohar Manna and Adi Shamir. The theoretical aspects of the optimal FixedPoint. SIAM Journal on Computing, 5(3): 414–426, September 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Munro:1976:SSM

[MS76b] Ian Munro and Philip M.

Spira. Sorting and searching in multisets. SIAM Journal on Computing, 5(1):1–8, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Megiddo:1984:CSC

[MS84] Nimrod Megiddo and Kenneth J. Supowit. On the complexity of some common geometric location problems. SIAM Journal on Computing, 13(1):182–196, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Maass:1987:STM

[MS87a] Wolfgang Maass and Amir Schorr. Speed-up of Turing machines with one work tape and a two-way input tape. SIAM Journal on Computing, 16(1):195–202, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1987:DMT

[MS87b] Russ Miller and Quentin F.

Stout. Data movement techniques for the pyramid computer. SIAM Journal on Computing, 16(1):38–60, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Munshi:1990:SSL

[MS90] Ashfaq A. Munshi and Barbara Simons. Scheduling sequential loops on parallel processors. SIAM Journal on Computing, 19(4):728–741, August 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Marcotte:1991:FMA

[MS91] Odile Marcotte and Subhash Suri. Fast matching algorithms for points on a polygon. SIAM Journal on Computing, 20(3): 405–422, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Melissaratos:1992:SPH

[MS92] Elefterios A. Melissaratos and Diane L. Souvaine. Shortest paths help solve geometric optimization problems in planar regions. SIAM Journal on Computing, 21(4):601–638, August 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mohaban:1997:RSA

[MS97] Shai Mohaban and Micha Sharir. Ray shooting amidst spheres in three dimensions and related problems. SIAM Journal on Computing, 26(3): 654–674, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
25208.

Maggs:1999:SAR

[MS99] Bruce M. Maggs and Ramesh K.

Sitaraman. Simple algorithms for routing on butterfly networks with bounded queues. SIAM Journal on Computing, 28(3):984–1003, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31181.

Mulmuley:2001:GCT

[MS01] Ketan D. Mulmuley and Milind Sohoni. Geometric complexity theory I: An approach to the P vs. NP and related problems. SIAM Journal on Computing, 31(2):496–526, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38715.

Morris:2004:RWT

[MS04] Ben Morris and Alistair Sinclair. Random walks on truncated cubes and sampling 0-1 knapsack solutions. SIAM Journal on Computing, 34 (1):195–226, February 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41191.

Madelaine:2007:CSL

[MS07a] Florent Madelaine and Iain A.

Stewart. Constraint satisfaction, logic and forbidden patterns. SIAM Journal on Computing, 37(1):132–163, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Molloy:2007:RCR

[MS07b] Michael Molloy and Mohammad R. Salavatipour. The resolution complexity of random constraint satisfaction problems. SIAM Journal on Computing, 37(3):895–922, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Markov:2008:SQC

[MS08a] Igor L. Markov and Yaoyun Shi. Simulating quantum computation by contracting tensor networks. SIAM Journal on Computing, 38(3):963–981, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mulmuley:2008:GCT

[MS08b] Ketan D. Mulmuley and Milind Sohoni. Geometric complexity theory II: Towards explicit obstructions for embeddings among class varieties. SIAM Journal on Computing, 38(3):1175–1206, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ma:2010:MEA

[MS10] Bin Ma and Xiaoming Sun. More efficient algorithms for closest string and substring problems. SIAM Journal on Computing, 39(4):1432–1443, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

McGregor:2016:SSF

[MS16] Andrew McGregor and Rahul Santhanam. Special section on the Forty-Fourth Annual ACM Symposium on Theory of Computing (STOC 2012). SIAM Journal on Computing, 45(4): 1448–1449, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:2017:USR

[MS17] Carl A. Miller and Yaoyun Shi.

Universal security for randomness expansion from the spotchecking protocol. SIAM Journal on Computing, 46(4):1304– 1335, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mohring:2004:SAP

[MSS04] Rolf H. Möhring, Martin Skutella, and Frederik Stork. Scheduling with AND/OR precedence constraints. SIAM Journal on Computing, 33(2): 393–415, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37727.

Magniez:2007:QAT

[MSS07] Frédéric Magniez, Miklos Santha, and Mario Szegedy. Quantum algorithms for the triangle problem. SIAM Journal on Computing, 37(2):413–424, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Marcus:2018:IFI

[MSS18] Adam W. Marcus, Daniel A. Spielman, and Nikhil Srivastava. Interlacing families IV: Bipartite Ramanujan graphs of all sizes. SIAM Journal on Computing, 47(6):2488–2509, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mansour:1991:LBC

[MST91] Yishay Mansour, Baruch Schieber, and Prasoon Tiwari. Lower bounds for computations with the floor operation. SIAM Journal on Computing, 20(2):315–327, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Morrison:1987:QAH

[MSV87] John A. Morrison, Larry A. Shepp, and Christopher J. Van Wyk. A queueing analysis of hashing with lazy deletion. SIAM Journal on Computing, 16(6):1155–1164, December 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Matias:2006:EBS

[MSV06] Yossi Matias, Eran Segal, and Jeffrey Scott Vitter. Efficient bundle sorting. SIAM Journal on Computing, 36(2):394–410, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Megow:2016:PRO

[MSVW16] Nicole Megow, Martin Skutella, José Verschae, and Andreas Wiese. The power of recourse for online MST and TSP. SIAM Journal on Computing, 45(3):859–880, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Maurer:1978:GEF

[MSW78] H. A. Maurer, A. Salomaa, and D. Wood. On good EOL forms. SIAM Journal on Computing, 7(2):158–166, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Maurer:1981:UIG

[MSW81] H. A. Maurer, A. Salomaa, and D. Wood. Uniform interpretations of grammar forms. SIAM Journal on Computing, 10(3): 483–502, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mertzios:2011:RTB

[MSZ11] George B. Mertzios, Ignasi Sau, and Shmuel Zaks. The recognition of tolerance and bounded tolerance graphs. SIAM Journal on Computing, 40(5):1234–1257, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i5/p1234_s1.

Megiddo:1978:ACM

[MT78] Nimrod Megiddo and Arie Tamir. An O(n · log n) algorithm for a class of matching problems. SIAM Journal on Computing, 7(2):154–157, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Megiddo:1983:NRC

[MT83] Nimrod Megiddo and Arie Tamir. New results on the complexity of p-centre problems. SIAM Journal on Computing, 12(4):751–758, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Manber:1984:ENH

[MT84] Udi Manber and Martin Tompa. The effect of number of Hamiltonian paths on the complexity of a vertexcoloring problem. SIAM Journal on Computing, 13(1):109– 115, ???? 1984. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mehlhorn:1986:AAI

[MT86] Kurt Mehlhorn and Athanasios Tsakalidis. An amortized analysis of insertions into AVL-trees. SIAM Journal on Computing, 15(1):22–33, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moffat:1987:APS

[MT87] Alistair Moffat and Tadao Takaoka. An all pairs shortest path algorithm with expected time O(n^{2} log n). SIAM Journal on Computing, 16(6):1023– 1031, December 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Melen:1989:NMN

[MT89] Riccardo Melen and Jonathan S.

Turner. Nonblocking multirate networks. SIAM Journal on Computing, 18(2):301–313, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miller:1999:DPC

[MT99] Gary L. Miller and Shang-Hua Teng. The dynamic parallel complexity of computational circuits. SIAM Journal on Computing, 28 (5):1664–1688, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28172.

McKenzie:2010:EUB

[MTV10] Pierre McKenzie, Michael Thomas, and Heribert Vollmer. Extensional uniformity for Boolean circuits. SIAM Journal on Computing, 39(7):3186– 3206, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Megiddo:1981:ATL

[MTZC81] N. Megiddo, A. Tamir, E. Zemel, and R. Chandrasekaran. An O(n log^{2} n) algorithm for the k th longest path in a tree with applications to location problems. SIAM Journal on Computing, 10(2):328–337, May 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Muchnick:1976:CCM

[Muc76] Steven S. Muchnick. Computational complexity of multiple recursive schemata. SIAM Journal on Computing, 5(3): 427–451, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mulmuley:1999:LBP

[Mul99]  Ketan Mulmuley.  Lower

bounds in a parallel model without bit operations. SIAM Journal on Computing, 28 (4):1460–1509, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28293.

Murgolo:1987:EAS

[Mur87] Frank D. Murgolo. An efficient approximation scheme for variable-sized bin packing. SIAM Journal on Computing, 16(1):149–161, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Murota:1990:CPS

[Mur90] Kazuo Murota. Computing puiseux-series solutions to determinantal equations via combinatorial relaxation. SIAM Journal on Computing, 19(6): 1132–1161, December 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Murota:1995:CDD

[Mur95] Kazuo Murota. Computing the degree of determinants via combinatorial relaxation. SIAM Journal on Computing, 24(4):765–796, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/20189.

Mandl:1975:CCI

[MV75] Robert Mandl and Thomas Vari. Computational complexity of inner products of vectors (and that of other bilinear forms) over a noncommutative ring (auxiliary functions allowed). SIAM Journal on Computing, 4(1):49–55, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mitchell:2000:QMG

[MV00] Scott A. Mitchell and Stephen A. Vavasis. Quality mesh generation in higher dimensions. SIAM Journal on Computing, 29(4):1334–1370, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31412.

Micciancio:2013:DSE

[MV13] Daniele Micciancio and Panagiotis Voulgaris. A deterministic single exponential time algorithm for most lattice problems based on Voronoi cell computations. SIAM Journal on Computing, 42(3):1364– 1391, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moore:2018:SSF

[MV18] Cris Moore and Santosh Vempala. Special section on the Fifty-Sixth Annual IEEE Symposium on Foundations of Computer Science (FOCS 2015). SIAM Journal on Computing, 47(6):2237, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Menezes:1992:SRA

[MvV92] A. J. Menezes, P. C. van Oorschot, and S. A. Vanstone. Subgroup refinement algorithms for root finding in GF(q). SIAM Journal on Computing, 21(2):228–239, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

McKenzie:2004:ACP

[MVW04] Pierre McKenzie, Heribert Vollmer, and Klaus W. Wagner. Arithmetic circuits and polynomial replacement systems. SIAM Journal on Computing, 33(6):1513–1531, December 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39207.

Morris:1988:STN

[MW88] Robert J. T. Morris and Wing Shing Wong. A shortterm neural network memory. SIAM Journal on Computing, 17(6):1103–1118, December 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moran:1993:GTD

[MW93] Shlomo Moran and Manfred K.

Warmuth. Gap theorems for distributed computation. SIAM Journal on Computing, 22(2):379–394, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mendelzon:1995:FRS

[MW95] Alberto O. Mendelzon and Peter T. Wood. Finding regular simple paths in graph databases. SIAM Journal on Computing, 24(6):1235–1258, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        22370.

Maurer:1999:RBB

[MW99a] Ueli M. Maurer and Stefan Wolf. The relationship between breaking the Diffie– Hellman protocol and computing discrete logarithms. SIAM Journal on Computing, 28(5):1689–1721, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30274.

Miyazawa:1999:AAO

[MW99b] F. K. Miyazawa and Y. Wakabayashi. Approximation algorithms for the orthogonal Z-oriented three-dimensional packing problem. SIAM Journal on Computing, 29(3):1008– 1029, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        31391.

Murray:2020:CLB

[MW20] Cody D. Murray and R. Ryan Williams. Circuit lower bounds for nondeterministic quasipolytime from a new easy witness lemma. SIAM Journal on Computing, 49(5):STOC18– 300–STOC18–322, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Machtey:1978:SGN

[MWY78] Michael Machtey, Karl Winklmann, and Paul Young. Simple Gödel numberings, isomorphisms, and programming properties. SIAM Journal on Computing, 7(1):39–60, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mehlhorn:1991:CWG

[MY91] Kurt Mehlhorn and Chee-Keng Yap. Constructive Whitney– Graustein theorem, or how to untangle closed planar curves. SIAM Journal on Computing, 20(4):603–621, August 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Myers:1985:ETA

[Mye85]  Eugene W. Myers.  An

O(E log E + I) expected time algorithm for the planar segment intersection problem. SIAM Journal on Computing, 14(3):625–637, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Miyakawa:1977:OST

[MYSH77] Masahiro Miyakawa, Toshitsugu Yuba, Yoshio Sugito, and Mamoru Hoshi. Optimum sequence trees. SIAM Journal on Computing, 6(2):201–234, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Mahdian:2006:AAM

[MYZ06] Mohammad Mahdian, Yinyu Ye, and Jiawei Zhang. Approximation algorithms for metric facility location problems. SIAM Journal on Computing, 36(2):411–432, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Maheshwari:2008:EPS

[MZ08] Anil Maheshwari and Norbert Zeh. I/O-efficient planar separators. SIAM Journal on Computing, 38(3):767–801, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Meka:2013:PGP

[MZ13] Raghu Meka and David Zuckerman. Pseudorandom generators for polynomial threshold functions. SIAM Journal on Computing, 42(3):1275– 1301, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Natarajan:1991:PAL

[Nat91] B. K. Natarajan. Probably approximate learning of sets and functions. SIAM Journal on Computing, 20(2):328–351, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Natarajan:1992:PAL

[Nat92] B. K. Natarajan. Probably approximate learning over classes of distributions. SIAM Journal on Computing, 21(3):438–449, June 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Natarajan:1995:SAS

[Nat95] B. K. Natarajan. Sparse approximate solutions to linear systems. SIAM Journal on Computing, 24(2):227–234, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24040.

Natarajan:1999:LFN

[Nat99] B. Natarajan. On learning functions from noise-free and noisy samples via Occam’s razor. SIAM Journal on Computing, 29(3):712–727, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27711.

Nedev:1999:FES

[Ned99] Zhivko Prodanov Nedev. Finding an even simple path in a directed planar graph. SIAM Journal on Computing, 29(2):

685–695, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
33034.

Newman:2002:TML

[New02] Ilan Newman. Testing membership in languages that have small width branching programs. SIAM Journal on Computing, 31(5):1557–1570, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38211.

Naor:1997:FAO

[NGM97] Dalit Naor, Dan Gusfield, and Charles Martel. A fast algorithm for optimally increasing the edge connectivity. SIAM Journal on Computing, 26(4):1139–1165, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23422.

Ngo:2005:WSN

[Ngo05] Hung Q. Ngo. WDM switching networks, rearrangeable and nonblocking [w,f]-connectors. SIAM Journal on Computing, 35(3):766–785, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ng:1990:LBS

[NH90] Cheng Ng and Daniel S.

Hirschberg. Lower bounds for the stable marriage problem and its variants. SIAM Journal on Computing, 19(1):71– 77, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nakajima:1982:CRS

[NHL82] K. Nakajima, S. L. Hakimi, and J. K. Lenstra. Complexity results for scheduling tasks in fixed intervals on two types of machines. SIAM Journal on Computing, 11(3):512–520, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nicol:1988:EPS

[Nic88] David M. Nicol. Expected performance of m-solution backtracking. SIAM Journal on Computing, 17(1):114–127, February 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nisan:1991:CPD

[Nis91] Noam Nisan. CREW PRAMs and decision trees. SIAM Journal on Computing, 20(6):999– 1007, December 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nishimura:1994:MAS

[Nis94] Naomi Nishimura. A model for asynchronous shared memory parallel computation. SIAM Journal on Computing, 23(6): 1231–1252, December 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/21967.

Naor:1993:SBP

[NN93] Joseph Naor and Moni Naor. Small-bias probability spaces: Efficient constructions and applications. SIAM Journal on Computing, 22(4):838–856, August 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Navarro:2014:ODS

[NN14] Gonzalo Navarro and Yakov Nekrich. Optimal dynamic sequence representations. SIAM Journal on Computing, 43(5): 1781–1806, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Navarro:2017:TOT

[NN17] Gonzalo Navarro and Yakov Nekrich. Time-optimal topk document retrieval. SIAM Journal on Computing, 46(1): 80–113, ???? 2017. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Naor:1989:FPA

[NNS89] Joseph Naor, Moni Naor, and Alejandro A. Schäffer. Fast parallel algorithms for chordal graphs. SIAM Journal on Computing, 18(2):327–349, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Norton:1989:PAR

[Nor89] G. H. Norton. Precise analyses of the right- and left-shift greatest common divisor algorithms for GF(q)[x]. SIAM Journal on Computing, 18(3): 608–624, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nordstrom:2009:NPM

[Nor09] Jakob Nordström. Narrow proofs may be spacious: Separating space and width in resolution. SIAM Journal on Computing, 39(1):59–121, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Noshita:1982:PND

[Nos82] Kohei Noshita. Predicting the number of distinct elements in a multiset. SIAM Journal on Computing, 11(4):611–619, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nivat:1997:MAD

[NP97] Maurice Nivat and Andreas Podelski. Minimal ascending and descending tree automata. SIAM Journal on Computing, 26(1):39–58, February 1997. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/16407.

Naor:2006:OPE

[NP06] Moni Naor and Benny Pinkas. Oblivious polynomial evaluation. SIAM Journal on Computing, 35(5):1254–1281, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nguyen:2022:SPA

[NP22] Danny Nguyen and Igor Pak. Short Presburger arithmetic is hard. SIAM Journal on Computing, 51(2):STOC17– 1–STOC17–30, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/17M1151146.

Nederlof:2023:FET

[NPSW23] Jesper Nederlof, Jakub Pawlewicz, Céline M. F. Swennenhuis, and Karol Wegrzycki. A faster exponential time algorithm for bin packing with a constant number of bins via additive combinatorics. SIAM Journal on Computing, 52(6):1369– 1412, November 2023. CODEN SMJCAT. ISSN 1095-7111.

Nievergelt:1973:BST

[NR73] J. Nievergelt and E. M. Reingold. Binary search trees of bounded balance. SIAM Journal on Computing, 2(1):33–43, March 1973. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Naor:1995:OFS

[NR95] Moni Naor and Ron M. Roth.

Optimal file sharing in distributed networks. SIAM Journal on Computing, 24 (1):158–183, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23746.

Newman:2013:MAS

[NR13] Ilan Newman and Yuri Rabinovich. On multiplicative λapproximations and some geometric applications. SIAM Journal on Computing, 42(3): 855–883, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Naor:2002:PFF

[NRR02] Moni Naor, Omer Reingold, and Alon Rosen. Pseudorandom functions and factoring. SIAM Journal on Computing, 31(5):1383–1404, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38925.

Nisan:1999:PHB

[NRS99] Noam Nisan, Steven Rudich, and Michael Saks. Products and help bits in decision trees. SIAM Journal on Computing, 28(3):1035–1050, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28244.

Nassimi:1980:FCC

[NS80] David Nassimi and Sartaj Sahni. Finding connected components and connected ones on a mesh-connected parallel computer. SIAM Journal on Computing, 9(4):744–757, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Niederreiter:1993:LRP

[NS93] H. Niederreiter and C. P. Schnorr. Local randomness in polynomial random number and random function generators. SIAM Journal on Computing, 22(4):684–694, August 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Naor:1995:WCC

[NS95] Moni Naor and Larry Stockmeyer. What can be computed locally? SIAM Journal on Computing, 24(6): 1259–1277, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25457.

Narasimhan:2000:ASF

[NS00] Giri Narasimhan and Michiel Smid. Approximating the stretch factor of Euclidean graphs. SIAM Journal on Computing, 30(3):978–989, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36167.

Naor:2007:PE

[NS07] Assaf Naor and Gideon Schechtman. Planar earthmover is not in L_{1}. SIAM Journal on Computing, 37(3):804–826, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nickerson:2008:DRS

[NS08] Bradford G. Nickerson and Qingxiu Shi. On k-d range search with Patricia tries. SIAM Journal on Computing, 37(5):1373–1386, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nguyen:2009:LAQ

[NS09] Phong Q. Nguyen and Damien Stehlé. An LLL algorithm with quadratic complexity. SIAM Journal on Computing, 39(3): 874–903, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Naor:2012:PKC

[NS12] Moni Naor and Gil Segev. Public-key cryptosystems resilient to key leakage. SIAM Journal on Computing, 41(4): 772–814, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Newman:2013:EPH

[NS13] Ilan Newman and Christian Sohler. Every property of hyperfinite graphs is testable. SIAM Journal on Computing, 42(3):1095–1112, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nowotka:2022:OBS

[NS22] Dirk Nowotka and Aleksi Saarela. An optimal bound on the solution sets of onevariable word equations and its consequences. SIAM Journal on Computing, 51(1):1–18, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1310448.

Nguyen:2008:ASS

[NSH^{+}08] C. Thach Nguyen, Jian Shen, Minmei Hou, Li Sheng, Webb Miller, and Louxin Zhang. Approximating the spanning star forest problem and its application to genomic sequence alignment. SIAM Journal on Computing, 38(3):946–962, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Natanzon:2000:PAA

[NSS00] Assaf Natanzon, Ron Shamir, and Roded Sharan. A polynomial approximation algorithm for the minimum fill-in problem. SIAM Journal on Computing, 30(4):1067–1079, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33607.

Narayanan:1994:RPA

[NSV94] H. Narayanan, Huzur Saran, and Vijay V. Vazirani. Randomized parallel algorithms for matroid union and intersection, with applications to arborescences and edge-disjoint spanning trees. SIAM Journal on Computing, 23(2):387–397, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19524.

Niyogi:2011:TVU

[NSW11] P. Niyogi, S. Smale, and S. Weinberger. A topological view of unsupervised learning from noisy data. SIAM Journal on Computing, 40(3): 646–663, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.

                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i3/p646_s1.

Nickelsen:2005:CFP

[NT05] Arfst Nickelsen and Till Tantau. The complexity of finding paths in graphs with bounded independence number. SIAM Journal on Computing, 34(5):1176–1195, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/44164.

Nikolov:2016:GDP

[NTZ16] Aleksandar Nikolov, Kunal Talwar, and Li Zhang. The geometry of differential privacy: The small database and approximate cases. SIAM Journal on Computing, 45(2):575– 616, ???? 2016. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Nutov:2010:ASN

[Nut10] Zeev Nutov. Approximating Steiner networks with nodeweights. SIAM Journal on Computing, 39(7):3001–3022, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ngo:2003:MRC

[NV03] Hung Q. Ngo and Van H. Vu. Multirate rearrangeable Clos networks and a generalized edge-coloring problem on bipartite graphs. SIAM Journal on Computing, 32 (4):1040–1049, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40823.

Nisan:1993:RCC

[NW93] Noam Nisan and Avi Wigderson. Rounds in communication complexity revisited. SIAM Journal on Computing, 22(1): 211–219, February 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Naor:1998:LCA

[NW98] Moni Naor and Avishai Wool.

The load, capacity, and availability of quorum systems. SIAM Journal on Computing, 27(2):423–447, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28123.

Niggl:2006:CPT

[NW06] Karl-Heinz Niggl and Henning Wunderlich. Certifying polynomial time and linear/ polynomial space for imperative programs. SIAM Journal on Computing, 35(5):1122– 1147, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Naor:2001:AAD

[NZ01] Joseph (Seffi) Naor and Leonid Zosin. A 2-approximation algorithm for the directed multiway cut problem. SIAM Journal on Computing, 31(2): 477–482, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32147.

Opderbeck:1975:RMP

[OC75] H. Opderbeck and W. W. Chu. The renewal model for program behavior. SIAM Journal on Computing, 4(3):356– 374, ???? 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ogihara:1995:PTM

[Ogi95] Mitsunori Ogihara. Polynomialtime membership comparable sets. SIAM Journal on Computing, 24(5):1068–1081, October 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25813.

Ogihara:1998:PHC

[Ogi98] Mitsunori Ogihara. The PL hierarchy collapses. SIAM Journal on Computing, 27 (5):1430–1437, October 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29592.

Oommen:1987:LOS

[OH87] B. John Oommen and E. R. Hansen. List organizing strategies using stochastic move-to-front and stochastic move-to-rear operations. SIAM Journal on Computing, 16(4):705–716, August 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ohtsuki:1976:FAF

[Oht76] Tatsuo Ohtsuki. A fast algorithm for finding an optimal ordering for vertex elimination on a graph. SIAM Journal on Computing, 5(1):133– 145, ???? 1976. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Osteen:1974:PSB

[OL74] R. E. Osteen and P. P. Lin.

Picture skeletons based on eccentricities of points of minimum spanning trees. SIAM Journal on Computing, 3(1): 23–40, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Opatrny:1979:TOP

[Opa79] J. Opatrný. Total ordering problem. SIAM Journal on Computing, 8(1):111–114, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ottmann:1984:MCB

[OPR^{+}84] Thomas Ottmann, D. Stott Parker, Arnold L. Rosenberg, Hans W. Six, and Derick Wood. Minimal-cost brother trees. SIAM Journal on Computing, 13(1):197–217, ????

1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Orlin:2004:ALS

[OPS04] James B. Orlin, Abraham P. Punnen, and Andreas S. Schulz. Approximate local search in combinatorial optimization. SIAM Journal on Computing, 33(5):1201– 1214, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        43100.

ORourke:1986:SPC

[O’R86] Joseph O’Rourke. The signature of a plane curve. SIAM Journal on Computing, 15(1): 34–51, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Okrasa:2021:FGC

[OR21] Karolina Okrasa and Paweł Rzazewski. Fine-grained complexity of the graph homomorphism problem for boundedtreewidth graphs. SIAM Journal on Computing, 50(2):487– 508, ???? 2021. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ordman:1989:MTS

[Ord89] Edward T. Ordman. Minimal threshold separators and memory requirements for synchronization. SIAM Journal on Computing, 18(1):152–165, February 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Orponen:1986:OAP

[ORS86] Pekka Orponen, David A.

Russo, and Uwe Schöning. Optimal approximations and polynomially levelable sets. SIAM Journal on Computing, 15(2):399–408, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ogden:1985:ILC

[ORW85] William Ogden, Rockford J.

Ross, and Karl Winklmann. An “interchange lemma” for context-free languages. SIAM Journal on Computing, 14(2): 410–415, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

ODonnell:2007:LMD

[OS07] Ryan O’Donnell and Rocco A.

Servedio. Learning monotone decision trees in polynomial time. SIAM Journal on Computing, 37(3):827–844, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

ODonnell:2011:CPP

[OS11] Ryan O’Donnell and Rocco A.

Servedio. The Chow parameters problem. SIAM Journal on Computing, 40(1):165–199, ???? 2011. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sicomp/
resource/1/smjcat/v40/i1/
p165_s1.

Otto:1986:CRT

[Ott86] Friedrich Otto. Church–Rosser Thue systems that present free monoids. SIAM Journal on Computing, 15(3):786–792, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ogiwara:1991:PTB

[OW91] Mitsunori Ogiwara and Osamu Watanabe. On polynomialtime bounded truth-table reducibility of NP sets to sparse sets. SIAM Journal on Computing, 20(3):471–483, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

ODonnell:2013:KKK

[OW13] Ryan O’Donnell and Karl Wimmer. KKL, Kruskal– Katona, and monotone nets. SIAM Journal on Computing, 42(6):2375–2399, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Oyamaguchi:1993:NSD

[Oya93]  Michio Oyamaguchi.  NV-

sequentiality: a decidable condition for call-by-need computations in term-rewriting systems. SIAM Journal on Computing, 22(1):114–135, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Probst:1979:FAP

[PA79] David K. Probst and Vangalur S. Alagar. A family of algorithms for powering sparse polynomials. SIAM Journal on Computing, 8(4):626–644, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See corrigendum [PA80].

Probst:1980:CFA

[PA80] David K. Probst and Vangalur S. Alagar. Corrigendum: “A family of algorithms for powering sparse polynomials”. SIAM Journal on Computing, 9(2):439, May 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [PA79].

Pacault:1974:CWC

[Pac74] Jean François Pacault. Computing the weak components of a directed graph. SIAM Journal on Computing, 3(1):56–61, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pagh:2001:LRS

[Pag01] Rasmus Pagh. Low redundancy in static dictionaries with constant query time. SIAM Journal on Computing, 31(2):353–363, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36990.

Pan:1980:NFA

[Pan80] V. Ya. Pan. New fast algorithms for matrix operations. SIAM Journal on Computing, 9(2):321–342, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pan:1994:NRI

[Pan94]  V. Y. Pan.  New resul-

tant inequalities and complex polynomial factorization. SIAM Journal on Computing, 23(5):934–950, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23571.

Pan:2000:PCC

[Pan00]  Victor Y. Pan.  Parallel

complexity of computations with general and Toeplitzlike matrices filled with integers and extensions. SIAM Journal on Computing, 30 (4):1080–1125, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34995.

Papadimitriou:1981:WCP

[Pap81] Christos H. Papadimitriou. Worst-case and probabilistic analysis of a geometric location problem. SIAM Journal on Computing, 10(3):542–557, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Papadimitriou:1983:CCL

[Pap83] Christos H. Papadimitriou.

Concurrency control by locking. SIAM Journal on Computing, 12(2):215–226, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Papadimitriou:1992:CLK

[Pap92] Christos H. Papadimitriou.

The complexity of the Lin– Kernighan heuristic for the traveling salesman problem. SIAM Journal on Computing, 21(3):450–465, June 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pardo:1977:SSM

[Par77] Luis Trabb Pardo. Stable sorting and merging with optimal space and time bounds. SIAM Journal on Computing, 6(2): 351–372, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Parker:1980:COH

[Par80] D. Stott Parker, Jr. Conditions for optimality of the Huffman algorithm. SIAM Journal on Computing, 9(3):470– 489, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [Par98].

Parker:1998:ECO

[Par98] D. Stott Parker, Jr. Erratum: Conditions for optimality of the Huffman algorithm. SIAM Journal on Computing, 27(1):317, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97404.  See

[Par80].

Patrascu:2011:ULC

[Păt11]  Mihai Pătraşcu.  Unifying

the landscape of cell-probe lower bounds. SIAM Journal on Computing, 40(3):827–847, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i3/
                                        p827_s1.

Patel:2013:DEE

[Pat13] Viresh Patel. Determining edge expansion and other connectivity measures of graphs of bounded genus. SIAM Journal on Computing, 42(3):1113– 1131, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Paul:1977:LBC

[Pau77] Wolfgang J. Paul. A 2.5nlower bound on the combinational complexity of Boolean functions. SIAM Journal on Computing, 6(3):427–443, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Provan:1983:CCC

[PB83a] J. Scott Provan and Michael O.

Ball. The complexity of counting cuts and of computing the probability that a graph is connected. SIAM Journal on Computing, 12(4):777– 788, November 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Purdom:1983:ABS

[PB83b] Paul Walton Purdom, Jr. and Cynthia A. Brown. An analysis of backtracking with search rearrangement. SIAM Journal on Computing, 12(4):717–733, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Purdom:1985:PLR

[PB85] Paul Walton Purdom, Jr. and Cynthia A. Brown. The pure literal rule and polynomial average time. SIAM Journal on Computing, 14(4):943–953, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Parikh:1985:EBR

[PCHM85] Rohit Parikh, Ashok Chandra, Joe Halpern, and Albert Meyer. Equations between regular terms and an application to process logic. SIAM Journal on Computing, 14(4):935– 942, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Patrascu:2006:LLB

[PD06] Mihai Patrascu and Erik D. Demaine. Logarithmic lower bounds in the cell-probe model. SIAM Journal on Computing, 35(4):932–963, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pang:1986:CCC

[PE86] King F. Pang and Abbas El Gamal. Communication complexity of computing the Hamming distance. SIAM Journal on Computing, 15(4):932–947, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pellegrini:1994:CFP

[Pel94] Marco Pellegrini. On collisionfree placements of simplices and the closest pair of lines in 3-space. SIAM Journal on Computing, 23(1):133– 153, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        22940.

Pellegrini:1996:PLM

[Pel96] Marco Pellegrini. On point location and motion planning among simplices. SIAM Journal on Computing, 25 (5):1061–1081, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25688.

Perlman:1974:IRP

[Per74] David M. Perlman. Isomorph rejection on power sets. SIAM Journal on Computing, 3(3): 177–183, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Peterson:1983:TEC

[Pet83] Gerald E. Peterson. A technique for establishing completeness results in theorem proving with equality. SIAM Journal on Computing, 12(1): 82–100, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pfaltz:1983:CCD

[Pfa83] John L. Pfaltz. Convex clusters in a discrete m-dimensional space. SIAM Journal on Computing, 12(4):734–750, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Purdom:1997:POB

[PH97] Paul Walton Purdom, Jr. and G. Neil Haven. Probe order backtracking. SIAM Journal on Computing, 26(2):456–483, April 1997. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25605.

Park:2009:PT

[PHNS09] Gahyun Park, Hsien-Kuei Hwang, Pierre Nicodème, and Wojciech Szpankowski. Profiles of tries. SIAM Journal on Computing, 38(5):1821–1880, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pigozzi:1991:ETI

[Pig91] Don Pigozzi. Equality-Test and If–Then–Else algebras: Axiomatization and specification. SIAM Journal on Computing, 20(4):766–805, August 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Piotrow:2004:DOS

[Pio04] Marek Piotrów. Depth optimal sorting networks resistant to k passive faults. SIAM Journal on Computing, 33(6):1484–1512, December 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35430.

Pippenger:1977:S

[Pip77] Nicholas Pippenger. Superconcentrators. SIAM Journal on Computing, 6(2):298–304, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pippenger:1978:GC

[Pip78] Nicholas Pippenger. Generalized connectors. SIAM Journal on Computing, 7(4):510– 514, ???? 1978. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pippenger:1980:EPM

[Pip80]  Nicholas Pippenger.  On

the evaluation of powers and monomials. SIAM Journal on Computing, 9(2):230–250, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pippenger:1987:SSR

[Pip87] Nicholas Pippenger. Sorting and selecting in rounds. SIAM Journal on Computing, 16(6): 1032–1038, December 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pippenger:1991:SN

[Pip91] Nicholas Pippenger. Selection networks. SIAM Journal on Computing, 20(5):878–887, October 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Peikert:2013:SSF

[PKS^{+}13] Chris Peikert, Robert Kleinberg, Aravind Srinivasan, Alan Frieze, Alexander Russell, and Leonard Schulman. Special section on the Forty-Second Annual ACM Symposium on Theory of Computing (STOC 2010). SIAM Journal on Computing, 42(3):1216–1217, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Plaisted:1978:SPI

[Pla78] David Plaisted. Some polynomial and integer divisibility problems are NP -hard. SIAM Journal on Computing, 7(4): 458–464, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Plaisted:1980:AMP

[Pla80] David A. Plaisted. The application of multivariate polynomials to inference rules and partial tests for unsatisfiability. SIAM Journal on Computing, 9(4):698–705, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Plaisted:1990:HAS

[Pla90] David A. Plaisted. A heuristic algorithm for small separators in arbitrary graphs. SIAM Journal on Computing, 19(2): 267–280, April 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Plotkin:1976:PC

[Plo76] G. D. Plotkin. A powerdomain construction. SIAM Journal on Computing, 5(3):452–487, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Puschel:2003:AAD

[PM03] Markus Püschel and José M. F.

Moura. The algebraic approach to the discrete cosine and sine transforms and their fast algorithms. SIAM Journal on Computing, 32 (5):1280–1316, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39272.

Poljak:1995:ILP

[Pol95] Svatopluk Poljak. Integer linear programs and local search for max-cut. SIAM Journal on Computing, 24(4):822– 839, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
24535.

Ponzio:1998:LBI

[Pon98] Stephen Ponzio. A lower bound for integer multiplication with read-once branching programs. SIAM Journal on Computing, 28(3):798–815, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29034.

Pan:1995:WPS

[PP95] Victor Y. Pan and Franco P. Preparata. Work-preserving speed-up of parallel matrix computations. SIAM Journal on Computing, 24(4):811– 821, August 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        23216.

Pedersen:1997:FSS

[PP97] Torben Pryds Pedersen and Birgit Pfitzmann. Fail-stop signatures. SIAM Journal on Computing, 26(2):291–330, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24557.

Pagh:2008:UHC

[PP08] Anna Pagh and Rasmus Pagh. Uniform hashing in constant time and optimal space. SIAM Journal on Computing, 38(1): 85–96, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pagh:2009:LPC

[PPR09] Anna Pagh, Rasmus Pagh, and Milan Ružić. Linear probing with constant independence. SIAM Journal on Computing, 39(3):1107–1120, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pietracaprina:2000:CDI

[PPS00] Andrea Pietracaprina, Geppino Pucci, and Jop F. Sibeyn. Constructive, deterministic implementation of shared memory on meshes. SIAM Journal on Computing, 30(2):625–648, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32712.

Peer:2004:IDP

[PPSS04] Itsik Pe’er, Tal Pupko, Ron Shamir, and Roded Sharan. Incomplete directed perfect phylogeny. SIAM Journal on Computing, 33(3):590–607, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40651.

Pan:1993:FEP

[PR93] Victor Pan and John Reif.

Fast and efficient parallel solution of sparse linear systems. SIAM Journal on Computing, 22(6):1227–1250, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pruesse:1994:GLE

[PR94] Gara Pruesse and Frank Ruskey. Generating linear extensions fast. SIAM Journal on Computing, 23(2):373–386, April 1994. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/20264.

Parker:1999:CHC

[PR99] D. Stott Parker and Prasad Ram. The construction of Huffman codes is a submodular (“convex”) optimization problem over a lattice of binary trees. SIAM Journal on Computing, 28(5):1875–1905, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31107.

Peleg:2000:NTL

[PR00] David Peleg and Vitaly Rubinovich. A near-tight lower bound on the time complexity of distributed minimumweight spanning tree construction. SIAM Journal on Computing, 30(5):1427–1442, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36974.

Pettie:2002:RTW

[PR02] Seth Pettie and Vijaya Ramachandran. A randomized time-work optimal parallel algorithm for finding a minimum spanning forest. SIAM Journal on Computing, 31(6): 1879–1895, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37106.

Pettie:2005:SPA

[PR05] Seth Pettie and Vijaya Ramachandran. A shortest path algorithm for real-weighted undirected graphs. SIAM Journal on Computing, 34(6): 1398–1431, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41965.

Petersen:2006:ESQ

[PR06] Holger Petersen and John Michael Robson. Efficient simulations by queue machines. SIAM Journal on Computing, 35(5): 1059–1069, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pass:2008:CNC

[PR08a] Rafael Pass and Alon Rosen.

Concurrent nonmalleable commitments. SIAM Journal on Computing, 37(6):1891–1925, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pass:2008:NIC

[PR08b] Rafael Pass and Alon Rosen.

New and improved constructions of nonmalleable cryptographic protocols. SIAM Journal on Computing, 38(2):702– 752, ???? 2008. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Patrascu:2014:DOB

[PR14] Mihai Pătrascu and Liam Roditty. Distance oracles beyond the Thorup–Zwick bound. SIAM Journal on Computing, 43(1):300–311, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Patel:2017:DPT

[PR17] Viresh Patel and Guus Regts. Deterministic polynomial-time approximation algorithms for partition functions and graph polynomials. SIAM Journal on Computing, 46(6):1893–1919, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Plaut:2020:CCD

[PR20] Benjamin Plaut and Tim Roughgarden. Communication complexity of discrete fair division. SIAM Journal on Computing, 49(1):206–243, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pratt:1975:EPS

[Pra75a] Vaughan R. Pratt. Every prime has a succinct certificate. SIAM Journal on Computing, 4(3):214–220, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pratt:1975:PNT

[Pra75b]  Vaughan R. Pratt.  The

power of negative thinking in multiplying Boolean matrices. SIAM Journal on Computing, 4(3):326–330, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Preparata:1979:NLS

[Pre79] F. P. Preparata. A note on locating a set of points in a planar subdivision. SIAM Journal on Computing, 8(4):542– 545, ???? 1979. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Preparata:1981:NAP

[Pre81] Franco P. Preparata. A new approach to planar point location. SIAM Journal on Computing, 10(3):473–482, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Priese:1979:TPC

[Pri79] Lutz Priese. Towards a precise characterization of the complexity of universal and nonuniversal Turing machines. SIAM Journal on Computing, 8(4):508–523, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Prill:1986:AIC

[Pri86] David Prill. On approximations and incidence in cylindrical algebraic decompositions.

SIAM Journal on Computing, 15(4):972–993, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Probert:1976:ACM

[Pro76] Robert L. Probert. On the additive complexity of matrix multiplication. SIAM Journal on Computing, 5(2):187– 203, ???? 1976. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Probert:1978:ECD

[Pro78] Robert L. Probert. An extension of computational duality to sequences of bilinear computations. SIAM Journal on Computing, 7(1):91–98, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Proskurowski:1981:RGR

[Pro81] Andrzej Proskurowski. Recursive graphs, recursive labelings and shortest paths. SIAM Journal on Computing, 10(2): 391–397, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Provan:1986:CRC

[Pro86] J. Scott Provan. The complexity of reliability computations in planar and acyclic graphs. SIAM Journal on Computing, 15(3):694–702, ???? 1986. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Provan:1988:ASF

[Pro88] J. Scott Provan. An approximation scheme for finding Steiner trees with obstacles. SIAM Journal on Computing, 17(5):920–934, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Prodinger:1992:EIN

[Pro92] Helmut Prodinger. External internal nodes in digital search trees via Mellin transforms. SIAM Journal on Computing, 21(6):1180–1183, December 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Parnas:2003:TCS

[PRR03] Michal Parnas, Dana Ron, and Ronitt Rubinfeld. On testing convexity and submodularity. SIAM Journal on Computing, 32(5):1158–1184, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41402.

Pestien:1994:PTN

[PRS94] Victor Pestien, S. Ramakrishnan, and Dilip Sarkar. Packet transmission in a noisy-channel ring network. SIAM Journal on Computing, 23(3):553–562, June 1994. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/16828.

Pudlak:1997:BCT

[PRS97] Pavel Pudlák, Vojtěch Rödl, and Jiří Sgall. Boolean circuits, tensor ranks, and communication complexity. SIAM Journal on Computing, 26(3):605–633, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26480.

Panagiotou:2023:ESS

[PRS23] Konstantinos Panagiotou, Leon Ramzews, and Benedikt Stufler. Exact-size sampling of enriched trees in linear time. SIAM Journal on Computing, 52(5):1097–1131, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1459733.

Paterson:1973:NNM

[PS73] Michael S. Paterson and Larry J. Stockmeyer. On the number of nonscalar multiplications necessary to evaluate polynomials. SIAM Journal on Computing, 2(1):60–66, March 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Cited in [?].

Papadimitriou:1977:CLS

[PS77] Christos H. Papadimitriou and Kenneth Steiglitz. On the complexity of local search for the traveling salesman problem. SIAM Journal on Computing, 6(1):76–83, March 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pnueli:1981:APF

[PS81] Amir Pnueli and Giora Slutzki.

Automatic programming of finite state linear programs. SIAM Journal on Computing, 10(3):519–535, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pach:1991:VVA

[PS91] János Pach and Micha Sharir.

On vertical visibility in arrangements of segments and the queue size in the Bentley– Ottmann line sweeping algorithm. SIAM Journal on Computing, 20(3):460–470, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Panconesi:1997:RDE

[PS97] Alessandro Panconesi and Aravind Srinivasan. Randomized distributed edge coloring via an extension of the Chernoff–Hoeffding bounds. SIAM Journal on Computing, 26(2):350–368, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25076.

Pavan:2002:SNC

[PS02] A. Pavan and Alan L. Selman. Separation of NP-completeness notions. SIAM Journal on Computing, 31(3):906–918, June 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38703.

Pitassi:2012:ELB

[PS12] Toniann Pitassi and Nathan Segerlind. Exponential lower bounds and integrality gaps for tree-like Lovász–Schrijver procedures. SIAM Journal on Computing, 41(1):128–159, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v41/i1/
                                        p128_s1.

Pomerance:1988:PAF

[PST88] Carl Pomerance, J. W. Smith, and Randy Tuler. A pipeline architecture for factoring large integers with the quadratic sieve algorithm. SIAM Journal on Computing, 17(2):387– 403, April 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Pacholski:2000:CRF

[PST00] Leszek Pacholski, Wiesław Szwast, and Lidia Tendera.

Complexity results for firstorder two-variable logic with counting. SIAM Journal on Computing, 29(4):1083– 1117, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
32300.

Palis:1990:OLT

[PSW90] Michael A. Palis, Sunil Shende, and David S. L. Wei. An optimal linear time parallel parser for tree adjoining languages. SIAM Journal on Computing, 19(1):1–31, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Poloczek:2017:GAM

[PSWvZ17] Matthias Poloczek, Georg Schnitger, David P. Williamson, and Anke van Zuylen. Greedy algorithms for the maximum satisfiability problem: Simple algorithms and inapproximability bounds. SIAM Journal on Computing, 46(3):1029– 1061, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Peng:2017:PWC

[PSZ17] Richard Peng, He Sun, and Luca Zanetti. Partitioning well-clustered graphs: Spectral clustering works! SIAM Journal on Computing, 46(2):710– 743, ???? 2017. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Paige:1987:TPR

[PT87a] Robert Paige and Robert E. Tarjan. Three partition refinement algorithms. SIAM Journal on Computing, 16(6): 973–989, December 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Papadimitriou:1987:SST

[PT87b] Christos H. Papadimitriou and John N. Tsitsiklis. On stochastic scheduling with in-tree precedence constraints. SIAM Journal on Computing, 16(1): 1–6, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Preparata:1989:FDP

[PT89] Franco P. Preparata and Roberto Tamassia. Fully dynamic point location in a monotone subdivision. SIAM Journal on Computing, 18(4): 811–830, August 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Preparata:1992:EPL

[PT92] Franco P. Preparata and Roberto Tamassia. Efficient point location in a convex spatial cell-complex. SIAM Journal on Computing, 21(2):267– 280, April 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pach:2002:BCU

[PT02] János Pach and Gábor Tardos.

On the boundary complexity of the union of fat triangles. SIAM Journal on Computing, 31(6):1745–1760, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38216.

Pavan:2007:REC

[PT07] A. Pavan and Srikanta Tirthapura. Range-efficient counting of distinct elements in a massive data stream. SIAM Journal on Computing, 37(2):359– 379, ???? 2007. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Patrascu:2009:HLB

[PT09] Mihai Pătraşcu and Mikkel Thorup. Higher lower bounds for near-neighbor and further rich problems. SIAM Journal on Computing, 39(2):730–741, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pass:2011:CPC

[PTW11] Rafael Pass, Wei-Lung Dustin Tseng, and Douglas Wikström. On the composition of publiccoin zero-knowledge protocols. SIAM Journal on Computing, 40(6):1529–1553, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.

                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i6/p1529_s1.

Papadimitriou:1987:CTT

[PU87] Christos H. Papadimitriou and Jeffrey D. Ullman. A communication-time tradeoff. SIAM Journal on Computing, 16(4):639–646, August 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Peleg:1989:OSH

[PU89a] David Peleg and Jeffrey D. Ullman. An optimal synchronizer for the hypercube. SIAM Journal on Computing, 18(4): 740–747, August 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Peleg:1989:TDP

[PU89b] David Peleg and Eli Upfal. The token distribution problem. SIAM Journal on Computing, 18(2):229–243, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Peleg:1990:TRT

[PU90] David Peleg and Eli Upfal. A time-randomness trade-off for oblivious routing. SIAM Journal on Computing, 19(2):256– 266, April 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pullman:1984:CCG

[Pul84] Norman J. Pullman. Clique covering of graphs. IV. algorithms. SIAM Journal on Computing, 13(1):57–75, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Purdom:1978:TSP

[Pur78] Paul W. Purdom. Tree size by partial backtracking. SIAM Journal on Computing, 7(4): 481–491, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Purdom:2004:ACP

[PVG04] Paul W. Purdom, Dirk Van Gucht, and Dennis P. Groth. Average-case performance of the apriori algorithm. SIAM Journal on Computing, 33 (5):1223–1260, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42288.

Paredaens:1998:FOQ

[PVV98] Jan Paredaens, Jan Van den Bussche, and Dirk Van Gucht. First-order queries on finite structures over the reals. SIAM Journal on Computing, 27(6): 1747–1763, December 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29796.

Phillips:2016:LBN

[PVZ16] Jeff M. Phillips, Elad Verbin, and Qin Zhang. Lower bounds for number-in-hand multiparty communication complexity, made easy. SIAM Journal on Computing, 45(1):174–196, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pan:2004:RNR

[PW04] Victor Y. Pan and Xinmao Wang. On rational number reconstruction and approximation. SIAM Journal on Computing, 33(2):502–503, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43718.

Peikert:2011:LTF

[PW11] Chris Peikert and Brent Waters. Lossy trapdoor functions and their applications. SIAM Journal on Computing, 40(6): 1803–1844, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i6/p1803_s1.

Papadimitriou:1979:SIO

[PY79] C. H. Papadimitriou and M. Yannakakis. Scheduling interval-ordered tasks. SIAM Journal on Computing, 8(3): 405–409, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Papadimitriou:1987:CRC

[PY87] Christos H. Papadimitriou and Mihalis Yannakakis. The complexity of reliable concurrency control. SIAM Journal on Computing, 16(3):538–553, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Pittel:1988:STE

[PY88] Boris Pittel and Jenn-Hwa Yu. On search times for early-insertion coalesced hashing. SIAM Journal on Computing, 17(3):492–503, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Papadimitriou:1990:TAI

[PY90] Christos H. Papadimitriou and Mihalis Yannakakis. Towards an architecture-independent analysis of parallel algorithms. SIAM Journal on Computing, 19(2):322–328, April 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Parberry:1991:IUL

[PY91] Ian Parberry and Pei Yuan Yan. Improved upper and lower time bounds for parallel random access machines without simultaneous writes. SIAM Journal on Computing, 20(1): 88–99, February 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Queyranne:2006:ABG

[QS06] Maurice Queyranne and Andreas S. Schulz. Approximation bounds for a general class of precedence constrained parallel machine scheduling problems. SIAM Journal on Computing, 35(5):1241–1253, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Qin:2021:NGN

[QY21] Minglong Qin and Penghui Yao. Nonlocal games with noisy maximally entangled states are decidable. SIAM Journal on Computing, 50(6): 1800–1891, ???? 2021. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reinhardt:2000:MNU

[RA00] Klaus Reinhardt and Eric Allender. Making nondeterminism unambiguous. SIAM Journal on Computing, 29 (4):1118–1131, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33904.

Rabin:1980:PAF

[Rab80] Michael O. Rabin. Probabilistic algorithms in finite fields. SIAM Journal on Computing, 9(2):273–280, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rajasekaran:1996:TAL

[Raj96] Sanguthevar Rajasekaran. Treeadjoining language parsing in O(n^{6}) time. SIAM Journal on Computing, 25(4):862– 873, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
26009.

Ramanan:1984:PPC

[Ram84] Prakash V. Ramanan. Pushdown permuter characterization theorem. SIAM Journal on Computing, 13(1):167–169, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ramanan:1994:NLB

[Ram94] Prakash Ramanan. A new lower bound technique and its application: Tight lower bound for a polygon triangulation problem. SIAM Journal on Computing, 23(4):834– 851, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
19007.

Ramanan:1996:EPA

[Ram96] Prakash Ramanan. An efficient parallel algorithm for the matrix-chain-product problem. SIAM Journal on Computing, 25(4):874–893, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rao:2009:ECN

[Rao09] Anup Rao. Extractors for a constant number of polynomially small min-entropy independent sources. SIAM Journal on Computing, 39(1):168– 194, ???? 2009. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rao:2011:PRP

[Rao11] Anup Rao. Parallel repetition in projection games and a concentration bound. SIAM Journal on Computing, 40(6): 1871–1891, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i6/p1871_s1.

Rappaport:1989:CSC

[Rap89] David Rappaport. Computing simple circuits from a set of line segments is NP-complete. SIAM Journal on Computing, 18(6):1128–1139, December 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Raz:1998:PRT

[Raz98] Ran Raz. A parallel repetition theorem. SIAM Journal on Computing, 27(3):763–803, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28089.

Raz:2003:CMP

[Raz03]  Ran Raz.  On the com-

plexity of matrix product. SIAM Journal on Computing, 32(5):1356–1369, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40214.

Raz:2011:CSP

[Raz11] Ran Raz. A counterexample to strong parallel repetition. SIAM Journal on Computing, 40(3):771–777, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i3/p771_s1.

Roth:1991:IAS

[RB91] Ron M. Roth and Gyora M.

Benedek. Interpolation and approximation of sparse multivariate polynomials over GF(2). SIAM Journal on Computing, 20(2):291–314, April 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rothstein:1979:STE

[RC79] Michael Rothstein and B. F.

Caviness. A structure theorem for exponential and primitive functions. SIAM Journal on Computing, 8(3):357–367, August 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rudin:2003:IBO

[RC03] John F. Rudin III and R. Chandrasekaran. Improved bounds for the online scheduling problem. SIAM Journal on Computing, 32(3):717–735, June 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40343.

Rowland:1981:STD

[RD81] John H. Rowland and Philip J. Davis. On the selection of test data for recursive mathematical subroutines. SIAM Journal on Computing, 10(1):59– 72, ???? 1981. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Regan:1996:LTM

[Reg96] Kenneth W. Regan. Linear time and memory-efficient computation. SIAM Journal on Computing, 25(1):133– 168, February 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        25188.

Regev:2004:QCL

[Reg04] Oded Regev. Quantum computation and lattice problems. SIAM Journal on Computing, 33(3):738–760, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44067.

Reingold:1972:IPT

[Rei72] Edward M. Reingold. Infix to prefix translation: The insufficiency of a pushdown stack. SIAM Journal on Computing, 1(4):350–353, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1980:CM

[Rei80] John H. Reif. Code motion.

SIAM Journal on Computing, 9(2):375–395, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1983:MCP

[Rei83] John H. Reif. Minimum st cut of a planar undirected network in O(n log^{2}(n)) time. SIAM Journal on Computing, 12(1):71–81, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1984:SPC

[Rei84] John H. Reif. On synchronous parallel computations with independent probabilistic choice. SIAM Journal on Computing, 13(1):46–56, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reischuk:1985:PPA

[Rei85] Rüdiger Reischuk. Probabilistic parallel algorithms for sorting and selection. SIAM Journal on Computing, 14(2):396– 409, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1986:LDC

[Rei86]  John H. Reif.  Logarith-

mic depth circuits for algebraic functions. SIAM Journal on Computing, 15(1):231–242, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1999:ACP

[Rei99] John H. Reif. Approximate complex polynomial evaluation in near constant work per point. SIAM Journal on Computing, 28(6):2059–2089, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        32429.

Reichardt:2014:SPE

[Rei14] Ben W. Reichardt. Span programs are equivalent to quantum query algorithms. SIAM Journal on Computing, 43(3): 1206–1219, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Renegar:1989:WCA

[Ren89] James Renegar. On the worstcase arithmetic complexity of approximating zeros of systems of polynomials. SIAM Journal on Computing, 18(2):350–370, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Renegar:1992:CCA

[Ren92] James Renegar. On the computational complexity of approximating solutions for real algebraic formulae. SIAM Journal on Computing, 21(6): 1008–1025, December 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Revah:1975:NMD

[Rev75] L. Revah. On the number of multiplications/divisions evaluating a polynomial with auxiliary functions. SIAM Journal on Computing, 4(3):381– 392, ???? 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Revesz:1985:ATL

[Rév́85] György Rév́esz. Axioms for the theory of lambda-conversion. SIAM Journal on Computing, 14(2):373–382, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reyner:1977:AGA

[Rey77] Steven W. Reyner. An analysis of a good algorithm for the subtree problem. SIAM Journal on Computing, 6(4):730– 732, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosenthal:1977:SAB

[RG77] Arnie Rosenthal and Anita Goldner. Smallest augmentations to biconnect a graph. SIAM Journal on Computing, 6(1):55–66, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ruskey:1977:GBT

[RH77] F. Ruskey and T. C. Hu. Generating binary trees lexicographically. SIAM Journal on Computing, 6(4):745–758, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosenkrantz:1978:PAD

[RH78] Daniel J. Rosenkrantz and Harry B. Hunt, III. Polynomial algorithms for deterministic pushdown automata. SIAM Journal on Computing, 7(4): 405–412, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosenkrantz:1993:CPH

[RH93] Daniel J. Rosenkrantz and Harry B. Hunt, III. The complexity of processing hierarchical specifications. SIAM Journal on Computing, 22(3):627– 649, June 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rhee:1990:NOP

[Rhe90] WanSoo T. Rhee. A note on optimal bin packing and optimal bin covering with items of random size. SIAM Journal on Computing, 19(4):705–710, August 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ravikumar:1989:RTA

[RI89] Bala Ravikumar and Oscar H.

Ibarra. Relating the type of ambiguity of finite automata to the succinctness of their representation. SIAM Journal on Computing, 18(6):1263– 1282, December 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rieger:1999:PAA

[Rie99]  J. H. Rieger.  Proximity

in arrangements of algebraic sets. SIAM Journal on Computing, 29(2):433–458, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30094. See cor-

rigendum [Rie02].

Rieger:2002:CPA

[Rie02] J. H. Rieger. Corrigendum:

Proximity in arrangements of algebraic sets. SIAM Journal on Computing, 31(3):987, June 2002. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38020. See [?].

Risler:1985:ACZ

[Ris85] J. J. Risler. Additive complexity and zeros of real polynomials. SIAM Journal on Computing, 14(1):178–183, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rivest:1976:PMR

[Riv76] Ronald L. Rivest. Partial match retrieval algorithms. SIAM Journal on Computing, 5(1):19–50, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rivest:1977:WCB

[Riv77] Ronald L. Rivest. On the worst-case behavior of stringsearching algorithms. SIAM Journal on Computing, 6(4): 669–674, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rubin:2012:IBG

[RKS12] Natan Rubin, Haim Kaplan, and Micha Sharir. Improved bounds for geometric permutations. SIAM Journal on Computing, 41(2):367–390, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ravi:1988:CNO

[RL88] S. S. Ravi and Errol L.

Lloyd. The complexity of nearoptimal programmable logic array folding. SIAM Journal on Computing, 17(4):696–710, August 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rhee:1996:TAF

[RLDL96] C. Rhee, Y. D. Liang, S. K.

Dhall, and S. Lakshmivarahan. An O(N + M)-time algorithm for finding a minimum-weight dominating set in a permutation graph. SIAM Journal on Computing, 25(2):404–419, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/20038.

Ripphausen-Lipa:1997:VDM

[RLWW97] Heike Ripphausen-Lipa, Dorothea Wagner, and Karsten Weihe. The vertex-disjoint Menger problem in planar graphs. SIAM Journal on Computing, 26(2):331–349, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25356.

Robson:1984:CEC

[Rob84] J. M. Robson. N by N checkers is Exptime complete. SIAM Journal on Computing, 13(2): 252–267, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Roditty:2010:SSP

[Rod10] Liam Roditty. On the k shortest simple paths problem in weighted directed graphs. SIAM Journal on Computing, 39(6):2363–2376, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Romani:1982:SPD

[Rom82] Francesco Romani. Some properties of disjoint sums of tensors related to matrix multiplication. SIAM Journal on Computing, 11(2):263–267, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ronse:1982:OCD

[Ron82] C. Ronse. Optimization of cost and delay in cellular permutation networks. SIAM Journal on Computing, 11(4):784–788, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosenberg:1972:SDG

[Ros72] Arnold L. Rosenberg. Symmetries in data graphs. SIAM Journal on Computing, 1(1): 40–65, ???? 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosenberg:1975:MSE

[Ros75a] Arnold L. Rosenberg. Managing storage for extendible arrays. SIAM Journal on Computing, 4(3):287–306, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosenberg:1975:PPA

[Ros75b] Arnold L. Rosenberg. Preserving proximity in arrays. SIAM Journal on Computing, 4(4): 443–460, December 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosen:1980:MRD

[Ros80] Barry K. Rosen. Monoids for rapid data flow analysis. SIAM Journal on Computing, 9(1): 159–196, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosen:1982:LDF

[Ros82a] Barry K. Rosen. A lubricant for data flow analysis. SIAM Journal on Computing, 11(3): 493–511, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosenthal:1982:DPO

[Ros82b] Arnon Rosenthal. Dynamic programming is optimal for nonserial optimization problems. SIAM Journal on Computing, 11(1):47–59, ????

1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rossman:2014:MCC

[Ros14]  Benjamin Rossman.  The

monotone complexity of kclique on random graphs. SIAM Journal on Computing, 43(1):256–279, ???? 2014. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rossman:2018:FVC

[Ros18] Benjamin Rossman. Formulas versus circuits for small distance connectivity. SIAM Journal on Computing, 47(5): 1986–2028, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rothvoss:2016:BPA

[Rot16]  Thomas Rothvoss.  Better

bin packing approximations via discrepancy theory. SIAM Journal on Computing, 45(3): 930–946, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rothvoss:2017:CDM

[Rot17] Thomas Rothvoss. Constructive discrepancy minimization for convex sets. SIAM Journal on Computing, 46(1):224–234, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Roughgarden:2004:SSS

[Rou04] Tim Roughgarden. Stackelberg scheduling strategies. SIAM Journal on Computing, 33(2):332–350, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39705.

Roughgarden:2015:SSF

[Rou15] Tim Roughgarden. Special section on the Fifty-Third IEEE Annual Symposium on Foundations of Computer Science (FOCS 2012). SIAM Journal on Computing, 44(5):1286, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rajasekaran:1989:OST

[RR89] Sanguthevar Rajasekaran and John H. Reif. Optimal and sublogarithmic time randomized parallel sorting algorithms. SIAM Journal on Computing, 18(3):594–607, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rao:2004:NAT

[RR04] Satish Rao and Andréa W.

Richa. New approximation techniques for some linear ordering problems. SIAM Journal on Computing, 34(2):388– 404, April 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41319.

Reingold:2021:CRI

[RRR21] Omer Reingold, Guy N. Rothblum, and Ron D. Rothblum. Constant-round interactive proofs for delegating computation. SIAM Journal on Computing, 50(3):STOC16– 255–STOC16–340, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Raskhodnikova:2009:SLB

[RRSS09] Sofya Raskhodnikova, Dana Ron, Amir Shpilka, and Adam Smith. Strong lower bounds for approximating distribution support size and the distinct elements problem. SIAM Journal on Computing, 39(3):813– 842, ???? 2009. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rodl:2001:MMQ

[RRW01] Vojtech Rödl, Andrzej Rucinski, and Michelle Wagner. Matchings meeting quotas and their impact on the blowup lemma. SIAM Journal on Computing, 31(2):428–446, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37105.

Rosenberg:1978:MCT

[RS78] Arnold L. Rosenberg and Lawrence Snyder. Minimalcomparison 2, 3-trees. SIAM Journal on Computing, 7(4): 465–480, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rackoff:1981:LSN

[RS81] Charles W. Rackoff and Joel I.

Seiferas. Limitations on separating nondeterministic complexity classes. SIAM Journal on Computing, 10(4):742–745, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rowland:1983:STD

[RS83] John H. Rowland and Leslie E.

Shader. On the selection of test data for vector-valued recursive subroutines. SIAM Journal on Computing, 12(3):526– 538, ???? 1983. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reeds:1985:SRS

[RS85a] J. A. Reeds and N. J. A.

Sloane. Shift register synthesis (modulo m). SIAM Journal on Computing, 14(3):505–513, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1985:USV

[RS85b] John H. Reif and Paul G. Spirakis. Unbounded speed variability in distributed communications systems. SIAM Journal on Computing, 14(1):75– 92, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reingold:1991:MNOb

[RS91a] Edward M. Reingold and Xiao Jun Shen. More nearly optimal algorithms for unbounded searching. II. the transfinite case. SIAM Journal on Computing, 20(1):184–208, February 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reingold:1991:MNOa

[RS91b] Edward M. Reingold and Xiao Jun Shen. More nearly optimal algorithms for unbounded searching, Part I: The finite case. SIAM Journal on Computing, 20(1):156–183, February 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reutenauer:1991:MRW

[RS91c] Christophe Reutenauer and Marcel-Paul Schützenberger. Minimization of rational word functions. SIAM Journal on Computing, 20(4):669–685, August 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1992:OPR

[RS92] John H. Reif and Sandeep Sen. Optimal parallel randomized algorithms for threedimensional convex hulls and related problems. SIAM Journal on Computing, 21(3):466– 485, June 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [RS94a].

Reif:1994:EOP

[RS94a] John H. Reif and Sandeep Sen. Erratum: “Optimal parallel randomized algorithms for three-dimensional convex hulls and related problems” [SIAM J. Comput. 21 (1992), no. 3, 466–485, MR 93f:68038]. SIAM Journal on Computing, 23(2):447–448, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [RS92].

Reif:1994:RAB

[RS94b] John H. Reif and Sandeep Sen.

Randomized algorithms for binary search and load balancing on fixed connection networks with geometric applications. SIAM Journal on Computing, 23(3):633–651, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18429.

Rubinfeld:1996:RCP

[RS96] Ronitt Rubinfeld and Madhu Sudan. Robust characterizations of polynomials with applications to program testing. SIAM Journal on Computing, 25(2):252–271, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25515.

Ruskey:1999:EAG

[RS99] Frank Ruskey and Joe Sawada. An efficient algorithm for generating necklaces with fixed density. SIAM Journal on Computing, 29(2):671–684, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34411.

Rajagopalan:2000:VI

[RS00] Sridhar Rajagopalan and Leonard J. Schulman. Verification of identities. SIAM Journal on Computing, 29 (4):1155–1163, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32538.

Raz:2003:LBM

[RS03a] Ran Raz and Amir Shpilka. Lower bounds for matrix product in bounded depth circuits with arbitrary gates. SIAM Journal on Computing, 32(2):488–513, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38462.

Reif:2003:FMS

[RS03b] John H. Reif and Zheng Sun. On frictional mechanical systems and their computational power. SIAM Journal on Computing, 32(6):1449–1474, December 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34665.

Roy:2007:DLG

[RS07] Amitabha Roy and Howard Straubing. Definability of languages by generalized firstorder formulas over (N, +). SIAM Journal on Computing, 37(2):502–521, ???? 2007. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rubin:2008:CFF

[RS08] K. Rubin and A. Silverberg.

Compression in finite fields and torus-based cryptography. SIAM Journal on Computing, 37(5):1401–1428, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rabani:2010:ECS

[RS10a] Yuval Rabani and Amir Shpilka. Explicit construction of a small ϵ-net for linear threshold functions. SIAM Journal on Computing, 39(8):3501– 3520, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Razborov:2010:SRA

[RS10b] Alexander A. Razborov and Alexander A. Sherstov. The sign-rank of AC^{0}. SIAM Journal on Computing, 39(5):1833– 1855, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosen:2010:CCS

[RS10c] Alon Rosen and Gil Segev. Chosen-ciphertext security via correlated products. SIAM Journal on Computing, 39(7): 3058–3088, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rabani:2022:CEC

[RS22] Yuval Rabani and Amir Shpilka. Corrigendum: Explicit construction of a small epsilonnet for linear threshold functions. SIAM Journal on Computing, 51(5):1692–1702, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1310321.

Rosenkrantz:1977:ASH

[RSL77] Daniel J. Rosenkrantz, Richard E. Stearns, and Philip M. Lewis, II. An analysis of several heuristics for the traveling salesman problem. SIAM Journal on Computing, 6(3): 563–581, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosenkrantz:1984:CSC

[RSL84] D. J. Rosenkrantz, R. E.

Stearns, and P. M. Lewis II. Consistency and serializability in concurrent database systems. SIAM Journal on Computing, 13(3):508–530, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reingold:2006:ERR

[RSW06] Omer Reingold, Ronen Shaltiel, and Avi Wigderson. Extracting randomness via repeated condensing. SIAM Journal on Computing, 35(5):1185–1209, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Raz:2008:LBS

[RSY08] Ran Raz, Amir Shpilka, and Amir Yehudayoff. A lower bound for the size of syntactically multilinear arithmetic circuits. SIAM Journal on Computing, 38(4):1624–1647, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Russell:2002:LBL

[RSZ02] Alexander Russell, Michael Saks, and David Zuckerman. Lower bounds for leader election and collective coinflipping in the perfect information model. SIAM Journal on Computing, 31(6): 1645–1662, December 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37600.

Reingold:1981:GHC

[RT81] Edward M. Reingold and Robert E. Tarjan. On a greedy heuristic for complete matching. SIAM Journal on Computing, 10(4):676–681, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1982:SPA

[RT82] John H. Reif and Robert E. Tarjan. Symbolic program analysis in almost-linear time. SIAM Journal on Computing, 11(1):81–93, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1988:EPP

[RT88] J. H. Reif and J. D. Tygar. Efficient parallel pseudorandom number generation. SIAM Journal on Computing, 17(2): 404–411, April 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Special issue on cryptography.

Rhee:1989:CCB

[RT89a] Wansoo T. Rhee and Michel Talagrand. The complete convergence of best fit decreasing. SIAM Journal on Computing, 18(5):909–918, October 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rhee:1989:CCF

[RT89b] Wansoo T. Rhee and Michel Talagrand. The complete convergence of first fit decreasing. SIAM Journal on Computing, 18(5):919–938, October 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rhee:1989:OCI

[RT89c] WanSoo T. Rhee and Michel Talagrand. Optimal bin covering with items of random size. SIAM Journal on Computing, 18(3):487–498, June 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rhee:1989:OPIa

[RT89d] Wansoo T. Rhee and Michel Talagrand. Optimal bin packing with items of random sizes. II. SIAM Journal on Computing, 18(1):139–151, February 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rhee:1989:OPIb

[RT89e] WanSoo T. Rhee and Michel Talagrand. Optimal bin packing with items of random sizes. III. SIAM Journal on Computing, 18(3):473–486, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1990:OSI

[RT90] John H. Reif and Stephen R.

Tate. Optimal size integer division circuits. SIAM Journal on Computing, 19(5):912–924, October 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Reif:1992:TCP

[RT92] John H. Reif and Stephen R. Tate. On threshold circuits and polynomial computation. SIAM Journal on Computing, 21(5):896–908, October 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rhee:1993:LPI

[RT93] WanSoo T. Rhee and Michel Talagrand. On-line bin packing of items of random sizes, II. SIAM Journal on Computing, 22(6):1251–1256, December 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rhee:1996:CDB

[RT96] Wansoo T. Rhee and Michel Talagrand. Convergence in distribution for best-fit decreasing. SIAM Journal on Computing, 25(4):894–906, August 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26718.

Regev:2010:SQC

[RT10] Oded Regev and Ben Toner. Simulating quantum correlations with finite communication. SIAM Journal on Computing, 39(4):1562–1580, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rose:1976:AAV

[RTL76] Donald J. Rose, R. Endre Tarjan, and George S. Lueker. Algorithmic aspects of vertex elimination on graphs. SIAM Journal on Computing, 5(2): 266–283, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Raiha:1981:MNE

[RU81] Kari-Jouko Räihä and Esko Ukkonen. Minimizing the number of evaluation passes for attribute grammars. SIAM Journal on Computing, 10(4):772– 786, November 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See also: report C-1979-121, —HELDCS— (1979).

Raghavan:1998:SCR

[RU98] Prabhakar Raghavan and Eli Upfal. Stochastic contention resolution with short delays. SIAM Journal on Computing, 28(2):709–719, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28533.

Rubinstein:1991:SPP

[Rub91]  Roy S. Rubinstein.  Self-

P-printability and polynomial time Turing equivalence to a tally set. SIAM Journal on Computing, 20(6):1021–1033, December 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rubinfeld:1999:RFE

[Rub99] Ronitt Rubinfeld. On the robustness of functional equations. SIAM Journal on Computing, 28(6):1972–1997, December 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        29862.

Rubinstein:2018:INE

[Rub18] Aviad Rubinstein. Inapproximability of Nash equilibrium. SIAM Journal on Computing, 47(3):917–959, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ruppert:2000:DCN

[Rup00] Eric Ruppert. Determining consensus numbers. SIAM Journal on Computing, 30 (4):1156–1168, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32943.

Russell:1977:ICI

[Rus77a] Bruce D. Russell. Implementation correctness involving a language with goto statements.

SIAM Journal on Computing, 6(3):403–415, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Russell:1977:IFC

[Rus77b] David L. Russell. Internal fragmentation in a class of buddy systems. SIAM Journal on Computing, 6(4):607–621, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ruskey:1978:GAT

[Rus78] Frank Ruskey. Generating t-ary trees lexicographically. SIAM Journal on Computing, 7(4):424–439, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ruskey:1981:LCS

[Rus81] Frank Ruskey. Listing and counting subtrees of a tree. SIAM Journal on Computing, 10(1):141–150, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rutten:1990:SCP

[Rut90] J. J. M. M. Rutten. Semantic correctness for a parallel object-oriented language. SIAM Journal on Computing, 19(2):341–383, April 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rajagopalan:1998:PDR

[RV98] Sridhar Rajagopalan and Vijay V. Vazirani. Primaldual RNC approximation algorithms for set cover and covering integer programs. SIAM Journal on Computing, 28(2):525–540, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26076.

Roychowdhury:2001:QFL

[RV01] Vwani P. Roychowdhury and Farrokh Vatan. Quantum formulas: a lower bound and simulation. SIAM Journal on Computing, 31(2):460–476, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37096.

Reif:2000:NDK

[RW00] John H. Reif and Hongyan Wang. Nonuniform discretization for kinodynamic motion planning and its applications. SIAM Journal on Computing, 30(1):161–190, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33197.

Rhee:2003:ITK

[RW03] Injong Rhee and Jennifer L. Welch. The impact of timing knowledge on the session problem. SIAM Journal on Computing, 32(4):1007–1039, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36094.

Rasala:2005:SNW

[RW05] April Rasala and Gordon Wilfong. Strictly nonblocking WDM cross-connects. SIAM Journal on Computing, 35(2): 449–485, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rosier:1987:LHP

[RY87] Louis E. Rosier and Hsu-Chun Yen. Logspace hierarchies, polynomial time and the complexity of fairness problems concerning ω-machines. SIAM Journal on Computing, 16(5): 779–807, October 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ramachandran:1996:EPA

[RY96] Vijaya Ramachandran and Honghua Yang. An efficient parallel algorithm for the general planar monotone circuit value problem. SIAM Journal on Computing, 25(2):312–339, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26077.

Rytter:1980:CPA

[Ryt80] Wojciech Rytter. A correct preprocessing algorithm for Boyer–Moore string-searching. SIAM Journal on Computing, 9(3):509–512, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Roditty:2008:IDR

[RZ08] Liam Roditty and Uri Zwick. Improved dynamic reachability algorithms for directed graphs. SIAM Journal on Computing, 37(5):1455–1471, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Rao:2010:EDP

[RZ10] Satish Rao and Shuheng Zhou. Edge disjoint paths in moderately connected graphs. SIAM Journal on Computing, 39(5): 1856–1887, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Roditty:2012:DAA

[RZ12] Liam Roditty and Uri Zwick. Dynamic approximate allpairs shortest paths in undirected graphs. SIAM Journal on Computing, 41(3):670–683, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Roditty:2016:FDR

[RZ16] Liam Roditty and Uri Zwick. A fully dynamic reachability algorithm for directed graphs with an almost linear update time. SIAM Journal on Computing, 45(3):712–733, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Safra:2006:EDA

[Saf06] Shmuel Safra. Exponential determinization for ω-automata with a strong fairness acceptance condition. SIAM Journal on Computing, 36(3):803–814, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sagiv:1983:QAM

[Sag83] Yehoshua Sagiv. Quadratic algorithms for minimizing joins in restricted relational expressions. SIAM Journal on Computing, 12(2):316–328, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sagiv:1988:BDS

[Sag88] Yehoshua Sagiv. On bounded database schemes and bounded Horn-clause programs. SIAM Journal on Computing, 17(1): 1–22, February 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sahni:1974:CRP

[Sah74] Sartaj Sahni. Computationally related problems. SIAM Journal on Computing, 3(4):262– 279, ???? 1974. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sakkalis:1990:EAD

[Sak90] Takis Sakkalis. The Euclidean algorithm and the degree of the Gauss map. SIAM Journal on Computing, 19(3):538–543, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Santos:1975:SSS

[San75] Eugene S. Santos. Statesplitting for stochastic machines. SIAM Journal on Computing, 4(1):85–96, ???? 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Santhanam:2009:CLB

[San09] Rahul Santhanam. Circuit lower bounds for Merlin– Arthur classes. SIAM Journal on Computing, 39(3):1038– 1061, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Savitch:1973:HMA

[Sav73]  Walter J. Savitch.  How

to make arbitrary grammars look like context-free grammars. SIAM Journal on Computing, 2(3):174–182, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Savage:1974:ACL

[Sav74] John E. Savage. An algorithm for the computation of linear forms. SIAM Journal on Computing, 3(2):150–158, ???? 1974. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sawada:2001:GBC

[Saw01]  Joe Sawada.  Generating

bracelets in constant amortized time. SIAM Journal on Computing, 31(1):259– 268, February 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37703.

Schoning:1984:IRN

[SB84] Uwe Schöning and Ronald V.

Book. Immunity, relativizations, and nondeterminism. SIAM Journal on Computing, 13(2):329–337, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Somma:2013:SGA

[SB13] R. D. Somma and S. Boixo.

Spectral gap amplification. SIAM Journal on Computing, 42(2):593–610, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Segerlind:2004:SLS

[SBI04] Nathan Segerlind, Sam Buss, and Russell Impagliazzo. A switching lemma for small restrictions and lower bounds for k-DNF resolution. SIAM Journal on Computing, 33 (5):1171–1200, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42855.

Sahni:1979:NLS

[SC79a] Sartaj Sahni and Yookun Cho. Nearly on line scheduling of a uniform processor system with release times. SIAM Journal on Computing, 8(2):275–285, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Stockmeyer:1979:PDC

[SC79b] Larry J. Stockmeyer and Ashok K. Chandra. Provably difficult combinatorial games. SIAM Journal on Computing, 8(2):151–174, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schnorr:1978:ATC

[Sch78] C. P. Schnorr. An algorithm for transitive closure with linear expected time. SIAM Journal on Computing, 7(2):127– 133, ???? 1978. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schnorr:1979:BEC

[Sch79] C. P. Schnorr. Bottlenecks and edge connectivity in unsymmetrical networks. SIAM Journal on Computing, 8(2):

265–274, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schonhage:1980:SMM

[Sch80] A. Schönhage. Storage modification machines. SIAM Journal on Computing, 9(3):490– 508, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schnorr:1981:ESD

[Sch81a] C. P. Schnorr. An extension of Strassen’s degree bound. SIAM Journal on Computing, 10(2):371–382, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schonhage:1981:PTM

[Sch81b] A. Schönhage. Partial and total matrix multiplication. SIAM Journal on Computing, 10(3):434–455, August 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Slater:1981:IDT

[SCH81c] P. J. Slater, E. J. Cockayne, and S. T. Hedetniemi. Information dissemination in trees. SIAM Journal on Computing, 10(4):692–701, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Scheinerman:1987:ASF

[Sch87] Edward R. Scheinerman. Almost sure fault tolerance in random graphs. SIAM Journal on Computing, 16(6):1124– 1134, December 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schaback:1988:ESB

[Sch88] R. Schaback. On the expected sublinearity of the Boyer– Moore algorithm. SIAM Journal on Computing, 17(4):648– 658, August 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Scheinerman:1990:ECB

[Sch90] Edward R. Scheinerman. On the expected capacity of binomial and random concentrators. SIAM Journal on Computing, 19(1):156–163, February 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shih:1992:AHC

[SCH92] Wei Kuan Shih, T. C. Chern, and Wen-Lian Hsu. An O(n^{2} log n) algorithm for the Hamiltonian cycle problem on circular-arc graphs. SIAM Journal on Computing, 21(6): 1026–1046, December 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schrijver:1994:FDP

[Sch94] Alexander Schrijver. Finding k disjoint paths in a directed planar graph. SIAM Journal on Computing, 23(4):780– 788, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
22406.

Schmidt:1998:AHS

[Sch98a] Jeanette P. Schmidt. All highest scoring paths in weighted grid graphs and their application to finding all approximate repeats in strings. SIAM Journal on Computing, 27(4):972–992, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28848.

Schrijver:1998:BEC

[Sch98b] Alexander Schrijver. Bipartite edge coloring in O(Δm) time. SIAM Journal on Computing, 28(3):841–846, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29926.

Schachinger:2004:DRC

[Sch04a] Werner Schachinger. Distributional results for costs of partial match queries in asymmetric K-dimensional tries.

SIAM Journal on Computing, 33(4):952–983, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42587.

Schwiegelshohn:2004:PWC

[Sch04b] Uwe Schwiegelshohn. Preemptive weighted completion time scheduling of parallel jobs. SIAM Journal on Computing, 33(6):1280–1308, December 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31501.

Schmidt:2013:CRC

[Sch13] Jens M. Schmidt. Contractions, removals, and certifying 3-connectivity in linear time. SIAM Journal on Computing, 42(2):494–535, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schmidt:2016:MSK

[Sch16] Jens M. Schmidt. Mondshein sequences (a.k.a. (2, 1)-orders). SIAM Journal on Computing, 45(6):1985–2003, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Scott:1976:DTL

[Sco76] Dana Scott. Data types as lattices. SIAM Journal on Computing, 5(3):522–587, ????

1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sellen:2000:PSE

[SCY00] Jürgen Sellen, Joonsoo Choi, and Chee-Keng Yap. Precisionsensitive Euclidean shortest path in 3-space. SIAM Journal on Computing, 29(5):1577– 1595, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
34020.

Siegel:1988:SGG

[SD88] Alan Siegel and Danny Dolev.

Some geometry for general river routing. SIAM Journal on Computing, 17(3):583–605, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sebo:1997:PUG

[Seb97] András Sebö. Potentials in undirected graphs and planar multiflows. SIAM Journal on Computing, 26(2):582–603, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18670.

Sedgewick:1977:QEK

[Sed77] Robert Sedgewick. Quicksort with equal keys. SIAM Journal on Computing, 6(2):240– 267, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sedgewick:1978:DMO

[Sed78] Robert Sedgewick. Data movement in odd-even merging. SIAM Journal on Computing, 7(3):239–272, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Seiferas:1977:LTC

[Sei77] Joel I. Seiferas. Linear-time computation by nondeterministic multidimensional iterative arrays. SIAM Journal on Computing, 6(3):487–504, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Seidl:1990:DEF

[Sei90] Helmut Seidl. Deciding equivalence of finite tree automata. SIAM Journal on Computing, 19(3):424–437, June 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Selman:1978:PTE

[Sel78] Alan L. Selman. Polynomial time enumeration reducibility. SIAM Journal on Computing, 7(4):440–457, ???? 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Selman:1988:NSR

[Sel88] Alan L. Selman. Natural selfreducible sets. SIAM Journal on Computing, 17(5):989–996, October 1988. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sellen:1996:LBG

[Sel96] Jürgen Sellen. Lower bounds for geometrical and physical problems. SIAM Journal on Computing, 25(6): 1231–1253, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24333.

Semaev:2009:SAE

[Sem09] Igor Semaev. Sparse algebraic equations over finite fields. SIAM Journal on Computing, 39(2):388–409, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Senizergues:2005:BPE

[Sén05]  Géraud Sénizergues.  The

bisimulation problem for equational graphs of finite outdegree. SIAM Journal on Computing, 34(5):1025–1106, October 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
37725.

Servedio:2002:PWP

[Ser02] Rocco A. Servedio. Perceptron, winnow, and PAC learning. SIAM Journal on Computing, 31(5):1358–1369, October 2002. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34092.

Sethi:1975:CRA

[Set75] Ravi Sethi. Complete register allocation problems. SIAM Journal on Computing, 4(3): 226–248, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sethi:1976:SGT

[Set76] Ravi Sethi. Scheduling graphs on two processors. SIAM Journal on Computing, 5(1):73–82, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sanchis:1990:EGL

[SF90] Laura A. Sanchis and Mark A. Fulk. On the efficient generation of language instances. SIAM Journal on Computing, 19(2):281–296, April 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Svitkina:2011:SAS

[SF11] Zoya Svitkina and Lisa Fleischer. Submodular approximation: Sampling-based algorithms and lower bounds. SIAM Journal on Computing, 40(6):1715–1737, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.

siam.org/sicomp/resource/
1/smjcat/v40/i6/p1715_s1.

Servedio:2004:ESB

[SG04] Rocco A. Servedio and Steven J.

Gortler. Equivalences and separations between quantum and classical learnability. SIAM Journal on Computing, 33(5):1067–1092, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41291.

Stearns:1985:ECP

[SH85] R. E. Stearns and H. B. Hunt, III. On the equivalence and containment problems for unambiguous regular expressions, regular grammars and finite automata. SIAM Journal on Computing, 14(3):598–611, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Stearns:1996:AMC

[SH96] Richard E. Stearns and Harry B.

Hunt, III. An algebraic model for combinatorial problems. SIAM Journal on Computing, 25(2):448–476, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24300.

Shamir:1979:LTA

[Sha79] Adi Shamir. A linear time algorithm for finding minimum cutsets in reducible graphs. SIAM Journal on Computing, 8(4):645–655, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sharir:1985:ICP

[Sha85] Micha Sharir. Intersection and closest-pair problems for a set of planar discs. SIAM Journal on Computing, 14(2):448–468, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sharir:1987:SPA

[Sha87] Micha Sharir. On shortest paths amidst convex polyhedra. SIAM Journal on Computing, 16(3):561–572, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherstov:2009:SDM

[She09] Alexander A. Sherstov. Separating AC^{0} from depth-2 majority circuits. SIAM Journal on Computing, 38(6):2113– 2129, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherstov:2011:PMM

[She11] Alexander A. Sherstov. The pattern matrix method. SIAM Journal on Computing, 40(6): 1969–2000, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.

siam.org/sicomp/resource/
1/smjcat/v40/i6/p1969_s1.

Sherstov:2012:SDP

[She12] Alexander A. Sherstov. Strong direct product theorems for quantum communication and query complexity. SIAM Journal on Computing, 41(5):1122– 1165, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherstov:2013:ITH

[She13] Alexander A. Sherstov. The intersection of two halfspaces has high threshold degree. SIAM Journal on Computing, 42(6): 2329–2374, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherstov:2016:MCC

[She16] Alexander A. Sherstov. The multiparty communication complexity of set disjointness. SIAM Journal on Computing, 45(4):1450–1489, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherstov:2018:BMP

[She18a] Alexander A. Sherstov. Breaking the Minsky–Papert barrier for constant-depth circuits. SIAM Journal on Computing, 47(5):1809–1857, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherstov:2018:CIC

[She18b] Alexander A. Sherstov. Compressing interactive communication under product distributions. SIAM Journal on Computing, 47(2):367–419, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherstov:2018:PAC

[She18c] Alexander A. Sherstov. The power of asymmetry in constantdepth circuits. SIAM Journal on Computing, 47(6):2362– 2434, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherstov:2020:AP

[She20] Alexander A. Sherstov. Algorithmic polynomials. SIAM Journal on Computing, 49(6): 1173–1231, ???? 2020. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shiloach:1979:MLA

[Shi79a] Yossi Shiloach. A minimum linear arrangement algorithm for undirected trees. SIAM Journal on Computing, 8(1): 15–32, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See correction [EiC17].

Shiloach:1979:MTF

[Shi79b] Yossi Shiloach. Multi-terminal 0–1 flow. SIAM Journal on Computing, 8(3):422–430, ????

1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shiloach:1980:MMC

[Shi80] Yossi Shiloach. A multiterminal minimum cut algorithm for planar graphs. SIAM Journal on Computing, 9(2):219– 224, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sarma:2012:DVH

[SHK^{+}12] Atish Das Sarma, Stephan Holzer, Liah Kor, Amos Korman, Danupon Nanongkai, Gopal Pandurangan, David Peleg, and Roger Wattenhofer. Distributed verification and hardness of distributed approximation. SIAM Journal on Computing, 41(5):1235– 1265, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shokrollahi:1992:OAM

[Sho92] Mohammad Amin Shokrollahi.

Optimal algorithms for multiplication in certain finite fields using elliptic curves. SIAM Journal on Computing, 21(6): 1193–1198, December 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shor:1997:PTA

[Sho97] Peter W. Shor. Polynomialtime algorithms for prime factorization and discrete logarithms on a quantum computer. SIAM Journal on Computing, 26(5):1484–1509, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29317.

Shpilka:2003:LBM

[Shp03] Amir Shpilka. Lower bounds for matrix product. SIAM Journal on Computing, 32 (5):1185–1200, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40595.

Shpilka:2009:IDA

[Shp09] Amir Shpilka. Interpolation of depth-3 arithmetic circuits with two multiplication gates. SIAM Journal on Computing, 38(6):2130–2161, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shmueli:1987:CVT

[SI87] Oded Shmueli and Alon Itai. Complexity of views: Tree and cyclic schemas. SIAM Journal on Computing, 16(1):17– 37, February 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sibeyn:1998:RMS

[Sib98] Jop F. Sibeyn. Row-major sorting on meshes. SIAM Journal on Computing, 28(3): 847–863, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
27011.

Siegel:2004:UCE

[Sie04]  Alan Siegel.  On univer-

sal classes of extremely random constant-time hash functions. SIAM Journal on Computing, 33(3):505–543, June 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38621.

Simons:1983:MSU

[Sim83] Barbara Simons. Multiprocessor scheduling of unit-time jobs with arbitrary release times and deadlines. SIAM Journal on Computing, 12(2):294–299, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Simon:1997:PQC

[Sim97a] Daniel R. Simon. On the power of quantum computation. SIAM Journal on Computing, 26(5):1474–1483, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29863.

Simon:1997:BNE

[Sim97b] Hans Ulrich Simon. Bounds on the number of examples needed for learning functions.

SIAM Journal on Computing, 26(3):751–763, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25918.

Singh:1997:LEC

[Sin97] Gurdip Singh. Leader election in complete networks. SIAM Journal on Computing, 26(3):772–785, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27686.

Sitters:2014:GWF

[Sit14] René Sitters. The generalized work function algorithm is competitive for the generalized 2-server problem. SIAM Journal on Computing, 43(1):96– 125, ???? 2014. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sitters:2021:PTA

[Sit21] René Sitters. Polynomial time approximation schemes for the traveling repairman and other minimum latency problems. SIAM Journal on Computing, 50(5):1580–1602, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Savage:1981:FEP

[SJ81] Carla Savage and Joseph Ja’Ja’. Fast, efficient parallel algorithms for some graph problems. SIAM Journal on Computing, 10(4):682–691, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shi:2005:NTT

[SJ05] Qingmin Shi and Joseph JaJa.

Novel transformation techniques using Q-heaps with applications to computational geometry. SIAM Journal on Computing, 34(6):1474–1492, December 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43572.

Smith:2018:SSF

[SK18] Adam Smith and Jochen Könemann. Special section on the Forty-Sixth Annual ACM Symposium on Theory of Computing (STOC 2014). SIAM Journal on Computing, 47(5): 1807–1808, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Skyum:1976:DTV

[Sky76] Sven Skyum. Decomposition theorems for various kinds of languages parallel in nature. SIAM Journal on Computing, 5(2):284–296, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Seroussi:1980:FSM

[SL80] Gadiel Seroussi and Abraham Lempel. Factorization of symmetric matrices and trace-orthogonal bases in finite fields. SIAM Journal on Computing, 9(4):758–767, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sheu:1994:ELH

[SL94] Ming-Jye Sheu and Timothy J. Long. The extended low hierarchy is an infinite hierarchy. SIAM Journal on Computing, 23(3):488–509, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/21864.

Shih:1996:LSI

[SL96] Wei-Kuan Shih and Jane W. S. Liu. On-line scheduling of imprecise computations to minimize error. SIAM Journal on Computing, 25(5):1105– 1121, October 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        23363.

Shih:1991:ASI

[SLC91] Wei Kuan Shih, Jane W. S. Liu, and Jen-Yao Chung. Algorithms for scheduling imprecise computations with timing constraints. SIAM Journal on Computing, 20(3):537–552, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schwiegelshohn:1998:SSB

[SLW^{+}98] Uwe Schwiegelshohn, Walter Ludwig, Joel L. Wolf, John Turek, and Philip S. Yu. Smart SMART bounds for weighted response time scheduling. SIAM Journal on Computing, 28(1):237–253, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28683.

Smith:1976:AOL

[Smi76] Alan Jay Smith. Analysis of the optimal, look-ahead demand paging algorithms. SIAM Journal on Computing, 5(4):743–757, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Smith:1986:PAD

[Smi86] Justin R. Smith. Parallel algorithms for depth-first searches. I. planar graphs. SIAM Journal on Computing, 15(3):814– 830, ???? 1986. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Strong:1975:RSS

[SMSR75] H. R. Strong, Jr., A. Maggiolo-Schettini, and B. K. Rosen. Recursion structure simplification. SIAM Journal on Computing, 4(3):307–320, ????

1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schulman:2007:PLH

[SMW07] Leonard J. Schulman, Tal Mor, and Yossi Weinstein. Physical limits of heat-bath algorithmic cooling. SIAM Journal on Computing, 36(6):1729– 1747, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Snir:1985:PS

[Sni85] Marc Snir. On parallel searching. SIAM Journal on Computing, 14(3):688–708, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Snyder:1992:ELS

[Sny92] Timothy Law Snyder. On the exact location of Steiner points in general dimension. SIAM Journal on Computing, 21(1): 163–180, February 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Straccia:2009:FPM

[SOAD09] Umberto Straccia, Manuel Ojeda-Aciego, and Carlos V. Damásio. On fixed-points of multivalued functions on complete lattices and their application to generalized logic programs. SIAM Journal on Computing, 38(5):1881–1911, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Solovay:1976:SCR

[Sol76] Robert M. Solovay. On sets cook-reducible to sparse sets. SIAM Journal on Computing, 5(4):646–652, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Soto:2013:MSP

[Sot13] José A. Soto. Matroid secretary problem in the randomassignment model. SIAM Journal on Computing, 42(1): 178–211, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Spira:1975:FUS

[SP75] P. M. Spira and A. Pan. On finding and updating spanning trees and shortest paths. SIAM Journal on Computing, 4(3): 375–380, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Smyth:1982:CTS

[SP82] Michael B. Smyth and Gordon D. Plotkin. The categorytheoretic solution of recursive domain equations. SIAM Journal on Computing, 11(4):761– 783, November 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). Also Report D.A.I. 60, University of Edinburgh, Department of Artificial Intelligence, December 1978.

Sharir:1984:VPP

[SPH84a] Micha Sharir, Amir Pnueli, and Sergiu Hart. Verification of probabilistic programs. SIAM Journal on Computing, 13(2):292–314, ???? 1984. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sherman:1984:IPP

[SPH84b] R. Sherman, A. Pnueli, and D. Harel. Is the interesting part of process logic uninteresting? A translation from PL to PDL. SIAM Journal on Computing, 13(4):825–839, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Spira:1973:NAF

[Spi73] P. M. Spira. A new algorithm for finding all shortest paths in a graph of positive arcs in average time O(n^{2} log^{2} n). SIAM Journal on Computing, 2(1):28–32, ???? 1973. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Spinrad:1985:CPG

[Spi85] Jeremy Spinrad. On comparability and permutation graphs. SIAM Journal on Computing, 14(3):658–670, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shallcross:1998:PIL

[SPLK98] D. F. Shallcross, V. Y. Pan, and Y. Lin-Kriz. Planar integer linear programming is NC equivalent to Euclidean GCD. SIAM Journal on Computing, 27(4):960–971, August 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27684.

Spreen:2002:SWM

[Spr02] Dieter Spreen. Safe weak minimization revisited. SIAM Journal on Computing, 31 (5):1542–1556, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38785.

Supowit:1983:DCH

[SR83] Kenneth J. Supowit and Edward M. Reingold. Divideand-conquer heuristics for minimum weighted Euclidean matching. SIAM Journal on Computing, 12(1):118–143, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Storer:1997:ERO

[SR97] James A. Storer and John H.

Reif. Error-resilient optimal data compression. SIAM Journal on Computing, 26(4):934– 949, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
24078.

Srinivasan:1999:IAG

[Sri99] Aravind Srinivasan. Improved approximation guarantees for packing and covering integer programs. SIAM Journal on Computing, 29(2):648–670, April 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31424.

Srinivasan:2006:ELL

[Sri06] Aravind Srinivasan. An extension of the Lovász Local Lemma, and its applications to integer programming. SIAM Journal on Computing, 36(3): 609–634, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Supowit:1983:TSP

[SRP83] Kenneth J. Supowit, Edward M. Reingold, and David A. Plaisted. The travelling salesman problem and minimum matching in the unit square. SIAM Journal on Computing, 12(1):144–156, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sekar:1995:APM

[SRR95] R. C. Sekar, R. Ramesh, and I. V. Ramakrishnan. Adaptive pattern matching. SIAM Journal on Computing, 24(6): 1207–1234, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24625.

Schmidt:1977:SDU

[SS77a] Erik Meineche Schmidt and Thomas G. Szymanski. Succinctness of descriptions of unambiguous context-free languages. SIAM Journal on Computing, 6(3):547–553, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Solovay:1977:FMC

[SS77b] R. Solovay and V. Strassen.

A fast Monte-Carlo test for primality. SIAM Journal on Computing, 6(1):84–85, March 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [SS78].

Solovay:1978:EFM

[SS78] R. Solovay and V. Strassen.

Erratum: a fast Monte-Carlo test for primality. SIAM Journal on Computing, 7(1):118, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [SS77b].

Schroeppel:1981:ACN

[SS81] Richard Schroeppel and Adi Shamir. A T = O(2^{n/2}), S = O(2^{n/4}) algorithm for certain NP-complete problems. SIAM Journal on Computing, 10(3): 456–464, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sharir:1986:SPP

[SS86a] Micha Sharir and Amir Schorr. On shortest paths in polyhedral spaces. SIAM Journal on Computing, 15(1):193–215, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shub:1986:CCG

[SS86b] M. Shub and S. Smale. Computational complexity: on the geometry of polynomials and a theory of cost. II. SIAM Journal on Computing, 15(1):145– 161, ???? 1986. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schimmler:1989:CNS

[SS89a] Manfred Schimmler and Christoph Starke. A correction network for N-sorters. SIAM Journal on Computing, 18(6):1179– 1187, December 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Steele:1989:WCG

[SS89b] J. Michael Steele and Timothy Law Snyder. Worst-case growth rates of some classical problems of combinatorial optimization. SIAM Journal on Computing, 18(2):278–287, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schmidt:1990:SCO

[SS90] Jeanette P. Schmidt and Alan Siegel. The spatial complexity of oblivious k-probe hash functions. SIAM Journal on Computing, 19(5):775–786, October 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Snyder:1995:PBE

[SS95] Timothy Law Snyder and J. Michael Steele. A priori bounds on the Euclidean traveling salesman. SIAM Journal on Computing, 24(3):665–671, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22677.

Seidel:2005:TAP

[SS05a] Raimund Seidel and Micha Sharir. Top-down analysis of path compression. SIAM Journal on Computing, 34(3): 515–525, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
43908.

Shi:2005:RSA

[SS05b] Weiping Shi and Chen Su. The rectilinear Steiner arborescence problem is NP-complete. SIAM Journal on Computing, 35(3):729–740, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Saks:2010:LMR

[SS10] Michael Saks and C. Seshadhri. Local monotonicity reconstruction. SIAM Journal on Computing, 39(7):2897– 2926, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Saxena:2011:AOR

[SS11a] Nitin Saxena and C. Seshadhri. An almost optimal rank bound for depth-3 identities. SIAM Journal on Computing, 40(1): 200–224, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i1/p200_s1.

Sharir:2011:SRR

[SS11b] Micha Sharir and Hayim Shaul. Semialgebraic range reporting and emptiness searching with applications. SIAM Journal on Computing, 40(4): 1045–1074, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i4/p1045_s1.

Spielman:2011:GSE

[SS11c] Daniel A. Spielman and Nikhil Srivastava. Graph sparsification by effective resistances. SIAM Journal on Computing, 40(6):1913–1926, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i6/p1913_s1.

Saxena:2012:BIT

[SS12a] Nitin Saxena and C. Seshadhri.

Blackbox identity testing for bounded top-fanin depth-3 circuits: The field doesn’t matter. SIAM Journal on Computing, 41(5):1285–1298, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Swamy:2012:SBA

[SS12b] Chaitanya Swamy and David B.

Shmoys. Sampling-based approximation algorithms for multistage stochastic optimization. SIAM Journal on Computing, 41(4):975–1004, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Saks:2017:ELI

[SS17] M. Saks and C. Seshadhri. Estimating the longest increasing sequence in polylogarithmic time. SIAM Journal on Computing, 46(2):774–823, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Singer:1985:ELT

[SSC85] M. F. Singer, B. D. Saunders, and B. F. Caviness. An extension of Liouville’s theorem on integration in finite terms. SIAM Journal on Computing, 14(4):966–990, November 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sutner:1991:CRN

[SSS91] K. Sutner, A. Satyanarayana, and C. Suffel. The complexity of the residual node connectedness reliability problem. SIAM Journal on Computing, 20(1): 149–155, February 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shalev-Shwartz:2011:LKB

[SSSS11] Shai Shalev-Shwartz, Ohad Shamir, and Karthik Sridharan. Learning kernel-based halfspaces with the 0-1 loss. SIAM Journal on Computing, 40(6):1623–1646, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i6/p1623_s1.

Shahrokhi:2001:BDL

[SSSV01] Farhad Shahrokhi, Ondrej Sýkora, László A. Székely, and Imrich Vrto. On bipartite drawings and the linear arrangement problem. SIAM Journal on Computing, 30(6):

1773–1789, December 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33167.

Shmoys:1994:IAA

[SSW94] David B. Shmoys, Clifford Stein, and Joel Wein. Improved approximation algorithms for shop scheduling problems. SIAM Journal on Computing, 23(3):617–632, June 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22676.

Shen:2003:SSS

[SSW03] Jian Shen, Li Sheng, and Jie Wu. Searching for sorted sequences of kings in tournaments. SIAM Journal on Computing, 32(5):1201–1209, October 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41005.

Sedgewick:1982:CFC

[SSY82] Robert Sedgewick, Thomas G.

Szymanski, and Andrew C. Yao. The complexity of finding cycles in periodic functions. SIAM Journal on Computing, 11(2):376–390, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sugihara:1990:SSP

[SSY90] Kazuo Sugihara, Ichiro Suzuki, and Masafumi Yamashita. The searchlight scheduling problem. SIAM Journal on Computing, 19(6):1024–1040, December 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shedler:1972:LPR

[ST72] G. S. Shedler and C. Tung. Locality in page reference strings. SIAM Journal on Computing, 1(3):218–241, ???? 1972. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sleator:1986:SAH

[ST86] Daniel Dominic Sleator and Robert Endre Tarjan. Self-Adjusting Heaps. SIAM Journal on Computing, 15(1):52– 69, February 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sundar:1994:UBS

[ST94] Rajamani Sundar and Robert E. Tarjan. Unique binary-search-tree representations and equality testing of sets and sequences. SIAM Journal on Computing, 23(1):24–44, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        18973.

Song:2000:CCF

[ST00] Nah-Oak Song and Demosthenis Teneketzis. On a conjecture by Coffman, Flatto, and Wright on stochastic machine minimization. SIAM Journal on Computing, 30(2):681–687, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32464.

Samorodnitsky:2009:GUI

[ST09a] Alex Samorodnitsky and Luca Trevisan. Gowers uniformity, influence of variables, and PCPs. SIAM Journal on Computing, 39(1):323–360, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Spakowski:2009:HU

[ST09b] Holger Spakowski and Rahul Tripathi. Hierarchical unambiguity. SIAM Journal on Computing, 38(5):2079–2112, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Spielman:2011:SSG

[ST11] Daniel A. Spielman and Shang-Hua Teng. Spectral sparsification of graphs. SIAM Journal on Computing, 40(4): 981–1025, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i4/p981_s1.

Spielman:2013:LCA

[ST13] Daniel A. Spielman and Shang-Hua Teng. A local clustering algorithm for massive graphs and its application to nearly linear time graph partitioning. SIAM Journal on Computing, 42(1):1–26, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shangguan:2023:GSB

[ST23] Chong Shangguan and Itzhak Tamo. Generalized singleton bound and list-decoding Reed– Solomon codes beyond the Johnson radius. SIAM Journal on Computing, 52(3):684–717, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M138795X.

Statman:1980:WCE

[Sta80] R. Statman. Worst case exponential lower bounds for input resolution with paramodulation. SIAM Journal on Computing, 9(1):104–110, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Steinberg:1997:SPA

[Ste97] A. Steinberg. A strip-packing algorithm with absolute performance bound 2. SIAM Journal on Computing, 26(2): 401–409, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
25580.

Stinson:1990:SOP

[Sti90] D. R. Stinson. Some observations on parallel algorithms for fast exponentiation in GF(2^{n}). SIAM Journal on Computing, 19(4):711–717, August 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Stoutemyer:1979:AAB

[Sto79] David R. Stoutemyer. Automatic asymptotic and big-O calculations via computer algebra. SIAM Journal on Computing, 8(3):287–299, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Stockmeyer:1985:AAP

[Sto85] Larry Stockmeyer. On approximation algorithms for #P. SIAM Journal on Computing, 14(4):849–861, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shawe-Taylor:1994:HCG

[STP94] John Shawe-Taylor and Tomaž Pisanski. Homeomorphism of 2-complexes is graph isomorphism complete. SIAM Journal on Computing, 23 (1):120–132, February 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/19890.

Strassen:1974:PRC

[Str74] Volker Strassen. Polynomials with rational coefficients which are hard to compute. SIAM Journal on Computing, 3(2): 128–149, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Strassen:1976:CCF

[Str76] Volker Strassen. Computational complexity over finite fields. SIAM Journal on Computing, 5(2):324–331, ???? 1976. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Strassen:1983:CCC

[Str83] V. Strassen. The computational complexity of continued fractions. SIAM Journal on Computing, 12(1):1–27, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shioura:1997:OAS

[STU97] Akiyoshi Shioura, Akihisa Tamura, and Takeaki Uno. An optimal algorithm for scanning all spanning trees of undirected graphs. SIAM Journal on Computing, 26(3):678–692, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27088.

Szymanski:1977:ERE

[SU77] T. G. Szymanski and J. D.

Ullman. Evaluating relational expressions with dense and sparse arguments. SIAM Journal on Computing, 6(1):109– 122, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Skutella:2005:SMS

[SU05] Martin Skutella and Marc Uetz. Stochastic machine scheduling with precedence constraints. SIAM Journal on Computing, 34(4):788– 802, August 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41500.

Shaltiel:2009:LEU

[SU09] Ronen Shaltiel and Christopher Umans. Low-end uniform hardness versus randomness tradeoffs for AM. SIAM Journal on Computing, 39(3): 1006–1037, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Subramanian:1994:NAS

[Sub94] Ashok Subramanian. A new approach to stable matching problems. SIAM Journal on Computing, 23(4):671– 700, August 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        16948.

Sutter:1991:FTF

[Sut91] Erich E. Sutter. The fast mtransform: a fast computation of cross-correlations with binary m-sequences. SIAM Journal on Computing, 20(4): 686–694, August 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Stockmeyer:1984:SPR

[SV84] Larry Stockmeyer and Uzi Vishkin. Simulation of parallel random access machines by circuits. SIAM Journal on Computing, 13(2):409–422, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schieber:1988:FLC

[SV88] Baruch Schieber and Uzi Vishkin. On finding lowest common ancestors: simplification and parallelization. SIAM Journal on Computing, 17(6): 1253–1262, December 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Saran:1995:FCW

[SV95] Huzur Saran and Vijay V. Vazirani. Finding k cuts within twice the optimal. SIAM Journal on Computing, 24 (1):101–108, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25173.

Scheideler:2000:SDR

[SV00a] Christian Scheideler and Berthold Vöcking. From static to dynamic routing: Efficient transformations of store-and-forward protocols. SIAM Journal on Computing, 30 (4):1126–1155, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35343.

Stacho:2000:VPL

[SV00b] Ladislav Stacho and Imrich Vrto. Virtual path layouts in ATM networks. SIAM Journal on Computing, 29 (5):1621–1629, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30815.

Sanghvi:2008:RCT

[SV08] Saurabh Sanghvi and Salil Vadhan. The round complexity of two-party random selection. SIAM Journal on Computing, 38(2):523–550, ???? 2008. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shaltiel:2010:HAP

[SV10] Ronen Shaltiel and Emanuele Viola. Hardness amplification proofs require majority. SIAM Journal on Computing, 39(7): 3122–3154, ???? 2010. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Singh:2018:AMC

[SV18] Mohit Singh and László A. Végh. Approximating minimum cost connectivity orientation and augmentation. SIAM Journal on Computing, 47(1): 270–293, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Svensson:2011:HPC

[Sve11]  Ola Svensson.  Hardness

of precedence constrained scheduling on identical machines. SIAM Journal on Computing, 40(5):1258–1274, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i5/
                                        p1258_s1.

Svensson:2012:SCS

[Sve12] Ola Svensson. Santa Claus schedules jobs on unrelated machines. SIAM Journal on Computing, 41(5):1318–1341, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Seiden:2003:NBV

[SvSE03] Steven S. Seiden, Rob van Stee, and Leah Epstein. New bounds for variable-sized online bin packing. SIAM Journal on Computing, 32(2):455– 469, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
41290.

Stefankovic:2012:DPT

[SVV12] Daniel Stefankovic, Santosh Vempala, and Eric Vigoda. A deterministic polynomialtime approximation scheme for counting knapsack solutions. SIAM Journal on Computing, 41(2):356–366, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Szymanski:1976:NEB

[SW76] Thomas G. Szymanski and John H. Williams. Noncanonical extensions of bottom-up parsing techniques. SIAM Journal on Computing, 5(2): 231–250, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Stockmeyer:1979:NCF

[SW79] L. J. Stockmeyer and C. K.

Wong. On the number of comparisons to find the intersection of two relations. SIAM Journal on Computing, 8(3):

388–404, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Satyanarayana:1985:LTA

[SW85] A. Satyanarayana and R. Kevin Wood. A linear-time algorithm for computing K-terminal reliability in series-parallel networks. SIAM Journal on Computing, 14(4):818–832, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Simons:1989:FAM

[SW89] Barbara B. Simons and Manfred K. Warmuth. A fast algorithm for multiprocessor scheduling of unit-length jobs. SIAM Journal on Computing, 18(4):690–710, August 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Santha:1993:LNC

[SW93] Miklós Sántha and Christopher Wilson. Limiting negations in constant depth circuits. SIAM Journal on Computing, 22(2):294–302, April 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Swaminathan:1994:CRP

[SW94] R. Swaminathan and Donald K. Wagner. On the consecutive-retrieval problem. SIAM Journal on Computing, 23(2):398–414, April 1994.

CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23548.

Shi:1999:DWN

[SW99] Weiping Shi and Douglas B.

West. Diagnosis of wiring networks: An optimal randomized algorithm for finding connected components of unknown graphs. SIAM Journal on Computing, 28(5):1541– 1551, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28811.

Sharir:2006:NCF

[SW06a] Micha Sharir and Emo Welzl.

On the number of crossingfree matchings, cycles, and partitions. SIAM Journal on Computing, 36(3):695–720, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shpilka:2006:DHT

[SW06b] Amir Shpilka and Avi Wigderson. Derandomizing homomorphism testing in general groups. SIAM Journal on Computing, 36(4):1215–1230, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Soloveichik:2007:CSA

[SW07] David Soloveichik and Erik Winfree. Complexity of selfassembled shapes. SIAM Journal on Computing, 36(6):1544– 1569, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schulman:2010:PCN

[SW10] Rebecca Schulman and Erik Winfree. Programmable control of nucleation for algorithmic self-assembly. SIAM Journal on Computing, 39(4):1581– 1616, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schwartz:2019:RCM

[SW19] Oded Schwartz and Elad Weiss. Revisiting “Computation of Matrix Chain Products”. SIAM Journal on Computing, 48(5):1481–1486, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [HS84].

Sahai:2021:HUI

[SW21] Amit Sahai and Brent Waters. How to use indistinguishability obfuscation: Deniable encryption, and more. SIAM Journal on Computing, 50(3):857–908, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Satyanarayana:1996:NLT

[SWCP96] A. Satyanarayana, R. K. Wood, L. Camarinopoulos, and G. Pampoukis. Note on: “A linear-time algorithm for computing K-terminal reliability in series-parallel networks” [SIAM J. Comput. 14 (1985), no. 4, 818–832, MR 86m:90069] by Satyanarayana and Wood. SIAM Journal on Computing, 25(2):290, April 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25798.

Sweedyk:1999:BAA

[Swe99] Z. Sweedyk. A2\frac{1}{2}-approximation algorithm for shortest superstring. SIAM Journal on Computing, 29(3):954–986, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32466.

Schmid:2009:IRL

[SWK09] Ulrich Schmid, Bettina Weiss, and Idit Keidar. Impossibility results and lower bounds for consensus under link failures. SIAM Journal on Computing, 38(5):1912–1951, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Shao:2011:MCC

[SWPL11] Cheng Shao, Jennifer L.

Welch, Evelyn Pierce, and Hyunyoung Lee. Multiwriter consistency conditions for shared memory registers. SIAM Journal on Computing, 40(1):28–62, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicomp/resource/
                                        1/smjcat/v40/i1/p28_s1.

Shmoys:1995:SPM

[SWW95] David B. Shmoys, Joel Wein, and David P. Williamson. Scheduling parallel machines on-line. SIAM Journal on Computing, 24(6):1313–1331, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        24831.

Selman:1983:PRC

[SXB83] Alan L. Selman, Mei-Rui Xu, and Ronald V. Book. Positive relativizations of complexity classes. SIAM Journal on Computing, 12(3):565–579, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Stockmeyer:1980:OLM

[SY80] Paul K. Stockmeyer and F. Frances Yao. On the optimality of linear merge. SIAM Journal on Computing, 9(1): 85–90, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Schaffer:1991:SLS

[SY91] Alejandro A. Schäffer and Mihalis Yannakakis. Simple local search problems that are hard to solve. SIAM Journal on Computing, 20(1):56– 87, February 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Suzuki:1992:SMI

[SY92] Ichiro Suzuki and Masafumi Yamashita. Searching for a mobile intruder in a polygonal region. SIAM Journal on Computing, 21(5):863–888, October 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Suzuki:1999:DAM

[SY99] Ichiro Suzuki and Masafumi Yamashita. Distributed anonymous mobile robots: Formation of geometric patterns. SIAM Journal on Computing, 28(4):1347–1363, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28292. See erra-

tum [SY06].

Suzuki:2006:EDA

[SY06] Ichiro Suzuki and Masafumi Yamashita. Erratum: Distributed anonymous mobile robots: Formation of geometric patterns. SIAM Journal on Computing, 36(1):279–280, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [SY99].

Snir:2012:RAP

[SY12] Sagi Snir and Raphael Yuster.

Reconstructing approximate phylogenetic trees from quartet samples. SIAM Journal on Computing, 41(6):1466–1480, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Syslo:1981:ECV

[Sys81] Maciej M. Sysło. An efficient cycle vector space algorithm for listing all cycles of a planar graph. SIAM Journal on Computing, 10(4):797–808, ???? 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Sudborough:1976:FLD

[SZ76] I. H. Sudborough and A. Zalcberg. On families of languages defined by time-bounded random access machines. SIAM Journal on Computing, 5(2): 217–230, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Srinivasan:1999:CVW

[SZ99] Aravind Srinivasan and David Zuckerman. Computing with very weak random sources. SIAM Journal on Computing, 28(4):1433–1459, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30091.

Saks:2000:WFS

[SZ00] Michael Saks and Fotios Zaharoglou. Wait-free k-set agreement is impossible: The topology of public knowledge.

SIAM Journal on Computing, 29(5):1449–1483, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30769.

Shi:2008:TNC

[SZ08] Yaoyun Shi and Yufan Zhu.

Tensor norms and the classical communication complexity of nonlocal quantum measurement. SIAM Journal on Computing, 38(3):753–766, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Szkaliczki:1999:RMW

[Szk99]  Tibor Szkaliczki.  Routing

with minimum wire length in the dogleg-free Manhattan model is N P-complete. SIAM Journal on Computing, 29(1):274–287, February 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30312.

Szpankowski:1993:GST

[Szp93] Wojciech Szpankowski. A generalized suffix tree and its (un)expected asymptotic behaviors. SIAM Journal on Computing, 22(6):1176–1198, December 1993. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Torlone:1999:EDU

[TA99] Riccardo Torlone and Paolo Atzeni. Efficient database updates with independent schemes. SIAM Journal on Computing, 28(3):1112–1135, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25799.

Tai:1980:PCF

[Tai80] Kuo Chung Tai. Predictors of context-free grammars. SIAM Journal on Computing, 9(3): 653–664, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tamassia:1987:EGG

[Tam87] Roberto Tamassia. On embedding a graph in the grid with the minimum number of bends. SIAM Journal on Computing, 16(3):421–444, June 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tamaki:1998:ESE

[Tam98] Hisao Tamaki. Efficient selfembedding of butterfly networks with random faults. SIAM Journal on Computing, 27(3):614–636, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27036.

Tanner:1978:MMS

[Tan78] R. Michael Tanner. Minimean merging and sorting: An algorithm. SIAM Journal on Computing, 7(1):18–38, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tarjan:1972:DFS

[Tar72] Robert E. Tarjan. Depth first search and linear graph algorithms. SIAM Journal on Computing, 1(2):146–160, June 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tarjan:1973:EEC

[Tar73] Robert Tarjan. Enumeration of the elementary circuits of a directed graph. SIAM Journal on Computing, 2(3):211– 216, ???? 1973. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tarjan:1974:FDD

[Tar74] Robert Tarjan. Finding dominators in directed graphs. SIAM Journal on Computing, 3(1):62–89, ???? 1974. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Toueg:1984:OCS

[TB84] Sam Toueg and Özalp Babaoğlu.

On the optimum checkpoint selection problem. SIAM Journal on Computing, 13(3):630– 649, ???? 1984. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tsin:1984:EPA

[TC84] Yung H. Tsin and Francis Y. Chin. Efficient parallel algorithms for a class of graph theoretic problems. SIAM Journal on Computing, 13(3):580–599, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tetali:1999:DLA

[Tet99] Prasad Tetali. Design of online algorithms using hitting times. SIAM Journal on Computing, 28(4):1232–1246, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33551.

Theobald:2002:EGF

[The02] Thorsten Theobald. An enumerative geometry framework for algorithmic line problems in R^{3}. SIAM Journal on Computing, 31(4):1212–1228, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38050.

Thomassen:1997:CFM

[Tho97] Carsten Thomassen. On the complexity of finding a minimum cycle cover of a graph. SIAM Journal on Computing, 26(3):675–677, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26725.

Thorup:2000:RPQ

[Tho00] Mikkel Thorup. On RAM priority queues. SIAM Journal on Computing, 30(1):86– 109, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28824.

Thorup:2005:QMC

[Tho05] Mikkel Thorup. Quick kmedian, k-center, and facility location for sparse graphs. SIAM Journal on Computing, 34(2):405–432, April 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38888.

Thorup:2018:DP

[Tho18] Mikkel Thorup. Sample(x) = (a ∗ x ≤ t) is a distinguisher with probability 1/8. SIAM Journal on Computing, 47(6): 2510–2526, ???? 2018. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Thurber:1999:EGM

[Thu99]  Edward G. Thurber.  Ef-

ficient generation of minimal length addition chains. SIAM Journal on Computing, 28(4):1247–1263, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29566.

Tsukiyama:1977:NAG

[TIAS77] Shuji Tsukiyama, Mikio Ide, Hiromu Ariyoshi, and Isao Shirakawa. A new algorithm for generating all the maximal independent sets. SIAM Journal on Computing, 6(3):505– 517, ???? 1977. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tenenbaum:1982:TSS

[TN82] Aaron M. Tenenbaum and Richard M. Nemes. Two spectra of self-organizing sequential search algorithms. SIAM Journal on Computing, 11(3):557– 566, ???? 1982. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tokuyama:1995:EAH

[TN95] Takeshi Tokuyama and Jun Nakano. Efficient algorithms for the Hitchcock transportation problem. SIAM Journal on Computing, 24(3):563–578, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23672.

Toda:1992:CCL

[TO92] Seinosuke Toda and Mitsunori Ogiwara. Counting classes are at least as hard as the polynomial-time hierarchy. SIAM Journal on Computing, 21(2):316–328, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Toda:1991:PHP

[Tod91] Seinosuke Toda. PP is as hard as the polynomial-time hierarchy. SIAM Journal on Computing, 20(5):865–877, October 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tompa:1982:TFT

[Tom82] Martin Tompa. Two familiar transitive closure algorithms which admit no polynomial time, sublinear space implementations. SIAM Journal on Computing, 11(1):130–137, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tompa:1992:LBU

[Tom92] Martin Tompa. Lower bounds on universal traversal sequences for cycles and other low degree graphs. SIAM Journal on Computing, 21(6):1153– 1160, December 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Toran:2004:HGI

[Tor04] Jacobo Torán. On the hardness of graph isomorphism. SIAM Journal on Computing, 33(5):1093–1108, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41096.

Toth:2003:BSP

[Tót03] Csaba D. Tóth. Binary space partitions for line segments with a limited number of directions. SIAM Journal on Computing, 32(2):307–325, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40378.

Toth:2008:BSP

[Tót08] Csaba D. Tóth. Binary space partitions for axis-aligned fat rectangles. SIAM Journal on Computing, 38(1):429–447, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Toueg:1987:FDA

[TPS87] Sam Toueg, Kenneth J. Perry, and T. K. Srikanth. Fast distributed agreement. SIAM Journal on Computing, 16(3): 445–457, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Traub:1972:CCI

[Tra72] J. F. Traub. Computational complexity of iterative processes. SIAM Journal on Computing, 1(2):167–179, ???? 1972. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Trevisan:2000:WHM

[Tre00] Luca Trevisan. When Hamming meets Euclid: The approximability of geometric TSP and Steiner tree. SIAM Journal on Computing, 30(2):475–485, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35273.

Trevisan:2012:MCS

[Tre12] Luca Trevisan. Max cut and the smallest eigenvalue. SIAM Journal on Computing, 41(6): 1769–1786, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Triesch:1994:SRE

[Tri94] Eberhard Triesch. Some results on elusive graph properties. SIAM Journal on Computing, 23(2):247–254, April 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19415.

Trifonov:2008:SAU

[Tri08]  Vladimir Trifonov.  An

O(log n log log n) space algorithm for undirected stconnectivity. SIAM Journal on Computing, 38(2):449–483, ???? 2008. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Trojanowski:1978:RLA

[Tro78] Anthony E. Trojanowski. Ranking and listing algorithms for k-ary trees. SIAM Journal on Computing, 7(4):492–509, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Toueg:1981:SCR

[TS81] Sam Toueg and Kenneth Steiglitz. Some complexity results in the design of deadlockfree packet switching networks. SIAM Journal on Computing, 10(4):702–712, November 1981. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ta-Shma:2011:SSE

[TS11]  Amnon Ta-Shma.  Short

seed extractors against quantum storage. SIAM Journal on Computing, 40(3):664–677, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sicomp/
                                        resource/1/smjcat/v40/i3/
                                        p664_s1.

Tsai:1992:AAA

[Tsa92]  Li-Hui Tsai.  Asymptotic

analysis of an algorithm for balanced parallel processor scheduling. SIAM Journal on Computing, 21(1):59–64, February 1992. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Trevisan:2000:GAL

[TSSW00] Luca Trevisan, Gregory B.

Sorkin, Madhu Sudan, and David P. Williamson. Gadgets, approximation, and linear programming. SIAM Journal on Computing, 29(6): 2074–2097, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32884.

Tarjan:1977:FMI

[TT77] Robert Endre Tarjan and Anthony E. Trojanowski. Finding a maximum independent set. SIAM Journal on Computing, 6(3):537–546, ???? 1977. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Therien:2001:TLS

[TT01] Denis Thérien and Thomas.

Temporal logic and semidirect products: An effective characterization of the until hierarchy. SIAM Journal on Computing, 31(3):777–798, June 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32277.

Tancer:2013:NGC

[TT13] Martin Tancer and Dmitry Tonkonog. Nerves of good covers are algorithmically unrecognizable. SIAM Journal on Computing, 42(4):1697–1719, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Toueg:1981:DFP

[TU81] Sam Toueg and Jeffrey D. Ullman. Deadlock-free packet switching networks. SIAM Journal on Computing, 10(3): 594–611, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tarhio:1993:ABM

[TU93] Jorma Tarhio and Esko Ukkonen. Approximate Boyer– Moore string matching. SIAM Journal on Computing, 22(2): 243–260, April 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tucker:1980:ETC

[Tuc80] Alan Tucker. An efficient test for circular-arc graphs. SIAM Journal on Computing, 9(1):1– 24, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tung:1991:CSN

[Tun91] Shih Ping Tung. Complexity of sentences over number rings. SIAM Journal on Computing, 20(1):126–143, February 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Turner:1986:PPH

[Tur86]  Jonathan S. Turner.  On

the probable performance of heuristics for bandwidth minimization. SIAM Journal on Computing, 15(2):561–580, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tarjan:1985:EPB

[TV85] Robert E. Tarjan and Uzi Vishkin. An efficient parallel biconnectivity algorithm. SIAM Journal on Computing, 14(4):862–874, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tarjan:1988:TAT

[TV88a] Robert E. Tarjan and Christopher J. Van Wyk. An O(n log log n)-time algorithm for triangulating a simple polygon. SIAM Journal on Computing, 17(1):143–178, February 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [TV88b].

Tarjan:1988:ETA

[TV88b] Robert E. Tarjan and Christopher J. Van Wyk. Erratum: “An O(n log log n)-time algorithm for triangulating a simple polygon”. SIAM Journal on Computing, 17(5):1061, October 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [TV88a].

Tamassia:1991:PTC

[TV91] Roberto Tamassia and Jeffrey S. Vitter. Parallel transitive closure and point location in planar structures. SIAM Journal on Computing, 20(4): 708–725, August 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Talamo:1999:EDS

[TV99] Maurizio Talamo and Paola Vocca. An efficient data structure for lattice operations. SIAM Journal on Computing, 28(5):1783–1805, October 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27440.

Traub:2020:BIR

[TV20] Vera Traub and Jens Vygen. Beating the integrality ratio for s–t-tours in graphs. SIAM Journal on Computing, 52 (6):FOCS18–37–FOCS18–84, February 2020. CODEN SMJ-CAT. ISSN 1095-7111.

Traub:2022:IAA

[TV22] Vera Traub and Jens Vygen. An improved approximation algorithm for the asymmetric traveling salesman problem. SIAM Journal on Computing, 51(1):139–173, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1339313.

Traub:2022:RPT

[TVZ22] Vera Traub, Jens Vygen, and Rico Zenklusen. Reducing path TSP to TSP. SIAM Journal on Computing, 51 (3):STOC20–24–STOC20–53, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M135594X.

Tang:1989:TRC

[TW89] Shou Wen Tang and Osamu Watanabe. On tally relativizations of BP-complexity classes. SIAM Journal on Computing, 18(3):449–462, June 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tulsiani:2014:QGL

[TW14] Madhur Tulsiani and Julia Wolf. Quadratic Goldreich– Levin theorems. SIAM Journal on Computing, 43(2):730–766, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tarjan:1984:SLT

[TY84] Robert E. Tarjan and Mihalis Yannakakis. Simple linear-time algorithms to test chordality of graphs, test acyclicity of hypergraphs, and selectively reduce acyclic hypergraphs. SIAM Journal on Computing, 13(3):566–579, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See addendum [TY85].

Tarjan:1985:ASL

[TY85] Robert E. Tarjan and Mihalis Yannakakis. Addendum: “Simple linear-time algorithms to test chordality of graphs, test acyclicity of hypergraphs, and selectively reduce acyclic hypergraphs”. SIAM Journal on Computing, 14(1):254–255, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [TY84].

Teng:1997:ASS

[TY97] Shang-Hua Teng and Frances F. Yao. Approximating shortest superstrings. SIAM Journal on Computing, 26(2):410–417, April 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28612.

Thorup:2012:TBI

[TZ12] Mikkel Thorup and Yin Zhang. Tabulation-based 5independent hashing with applications to linear probing and second moment estimation. SIAM Journal on Computing, 41(2):293–331, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Thapper:2017:PSA

[TZ17] Johan Thapper and Stanislav Zivný. The power of Sherali– Adams relaxations for generalvalued CSPs. SIAM Journal on Computing, 46(4):1241– 1279, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Tidor:2022:TLI

[TZ22] Jonathan Tidor and Yufei Zhao. Testing linear-invariant properties. SIAM Journal on Computing, 51(4):1230–1279, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1397246.

Tzeng:1992:PTA

[Tze92] Wen-Guey Tzeng. A polynomialtime algorithm for the equivalence of probabilistic automata. SIAM Journal on Computing, 21(2):216–227, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ukkonen:1983:TRP

[Ukk83]  Esko Ukkonen.  Two re-

sults on polynomial time truthtable reductions to sparse sets. SIAM Journal on Computing, 12(3):580–587, ???? 1983. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ullman:2016:ACQ

[Ull16] Jonathan Ullman. Answering n^{2} +o(1) counting queries with differential privacy is hard. SIAM Journal on Computing, 45(2):473–496, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Urquhart:2011:NOS

[Urq11] Alasdair Urquhart. A nearoptimal separation of regular and general resolution. SIAM Journal on Computing, 40(1): 107–121, ???? 2011. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i1/p107_s1.

Ullman:1991:HPP

[UY91] Jeffrey D. Ullman and Mihalis Yannakakis. High-probability parallel transitive-closure algorithms. SIAM Journal on Computing, 20(1):100–125, February 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Varadarajan:2000:ASP

[VA00] Kasturi R. Varadarajan and Pankaj K. Agarwal. Approximating shortest paths on a nonconvex polyhedron. SIAM Journal on Computing, 30(4):1321–1340, August 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35275.

Vadhan:2001:CCS

[Vad01] Salil P. Vadhan. The complexity of counting in sparse, regular, and planar graphs. SIAM Journal on Computing, 31(2):398–427, April 2001. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32160.

Vadhan:2006:USC

[Vad06] Salil P. Vadhan. An unconditional study of computational zero knowledge. SIAM Journal on Computing, 36(4):1160– 1214, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vaishnavi:1987:WLA

[Vai87] Vijay K. Vaishnavi. Weighted leaf AVL-trees. SIAM Journal on Computing, 16(3):503–537, June 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [Vai90].

Vaidya:1988:MST

[Vai88] Pravin M. Vaidya. Minimum spanning trees in kdimensional space. SIAM Journal on Computing, 17(3): 572–582, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vaidya:1989:GHM

[Vai89a] Pravin M. Vaidya. Geometry helps in matching. SIAM Journal on Computing, 18(6):1201– 1225, December 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vaidya:1989:STT

[Vai89b] Pravin M. Vaidya. Spacetime trade-offs for orthogonal range queries. SIAM Journal on Computing, 18(4):748–758, August 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vaishnavi:1990:EWL

[Vai90] Vijay K. Vaishnavi. Erratum: “Weighted leaf AVLtrees” [SIAM J. Comput. 16 (1987), no. 3, 503–537, MR 88c:68012]. SIAM Journal on Computing, 19(3):591, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Vai87].

Valiant:1975:PCP

[Val75] Leslie G. Valiant. Parallelism in comparison problems. SIAM Journal on Computing, 4(3): 348–355, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Valiant:1979:CER

[Val79] Leslie G. Valiant. The complexity of enumeration and reliability problems. SIAM Journal on Computing, 8(3):410– 421, ???? 1979. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Valiant:1982:SFP

[Val82] L. G. Valiant. A scheme for fast parallel communication. SIAM Journal on Computing, 11(2):350–361, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Valiant:1986:NPB

[Val86] L. G. Valiant. Negation is powerless for Boolean Slice functions. SIAM Journal on Computing, 15(2):531–535, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Valiant:2002:QCC

[Val02] Leslie G. Valiant. Quantum circuits that can be simulated classically in polynomial time. SIAM Journal on Computing, 31(4):1229–1254, August 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37702.

Valiant:2008:HA

[Val08] Leslie G. Valiant. Holographic algorithms. SIAM Journal on Computing, 37(5):1565–1594, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Valiant:2011:TSP

[Val11] Paul Valiant. Testing symmetric properties of distributions. SIAM Journal on Computing, 40(6):1927–1968, ???? 2011. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v40/i6/p1927_s1.

Varricchio:1997:PCR

[Var97] Stefano Varricchio. A pumping condition for regular sets. SIAM Journal on Computing, 26(3):764–771, June 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/17994.

Varghese:2000:SSC

[Var00]  George Varghese.  Self-

stabilization by counter flushing. SIAM Journal on Computing, 30(2):486–510, April 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32760.

Vazirani:1997:ISS

[Vaz97] Umesh Vazirani. Introduction to special section on quantum computation. SIAM Journal on Computing, 26 (5):1409–1410, October 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97402.

Vitter:1985:OAM

[VC85] Jeffrey Scott Vitter and Wen Chin Chen. Optimum algorithms for a model of direct chaining. SIAM Journal on Computing, 14(2):490–499, May 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vesely:2022:CAS

[VCJS22] Pavel Veselý, Marek Chrobak, Lukasz Jez, and Jirí Sgall. A ϕ-competitive algorithm for scheduling packets with deadlines. SIAM Journal on Computing, 51(5):1626–1691, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1469753.

vanDam:2006:QAS

[vDHI06] Wim van Dam, Sean Hallgren, and Lawrence Ip. Quantum algorithms for some hidden shift problems. SIAM Journal on Computing, 36(3): 763–778, ???? 2006. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vanDam:2007:STU

[vDMMS07] Wim van Dam, Frédéric Magniez, Michele Mosca, and Miklos Santha. Self-testing of universal and fault-tolerant sets of quantum gates. SIAM Journal on Computing, 37(2):611–629, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vegh:2016:SPA

[Vég16] László A. Végh. A strongly polynomial algorithm for a class of minimum-cost flow problems with separable convex objectives. SIAM Journal on Computing, 45(5):1729– 1761, ???? 2016. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Veldhorst:1985:ACO

[Vel85] Marinus Veldhorst. Approximation of the consecutive ones matrix augmentation problem. SIAM Journal on Computing, 14(3):709–729, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Venkateswaran:1992:CDN

[Ven92] H. Venkateswaran. Circuit definitions of nondeterministic complexity classes. SIAM Journal on Computing, 21(4): 655–670, August 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Verma:1997:GTA

[Ver97] Rakesh M. Verma. General techniques for analyzing recursive algorithms with applications. SIAM Journal on Computing, 26(2):568–581, April 1997. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24058.

Vertigan:2005:CCT

[Ver05] Dirk Vertigan. The computational complexity of Tutte invariants for planar graphs. SIAM Journal on Computing, 35(3):690–712, ???? 2005. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vershynin:2009:BHC

[Ver09] Roman Vershynin. Beyond Hirsch conjecture: Walks on random polytopes and smoothed complexity of the simplex method. SIAM Journal on Computing, 39(2):646– 678, ???? 2009. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Villanger:2009:ICF

[VHPT09] Yngve Villanger, Pinar Heggernes, Christophe Paul, and Jan Arne Telle. Interval completion is fixed parameter tractable. SIAM Journal on Computing, 38(5):2007– 2020, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vidick:2016:TPE

[Vid16] Thomas Vidick. Three-player entangled XOR games are NPhard to approximate. SIAM Journal on Computing, 45(3):

1007–1063, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See erratum [Vid20].

Vidick:2020:ETP

[Vid20] Thomas Vidick. Erratum: Three-Player Entangled XOR Games are NP-hard to Approximate. SIAM Journal on Computing, 49(6):1423–1427, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). See [Vid16].

Vikas:2003:CCC

[Vik03] Narayan Vikas. Computational complexity of compaction to reflexive cycles. SIAM Journal on Computing, 32(1):253–280, February 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38352.

Vikas:2004:CRC

[Vik04] Narayan Vikas. Compaction, retraction, and constraint satisfaction. SIAM Journal on Computing, 33(4):761– 782, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
39780.

Vinodchandran:2004:CCS

[Vin04] N. V. Vinodchandran. Counting complexity of solvable black-box group problems.

SIAM Journal on Computing, 33(4):852–869, August 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42065.

Viola:2007:PBC

[Vio07] Emanuele Viola. Pseudorandom bits for constant-depth circuits with few arbitrary symmetric gates. SIAM Journal on Computing, 36(5):1387– 1403, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Viola:2012:BPL

[Vio12a] Emanuele Viola. Bit-probe lower bounds for succinct data structures. SIAM Journal on Computing, 41(6):1593–1604, ???? 2012. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Viola:2012:CD

[Vio12b] Emanuele Viola. The complexity of distributions. SIAM Journal on Computing, 41(1): 191–218, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (elec-

tronic). URL http://epubs.
siam.org/sicomp/resource/
1/smjcat/v41/i1/p191_s1.

Viola:2014:ECS

[Vio14] Emanuele Viola. Extractors for circuit sources. SIAM Journal on Computing, 43(2):655–672, ???? 2014. CODEN SMJCAT.

ISSN 0097-5397 (print), 1095-7111 (electronic).

Viola:2020:SLB

[Vio20]  Emanuele Viola.  Sam-

pling lower bounds: Boolean average-case and permutations. SIAM Journal on Computing, 49(1):119–137, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vishkin:1991:DSN

[Vis91]  Uzi Vishkin.  Determinis-

tic sampling — a new technique for fast pattern matching. SIAM Journal on Computing, 20(1):22–40, February 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vitanyi:1985:OSCa

[Vit85a] Paul M. B. Vitányi. An optimal simulation of counter machines. SIAM Journal on Computing, 14(1):1–33, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vitanyi:1985:OSCb

[Vit85b] Paul M. B. Vitányi. An optimal simulation of counter machines: the ACM case. SIAM Journal on Computing, 14(1): 34–40, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vitanyi:1988:LCI

[Vit88] Paul M. B. Vitányi. Locality, communication, and interconnect length in multicomputers. SIAM Journal on Computing, 17(4):659–672, August 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vanKreveld:2010:PIP

[vKLM10] Marc van Kreveld, Maarten Löffler, and Joseph S. B. Mitchell. Preprocessing imprecise points and splitting triangulations. SIAM Journal on Computing, 39(7):2990–3000, ???? 2010. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vincent:2000:RPN

[VL00] Millist W. Vincent and Mark Levene. Restructuring partitioned normal form relations without information loss. SIAM Journal on Computing, 29(5):1550–1567, October 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32631.

vanMelkebeek:2019:DIS

[vMP19] Dieter van Melkebeek and Gautam Prakriya. Derandomizing isolation in spacebounded settings. SIAM Journal on Computing, 48(3):979– 1021, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vanMelkebeek:2005:HPD

[vMS05] Dieter van Melkebeek and Rahul Santhanam. Holographic proofs and derandomization. SIAM Journal on Computing, 35(1):59–90, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vidick:2020:SSF

[VNA20] Thomas Vidick, Danupon Nanongkai, and Dimitris Achlioptas. Special section on the Fiftieth Annual ACM Symposium on Theory of Computing (STOC 2018). SIAM Journal on Computing, 49(5): STOC18–i–STOC18–ii, ???? 2020. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vondrak:2013:SAS

[Von13] Jan Vondrák. Symmetry and approximability of submodular maximization problems. SIAM Journal on Computing, 42(1): 265–304, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Valencia-Pabon:2003:RTA

[VP03] Mario Valencia-Pabon. Revisiting Tucker’s algorithm to color circular arc graphs. SIAM Journal on Computing, 32(4):1067–1072, August 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38215.

Verma:1989:AGA

[VR89] Rakesh M. Verma and Steven W.

Reyner. An analysis of a good algorithm for the subtree problem, corrected. SIAM Journal on Computing, 18(5):906–908, October 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Valiant:1983:FPC

[VSBR83] L. G. Valiant, S. Skyum, S. Berkowitz, and C. Rackoff. Fast parallel computation of polynomials using few processors. SIAM Journal on Computing, 12(4):641–644, November 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Venkateswaran:1989:NPG

[VT89] H. Venkateswaran and Martin Tompa. A new pebble game that characterizes parallel complexity classes. SIAM Journal on Computing, 18(3): 533–549, June 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vanTonder:2004:LCQ

[vT04] André van Tonder. A lambda calculus for quantum computation. SIAM Journal on Computing, 33(5):1109–1135, October 2004. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43216.

Valdes:1982:RSP

[VTL82] Jacobo Valdes, Robert E. Tarjan, and Eugene L. Lawler. The recognition of series parallel digraphs. SIAM Journal on Computing, 11(2):298–313, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vazirani:1989:TPS

[VV89] Umesh V. Vazirani and Vijay V. Vazirani. The twoprocessor scheduling problem is in random NC. SIAM Journal on Computing, 18(6):1140– 1148, December 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Valiant:2017:AIP

[VV17] Gregory Valiant and Paul Valiant. An automatic inequality prover and instance optimal identity testing. SIAM Journal on Computing, 46(1):429–455, ???? 2017. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Varadarajan:2007:ARP

[VVYZ07] Kasturi Varadarajan, S. Venkatesh, Yinyu Ye, and Jiawei Zhang. Approximating the radii of point sets. SIAM Journal on Computing, 36(6):1764–1776, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vijayan:1985:RGT

[VW85a] Gopalakrishnan Vijayan and Avi Wigderson. Rectilinear graphs and their embeddings. SIAM Journal on Computing, 14(2):355–372, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Vishkin:1985:TOB

[VW85b] Uzi Vishkin and Avi Wigderson. Trade-offs between depth and width in parallel computation. SIAM Journal on Computing, 14(2):303–314, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Varma:2023:ASG

[VY23] Nithin Varma and Yuichi Yoshida. Average sensitivity of graph algorithms. SIAM Journal on Computing, 52 (4):1039–1081, ???? 2023. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1399592.

Verbin:2013:LBT

[VZ13] Elad Verbin and Qin Zhang.

The limits of buffering: a tight lower bound for dynamic membership in the external memory model. SIAM Journal on Computing, 42(1):212–229, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vonzurGathen:1984:PAA

[vzG84] Joachim von zur Gathen. Parallel algorithms for algebraic problems. SIAM Journal on Computing, 13(4):802–824, ???? 1984. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vonzurGathen:1986:RPC

[vzG86] Joachim von zur Gathen. Representations and parallel computations for rational functions. SIAM Journal on Computing, 15(2):432–452, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vonzurGathen:1987:CPP

[vzG87] Joachim von zur Gathen. Computing powers in parallel. SIAM Journal on Computing, 16(5):930–945, October 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vonzurGathen:1991:TPP

[vzG91] Joachim von zur Gathen. Tests for permutation polynomials. SIAM Journal on Computing, 20(3):591–602, June 1991. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

vonzurGathen:1998:CCP

[vzGS98] Joachim von zur Gathen and Igor Shparlinski. Computing components and projections of curves over finite fields. SIAM Journal on Computing, 28(3):822–840, June 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27741.

vonzurGathen:2000:CPC

[vzGS00] Joachim von zur Gathen and Igor E. Shparlinski. The CREW PRAM complexity of modular inversion. SIAM Journal on Computing, 29(6): 1839–1857, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32807.

vonzurGathen:2003:FPC

[vzGSS03] Joachim von zur Gathen, Igor Shparlinski, and Alistair Sinclair. Finding points on curves over finite fields. SIAM Journal on Computing, 32(6):1436–1448, December 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35101.

Wadsworth:1976:RBC

[Wad76] Christopher P. Wadsworth.

The relation between computational and denotational properties for Scott’s D_{∞}-models of the lambda-calculus. SIAM Journal on Computing, 5(3): 488–521, September 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wadsworth:1978:ARL

[Wad78] Christopher P. Wadsworth. Approximate reduction and lambda calculus models. SIAM Journal on Computing, 7(3): 337–356, August 1978. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wagner:1990:BQC

[Wag90] Klaus W. Wagner. Bounded query classes. SIAM Journal on Computing, 19(5):833–846, October 1990. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Walkup:1979:EVR

[Wal79] David W. Walkup. On the expected value of a random assignment problem. SIAM Journal on Computing, 8(3): 440–442, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Walsh:1998:PTC

[Wal98] P. G. Walsh. A polynomial time complexity bound for computations on curves. SIAM Journal on Computing, 28(2):704–708, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30096.

Wang:1992:PTP

[Wan92] Jie Wang. Polynomial time productivity, approximations, and levelability. SIAM Journal on Computing, 21(6):1100– 1111, December 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wang:1998:GRP

[Wan98] Yongge Wang. Genericity, randomness, and polynomial-time approximations. SIAM Journal on Computing, 28(2):394– 408, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
30235.

Wang:1999:DWP

[Wan99] Jie Wang. Distributional word problem for groups. SIAM Journal on Computing, 28 (4):1264–1283, August 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33226.

Watrous:2009:ZKA

[Wat09] John Watrous. Zero-knowledge against quantum attacks. SIAM Journal on Computing, 39(1):25–58, ???? 2009. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Watson:2014:THS

[Wat14] Thomas Watson. Time hierarchies for sampling distributions. SIAM Journal on Computing, 43(5):1709–1727, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Whidden:2013:FPA

[WBZ13] Chris Whidden, Robert G. Beiko, and Norbert Zeh. Fixed-parameter algorithms for maximum agreement forests. SIAM Journal on Computing, 42(4):1431–1466, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wagner:1990:ETH

[WC90] A. Wagner and D. G. Corneil. Embedding trees in a hypercube is NP-complete. SIAM Journal on Computing, 19(3): 570–590, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wang:1992:HSP

[WC92] Qingzhou Wang and Kam-Hoi Cheng. A heuristic of scheduling parallel tasks and its analysis. SIAM Journal on Computing, 21(2):281–294, April 1992. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wong:1980:EMW

[WE80] C. K. Wong and M. C. Easton. An efficient method for weighted sampling without replacement. SIAM Journal on Computing, 9(1):111–113, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Weber:1993:DFV

[Web93] Andreas Weber. Decomposing finite-valued transducers and deciding their equivalence. SIAM Journal on Computing, 22(1):175–202, February 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Weide:1980:RGG

[Wei80] Bruce W. Weide. Random graphs and graph optimization problems. SIAM Journal on Computing, 9(3):552–557, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Westbrook:1994:RAM

[Wes94]  Jeffery Westbrook.  Ran-

domized algorithms for multiprocessor page migration. SIAM Journal on Computing, 23(5):951–965, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19979.

Weyuker:1979:TDQ

[Wey79] Elaine J. Weyuker. Translatability and decidability questions for restricted classes of program schemas. SIAM Journal on Computing, 8(4):587– 598, ???? 1979. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wagh:2005:MCT

[WG05] Meghanad D. Wagh and Osman Guzide. Mapping cycles and trees on wrap-around butterfly graphs. SIAM Journal on Computing, 35(3):741–765, ???? 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Whitehead:1990:CFT

[Whi90] Jennifer Whitehead. The complexity of file transfer scheduling with forwarding. SIAM Journal on Computing, 19(2): 222–245, April 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Williamson:1973:IRT

[Wil73] S. G. Williamson. Isomorph rejection and a theorem of de Bruijn. SIAM Journal on Computing, 2(1):44–59, ???? 1973. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Williamson:1976:RAL

[Wil76] S. G. Williamson. Ranking algorithms for lists of partitions. SIAM Journal on Computing, 5(4):602–617, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Willard:1982:PR

[Wil82] Dan E. Willard. Polygon retrieval. SIAM Journal on Computing, 11(1):149–165, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Willard:1985:NDS

[Wil85a] Dan E. Willard. New data structures for orthogonal range queries. SIAM Journal on Computing, 14(1):232–253, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Willard:1985:SUN

[Wil85b] Dan E. Willard. Searching unindexed and nonuniformly generated files in log log N time. SIAM Journal on Computing, 14(4):1013–1029, November 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Willard:1986:LLS

[Wil86]  Dan E. Willard.  Log-

logarithmic selection resolution protocols in a multiple access channel. SIAM Journal on Computing, 15(2):468–477, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wilber:1989:LBA

[Wil89] Robert Wilber. Lower bounds for accessing binary search trees with rotations. SIAM Journal on Computing, 18(1):

56–67, February 1989. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wilson:1990:D

[Wil90] Christopher B. Wilson. On the decomposability of NC and AC. SIAM Journal on Computing, 19(2):384–396, April 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Willard:2000:ECG

[Wil00] Dan E. Willard. Examining computational geometry, Van Emde Boas trees, and hashing from the perspective of the fusion tree. SIAM Journal on Computing, 29(3):1030–1049, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32242.

Williams:2013:IES

[Wil13] Ryan Williams. Improving exhaustive search implies superpolynomial lower bounds. SIAM Journal on Computing, 42(3):1218–1244, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Williams:2016:NPV

[Wil16] R. Ryan Williams. Natural proofs versus derandomization. SIAM Journal on Computing, 45(2):497–529, ???? 2016. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Williams:2018:FAP

[Wil18] R. Ryan Williams. Faster allpairs shortest paths via circuit complexity. SIAM Journal on Computing, 47(5):1965– 1985, ???? 2018. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Winograd:1980:MPM

[Win80] S. Winograd. On multiplication of polynomials modulo a polynomial. SIAM Journal on Computing, 9(2):225–229, ???? 1980. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wirth:1979:SVD

[Wir79] Michael C. Wirth. Symbolic vector and dyadic analysis. SIAM Journal on Computing, 8(3):306–319, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wang:2000:MEA

[WJG00] Lusheng Wang, Tao Jiang, and Dan Gusfield. A more efficient approximation scheme for tree alignment. SIAM Journal on Computing, 30 (1):283–299, February 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31350.

Wang:1995:HSI

[WL95] Qingzhou Wang and Ernst L. Leiss. A heuristic scheduling of independent tasks with bottleneck resource constraints. SIAM Journal on Computing, 24(2):235–241, April 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24331.

Wu:1999:PTA

[WLB^{+}99] Bang Ye Wu, Giuseppe Lancia, Vineet Bafna, Kun-Mao Chao, R. Ravi, and Chuan Yi Tang. A polynomial-time approximation scheme for minimum routing cost spanning trees. SIAM Journal on Computing, 29(3):761–778, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32253.

Wang:2008:AAB

[WLL08] Lusheng Wang, Yu Lin, and Xiaowen Liu. Approximation algorithms for biclustering problems. SIAM Journal on Computing, 38(4):1504– 1518, ???? 2008. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Walicki:1997:SPN

[WM97] Michal Walicki and Sigurd Meldal. Singular and plural nondeterministic parameters. SIAM Journal on Computing, 26(4):991–1005, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26431.

Wormald:1987:GRU

[Wor87] Nicholas C. Wormald. Generating random unlabelled graphs. SIAM Journal on Computing, 16(4):717–727, August 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wang:2003:AEA

[WP03] Xinmao Wang and Victor Y.

Pan. Acceleration of Euclidean algorithm and rational number reconstruction. SIAM Journal on Computing, 32(2):548–556, April 2003. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40863.

Wrathall:1978:RPR

[Wra78] Celia Wrathall. Rudimentary predicates and relative computation. SIAM Journal on Computing, 7(2):194–209, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wright:1986:CTG

[WROM86] Robert Alan Wright, Bruce Richmond, Andrew Odlyzko, and Brendan D. McKay. Constant time generation of free trees. SIAM Journal on Computing, 15(2):540–548, ???? 1986. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wagner:1978:CCA

[WS78] Robert A. Wagner and Joel I. Seiferas. Correcting counterautomaton-recognizable languages. SIAM Journal on Computing, 7(3):357–375, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wolfson:1991:CCA

[WS91] Ouri Wolfson and Adrian Segall. The communication complexity of atomic commitment and of gossiping. SIAM Journal on Computing, 20(3): 423–450, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wang:1979:NAP

[WT79] Paul S. Wang and Barry M. Trager. New algorithms for polynomial squarefree decomposition over the integers. SIAM Journal on Computing, 8(3):300–305, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Westbrook:1989:AAA

[WT89] Jeffery Westbrook and Robert E. Tarjan. Amortized analysis of algorithms for set union with backtracking. SIAM Journal on Computing, 18(1):1–11, February 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

White:1975:CAE

[WW75] D. E. White and S. G.

Williamson. Computational algorithms for the enumeration of group invariant partitions. SIAM Journal on Computing, 4(3):209–213, ???? 1975. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wang:2006:RCE

[WW06] Li-San Wang and Tandy Warnow. Reconstructing chromosomal evolution. SIAM Journal on Computing, 36(1): 99–131, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Williams:2013:FMC

[WW13] Virginia Vassilevska Williams and Ryan Williams. Finding, minimizing, and counting weighted subgraphs. SIAM Journal on Computing, 42(3): 831–854, ???? 2013. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Widmayer:1987:SDP

[WWW87] P. Widmayer, Y. F. Wu, and C. K. Wong. On some distance problems in fixed orientations. SIAM Journal on Computing, 16(4):728–746, August 1987. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Woodruff:2007:GAI

[WY07] David Woodruff and Sergey Yekhanin. A geometric approach to informationtheoretic private information retrieval. SIAM Journal on Computing, 37(4):1046–1056, ???? 2007. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Weihrauch:2006:ACF

[WZ06] Klaus Weihrauch and Ning Zhong. An algorithm for computing fundamental solutions. SIAM Journal on Computing, 35(6):1283–1294, ???? 2006. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wilson:2015:FBS

[WZ15] David B. Wilson and Uri Zwick. A forward-backward single-source shortest paths algorithm. SIAM Journal on Computing, 44(3):698–739, ???? 2015. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Wang:2021:TAC

[WZ21] Haitao Wang and Jingru Zhang. An O(n log n)-time algorithm for the k-center problem in trees. SIAM Journal on Computing, 50(2):602–635, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Xia:2013:SFD

[Xia13] Ge Xia. The stretch factor of the Delaunay triangulation is less than 1.998. SIAM Journal on Computing, 42(4):1620– 1659, ???? 2013. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Xiao:2009:EAR

[XLXJ09] Jing Xiao, Lan Liu, Lirong Xia, and Tao Jiang. Efficient algorithms for reconstructing zero-recombinant haplotypes on a pedigree based on fast elimination of redundant linear equations. SIAM Journal on Computing, 38(6):2198–2219, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yacobi:1998:FEU

[Yac98] Yacov Yacobi. Fast exponentiation using data compression. SIAM Journal on Computing, 28(2):700–703, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23497.

Yannakakis:1981:NDP

[Yan81a] M. Yannakakis. Node-deletion problems on bipartite graphs. SIAM Journal on Computing, 10(2):310–327, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yannakakis:1981:EDP

[Yan81b] Mihalis Yannakakis. Edgedeletion problems. SIAM Journal on Computing, 10(2):297– 309, ???? 1981. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yannakakis:1982:FDS

[Yan82] Mihalis Yannakakis. Freedom from deadlock of safe locking policies. SIAM Journal on Computing, 11(2):391–408, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1976:EP

[Yao76] Andrew Chi-Chih Yao. On the evaluation of powers. SIAM Journal on Computing, 5(1): 100–103, ???? 1976. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1978:LSA

[Yao78] Andrew Chi Chih Yao. On the loop switching addressing problem. SIAM Journal on Computing, 7(4):515–523, ???? 1978. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1979:CPM

[Yao79] Andrew Chi Chih Yao. The complexity of pattern matching for a random string. SIAM Journal on Computing, 8(3): 368–387, ???? 1979. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1980:BSN

[Yao80] Andrew Chi Chih Yao. Bounds on selection networks. SIAM Journal on Computing, 9(3): 566–582, ???? 1980. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1981:AMA

[Yao81] Andrew C. Yao. An analysis of a memory allocation scheme for implementing stacks. SIAM Journal on Computing, 10(2): 398–403, ???? 1981. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1982:CMS

[Yao82] Andrew Chi Chih Yao. On constructing minimum spanning trees in k-dimensional spaces and related problems. SIAM Journal on Computing, 11(4): 721–736, ???? 1982. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1985:CMP

[Yao85a] Andrew C. Yao. On the complexity of maintaining partial sums. SIAM Journal on Computing, 14(2):277–288, ???? 1985. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1985:EPP

[Yao85b] Andrew C. Yao. On the expected performance of path compression algorithms. SIAM Journal on Computing, 14(1): 129–133, ???? 1985. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1988:MBG

[Yao88] Andrew Chi Chih Yao. Monotone bipartite graph properties are evasive. SIAM Journal on Computing, 17(3):517–520, June 1988. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1989:CPO

[Yao89] Andrew Chi Chih Yao. On the complexity of partial order productions. SIAM Journal on Computing, 18(4):679–689, August 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1991:LBA

[Yao91] Andrew Chi Chih Yao. Lower bounds for algebraic computation trees of functions with finite domains. SIAM Journal on Computing, 20(4):655–668, August 1991. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1994:NOT

[Yao94] Andrew Chi Chih Yao. Nearoptimal time-space tradeoff for element distinctness. SIAM Journal on Computing, 23 (5):966–975, October 1994. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/14895.

Cai:2000:CP

[yCLZ00] Jin yi Cai, Richard J. Lipton, and Yechezkel Zalcstein. The complexity of the ABC problem. SIAM Journal on Computing, 29(6):1878–1888, December 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
27685.

Yao:1989:PSR

[YDEP89] F. Frances Yao, David P.

Dobkin, Herbert Edelsbrunner, and Michael S. Paterson. Partitioning space for range queries. SIAM Journal on Computing, 18(2):371–384, April 1989. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Ye:1990:CPT

[Ye90] Yin Yu Ye. A class of projective transformations for linear programming. SIAM Journal on Computing, 19(3):457–466, June 1990. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yesha:1983:CPT

[Yes83] Yaacov Yesha. On certain polynomial-time truth-table reducibilities of complete sets to sparse sets. SIAM Journal on Computing, 12(3):411–425, ???? 1983. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yokouchi:1990:RSC

[YH90] Hirofumi Yokouchi and Teruo Hikita. A rewriting system for categorical combinators with multiple arguments. SIAM Journal on Computing, 19(1): 78–97, February 1990. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yang:1987:OPS

[YHC87] Mark C. K. Yang, Jun S. Huang, and Yuan-Chieh Chow. Optimal parallel sorting scheme by order statistics. SIAM Journal on Computing, 16(6):990– 1003, December 1987. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yamada:1983:OST

[YI83] Hisao Yamada and Masatosi Imori. One step transformation of periodic sequences by cellular automata. SIAM Journal on Computing, 12(3):539– 550, ???? 1983. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yang:1995:RPP

[YLW95] Chung-Do Yang, D. T. Lee, and C. K. Wong. Rectilinear path problems among rectilinear obstacles revisited. SIAM Journal on Computing, 24(3):457–472, June 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22967.

Yao:1980:PDP

[YR80] Andrew C. Yao and Ronald L.

Rivest. On the polyhedral decision problem. SIAM Journal on Computing, 9(2):343– 347, ???? 1980. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yu:1992:RDO

[YR92] Lin Yu and Daniel J. Rosenkrantz.

Representability of design objects by ancestor-controlled hierarchical specifications. SIAM Journal on Computing, 21(5): 824–855, October 1992. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yeap:1993:FPG

[YS93] Kok-Hoo Yeap and Majid Sarrafzadeh. Floor-planning by graph dualization: 2-concave rectilinear modules. SIAM Journal on Computing, 22(3): 500–526, June 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yu:2022:NOS

[Yu22] Huacheng Yu. Nearly optimal static Las Vegas succinct dictionary. SIAM Journal on Computing, 51(3):STOC20– 174–STOC20–249, ???? 2022. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1363649.

Yun:1979:UBC

[Yun79] David Y. Y. Yun. Uniform bounds for a class of algebraic mappings. SIAM Journal on Computing, 8(3):348–356, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1982:ACC

[YY82] Andrew C. Yao and F. Frances Yao. On the average-case complexity of selecting the k th best. SIAM Journal on Computing, 11(3):428–447, ???? 1982. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yao:1985:FTN

[YY85] Andrew C. Yao and F. Frances Yao. On fault-tolerant networks for sorting. SIAM Journal on Computing, 14(1):120– 128, ???? 1985. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yousefi:2014:LPM

[YY14] Arman Yousefi and Neal E. Young. On a linear program for minimum-weight triangulation. SIAM Journal on Computing, 43(1):25–51, ???? 2014. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Yoshida:2012:ICT

[YYI12] Yuichi Yoshida, Masaki Yamamoto, and Hiro Ito. Improved constant-time approximation algorithms for maximum matchings and other optimization problems. SIAM Journal on Computing, 41(4): 1074–1093, ???? 2012. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Zave:1977:OPS

[Zav77]  Derek A. Zave.  Optimal

polyphase sorting. SIAM Journal on Computing, 6(1):1–39, ???? 1977. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Zellini:1988:OBS

[Zel88] Paolo Zellini. Optimal bounds for solving tridiagonal systems with preconditioning. SIAM Journal on Computing, 17(5): 1036–1043, October 1988. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Zhang:2005:ETC

[ZH05] Huaming Zhang and Xin He.

On even triangulations of 2connected embedded graphs. SIAM Journal on Computing, 34(3):683–696, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40824.

Zhang:1995:ORS

[Zha95] Yanjun Zhang. On the optimality of randomized αβ search. SIAM Journal on Computing, 24(1):138–147, February 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
23037.

Zhang:2009:TBR

[Zha09] Shengyu Zhang. Tight bounds for randomized and quantum local search. SIAM Journal on Computing, 39(3):948–977, ???? 2009. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Zhu:2021:TSB

[Zhu21] Leqi Zhu. A tight space bound for consensus. SIAM Journal on Computing, 50 (3):STOC16–18–STOC16–29, ???? 2021. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Zimand:1998:WNO

[Zim98] Marius Zimand. Weighted NP optimization problems: Logical definability and approximation properties. SIAM Journal on Computing, 28(1):36– 56, February 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
28510.

Zaks:1979:GTO

[ZR79] S. Zaks and D. Richards. Generating trees and other combinatorial objects lexicographically. SIAM Journal on Computing, 8(1):73–81, ???? 1979. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Zhang:1989:SFA

[ZS89] Kaizhong Zhang and Dennis Shasha. Simple fast algorithms for the editing distance between trees and related problems. SIAM Journal on Computing, 18(6):1245–1262, December 1989. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Zhou:2002:AMV

[ZS02] Yunhong Zhou and Subhash Suri. Algorithms for a minimum volume enclosing simplex in three dimensions. SIAM Journal on Computing, 31 (5):1339–1357, October 2002. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36399.

Zuckerman:1996:UVN

[Zuc96] David Zuckerman. On unapproximable versions of NPcomplete problems. SIAM Journal on Computing, 25(6):

1293–1304, December 1996. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26640.

Zuckerman:2019:CPF

[Zuc19] David Zuckerman. Certifiably pseudorandom financial derivatives. SIAM Journal on Computing, 48(6):1711–1726, ???? 2019. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Zwick:1991:LBC

[Zwi91] Uri Zwick. A 4n lower bound on the combinational complexity of certain symmetric Boolean functions over the basis of unate dyadic Boolean functions. SIAM Journal on Computing, 20(3):499–505, June 1991. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).